A Complete Bibliography of Publications in IEEE Transactions on Visualization and Computer Graphics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 2.01

Title word cross-reference

#FluxFlow [ZCW^{+}14].

(X, Y) [BBD^{+}11]. + [CMP09, FWG09, QWC^{+}09, WC11, ZZG^{+}12]. 1 [SZB^{+}09]. 1.5 [SWW^{+}15]. 2 [AL06, AJDL08, AHKMF11, ARH^{+}15, AJ19, BBiA09, BHSH15, CKW^{+}12, DNP07, DHM13b, DMC15, GRT17, GSS^{+}19, GJR^{+}14, JDL09, LKJ^{+}05, LWS^{+}17, LRN96, LJL04, LKR^{+}18, LLCD11, LCS^{+}12, LFR03, MSA17, NHYY18, PQMCR17, PGT^{+}08, PMW13, QCT13, RKG^{+}11, RF11, SZB^{+}09, SHM^{+}07, SMN12, SWCR15, SS13b, TWHS05, TKAM06, TIK15, WWYS04, WLL^{+}05, WTL^{+}09, Won16, YHH^{+}19, YEII12, ZWZ^{+}13, ZWR14, vWN04]. 2.5 [APV^{+}15, YSL^{+}13, YJL^{+}15]. 3 [ARH^{+}15, AHR^{+}11, AVP19, BSB^{+}18, BGCS17, BF01, BIAI17, BTB10, BTHD11, BC12, CWZ^{+}14, CWL12, CPW^{+}18, ChLYL09, CC08, CZN^{+}11, CZZ^{+}19, CMK15, CCM11, DDW14, DHL09, DMR04, DBW11, DRHK07, ERL^{+}13, ET08, FW08, FYTL19, FCL09, GCL^{+}15, GW13, GRT17, GCL^{+}18, GIS03, GJR^{+}14, GPP^{+}16, HDBC15, HE06, HLL97, HLRC^{+}12, HM10, HLYL18, IFP97, IFM14, JSG03, JK16, JNC^{+}15, JDL09, JBS06, JCWD14, KRW19, KSG^{+}16, KZL07, KHS^{+}19, KCA16, KLG^{+}16, KCK^{+}19a, KYT^{+}18, KKKW05, LYY^{+}16a, LS07a, LH03, LDH09, LKL^{+}15, LYL19, LKR^{+}18, LZZ^{+}19, LDN11, LJZ12, LL05, LY06, LB17, LGS12, LODI16, LSM03, MS08, MWSJ14, MW99, MCHM10, MAST16, MCG12, MH10, MY14, MJK06, MDS16, MES^{+}11, MWC^{+}12, NWHWD16, NB95, NQX^{+}05, NHYY18, OA11, dJOBNM17, OHH06, PDRK19, PBO^{+}14, PK08, PLW11, PLW12, PJ03, PZ07, PH07]. 3 [PUNI11, PMT^{+}19, PGI^{+}17, QY07, QWC^{+}09, RSBB17, RHZN11, RT12, RKZZ19, SK16a, SW12, SWB^{+}00, SBV^{+}11, Sel15, SKYS14, SPP^{+}14, SPB96, SXX^{+}19, SW97, SRW^{+}16, SGJM18, SKW^{+}11, SBW17, SKH^{+}19, SW17, TKTN09, TCL^{+}13, TNT17, TMWS13, TC09b, TZL^{+}12, TKAM06, TLC^{+}10, VCP08, VAB12, VW12, Vas16, WH09, WWC^{+}14, WM18, WTL^{+}09, WQZ^{+}18, WB08, WSE07, WFG^{+}19, WXY17, WCB^{+}12, WLDW11, XCZ^{+}19b, XFZ^{+}19, XLZ19c, YOS13, YPRI17, YFC^{+}19, YLSL11, YL95, YHH^{+}19, YSI^{+}10, YEII12, YEII16, ZTP05, ZTZ17, ZPP05, ZCFL15, vLBB16, zBBKN14]. 360^{∘} [LXRY18, MS18a, ZLK^{+}18]. 360 [BCR19, SKC^{+}19]. 4 [BB09, BPM^{+}13, CTGH13, CFHH09, HDJ05, INCB18, KGP^{+}13, LHC16, ZH07, ZWR14, vPBB^{+}10, vPBB^{+}11]. 6 [GL17]. 8 [HDJ05]. ^{2} [BJK^{+}16, SB17, XCZ^{+}19b]. ^{∘} [RPAC17]. ^{N−D} [NM13]. f: R^{3} → R^{2} [SSC^{+}16]. hp [NK06, KDBB17]. K [Zha14, CK10, KHM^{+}98, VPF15]. k^{+} [VPF15]. L_{0} [GXW^{+}18a, GXW^{+}18b, SSW18]. μ [BWS^{+}19]. N [IHR01, OHWS13]. ν [Nie04]. p [ZK14a]. R [FGF^{+}05]. \sqrt{3} [WQS07]. ⊇ [OBLN17]. Z [KYK11, WHL16]. || [BJK^{+}16].

-Adaptive [KDBB17]. -buffer [VPF15, Zha14]. -Buffers [WHL16]. -Clustering [BBD^{+}11]. -D [MS08]. -dimensional [ZWR14, IHR01]. -DOF [GL17, TIK15]. -DOPs [KHM^{+}98]. -Functions [FGF^{+}05]. -Nearest [CK10]. -Partition [ZK14a]. -Quaternion [Nie04]. -Subdivision-Based [WQS07]. -Test [KYK11].

1 [ZCD19]. 11 [RPHI08]. 1115 [CWDH09]. 15th [KHE09].

2 [CD19]. 2000 [Var01]. 2005 [CLS07, SGR06, SW06]. 2008 [CR08, EDF08]. 2009 [Ano09e, Ano09d]. 2010 [Ano10c, Ano10d]. 2011 [Ano11i, Ano11c, Ano11d]. 2012 [Ano12e, Ano13c, Ano13d, HKQ13, Ros13, Sch13]. 2013 [Ano13t, Ano13u, Ano13e, Ano13f, Bil13, CLS13a, Fuc13, GJK15, MY14]. 2014 [Ano14e, Ano14g, Ano14f, BHTY15, JLS15]. 2015 [Ano15c, Ano16b, Ano16c, LST^{+}16, WLW17]. 2016 [Ano16d, Ano16e, Dil17, Ebe17, Meh17, Ota17]. 2017 [DS17a, De 18b, DW17, Han18, Hee18]. 2018 [Ano19a, BKL18, Car19, DS18a, Ynn19]. 2019 [MSW19]. 2nd [KBB^{+}18].

3D [PKMR15]. 3DCT [AHRG10].

50 [BWS^{+}19].

6 [BDF16]. 6-DoF [BDF16].

7DOF [CLW18].

AABBs [ZK07]. Abdominal [SG09, Sel15]. Abilities [FTES13]. Ability [DBP14, OPH^{+}16]. Ablation [RKSH11]. Abstract [BM13, CL06, MHD^{+}18, PTC10, SPL^{+}13]. Abstracting [MJW^{+}13]. Abstraction [AAFW17, CCM^{+}14, CG07, CWRY06, EBB^{+}15, JER16, KLC09, KMM^{+}13, LWCC18, SMER06, VI18, ZMT^{+}19]. Abstractocyte [MAAB^{+}18]. ABySS [NJBJ09]. ABySS-Explorer [NJBJ09]. Accelerated [BC12, KMH11, LM05, MK09, PSR17, QMK^{+}06, SH00b, FM07]. Accelerating [LCDP13]. Acceleration [KRHH11, LQLX14, RGC^{+}14]. Acceptability [KLD^{+}09]. Accessibility [KKMS11, SR00]. Accessible [CKLL09, KMKY10, eYL07]. Accessory [KW14]. Accident [UKW19]. Accommodation [Kra16, XGS^{+}19]. Accommodation-Vergence [XGS^{+}19]. Accommodative [YBW^{+}19]. Accrual [MDL^{+}19]. Accumulation [MDL^{+}19]. Accuracy [GBP19, HV00, LB03, MKW07, SCKR08, WMS98]. Accuracy-Conserving [SCKR08]. Accurate [AT05, AGDJ10, BES12, BN12, DWB^{+}06, FH16, GKT^{+}08, GBP12, HSSK16, HR11, LDW^{+}15, MCP^{+}06, MNKW07, NLKH12, PMT^{+}19, RNL09, SSIF09, SCYW16, Vis15, WHM14, WHL16]. Achievement [Ano10d, Ano11d, Ano13d, Ano13f, Ano14e, Ano14g, Ano14f, Ano16c, Ano16e, Bil13, Ebe17, Hee18, Sch13, Ynn19]. Achieving [PMvWC05]. ACM [BvdP12, BDC17, CLS07, GW13, HK10, KS14a, KPGL12, KL14b, LS06, MY14, Ota17, SK16a, SK15, SW17, TL11, WLW17]. ACM/Eurographics [Ota17]. Acoustic [ACTM12, FM04, MAST16, MDHB^{+}07, SLM18]. Acoustics [DBM^{+}06]. Acquired [BSS^{+}13, BAW16]. Acquiring [SSI99, YL16]. Acquisition [AVP19, LZD13, MWC^{+}12, OWS15]. Across [GDST16, GGL^{+}14a, HTC09, KHSS14, RGFLL14, BLLS17, KMDH11, YQK^{+}17]. Action [BBS^{+}08, JNC^{+}15, JOR^{+}19, MI13, NTA^{+}19, WCR^{+}11]. Action-Based [BBS^{+}08]. Actionable [MWN^{+}19]. Activation [AS11]. Active [BHZ^{+}18, CVC10, GCL^{+}15, RRJH18, WHP^{+}18]. Actively [SFMB12]. ActiVis [KAKC18]. Activities [KFS^{+}19]. ActiviTree [VJC09]. Activity [AAM^{+}12, ERLW18, Mar18, MMAM14, OSBM14, RFFT17, STS^{+}14, WS09]. Activity-Centered [Mar18]. Actuation [KJH^{+}18]. Acuity [PK13]. Acuity-Driven [PK13]. AD [CLT^{+}08]. AD-Frustum [CLT^{+}08]. ADAPT [KMSB14]. Adaptation [ANR^{+}18, Bac07, BSSL19, HDJ05]. Adaptive [BT13, CLCQ12, CL18, CZZ17a, CSG^{+}19, CLT^{+}08, DK11a, Dic14, DWB^{+}06, EDF11, FM12a, GSA^{+}09, GABJ08, GSDJ04, GB08b, Jen12, KSH03, KL07, KL14a, KNO15, KDBB17, LPKK12, LPLT11, LPQF14, MS08, MDS^{+}18, NW10, NOB16, OR98, OR99, PBPP11, PMvWC05, RNL09, RLNN11, STH13, SRML09, TSLR07, WWW^{+}19, WL16, WSC^{+}95, WPB^{+}11, XESV97, XA09, YWV^{+}19, YL08, ZG12, ZHZ15, ZPS04, Zhu05]. Adaptively [ATT12, WWC^{+}14]. AdaptiviTree [TSLR07]. Adding [SRCP02, XST^{+}18, SRCP03]. Additive [WAWS18]. Address [Cox11, Heg10, SDW09, Sza10, Tha11]. Addressable [LHC10]. Adjacency [DWvW12, HBW14, Hol06, OJK19]. Adjoints [Chr03]. Adjustment [CLW18, DL12, HWHK16, LWZQ17, LGY19, SSEW19]. Admissible [HQ13]. ADMM [OBLN17]. Adolescents [BZS^{+}13]. Adoption [BISM14]. ADR [KST^{+}14]. Adults [JBS^{+}18]. Advanced [HFL18, KKPS08, LS13a, LMG06, NWF^{+}05]. Advancement [KMT14]. Advances [BSS^{+}19, LMW^{+}17, PKS^{+}08]. Advection [GZL^{+}14, Har16, HZ13, JEH02, WSE07, YNBH11, ZHGH11]. Advection-Based [GZL^{+}14]. Advections [KLLR07]. Adversarial [SKY12, WGYS18]. Adverse [BSR^{+}14]. Aerial [VABW09, ZHLR14, ZBMY14]. Aesthetic [ASG15]. Aesthetics [BW08c, PPP12]. Affect [RKC^{+}16, VSS08, ZGC^{+}17]. Affective [SPW07, VBC^{+}16]. Affine [JCC^{+}11, LQLX14]. Affordable [LPG^{+}18]. Affordance [CC12]. Affordance-Based [CC12]. Affordances [BEDF16, BRNB19, GPK14]. After [LAP19]. Age [NC07]. Agency [JAAL18, KPV^{+}18, RKC^{+}16]. Agent [GCL^{+}18, JOR^{+}19, KMSB14, NBM19]. Agents [DV95, GSA^{+}09, PUNI11, RBK^{+}19, WKW06]. Aggregate [CMFL16, TMDO15]. Aggregation [AA11, BWK^{+}13, EF10, JER16, WPS^{+}09, Wil18]. Aggregations [CvW18, YEB16, vdEvW14]. AggreSet [YEB16]. Aging [GdBG12]. Agnostic [ZWM^{+}19]. Aid [KA12, SSMG13]. Aided [FYZ^{+}17, HQK06, Yan18]. Aiden [Ano13u]. Air [AAFG18, EGG^{+}12, KZL07, QCX^{+}07]. Aircraft [Chi16, HTC09]. Airflow [ZMH^{+}09]. Algebraic [AG16a, KS14b]. Algorithm [AGY^{+}17, Ano96b, BMR^{+}99, BKS01, BW00, BLW14, CL18, ChLYL09, CMF^{+}18, CL09, DBD13, DSF^{+}14, GH00, Gor02, HK99, HSL19, IHR01, IYK01, IYIK04, KNP04, KS00a, KS00b, KS01, KDBB17, LKHW04, LMG06, LDX10, LYY^{+}16b, LSJ96, LB03, MS08, MM11, MPG^{+}14, SK98, SPB96, Sil95, SM97, TLH10, WTL^{+}09, XA09, YYSZ06, YNM15, YPRI17, YXSH13, Zha14]. Algorithms [AZM12, BLS04, CSM07, FHS^{+}12, HBT14, KWP01, LLG17, PML97, PPP12, SJL^{+}18a, TD95, TWSM^{+}11, WCC^{+}19, WHR02, XXM^{+}19b, vLBB16]. Alias [SK99]. Alias-Free [SK99]. Aliased [SF19]. Aliasing [CMFL16, MT05]. Aligned [CS08, CSC06]. Alignment [ADG11, FA15, JKM06, KH19]. Alignments [WWFT03]. All-Frequency [HXF^{+}15, LHLW10, SHR^{+}11, WPC^{+}13, XJF^{+}08]. AllAboard [DSC^{+}16]. Allocation [DDBR^{+}19, HLG^{+}14, MMT^{+}14]. Almost [PTC10]. Alone [LKS^{+}19]. along [WM18]. Alpha [WM19]. Alphabetically [WBDS11]. Altering [SMP17]. Alternate [CG14]. Alternative [PDW^{+}14, SAB^{+}16, HSKIH07]. Ambient [ASW13, AD16, HLY10, LKS^{+}19, PMP10, RB11, SMG^{+}13, SVGR16, TCM06].

Ambiguity [AJ19, LLCM12, WSA^{+}16].

Ambiguity-Free [LLCM12].

AmbiguityVis [WSA^{+}16]. Ambiguous [KNKH19, KCC^{+}17]. Ameliorating [RMW09]. Among [KGS98, AL06]. Amongst [CC07]. Amplified [RSBB17]. AMR [Ano09b, ME11b, SP07]. AnaFe [GDKB17]. Analogy [RM15, SVK^{+}07]. Analyses [GLK^{+}13, JBS^{+}18, JA18, KTB^{+}18]. Analysis [AHSS14, wAPS14, AHK^{+}17, AAMH13, AHRG10, AAH^{+}13, AAFG18, AAGS19, ABC^{+}19, ASG15, BSS^{+}13, BMJK09, BRT12, BBD^{+}11, BE18, BKW16, BKL^{+}11, BLM96, BMLC19, BMWM06, BAAK^{+}13, BPB14, BCB10, BAF^{+}13, Bon98, BPM^{+}13, BISM14, BNTM16, BWT^{+}11, BDW^{+}08, CGSQ11, CSL^{+}16, CLG16, CM10, CD19, CZZ17b, CHW^{+}18, CGJM19, CWQ^{+}07, CDL^{+}16, CRT04, CDK^{+}17, DVP^{+}18, DLW^{+}17, DvVH^{+}19, DSG^{+}17, DFD^{+}14, DTW^{+}15, DCK^{+}12, DS16b, DCH^{+}17, DB07, Eic00, EJR^{+}14, FPB17, FCZ15, FMH08, FHG^{+}09, FZCQ17, GFG^{+}14, GKL^{+}13, GMS^{+}07, GMD13, GSL^{+}17, GJC^{+}17, GGA^{+}11, GHL18, GLG^{+}13, GRVE07, GQM^{+}18, GXY12, GXZ^{+}18, GHS^{+}19, GJG^{+}19, GWK12, HSCW13, HBG11, HEFR18, HDSC19, HMSA08, HTP^{+}08, HKB^{+}19, HV00, HMZ^{+}14, HPvU^{+}18, HLG^{+}14, HZ13, HOGJ13, JFSK16, JBMS09, JCG08, KPHH12, aKGS11, KBE^{+}18, KMDH11, KH13, KHS^{+}19, KJW^{+}18, KCS^{+}16, KOJL^{+}14, KLG^{+}16, KBGE11]. Analysis [KJW^{+}14, KMG^{+}06, KBH06, KFS^{+}19, KGG^{+}12, KRRW19, KCPS08, Lac96, LSSB12, LBS13, LTM18, LBW19, LS13b, LSPS10, LGYG12, LPLT11, LLC15, LLB^{+}06, Lin16b, LSS09, LH14, LS16, LSL^{+}17, LXR19, ME18, MHS07, MEV^{+}14, MS04, MKN^{+}07, MMB^{+}19, MGKH09, MGJ^{+}10, MBL^{+}06, MGB^{+}19, MMDP10, MBH^{+}12, MDHB^{+}07, MGM09, MZC^{+}16, NMGK17, NDR96, NTT^{+}19, OJ15, OBJ16, OSSK12, OJCJP16, ODH^{+}07, OHWS13, PVF13, PYHZ14, PH11, POM^{+}09, PV06, PBCR11, QCX^{+}07, RESC16, RBS^{+}18, RS12, RAL^{+}17, RWF^{+}13, RML12, RFLL18, RKC^{+}16, RGC^{+}14, RB18, SZS^{+}17, SKB^{+}18, SPS06, SJJ^{+}17, SCT^{+}10, STM17, SHVV16, SFB^{+}12, SHM^{+}07, SGB13, SF14, SK13, SHB^{+}14, SW13, SOL^{+}13, SMER06, SYS^{+}06, SLK^{+}17b, SR00, SS18, SJL^{+}18b, STH02, SSL^{+}12, SGPR18, SWL^{+}14b, TKC17, TDLG19, TKW08, TFH11, TLLH12, TKBH17, USKD12, VP04a, WSH^{+}19, WZvdW13, WLSL17]. Analysis [WLS^{+}19, WHLS19, WK06, Wea10, Wen14, WGS^{+}13, WG16, WMA^{+}16, WPS^{+}16, WCQ^{+}09, WLY^{+}14, WPZ^{+}16, XYC^{+}18, XWW^{+}13, YSZ04, YRWG13, ZZD14, ZGC^{+}17, ZYLL09, ZYM^{+}14, ZMZM15, ZCPB11, ZCCB12, ZCW^{+}14, ZGB^{+}17, ZWC^{+}18, ZTZX13, ZZBW08, ZCW19, cKJG^{+}12, dLVvL06, tCMR08, vLBR^{+}16, WTS^{+}07]. Analyst [WM16]. Analysts [AZL^{+}19, GDJ^{+}13, GGZL16]. Analytic [CFEC17, DSP^{+}17, KGS^{+}08, LDSM17, MYM16, NXW^{+}16, PLK12, SKB^{+}18, WPC^{+}13]. Analytical [AJ97, Ano09b, EFN12, TAE^{+}11, WYM12]. Analytics [AAMG12, AABW12, Ano12g, BFE15, BMR^{+}19, BTC13, BSS^{+}19, BJK^{+}16, BAF^{+}13, CCSK19, CGM^{+}17, CCL^{+}16, CDW^{+}16, DLW^{+}17, ERLW18, EASS^{+}18, EASD^{+}19, Ert10c, GKL^{+}13, GGJ^{+}18, GS14, GDKB17, HHKE16, HKPC19, HSTD18, HTA^{+}15, HZM^{+}16, IFP^{+}12, aKGS11, aKS12, KLYE13, KRTvW06, KSDD14, KFS^{+}19, KTE15, KW13, KHSW17, KKL^{+}16, KCK^{+}19b, LCP^{+}13, LGM^{+}18, LXC^{+}17, LYW^{+}16, LWL^{+}17, LDM^{+}18, LFA^{+}16, LWLM18, MWSJ14, MRH^{+}10, MHR^{+}11, MMT^{+}14, MGS^{+}14, MW13, NA19, PBN^{+}13, PKL^{+}18, PGU^{+}13, PLvdM^{+}17, PHV^{+}18, PSBS12, PSM12, RLA^{+}13, SSS^{+}14, SSK^{+}16, SKKC19, SNLD06, SWY^{+}17, SKU^{+}12, SLK^{+}17b, SPG14, TKC17, TKE16, TKBH17, WSH^{+}19, WSD^{+}13, WDC^{+}18, WGYS18, WGSY19, WLS^{+}19, WCR^{+}18, WAG06, WFM^{+}06, WFC^{+}06, WSM^{+}09, XCH^{+}14, XZM17, XXM19a, YXG^{+}10, ZWLC19, vW11, vdEHBV16]. Analyze [VP09]. Analyzer [PGT^{+}08]. Analyzing [AMM^{+}08, AJ17, AHH^{+}14, BWP^{+}10, BvL06, CRI06, Dac11, FHKM17, KSDD14, LBS14, LSC^{+}18, LLL^{+}19b, MK13a, ORRL10, RC06, SKK^{+}14, TSH^{+}14]. Anatomically [WHM14]. Anatomy [DCKY02, ERL^{+}13, KST^{+}14, TBB^{+}08, XSZ^{+}17]. Anatomy-Based [DCKY02]. Anatomy-Driven [KST^{+}14]. Ancestral [FDC^{+}18]. Anchor [SZ11]. Anchoring [VZS18]. Aneurysm [JPLŠ16, MGB^{+}19, OJCJP16]. Aneurysms [GNBP11, GLvP^{+}12, GLH^{+}14, MVB^{+}17, MOJB^{+}19]. Angle [BA05, BHB04, HV13, HYB^{+}17, NSZ^{+}17]. Angles [LSK^{+}18, WYL^{+}19]. Angular [BZGV14, DMC15, GPL^{+}11, vAPP^{+}11]. Animal [KBH^{+}10]. Animated [BFP14, BGB15, CDF14, CMP14, HR07b, KH19, KKSM19, LWW^{+}07, NKP^{+}15, NDR96, TDM^{+}18, WASQ18, vRKEE17]. Animating [Dan16, LSY^{+}18, Ney98, ZBO13]. Animation [APP11, BW08a, BvdP12, BBiA09, BDC17, BB19, CGD97, CTGH13, CDR^{+}18, CLZ^{+}03, DNL^{+}06, EGS03, FvdPT97, FC95, HZM13, JAO^{+}14, KL07, KS14a, KP05, KL14b, LPG^{+}18, LGV^{+}16, LHBF19, LJL04, LG15, LAP19, LLPY07, MD12, MPT03, NB95, NT99, OH12, Ota17, Per95, QYH^{+}18, RLNN11, RFF^{+}08, RM13, SRCP02, SK15, TYSN06, TLC^{+}10, VMT06, WBA^{+}14, YN03, ZPS04, ZWW^{+}12, ZHF12, SRCP03]. Animations [CRH05, HE06, KIL07, LXB17, YLL^{+}12, ZCZ^{+}09]. AnimoAminoMiner [BLG^{+}16]. AniPaint [OH12]. Anisotropic [AD16, ATT12, DPR00, FHHJ08, FT09, HM10, KBH13, LB15, QLLM13, YSS^{+}12, ZHZ15]. Anisotropically [LKC09a]. Annealing [KH01]. Anniversary [KHE09]. Annotation [BMW17, GS06, NHYY18, PWHK16, ZGB^{+}17]. Annotations [LCL^{+}19, YHH^{+}19, ZGB^{+}17]. Annual [Ano01a, Ano03b, Ano04a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a, Ano14a]. Anomalous [CSL^{+}16, XXM19a, ZCW^{+}14]. Anomaly [CLZ^{+}18, XXM^{+}19b]. Anonymizing [WCC^{+}18]. Anthropomorphic [TAK^{+}05]. Anti [CMFL16, SF19]. Anti-Aliased [SF19]. Anti-Aliasing [CMFL16]. Antialiasing [MMS^{+}98, MRT00, SPW02, VT08]. Anticipation [MZH^{+}08]. Any [BWC04]. Aorta [RHR^{+}09]. Aortic [BSR^{+}14]. Apart [FFB18]. Apertures [BRNB19]. Appearance [JWD^{+}14, MSM^{+}11, SHM10, SLB04, Tay02, VBC^{+}16, Zho16, ZKM18]. Appearance-Based [MSM^{+}11]. Application [ACR^{+}19, BTB^{+}04, BGM^{+}17, DMR04, GGA^{+}11, HLRC^{+}12, HFG^{+}12, KKPS08, LLBS17, LKC09b, LRF^{+}11, MTB17, OMD^{+}12, PFP^{+}11, PM08, PNML08, PPM^{+}11, STS10, SD11, SGM08, YNYH06, ZWA^{+}13]. Application-Driven [ACR^{+}19]. Applications [AB01, BI12, BSO^{+}12, BJK^{+}16, CCSK19, COCSD03, CSM07, FDFR10, GXH^{+}13, GGTH07, GMM05, Gué01, GGLQ19, HQ13, Hub95, ISC07, JKJTM06, JBS06, Kei00, LLT04, LL04, MGPH06, MK13c, NTS11, QMK^{+}06, RF11, RSD^{+}13, SJL^{+}18a, SLA^{+}09, SXM17, SWC^{+}08, SK13, Ste98, WGS07b, WDC08, XXM19a, YL16, YML^{+}17, YXM^{+}15]. Applied [Kin10, SCT06, dLVvL06, vAPP^{+}11]. Applying [CLB13, HSTD18, SKK^{+}14, TFO09]. Approach [ADWK^{+}17, AAGS19, BMTD05, BMA^{+}19, BJK^{+}16, BW01, BMW17, CBL07, CCB11, CHW^{+}18, CGJM19, CGH^{+}19, CXM19, CSWP18, DC17, DGW11, DBB10, DSC^{+}08, DCKY02, FM12a, FHG^{+}09, FBLS05, GNBP11, GIK^{+}07, GNP^{+}06, GBHP08, HBJP12, HM95, HBW06, HZM13, HNR^{+}06, HZM^{+}16, KHA12, KLYE13, KS02, KTE15, KGZ^{+}12, KY06, KCM18, LBS13, LVRH07, LAK^{+}11, LT18, LXC^{+}17, LLZ^{+}16, LSPW12, MY96, MRH^{+}10, MHR^{+}11, MB19a, NR18, NSL19, OJ15, OHH06, PBN^{+}13, PSPM15, PSF09, PGL^{+}12, PSR17, RKK16, RE14, RFLL18, STS06, SS08, SS16, SAS16, SKU^{+}12, SPB08, SC15, SA19, SSBC19, TMWS13, TWSS16, TLM05, VFR13, WAM^{+}19, WM08, WGYS18, WFC^{+}18, WCC^{+}18, WASQ18, WGSY19, WQZ^{+}18, WBH04, XCH^{+}14, XZM17, YHH^{+}19, ZHF12, ZHLB13, ZGM18, vdEHBV16, WBD14]. Approaches [CK05a, FHKM17, FTES13, HB13, JAAL18, LD11b, MKT^{+}18]. Approaching [WAWS18]. Approximate [HMTR19, HQ07, JWD^{+}14, KMDZ10, RMCW19, TFO09, Wan11, Wil12]. Approximated [PLvdM^{+}17]. Approximating [HTF97, JSG03]. Approximation [BYA15, CGL^{+}17, GGZ^{+}18, LRZM11, PA06, RKSH11, SCOIT05, WXC^{+}08]. Approximations [GSG96, HJW99a, HJW99b, KH01, LDSM17, WBH04]. AR/VR [XGS^{+}19]. ARbis [IHD^{+}18]. Arbitrarily [PHF07]. Arbitrary [BWW^{+}12, HB03, KLS^{+}18, LLLF08, LTWH08, LBG^{+}08, LJWH08, LHZ^{+}04, NPPZ12, RNK^{+}15, RE14, VPB^{+}11, WDC08, YXSH13, ZC06]. Architectural [SYYC11]. Architecture [ARB07, BWW^{+}12, BGM^{+}07, BGM^{+}08, CGC^{+}11, CFHH09, FH07, HKC^{+}12, LLW15, MB03, MGPH06, MFS^{+}09, NKP^{+}15, PTMB09, SRHH16, SS95, SLF^{+}12, Wal12, ZPS04, LLLN^{+}14]. Architectures [HWZ^{+}19, HQQ12, RPSC99, SHC^{+}09, SvLF10]. Area [CDK04, LDSM17, LDN11, NSZ^{+}17, RSFH14, SNM16, TGS11, Wil12, XHL18, ZSG^{+}13]. Area-Based [LDN11]. Area-Preservation [ZSG^{+}13].

Area-Preserving [TGS11]. Area-Proportional [RSFH14, Wil12]. Areas [EIKS18, MVN^{+}19]. Arithmetic [LQLX14]. Arithmetic-Based [LQLX14]. Arm [TAK^{+}05]. Arrangements [GNCM^{+}16, MSSH14, SDMT16]. Array [TKTN09]. Arrays [Lin14f]. Arrive [GGZL16]. Art [Elb95, Elb98, KCWI13, NQX^{+}05, SPL^{+}13, SEH08, WFS^{+}19, WWY14, XZX^{+}17, LWP^{+}06]. Arterial [IPD^{+}07]. Artery [BGP^{+}11, MSSD^{+}08, TBB^{+}07]. Arthrodial [MGL07]. Arthroscopy [STH13]. Articulated [Fau99, KMT14, KBH^{+}10, KWC^{+}10, RLNN11, TF06]. Artifact [AHR^{+}11, MSSH14]. Artifacts [CMS06]. Artificial [BHHM19, RFFT17]. Artist [SK16b]. Artistic [CCM11, CRH05, KCWI13, RB11, Yon14]. Arts [LML^{+}18]. Artwork [PMCS11]. Artworks [ZJX^{+}15]. As-Rigid-As-Possible [LG15]. As-Similar-As-Possible [LG13]. Ascending [NMN^{+}18]. ASCII [IG19, XZX^{+}17]. ASK [AvHK06]. ASK-GraphView [AvHK06]. Aspect [FHSW13, JE13, WWZ^{+}18, WWF^{+}19]. Aspects [CZ11]. Assemblies [LSZ^{+}18, NJBJ09]. Assembly [CPG^{+}15, DBP14, MLS18, SLMA06, XMRC17]. Assembly/Maintenance [SLMA06]. Assessing [BRBF14, CB15, DBP14, MDF12, WCC^{+}19, YAE07]. Assessment [ASMP17, BTS^{+}18, GO15, HWZ^{+}19, LKK17, MMK^{+}17, WM08, ZZH19, vLBB16, HSKIH07]. Asset [KOBH19]. Assignment [RC06, SS16, WCG^{+}19]. Assist [HTP^{+}08, aKGS11, SBB^{+}18]. Assistance [FM06, MSA17, WWYP19]. Assisted [CICS05, CCJ^{+}19, HKR^{+}08, IV11, LAK^{+}11, LMC02, RLW^{+}11, SKKC19, WSH^{+}19]. Associate [De 18b, Lin14c, DS16a, DS17b, DS18b, MB19b]. associated [HSCW13]. Association [LS16, ZGW^{+}19].

Associations [HV13]. Assurance [TTR10]. Astroglial [MAAB^{+}18]. Astronomical [WAG^{+}12]. Astrophysical [LFH06, LFLH07, LFH08]. Astrophysics [SKW^{+}11]. Asymmetric [BN11, PLC^{+}11a, ZYLL09]. Asymmetrically [KHSS14]. Asymptotic [AJ19, CGL^{+}17]. Asynchronous [BSM06, KZX^{+}14, ZGI^{+}18]. Athletes [SB14]. Atlas [FFST19, KSI^{+}96, SLK^{+}17a]. Atlas-based [SLK^{+}17a]. Atlases [TDN^{+}12]. Atmospheric [KHS^{+}18, KHS^{+}19, SYS^{+}06]. Atom [PDFE18]. Attachment [KHL99, WLHD17]. Attack [RPHI08]. Attempt [MNKT01]. Attended [LKC09b]. Attention [CZZ^{+}19, Fis07, HW12, HE12, HLRC^{+}12, KV08, VFSG06, WLB^{+}14]. Attenuation [SCT06]. Attraction [DBD17, DBBF19, ZK10]. Attractive [WLB^{+}14]. Attribute [GLG^{+}13, HTE11, JBS08, LSS09, MJW^{+}13, MCK12, RSS14, SPEB18, SGAS16, TSAA12, TSH^{+}14, WZK12, WCC^{+}18]. Attribute-Based [SPEB18]. Attributes [AERA14, BMLC19, CM16, HSCW13, LWS^{+}17, LPCC17, LDFZ14, SPM^{+}13]. Audiences [DGAB16]. Audio [DDBR^{+}19, SNM16, SRKL19]. Audio-Material [SRKL19]. Audio-Visual-Olfactory [DDBR^{+}19]. Auditory [AGN^{+}19, RYKL13]. Augmentation [HDBC15, IAS19, TYL^{+}18]. Augmentations [OTKS15]. Augmented [BSB^{+}18, BBC15, BBBM18, BDB^{+}16, BSE^{+}17, BLO^{+}05, BLRW05, CDAF18, CMPC06, EPS^{+}15, EIKS18, FM04, GS08, GLM^{+}17, GKR14, GJK15, GLZR17, HBESB11, HIH^{+}18, HCP^{+}15, HBKS09, HAGS16, HJLH19, HF10, HF11, HV00, IHD^{+}18, JLS15, KMS09, KSNY17, KGAM18, KBB^{+}18, KM10, KHSB11, KYT^{+}18, KOK^{+}19, KV98, LH09, LBS^{+}16, LG12, LBKD09, LABS10, LDFZ14, MTM^{+}16, MNZ^{+}15, MUS16, MK13c, MKT^{+}18, NJJ11, OSB^{+}15, PLW12, PTM^{+}18, PIS15, RHJ^{+}16, RWBF18, RWF19, RBDG15, RJG17, SH00a, SYYC11, SJK^{+}07, SSE15, TN13, TGW^{+}95, UMW^{+}12, VGKS12, VBV^{+}18, WRM^{+}10, YC14, YON05, YON06, ZHLR14]. Augmenting [SDES19]. Aura [QY07]. Aural [AM13]. Auralization [TCM^{+}12]. Authalic [ZHGH11]. Authentically [BBBM18]. Author [Ano09d, Ano10e, Ano11e, Ano12c, Ano12d, Ano13g, Ano13h, Ano15e, Ano16f, Ano16g, Ano17c, Ano17d, LB19, SWY^{+}17]. Authored [FDFR10, ZGB^{+}17]. Authoring [ARL^{+}17, DLA^{+}09, FBM16, QH18, WLF^{+}19]. authors [Ano11l]. Autism [BBC15, BZS^{+}13, PMvWC05]. Auto [CDAF18, FCZ15, KLS^{+}18, SXX^{+}19]. Auto-Calibration [KLS^{+}18]. Auto-encoder [SXX^{+}19]. Auto-focus [CDAF18]. Auto-Harmonization [FCZ15]. Autocalibrating [SM11]. Autocalibration [SM12]. Automated [ASG15, BPG12, FH06, GHL18, MPK^{+}13, MRSS^{+}13, OJ15, PS12, PPM^{+}11, TAE^{+}11, ZCL08]. Automatic [AMJ^{+}12, AT16, BKM13, CCM11, DLF^{+}09, ED06, FXG12, GLvP^{+}12, HLB^{+}18, IK95, JHW^{+}14, KGP^{+}13, KWS^{+}14, LLL^{+}12, MHS07, MIO^{+}15, PPT^{+}11, RNK^{+}15, SRMOW11, SWF^{+}16, TW18, VAB12, WWYP19, Won16, ZT09, WHZ^{+}18]. Automating [Koo08]. Automultiscopic [LSK^{+}18]. Autonomous [WKW06, XLC^{+}18]. Autopsies [LWP^{+}06]. Autoregressive [LFA^{+}16]. Autostereoscopic [LDN11, MT05, PKS^{+}08, YHLJ08]. Available [BVW^{+}07]. Avatar [BPS13, JAO^{+}14, NBM19, RBLW07, WGR^{+}18]. Avatar-Agent [NBM19]. Avatars [Bai13, CK16, JAO^{+}14, KPBL16, LBLE^{+}14, PMT^{+}19, SSS13]. Average [GCNF13, RLK19, WC13]. Averaging [GBP^{+}13]. Aviation [CB15]. Avoid [DWA10]. Avoidance [BOP15, LCNG14, LKM^{+}18]. Award [Ano14d, Ano14e, Lin14c, Ano10c, Ano10d, Ano11c, Ano11d, Ano13c, Ano13e, Ano13d, Ano13f, Ano14g, Ano14f, Ano15c, Ano16b, Ano16c, Ano16d, Ano16e, Bil13, Car19, De 18b, Dil17, Ebe17, Fuc13, Han18, Hee18, Meh17, Ros13, Sch13, Ynn19]. Aware [AZC^{+}12, BYA15, BBBM18, BDSW13, BL07, CL18, CQC^{+}08, CBL07, CZC^{+}15, CLB^{+}16, DZMQ16, DWK^{+}16, HM10, JFZ^{+}18, KEJK19, KOBH19, LLY^{+}13, LJHY14, LLZ^{+}16, LSV^{+}18, MLKS18, MLMP18, NN11a, NN11b, OSS^{+}17, PRH10, RDB^{+}12, SCOIT05, TLD^{+}12, VF13, WCC^{+}18, WLZM10, YXG^{+}13, Yan18, YEII12, YLG^{+}14, YEII16, Zha14, ZJX^{+}15, WWZ^{+}19]. Awareness [GLZR17, HTL13, dJOBNM17, QPNK18, SSK^{+}16, IPJT19]. Axes [CvW11, KCPE16, LYY^{+}16b, TLH10]. Axis [AG16a, FGS19, KKW^{+}17, RSS14, SPB96, SCYW16]. AxiSketcher [KKW^{+}17].

B [BDHJ04, Cse10, Cse13, HHQH17, LWS97, LQLX14, ONL^{+}12, PH07, RE01a, XF04]. B-Mode [PH07]. B-Spline [BDHJ04, Cse10, Cse13, LQLX14, ONL^{+}12, RE01a]. B-Splines [LWS97, XF04, HHQH17]. Back [Ano08b, Ano11f]. Backchannel [DGWC10]. Background [KSY16]. Backgrounds [SPP^{+}14]. Backward [GT17]. Bacterial [WWFT03]. Bad [BTS^{+}18]. Bag [EHBA11]. Bag-of-Features [EHBA11]. Balanced [NLS11]. Balancing [DL12, NSZ^{+}17, PS06, ZGH^{+}18]. Ball [BMR^{+}99, WR11]. Ball-Morph [WR11]. Ball-Pivoting [BMR^{+}99]. Balloon [TS08]. Ballot [WBDS11]. BallotMaps [WBDS11]. Band [LKHW04]. Bandwidth [SCL08, WSS09]. Banking [HA06a]. Banquet [Kas12]. Bar [KHD02, KHDL07, TJW^{+}17, TSA14].

Baroclinic [KLySK12]. Barrier [HSF^{+}06, PKS^{+}08]. Bars [CG14]. Barycentric [SCT06, SAS05]. Bas [SRML09, SJM14, WTP^{+}19, ZZL^{+}15]. Bas-Relief [SRML09, SJM14, WTP^{+}19, ZZL^{+}15]. Base [KZW12]. Based [AJDL08, AHR^{+}11, Ano14h, ALMF19, AW03, ADP02, BW03, BW04, BRT12, BF01, BHST17, BKRE19, BBH^{+}17, BKDE00, BKL^{+}11, BWK^{+}13, BCB10, BK12, BSL^{+}12, BDSS18, BBS^{+}08, BMST97, BBB^{+}12, BLIC19, BWT^{+}11, Bro06, BMY05, BC12, BPS^{+}11, BKW10, CWL12, CWT^{+}08, CWM^{+}09a, COJ15, CBL07, CZZ17b, CMF^{+}18, CGH^{+}19, CTT^{+}16, CK11, CLRP13, CAN14, CRT04, CCM11, CMN13, CLB^{+}16, CAH^{+}13, CCB^{+}18, CC12, CZQ^{+}08, DVP^{+}18, DMC^{+}12, DH08, DCKY02, DZL^{+}14, DTW^{+}15, DBP14, EHBA11, EHP^{+}11, EMdSP^{+}15, FYP10, FXG12, FWZQ13, FYWY16, FHSW13, FR13, FBL^{+}18, FM12a, FGBB09, FWT^{+}04, FWR00, FBLS05, FST^{+}14, GS08, GXH^{+}13, GPL^{+}11, GNSP^{+}14, GIK^{+}07, GR04, GKM^{+}15, GHK97, Guo95, GLB^{+}06, GXY12, GZL^{+}14, GGZ^{+}18, GGLQ19, HWM95, HSR13a, HSR13b, HA17, HLM10, HEG^{+}17, HK99, HLG^{+}14, HQQ12, HQ04, HNR^{+}06, HPC^{+}13, HZM^{+}16, HTE11]. Based [IZM18, JWS04, KLC09, KRW19, KSG^{+}16, KRHH11, KKKT18, KSY16, KPBG13, KKSS13, KJ12, KEJK19, KBKG07, KWS^{+}14, KK19, KFN06, KTB^{+}18, KHH^{+}16, Lac96, LS07a, LPG^{+}18, LA11, LvWJH04, LMZ^{+}14, LCP^{+}13, LFP07, LS07b, LWZ^{+}16, LXB17, LDN11, LHFY12, LWCC18, LQLX14, LBH11, LDX10, LY12, LPG12, LDW^{+}15, LZLS16, LC10, LSS^{+}15, LXW^{+}18, LJH^{+}18, MCP^{+}06, MS08, MWSJ14, ME18, MOF10, MRSS^{+}12, MK13a, MB03, MS04, MAST16, MCK12, MTL18, MAKM14, MOG11, MSSD^{+}08, MSM^{+}11, MP13, MRT00, NJB07, NKH11, NKH14, NDR96, NOB16, NT99, NSS03, PDRK19, PYW^{+}16, PZLZ17, PPT^{+}11, PMP10, PJ03, PA06, PGL^{+}12, PFG08, PSM06, POD^{+}13, PMD^{+}07, QT96, QXF^{+}07, QK04, RGE19, RKG^{+}11, RKK16, RLNN11, RML12, RGC^{+}14, STB18, SED19, SVAC12, SND05, SNLD06, SWB^{+}00, SYM14, SMP11, SWC^{+}08, SF14, SZK15, SHS11b, SSIF09]. Based [Sel15, SAM^{+}07, SB04, SZN^{+}18, SZY^{+}18, SPN^{+}16, Sim07, SWCR15, SOL^{+}16, SPEB18, SO17, SGP^{+}19, SPW07, SSBC19, TL07, TDLG19, TCYM09, TWHS05, TSAA12, TLC^{+}10, TGW^{+}95, USE13, VP04a, VP04b, VJC09, VN19, WWYS04, WZW^{+}05, WW07, WGS07a, WQS07, WLSW08, WSS09, WH09, WM13b, WZvdW13, WFS^{+}16, WH18, WWF^{+}19, WZF^{+}04, WB08, WKME03, WGF08, WEE03, WSE07, WLD^{+}19, WQ07, WXZ^{+}16, XESV97, XLND11, XHT^{+}07, XCZ^{+}19a, XZB14, YRP18, YGV^{+}13, YLZ^{+}13, YNM15, YCLJ12, YLSL11, YL95, YL06, YLK12, YLY^{+}12, YON05, ZCL08, ZZL^{+}15, ZSL^{+}16, ZWS^{+}17, ZGH^{+}18, ZLK^{+}18, ZGW^{+}19, ZGB^{+}17, ZXM10, ZWM13, ZG06, ZCFL15, vAPP^{+}11, vPVvdW10, vLBB16, zBBKN14, BPC^{+}10, CJR07, CR08, CCSK19, CWDH09, CFHH09, CM08, DC17, DBM^{+}06, DCCW08, DMS^{+}08, EASD^{+}19, GHE19, HDSC19, JKM06, JM10, KLK^{+}09, LSC08, MOF09, MEV^{+}14, McK09, MRG^{+}15, MWK^{+}08, MDHB^{+}07]. based [NC07, PFK07, QMK^{+}06, RWF19, RKSH11, RL08, SNM16, SLK^{+}17a, WSW16, XXM^{+}19b, YDC^{+}14, YS17, YXG^{+}10, ZK06]. Baseline [NZS^{+}17]. Bases [SCOIT05]. Basic [HA18, SED19]. Basis [LHLW10, SG09, XJF^{+}08]. Basketball [WCW^{+}16]. BASTet [RB18]. Battery [GKL^{+}16]. Bayesian [CH17, GBP^{+}13, MDF12, NBM19, OPH^{+}16]. BCC [Kim13]. Be [RZP12]. Beam [AS11, MAST16]. Beautification [OK11]. Before [DBN06, LAP19].

Before-and-After [LAP19]. Behavior [BGC^{+}11, DV95, FPH19, GRS^{+}19, KO12, KGAM18, KLD^{+}09, KTE15, LKD19, LKM^{+}18, MGPH06, RCL^{+}15, RKC^{+}16, SPO^{+}12, SPCJL06, TFJ12, ZBMY14]. Behavioral [MZC^{+}16, NT99, VBK17]. behaviorism [FHL10]. Behaviors [CSL^{+}16, GGJ^{+}18]. Behaviour [NTA^{+}19, SMP17]. behind [LXC^{+}17]. Ben [Ano13c]. Benchmark [EHBA11, PFG08]. Benchmarking [WCQ^{+}09]. Bend [AFRS05]. Bending [LLBS17]. benefit [CGJM19]. Benefits [BE09, HF11, WB08]. Benefitting [HAS11]. Bernstein [HWM95]. Bertin [PDF14, PGU12]. Bespoke [RLB19]. Best [De 18b, HJW99a, HJW99b, Lin14c, Meh17, MFS^{+}09, WBH04, De 18b]. Better [CLT^{+}11, LSL^{+}17, LBLE^{+}14, WCD^{+}19, WWS^{+}16]. Between [HGWW18, HKB^{+}19, LDM^{+}18, PML97, BDFM17, GJZ^{+}12, GSL14, KLM04, LPK^{+}13, LKT13, LKM^{+}18, RSRDS16, SH12, TCYM09, WB05, YPI13]. Beyond [BBK^{+}16, BAW16, ERLW18, JD13, KH16, LIRC12, TNT17, WBD14]. Bézier [HWM95, HTF97]. Bezigons [YCZ^{+}16]. Bi [WYP^{+}15]. Bi-Normal [WYP^{+}15]. Bias [WBDS11]. Biases [ACS^{+}18, HUPS14]. Biasing [CH17]. Bicluster [SNR14]. Biclustering [SHR^{+}11]. Biclusters [SMNR16, SZW^{+}19, ZSCC18]. Biconnected [AMA06]. Bidirectional [JWD^{+}14, LHZ^{+}04, ZDW^{+}05]. BiDots [ZSCC18]. Big [BSS^{+}19, FPV^{+}13, NHEM17, PSSC17, Wil18]. Biharmonic [HHQH17, XYGL13]. Bilateral [Wan06, ZFAT11]. Billboard [LWL^{+}17]. Bimanual [MLS18, HWA15]. Binaries [YC14]. Binary [FWT^{+}04, KCOY03, MS08, SMDS14, TDR10, YRP18]. Binary-Space-Partitioned [FWT^{+}04]. Binding [DdL14]. Binning [PQMCR17]. Binoculars [OSB^{+}15]. Bintree [LFR03]. Bioinformatics [SND05]. Biological [BMGK08, KAK^{+}18, KCK^{+}19a, LPK^{+}13, MRSS^{+}12, MCS^{+}08]. Biology [MiS^{+}18]. Biomechanical [KERC09, MGL07, PBO^{+}14, WHM14]. Biomedical [HSSK16, HNR^{+}06, JST^{+}10, NGK18]. Biorthogonal [GGLQ19, WQS07]. Bipartite [CXDR19, SZW^{+}19]. Biped [TLC^{+}10]. Bipolar [BFL06]. Bird [FLF^{+}11]. BirdVis [FLF^{+}11]. Biscale [SHR^{+}11]. Bisection [HJW99b]. Bisections [HJW99a]. BiSet [SMNR16]. Bitcoin [YSZ^{+}19]. BitExTract [YSZ^{+}19]. Bitmap [PMH18]. Bivariate [NASK18, TC17, War09]. Black [MPG^{+}14]. Blended [ATK16, MB19a]. Blending [CWM09b, HSKIH07, HLYL18, KGZ^{+}12, PSG04, PB13, WXJD17]. Blind [ZTP05]. Blinded [DJ18]. Blinding [SFC^{+}07]. Blindness [MXW^{+}13, SBHW11]. Blob [JPLŠ16]. Block [MWC^{+}12, YL06, BPC^{+}10]. Block-Based [YL06]. Blockwise [YSD^{+}17]. Blood [BPM^{+}13, GNBP11, HTP^{+}08, KGP^{+}13, LLL^{+}12, LGV^{+}16, MOJB^{+}19, PH07, vPBB^{+}10, vPBB^{+}11]. Blood-Flow [vPBB^{+}11]. Blowups [SS13a]. Blue [AGY^{+}17, CYC^{+}12, CCS12]. Blue-Noise [AGY^{+}17]. Blur [IAIK16, PLW12, SYK^{+}18, SBE^{+}15]. Bodies [ELF13, GSM^{+}14, KWC^{+}10, ORC07, SSF13, TF06, TZL^{+}12]. Body [AdLH13, CPW^{+}18, CTT^{+}16, Cse10, Cse13, Dan16, Dru08, EVM08, KBS13, KFL^{+}15, LWP^{+}06, MZH^{+}08, MDB18, PDBG18, SM04, SL08, WGR^{+}18]. Body-Centered [Cse10, Cse13]. Body-Swap [PDBG18]. Boltzmann [AEM09, GLX17, LPQF14, WLMK04]. Bone [KSS09, WAWS18]. Bone-Like [WAWS18]. booc.io [SST^{+}17]. Boolean [Ros11, FGF^{+}05, Ros11, Wan11]. Boosting [LXL^{+}18]. Bootstrapping [SKLU^{+}11]. both [CDAF18, YNBH11]. Boundaries [SBSG06, WH18, ZM17]. Boundary [BPB14, CMF12, DBTH07, IK95, LJWH08, LJHY14, RDB^{+}12, SAS16, Wan08, WM13b, WWC^{+}14, WTS^{+}07, YXG^{+}13]. Boundary-Aware [RDB^{+}12, YXG^{+}13]. Bounded [HPAW07, HYB^{+}17, LSY^{+}18]. Bounding [KMKY10, KHM^{+}98, MB18b, SE18]. Bounds [THJ99]. Bowman [Ano14e]. Box [EM06, EVM08, KEP08, Kim13, MPG^{+}14, SK13, MPK^{+}13]. Box-Spline [Kim13]. Boxplot [MWK14]. Boxplots [WMK13]. Boy [DHL09]. BPH [ZZSS10]. Brace [Chi16]. Brachytherapy [LDM^{+}18]. Brackets [TSLR07]. Brain [BZGV14, BJA^{+}19, BSL^{+}14, BTJ^{+}13, EBB^{+}15, JDL12, KSI^{+}96, PFK07, STS07, SAM^{+}05, YSD^{+}17]. BrainGazer [BŠG^{+}09]. Branched [RKG^{+}18]. Branching [WSPVJ11]. BRDF [MDG00, PCS^{+}12]. BRDFs [SAS05, SSR^{+}07]. Breadth [NO97]. Breadth-First [NO97]. Break [PPM^{+}11]. Break-and-Repair [PPM^{+}11]. Breakdown [RC06]. Breast [KLL12]. BReps [HR07a]. Brick [LWYM12]. Bridging [DPW^{+}15, LTM18, LWC^{+}19]. Bring [SJL^{+}18b]. Bringing [HSC08]. Bristle [KMM^{+}13]. Brittle [GMD13]. Broad [TKE16]. Broad-Scale [TKE16]. Broadcast [WCW^{+}16]. Bronchial [STM08]. Browse [SCM^{+}06]. Browser [MMP09]. Browsing [AMA11, BAB^{+}18, BL07, BTC10, JhR10, WMA^{+}16]. Brush [FYY^{+}19, TDM^{+}18]. Brushes [FWR00]. Brushing [HTL13, JBS08, KPV^{+}18, RLS^{+}19, TFH11]. Brushlet [Sel15]. BSP [WM13b]. BTF [LPF^{+}07]. Bubble [CPC09, GSWD18, ZFS^{+}19]. Buckling [KNO15]. Budget [WLHD17]. Buffer [CMFL16, QK04, Zha14, VPF15]. Buffers [WHL16]. Building [MAF11, MWC^{+}12, MP13, PKL^{+}18, SLC^{+}19, SWF^{+}16, VAB12, WFS^{+}16].

Bundled [HEF^{+}14, HPNT18]. Bundles [Hol06, YWSC12, ZCL08]. Bundling [BRH^{+}17, BSL^{+}14, EHP^{+}11, LLCM12, SHH11, SMNR16, SZW^{+}19, vdZCT16]. Burning [BWP^{+}10, LIGF06]. Business [BCH^{+}13, VvWvdL06]. Butterfly [MQV00]. BVHs [Wal12]. by-Example [DV95].

Cable [SL08]. Cache [SJH^{+}07, TDR10]. Cache-Efficient [TDR10]. Caches [YL06]. Caching [GBP07, IWR^{+}18, KGPB05, LY06]. CAD [HC05]. Calculations [SS95]. Calculus [BBK07]. Calibrated [BJM07]. Calibration [BSM06, BMY05, GIMS18, HBKS09, IK15, KBB^{+}12, KLS^{+}18, KV98, MIO^{+}15, PIN^{+}15, PGRS13, RNK^{+}15, RSS14, SH00a, TGW^{+}95, WCW^{+}16]. Calibration-Free [KV98, SH00a]. Caliper [PMT^{+}19]. Call [Ano08c, Ano12e, XXM19a]. Camera [BMY05, CRPH10, Dan16, DH08, GSCI15, HZM13, JWS04, KNR17, LL04, PRA^{+}10, RK17, RNK^{+}15, SM09, TKTN09, TMM^{+}13, WLHD17, WCW^{+}16, XCZ^{+}19b, YLL^{+}12, YGFX19, ZJH^{+}11, Zho16]. Camera-Based [BMY05, JWS04]. Camera-Sampling [LL04]. Cameras [BDF16, CPW^{+}18, KM10, LG12, PSM06, SLG^{+}17, WLT^{+}18b, WXY17, XLC^{+}18]. Can [JAO^{+}14, aKGS11, TKE16]. Cancellation [SRW^{+}16]. Cancer [BSKR19, GHL18, HHO^{+}17, MMB^{+}19]. Canopies [CBLD11]. Cap [XCZ^{+}19b]. Cap-mounted [XCZ^{+}19b]. Capable [RWF19]. Capacities [SLW^{+}10, HSKIH07]. Capacity [ChLYL09, HW12]. CAPE [JNC^{+}15]. Capstone [Ano13t, Cox11, Fra12, Sza10, Bla12]. Capture [BWK^{+}13, BB19, CTGH13, CPW^{+}18, CBL07, HCMTH15, LPG^{+}18, LJH^{+}18, MCP^{+}06, NQX^{+}05, RK17, SLG^{+}17, TAL^{+}07, WLT^{+}18b, XLC^{+}18, XCZ^{+}19b]. Capturing [BB12, HHH16, MDL^{+}17, Tay02]. Car [MF11, TYSN06]. Carbon [BWW^{+}17]. Card [HA17]. Card-Based [HA17]. Cardiac [BPM^{+}13, KGP^{+}13, XSZ^{+}17]. Cards [SOR^{+}09]. Career [Ano10c, Ano11c, Ano13c, Ano13e, Ano14d, Ano15c, Ano16b, Ano16d, Car19, Dil17, Fuc13, Han18, Ros13]. Caricatures [CCM11]. Caricaturistic [RVG06]. Carlo [HKL17, LSPW12, RH19, Sbe97]. carrying [HSKIH07]. Cars [RHJ^{+}16]. Cartesian [EM06, KBVH17]. CartoDraw [KNP04]. Cartogram [NASK18, NAK18]. Cartograms [KNP04]. Cartography [Rot13]. Cartoon [YJL^{+}15, ZCZ^{+}09, ZHF12]. Carving [DZL^{+}14, FG99]. Carvings [LTPH17]. CasCADe [INCB18]. Cascading [AS98]. Case [BvL06, FWD^{+}17, GGZL16, JCRS09, aKS12, LD11b, MRSS^{+}12, PLC^{+}11b, SS06b, dLVvL06]. Cases [BLS04]. CAST [YEII16]. Casting [BPL^{+}19, HBAB14, HK99, KHW^{+}09, LGM^{+}08, LYS^{+}10, RPSC99, SMP11, SF14, SM97, WJ08, WKI^{+}17, vAPP^{+}11]. Casual [BCR19, PSM07]. Catalog [SJMS19]. Categorical [AAMG12, FWD^{+}17, HV13, KBH06, LSS13, SS16, UDSL18, WPS^{+}09, XZM17, ZMZM15]. Categories [BSG18, KPBG13]. Categorization [JF16, KHSI04]. Category [LGG^{+}18, VFR13]. Catmull [GGLQ19, QMV98]. Causal [KIL07, WM16]. Causality [WM16]. Caused [CMS06]. Caustics [HQ07, MB18a, SKP07]. Cave [PFK^{+}08, CDK^{+}17, DJK^{+}06, FWK16, PWHK16, SM12]. CAVE-Like [SM12, FWK16, PWHK16]. CAVE-style [CDK^{+}17]. Celebrating [KHE09]. Cell [CMSW04, GJ10, HPvU^{+}18, IK95, KSG^{+}16, KHDL07, WKME03]. Cellphone [DSC^{+}16]. Cells [MAAB^{+}18]. Cellular [DSG^{+}17, MAK08, MCS^{+}08]. Center [RPHI08]. Centered [Cse10, Cse13, EVM08, GDJ^{+}13, HHO^{+}17, KEP08, LD11b, Mar18, PQF^{+}09]. Centralities [HZM^{+}16]. Centrality [BKW03, CCM12]. Centralized [LWZQ17]. Centric [KOJL^{+}14, LCMH09, SKK^{+}14]. Centroidal [LSPW12, RLW^{+}11, YW16]. Cerebral [GNBP11, GLvP^{+}12, MVB^{+}17, MOJB^{+}19, OJCJP16, BMGK08]. CFD [CMN13, HOGJ13]. CG2Real [JDA^{+}11]. CGLX [DK11b]. Chains [RS12]. Chairs [Ano14o, Ano16p, Ano16q, BSS18, BRS18, CKSB14, IKLW14, Ano13o, Ano15m, CGGR18, CFK12, CLS13b, DLM^{+}12, GGHZ19, vHMM^{+}11]. Challenges [Nie96, WFS^{+}19]. Challenging [PK08]. Change [GGPPS13, KBH^{+}10, LHH16, MXW^{+}13, RCL^{+}15, SBHW11]. Changes [HHWN02, JBMS09, LMZ06, NSH^{+}18, SCL^{+}12, TS07]. Changing [BS16, DHM13a, SFMB12]. channel [WOCH09, WBD14]. Character [ASvdP14, BMST97, DdL14, KJ12, KCPS08, yKL12, LPG^{+}18, MAF11, SKY12, TLC^{+}10, VBK17, VSS08, ZKM18]. Characterise [JAAL18]. Characteristic [HVSW11]. Characteristics [BS02, GLvP^{+}12, RBK^{+}19, YAE07]. Characterization [HEWK03, HEG^{+}17, MY96, Mar18]. Characterizing [CGM^{+}17, FBW16, GGJ^{+}18, GBP^{+}13, GBCG^{+}14, RESC16, WMK13]. Characters [FS19, HK09, MNZ^{+}15, SPW07, ZKM18]. CHARM [FYZ^{+}17]. Chart [KA12, RLB19, RMCW19, TJW^{+}17, WBJ16]. Charticulator [RLB19]. Chartification [CKLL09]. Charts [DDW14, DJ18, HA18, KHD02, KHDL07, TSA14, WWZ^{+}18, LTWH08]. Chase [vW14]. Checks [ASMP17, CLKS19]. Chemical [GBCG^{+}14]. Chemistry [JV09]. Chess [LWL14]. Chief [Lin14d, Ano13n, Ano14n, Ano15l, Ano16n, Ano16o, De 15d, DS16a, DS17b, De 18a, DS18b, De 19b, Ert07a, Ert10d, De 16a, Flo17, Hag03, Kau98, Lin11d, Lin11c, Lin12e, Lin12a, Lin13e, MB19b, Mue19]. Child [BGC^{+}11]. Childhood [JAM^{+}14]. Children [ANR^{+}18, BBC15, PMvWC05]. Chinese [HQQ12, LCC^{+}17, TDM^{+}18]. Choices [SMT13]. Choking [FT13]. Choropleth [ZM17]. Chromium [PAB^{+}08]. ChronoLenses [ZCPB11]. Chunked [SE18]. Circle [YLG^{+}14]. Circles [BSSL19, SZHR11, YR95]. Circular [DBD13, WSPVJ11, Wil12]. CiSE [DBD13]. Citation [BMS17, HHKE16]. Citation-Driven [BMS17]. cite2vec [BMS17]. CiteRivers [HHKE16]. Cities [CWK^{+}07, MDL^{+}17, MPBM^{+}18]. City [AERA14, MDL^{+}19, PY09, ZZM06, FPV^{+}13]. City-Scale [MDL^{+}19]. Clark [GGLQ19, QMV98]. Class [BJY^{+}18, PS12, WCG^{+}19, CCM^{+}14]. Class-Optimal [PS12]. Classical [CQM10]. Classification [AHH^{+}14, BGOJ16, BS02, CDS^{+}12, CZZ^{+}19, CRT04, CM09, GHL18, KKCS98, LZH^{+}07, ME18, MOJB^{+}19, PFP^{+}11, PSPM15, SWB^{+}00, SLM18, SPL^{+}13, TKC17, TLM05, ZM17, vdCvW14]. Classifier [HKBE12]. Classifiers [MQB19, RAL^{+}17, WGZ^{+}19]. Claudio [Ano14f]. Clean [GDN^{+}07]. Clear [Bru17]. Clearance [KMH11]. ClearView [KSW06]. Cleaving [BLW14]. Click [ZTA12]. Clicking [JNC^{+}15]. Clickstreams [CCL^{+}16, LWD^{+}17]. Client [PJ03, SKH^{+}19]. Client-Server [PJ03]. Clifford [ES05]. Climate [DPW^{+}15, JBMS09, KBL19, KLM^{+}08, PDW^{+}14, WLSL17, WPS^{+}16]. Clinical [CLB13, KLL12, NGCL19, SSBC19, WSH^{+}19]. Clip [SEH08]. Clipart [YCZ^{+}16]. Clipped [XDN11]. Clipping [SvW98, WM13b, WEE03]. Clips [SJK^{+}12]. Clique [RFLL18]. clock [SBS16]. Cloning [ZJX^{+}15]. Close [HK09]. Closed [BKA^{+}11, BSG18, KH19, WS01]. Closed-Loop [BKA^{+}11, KH19]. Closely [YHH^{+}19]. Cloth [BW04, CLMO17, KNO15, SZK15, SSIF09, WLT18a, WY19a, ZBO13]. Cloud [APS^{+}14, DNN13, HPJG08, LDX10, LZH^{+}13, MZC^{+}16, WCB^{+}18, YEII12]. CloudLines [KBK11]. Clouds [BHS12, CLC^{+}15, CK10, FFB18, JBS08, LRKC10, MOG11, OHJ^{+}11, OHWS13, TCL^{+}13, YEII16]. Cluster [BDF16, CAN14, GRVE07, KTB^{+}18, LWCC18, LLRR08, NWI17, OJCJP16, SKB^{+}18, SGM08, YHLJ08]. Cluster-Based [CAN14, KTB^{+}18, LWCC18]. Clustered [HBF08]. Clustering [AAFG18, BBD^{+}11, CD19, CK11, CL09, CZQ^{+}08, DDGL07, FKRW17, GPR^{+}01, HSCW13, HOG^{+}12, JNK19, KSDD14, KEV^{+}18, MB18b, NOB16, OK11, PBN^{+}13, PGL^{+}12, SK13, TN14, TRd12, WCR^{+}18, WS09, ZCL08, ZAM11, vLBR^{+}16]. Clusterings [PGU12]. Clusters [CGSQ11, MFS^{+}09, Sil95, ZLC^{+}19]. Clustervision [KEV^{+}18]. Clustrophile [CD19]. Clutter [ED06, ED07]. Cluttered [BK12, SPP^{+}14]. Clutterpalette [YYT16]. CNNs [HTP19]. Co [EPS^{+}15, IC07, IFP^{+}12, LHD18, LCZ^{+}19, TIC09, WHX^{+}19, WXZ^{+}16]. Co-Filtering [WHX^{+}19]. Co-Located [IFP^{+}12, IC07, LHD18, TIC09]. Co-Location [EPS^{+}15]. Co-Occurrence [LCZ^{+}19, WXZ^{+}16]. Coarticulation [DNL^{+}06]. Coating [HL02]. CoDDA [HDSC19]. code swarm [OM09]. Codes [BAW16]. Coding [FM07, HCMTH15, HPJG08, MCA^{+}10]. Coefficient [YYSZ06]. Coefficient-Optimizing [YYSZ06]. Coexpression [NKHC08]. Cognition [LNS08, MWCR06]. Cognitive [BSE^{+}17, BLS15, SBS16, SFR^{+}10, TTR10, ZLB^{+}05]. Coherence [HSR13a, MTM^{+}16, tCMR08].

Coherency [BSL^{+}12]. Coherency-Based [BSL^{+}12]. Coherent [CRH05, FWSL12, GGTH07, GIK^{+}07, GHP^{+}16, Har16, HVSW11, HPC^{+}13, MPT03, RJG17, SP07, SFB^{+}12]. Cohort [GGC^{+}17, KOJL^{+}14, KPS16]. CoLa [DKM06b]. Collaboration [CC12, GLB16, NJJ11, TIC09, YLC^{+}19]. collaborations [IPJT19]. Collaborative [BCB10, BE09, BRNB19, CDK^{+}17, GCL^{+}18, HTL13, IC07, IFP^{+}12, JWL05, KAM^{+}08, LH11, LLL06, MT14, MGPH06, RZP^{+}07, SKY12, ZGI^{+}18]. Collage [FYF^{+}18, LZJ^{+}18, WHFL14, YLG^{+}14]. Collection [IHK^{+}17, JDA^{+}11, WFS^{+}19]. Collections [BKW16, DYW^{+}13, FYF^{+}18, HHWN02, HFM16, KKP^{+}17, LYK^{+}12, PM08, RSOW18, SKK^{+}14, XCZ^{+}19a, ZGW^{+}14, ZLDM16, cKJG^{+}12, dLVvL06]. Collective [BJC^{+}19, HZH14]. Collision [BW03, BW04, CL18, CCW^{+}09, CS08, FS19, GSM^{+}14, GLM06, Hub95, KGS98, KMT14, KGAM18, KHM^{+}98, LCNG14, LKM^{+}18, MBT07, MTS07, PML97, RGF^{+}04, SH12, TCYM09, TF06, ZK07, ZK12, ZK14a]. Collisions [FG99, MNC14, SSIF09]. Colon [GSZ^{+}13, NMGK17, ZMG^{+}10]. Colonoscopy [MGO^{+}19, ZBB^{+}06]. Color [CXM19, CWM09b, DFG^{+}14, DSKA19, EDF11, GR15, HRISI15, IDAK15, KGZ^{+}12, MOF09, MOF10, MS04, MK13c, PH07, PFC18, PMH18, RG95, SLGM09, SMO^{+}13, SGS^{+}19, SS16, Sza18, TdJ14, TdJ15, WW07, WGM^{+}08, WCG^{+}19, XDN11, ZT09, GLS17, HSKIH07, PFC18]. Color-Blending [KGZ^{+}12]. Colored [LAM10, MNZ^{+}15]. Colorgorical [GLS17]. Coloring [BKW10, DHL09, HSL19, MNKT01, ZTZ17]. Colorization [ZGM18]. Colormap [FWD^{+}17, SGS^{+}19, CXM19]. Colormaps [BTS^{+}18, Han16, WTB^{+}19]. Colors [HE99, RC06, TdJ15, TdJ14]. Colour [BCS11]. Column [Gor02]. Combination [TKAM06, WCR^{+}18]. Combinations [JJ09]. Combinatorial [KY06, RLH11, RKWH12]. Combined [GLH^{+}14, MVB^{+}17, SXM17, SM17, SMG^{+}13]. Combiner [LH16, LHH16]. Combing [IBJ^{+}14]. Combining [BWW^{+}12, CHK04, GLK^{+}13, HE99, LHD18, LB19, LGYG12, LNHS16, NSS14, PH07, SJL^{+}18b, TNT17, WKCB07]. Combustibles [YL18]. Combustion [BKL^{+}11]. Comeback [LXC^{+}17]. Command [IG19]. Comments [BNPB13a]. Commercial [BSS^{+}19]. Committee [Ano12h, Ano13l, Ano13j, Ano13k, Ano14m, Ano14l, Ano16h, Ano16k, Ano16l, Ano17f, Ano17g, Ano17p, Ano18c, Ano18f, Ano19d, Ano19e, Ano19m, Sil17a, Sil17b, Sil18a, Sil18b, Sil19, Ano12q, Ano13i, Ano13l, Ano13v, Ano14k, Ano14w, Ano14s, Ano15h, Ano15k, Ano16v, Ano17h, Ano17l, Ano18k, Ano18h, Ano18j, Ano19g, Ano19i]. Committees [Ano11g, Ano12f, Ano12i, Ano12l, Ano15d, Ano16m, Ano13m, Ano13r, Ano14t, Ano14v, Ano16w, Ano16x]. Commodity [NHN07]. Common [HV13, KZW12, LWD^{+}17, MGJH08]. Communicate [BW14]. Communicating [BKW03, MiS^{+}18]. Communication [CSL^{+}16, CLB13, HHO^{+}17, HMSA08, IBJ^{+}14, MT14, MGM09, SDES19, VSS08]. Communities [KKL^{+}16, VRW13]. Community [MMT^{+}14, Nie95, RFLL18, YLZ^{+}13]. Compact [KBK11, LGLR14, NXSL13, PW95, SCYW16, YDG^{+}16]. Comparability [MES^{+}11]. Comparative [ALBR16, BSR^{+}14, BGR06, DLW^{+}17, DBM^{+}06, DDW14, GBFM16, HOGJ13, KBVH17, LCMH09, LSPS10, MHG10, MMDP10, NBW14, PW95, PFK^{+}08, SLK^{+}17b, VP04c, WR11, WS06b, YHR^{+}19, ZSL^{+}16, vLBB16, RSRDS16]. Compare [SZB^{+}09]. Comparing [BHZ^{+}18, FCL09, GTPB19, GGL^{+}14a, GR15, HB13, LKJ^{+}05, LB17, PJ03, PGU12, QM16, RJD^{+}07, TSW^{+}07, TSD09, WOO17, YDK^{+}18]. Comparison [AG16b, BVW^{+}07, BBB^{+}19, CHM11, DPW^{+}15, DJK^{+}06, DLF^{+}09, Gle18, GHC^{+}16, GGC^{+}17, GNDV^{+}18, HJLH19, HFMC12, IC07, KPBG13, KSDD14, KWH19, MMH^{+}13, MLS18, OJEF19, SSMG13, SGB13, SWC^{+}08, TFJ12, VS11, VBV^{+}18, WPS^{+}09, WHR02, WB08, WFG^{+}19, YSD^{+}17, ZLB^{+}05]. Comparisons [KWP01, LBW19, TGH12]. Compatible [BBiA09, BMST97, ER97, YFC^{+}19]. Compensation [GB08b, IHS17, KNO15, LCR16, TIS16, TIK15]. Competent [PW95]. Competing [HFMC12]. Competition [XWW^{+}13]. Compiled [LBG^{+}16]. Complete [HHCL01, ZHF12]. Completion [Hu16a, LSWZ17, LLR18]. Complex [AdLH13, Ano05d, BW14, BBD06, BMST97, BvL06, CFM^{+}13, CVG13, GGJ^{+}18, GK95, GPP^{+}16, GBHP08, HSK14, HBW06, KTC^{+}19, KG06, MGKH09, MGS^{+}14, MK09, Ney98, RFLL18, SFR^{+}10, VMT06, JQD^{+}08, LCM07]. Complex-Valued [GK95, GPP^{+}16]. Complexes [GNPH07, GBP12, GKK^{+}12, GGL^{+}14b, GBP19, SMN12]. Complexity [JWSK07, RBK^{+}15, RMCW19, Sbe97]. Compliant [ATK16]. Component [BMR^{+}19, FWL17]. Components [AMA06, HKG07, SNB^{+}17]. Composable [BTB10, BTHD11]. Composite [ALBR16, Jen12, LPF^{+}19, SWvdW^{+}11, SM06]. Composites [BWW^{+}17]. Compositing [GPC^{+}17, KM10, PMD^{+}07]. Composition [BMWM06, KBVH17, KLJ^{+}09, LRN96, LJWH08, TZC13, TRd12]. Comprehending [ARH^{+}15]. Comprehensible [KMS09]. Comprehension [LR11]. Comprehensive [AR17, TBB^{+}07, WRF^{+}11, XSZ^{+}17]. Compressed [AYRW09, BRSP18, DWvW12, KMKY10, KKSM19, KM16, Lin14f, PR00a, SD11, WAG^{+}12, YL95, eYL07, PR00b]. Compressing [DSKA19]. Compression [BRP19, BDHJ04, BSL^{+}12, CBL07, CKLL09, CM02, CHF95, CHF96, DRMM13, FM12a, GO15, LLPP19, LLT04, LI06, LFR03, MRSS^{+}13, PA06, Ros99, SLW^{+}10, VP04b]. Compressive [SD11, ZFSL19]. Computation [AZM12, BA05, BGT12, CC12, DF96, GGTH07, GNPH07, GBHP08, GBP19, KL14a, KGPB05, LKHW04, PSF09, RKWH12, RLW^{+}11, SMN12, Ste98, TC17, WGS07b, YSS^{+}12]. Computational [BSO^{+}12, CJTM05, GLK^{+}13, KZX^{+}14, PW12, YESK95]. Compute [MFS^{+}09]. Computed [AHR^{+}11, MBW^{+}07, SA19]. Computer [Ano14c, Ano14j, Ano15a, Ano16f, Ano16a, Ano17e, Ano18a, Ano18b, Ano19a, Ano19f, BvdP12, BSM^{+}13, BDC17, CC12, HE12, HQK06, JDA^{+}11, Kas12, KHE09, KS14a, KL14b, LWZQ17, MMCE09, Ota17, PBO^{+}14, RSSA08, SK15, SR00, XSZ^{+}17, Yan18, ZLZY18]. Computing [ABCO^{+}03, BEK10, BTJ^{+}13, CWZ^{+}14, DN13, GBP12, JZLG09, KLL^{+}13, KSY14, MGM09, MZC^{+}16, MTRP10, NJ99, QCT13, RvWT08, RRJH18, Wan08, WTL^{+}09, XXM19a, XHF12, XWL^{+}15]. Concatenating [MCP^{+}06]. Concave [JFBB10]. Concavity [AZC^{+}12]. Concavity-Aware [AZC^{+}12]. Concentration [OKSK16]. Concept [SST^{+}17, USKD12, VI18]. Concepts [HJLH19, Ohl18]. Conceptual [ASW12, DBD18, SHB^{+}14, ZK10]. ConceptVector [PKL^{+}18]. Concordance [WV08]. Concrete [CVG13, FHG^{+}09]. Concurrent [EGH^{+}06]. Condition [PH11]. Conditioning [LLL^{+}10]. Conditions [BMGK08, CMF12, LH11, TNB11]. Conduct [LBW19]. Cone [JWC05, QCH^{+}14]. Conference [Ano11i, Ano12g, Ano12q, Ano13v, Ano14p, Ano14w, Ano14s, Ano15e, Ano15h, Ano16g, Ano16h, Ano16v, Ano17p, Ano18c, Ano18k, Ano19m, BGK11, CR08, CLS13a, EDF08, ILMH12, KKL11, LSCN09, MW13, SL11, Ano11i, Ano13i, Ano14j, Ano17h]. Confetti [PSG04]. Confidence [FKLT10, KTB^{+}18]. Configuration [LH16]. Configurations [Dac11]. Configuring [SDW09]. Conflict [BFL06, Kra16]. Confluent [BRH^{+}17]. Confocal [PFK^{+}08, dLVvL06, WOCH09]. Conformal [GLB^{+}06, HAT^{+}00, PPZ^{+}12, ZMG^{+}10, ZZG^{+}12]. Conforming [GGL^{+}14b]. Conical [AG17]. Conjoint [GMS^{+}07]. Conjunctive [SW13, Wea09]. Conley [CDS^{+}12]. Connected [HKF16]. Connection [SS13b]. Connectivity [AABS^{+}14, BSL^{+}14, CDK^{+}17, JFTW07, JDL12, LL05, MN07, Ros99, SAM^{+}05, TCYM09, TDN^{+}12, VB13]. Connectivity-Based [TCYM09]. ConnectomeExplorer [BAAK^{+}13]. Connectomics [AABH^{+}16, HKBR^{+}14]. Conquer [LYY^{+}16b, ZHLB13]. Consensus [YNM15]. Conservation [LTPH17, SSF13]. Conservative [KS01]. Conserving [CM14, SCKR08]. Considerations [BLB^{+}17, Gle18, LIRC12, TM12]. Considered [CG14]. Considering [KSY16, WP19]. Consistency [PWG17, Wah14, WCB^{+}18, ZCL^{+}19]. Consistency-Preserving [WCB^{+}18]. Consistent [FYP10, IWR^{+}18, JFZ^{+}18, LGQ09, LSS^{+}15, MGPH06, QH18, SKMH14]. Constancy [SSG16]. Constant [HUPS14, SZ12, XJF^{+}08, ZWZD15]. Constrained [HBT14, MZX15, QCT13, RB07, SLGM09, VML97, WWS^{+}18, YLY^{+}12, ZZC11, ZGH^{+}18, ZWBH13]. Constrained-Environment [ZWBH13]. Constraining [YN03]. Constraint [BLS12, DKM06b, GMD^{+}17, Won16, DMS^{+}08]. constraint-based [DMS^{+}08]. Constraints [CCM^{+}13b, LM96, LWCS96, LYY08, MDB18, MWN^{+}19, OBLN17, QH18, TMDO15]. Construct [GTS10, GTS11]. Constructing [CMP09, GHJ^{+}98, HJC14]. Construction [BWC04, CTM^{+}13, CK10, DBTH07, DN12, FSTG16, HJW99a, HJW99b, HSZ^{+}11, INCB18, KSDA16, KCH11, KAK^{+}18, KPS16, KH01, LS02, MB18b, PPT^{+}11, PS12, RLB19, SAR96, TDM^{+}18, VHBS16, Wal12, WBH04, YXM^{+}15]. Constructions [USM96]. Constructive [HR07a]. Consumer [SFL^{+}16, Zho16]. Contact [BDH^{+}18, HK09, IAS19, KFN06, LA11, LL19, MBT07, SM04, SLMA06, TF06, XZB14]. Contacts [AdLH13, CLMO17, SLNB11]. Contagion [VBC^{+}16, WBA^{+}14]. Containment [WLSW08]. Contemporary [VHBS16]. Content [Ano19k, ADP02, CTM^{+}13, CLB^{+}16, GQM^{+}18, HWS17, KEJK19, LLY^{+}13, LJH^{+}18, SBB^{+}18, YLG^{+}14]. Content-Aware [CLB^{+}16, KEJK19, LLY^{+}13, YLG^{+}14]. Content-Based [ADP02, LJH^{+}18]. Contents [Ano10f, Ano11k, Ano12n, Ano12o, Ano13s, Ano15o, Ano16s, Ano16t, Ano17i, Ano17j, Ano18d, Ano18e, Ano14q, Ano14r, CLS13a]. Contest [PFG08]. Context [BBBM18, BMGK08, BSSB10, BGKG06, CL18, CM16, DZMQ16, FWSL12, FNM13, GMS19, GNBP11, GLZR17, KOBH19, KSW06, LDM^{+}18, LBS^{+}19, MB19a, PH07, QWC^{+}09, RGP^{+}12, RMW09, SBB^{+}18, SWS^{+}11, TWSK14, TS08, WBK^{+}08, WLT08, WC11, WWLM11, WP16b, Yan18, YEII16, ZZG^{+}12, vHP09, zBBKN14, Kin10, NH06]. Context-Aware [BBBM18, CL18, DZMQ16, KOBH19, Yan18, YEII16]. Context-Awareness [GLZR17]. Context-Preserving [BGKG06, SWS^{+}11]. Contexts [BBD06]. Contextual [BLE19, LPK^{+}13, WS06b]. Contextualized [BAB^{+}18, WKCB07, WBK^{+}08]. Contextualizing [OKB^{+}19]. Contiguous [KNP04, SKYS14]. Contingency [AAMG12]. Continues [vW14]. Continuity [JDSR^{+}18]. Continuous [BW08b, BTS^{+}18, CCW^{+}09, CRPH10, GPR^{+}01, HW09, KMT14, LT10, LT11, LBL19, OKB^{+}19, RH19, TCYM09, ZK12, vLdL03]. Continuous-Scale [LBL19]. Continuum [BBG^{+}18, EGS03, WM18]. Contour [CD14, DN13, FIB^{+}14, HSCS11, SB06, TFO09, WMK13, ZT09, LLLN^{+}14, TGSP09, PLS^{+}14]. contour-preserving [LLLN^{+}14]. Contouring [PSF09, SJW07, Wu16]. Contours [HTF97, SWC^{+}08, TN14]. Contraction [EBB^{+}15, YSZ04]. Contrast [GUFM15, HRISI15]. Control [EIKS18, FM06, FSME14, GGLQ19, HSK14, HSH10, JAAL18, KLKS10, LGS^{+}11, MY96, RH19, SWR^{+}13, SLB04, SO17, SBE^{+}15, SSS13, SPW07, TKTN09, TIK15, WHR02, WRF^{+}11, WGF08, YWV^{+}19, YLL^{+}12, ZOC^{+}13, ZBMY14, vdCvW14]. Controllable [HTZ^{+}11, HYZ^{+}12]. Controlled [AVP19, BDK98, HHVW96, KZL07, KMT14, OTKS15, WDC^{+}07, ZT09]. Controller [VBK17]. Controls [LNHS16]. Convection [SPO^{+}12]. Conventional [DLW^{+}17, FXG12]. Convergence [GCZL14, LHH16, TRL^{+}19]. Conversion [LKL^{+}15, LKR^{+}18, LLWQ13, YR95]. Converting [GTLH01, HA18, SN97]. Convex [BWC04, FYZ^{+}17, GCT17, KLM04, WPZ^{+}11]. Convey [PBE19]. Conveying [IFP97, IPJT19, KHSI04, HSKIH07]. Convolution [FW08, FC95, LM05, SK98, Sun03]. Convolutional [BJY^{+}18, HKYM17, LYL19, LSL^{+}17, XLZ19c]. Cooperation [LWZQ17]. Cooperative [TIK15, DMS^{+}08]. Coopetition [SWL^{+}14b]. Coordinate [ED06, REB^{+}16, tCMR07]. Coordinated [DCCW08, GR15, KSG^{+}16, KERC09, LKD19]. Coordinates [BBP08, DK10, DK11a, GXY12, HW09, JF16, LT11, LT13, LH13, LTP^{+}05, ML19, NR18, NH06, RLS^{+}19, RSS14, RSRDS16, WLSL17, YGX^{+}09, ZW18]. Coordinating [LHD18, TIC09]. Coordination [HZL^{+}19, MT14]. COPE [LCZ^{+}19]. Coping [DSP^{+}17]. Copula [HDSC19]. Copula-based [HDSC19]. Copyright [ZTP05]. Core [AGL06, HKC^{+}12, HEFR18, HBC12, KHS^{+}18, LP02, LCDP13, SHC^{+}09, USM97, Wal12, YSGM05, ZK14a]. Corelines [GT19]. Cores [GT14, UDSL18, WSTH07]. Corneal [PIN^{+}15]. Corneal-Imaging [PIN^{+}15]. Coronary [MSSD^{+}08, TBB^{+}07, TBB^{+}08]. Corpora [CSL^{+}10, CLWW14]. Corpus [CWDH09]. Correct [FST^{+}14, HN13]. Correcting [MT05]. Correction [Ano96a, Ano96b, HRISI15, IK15, JHW^{+}14, LL19, SOS^{+}17, SRCP03, WLH^{+}12, XDN11]. Corrections [BGM^{+}08, HJW99a, KS00a, MOF10, PR00b]. Corrector [BS95, SFBP09]. Correlated [JY17, VS11]. Correlation [HYFC14, KH16, LZJ^{+}18, YHR^{+}19, ZMZM15]. Correlation-Preserving [LZJ^{+}18]. Correlations [STS06, YPI13, ZW18]. Correspondence [AW03, ADDG12, LNHS16, XA09]. Corresponding [NMGK17]. Cosine [Cse13]. Cosine-Weighted [Cse13]. Cosmic [SPN^{+}16]. Cosmological [FSW09, NJB07]. Cosserat [ST09]. Cost [BSE^{+}17, CGJM19, DRHK07, KM10, RZP^{+}07]. Cost-benefit [CGJM19]. Costs [Lam08]. Coulomb [JR07]. Countermeasure [BDB^{+}16]. Counting [BLS04]. Coupled [HZ13, LTKF08, RSM^{+}16]. Coupling [BLE19, BTT09, CMK15, VHLL14]. Courses [CCL^{+}16]. Court [WCW^{+}16]. Cover [Ano10e, Ano15f, Ano15g, Ano16i, Ano16j, Ano11f, Ano11h, Ano11j, Ano14i]. Coverage [Elb95, IDA^{+}14, KM16, STM17]. Covering [RKG^{+}18]. CoViCAD [TBB^{+}07]. Cox [MPK^{+}13]. CPU [SHC^{+}09, WJA^{+}17, WWW^{+}19]. Crafted [Gle13]. Crafting [PDF14]. Crease [STS10, TKW08]. Creased [KMDZ10]. Creases [KSSW09]. Create [BDFM17, JK16]. Created [PPP12]. Creating [GLS17, KP05, KHA10, SVK^{+}07, SK16b, SJK^{+}12, TSB^{+}05]. Creation [PMT^{+}19, SLA^{+}09, WSYM17, YT02]. Creative [CLEK13, GDJ^{+}13, KGD^{+}19, RRJH18, WXJD17]. Creativity [Ger17]. Creature [LGS^{+}11]. Credits [Ano10e]. Crepuscular [KKMS11]. Crest [SF04]. Criteria [FPH^{+}08, HZM13, LFA^{+}16, PSTW^{+}17]. Critical [BDH^{+}18, FFST19, RKG^{+}11, SRW^{+}16]. Critique [AJDL08]. Critique-Based [AJDL08]. Cropping [LWZ^{+}18]. Cross [CPW^{+}15, GGC^{+}17, KZW12, LB17, MGL07, NJJ11, SKYS14, Wea10, WPZ^{+}11]. Cross-Cutting [CPW^{+}15]. Cross-Filtered [Wea10]. Cross-Organizational [NJJ11]. Cross-Parameterization [MGL07, WPZ^{+}11]. Cross-Sectional [GGC^{+}17, SKYS14]. Cross-Sections [LB17]. Crossing [BGC^{+}11, JOR^{+}19, SS08]. Crowd [DGAB16, FR13, LCNG14, PvdBC^{+}11, RE14, WOO17]. Crowded [KAK^{+}18, KCK^{+}19a]. Crowds [LKS^{+}19, MT01, SMTT^{+}17]. Crowdsourced [BLIC19, LCL^{+}19, WGS^{+}13]. Crowdsourcing [GJC^{+}17, MDF12]. Crystal [JKM06]. Crystalline [DK13]. Crystals [MJK06, SPCJL06]. CSG [HR07a]. CST [HR07a]. CT [HLRS^{+}08, HKG07, RHR^{+}09, XSZ^{+}17]. Cube [CVC^{+}12, KDA^{+}09, MLKS18, XHL18]. Cubemap [WWL07]. Cubes [DSS^{+}09, LB03, Nie03, STH03, WFW^{+}17]. Cubic [Cse10, Cse13, EVM08, KEP08, PQF^{+}09, POD^{+}13]. Cubical [SVAC12]. Cubist [CH03]. CUBu [vdZCT16]. Cue [LKR^{+}18]. Cueing [LSB^{+}16, TCM06]. Cues [BW14, KOCC14, LHC10, LDFZ14, LKM^{+}18, PB16, SLMA06, VBV^{+}18, WHA07, WS06b, IPJT19]. Culling [BMA^{+}19, CL18, GLM06, KS01, MSHC99, SG99, TCYM09, ZK12]. Cultural [WFS^{+}19]. Cumulative [GWK12]. Curation [SSBC19]. Curation-Based [SSBC19]. Current [Sat13]. Curvature [AT05, BSSL19, GWBO12, GCZL14, JFBB10, KWH19, MOG11, NSE^{+}12, ONL^{+}12, WT10b, YKL^{+}08, ZG12, ZBB^{+}06]. Curvature-Adaptive [ZG12]. Curve [AHL^{+}13, BSL^{+}12, CSM07, JM10, LCMH09, LLBS17, LGS12, MWK14, ONL^{+}12, OK11, RvWT08, WL08, ZWR14]. Curve-Centric [LCMH09]. Curve-Skeleton [CSM07, WL08]. Curve-Skeletons [LGS12]. Curved [AFRS05, AMB^{+}13, CRPH10, LLBS17, WGC^{+}08, XRP^{+}12, vGMSW15]. Curves [BSH^{+}16, DMC15, Elb95, HVSW11, KBI^{+}18, MWK14, MYM16, MM08, NJ99, WT10a, WFG^{+}19, ZDZ18]. Curvilinear [FC95, HHM14, HK99, STYC12, ZCW19]. Curvilinear-Grid [HHM14]. Curving [IFP97]. Custom [BLO^{+}05, SLA^{+}09]. Custom-Tailored [BLO^{+}05]. Customer [WWL^{+}10]. Cut [NKH11, LLC15]. Cut-Surface [NKH11]. Cutouts [LSR^{+}13]. Cuts [DGW11, WWB^{+}13, XTY^{+}11, ZZBW08]. Cutting [CPW^{+}15, FG99, GTLH01, SBW17, WZW^{+}05, ZZC11]. Cyberinfrastructure [MEV^{+}14]. Cybersecurity [Ano14p]. Cyclist [BGC^{+}11]. Cyclone [LPCRH19, VMN^{+}19]. Cylinder [LZH^{+}13]. Cylindrical [AL11, AG17]. CyteGuide [HPvU^{+}18].

D [CMK15, GW13, MY14, NHYY18, SK16a, SW17, AL06, AJDL08, AHKMF11, ARH^{+}15, AHR^{+}11, AVP19, APV^{+}15, AJ19, BSB^{+}18, BGCS17, BB09, BF01, BBiA09, BIAI17, BPM^{+}13, BTB10, BTHD11, BC12, BHSH15, CWZ^{+}14, CTGH13, CWL12, CPW^{+}18, ChLYL09, CC08, CZN^{+}11, CKW^{+}12, CZZ^{+}19, CFHH09, CCM11, DDW14, DHL09, DMR04, DBW11, DRHK07, DNP07, DHM13b, DMC15, ERL^{+}13, ET08, FW08, FYTL19, FCL09, GCL^{+}15, GRT17, GSCI15, GCL^{+}18, GIS03, GSS^{+}19, GJR^{+}14, GPP^{+}16, HDBC15, HE06, HLL97, HLRC^{+}12, HM10, HLYL18, IFP97, INCB18, IFM14, JSG03, JK16, JNC^{+}15, JDL09, JBS06, JCWD14, KRW19, KSG^{+}16, KZL07, KHS^{+}19, KCA16, KLG^{+}16, KGP^{+}13, KCK^{+}19a, KYT^{+}18, KKKW05, LYY^{+}16a, LKJ^{+}05, LS07a, LWS^{+}17, LRN96, LH03, LDH09, LKL^{+}15, LHC16, LYL19, LJL04, LKR^{+}18, LZZ^{+}19, LG13, LDN11, LJZ12, LL05, LY06, LB17, LLCD11]. D [LCS^{+}12, LGS12, LFR03, LODI16, LSM03, MS08, MWSJ14, MW99, MCHM10, MAST16, MCG12, MSA17, MH10, MJK06, MDS16, MES^{+}11, MWC^{+}12, NWHWD16, NB95, NQX^{+}05, NHYY18, OHWS13, OA11, dJOBNM17, OHH06, PDRK19, PQMCR17, PBO^{+}14, PZLZ17, PK08, PLW11, PLW12, PJ03, PGT^{+}08, PZ07, PH07, PMW13, PUNI11, PMT^{+}19, PGI^{+}17, QCT13, QT96, QY07, QWC^{+}09, RSBB17, RHZN11, RKG^{+}11, RT12, RKZZ19, RF11, SW12, SZB^{+}09, SWB^{+}00, SBV^{+}11, SHM^{+}07, Sel15, SKYS14, SPP^{+}14, SPB96, SWW^{+}15, SMN12, SXX^{+}19, SW97, SWCR15, SRW^{+}16, SGJM18, SKW^{+}11, SBW17, SKH^{+}19, SS13b, TKTN09, TCL^{+}13, TNT17, TMWS13, TC09b, TWHS05, TZL^{+}12, TKAM06, TLC^{+}10, VCP08, VAB12, VW12, Vas16, WWYS04, WLL^{+}05, WH09, WM13a, WWC^{+}14, WM18, WTL^{+}09, WQZ^{+}18, WB08, WSE07, WFG^{+}19, WXY17, WCB^{+}12, Won16, WLDW11, XCZ^{+}19b, XFZ^{+}19, XLZ19c, YOS13]. D [YPRI17, YFC^{+}19, YLSL11, YSL^{+}13, YJL^{+}15, YL95, YHH^{+}19, YSI^{+}10, YEII12, YEII16, ZTP05, ZH07, ZWZ^{+}13, ZWR14, ZTZ17, ZGH^{+}18, ZPP05, Zho16, ZCFL15, vPBB^{+}10, vPBB^{+}11, vWN04, vLBB16, zBBKN14]. D-NURBS [QT96]. D-Snake [LG13]. D [?]to-Leimkuhler:2018:PRT. D.O.F. [TAK^{+}05]. D3 [HA18]. DAG [GK07, IG19]. Damage [HTP^{+}08]. Damping [SRKL19]. Dance [FXG12]. Daniel [Ano11d]. Dark [KHA12]. Dashboard [BSKR19]. Dashboards [McK09, SCB^{+}19]. Data [AMM^{+}08, AJ17, ACS^{+}18, AAMG12, AHH^{+}14, AZL^{+}19, ARH^{+}15, ARL^{+}17, AA11, AAH^{+}13, AAFW17, Ano09b, ASMP17, ASE16, AJ19, AHL^{+}13, BSH^{+}16, BPP^{+}16, BE18, BEW95, BSS^{+}19, BKL^{+}11, BWK^{+}13, BIAI17, BAAK^{+}13, BGB15, BPL^{+}19, BDSW13, BBP08, BBB^{+}19, BDD^{+}16, BBB^{+}12, BWT^{+}11, BG06, BGKG06, BAW16, CDC^{+}07, CLG16, CLZ^{+}18, CWT^{+}08, CDR^{+}18, CBL07, CZC^{+}15, CYW^{+}16, CXR18, CHW^{+}18, CM16, CXM19, CLC^{+}15, CM02, CMPS97, CvW11, CMN13, CML^{+}12, CMSW04, CSC07, CL11, CJTM05, CWRY06, CMP14, DC17, DSKA19, DFD^{+}14, DR08, DS16b, EGG^{+}12, EM06, EMdSP^{+}15, EBRI09, FPB17, FKLT10, FPV^{+}13, FM12a, FAW10, FMH08, FSE12, FWR00, FIBK17, FMST96, FZC^{+}07, GPL^{+}11, GHGM06, GHA^{+}08, GBWI17, GABJ08, GS14, GIK^{+}07, GWBO12, GK95, GSG96, GW11, GWP^{+}16, Guo09, GZL^{+}14, GZ14, GPP^{+}16, GXZ^{+}18, GJG^{+}19, HLRS^{+}08, HBJP12, HTL13]. Data [HEFR18, HDSC19, HR07b, HHM14, HKB^{+}19, HB13, HHH16, Hol06, HM10, HCMTH15, HZM13, HZH14, HZL^{+}19, HZM^{+}16, HHG14, HEF^{+}14, IWR^{+}18, IML13, IYIK04, JFSK16, JEG12, JER16, JBS08, JYC^{+}10, JA18, JhR10, JH09, JFY16, JY17, KSH03, KNKH19, KPHH12, KGS^{+}08, KERC09, KLM^{+}08, KMDH11, KH13, Kei02, KHDL07, KKSS13, KLC08, KWP01, KLK^{+}09, KSL^{+}17, KJW^{+}18, KRH18, KOJL^{+}14, KLG^{+}16, KM96, KHPS07, KGP^{+}13, KBH06, KSDD14, KPB14, KS02, KH01, KGJ09, KGG^{+}12, KGZ^{+}12, KW13, KHH^{+}16, KXW^{+}18, LBG^{+}16, LSSB12, LBS13, LSJ^{+}15, LCC^{+}17, LWG05, LHD18, LBW19, LPF^{+}19, LWS97, LS02, LS09, LSS13, LKS13a, LT16, LSPS10, LSS^{+}11, LFP07, LFLH07, LFH08, LPLT11, LI06, LLB^{+}06, LLRR08, LLCD11, LSS09, LS16, LMW^{+}17, LXR19, LPK^{+}16, LPCC17, LODI16, DSC^{+}16, LGLR14, LMC02, LJH^{+}18, MCP^{+}06, MS08, MD12, MV06, MMH^{+}13, MKH12, MGKH09, MHD^{+}18]. Data [MMYK06, MGB^{+}19, MMDP10, MBH^{+}12, MLKS18, ME11b, MCS^{+}08, MFS^{+}09, MNKT01, MVN^{+}19, NN11a, NN11b, NKHC08, NR18, NGCL19, NASK18, OJ15, ODH^{+}07, OHWS13, OKB^{+}19, PSSC17, PFP^{+}11, PSPM15, PBO^{+}14, PCY08, PMN^{+}14, PLS^{+}14, PGT^{+}08, PBE19, PSN10, PH07, PHJ^{+}10, PWHK16, PW12, PSM07, PFK^{+}08, POM^{+}09, RESC16, RBS^{+}18, RGK^{+}13, RML12, RLA^{+}13, Rob98, RP12, RL08, RZNS04, RGC^{+}14, RB18, RSS14, SMDS14, SKBE17, SSRE18, SE17, SSEW19, STS06, SXM17, SFBP09, SWR^{+}13, SKS12, SGS^{+}19, SFP^{+}19, SR17, SK16b, STS07, SH10, SSH14, SHH11, SS06b, SBSG06, SS16, SvdBLM11, SW13, SCL^{+}12, SLC^{+}19, SS19, SK00, SDES19, SJL^{+}18b, SGAS16, STH03, SSL^{+}12, TJW^{+}17, THM15, TIW^{+}19, TAE^{+}11, TTR10, TdJ14, TdJ15, TSB^{+}05, TBHC16, TCM10, TSAA12, TS07, TFH11, TLLH12, TSH^{+}14, TKBH17, TLM05, VH16]. Data [VvWvdL06, VJC09, WBJ16, WDC^{+}18, WLJ^{+}12, WYM08, WM08, WSPVJ11, WWLM11, WLL^{+}12, WZvdW13, WYL^{+}14, WX17, WOO17, WM18, WCC^{+}18, WWW^{+}19, WLF^{+}19, WHLS19, WRF^{+}11, WK06, WBP07, WZ08, WKZL04, WSLL12, Wil18, WGS^{+}13, WJD17, WMS98, WFS^{+}19, WMGE12, WS09, WXC^{+}08, WXZ^{+}16, WLS^{+}18, XNT11, XXM^{+}19b, YEB18, YHW^{+}07, YLX^{+}12, YL95, YS17, YXG^{+}10, YRWG13, ZZD14, ZCL08, ZLD^{+}14, ZMZM15, ZGB^{+}17, ZWC^{+}18, ZXM10, ZG06, ZGHG11, ZMT^{+}19, ZGM18, ZD18, dLVvL06, dOL03, tCMR07, tCMR08, vTRvdM97, vLFR17, CJR07, FHL10, HSKIH07, THT19, WOCH09, WBD14]. Data-Driven [ARL^{+}17, CDR^{+}18, CXM19, HZL^{+}19, KSL^{+}17, KGZ^{+}12, LCC^{+}17, LFP07, MD12, PLS^{+}14, SSH14, ZLD^{+}14, ZGM18, tCMR08]. Data-Flow [SWR^{+}13]. Data-Parallel [ZGHG11]. Data-Rich [GBWI17]. Database [CWZ^{+}14, CTM^{+}13, MSA17]. Databases [STH02]. DataClips [ARL^{+}17]. Dataflow [SRHH16, WSW^{+}18]. Dataset [MHG10, WDSC07]. Datasets [ALM11, BRT12, BTC10, FSW09, GGL^{+}14a, Har16, HE99, JBH^{+}09, LBS14, LKS13b, MZFM98, NJB07, PLS^{+}14, PSR17, PHF07, SZB^{+}09, SYM14, SGM08, WFW^{+}17, XYC^{+}18, ZCCB13]. Dawn [Bai13]. Daylighting [SYYC11]. DCB [CLCQ12]. DCB-Spline [CLCQ12]. DCT [YL95]. DCT-Based [YL95]. De-Biasing [CH17]. Deadeye [KK19]. Dealing [GNCM^{+}16]. Debugging [SGB^{+}19, WTL^{+}09]. Decadal [KBL19]. Decade [KBB^{+}18, LMW^{+}17]. Decal [RASS17, RSA^{+}19]. Decal-Lenses [RSA^{+}19]. Decal-Maps [RASS17]. Decals [RASS17]. Decider [AJ19]. Decimation [CE01]. Decision [BTC13, DBD18, KDX^{+}12, KWS^{+}14, KHZR18, MLMP18, PSTW^{+}17, SOL^{+}16, ZCD19]. DecisionFlow [GS14]. Decisions [UKW19]. Declarative [HB10, JVDF19, SRHH16, SRM19]. Decomposition [BNPB13a, BNPB13b, BPB14, CML^{+}07, GMD^{+}16, GH95, HHG14, HPNT18, KL96, LBD13, LGQ09, PPL^{+}10, RNL09, Rob98, RF11, SS08, YBW^{+}19, ZGH^{+}18, ZCFL15]. Decompositions [BRP19, CMLZ08, ER97, SZ12, Szy13]. Decompressible [KXW^{+}18]. Deconstraining [YN03]. Decorating [ZDW^{+}05]. Decoration [BKRE19]. Decoupling [BB09]. Dedicated [GUO00]. Deep [BTJ^{+}13, CCJ^{+}19, GL17, HKPC19, KAKC18, KTC^{+}19, LSL^{+}17, LSC^{+}18, PHV^{+}18, SLGM19, WGZ^{+}19, WGSY19, WSW^{+}18, XLZ19c]. Deep-Learning-Assisted [CCJ^{+}19]. Deeper [HDR^{+}13, YaKSJ07]. DeepEyes [PHV^{+}18]. DeepTree [BHP^{+}12]. DeepVID [WGZ^{+}19]. Defect [SPCJL06]. Defects [MJK06, RGK^{+}13]. Deferred [TCM10]. Deficiency [MOF09, MOF10]. defined [JJ09]. Defines [AG17]. Definition [WR11]. Deformable [ADDG12, BW04, CGD97, DGW11, DTT^{+}17, DDKA06, EGS03, FS19, FST^{+}14, HBAB14, KL14a, LSH07, LKT13, NWI17, RLK19, SO17, SSF13, TL07, TCYM09, WPG05, WB05, YGV^{+}13, ZK07, ZWW^{+}12]. Deformation [AL11, BS08, BDD^{+}16, CK05b, CCM11, CSC07, DCKY02, FG99, GD01, HHM14, HTZ^{+}11, HYZ^{+}12, KSNY17, LZD13, LG13, LL14, LXW^{+}18, MVB^{+}17, SVAC12, SM06, TC09a, TWSK14, TEC^{+}16, WCB^{+}12, YHMY08, YXG^{+}13, ZZL^{+}15]. Deformations [CDA99, FvdPT97, HSSK16, HHZH17, KJ12, SSH14]. Deformed [KBKG07]. Deforming [GWBO12, MBT07, NWI17, WLT18a, ZK12]. Degenerate [DdL14]. Degradations [KLYE13]. Degree [AHSS14, JWC05, LHJ^{+}18, ORC07, vHP09]. Degree-of-Freedom [JWC05, ORC07]. Degree-of-Interest [AHSS14, vHP09]. Degrees [CMHL11, HA06a, OHH06]. Degrees-of-Freedom [CMHL11]. Degrees-Viewable [OHH06]. Delaunay [FHSW13, MS08, QCT13]. Delay [DL12]. Delivery [CYB08]. Demand [SA19, VHBS16, YAE07, vHP09]. Demands [BLS15, BPC^{+}10]. Demonstration [SKBE17]. Demonstrators [HK16]. Dendrogram [CMP09, WZK12]. Dendrogram-Matrix [CMP09]. Denoising [ALMF19, CS18, FYP10, GJR^{+}14, LC10, LDC16, SRML07, WYP^{+}15, WHX^{+}19, YRP18, YRP19, ZWZD15, ZFAT11]. Dense [BHST17, CWL12, DMC15, DRMM13, EBRI09, GCL^{+}18, HQ12, IFM14, MK13b, OKI15, OKSK16, RGE19, WB08, YYR17]. Density [JVDF19, KLySK12, LBH18, NE04, OHJ^{+}11, SWvdW^{+}11]. Dental [WZW^{+}05]. Dependent [BGM^{+}07, BGM^{+}08, CWK^{+}07, CC08, EBRI09, GLRH13, GKT^{+}08, GY02, HMTR19, HSH10, IAIK16, KRHH11, KL96, KNO15, KKSM19, KGG^{+}12, LHV06, LP02, LPG15, MCHM10, NSE^{+}12, PD04, SBV^{+}11, TLQ^{+}08, TWHS05, TEC^{+}16, TLS17, USE13, VCP08, WC09, WC10, YSGM05, ZEC08, Zhu05]. Depiction [BW11, BBIF12, CVG13, VPB^{+}11]. Depictions [PSM07]. Deployable [ZWC^{+}16]. Deprivation [BDB^{+}16]. Depth [APW16, BE06, BG07, CMP14, Dan16, EBRI09, IHK05, IMS15, KPR^{+}15, KOCC14, KSTE06, KLM04, KJW^{+}14, LDH09, LKC09a, LKL^{+}15, LH16, LHH16, LDN11, NRS15, PBK^{+}12, RWBF18, SLG^{+}17, SEA09, SJK^{+}07, SSE15, VF13, YYFX18, YGFX19, ZDJ^{+}09, ZWM13, ZHF^{+}19]. Depth-at-Field [RWBF18]. Depth-Dependent [EBRI09]. Depth-Fighting [VF13]. Depth-Fused [LDH09]. Depth-Inferred [ZDJ^{+}09]. Depth-of-Field [IMS15, LKC09a]. Depth-Ordering [ZWM13]. Depth-Perception [BG07]. Derivative [ySKK07]. Derived [OJ12, Rot13, TBB^{+}08, WT10a, GDKB17]. Deriving [EGS03]. Descending [NMN^{+}18]. Descent [ZPG19]. Description [CXDR19, SAR96, VC17]. Descriptions [LB19, MB01]. Descriptor [CZN^{+}11]. Descriptors [CPG09, DNP07, EHBA11, UMW^{+}12]. Design [AJDL08, ARRC11, AS05, BB12, BSKR19, BEJK12, BISM14, BLB^{+}17, BGK11, BM10, CKW^{+}12, CZZ17b, CWQ^{+}07, CLEK13, CJTM05, DPW^{+}15, DBH14, EF10, FFB18, FCZ15, GS08, GKL^{+}13, GWFI14, GJZ^{+}12, GDJ^{+}13, HKBR^{+}14, HHO^{+}17, HA17, HKR^{+}08, HA06b, HB10, HLRC^{+}12, HZM13, HLB^{+}18, JHKH13, aKGS11, KW14, KAM^{+}08, KSL^{+}17, KS14b, KBVH17, KHLM17, KOBH19, KPR^{+}14, KOK^{+}19, LJX^{+}10, LTM18, LWS^{+}17, LIRC12, LIM^{+}12, LCS^{+}12, LD11b, LLY06, MLMF12, ME18, MJW^{+}13, MRSS^{+}12, MMAM14, MKT^{+}18, MPOW17, MMMY97, MWN^{+}19, Mun09, OM09, OK11, PFW12, PWHK16, PB16, QT96, RHY14, RHR16, RKG^{+}18, SKK06, SVAC12, SFB^{+}12, SHS11a, SNHS13, SMM12, SYYC11, SOL^{+}16, SD12, SSL^{+}12, SNR14, Sza18, TM12, TLD^{+}12, TSW^{+}07, WGM^{+}08, WZK12, WASQ18, WKD19, WQ07, YCLJ12, Zag96, ZHT07, ZM13, tCMR07, vLFR17, WBD14]. Design-Sheet [RHR16]. Designing [Chi16, EMRY02, HN13, Kei00, KKL^{+}16, dJOBNM17, PHV^{+}18, SKK^{+}14, TKBH17, WK06, ZWC^{+}16]. Designs [HFMC12, OAH14, RHR16, WLF^{+}19]. Desktop [JD13, LvL12, PFK^{+}08, TMM^{+}13, TBR^{+}12]. Destination [AAFW17, GZ14, YDJ^{+}19, ZFL17, ZMT^{+}19]. Detail [AWHS16, BHSH15, CMP09, GLX17, HSH10, HYZ^{+}12, KCJ^{+}10, MB01, NM13, NC07, WGS07a, WY19a, XESV97, Zhu05, ZBDS12, vdEvW14, DMS^{+}08]. Detail-Preserving [GLX17, HYZ^{+}12]. Detailed [KLCK17, SSH14, ySKK07]. Detailing [YYT16]. Details [BBD06, LBS^{+}19, NDS10, SA19]. Details-First [LBS^{+}19]. Details-on-Demand [SA19]. Detecting [AG17, KSDD14, SXX^{+}19, TN13, WBDS11]. Detection [BW04, BDJ14, CLZ^{+}18, CCW^{+}09, CRT04, CS08, FS19, GLvP^{+}12, GSM^{+}14, GSL^{+}17, GRVE07, HLRS^{+}08, HQK06, Hub95, JWS04, JPLŠ16, KHS^{+}18, KGS98, KMT14, KHM^{+}98, LWZ^{+}16, LZH^{+}13, MBT07, MRSS^{+}13, MJK06, MTS07, OSBM14, PLW11, PML97, SBJ^{+}10, TCYM09, TN14, WRM^{+}10, WWC^{+}14, WX17, WSYM17, WTB^{+}19, WS01, Won16, XXM19a, XXM^{+}19b, ZK07, ZK12, ZK14a, ZLK^{+}18, ZBB^{+}06, JQD^{+}08]. Detector [BW03, DNP07]. Deterministic [BMA^{+}19]. Developable [TC09a, Wan08]. Developing [BW11, KFS^{+}19, NTT^{+}19]. Development [KMSB14, LKK17, LFA^{+}16, MYI13, TAK^{+}05]. Device [EPS^{+}15]. Devices [BJM07, BMWW18, DFG^{+}14, HTP^{+}08, KPR^{+}15, LS07a, LHD18, RBDG15, RJG17, SOS^{+}17, YC14, YLC^{+}19]. DIA2 [MEV^{+}14]. Diabetes [ZCD19]. Diagnosis [BGP^{+}11, KHZR18, LXL^{+}18, MSSD^{+}08, OIR^{+}17, SSMG13, WGZ^{+}19, ZWM^{+}19]. Diagnostics [XMRC17]. Diagram [CdOKRV09, YBZW14]. Diagrammatic [JhR10]. Diagramming [LFW^{+}19]. Diagrams [BTC10, BKH^{+}11, CLB^{+}16, CRB^{+}05, JRHT14, KLMA10, MPWG12, RD10, RZP12, RSFH14, SSKB14, SAS16, SZHR11, SRHZ11, VCP08, Wil12]. Dialogue [PMvWC05]. Diamond [WD09]. Dice [EDF08]. Dichoptic [KK19]. Dichromats [CCB11, KOF08, SMO^{+}13]. DICON [CGSQ11]. Diderot [KCS^{+}16]. DiffAni [RM13]. Difference [RM13, Sza18, WCJ06, YC14]. Differences [TSD09]. Different [BC18b, FTES13, GJZ^{+}12, HOT98, SLW^{+}10, TAL^{+}07, YL18]. Differential [KHW^{+}09]. Difficulties [HAS11]. Diffraction [RSR^{+}18]. Diffuse [NSS14, SVGR16]. Diffusion [CLS^{+}12, CYZ^{+}09, DPR00, FT09, GWK12, GSZ^{+}13, GKL^{+}16, HM10, HQ13, HLD^{+}08, KLLR07, KWH00, KW06, PT17, STS07, STS10, WW07, WDC08, WLY^{+}14, ZDL03, ZSL^{+}16, ZDZ18, ZHL^{+}09, vAPP^{+}11]. Digital [AB06, DiV15, JSG03, KSI^{+}96, MBW^{+}07, PMCS11, SCT06, SE18, ZSTR07]. Dimension [BWC04, DWF^{+}19, GXY12, KW10, STM17, SMT13, WCR^{+}18, YRWG13]. Dimensional [ALM11, ADP02, BSL^{+}14, CCAL12, DAW13, EHS13, GPL^{+}11, GBPW10, GS14, GNP^{+}06, HEWK03, HHM14, HNR^{+}06, HRD^{+}19, IHR01, JDL12, KRHH11, KPB14, LMZ^{+}14, LT16, LMW^{+}17, MBS^{+}04, NM13, NE04, OHJ^{+}11, SS06b, SLB04, TAE^{+}11, TFH11, TLLH12, TKBH17, TLM05, WSPVJ11, WM13a, WM18, WAG06, WXC^{+}08, XYC^{+}18, YRWG13, ZWC^{+}18, ZLC^{+}19, FZC^{+}07, GNPH07, LLRR08, LPK^{+}16, ZWR14]. Dimensionality [JJ09, KC04, SZS^{+}17, SDMT16, WFC^{+}18]. Dimensions [LBGV13, TFH11, YHW^{+}07]. Diminishable [AIS18]. Diminished [KSY16, STB18]. DimpVis [KC14]. DimReader [FGS19]. Dinosaur [NTT^{+}19]. Dipoles [GBM^{+}12]. Direct [AWC10, ACTM12, BTT09, BMWM06, BKW08, CR08, CWM^{+}09a, CCB11, CWC^{+}06, DWB^{+}06, FDPH17, FE17, GPC^{+}17, GKK^{+}12, HLY10, HXF^{+}15, JKRY12, KH19, KWP01, KWH00, KC14, KJL^{+}12, LR11, LLY06, MCK12, Max95a, OKB^{+}19, RE01a, SE17, SF14, Sel15, WQ07, WKI^{+}17, YCZ^{+}16, YSI^{+}10, ZWM13]. Direct-to-Indirect [ACTM12]. Direct-Touch [YSI^{+}10]. Directable [JFBB10]. Directed [CHK04, DKM06a, DRMM13, MKT^{+}18]. Directing [PvdBC^{+}11]. Directional [SHH11, WWYS04, WC13, ZHZ15]. Directionally [AM13]. Directionally-Varying [AM13]. Directivity [MAKM14]. Directly [HBM^{+}13]. Directness [BIRW19]. Dirichlet [HLG^{+}14]. Disambiguating [HSL19, SC15]. Disambiguation [SWY^{+}17]. Discerning [LZZ^{+}19]. Disciplines [GJZ^{+}12]. Discontinuities [LT10]. Discontinuous [SCKR08]. Discourse [ASW12, LSB^{+}16, ZCCB12]. Discovering [CYW^{+}16, TWC^{+}18]. Discovery [BVV^{+}19, DCCW08, Eic00, IDA^{+}14, PLS^{+}14, SS06b]. Discrete [AT05, AHK^{+}17, AJ97, CS08, EDvW19, HKKS18, IHR01, JKLG08, KCOY03, LS13b, LLT04, LCS06, PLK^{+}06, RS12, SvdBLM11, SN97, VCP08, Zag96]. Discretization [EJR^{+}14, KDBB17]. discriminable [GLS17]. Discriminant [ZPP05]. Discriminative [LZZ^{+}19, MGMP18]. Discussion [SRW^{+}16]. Disease [BGP^{+}11, GNDV^{+}18, MSSD^{+}08, OIR^{+}17, TBB^{+}07]. Disentanglement [NOB16]. Disk [LLLF08, YXSH13]. Disney [Min13]. Disocclusion [DZMQ16]. Disparity [WZC^{+}15]. Displacement [AYRW09]. Displacements [KMDZ10, WCB^{+}12]. Display [BS02, BC18b, CK16, CM16, CDK^{+}17, DK11b, DTT^{+}17, ETO^{+}10, GMS19, Gor02, GPK14, HIH^{+}18, IFP^{+}12, JH13, JRHT14, KKKT18, KBK11, KFL^{+}15, KML96, LSJ^{+}15, LDH09, LDN11, LHC10, MYI13, MBZB12, MIO^{+}15, NHN07, dJOBNM17, ORC07, OHH06, PKS^{+}08, PGRS13, RWBF18, RWF19, RLM10, Sim07, SvLF10, SZH97, SBHW11, TKTN09, TIS16, VB18, WZW^{+}05, WB08, WPSH06, XGS^{+}19, YNYH06, YHW^{+}07, YHLJ08, ZHF^{+}19]. Displaying [JS98]. Displays [ACR^{+}19, AHKMF11, AT16, BSE^{+}17, BI12, BSM06, BJM07, BSEN18, BC18a, BMY05, CAN14, DVCD07, DFG^{+}14, DJK^{+}06, GIMS18, HN13, HRISI15, IK15, IDAK15, IAIK16, IHS17, JH13, JWS04, KPBG13, KBB^{+}12, Kra16, KOK^{+}19, LCR16, LKD19, LH16, LSK^{+}18, MS04, MT05, MMCE09, OTKS15, PIN^{+}15, PBC17, PGI^{+}17, QPNK18, RHZN11, RM17, SLGM09, SBK^{+}11, SFC^{+}07, TKAM06, VH16, WRHR19, XLZ^{+}19a, XLZ^{+}19b, YBW^{+}19, ZFSL19]. Dissemination [BKW16]. Dissertation [Meh17]. Dissimilar [HGWW18, ZCJH12]. Dissimilarity [BDD^{+}16]. Dissipation [KLLR07]. Distance [BA05, BKS01, CZN^{+}11, DBW11, GMS19, Gué01, GDN^{+}07, HLD^{+}08, JBS06, KBB^{+}12, KHSS14, KL14a, KW11, KDBB17, KMH11, LPKK12, LMZ^{+}14, NDS10, SFH06, SK00, YXM^{+}15].

Distance-weighted [HLD^{+}08]. Distances [LKD19, XWL^{+}15]. Distant [WHK15]. Distillation [WGZ^{+}19]. Distorted [SM09]. Distortion [BGR06, CVS^{+}19, GHE19, NWHWD16, WLT08, YYSZ06]. Distortion-based [GHE19]. Distortions [NSZ^{+}17, NA19]. Distractors [PFW09, WP19]. Distributed [BSM06, BSO^{+}12, CMCL06, CLL08, CCQ^{+}14, DL12, LSJ^{+}15, LKS^{+}19, LNS08, MB03, NSS03, OWS15, PST^{+}15, PMD^{+}07, RLM10, SSMG13, SKS12, TIK15, Vis15, Wah14, WAG^{+}12, Wil18, XZB14]. Distributing [LLLF08]. Distribution [DS16b, DCH^{+}17, GQGP17, HDSC19, HLCB18, IWR^{+}18, JH09, KBH13, OHJ^{+}11, SSI99, XZM17]. Distribution-Driven [JH09]. Distributions [AWHS16, BWW^{+}12, LLL^{+}10, NSW^{+}17, WAG06]. Divergence [BK17]. Divergence-Free [BK17]. Diverse [ASvdP14, FPH19, MSSH14]. Diversity [PHJ^{+}10]. Divide [LYY^{+}16b, ZHLB13]. Divide-and-Conquer [ZHLB13]. Divided [SHH11]. Dividing [VB18]. DNA [LBLH19, MDS^{+}18]. Do [BJY^{+}18, FDC^{+}18, GLB16, HGWW18, RKC^{+}16, SGQ16, SCB^{+}19, GMD^{+}17, LKH^{+}16, JAAL18]. Doccurate [SSBC19]. Document [BLE19, BMS17, BISM14, CWDH09, GLK^{+}13, HHWN02, HKBE12, JSR^{+}19, KKP^{+}17, MDL07, PM08, PNML08, SOR^{+}09, cKJG^{+}12]. Documentation [HF11]. Documents [BLE19, GBWI17, KZD^{+}10, KJW^{+}14, SOR^{+}09]. Does [KDX^{+}12, YAE07, ZKM18]. DOF [GL17, TIK15, YEII12, BDF16]. DoG [Mur95]. DOI [JA18]. Domain [BRSP18, CCQ^{+}14, CJTM05, DLW^{+}17, DNP07, DL03, GHL18, aKS12, KCS^{+}16, LS02, Mar18, RBGH14, YLSL11, Zhu05]. Domain-Specific [CCQ^{+}14, KCS^{+}16, RBGH14]. Domains [DPR00, KBH13, KZW12, NJ99, PTC10, WPZ^{+}11]. Domino [GGL^{+}14a]. Don’t [BS11]. Doors [SFB^{+}12]. Doppler [PH07]. DOPs [KHM^{+}98]. Dot [IFM14, NWI17, RW18, TSD09, ZTA12]. Double [WPC^{+}13, YSL^{+}13, YHJ^{+}17]. Double-Sided [YSL^{+}13, YHJ^{+}17]. Doug [Ano14e]. Down [DH08, LFR03, LS10]. DQNViz [WGSY19]. Dr. [Han95]. Draco [MWN^{+}19]. Drag [OKB^{+}19]. Dragging [CLEK13]. DrawFromDrawings [MSA17]. Drawing [AMA06, CHK04, DKM06a, FWSL12, FT08, HS11, HSCS11, HQQ12, KZL07, LLHL14, MSA17, MNS18, NW11, PPP12, RSFH14, SZHR11, ZPG19, ZCFL15]. Drawing-Based [HQQ12]. Drawings [BRH^{+}17, BKW03, HSL19, JFZ^{+}18, KKW^{+}17, MNZ^{+}15]. Drawn [CLAL12, TRL^{+}19]. DRF [ZCFL15]. DRF-Based [ZCFL15]. Drills [LLQ^{+}17]. Drive [YHMY08]. Driven [AZM12, ACR^{+}19, AG16b, ARL^{+}17, BGOJ16, BMS17, BM10, CDBR14, CDR^{+}18, CXM19, CLRP13, CM11, DS16b, FXG12, GH00, GABJ07, GABJ08, GGA^{+}11, HBJP12, HDSC19, HD12, HZL^{+}19, JH09, KW14, KSI^{+}96, KSL^{+}17, KST^{+}14, KGZ^{+}12, LJX^{+}10, LCC^{+}17, LHJ^{+}18, LMD12, LSS13, LFP07, LWC^{+}19, LLL^{+}19b, MD12, Mar18, PZLZ17, PK13, PLS^{+}14, PGL^{+}12, RLNN11, SSH14, SPG14, SSL^{+}12, VKG05, VFSG06, WAM^{+}19, WYM08, WHK15, WFC^{+}18, WMGE12, YLL^{+}12, ZLG^{+}06, ZH07, ZLD^{+}14, ZNZX16, ZGM18, tCMR08, MKT^{+}18, YBW^{+}19, ZZBW08]. Driver [KGAM18]. Drivers [LWLM18]. Driving [BIPS12, MKT^{+}18, TYSN06]. Drone [EIKS18]. Drone-Augmented [EIKS18]. Drones [EIKS18, SEB19]. Drop [ZWW^{+}12]. Drug [PLS^{+}14]. Drumming [KBS13]. DSLR [WLHD17]. DSPCP [NR18]. DT [AS19, JFTW07, MN07]. DT-MRI [AS19, JFTW07, MN07]. DTI [BVPtHR09, CDZ^{+}09, JDL09, PFK07, PPvA^{+}11, ZCL08]. DTI/HARDI [PPvA^{+}11]. Dual [ATLF06, DWF^{+}19, HKG07, HF06, KLM04, KM96, LLLF08, LML^{+}18, ME11b, SJW07, TFH11, WLLC15, YLSL11, YFZ^{+}18]. Dual-Domain [YLSL11]. Dual-Matrix [WLLC15]. Dual-Mesh [ME11b]. Dual-Representation [HF06]. Dual-Space [KLM04]. Ductile [FHG^{+}09]. Duet [LBW19]. Duplication [HBF08]. Duration [MVN^{+}19]. During [DC17, JBS^{+}18, KBGE11]. DVV [KOJC12]. Dwell [XLZ^{+}19a, XLZ^{+}19b]. Dwell-Free [XLZ^{+}19a, XLZ^{+}19b]. DXR [SLC^{+}19]. Dye [JEH02]. Dynallax [PKS^{+}08]. Dynamic [AHSS14, AL11, APP11, BFP14, BSS^{+}13, BDY06, BMWW18, BFL06, BVB^{+}11, CWL12, CS08, DVP^{+}18, DL12, FvdPT97, FBL^{+}18, FM06, FT08, FTES13, GJG^{+}15, GPK14, GGPPS13, Gué01, GN12, HL02, HSS11, HSCW13, HBW14, HWL^{+}11, HQ04, HTZ^{+}11, HEF^{+}14, JS06, JFY16, JWL05, KJL^{+}12, KLS^{+}18, KW13, LRP97, LPF^{+}19, LBH18, MB18a, MQV00, MKH12, MZH^{+}08, MKW07, MTB17, NWI17, NSE^{+}12, NKHC08, OBKP18, OBLN17, PCY08, PKS^{+}08, PH08, QMV98, RD05, RGFLL14, RP12, RG95, RM13, RSR^{+}18, SST^{+}17, SAM^{+}05, SWW^{+}15, SAC^{+}08, SG99, SLW^{+}10, Sun03, TJW^{+}17, TF06, TSH^{+}14, VS11, VPF15, VBW16, WLB^{+}14, WMWL11, WLMP19, WFG^{+}19, WB05, YNCP06, ZK17, ZPS04, ZDM13, ZGH^{+}18, ZX18, ZDH^{+}19, vdEHBvW14, vdEHBV16, FHL10, LCM07]. Dynamically [BS16, LYY^{+}16a, SA19]. Dynamics [BRNB19, BLG^{+}16, Fau99, GRVE07, GBM^{+}12, KBE^{+}18, KBE09, LVRH07, OBLN17, SJJ^{+}17, SWL^{+}14a, SGPR18, YESK95, ZBMY14, vLBR^{+}16].

E-SD [HRN^{+}03]. E-Transactions [XCH^{+}14]. Early [KYK11, Zag96]. Earth [SPO^{+}12]. Earthquake [LLQ^{+}17]. Easier [OSSK12]. Echoes [AGN^{+}19]. Ecosystem [McK09]. Ecosystems [BCH^{+}13]. Eddy [WPS^{+}16]. Edge [BRH^{+}17, BYA15, BJB^{+}12, BSL^{+}14, BVB^{+}11, CZQ^{+}08, DSC^{+}08, DSS^{+}09, DRMM13, EHP^{+}11, Hol06, HM10, KEJK19, LLCM12, OKSK16, SHH11, SMNR16, SZW^{+}19, TCM06, YSZ04]. Edge-Aware [BYA15]. Edge-Based [KEJK19]. Edge-Bundling [LLCM12]. Edgebreaker [Ros99]. Edges [AFRS05, GHL15, OKSK16, SPEB18, Wan06, XRP^{+}12]. Edit [BDFM17, XNT11]. Editable [GXH^{+}13]. Editing [ATLF06, BKW08, CML^{+}07, HLCB18, LZH^{+}07, LHFY12, LZW^{+}13, LSS^{+}15, TDLG19, WM13a, WP16a, WCB^{+}18, WQ07, XLND11, ZZL^{+}15]. Editor [De 18b, Lin11a, Lin14c, Lin14d, MY14, MFS^{+}09, AD12, Ano05a, Ano05b, Ano13n, Ano14n, Ano15l, Ano16n, Ano16o, BvdP12, BDC17, BGK11, CW11, De 15b, De 15a, De 15c, De 15d, De 16b, DS16a, De 17b, DS17b, De 18c, De 18a, DS18b, De 19b, DFQ12, Ebe00, Ebe03a, Ebe03b, Ebe04a, Ebe04b, Ebe06a, Ebe06b, Ert07a, Ert07b, Ert07c, Ert08, Ert09a, Ert09b, Ert10a, Ert10b, Ert10c, Ert10d, Ert11b, Flo16, De 16a, De 17a, Flo17, GKR14, GJK15, HK10, Hag99, Hag00, HE02, Hag03, HVY16, IKT15, Joy02, JLS15, Kau98, KL14b, LSCN09, Lin11b, Lin11d, Lin11c, Lin12b, Lin12c, Lin12d, Lin12e, Lin12a, Lin13a, Lin13b, Lin13c, Lin13d, Lin13e, Lin14c, Lin14a, Lin14b, Lin16a, MH10, Moo03, MB19b, Mue19, MYM08, NSvW11, OO15, OA11, PZ12, Pur09, Qin09, Rus99, SK16a, SK15, SSL08, SL11]. Editor [SW17, Var01, WY19b]. Editor-in [Lin14d]. Editor-in-Chief [Ano13n, Ano14n, Ano15l, Ano16n, Ano16o, De 15d, DS16a, DS17b, De 18a, DS18b, De 19b, Ert07a, Ert10d, De 16a, Flo17, Hag03, Kau98, Lin11d, Lin11c, Lin12e, Lin12a, Lin13e, MB19b, Mue19]. Editorial [Ano97b, Ebe07, Ert07a, Ert09a, Ert11a, Kau95, Kau96a, Kau96b, Kau96c, Kau97, Kau98, KRTvW06, SGR06, SW06, BLRW05].

Editorials [KN95]. Editors [CKSB14, vW11, Ano12g, Ano13o, Ano14o, Ano15m, Ano16p, Ano16q, BHTY15, BSS18, BRS18, BKL18, CCH14, CLS07, CGGR18, CFK12, CLS13b, DW17, DLM^{+}12, FBI07, GGHZ19, GW13, GMM05, HKQ13, HP04, HLM10, ILMH12, KMN04, KS14a, KKL11, KHSB11, KPGL12, LS06, LST^{+}16, LBKD09, LABS10, MSW19, MW13, RvWT05, TL11, VW12, WLW17, WM05, vHMM^{+}11, vWMT04]. Education [CB15, Chi16, HA17, SST^{+}17]. EdWordle [WCB^{+}18]. EEG [tCMR07, tCMR08]. Effect [AS11, APP11, BCS11, CDF14, DBD17, DBBF19, GUFM15, GLB16, HKKS18, JSB13, MS18a, MWCR06, MDF12, MTW^{+}12, PDBG18, SZW^{+}19, UBH19, ZKM18]. Effective [BSB^{+}18, CGH^{+}19, CVG13, CG08, DSSK08, EMRY02, HHO^{+}17, HHB16, JQD^{+}08, LFH06, PKG12, PRA^{+}10, SOK^{+}16a, SOK^{+}16b, SRML07, TRL^{+}19, WBJ16, XHT^{+}07, YCZ^{+}16, YEII16, CJR07]. Effectively [RKA^{+}13]. Effectiveness [BS16, HW12, NAK18, RBK^{+}15, RFF^{+}08, SED19, SBW17, TTR10, BPC^{+}10]. Effects [AW14, BDM^{+}17, BM10, BC18b, CK16, CCAL12, FKS16, GMS19, GO15, HD11, JWD^{+}14, KBB^{+}12, KHSI04, LSSB12, LBS13, LBS14, LLPP19, LH11, LRM^{+}13, LH16, LHH16, LHC16, LBHW18, LT18, LH14, OPH^{+}16, RKSB13, RBK^{+}15, RSBB17, RGFLL14, RMW09, SMG^{+}13, SDHH12, SFR^{+}10, VZS18, VBC^{+}16, WBK^{+}08, WX13, WBA^{+}14, vdCvW14, SBS16]. Efficiency [Vas16]. Efficient [APW16, BES12, BSEN18, BN12, CMLZ08, CCB11, CCS12, EG09, FAW10, FHG^{+}09, GGTH07, GJ10, GH00, GSG96, GJR^{+}14, GNPH07, HAAB^{+}18, HAGS16, JR07, KSDA16, KYK11, KLCK17, KKP^{+}17, KHM^{+}98, KS01, KKM^{+}09, KGPB05, KJL^{+}12, KOF08, KZW12, KKCS98, LK11, LVRH07, LS13b, LLLF08, LI06, MBT07, MNKW07, MDL^{+}19, MTS07, MSHC99, MTRP10, OBS^{+}15, PD04, PQF^{+}09, PRA^{+}10, RNL09, RN19, RKWH12, SP07, SNM16, SRK^{+}11, SvW98, SR17, SSB^{+}17, SF19, THM15, TDR10, TNS10, USM96, VPF15, WWYS04, WCZ^{+}11, WM13b, WX13, WSYM17, WWZ^{+}19, WD10, WSC^{+}95, WP18, XTY^{+}11, XNT11, YEII12, YEII16, ZT99, ZCW19, ZD18, vRKEE17, DMAM19]. Efficiently [AMA11, NSS14, XHF12]. Effort [CFEC17]. Egocentric [BDF16, JNC^{+}15, MDB18, RFSP19, SWW^{+}15, SJK^{+}07, WPZ^{+}16]. egoSlider [WPZ^{+}16]. EGPGV [AD12, CW11, PZ12, MFS^{+}09]. EIC [De 19a, Ebe07, Ert11a, Lin15, Ano15i]. Eigenfunction [NMGK17]. Eigenvalue [PYW^{+}16]. EL-REP [RF11]. Elastic [BLE19, CDA99, HDBC15, LXW^{+}18]. Elasticity [YL16]. Elastodynamic [CAP18]. Elastoplastic [HWW18]. Electric [CVC10, WSM^{+}09]. Electronic [Kin10, KCK^{+}19b, MGJ^{+}10]. Electrophysiological [LHH^{+}12]. Electrostatic [GBM^{+}12, YNYH06]. Element [AWB11, BTB^{+}04, BSL^{+}12, FSHH12, HTE11, MNKW07, NKH11, NLKH12, NKH14, WBH04, YEB16]. Element-Based [HTE11]. Elemental [JDSR^{+}18]. Elements [AERA14, DWA10, KA12, NK06, SBM^{+}06, ZQS11]. Elevation [ZSTR07]. Elicit [BBC15]. Elicitation [HHH16]. Eliciting [CdOKRV09]. Elimination [SFC^{+}07]. Ellipsoidal [SLG^{+}17]. Ellipsoids [CCW^{+}09]. ElVis [NLKH12]. Embeddable [FDFR10]. Embedded [SSEW19, TNB11, WHA07, WJD17, YDGM17]. Embedder [DBD13]. Embedders [KW05]. Embedding [BBiA09, CVS^{+}19, JSR^{+}19, KHZR18, LPCC17, PKL^{+}18, RT12, SLQW17]. Embeddings [BMS17, LBT^{+}18]. Embellishments [BARM^{+}12]. Embodied [GPK14, LKS^{+}19]. Embodiment [BPS13, BDH^{+}18, SFL^{+}16]. EMDialog [HSC08]. Emissive [MMCE09]. Emotion [SPW07, VBC^{+}16, WBA^{+}14]. Emotion-Based [SPW07]. Emotional [WGR^{+}18]. Empathy [Ger17]. Empirical [BDJ14, BARM^{+}12, DBB10, FFB18, JHKH13, KWH19, LBI^{+}12, LKD19, MRO^{+}12, MGMP18, PGK16, SMT13, TGH12]. Empirically [Rot13]. Empirically-Derived [Rot13]. Employing [GS08, WCC^{+}19]. Empty [HAAB^{+}18]. Emulation [OIR^{+}17]. Enabling [ED06, JKRY12]. Encapsulation [JKM01]. Encode [SMO^{+}13]. encoder [SXX^{+}19]. Encoding [ACS^{+}18, BDD^{+}16, FDPH17, GSCI15, KEJK19, KJL^{+}12, MPK^{+}13, WKB^{+}13, ZBG^{+}17]. Encodings [CG14, GBFM16, SSRE18]. Encouraging [FDPH17]. Endoscopic [NWF^{+}05]. Endoscopy [HHCL01, KKPS08, WGC^{+}08]. Energy [BNTM16, CHK04, GDJ^{+}13, HKG07, RB07, RB11, SSF13, WHK15]. Energy-Driven [WHK15]. Enforcement [MMT^{+}14]. Engaging [LKS13a]. Engine [CDL^{+}16, DKMI13, DCH^{+}17]. Engineering [GS08, JKJTM06, MGS^{+}14, SZK15, Zag96]. Engines [DDGL07, SWL^{+}14a]. Enhance [FWL17, STH13, SJL^{+}18b, ZK17]. Enhanced [BLE19, CZZ^{+}19, ETO^{+}10, LPP^{+}06, LG15, NDS10, REB^{+}16, SASS16, SEA09, TC13, YESK95, YFC^{+}19]. Enhancement [JPLŠ16, KV06, MDM10, VKG05, VJN^{+}15, WK13, ZGW^{+}14, ZNZX16, ZWM13]. Enhancements [KWL14]. Enhancing [BG07, CCSK19, GYK^{+}16, KW14, KOCC14, KSTE06, KSS09, LJZ12, MJL^{+}13, TCM06]. Enriching [LXB17]. Enrichment [KTE15, NA19]. Ensemble [BGOJ16, CZC^{+}15, DDW14, HZ13, KTB^{+}18, KRRW19, LBS^{+}19, SZD^{+}10, WLSL17, WHLS19, WRF^{+}11, XXM^{+}19b]. Ensemble-based [XXM^{+}19b]. EnsembleLens [XXM^{+}19b]. Ensembles [BLLS17, CLEK13, FFST19, FBW16, FKRW17, GBP^{+}13, HHB16, HSSK16, HMZ^{+}14, HOGJ13, KBL19, KSDD14, LBR^{+}17, LPCRH19, ME19, MGS^{+}14, MWK14, OBJ16, WMK13]. Ensuring [HHCL01]. Enterprise [KPHH12, PGU^{+}13]. Entertainment [NQX^{+}05]. Entity [BCB10, FWL17, KGS^{+}08]. Entity-Based [BCB10]. Entity-Component [FWL17]. Entourage [LPK^{+}13]. Entropy [RMCW19]. Entry [XLZ^{+}19a, XLZ^{+}19b, YFZ^{+}18]. Environment [AB01, CPW^{+}18, CDBR14, EGH^{+}06, FYWY16, HWHK16, HWL^{+}11, JOR^{+}19, JPD^{+}18, KCH11, KLL12, LH11, LGS^{+}11, LIM^{+}12, LODI16, MZS^{+}19, RCW^{+}18, RJG17, RSD^{+}13, VSS08, WMWL11, WKCB07, YAE07, ZLB^{+}05, ZWBH13]. Environmental [MBH^{+}12, NTS11, SKU^{+}12]. Environments [ANR^{+}18, AGN^{+}19, AM13, BZS^{+}13, BK12, BL15, BIPS12, BRNB19, CWL12, CLL08, CGJM19, CPK^{+}05, CRI06, CVC^{+}12, DL12, DK11b, DDBR^{+}19, DDKA06, ESV98, FWK16, FS14, GPL^{+}13, HSK14, HLRC^{+}12, IDW^{+}13, JAAL18, JSB13, JHBK19, JWL05, KHSS14, KLL^{+}13, KCT^{+}17, KAK^{+}18, KCK^{+}19a, LWG05, LKC09b, LFH06, MWCR06, MRG^{+}15, NBM19, NDR96, NSS03, OBKP18, PFW09, PWHK16, PBK^{+}12, PGRS13, QWC^{+}09, RZP^{+}07, RSR^{+}18, SM12, SGQ16, SL08, SMP17, SAC^{+}08, SFR^{+}10, SLF^{+}12, Tay02, TWMZ19, UK12, Wah14, WZQK04, WKW06, YDJ^{+}19, ZDM13, ZKM18]. Episodes [KBK11]. Equal [BHB04]. Equalization [SRML09]. Equalizer [EMP09]. Equation [AG17, DRW16]. Equations [WDC08]. Equipment [SFL^{+}16]. Erez [Ano13u]. Ergonomics [ZLDM16]. Ergonomics-Inspired [ZLDM16]. Errata [GXW^{+}18a, NN11a, SOK^{+}16a, TdJ15, WC10, XLZ^{+}19a]. Erratum [Ano14h, Gov18, GTS11]. Error [CCK07, CG14, EJR^{+}14, FCZ15, FSME14, HYB^{+}17, HQC^{+}19, MY96, MGM19, PA06, RMW09, Sbe97, THJ99]. Error-Bounded [HYB^{+}17]. Errors [BJB^{+}12, MMS^{+}98, SDMT16]. Escape [MSME14]. Escher [LML^{+}18]. eSeeTrack [TTS10]. Establishing [NSN14]. Estimates [GJC^{+}17]. Estimating [RRD^{+}13, SS08, Zag96]. Estimation [AT05, AMC10, CHM11, GWBO12, HBESB11, HAM11, HKKS18, IAIK16, IFM14, JNC^{+}15, JFBB10, KBB^{+}12, KLM04, KS00b, LKL^{+}15, LDW^{+}15, MUS16, MDS16, MOG11, MSM^{+}11, RSS14, SBS16, SBJ^{+}10, THV^{+}14, WP19, ZCL^{+}19, Zho16, KS00a]. Estimators [BLM96, KWH19]. Eternal [Bai13]. Euclidean [DPR00, KW05]. Euler [CAP18, RD10, RZP12, RSFH14, SAS16, SZHR11, SRHZ11, Wil12]. Eulerian [CM14, CMK15, JEH02, SWTH07, SXM17]. Eurographics [BvdP12, BDC17, KS14a, KL14b, MFS^{+}09, SK15, AD12, CW11, Ota17, PZ12]. EuroVis [MYM08]. EVA [LGM^{+}18]. Evaluate [DBP14, TTR10]. Evaluating [BW11, BBG^{+}18, BJK^{+}16, BPC^{+}10, BBIF12, DBD18, GO15, GQM^{+}18, HTP19, HSR18, IIC^{+}13, KO12, KLD^{+}09, KPR^{+}14, LBS13, MHD^{+}18, MBZB12, MTW^{+}12, NAK18, OJEF19, PQMCR17, PBC17, Ros11, SSRE18, SND05, SBW17, Vas16, WZ08, ZLC^{+}19, cKJG^{+}12, SS19]. Evaluation [AJDL08, AZM12, ALBR16, AS05, BWS^{+}19, BGP^{+}11, BYB^{+}13, BPS13, BTB10, BLIC19, BE09, Bru17, BGR06, BKH^{+}11, BTJ^{+}13, CCK07, COMP13, Csé08, Cse10, DRW16, DRHK07, EMdSP^{+}15, FSTG16, FTES13, GLH^{+}14, GHL15, HKBR^{+}14, HMSA08, HKH^{+}12, HQC^{+}19, JRHT14, JF16, KGS^{+}08, aKGS11, KKKT18, KOCC14, KCH11, KLL12, KDA^{+}09, LLPP19, LT99, LD11a, LDA12, LFA^{+}16, MLMF12, MYI13, MK13b, MMMY97, MIO^{+}15, NDM^{+}97, NHB^{+}17, NTS11, OBKP18, PFW09, PFW12, PWHK16, PFG08, PB16, RMW09, SSKB14, SD12, SFR^{+}10, WAM^{+}19, WR11, WKSS05, WCA^{+}17, XC19, YDGM17, ZHF^{+}19, tCMR07]. Evenly [LMG06, WLZM10]. Evenly-Spaced [LMG06]. Event [CvW18, CXR18, DFD^{+}14, DSP^{+}17, GS14, GXZ^{+}18, GJG^{+}19, KBK11, KFN06, LLMB19, LSB^{+}16, MLL^{+}13, PLC^{+}11b, VJC09, WG12, YLL^{+}12, ZCD19]. Event-Based [KFN06, VJC09]. Event-Guided [DFD^{+}14]. EventRiver [LYK^{+}12]. Events [BSR^{+}14, CDW^{+}16, DGWC10, LGM^{+}18, LWLM18, WCS^{+}18]. EventThread [GXZ^{+}18]. Eversion [FS04]. Everyday [PSM07]. Evidence [BE09]. Evolution [wAPS14, BSH^{+}16, BGM^{+}17, DMR04, KRRW19, KZW^{+}16, LWW^{+}13, LWL14, SBV^{+}11, Sun03, WPZ^{+}16]. Evolutionary [GdBG12]. Evolve [CLWW14]. Evolving [CLT^{+}11, KW19, LSB^{+}16, WLWL10]. EvoRiver [SWL^{+}14b]. EWA [ZPvBG02]. Exact [KL14a, KTCG17, Wil12, XLND11, XHF12, XWL^{+}15]. Examination [HHCL01, KLL12]. Examining [FWK16, aKS12, YQK^{+}17]. Example [ADP02, BGCS17, CZZ17b, DV95, FXG12, LC10, TTR10, WWYS04, ZBO13]. Example-Based [CZZ17b, FXG12, LC10, WWYS04]. Examples [HYZ^{+}12, SZ11, XYS^{+}16]. Exceptions [QH18]. Exchange [LBLE^{+}14, YSZ^{+}19]. Excluded [LBH14]. EXCOL [ZHF12]. Execution [EASD^{+}19, IBJ^{+}14]. Executive [Ano14v, Ano16x]. Exemplar [CWDH09, LKL^{+}15]. Exemplar-based [CWDH09]. Exemplars [DHM13b]. Exhaustive [JNK19]. Exhibit [RGFLL14, BHP^{+}12]. Existing [CPC09, MPG^{+}14]. Exocentric [EIKS18]. Expand [vHP09]. Expanding [LIRC12, Mal05, WWC^{+}14]. Expansion [Hau97, KLM04, MMMY97]. Expectations [KRH18]. Experience [LKS^{+}19, TUG17]. Experiences [HHH16, LLKN17, LLL^{+}19a, LSV^{+}18, PKG12, RKA^{+}13, WBA^{+}14]. Experiment [SNB^{+}17, SFL^{+}16]. Experimental [BMGK08, BHZ^{+}18, BE09, BPS^{+}11, FIBK17, KHSI04]. Experimentation [KTC^{+}19]. Experiments [MRSS^{+}12, PGK16, TSA14, vFWTS08]. Expert [Gle13]. Experts [AJDL08, GHL18, aKS12, RH19, TvET14]. explain [FGS19]. Explainers [Gle13]. Explaining [GRS^{+}19, PSM12]. Explanation [BW00]. Explanatory [RRJH18, WBE^{+}06]. Explicit [FSHH12, LCS^{+}12, YYD^{+}19, ZCCB13]. Exploded [BG06, KLMA10]. Exploiting [DMAM19, GH95, VB13, WWL07]. Exploration [AZL^{+}19, AMA08, BSB^{+}18, BTC13, BBH^{+}17, BMS17, BIAI17, BWW^{+}17, BMWW18, BBP08, BBG^{+}09, BVV^{+}19, BWT^{+}11, BGKG06, BM10, BLG^{+}16, CDDS18, CQC^{+}08, CWT^{+}08, CDZ^{+}09, CCM^{+}14, CZC^{+}15, CGH^{+}19, CSC07, CM08, DC17, DvVH^{+}19, DNN13, DFD^{+}14, DCCW08, DMS^{+}08, EDF08, ERHRF10, EDF11, EBB^{+}15, FDPH17, FPH19, FPV^{+}13, FMST96, GCL^{+}15, GNBP11, GBPW10, GLK^{+}13, GS06, GW11, HSS11, HLRS^{+}08, HBJP12, HLRC^{+}12, IVJ12, JNK19, JBMS09, JKM01, JKMG07, JSR^{+}19, JFY16, KAKC18, KLM^{+}08, KKP^{+}17, KBH06, KG06, LHD18, LLMB19, LPP^{+}06, LS09, LMZ^{+}14, LBK^{+}18, LFH06, LCZ^{+}19, LGG^{+}18, LKS13b, LS16, LWD^{+}17, LBT^{+}18, LODI16, DSC^{+}16, LBS^{+}19, MJW^{+}13, MV06, MI13, MLKS18, MTRP10, NM13, PBN^{+}13, PSSC17, PSTW^{+}17, PLS^{+}14, PM08, PS06, PHE^{+}18, PTMB09, PFK^{+}08, POM^{+}09, RGK^{+}13, RT12, SHM10, SKBE17, SM17, SZY^{+}18, SPN^{+}16, SKU^{+}12]. Exploration [SYS^{+}06, SS18, SPG14, TJW^{+}17, TAE^{+}11, TEC^{+}16, TWSM^{+}11, UDSL18, VMCJ10, VJC09, WLJ^{+}12, WCZ^{+}11, WYL^{+}14, WFW^{+}17, WWZ^{+}19, WWYP19, WAG06, WMGE12, WDSC07, WS09, WXZ^{+}16, XXM^{+}19b, YEB16, YEB18, YHW^{+}07, YSI^{+}10, YDC^{+}14, YRWG13, ZMH^{+}09, ZK06, ZCCB13, ZSCC18, ZLDM16, dOL03, vHP09, vPBB^{+}10, vdEvW14, vdEHBV16, SKL^{+}14]. Explorations [Gle13]. Exploratory [BE18, BWK^{+}13, DPW^{+}15, DB07, FPB17, GBWI17, LH14, LLZ^{+}16, LTP^{+}05, MLMF12, SKB^{+}18, STM17, TWSS16, VMN^{+}19, VN19, WMA^{+}16, XYC^{+}18, ZGC^{+}17, ZCPB11]. Explore [CSWP18, HF06, HFM16, LPK^{+}16]. Explorer [PGK16, BGM^{+}17, GFG^{+}14, NJBJ09]. Exploring [AAH^{+}13, BDF^{+}10, BDSW13, CvW18, CHW^{+}18, CG14, DYW^{+}13, EIKS18, FSW09, FH07, GWP^{+}18, GWFI14, GK07, HF11, JDL09, JDL12, KBL19, KC14, LPCC17, LSB^{+}16, LYK^{+}12, MB19a, MAAB^{+}18, PGSF16, PBC17, RGFLL14, RB11, Sel15, SAM^{+}05, SDW11, SSGM19, STS^{+}14, TIW^{+}19, WM18, Wen14, WG12, WCS^{+}18, YLZ^{+}13, ZWJZ12, vGMSW15, GMD^{+}17]. Exponent [BGT12]. Exponential [CAP18]. Exponents [GT17, GHP^{+}16]. Exposing [WGS^{+}13]. Exposition [WKD19]. Exposure [SBE^{+}15, YWV^{+}19]. Expression [CWZ^{+}14, DNL^{+}06, EG09, KLK^{+}09, ZPS04, ZLG^{+}06]. Expressions [BZS^{+}13, Ros11, WHM14]. Expressive [BLB^{+}17, DNL^{+}06, KSL^{+}17, KKSE17, RHY14, YEB18]. Extended [CMFL16, GPL^{+}13, GQGP17, IMS15, KWP01, PLW11, PGI^{+}17, RWBF18, SSS13, VGKS12, vdEHBvW14]. Extending [FDFR10]. Extensible [MWN^{+}19, WSD^{+}13]. Extension [DCK^{+}12, TLH10, TUG17]. Extensions [BBP08, EM06]. External [CMRS03, ČB19].

Externalizing [MGM19]. Extinction [AZD17, SMP11]. Extinction-Based [SMP11]. Extinction-Optimized [AZD17]. EXtract [ZHF12]. EXtract-and-COmplete [ZHF12]. extracted [BWW^{+}17]. Extracting [AAH^{+}13, GGL^{+}14a, GPP^{+}16, JCWD14, PMH18, YFM01, YSZ^{+}19]. Extraction [Ano96b, AE13, AJ19, BDSS18, CML^{+}07, CMM^{+}97, DS16b, ESN^{+}09, ENS^{+}12, GSDJ04, GH95, GKL^{+}16, HKG07, KHL99, KGP^{+}13, LTPH17, LLC15, LBH11, LLRR08, LXR19, LSJ96, MS08, NN11a, NN11b, NKH11, OMD^{+}12, RL08, RKZZ19, SP07, SSS06, STS10, SJ06, SWF^{+}16, SH00b, VAB12, WL08, WC09, WC10, WM13b, WPB^{+}11, vWPSP96]. Extrafoveal [TUG17]. Extrema [IK95]. Extreme [BMA^{+}19, GHS^{+}19]. Extreme-Scale [BMA^{+}19, GHS^{+}19]. Extremum [TN13, XHT^{+}07]. Extruded [SM09, SM11]. Eye [ACR^{+}19, AJ17, AAM^{+}12, AABW12, BPS13, BKH^{+}11, DTT^{+}17, FGBB09, HZL^{+}19, JA18, KDX^{+}12, KW13, KHH^{+}16, KHSW17, LMD12, MSM^{+}11, MVN^{+}19, NBW14, OTKS15, OIR^{+}17, RWBF18, WXY17, XGS^{+}19]. Eye-Controlled [OTKS15]. Eye-Head [HZL^{+}19]. Eye-Mark [MSM^{+}11]. Eye-Tracked [OIR^{+}17]. Eye-Tracking [AJ17, JA18, KW13, MVN^{+}19]. Eyeglass [XGS^{+}19]. Eyeglasses [CDAF18]. Eyes [KRH18].

Fabricating [AIS18]. Fabrication [PDRK19]. Fabrics [ZBZ^{+}13]. Facade [AYRW09]. Facades [FWZQ13]. Face [CPW^{+}18, KEP08, LJZ12, OJEF19, PZ07, PQF^{+}09, SLGM19, SLS^{+}17, WRHR19]. Face-Centered [KEP08, PQF^{+}09]. Face/On [WRHR19]. Faced [ZCFL15]. FaceForge [SLS^{+}17]. Faces [FBLS05]. FacetAtlas [CSL^{+}10]. Faceted [DRRD12, WMA^{+}16, ZCCB13]. FacetMap [SCM^{+}06]. FaceWarehouse [CWZ^{+}14]. Facial [ADDG12, BZS^{+}13, CWZ^{+}14, DNL^{+}06, LZD13, WHM14, ZPS04, ZLG^{+}06]. Facilitate [BHP^{+}12, SDES19, SPB08]. Facilitating [LDFZ14, SS18, ZCCB12]. Factor [TLLH12]. Factories [XMRC17]. Factorization [CLRP13, EGG^{+}12, Lac96, ZPP05]. Factors [DBP14, TM04, WG12, ZOC^{+}13]. Facts [SDES19]. Faithful [Hu16a]. Familiar [LPCC17]. Familiarity [DLW^{+}17]. Families [KMG^{+}06, MGKH09]. Family [FDC^{+}18]. Farewell [De 19a, Ebe07, Ert11a, Hag03, Lin15]. Farms [SOL^{+}13]. Fast [BYA15, CF10, CK10, Cse13, DMR04, DL03, Dru08, ERL^{+}13, Fau99, GJ10, GSM^{+}14, GLM06, GJR^{+}14, Har16, HKH^{+}12, HK99, IYK01, IYIK04, IMS15, JvdLR13, KH19, KSDA16, KNP04, KGS98, KL07, KTCG17, LH03, LYS^{+}10, LRZM11, LGY19, LI06, LSC08, MRT00, NB95, OBLN17, PB13, PNML08, RLH11, RKZZ19, SM04, SE18, SPP^{+}14, SK00, SZH97, Ste98, SRML07, TCM^{+}12, VML97, Vis15, Wal12, WLMP19, WLD^{+}19, Wu16, WHL16, XLND11, YYY16, Zag96, XWL^{+}15]. Fast-Forward [HKH^{+}12]. Fast-Turnaround [NB95]. Faster [AMA11, BHB04, GUO00, HH10, HL09, WFM^{+}05]. Fault [WCB^{+}12]. Fauxvea [GJC^{+}17]. Favre [MFS^{+}09]. Feasibility [SB14]. Feature [ALMF19, AFRS05, BKRE19, BKL^{+}11, CCM^{+}13b, CRT04, CMN13, CSC06, CJTM05, DANS10, FYP10, GKK^{+}12, HLRS^{+}08, HYB^{+}17, JSR^{+}19, KHL99, KPB14, LZH^{+}07, LWCS96, LH09, LWZ^{+}16, LDC16, LWC^{+}18, ME18, MWCE09, MK13a, MDM10, MOG11, PYW^{+}16, PSR17, RKWH12, RGC^{+}14, SYM14, SM17, SS06b, SRML07, VKG05, Wan06, WWLM11, WSW16, WTB^{+}19, WTVP11, WMK13, WMGE12, ZAM11, vWPSP96, CJR07, LSC08]. Feature-Based [BKRE19, BKL^{+}11, ME18, MK13a, RGC^{+}14]. Feature-Driven [WMGE12]. Feature-Insensitive [AFRS05, Wan06]. Feature-Preserving [FYP10, LDC16, LWC^{+}18, SRML07, WWLM11]. FeatureLego [JNK19]. Features [AMA07, CC08, DW14, DS16b, EHBA11, GK95, GDKB17, KY06, LT11, LLLF08, MPWG12, QM16, SXX^{+}19, SW97, WLL^{+}05, WSPVJ11, WFS^{+}16, YHR^{+}19, YON06, LSC08]. Feed [PSM06, WXKP14]. Feed-Forward [WXKP14]. Feedback [AGN^{+}19, BKA^{+}11, EASB19, HOG^{+}12, KFN06, LBHW18, RB11, WRHR19, WWL^{+}10]. Feel [JAAL18]. Feels [GLM^{+}17]. Feiner [Ano14d]. Felix [SPN^{+}16]. FEM [JDSR^{+}18]. Ferns [GSCI15]. Fetus [MMK^{+}17]. FI3D [YSI^{+}10]. Fiber [BWW^{+}17, BVPtHR09, EBB^{+}15, JDL09, KTCG17, RBN^{+}19, TC17, WKI^{+}17, WY19a, ZCL08]. Fiber-Clustering [ZCL08]. Fiber-Level [WY19a]. FiberClay [HRD^{+}19]. Fibers [CDZ^{+}09, SSC^{+}16, SS08]. Fibrous [MMYK06]. Fidelity [BBG^{+}18, BZGV14, CS18, CPK^{+}05, HWA15, LBS14, MWCR06, MBZB12, MZS^{+}19, RYL^{+}18, RPHI08, SHC^{+}09, VBC^{+}16, YRP19]. Fiducials [HR11]. Field [Ano05c, Ano14h, BES12, BW08a, BSS^{+}19, BBG^{+}18, BJA^{+}19, BSEN18, CML^{+}07, DANS10, DGBW09, DPR00, DTT^{+}17, FBW16, FCL09, FSE12, GPR^{+}01, IK15, IAIK16, IMS15, JDSR^{+}18, JHW^{+}14, KDBB17, KOK^{+}19, LJX^{+}10, LKJ^{+}05, LKC09a, LH16, LHH16, LBZ^{+}11, LL04, LPG15, MK13b, PW95, PLC^{+}11a, PBL10, PGL^{+}12, QPNK18, RKSB13, RBK^{+}15, RBN^{+}19, RWBF18, RLH11, RLL^{+}13, RBDG15, RLK19, RT12, RKZZ19, SKMR98, SZH97, TN11, USE13, WLL^{+}05, WSY07, WLHD17, WASQ18, WZF^{+}04, WT10a, YHLJ08, YCLJ12, YLK12, YNBH11, YXM^{+}15, Zag96, ZHT07, ZFSL19, LLRR08, WSW16]. Field-Guided [LBZ^{+}11]. Field-of-View [BSEN18]. Fields [AZC^{+}12, BHHM19, BGT12, BZGV14, BHSH15, CMLZ08, CKW^{+}12, CLB11, DHL09, ES05, FS19, FA15, GPR^{+}01, GKT^{+}08, GSS^{+}19, GY02, GJR^{+}14, GYK^{+}16, GDN^{+}07, HMTR19, HA04, HE06, HLL97, HVSW11, HRN^{+}03, JBS06, KL14a, KWH00, KKSM19, LPKK12, LJL04, LVRL06, LCS06, LTWH08, LJHY14, LFR03, MBS^{+}04, MCHM10, NKH11, NKH14, NJ99, NLS11, NDS10, PQMCR17, PZ11, PYW^{+}16, PvdBC^{+}11, PMW13, RKG^{+}11, RKWH12, SCT^{+}10, SHM^{+}07, SWC^{+}08, SK98, SLB04, SWCR15, SRW^{+}16, SCKR08, SBW17, SH00b, SZ12, SS13b, TWC^{+}18, TWHS05, TN13, TN14, TP12, TKW08, WC09, WC10, WTVP11, Wen14, WJE01, WPG05, WGS07b, WPL96, YQK^{+}17, ZWZ^{+}13, ZSL^{+}16, ZBG^{+}17, ZPP05, vLdL03]. Fighting [VF13]. Figure [KCC^{+}17]. Figure-Ground [KCC^{+}17]. Figures [LYY^{+}16a]. Filament [LSY^{+}18, MK09]. Filament-Mesh [LSY^{+}18]. Filaments [SPN^{+}16, ZW07]. Files [BSO^{+}12]. Filesystem [BYB^{+}13]. Filling [BB12, MSSH14, MM08, Wen14, WFM^{+}12, DSF^{+}14, WFG^{+}19]. Film [YNYH06]. Films [LHC16, VAW^{+}17]. Filter [BC12, CC08]. Filtered [ASDW14, LKC09a, SP07, Wea10]. Filtering [BRSP18, BTH^{+}13, DMR04, JWD^{+}14, LWC^{+}18, NHPN14, RML12, SJ06, SBE^{+}15, TLQ^{+}08, WC09, WC10, WYP^{+}15, WHX^{+}19, WDC08, ZDH^{+}19, ZFAT11]. Filters [LS02, MMMY97, SCKR08]. Finding [AG16a, BOZ^{+}14, KHW^{+}09, TWSM^{+}11]. Findings [JHKH13, KDX^{+}12, LKD19]. Fine [DWBR06, UBH19]. Fine-grained [DWBR06]. Finger [JNC^{+}15]. Fingers [LBS^{+}19, TMDO15]. Finite [AWB11, BGT12, BTB^{+}04, BSL^{+}12, FSHH12, GT17, GT19, GHP^{+}16, MNKW07, NKH11, NLKH12, NKH14, SBM^{+}06, WBH04, XHF12]. Finite-sized [GT19]. Finite-Time [BGT12, GT17, GHP^{+}16].

Firefly [SEB19]. First [BYA15, LBS^{+}19, MAWM11, NO97, TMM^{+}13]. First-Person [TMM^{+}13]. Fish [ZHF^{+}19]. Fishes [BDSS18]. Fisheye [BGR06, CLB^{+}16, GKN05, WWZ^{+}19, WLH^{+}12, XCZ^{+}19b]. Fishtank [DJK^{+}06, PFK^{+}08]. Fit [TTR10]. Fitmaps [PPT^{+}11]. Fitting [CWQ^{+}07, KZW^{+}16, LLW06, OK11, YLSL11]. Five [RHR16, SNR14, ZWA^{+}13]. Five-Level [SNR14]. Fixation [TTS10]. Fixed [Lin14f]. Fixed-Rate [Lin14f]. Flame [SZN^{+}18, YL18]. Flames [BWP^{+}10]. Flat [NWHWD16]. Flat-Shaded [NWHWD16]. Flattening [GSZ^{+}13, PMCS11, SFH06, SF04, ZKK02]. FLDA [HLG^{+}14]. Flexibility [BPG12]. Flexible [BG07, CvW11, CCS12, HDSC19, KLKS10, KS02, PS06, PHF07, SLB04, SRM19]. Flexibly [HOG^{+}12]. Flicker [WLB^{+}14]. Flight [AAGS19, KWDG11]. Flip [GCT17]. Floating [FM12a, Gor02, LI06, Lin14f]. Floating-Point [FM12a, LI06, Lin14f]. Flood [KWS^{+}14]. Flooding [RWF^{+}13]. Floorplan [SWF^{+}16]. Flow [AHK^{+}17, AH11, Ano14h, BS95, BT13, BB07, BMLC19, BJB^{+}12, BPB14, BPM^{+}13, BMST97, BPS^{+}11, BSV11, BHSH15, CFM^{+}13, COJ15, CDR^{+}18, DMC^{+}12, ES01, FC95, FPH^{+}08, GGTH07, GNBP11, GST16, Guo09, GZL^{+}14, GZ14, GHS^{+}19, HMTR19, Har16, HEWK03, HLNW11, HLG^{+}14, HZ13, HPC^{+}13, JKLG08, JEH02, JCG08, JM10, KLC09, KSG^{+}16, KBE^{+}18, KHS^{+}18, KJW^{+}18, KM96, KGP^{+}13, KGJ09, KGG^{+}12, LvWJH04, LGV^{+}16, LM05, LCS^{+}12, MOJB^{+}19, MEB^{+}14, NTT^{+}19, OJ12, OJCJP16, PW95, PH07, PW13, RKWH12, RLL^{+}13, RYL^{+}18, RC06, SPS06, SSMG13, SYM14, SBV^{+}11, SWR^{+}13, SHM^{+}07, SWC^{+}08, SPO^{+}12, SK98, STM08, TMWS13, TWSK14, TWSS16, TWC^{+}18, VP04c, WRF^{+}11, WZF^{+}04, WTVP11, WSE07, Wen14, WGS07b, WRT19, WG12, WT10b, WYM12, XLS10, YDGM17, YDJ^{+}19, YLK12, YS17, ZYLL09, ZWZ^{+}13, ZT09, vPBB^{+}10, vPBB^{+}11, vFWTS08, CRI06]. Flow [LTWH08]. Flow-Based [DMC^{+}12, KLC09]. Flow-Sensitive [KGG^{+}12]. FLOWLENS [GNBP11]. Flows [CVC10, FC95, GT17, GT19, GHP^{+}16, HPAW07, KKKW05, LCC^{+}17, LXT18, SHVV16, SvdBLM11, VHLL14, WSTH07, WS01]. FlowVizMenu [VMCJ10]. Fluid [AB06, ATT12, BTT09, CMF12, DRW16, GLRH13, GGTH07, GT19, KWC^{+}10, LGS^{+}11, LBL19, LTKF08, VHLL14, YESK95, ZWJZ12]. Fluids [BB07, BK17, BKKW19, BLS12, CMHL11, HMT10, KAK^{+}07, MEB^{+}14, PT17, ySKK07]. Flux [GPP^{+}16]. Fly [ZD18]. FlyAR [ZHLR14]. FlyCap [XLC^{+}18]. Flying [PLE^{+}18, XLC^{+}18]. Flythrough [CORLS96]. Foam [BKKW19, KLCK17, LLCD11]. FoamVis [LLCD11]. Focal [BE06, GMS19, IMS15]. Focus [CVS^{+}19, CWK^{+}07, DSP^{+}17, FHS^{+}12, GNBP11, HS11, LHC10, QWC^{+}09, RWF19, TWSK14, TS08, VFSG06, WLT08, WC11, WWLM11, WP16b, ZZG^{+}12, vLBB16, zBBKN14, CDAF18, Kin10, NH06]. Focus-and-Context [GNBP11, zBBKN14]. Focus-Dependent [CWK^{+}07]. Focus-tunable [RWF19]. FocusAR [CDAF18]. Focused [BDW^{+}08]. Focusing [BW00, GDKB17]. Fold [NMGK17]. Folding [BSH^{+}16, ERHRF10]. Foldover [MZX15]. Foldover-Free [MZX15]. Folks [GJZ^{+}12]. Font [ACS^{+}18]. Footprint [CDK04, Ros11]. Footprints [IDA^{+}14]. Footsteps [NTS11]. Force [KKKT18, YML^{+}17]. Forces [HEG^{+}17, LA11, ZK10]. Forcing [BTT09]. Forecast [FKRW17, HMZ^{+}14, KTB^{+}18, KRRW19, LPCRH19, ME19]. Forecasting [MHR^{+}11]. Forecasts [QM16]. Foreground [LLC15]. Forensics [AERA14]. Forests [ZWLC19]. Form [AL11, DQ07, FvdPT97, GD01, RSB96, Stü98]. Formal [VFR13]. Formalizing [MWN^{+}19]. Forman [LLT04]. Formation [HSK14, NTT^{+}19, WLT18a, ZFS^{+}19]. Formations [WXW^{+}19]. Forms [Elb98, SZY^{+}18, Wea09]. Formulation [SCT06, WT10b]. Forums [FZCQ17]. ForVizor [WXW^{+}19]. Forward [CLEK13, HKH^{+}12, PSM06, WXKP14]. Four [BVW^{+}07, HB13, SZB^{+}09, TSA14]. Fourier [ES05, DNP07]. Foursquare [KTE15]. FOV [WP19]. Foveated [ACR^{+}19]. Fractal [CHF95, CHF96, KW10]. Fractals [VML97]. Fractional [KBVH17]. Fracture [AHK^{+}17, GMD13, Vis15]. Fracturing [BHTF07, GSM^{+}14]. Fragrance [HQS18]. Frame [HM95, JHW^{+}14, KPR^{+}15, MPT03, SSB^{+}17]. Frame-Rate [SSB^{+}17]. Frame-to-Frame [MPT03]. Frameless [FSTG16, PK16]. Frames [SWTH07, SM17]. Framework [AT05, Bac07, BKDE00, BDSW13, BC12, BTS^{+}18, CJ10, CVG13, CMPC06, CC12, DiV15, DK11b, DNN13, EMP09, EASS^{+}18, GKM^{+}15, HDSC19, HMZ^{+}14, JKMG07, KGD^{+}19, Lam08, LP02, LNS08, LWLM18, ME19, MMH^{+}13, MMAM14, MGM19, MP13, NM13, NDR96, PHF07, QT96, QK04, RESC16, RRJH18, SE17, SBM^{+}06, SHB^{+}14, SS06b, SPN^{+}16, SLK^{+}17b, SNR14, THM15, TWSK14, TWC^{+}18, TCM10, VMN^{+}19, VPF15, WJA^{+}17, WSD^{+}13, WWS^{+}18, WLS^{+}19, WLLM13, XHT^{+}07, XXM19a, XLS10, YYY16, YDC^{+}14, YS17, ZWM^{+}19, ZAM11, ZKG07, FHL10]. Frameworks [BVW^{+}07, CAN14]. Framing [HD11]. Frankenrigs [MAF11]. Fraudulent [LGM^{+}18]. Free [AL11, BK17, BC18a, DQ07, ELF13, FvdPT97, GD01, JHGH08, KV98, LLB^{+}06, LDX10, LLCM12, MZX15, MSM^{+}11, OMD^{+}12, RSB96, SH00a, SK99, Stü98, TYSN06, Wan08, WLW^{+}18, WQZ^{+}18, XLZ^{+}19a, LGV^{+}16, THT19, XLZ^{+}19b].

Free-Form [AL11, DQ07, FvdPT97, GD01, RSB96, Stü98]. Free-Viewpoint [LDX10]. Freedom [CMHL11, JWC05, ORC07]. Freeform [LKS13a, ZK14b]. Freehand [JK16]. Freely [BVW^{+}07]. Frequency [BSS^{+}13, BLM96, GPL^{+}11, HXF^{+}15, HQ13, LHLW10, OMD^{+}12, SHR^{+}11, WPC^{+}13, XJF^{+}08]. Frequency-Based [GPL^{+}11]. Friction [KKKT18, SSIF09]. Friction-Based [KKKT18]. Frictional [SLNB11]. Friendliness [RBK^{+}19]. Friendly [VPF15]. Frits [Ano11c]. FromDaDy [HTC09]. Front [Ano09e, Ano11h, Ano14i, Ano15f, Ano15g, Ano16i, Ano16j, Ano11j]. Frontiers [HKPC19]. Fronts [KHS^{+}19, ZK14a]. Frustum [CLT^{+}08, WHL16, LCM07, CLT^{+}08]. Frustum-Traced [WHL16]. FTMS [BvL06]. Full [AdLH13, KFL^{+}15, LWP^{+}06, TZL^{+}12, WLL^{+}16]. Full-Body [AdLH13]. Fully [CPW^{+}18]. Function [BMWM06, Gor02, IWR^{+}18, KMG^{+}06, LWS^{+}17, ME18, MWCE09, MJW^{+}13, SKK06, SG09, Sel15, SP96, WZK12, WQ07, ZWZD15, ZT09]. Functional [BKDE00, BSL^{+}14, DWBR06, FCSF17, HPNT18, JV09, JEG12, LLL^{+}19a, LSZ^{+}18, MF11, VAW^{+}17, WSLL12, tCMR08]. Functions [BEHP04, CR08, CM08, CM11, EMRY02, EDvW19, FGF^{+}05, GBPW10, GQGP17, GNP^{+}06, GNPH07, JWD^{+}14, KM96, KKH02, LF97, LHLW10, LLL^{+}12, LLL^{+}10, LHZ^{+}04, LLY06, NHPN14, RE01a, SCT06, SSC^{+}16, WTB^{+}19, XJF^{+}08, YCLJ12, ZDW^{+}05, vAPP^{+}11]. Fundamental [XYGL13]. Funding [MEV^{+}14]. Furniture [KHLM17, ZGM18]. Furry [QCH^{+}14]. Further [AMA06]. Fused [LDH09, PPvA^{+}11]. Fusing [CM16, WLDW11, YYFX18]. Fusion [BBB^{+}12, LKR^{+}18, WHR02, ZNZX16]. Future [JF16, KJH^{+}18, Min13, NJ16, Sat13, WFS^{+}19]. Futzing [AZL^{+}19]. Fuzzy [FM12b, SJL^{+}18a, SB04, VRW13, ZLC^{+}19]. FVA [RBK^{+}19]. FWP [XWL^{+}15].

G [CMFL16]. G-Buffer [CMFL16]. Gabor [GK95, LLD11]. Gaia [SJMS19]. Gain [BSSL19, FWK16, NSE^{+}12, WP19]. Gains [SHV^{+}18, ZLK^{+}18]. Gait [JBS^{+}18, WSH^{+}19]. Galaxy [MQF06]. Galleries [JFY16]. Galvanoscopic [PWG17, WG16]. Game [CWL12, CB15, MBZB12, WGYS18]. Game-Based [CWL12]. Games [AZM12, BEJK12, Chi16, GW13, LXC^{+}17, MH10, MY14, OA11, SK16a, SW17, VW12]. Gamut [SLGM09]. GAN [KTC^{+}19]. GANViz [WGYS18]. Gap [BE18]. Gaps [MDS^{+}17]. Garments [KZW^{+}16]. Garuda [NHN07]. Gaseous [WLMK04]. Gases [LIGF06]. Gather [HOG^{+}12]. Gaussian [HLCB18, IAIK16, JSG03, LZLS16, TWBBM17, WCZ^{+}11, WFW^{+}17]. Gaze [GJC^{+}17, GLB16, HZL^{+}19, KHH^{+}16, LSV^{+}18, MSM^{+}11, RKC^{+}16, RRD^{+}13, SKYS14, SLK^{+}17b]. Gaze-Aware [LSV^{+}18]. GazeDx [SLK^{+}17b]. GazeVis [SKYS14]. GDSPM [BGK11]. Gel [LLB^{+}06]. Gel-Free [LLB^{+}06]. Gender [PDBG18, WP19]. Gene [DWvW12, FNM13, KLK^{+}09, YDC^{+}14]. Genealogies [BDF^{+}10]. Genealogy [NGCL19]. GeneaQuilts [BDF^{+}10]. General [DLA^{+}09, GRT17, GUO00, IKLW14, LS02, LP02, MM11, MHDG11, PRA^{+}10, WBE^{+}06, YNM15, ZHGH11]. Generality [GBHP08]. Generalization [AA11, MWK14]. Generalized [BDHJ04, CCM13a, GMD^{+}16, GGLQ19, HB13, HLCB18, IHR01, IML13, JR07, LLWQ13, LWZ^{+}16, QLLM13, TP12, VvWvdL06, WTS^{+}07]. Generate [IHK05]. Generated [DLF^{+}09, JDA^{+}11, SGJM18, vHR08]. Generating [DC17, KNP04, KCC^{+}17, SG09, SRHZ11, THM15, WFM^{+}06, WHM14, YHH^{+}19, ZFL17]. Generation [ALM11, BB09, BKRE19, BMW17, CK05a, CCM11, ERL^{+}13, GKT^{+}08, HWM95, IYK01, KLM^{+}08, KDBB17, KBH13, LHJ^{+}18, LS07b, LML^{+}18, LSV^{+}18, MWCE09, NLS11, SSS^{+}14, SRML09, TW18, TLLH12, VABW09, WZC^{+}15, ZZL^{+}15, ZLZY18, ZT09]. Generations [FDC^{+}18]. Generative [BLL19, KTC^{+}19, LSC^{+}18]. Generators [LMD12]. Generic [DH02, GW06, HPJG08, LJH^{+}18, VCP08]. Genes [SAB^{+}16]. GeneShelf [KLK^{+}09]. Genetics [NKHC08]. Genome [BGM^{+}17, LBK^{+}18, NJBJ09]. Genomes [WWFT03]. Genomic [ADG11, ORRL10]. Genomics [GHL18]. Genotet [YDC^{+}14]. Genus [RvWT08]. Geo [CS18, CYW^{+}16, NASK18, PSKN06]. Geo-spatial [PSKN06]. Geo-Statistical [NASK18]. Geo-tagged [CYW^{+}16]. GeoBuilder [WTL^{+}09]. Geodemographics [SDW11]. Geodesic [DdL14, HLD^{+}08, PZLZ17, WT10b, XHF12, XWL^{+}15, YSS^{+}12]. Geographic [Fis07]. Geographical [TSH^{+}14]. Geographically [DB07, YDGM17]. Geographically-Embedded [YDGM17]. Geography [GDST16]. Geologic [CRB^{+}05]. Geology [CCB^{+}18]. Geometric [AIS18, BDK98, BMPB08, BGK11, BGR06, CGL^{+}17, CZN^{+}11, CCM^{+}13b, DBM^{+}06, DVC18, DQ07, GNCM^{+}16, HOT98, HHZH17, HLYL18, KBB^{+}12, LM96, MTB17, MTB18, PWG17, QT96, RF11, SK99, TD95, TL07, WHR02, WTL^{+}09, YNM15, ZHL^{+}09]. Geometric-Semantic [BGR06]. Geometrically [FST^{+}14, LSS^{+}15]. Geometrically-Correct [FST^{+}14]. Geometry [BSS^{+}13, BDHJ04, BBK07, BW08c, CZQ^{+}08, DMC^{+}12, DdL14, DSKA19, FH16, GBP12, GKL^{+}16, HDJ05, KV03, KSY16, KV08, LHV06, LBZ^{+}11, LJZ12, MCK12, MDS16, PH08, RYKL13, RNK^{+}15, SMG^{+}13, SCOIT05, VB13, YL08, ZLG^{+}06, ZDH^{+}19]. Geometry-Aware [SCOIT05]. Geometry-Based [CZQ^{+}08]. Geometry-Dependent [LHV06]. Geometry-Driven [ZLG^{+}06]. Geometry-Invariant [RYKL13]. Geophysical [KSDD14, WPB^{+}11]. Geoscientific [USKD12]. Geospatial [BDW^{+}08, TMH^{+}10, ZMT^{+}19]. Geotagged [JhR10]. Geovisualization [KMM^{+}13, LD11b, Rot13, WDSC07]. Geovisualizations [GR15]. Geriatric [MMH^{+}13]. Gestures [EASB19]. Get [GMY11]. Getting [RGFLL14]. Ghosted [ČB19]. Giant [PLE^{+}18]. Gibbs [VML97]. Gigantic [YSGM05]. Gigapixel [PK13, PST^{+}15]. GIS [WSLL12]. Given [KCA16]. GL4D [CFHH09]. Glance [HE99, RMCW19, YDK^{+}18]. Glanceable [BBB^{+}19]. Glass [Bro06]. Glimmer [IMO09]. GLO [SKL^{+}14]. GLO-STIX [SKL^{+}14]. Global [GLB^{+}06, JKRY12, JY17, KSY14, KGPB05, KPR^{+}14, LWZ^{+}16, LGY19, LSPW12, LKM^{+}18, NXSL13, NPPZ12, NDR96, PSKN06, QXF^{+}07, RvWT08, Sil95, SS95, VARS14, WWFT03, WPSH06]. Globally [JFZ^{+}18, LBG^{+}08, SJB10, ZM13]. Globbing [GHA^{+}08]. Globe [BAB^{+}18]. Glossy [WXKP14]. Glyph [DTW^{+}15, KW06, KBH13, MRSS^{+}12, MSSD^{+}08, TLS17, ZSL^{+}16]. Glyph-Based [DTW^{+}15, MSSD^{+}08, ZSL^{+}16]. GlyphLens [TLS17]. Glyphs [FIB^{+}14, FIBK17, GRT17, HLNW11, JKM06, PW13, SK10, SBW17, WPL96]. GMM [LWC^{+}18]. GMM-Inspired [LWC^{+}18]. Goal [MKT^{+}18]. Goal-directed [MKT^{+}18]. Goals [LTM18, SKY12]. God [ORC07]. God-Object [ORC07]. Gödel [GB08a]. Going [BOP15]. Good [BTS^{+}18, CLKS19, MS18b]. Gosper [AHL^{+}13]. GosperMap [AHL^{+}13]. GPF [LWC^{+}18]. GPLOM [IML13]. GPU [BC12, BFTW09, CK11, CFHH09, FT07, GW06, GWBO12, GN12, HSZ^{+}11, IMO09, KKM^{+}09, KMH11, KKPS08, LBG^{+}16, LQLX14, LCDP13, MN07, NKH11, NKH14, PSR17, PFK07, QCT13, RZHB^{+}08, RKSH11, RLW^{+}11, SMP11, SR17, SN10, SHC^{+}09, WX13, WZC^{+}15, WAG^{+}12, WJR^{+}13, vAPP^{+}11, vPVvdW10]. GPU-Accelerated [BC12, KMH11, PSR17]. GPU-Assisted [RLW^{+}11]. GPU-Based [CK11, NKH11, NKH14, vAPP^{+}11, vPVvdW10, CFHH09, PFK07, RKSH11]. GPUs [GPC^{+}17, HL09, KL14a, KM16, ME09, MGM14, VHBS16]. GRACE [MMH^{+}13]. Gradient [AMC10, BHST17, BLM96, CHM11, GAMD10, GWK12, HAM11, IVJ12, POD^{+}13, ZZL^{+}15, ZPG19]. Gradient-Based [BHST17, POD^{+}13, ZZL^{+}15]. Gradients [PT17, PMW13]. grained [DWBR06]. Grammar [DC17, PDFE18, SMWH17, SRM19]. Grammar-based [DC17]. Grammars [ARB07, LBZ^{+}11]. Granular [GSL^{+}17, MSW^{+}08]. Graph [AvHK06, AHK^{+}17, AHKMF11, AMA07, AMA08, ALMF19, BMGK08, BVB^{+}11, CZQ^{+}08, DDGL07, DLF^{+}09, EHP^{+}11, FYF^{+}18, FWSL12, FT07, FT08, FT09, GHN13, GHL15, HMM00, HSL19, HZM^{+}16, HEF^{+}14, IK95, JHH^{+}10, KKC15, KHZR18, KMLM16, KCM18, LWZ^{+}16, LLCM12, MWSJ14, MM08, NHEM17, NSL19, PSF09, PHE^{+}18, PBC17, PPP12, RSOW18, SNG^{+}17, SD12, SKL^{+}14, TWC^{+}18, VBW16, VJC09, WWB^{+}13, WSA^{+}16, WHZ^{+}18, WWS^{+}18, WCC^{+}19, WWZ^{+}19, WFC^{+}06, WCA^{+}17, XRP^{+}12, YDK^{+}18, YCHZ12, ZGB^{+}17, ZPG19, ZZBW08, vHR08, vHP09]. Graph-Based [HZM^{+}16, MWSJ14, ZGB^{+}17]. Graph-Cuts [WWB^{+}13]. Graph-Level [SKL^{+}14]. GraphDiaries [BFP14]. Graphic [DWA10, KOBH19, OAH14].

Graphical [BH09, HTP19, HMSA08, HFMC12, HKKS18, IIS14, JME10, KA12, MB01, RMW09, SSRE18, VB18, WCHB10, BDM^{+}17]. Graphicle [MB19a]. Graphics [AD12, Ano12h, Ano13j, Ano13k, Ano14j, Ano14k, Ano14l, Ano16k, Ano16l, BW17, CW11, DNP07, DB07, GW13, GBWI17, GLM06, HE12, HR07b, HLM10, Hub95, JvdLR13, Kas12, KSL^{+}17, LS07a, MH10, MY14, MGB^{+}19, MMCE09, MFS^{+}09, OA11, PZ12, QMK^{+}06, QK04, SK16a, SMWH17, SvW98, Sil17a, Sil17b, Sil18a, Sil18b, Sil19, SDW11, SR00, SW17, VW12, YSL^{+}13, Ano14c, Ano15a, Ano16f, Ano16a, Ano19a, Ano19f, KHE09, Ano17e, Ano18a, Ano18b]. Graphiti [SPEB18]. Graphitic [GKL^{+}16]. GraphProtector [WCC^{+}19]. Graphs [APP11, AMA11, BW11, BBD^{+}11, BBG^{+}09, BW08c, CHK04, CK10, DN12, DN13, DKM06a, DKM06b, DRMM13, GKN05, GWP^{+}16, HBW14, KMG^{+}06, KG06, MK13a, MNS18, NGCL19, OKSK16, PV06, RM13, STS06, SAC^{+}08, SZW^{+}19, SS13b, TN13, WFM^{+}06, WFM^{+}12, WSW^{+}18, WWS^{+}16, YLZ^{+}13, YDK^{+}18, ZGB^{+}17, ZGI^{+}18, ZBDS12, vLdL03, vdZCT16, vLBR^{+}16, TGSP09]. GraphSplatting [vLdL03]. GraphView [AvHK06]. Grasp [LFP07, TWMZ19]. Grasping [PB16]. Gravity [KCT^{+}17, YQK^{+}17]. Gravity-Reduced [KCT^{+}17]. Gregory [Ano13e]. Gremlin [ORRL10]. Grid [BH07, GW06, GIK^{+}07, HHM14, KEJK19, MHDH07, WSM^{+}09, YDG^{+}16]. Grids [BCS11, BS11, CBPS06, Dic14, FC95, GJ10, IHR01, MSHC99, MHDG11, SSS06, SM97, USM96, ZCW19, GN12]. Grotto [KLYE13]. Ground [KCC^{+}17, LB17, PGI^{+}17, ZBMY14]. Grounded [LKH^{+}16]. Group [AAB^{+}13, BKKF13, GFG^{+}14, JRHT14, SSKB14, tCMR08]. Group-to-Group [BKKF13]. Grouped [YDG^{+}16]. Grouper [LGLR14]. Grouping [BW14, SSEW19, YESK95, GSL14]. Groups [BOP15, DSC^{+}08, KO12, ZCJH12]. GrouseFlocks [AMA08]. Growing [MDS16]. Growth [BMJK09, VN19]. Guarantees [BLW14]. Guest [BSS18, BRS18, CGGR18, CKSB14, GGHZ19, KRTvW06, MY14, OO15, SGR06, SW06, AD12, Ano12g, Ano13o, Ano14o, Ano15m, Ano16p, Ano16q, BvdP12, BDC17, BHTY15, BGK11, BKL18, BLRW05, CCH14, CLS07, CW11, CFK12, CLS13b, DFQ12, DW17, DLM^{+}12, Ebe00, Ert10c, FBI07, GKR14, GJK15, GW13, GMM05, HK10, HVY16, HKQ13, HP04, HLM10, ILMH12, Joy02, JLS15, KMN04, KS14a, KKL11, KHSB11, KPGL12, KL14b, LS06, LSCN09, LST^{+}16, LBKD09, LABS10, MSW19, MH10, MW13, Moo03, MFS^{+}09, MYM08, NSvW11, OA11, PZ12, Pur09, Qin09, Rus99, RvWT05, SK16a, SK15, SSL08, SL11, SW17, TL11, Var01, VW12, WLW17, WM05, WY19b, vHMM^{+}11, vWMT04, vW11]. Guidance [AVP19, BSEN18, CGM^{+}17, FM06, HPvU^{+}18, SMT13]. Guide [OHWS13]. Guided [BBH^{+}17, BAAK^{+}13, BTH^{+}13, CD19, CCK07, DFD^{+}14, DCH^{+}17, GSPJ08, KOJC12, LBZ^{+}11, MFZ^{+}17, PBN^{+}13, PRH10, QM08, SKB^{+}18, SLGM19, TCM^{+}12, WHX^{+}19, WAG06, XFZ^{+}19, ZHZ15, KV06]. Guidelines [EF10, JF16, KHA10, SOK^{+}16a, SOK^{+}16b]. Guides [KSL^{+}17, PTM^{+}18]. Guiding [WLB^{+}14].

Hair

[CZZ17a, LHBF19, WBK^{+}07, YJL^{+}15]. Hair-Solid [CZZ17a]. Hair-Water [LHBF19]. Hairball [IBJ^{+}14]. Hairstyle [CK05a]. Hairstyles [VMT06]. Hairy [SBW17]. Half [MSM^{+}11]. Half-Silvered [MSM^{+}11]. Halos [BG07, EBRI09]. Hamilton [JFTW07]. Hand [AVP19, BSB^{+}18, CLL08, CLAL12, CGB^{+}13, HSR18, TRL^{+}19, TWMZ19].

Hand-Drawn [CLAL12, TRL^{+}19]. Hand-Held [AVP19]. Hand-Object [TWMZ19]. Handheld [PTM^{+}18, WLT^{+}18b, ZJH^{+}11]. Handles [ZWR14]. Handling [CMF12, CRPH10, HDBC15, PLW12, SH12]. Handoff [ZGI^{+}18]. Hands [MUS16, XLZ^{+}19a, XLZ^{+}19b]. Hands-Free [XLZ^{+}19a, XLZ^{+}19b]. Hands-On [MUS16]. HapTable [EASB19]. Haptic [AB01, CMHL11, CGB^{+}13, DDKA06, EPS^{+}15, EASB19, FM06, GLM^{+}17, JAO^{+}14, JWC05, KFN06, LA11, MTS07, ORC07, PCY08, PIS15, SO17, TYL^{+}18, UK12, WZW^{+}05, WRHR19, ZK17, ZH07, UBH19]. Haptics [BLRW05, HSR18, HBKS09, HQ04, NMN^{+}18, STH13]. Haptics-Based [HQ04]. Hard [LYY08, WHL16]. HARDI [PPvA^{+}11]. Hardware [CICS05, FM07, GLM06, GUO00, JvdLR13, JFTW07, KXW^{+}18, LMC02, MK09, QMK^{+}06, SPM^{+}13, SR00, WWL07, WKME03]. Hardware-accelerated [FM07]. Hardware-Assisted [CICS05, LMC02]. Hardware-Based [WKME03]. Hardware-Decompressible [KXW^{+}18]. Harmful [CG14]. Harmonic [YCLJ12]. Harmonics [BZGV14, LLW06, LPG12, ZTZX13, ZT09]. Harmonization [FCZ15]. Harnessing [McK09, PBE19]. HART [NKP^{+}15]. Hashed [WM19]. Hashedcubes [PSSC17]. Hasse [CdOKRV09]. Hazard [Zha14]. Hazy [ZHC18]. HDF5 [BSO^{+}12]. HDR [YNCP06]. Head [BWS^{+}19, BSEN18, CPW^{+}18, CDK^{+}17, DFG^{+}14, GIMS18, HIH^{+}18, HWHK16, HRISI15, HV00, HZL^{+}19, IK15, IDAK15, IAIK16, KBB^{+}12, Kra16, LCR16, LMD12, LH16, LHC10, LXRY18, MZH^{+}08, MIO^{+}15, dJOBNM17, OTKS15, PIN^{+}15, PGI^{+}17, QPNK18, RKSB13, RSBB17, SGQ16, SHV^{+}18, SBK^{+}11, WRHR19, XST^{+}18, XLZ^{+}19a, XLZ^{+}19b, KNR17].

Head-and-Eye [LMD12]. Head-Motion [LXRY18]. Head-Mounted [BSEN18, CDK^{+}17, GIMS18, HIH^{+}18, HRISI15, IK15, IDAK15, IAIK16, KBB^{+}12, Kra16, LCR16, LH16, LHC10, MIO^{+}15, dJOBNM17, PIN^{+}15, PGI^{+}17, QPNK18, SBK^{+}11, WRHR19, XST^{+}18, XLZ^{+}19a, XLZ^{+}19b]. Head-Tracked [HWHK16]. Head-Worn [DFG^{+}14, CPW^{+}18]. Heads [DH08]. Health [HHO^{+}17, KKL^{+}16, MMK^{+}17]. Healthcare [ZWA^{+}13]. Heart [BGP^{+}11]. Heat [GWK12, GSZ^{+}13, KLG^{+}16, ZHF^{+}07]. Height [BES12, GY02]. Height-Field [BES12]. Heightfield [CMK15]. Held [AVP19]. Helmholtz [BNPB13a, BNPB13b, BPB14, PPL^{+}10]. Help [KNKH19]. Helping [LBW19]. Hemicubes [Max95b]. Hemodynamic [GLvP^{+}12]. Hemodynamics [MVB^{+}17, OJCJP16]. Here [LKS^{+}19]. Heritage [LTPH17, WFS^{+}19]. Hertz [BDF16]. Hessian [ZPP05]. Heterogeneity [HCP^{+}15]. Heterogeneous [AWB11, CDC^{+}07, CCQ^{+}14, GSA^{+}09, GWE^{+}19, HKC^{+}12, KMDH11, KLL^{+}13, KSY14, LPCC17, SMER06, SSL^{+}12, Vis15]. Heuristic [PBN^{+}13, WAM^{+}19]. Hex [MCK12]. Hex-Based [MCK12]. Hexagonal [LLW15, NPPZ12]. Hexahedral [DGW11, Dic14, XC19]. Hexahedralization [JHW^{+}14]. Hi [ETO^{+}10, MSvG^{+}11]. Hi-Res [ETO^{+}10]. Hi-Trees [MSvG^{+}11]. Hidden [DW14, EIKS18, RFFT17, SGPR18]. Hiding [KSNY17]. Hierarchical [AHL^{+}13, Bac07, BKM13, BWK^{+}13, CLCQ12, CDS^{+}12, CSWP18, CLWW14, EF10, FKRW17, GW11, HSW11, HPvU^{+}18, Hol06, HFL18, IVJ12, IYIK04, KHD02, KMT14, LBGV13, MV06, MB01, MT01, PLC^{+}11b, PM08, SST^{+}17, SG09, Sil95, SDW09, SJ09, TS07, VC17, WXJD17, WXC^{+}08, XNT11, XF04, YSZ04, YWSC12, ZGW^{+}19].

Hierarchically [FWR00]. HierarchicalTopics [DYW^{+}13]. Hierarchies [AMA11, DYW^{+}13, GHJ^{+}98, GGPPS13, HDJ05, KSH03, KMKY10, KHM^{+}98, SE18, SHS11b, VBW16, WD09, WD10]. Hierarchy [AMA08, BJY^{+}18, BEHP04, CHK04, CL06, HJW99a, HJW99b, HSZ^{+}11, LSJ^{+}15, MB18b, SHS11a, Szy13, VHBS16, WBH04]. High [ALM11, Ano09b, BMR01, BHWB07, BZGV14, BSL^{+}12, BvL06, ChLYL09, CPK^{+}05, CAN14, CF10, Cse13, DAW13, DK11b, FAW10, FYY^{+}19, GPL^{+}11, GBPW10, GS14, GLX^{+}18, HBKS09, HE06, HB14, HAM11, KPR^{+}15, KPB14, KHZR18, LRP97, LAM10, LYS^{+}10, LMZ^{+}14, LKL^{+}15, LJL04, LT16, LBH18, LWW^{+}07, LCDP13, LMW^{+}17, MCK12, MNKW07, MZS^{+}19, MSHC99, NB95, NM13, NKH11, NLKH12, NKH14, NW11, OHJ^{+}11, PNML08, RZHB^{+}08, RYL^{+}18, RPHI08, SKK06, SSB^{+}17, SSS06, SSIF09, SS06b, SRCP02, SRCP03, SHC^{+}09, SB14, SPW07, SLW^{+}10, SJK^{+}12, TAE^{+}11, TFH11, TLLH12, TKBH17, WJ08, WSPVJ11, WM13a, WM18, WD09, WAG06, WMS98, WPS^{+}16, WDW16, XYC^{+}18, XXM19a, YRP19, YHJ^{+}17, YDG^{+}16, YNCP06, YRWG13, ZK06, vAPP^{+}11]. High-Density [LBH18]. High-Dimensional [ALM11, DAW13, GS14, LMZ^{+}14, LT16, LMW^{+}17, NM13, OHJ^{+}11, SS06b, TAE^{+}11, TFH11, TLLH12, TKBH17, WM13a, WM18, WAG06, XYC^{+}18]. High-Dynamic-Range [SLW^{+}10]. High-Fidelity [CPK^{+}05, MZS^{+}19, RYL^{+}18, RPHI08, SHC^{+}09]. High-Level [BvL06, NB95, SKK06, SRCP02, SRCP03, SPW07, WD09]. High-Order [BSL^{+}12, MCK12, MNKW07, NKH11, NLKH12, NKH14]. High-Performance [DK11b]. High-Precision [PNML08]. High-Pressure [SB14]. High-Quality [Ano09b, BMR01, BHWB07, CF10, Cse13, FAW10, GLX^{+}18, HE06, HB14, HAM11, LYS^{+}10, LKL^{+}15, LJL04, LCDP13, MSHC99, NW11, RZHB^{+}08, RPHI08, SSS06, SJK^{+}12, YDG^{+}16, ZK06]. High-Relief [YHJ^{+}17]. High-Resolution [CAN14, KHZR18, SSIF09, WJ08, WPS^{+}16, WDW16]. Higher [BBG^{+}09, LVRL06, SCT06, SBM^{+}06, SW13, TLM05, ZG06]. Higher-Dimensional [TLM05]. Higher-Order [BBG^{+}09, LVRL06, SCT06, SBM^{+}06, SW13, ZG06]. Highlighting [AHKMF11, GR15, SOK^{+}16a, SOK^{+}16b]. Highly [PT17, SPP^{+}14]. Hilbert [TC13]. HindSight [FDPH17]. HiPiler [LBK^{+}18]. HiPP [PM08]. Histogram [BRP19, SRML09]. Histograms [CBB06, CM11, DCM13, GPL^{+}11, IVJ12, LS13b, LLY06, NSW^{+}17, SSD^{+}08, SBSG06, ZCW19]. Histology [JST^{+}10]. Histopathology [FYTL19]. Historical [KZD^{+}10]. Histories [AAM^{+}12, BSSB10, BSKR19, HMSA08, TRd12]. History [CDW^{+}16, FDPH17, IWSK07, SSIF09]. History-Based [SSIF09]. Historygrams [JKRY12]. HMD [AVP19, LHH16]. Hockey [PSBS12]. Hodge [BNPB13a, BNPB13b, BPB14, PPL^{+}10]. HOLA [KDMW16]. Hollowing [WLW^{+}18]. Hologram [BSB^{+}18]. Holograms [HAGS16]. Holographic [ZKG07]. Homogeneous [WH18]. Homography [NZS^{+}17]. Homotopy [LM96]. Hong [QCX^{+}07]. Horizon [Ste98]. Horologium [MQF06]. Horologium-Reticulum [MQF06]. Hotel [WWL^{+}10]. Hotmap [Fis07]. Hotspot [KSW06]. Hotspots [MRH^{+}10, MHR^{+}11]. HPC [BSO^{+}12]. HRTF [SNM16]. HRTF-based [SNM16]. HTDs [WHR02]. Hub [QMK^{+}06]. Hub-based [QMK^{+}06]. Hue [CWM09b, KGZ^{+}12]. Hue-Preserving [CWM09b, KGZ^{+}12]. Huge [CMRS03]. Hull [FYZ^{+}17, GCT17, HSR13a, LGS12, WPZ^{+}11]. Hulls [BWC04]. Human [AZM12, ASW12, BWK^{+}13, CLAL12, CBL07, CZN^{+}11, CTM^{+}13, CTT^{+}16, CC12, CFEC17, Dan16, DCKY02, EIKS18, FGBB09, FBLS05, FWG09, GMS19, HHCL01, HOT98, HCMTH15, IPD^{+}07, JER16, JOR^{+}19, KDMW16, KPBL16, LD11b, LC10, MMK^{+}17, MT01, NQX^{+}05, PBO^{+}14, RJD^{+}07, RRD^{+}13, SAM^{+}07, SZY^{+}18, SSS13, TKC17, TSB^{+}05, TZL^{+}12, TM04, VBC^{+}16, WSD^{+}13, WBA^{+}14, WXZ^{+}16, YSD^{+}17, YQK^{+}17]. Human-Analytics [TKC17]. Human-Centered [LD11b]. Human-Computer [CC12, PBO^{+}14]. Human-like [KDMW16]. Human-Object [KPBL16]. Human-Scale [SZY^{+}18]. Humanoid [RBLW07, SSS13]. Humanoids [NT99]. Humans [KLD^{+}09, LBHW18, RGFLL14, RKA^{+}13, RCL^{+}15]. Hundreds [YHW^{+}07]. Hurricanes [JCRS09]. Hybrid [BW03, BBD^{+}11, BIAI17, BMA^{+}19, BPL^{+}19, CGC^{+}11, FCSF17, GPC^{+}17, HIH^{+}18, HFM07, HBC12, IDW^{+}13, LH09, LCNG14, MGS^{+}14, MSE^{+}06, MRS^{+}13, MHDH07, NKP^{+}15, RM13, SSB^{+}17, SLNB11, TK14, VMCJ10, YLX^{+}12, YR95, YBW^{+}19, YGFX19]. Hybrid-Image [IDW^{+}13]. Hydrodynamic [LBM^{+}06]. Hydrodynamics [SFBP09]. Hydrogen [BWP^{+}10]. Hyperboloidal [MSM^{+}11]. Hyperelastic [OBLN17]. Hypergraph [SW13]. Hyperlapse [LHJ^{+}18]. HyperSlice [TWBBM17]. Hyperspectral [KZD^{+}10]. Hyperspheres [AJ97]. Hyperstreamline [FA15]. Hypothesis [KLM^{+}08]. Hypothetical [KNKH19]. HySAR [HIH^{+}18].

I3D [GW13, SK16a, SW17, MH10, OO15, OA11, WY19b]. IBFVS [LvWJH04]. IBR [BHST17]. ICCD [TCYM09]. Ice [PSBS12]. Iconic [vWPSP96]. Idea [PGU12]. Identification [DANS10, LGG^{+}18, LSH07, MVN^{+}19, SZ11, VMN^{+}19]. Identify [DNN13, LGM^{+}18]. Identifying [AAM^{+}12, GRS^{+}19, LWLM18, NR18, WGS^{+}13, YDK^{+}18, ZCL08]. IDMVis [ZCD19]. IDSS [ZBZ^{+}13]. IEEE [Ano14j, Ano19a, DFQ12, ILMH12, Ano11i, Ano12e, Ano12g, Ano12h, Ano13j, Ano13k, Ano14c, Ano14j, Ano14k, Ano14l, Ano15a, Ano16f, Ano16a, Ano16k, Ano16l, Ano17k, Ano19f, BHGK14, BHTY15, BKL18, CR08, CCH14, CLS13a, DS17a, DS18a, DW17, EDF08, Ert10c, GMM05, HVY16, HKQ13, IHK^{+}17, Joy02, JLS15, KHE09, KKL11, KHSB11, LSCN09, LST^{+}16, MSW19, MW13, NSvW11, RvWT05, Sil17a, Sil17b, Sil18a, Sil18b, Sil19, vWMT04, vW11, Ano17e, Ano18a, Ano18b]. iForest [ZWLC19]. iForum [FZCQ17]. Illuminated [SZH97, ZM13]. Illuminating [EDK10]. Illumination [AZD17, BB09, BKA^{+}11, BZGV14, DWB^{+}06, HXF^{+}15, JKRY12, JY17, KSY14, KGPB05, KPR^{+}14, KJL^{+}12, LR11, MB18a, NW10, NDR96, QXF^{+}07, RBDG15, RJG17, SMP11, Sil95, SS95, SEB19, SYR11, WB08, WSE07, WLDW11, WPSH06]. Illusion [SNB^{+}17]. Illusions [BSWL12, PDBG18]. Illustrate [LSM03]. Illustrating [PGT^{+}08]. Illustration [BPG12, CYZ^{+}09, CSC06, JCRS09, KZL07, RE01b, RHD^{+}06, SEA09, ZHX^{+}11]. Illustration-Inspired [JCRS09, SEA09]. Illustrations [Elb98, LCC^{+}17]. Illustrative [BWF^{+}10, BGKG06, CFM^{+}13, CSC07, EBRI09, HGH^{+}10, LMT^{+}03, RBG07, VI18, WGM^{+}08, WBE^{+}06, XHT^{+}07, vPVvdW10]. Image [Ano10e, ADP02, BW03, BW04, BBH^{+}17, BVW^{+}07, BS02, BTC10, Bro06, CTM^{+}13, CMF^{+}18, CAH^{+}13, CMSW04, DMC^{+}12, DMR04, DZL^{+}14, DWK^{+}16, DL03, DHM13b, EHBA11, FWZQ13, FTB^{+}13, FYF^{+}18, FGBB09, FWT^{+}04, FZC^{+}07, GO15, GPC^{+}17, GSPJ08, GGZ^{+}18, GDKB17, HKC^{+}12, HSR13a, HSR13b, HSSK16, HLCB18, HLD^{+}08, HPC^{+}13, IHK05, IDW^{+}13, JST^{+}10, KLC09, KSY16, KOJC12, KEJK19, KCS^{+}16, KOJL^{+}14, KHPS07, KOF08, KHH^{+}16, LBD13, LvWJH04, LRN96, LWCS96, LG13, LS07b, LSWZ17, LLG17, LHFY12, LLC15, LR11, LY12, LSS^{+}15, MS08, ME09, NZS^{+}17, NDR96, PJ03, PGL^{+}12, PBCR11, PPM^{+}11, PMD^{+}07, STB18, SGB13, SSB^{+}17, SZK15, SK13, SKP07, SSW18, SYR11, TDLG19, TWSM^{+}11, Vas16, WW07, WGS07a, WWC^{+}14, WFS^{+}16, WGZ^{+}19, WWF^{+}19, WJR^{+}13, XLND11, XF04, XCZ^{+}19a, YCZ^{+}16, YL08, YCLL08, ZHZ15, ZWS^{+}17, ZJX^{+}15, ZKG07, vdCvW14, vLBB16]. Image-Based [BW03, BW04, BBH^{+}17, Bro06, CAH^{+}13, FWZQ13, FGBB09, FWT^{+}04, HSR13a, HSR13b, KHH^{+}16, LS07b, NDR96, PJ03, PGL^{+}12, STB18, SZK15, WW07, WGS07a, WFS^{+}16, WWF^{+}19]. Image-Centric [KOJL^{+}14]. Image-derived [GDKB17]. Image-Recoloring [KOF08]. Image-Set [HKC^{+}12]. Image-Space [CMSW04, HPC^{+}13, ME09, SKP07]. Image-to-Geometry [DMC^{+}12]. ImageAdmixture [ZCJH12]. Imagery [CDAF18, CYB08, JK16, LDH09, PSR17, VABW09]. Images [ASG15, BEK10, BJA^{+}19, BMW17, CYZ^{+}09, CWC^{+}06, CMP14, DMR04, DHM13a, ERL^{+}13, FWT^{+}04, FBLS05, GQM^{+}18, Hu16a, HNR^{+}06, IV11, JDA^{+}11, KPR^{+}15, KZD^{+}10, KWH19, KCC^{+}17, KCWI13, LAK^{+}11, LWYM12, LNHS16, LLR18, LAP19, MXW^{+}13, NRS15, OHH06, PZLZ17, PMH18, SKYS14, Sim07, SLK^{+}17b, SN97, SJM14, SJK^{+}12, TZC13, WLWL10, WQ07, XDN11, YL16, YYFX18, YFM01, ZDL03, ZHC18]. Imaging [KBH^{+}10, MNKT01, PIN^{+}15, RB18, SHC^{+}09, TT05, XSZ^{+}17, vAPP^{+}11]. Imagining [HKKS18]. Imitation [ZJH^{+}11]. Immaterial [LDH09]. Immersion [DHL09, LSSB12, LBS13, Ohl18, SBS16, SHV^{+}18, WGR^{+}18]. Immersive [ANR^{+}18, BGC^{+}11, BSB^{+}18, BKKF13, BL15, BIPS12, BRNB19, CB15, CDK^{+}17, DRHK07, ES01, GPL^{+}13, GS16, KNR17, KBS13, KMLM16, LLKN17, MWCR06, MGO^{+}19, PGRS13, RBLW07, SM12, SLC^{+}19, SKH^{+}19, TIS16, Tay02, TUG17, VSS08, YDJ^{+}19, YLC^{+}19, ZKM18]. Immiscible [MEB^{+}14, YYD^{+}19]. Impact [ARH^{+}15, AN13, AAM^{+}12, HCP^{+}15, KRH18, PQMCR17, STS^{+}14, VSS08, WGR^{+}18, ZM17, BPC^{+}10]. Impacts [MZH^{+}08]. Imperfect [BJM07]. Implant [DGBW09, DBW11]. Implantations [BSR^{+}14]. Implementation [BWS^{+}19, BC18a, JPD^{+}18, ME09, NDR96, PD04, WZC^{+}15, WR11]. Implementations [MPG^{+}14]. Implications [BMWM06, aKGS11]. Implicit [AG17, BHS12, CGD97, DQ07, Elb98, FGF^{+}05, HQ04, ICS^{+}14, LCS^{+}12, MGM19, SHS11a, SN10, WFM^{+}05, WWC^{+}14, XZB14, YT02, ZG12, ZCCB13]. Importance [CDBR14, Chr03, KW14, KDM^{+}16, LPG15, MDHB^{+}07, PBPP11, RvWT08, VKG05, VFSG06, WYM08]. Importance-Driven [KW14, VKG05, VFSG06, WYM08]. Impossible [LYY^{+}16a, SLF^{+}12]. Improve [HKKS18, LCL^{+}19, SLMA06, VB13]. Improved [EHS13, GBP19, JPLŠ16, LY06, MPK^{+}13, PFP^{+}11, WM19]. Improvement [HYB^{+}17, SAS16]. Improvements [DRW16]. Improving [DSS^{+}09, HBF08, JDA^{+}11, KFN06, LLD11, LWS^{+}17, LB03, OPH^{+}16, RSS14, SCKR08, VPB^{+}11, WHA07]. Impulse [WGF08]. Impulse-Based [WGF08]. Impulsive [SM04]. In-Depth [KJW^{+}14]. In-Place [PBA10]. In-Progress [SPG14]. In-Situ [WKSS05]. Inbetweening [Yan18]. Inclusion [KBE^{+}18]. Incomplete [KLC08, LLR18]. Incompressible [BK17, ICS^{+}14, SPS06]. Inconsistency [LLC15]. Incorporating [HOG^{+}12, KLCK17, MFS^{+}09]. Increased [BSSL19, MPG^{+}14]. Increasing [BE06, SCKR08]. Incremental [DKM06b, KLM04, PSPM15, PML97, SASS16]. Independent [PBA10, SM09, WKME03, Zha14]. Index [Ano97a, Ano98, Ano99, Ano00b, Ano01a, Ano02, Ano03b, Ano04a, Ano09a, Ano09d, Ano10a, Ano10e, Ano11a, Ano11e, Ano12a, Ano12c, Ano12d, Ano13h, Ano13a, Ano14a, Ano14c, Ano15e, Ano15a, Ano16f, Ano16g, Ano16a, Ano17c, Ano17d, Ano18a, Ano19a, CDS^{+}12, Ano13g]. Indexed [ZW18]. Indexed-Points [ZW18]. Indexing [BG04, CBL07, CdOKRV09, HK09, SR17]. Indicator [EDvW19]. Indicators [ALBR16, FRG^{+}19]. Indirect [ACTM12, NW10]. Indirectly [RKC^{+}16]. Individual [BOP15]. Indoor [PZLZ17, YYT16, ZX18, ZGM18]. Induced [Hu16b, WTS^{+}07]. Inductively [SRHZ11]. Industrial [AHR^{+}11, GUFM15, HLRS^{+}08, RGK^{+}13, SLMA06]. Industry [KAKC18]. Industry-Scale [KAKC18]. Inelastic [WLT18a]. Inertial [GT14, GT17, SJJ^{+}17]. Inference [BK12, GHL18, LLL^{+}19b, BDM^{+}17, WCHB10]. Inference-Based [BK12]. Inferred [ZDJ^{+}09]. Inferring [EFN12, NBM19, VABW09]. Infill [WAWS18]. Infinite [Bai13, MCA^{+}10]. Infinity [RS12]. Influence [FBL^{+}18, FIB^{+}14, HW12, KOK^{+}19, LR11]. Influences [BGC^{+}11]. Infographics [BAW16]. Inform [GS08]. Informal [BHP^{+}12]. Informatics [ZWA^{+}13]. Information [AJ17, AS05, Ano11i, Ano11l, Bac07, BKDE00, BI12, BDSW13, BEDF16, BBB^{+}12, BE09, CLS^{+}12, CJ10, CGJM19, DBD17, DCCW08, DRRD12, DLR09, ED07, EDF08, EF10, FWR00, GTS10, GTS11, HW12, HA06b, HMM00, HSC08, IC07, IIS14, JH13, JHKH13, JS98, JSR^{+}19, JRHT14, Kei02, KCA16, KSL^{+}17, KJW^{+}18, KC14, LMK07, Lam08, LBI^{+}12, LIRC12, LKH^{+}16, LNS08, LS10, ME09, McK09, MTW^{+}12, PBE19, PSM07, QWC^{+}09, RHY14, RNE^{+}17, SHM10, SMO^{+}13, SPB08, SLQW17, TTR10, TIC09, WZ08, WII^{+}12, XLS10, YaKSJ07, ZWA^{+}13, ZCW^{+}14, ZK08, vWN04, GLS17, HSKIH07]. Information-Aware [BDSW13]. Information-Theoretic [CGJM19, XLS10, CJ10]. Informational [BW14]. Informative [ZAM11]. Informed [DJ18, FA15, RH19]. InfoVis [Ano09e, WCHB10, Ano09d, Ano19g, Ano19h, HFM16, HAS11, KMN04, SW06, WM05, HA17]. InfoVis2009 [CWDH09]. InfoVis2009-1115 [CWDH09]. Infrastructure [BLO^{+}05, KZX^{+}14, PAB^{+}08]. INFUSE [KPB14]. Inhomogeneous [LSS^{+}11, SKLU^{+}11]. Initial [BE18]. Initialization [APV^{+}15, SG09]. Initiative [HKR^{+}08, WDC^{+}18]. Injection [MGJH08]. Injector [MGJ^{+}10]. Ink [WW07]. InkPlanner [LFW^{+}19]. Innovation [Min13]. Inpainting [GGZ^{+}18, HB14, KSY16]. Input [KZL07, XST^{+}18, YEII12, YCHZ12]. Insect [KWDG11]. Insensitive [AFRS05, Wan06]. Insertion [CLCQ12, HJLH19]. Inserts [LSK^{+}18]. Insight [GGZL16, PFG08, SND05, SNLD06]. Insight-Based [PFG08, SND05, SNLD06]. Insights [GGZL16, HRD^{+}19, KBGE11, OJK19]. Inspection [PTM^{+}18]. Inspired [FCSF17, JCRS09, LWC^{+}18, MZH^{+}08, NM13, RD05, SEA09, TFJ12, ZLDM16, vFWTS08]. Instabilities [LBM^{+}06]. Instance [KCK^{+}19a]. Instances [OKB^{+}19]. Instant [APV^{+}15, HKL17, KAK^{+}18]. Instructions [ZWBH13]. Instrument [BWS^{+}19]. Integer [NW11, WJR^{+}13]. Integral [BRP19, CMF^{+}18, FW08, FC95, GKT^{+}08, HGH^{+}10, LS13b, LDN11, LM05, MBH^{+}12, PLK12, SK98, Sun03, TKTN09, WH09]. Integrality [MCG12]. Integrals [WPC^{+}13]. Integrate [MDG00]. Integrated [GAMD10, MMH^{+}13, RSD^{+}13, SOL^{+}16, Wal12]. Integrating [DQ07, DCM13, OSS^{+}17]. Integration [CGC^{+}11, CWT^{+}08, CGB^{+}13, FM04, HSW11, JYC^{+}10, KPR^{+}15, KBGE11, RLK19, SCKR08, WCD^{+}19]. Integrator [GSS^{+}15]. Integrators [CAP18]. Intel [BWW^{+}12, Wal12]. Intellectual [ZCL09]. Intelligence [BCB10, YSZ^{+}19]. Intelligent [DZL^{+}14, FM06, LFW^{+}19, TLM05, YCHZ12]. Intensity [IVJ12, SWB^{+}00]. Intensity-Gradient [IVJ12]. Intent [NBM19]. Inter [GCL^{+}18, LSS09, MGM09, NSS14, TIS16, VBC^{+}16, WBA^{+}14]. Inter-Agent [GCL^{+}18]. Inter-Attribute [LSS09]. Inter-Personal [VBC^{+}16, WBA^{+}14]. Inter-Process [MGM09]. Inter-Reflection [NSS14, TIS16]. Interacting [ADWK^{+}17, BPP^{+}16, DDKA06, YHH^{+}19]. Interaction [BIAI17, BBG^{+}18, BGR06, CMHL11, DK13, EFN12, FWSL12, FDPH17, Guo09, GGZL16, HKR^{+}08, HEG^{+}17, HSTD18, JD13, JE13, KCPE16, KPBL16, KMLM16, Lam08, LHBF19, LBK^{+}18, LGYG12, LS10, MWCR06, MCG12, MJ09, MBZB12, MF11, NW15, PBO^{+}14, PGK16, PMvWC05, PSM12, RLM10, Rot13, RBLW07, SZS^{+}17, SKBE17, SBV^{+}11, Sim07, SS18, SDMT16, TWMZ19, TCM10, TFJ12, WBJ16, WLJ^{+}12, WB05, YML^{+}17, YaKSJ07, YSI^{+}10, YEII16, HWA15, JSV^{+}08]. Interactions [AL06, BOZ^{+}14, CPW^{+}15, DC17, GABJ07, GBCG^{+}14, HSR18, JYC^{+}10, JWL05, LIRC12, MSW^{+}08, PDF14, RJD^{+}07, SKY12, STM08, VBK17]. Interactive [ARB07, AAMH13, AM13, BSB^{+}18, BSS^{+}13, BGT12, BCH^{+}13, BE18, BSM^{+}13, BHZ^{+}18, BWW^{+}17, BBP08, BDSS18, BTB10, BJNN98, BWT^{+}11, BPS^{+}11, BIRW19, BGM^{+}17, BMW17, BFTW09, BTJ^{+}13, CGSQ11, CTGH13, CZZ17a, CWT^{+}08, CLT^{+}08, CA00, CYB08, CDZ^{+}09, CYW^{+}16, CZZ17b, CMK15, CK05a, CLRP13, CFHH09, CML^{+}12, CCB^{+}18, CMP14, DANS10, DZMQ16, DBW11, DvVH^{+}19, DWF^{+}19, DB07, EASB19, FYTL19, FPB17, FBTW10, FMH08, FSME14, FH16, FBM16, GFG^{+}14, GXH^{+}13, GW13, GSA^{+}09, GLK^{+}13, GHL18, GPK14, HLRS^{+}08, HBJP12, HKBR^{+}14, HB10, HE06, HSK14, HTP^{+}08, HSSK16, HFG^{+}12, HQ07, HC05, HTZ^{+}11, Hub95, IC07, IIS14, JH13, JV09, JFS16, JWS04, JFTW07, JBH^{+}09, JST^{+}10, JJ09, JCG08, JKRY12, JY17, KSH03, KTC^{+}19, KGS^{+}08, Kas12, KERC09, KLM^{+}08, KMDH11, KL96, KHS^{+}19, KZD^{+}10, KSY14, KOJL^{+}14, KKH02, KZX^{+}14, KMG^{+}06, KBH06]. Interactive [KPB14, KGPS13, KBE09, KKKW05, KSW06, KSBE18, KML96, KKW^{+}17, KCK^{+}19b, LSJ^{+}15, LB19, LCM07, LPP^{+}06, LS13a, LKHW04, LBK^{+}18, LDC96, LSR^{+}13, LCZ^{+}19, LWZ^{+}18, LGG^{+}18, LBLH19, LY12, LH14, LWD^{+}17, LDM^{+}18, LCL^{+}19, LDR00, LMT^{+}03, LMC02, LLCM12, ME19, MB18a, MV06, MKN^{+}07, MMB^{+}19, MAST16, MGPH06, MGJH08, MGKH09, MGJ^{+}10, MH10, MJL^{+}13, MY14, MI13, MAKM14, MRG^{+}15, MQF06, MWC^{+}12, MNC14, MHDG11, MPBM^{+}18, NKH11, NLKH12, NHYY18, NW10, NTS11, NT99, ORRL10, OH12, ODH^{+}07, OA11, PZ11, PKL^{+}18, PPL^{+}99, PPZ^{+}12, PHE^{+}18, PMCS11, PTMB09, PH08, QYH^{+}18, RE01a, RGK^{+}13, RHY14, RAL^{+}17, RLB19, RKK16, RGFLL14, RSA^{+}19, RBDG15, RJG17, RDB^{+}12, RP12, Rot13, RKG^{+}18, RSM^{+}16, RSR^{+}18, SSRE18, SSC^{+}16, SK16a, SRHH16, SMWH17, SS13a, SWC^{+}08, SKYS14, SLB04, STYC12, SDW11, SLK^{+}17b, SH12]. Interactive [SLMA06, SPEB18, SDES19, SKL^{+}14, SB17, SEB19, SW17, SFC^{+}07, SGM^{+}11, TKTN09, TCYM09, TDN^{+}12, TZC13, TLS17, TWBBM17, THT19, TLLH12, TKBH17, VW12, VJC09, WFKH07, WLJ^{+}12, WZQK04, WGS07a, WJ08, WP16a, WM16, WWS^{+}18, WLMP19, WV08, WEE03, WKZL04, WCB^{+}12, WAG06, WFM^{+}06, WDSC07, WQ07, WCQ^{+}09, WWL^{+}10, WLS^{+}18, WPSH06, YHW^{+}07, YHJ^{+}17, YDC^{+}14, YYT16, YRWG13, YSZ^{+}19, ZK06, ZK07, ZHT07, ZZC11, ZCCB12, ZCCB13, ZEC08, ZG06, ZBDS12, vHvdWvW02, vPBB^{+}11, vFWTS08, WOCH09]. Interactively [JHP^{+}14]. Interactivity [ARH^{+}15, BVV^{+}19, BEDF16]. InterAxis [KCPE16]. Interception [RFSP19]. Intercluster [MQF06]. Interest [AHSS14, HD12, JA18, MVN^{+}19, SXX^{+}19, vHP09]. Interesting [BBH^{+}17]. Interface [AGDJ10, BDS^{+}03, CDZ^{+}09, CML^{+}12, CLEK13, JKM01, KMDH11, KLK^{+}09, KFL^{+}15, LIM^{+}12, LTP^{+}05, MWSJ14, MKT^{+}18, OKB^{+}19, PFW12, PLE^{+}18, SK16b, SYYC11, WS06a, WM16, WCC^{+}19, Wu16, ZTA12]. Interfaced [MBT^{+}18]. Interfaces [ADP02, BLO^{+}05, CDF09, HF10, LMK07, LPG^{+}18, OBKP18, POD^{+}13, SKK06]. Interferometry [BPS^{+}11]. Interfirm [BCH^{+}13]. Interior [LL14]. Interlayer [GKL^{+}16]. International [Ano12i, Ano14m, Ano14t, Ano16m, Ano16w, Ano18f, Ano18h, Ano19g, Ano19i, GKR14, JLS15, KHSB11, LBKD09, LABS10, Ano13l, Ano13m, Ano17l, GJK15]. Internet [CYB08, CDM^{+}06, Hu16a, KJH^{+}18, LMZ06, PUNI11, VWvH^{+}07, ZGW^{+}14]. Interoperability [LKT13]. Interpenetration [ELF13, PB16]. Interpersonal [BRNB19, RJD^{+}07, RKA^{+}13]. Interperspective [MT05]. Interpolant [CM10, HTF97]. Interpolating [HB03]. Interpolation [AE13, BHB04, COJ15, CL09, Cse10, Cse13, GAMD10, HWM95, HTZ^{+}11, KCOY03, LWS97, LKC09a, MSA17, ME11a, ME11b, NZS^{+}17, Nie04, ONL^{+}12, PLK^{+}06, PFC18, SFA^{+}15, SKS12, TC09a, ZC06]. Interpolation-Based [COJ15]. Interpretable [KCK^{+}19b]. Interpretation [KRH18, MKN^{+}07, SDES19, WGZ^{+}19, ZWM^{+}19]. Interpreted [RBGH14]. Interpreting [LLL^{+}19b, PGT^{+}08, SDMT16, ZWLC19]. Interreflections [SVGR16, XA09]. Interrelation [JH13]. Interrogating [CML^{+}12]. Interrogation [LLL^{+}19b]. Interrogative [HKPC19]. Interruptible [RN19]. Interruptions [GWP^{+}18]. Intersecting [LGS^{+}14, PHF07]. Intersection [AL11, GD01, HH10, LQLX14, WLMP19]. Intersections [AR17, SOL^{+}13]. Interstitial [GKL^{+}16]. Interval [AWC10, CMM^{+}97, FMST96, SYK^{+}18, WD10]. Intervals [WCJ06]. Interventions [BHWB07]. Interview [HHH16, KPHH12]. Interviews [AZL^{+}19]. Interweaving [LLMB19]. Intestinal [RGF^{+}04]. Intracranial [JPLŠ16]. Intrinsic [LBD13, MFZ^{+}17, YXSH13, ZZH19, ZHL^{+}09]. Introducing [DS17a, DS18a, WLF^{+}19]. Introduction [AD12, Ano12g, BvdP12, BDC17, BHTY15, BGK11, BKL18, CCH14, CLS07, CW11, De 19a, DFQ12, DW17, Ebe00, Ebe07, Ert10c, Ert11a, FBI07, GKR14, GJK15, GW13, GMM05, HK10, HVY16, HKQ13, HP04, HLM10, ILMH12, Joy02, JLS15, KMN04, KS14a, KKL11, KHSB11, KPGL12, KL14b, LS06, LSCN09, Lin15, LST^{+}16, LBKD09, LABS10, MSW19, MH10, MY14, MW13, Moo03, MYM08, Nie95, Nie96, NSvW11, OO15, OA11, Ota17, PZ12, Pur09, Qin09, Rus99, RvWT05, SK16a, SK15, SSL08, SL11, SW17, TL11, Var01, VW12, WLW17, WM05, WY19b, vWMT04, vW11]. Intuitive [JFY16, LH03, SJM14].

Invariance [LJWF12]. Invariant [GST16, RBN^{+}19, RYKL13, TKW08, YSS^{+}12]. Invariants [BHSH15, RC06, SHM^{+}07, WH09]. Inverse [BKA^{+}11, CLEK13, HSK17, YKL^{+}08, ZDZ18]. Inverted [TLC^{+}10]. Investigating [BRH^{+}17, HJC14, KDX^{+}12, LD11b, TJW^{+}17]. Investigation [FFB18, JH16, SCKR08]. Investigations [KHSI04, PMCS11]. Investigative [BISM14, aKGS11, KFS^{+}19, cKJG^{+}12]. Inviscid [VHLL14]. Involvement [MPG^{+}14]. Ion [BvL06, GKL^{+}16]. IPD [LHH16]. IPSep [DKM06b]. IPSep-CoLa [DKM06b]. IR [IAS19]. IR-RGB [IAS19]. IRIS [HGH^{+}10]. Iron [FHG^{+}09]. Irradiance [DF96]. Irregular [BG04, Bon98, HCP^{+}16, SSS06, SM97, VP04a, WHL16]. Irregularly [ZG12]. ISA [LvWJH04]. ISMAR [KHSB11, LBKD09, LABS10, Ano15h, Ano15i, Ano15j, Ano15k, BSS18, BRS18, CGGR18, GGHZ19, KBB^{+}18]. Isocontour [Lin16b]. Isocontours [CPC09, PH11]. Isocube [WWL07]. Isodiamond [WD10]. Isogeometric [FH16, SF14]. Isometric [JBS^{+}18, PTC10]. Isoparametric [Elb95]. Isosurface [Ano96b, AE13, ASE16, BB09, BWC04, CBB06, CMM^{+}97, DCM13, ESN^{+}09, ENS^{+}12, GUO00, HL09, IK95, IYK01, LBS14, LSJ96, MCK12, MKW07, NN11a, NN11b, NK06, SSD^{+}08, SF14, SJ06, SH00b, TIW^{+}19, WFM^{+}05, WFKH07, WJ08, WC09, WC10, WWW^{+}19, WCJ06, ZK06]. Isosurfaces [GSDJ04, KW10, SSS06, SEH08, WD10, WPSH06, YWW14]. Isosurfacing [LB03]. Isotropic [SAS05, WYL^{+}19]. Isotropically [MMCE09]. ISP [HKC^{+}12]. Issue [BLRW05, DS17a, DS18a, Hag98, MY14, Nie95, Nie96, Var01, Joy02]. Issues [DBD18]. IStar [KHPS07]. iStoryline [TRL^{+}19]. Iterated [YYY16]. Iterating [BDFM17]. Iteration [SKLU^{+}11]. Iterative [DWB^{+}06, Fau99, HHO^{+}17, HTC09, KBGE11, KPS16, LDW^{+}15, WL08, YSZ04]. ITK [BVW^{+}07]. iTTVis [WLS^{+}18]. iView [ZAM11]. iVisDesigner [RHY14].

JackIn [KNR17]. Jacobi [JFTW07, TC17]. Jam [WZvdW13]. James [EDK10]. Jamming [SO17]. Jane [Bar05]. JanusVF [HR11]. Jarke [Ano13t]. Java [MB03]. Java-Based [MB03]. Jean [MFS^{+}09]. Jet [CDL^{+}16, DCH^{+}17, KHS^{+}18]. Jet-Stream [KHS^{+}18]. Jigsaw [GLK^{+}13]. Jim [EDK10]. JiTTree [LBG^{+}16]. John [Ano13d]. Joining [FG99]. Joint [CD14, JOR^{+}19, MGL07, RSOW18, SYM14, TAL^{+}07, XZM17, BGK11]. Joints [WGF08]. Joseph [EDK10]. Journal [Ano18f, De 17c, Flo18, Lin14e]. Journalists [BISM14]. Joy [Ano14g]. Judge [KNKH19]. Judgment [RKSB13, YHR^{+}19]. Judgments [KHSI04, SJK^{+}07, SSE15, YQK^{+}17]. Jump [KCT^{+}17, HL09]. Juniper [NSL19]. Jury [RCW^{+}18]. Just [LBG^{+}16, NWHWD16]. Just-in-Time [LBG^{+}16].

KAVAGait [WSH^{+}19]. KD

[WFM^{+}05, HL09]. Kd-Jump [HL09]. KD-Trees [WFM^{+}05]. Keeping [QH18]. Keim [Ano11d]. KelpFusion [MRS^{+}13]. Ken [Ano14g]. Kernel [WHX^{+}19]. Kernelized [BPP^{+}16]. Kernels [DBH14, HUPS14, RSR^{+}18]. Keshif [YEB18]. Key [CTT^{+}16]. Keyboard [LIRC12]. Keyframe [GSCI15, PLW11]. Keynote [Ano13u, Bai13, Cze12, Heg10, Min13, Sat13, Seq12, Tha11]. Keyword [FFB18, RGP^{+}12]. Keyword-in-Context [RGP^{+}12]. Keywords [IIS^{+}17]. kHz [BWS^{+}19]. Kick [PVF13]. Kick-off [PVF13]. Killing [HMTR19]. Kilo [BDF16]. Kilo-Hertz [BDF16]. Kinect [LZLS16]. Kinectrack [IFM14]. Kinects [TZL^{+}12].

Kinematic [COMP13]. Kinetic [LBL19, LSM03]. Kirchhoff [RLK19]. KL [Rob98]. Knitwear [CLZ^{+}03, GRS95]. Knot [CLCQ12, HHQH17, ZWJZ12]. KnotPad [ZWJZ12]. Knots [GHK97]. Knowledge [AS05, CB15, KBKG07, MWN^{+}19, SSS^{+}14, SS06b, WSH^{+}19, WGZ^{+}19, ZGI^{+}18]. Knowledge-Assisted [WSH^{+}19]. Knowledge-Based [KBKG07]. Knowledge-Transfer [ZGI^{+}18]. KnowledgePearls [SGP^{+}19]. Known [RNK^{+}15]. Kong [QCX^{+}07]. Krueger [Han95]. Kwan [Ano13f]. Kwan-Liu [Ano13f].

L [MCA^{+}10]. L-Infinite [MCA^{+}10]. Lab [Ano05c, BTC13, KTC^{+}19]. Label [CLG16, KOK^{+}19]. Labeling [BDY06, BHZ^{+}18, CG08, ČB19, FHS^{+}12, KCK^{+}19a, MTM^{+}16, NW11, XLZ19c, LSC08]. Labels [IYS13, KCK^{+}19a, TLH10]. Lagrangian [BTT09, GHP^{+}16, HOGJ13, JEH02, SP07, SWTH07, SXM17, SFB^{+}12, Wu16, YNBH11]. Lake [UDSL18]. Lambertian [MBT^{+}18]. Lamp [GHE19]. Lamps [RKK16]. Landing [GS16]. Landmarks [Hu16a]. Landscape [IV11, OHWS13, TSD09]. Landscapes [TSW^{+}07, WBP07]. Language [CCQ^{+}14, DR08, HB10, KCS^{+}16, LLL^{+}19b, NW15]. Languages [DWBR06, RBGH14]. Laplacian [ATLF06, ZHX^{+}11]. Large [AvHK06, AHSS14, AGL06, AABH^{+}16, AAMG12, AAMH13, APS^{+}14, APW16, BW11, BDJ14, BBD^{+}11, BAAK^{+}13, BDF^{+}10, BBP08, BBD06, BWT^{+}11, BTC10, CBPS06, CGC^{+}11, CMCL06, CWDH09, CGH^{+}19, CMK15, CPK^{+}05, CK05b, CAN14, CVC^{+}12, CLB^{+}16, CLWW14, Dic14, DGWC10, DYW^{+}13, DHR^{+}19, FSHH12, FSW09, GKN05, GCL^{+}15, GSS^{+}15, GPL^{+}11, GHGM06, GHA^{+}08, GCL^{+}18, Guo09, HSS11, HAAB^{+}18, HHWN02, HDSC19, HE99, HSSK16, HBC12, HC05, HTE11, IWR^{+}18, IV11, IDW^{+}13, JH13, JST^{+}10, JS98, JHP^{+}14, JDA^{+}11, KSH03, KRW19, KHDL07, KLK^{+}09, KKP^{+}17, KZX^{+}14, KJW^{+}14, KGJ09, KML96, KCM18, LKD19, LBK^{+}18, LSS^{+}11, LFH06, LFLH07, LZH^{+}13, LWL^{+}17, LXR19, LBS^{+}19, MS08, MGM14, MNS18, MGM09, MOC^{+}14, NLS11, PFW09, PFW12, PHJ^{+}10, PY09, PGI^{+}17, SMDS14, SXM17, SF19, SHS11b, SMER06, SCL^{+}12, SKH^{+}19, USM97, WGS07a, Wan11, WFW^{+}17, WYL^{+}19, WWZ^{+}19]. Large [WSLL12, WDSC07, XCZ^{+}19a, YHW^{+}07, YXM^{+}15, ZBG^{+}17, ZMT^{+}19, ZBDS12, dLVvL06, vHP09, vWN04, vdZCT16]. Large- [HBC12]. Large-Magnitude-Range [ZBG^{+}17]. Large-Scale [APS^{+}14, BWT^{+}11, Dic14, DGWC10, FSW09, GCL^{+}18, HAAB^{+}18, HDSC19, JST^{+}10, KKP^{+}17, LFLH07, LZH^{+}13, LXR19, LBS^{+}19, MGM09, MOC^{+}14, PFW12, PY09, SXM17, SKH^{+}19, WSLL12, YXM^{+}15, LFH06, LWL^{+}17, SMDS14]. Large-Scene [APW16]. Largest [SWC^{+}08]. Lark [TIC09]. Laser [GK95, KIS17, PWG17, VAB12, WG16, ZZSS10]. Last [LRN96, LBS^{+}19]. Latency [BWS^{+}19, FS14, FSTG16, FKS16, LBS^{+}16, LH14, OWS15, SGQ16]. Latent [HLG^{+}14]. Lattice [AEM09, BLW14, Cse10, Cse13, EM06, EVM08, GLX17, KEP08, Kim13, LPQF14, QXF^{+}07, WZF^{+}04, WLMK04]. Lattice-Based [QXF^{+}07, WZF^{+}04]. Lattice-Boltzmann [AEM09, WLMK04]. Lattices [AEM09, HAM11, PQF^{+}09]. Law [KH16, MMT^{+}14, HYFC14]. Laws [ZK10]. Layer [IHS17, LBM^{+}06, SvLF10, SWF^{+}16, WLSW08, YBW^{+}19]. Layer-Based [WLSW08]. Layer-Wise [SWF^{+}16]. Layered [BW11, BHW06, BH07, IHK05, KS00a, KS00b, KS01, KA12, MTS07, OKSK16, WWY14, ZHC18]. Layering [RASS17, ZHF12]. Layers [GQGP17, RBG07, SBV^{+}11, WTP^{+}19].

Laying [AHL^{+}13]. Layout [AMA07, BKM13, BSV11, DBD13, DLA^{+}09, DKM06b, FT07, GHN13, KDMW16, KMLM16, KCM18, MSvG^{+}11, MM08, NA19, SNG^{+}17, SRMOW11, WLD^{+}19, Won16, YDG^{+}16, YCHZ12, ZOC^{+}13, DMS^{+}08]. Layouts [BB12, DLF^{+}09, FT09, LPP^{+}06, OAH14, RSOW18, RLB19, SDW09, WSA^{+}16, YL06, vHR08]. Lazy [DWBR06, SM97]. LazyNav [PGI^{+}17]. LBM [PQF^{+}09]. LCDs [ZFSL19]. LDSScanner [XYC^{+}18]. Leader [GR15]. Leak [IHS17]. Lean [BWP^{+}10]. Leap [GS16]. Learn [BJY^{+}18, Chi16]. Learned [KFS^{+}19, TWMZ19]. Learner [KLL12]. Learners [RKA^{+}13, TNT17]. Learning [BJEYLW01, BHZ^{+}18, BBG^{+}18, BHP^{+}12, BOZ^{+}14, BC18b, CPG^{+}15, CZN^{+}11, CCJ^{+}19, DBH14, DNL^{+}06, EASS^{+}18, FWG09, HKPC19, HSK17, IHD^{+}18, KCM18, LZZ^{+}19, LLL^{+}19a, OAH14, RRJH18, RM15, SKKC19, SST^{+}17, TFO09, WSW^{+}18, YC14, YWV^{+}19, ZWM^{+}19]. Least [MDS16, MGM14, PLK12, PNML08, SZ11, WL08]. Least-Squares [MGM14, SZ11]. Legends [DWS10]. Legibility [CCAL12]. Legible [CWK^{+}07, War09]. Length [CXDR19, MCP^{+}06, VC17, Wan08]. Length-Preserved [Wan08]. Lens [BHST17, BTHD11, CMF^{+}18, EDF11, HTE11]. Lenses [BTB10, BTHD11, KCJ^{+}10, PBA10, RSA^{+}19]. lensing [THT19]. Less [SGQ16]. Lessons [KFS^{+}19]. Level [AJ19, BGB15, BW01, BM13, BMPB08, BvL06, CLMO17, EWWL98, FT07, HMBG01, HSH10, Hu16b, JvdLR13, KWP01, KCPE16, KKP^{+}17, LVRH07, LKHW04, LY12, LTKF08, LKT13, MB01, ME11b, NB95, NC07, SKK06, SRCP02, SRCP03, SJH^{+}07, SKL^{+}14, SPW07, SNR14, VB18, WGS07a, WOO17, WD09, WT10b, WY19a, XESV97, Zhu05, ZBDS12]. Level-of-Detail [HSH10, MB01, WGS07a, Zhu05, ZBDS12].

Level-of-Detail-Based [XESV97].

Level-Set [AJ19, BGB15, BW01]. Levels [AWHS16, BBG^{+}18, KCK^{+}19a]. Leveraging [ADG11, RKA^{+}13]. Levitation [TYL^{+}18]. Lexicon [PKL^{+}18]. LH [SBSG06]. Library [RLA^{+}13]. LIC [HA04]. LiDAR [WX17]. Lie [ZHGH11]. Lieberman [Ano13u]. Life [Bai13, BHP^{+}12, PSM07, RKKF19]. Lift [JK16]. Lift-Off [JK16]. Lifting [JAO^{+}14]. Ligand [LBH14]. Light [BB09, BKA^{+}11, DK13, GYK^{+}16, Hau97, IK15, IAIK16, IHS17, KKSE17, KKSM19, LDSM17, LPG15, MYM16, MTB18, SFC^{+}07, TAL^{+}07, WLL^{+}05, WSY07, WLHD17, WC13, XA10, YHLJ08, ZFSL19, ZW07]. Light-Field [IK15]. Light-Joint [TAL^{+}07]. Lighting [GUFM15, KISE14, KPH^{+}03, KPR^{+}14, LHLW10, LHV06, LG12, MBW^{+}07, SOL^{+}16, TLD^{+}12, WK13, ZM13, ZDM13]. Lightness [SSG16]. Lightning [KL07]. Lights [KW14, OBS^{+}15, WLMP19]. Like [JKM01, KCM18, SM12, WAWS18, BS16, FWK16, KDMW16, MMYK06, PWHK16]. limitations [DMAM19]. Limited [BSEN18, SHV^{+}18]. Limits [HW12]. Line [BPM^{+}13, CF10, DHL09, EHS13, Elb95, Elb98, EBRI09, FW08, FC95, GR15, HSW11, HZ13, IG19, JFZ^{+}18, KRW19, KZL07, KGP^{+}13, LM05, RMCW19, SK98, Sun03, WHZ^{+}18, WWZ^{+}18, XMRC17, YCLJ12, ZCFL15]. Lineage [NGCL19]. LineAO [EHS13]. Linear [ATK16, AE13, BS08, GAMD10, HJW99a, HJW99b, KC04, KH01, LVRL06, MYM16, NHPN14, PCG15, SJB10, WFS^{+}16, WB08, WBH04, YPI13, FGS19, SST^{+}17]. Linearly [NJ99]. Lines [IHR01, KHL99, KHS^{+}18, MZC^{+}16, SZH97, SF04, TWHS05, TKW08, WFR^{+}10, WT10a, WFG^{+}19, WTS^{+}07, XHT^{+}07, ZHX^{+}11, ZBB^{+}06, ZPP05, FGS19]. LineSets [ARRC11]. LineUp [GLG^{+}13]. LineUps [BDM^{+}17]. Linguistic [SS16]. Link [JRHT14, NBW14, OJK19, SSKB14, TNB11]. Linked [CvW11, FG99, JDL09]. Linking [HTL13, KPV^{+}18, NA19]. Links [SWS^{+}11]. Liquid [CM14, CMK15, Dic14, GLX17, JKM06, MLMF12, MJK06, SPCJL06]. Liquids [LIGF06, YYD^{+}19]. List [Ano00a, Ano01b, Ano03a, Ano04b, Ano05e, Ano06, Ano07, Ano08a, Ano09c, Ano10b, Ano11b, Ano12b, Ano13b, Ano15b, Ano15n, Ano17a, Ano17b, Ano19b, Ano19c, Ros11, Ano14b]. Listener [MAKM14, MDHB^{+}07]. Listener-based [MDHB^{+}07]. Listening [BSSB10]. Lists [HBW14, IK95]. Lite [SMWH17]. Literacy [BRBF14, LKK17, RM15]. Literary [HFM16]. Literate [WKD19]. Literature [BKW16, FHKM17, SZS^{+}17]. LiteVis [SOL^{+}16]. Liu [Ano13f]. Live [HB13, LMD12, MNZ^{+}15, MFZ^{+}17, SKH^{+}19, TKTN09]. LiveGantt [JHP^{+}14]. Liver [HCP^{+}15]. Lives [BSSB10, TKE16]. LiveSync [KBKG07]. Living [IWSK07, MLMF12]. LloydRelaxer [LT18]. LMap [NGK18]. Load [DL12, NLS11, SBS16, ZGH^{+}18]. Load-Balanced [NLS11]. Local [CRT04, DK13, DMC^{+}12, DVC18, EBB^{+}15, HLY10, JWSK07, JCC^{+}11, KV03, KO12, KW14, LS13b, LWZ^{+}16, LGY19, LLL^{+}10, LZLS16, LLY06, LKM^{+}18, MGW10, NGK18, NOB16, OHWS13, PSKN06, PMCS11, SWB^{+}00, SPP^{+}14, SSV18, WZW^{+}05, YC14, YNM15, ZZG^{+}12, ZCW19, vLBB16]. Local-to-Global [LWZ^{+}16]. Local/Global [LGY19]. Locality [JM10, MFS^{+}09]. Locality-based [JM10]. Localization [APV^{+}15, LD11a, VARS14, Zho16]. Localized [APS^{+}14, WGS07b, YBZW14]. Locally [Gué99, MB18b, RLNN11]. Locally-Ordered [MB18b]. Locate [VP09]. Located [IFP^{+}12, IC07, LHD18, TIC09]. Location [EPS^{+}15, GJ10, GJC^{+}17, PBK^{+}12].

Locations [LWL^{+}17]. Locomotion [GPK14, KFL^{+}15, MK13a, NSN14, OBKP18, PFW12]. Locomotive [ANR^{+}18]. Locus [ZOC^{+}13]. LOD [LKC09b, WS06a, PFK07]. Logarithmic [BBD06]. Logical [IBJ^{+}14]. Logistic [DvVH^{+}19]. Logs [BMJK09, GGZL16]. Long [LDN11, LCNG14, LD11b, MVN^{+}19]. Long-Duration [MVN^{+}19]. Long-Range [LCNG14]. Long-Term [LD11b]. Longitudinal [BN11, PGU^{+}13, SNLD06]. Look [BRP19, KH16, KCM18]. Look-Up [BRP19]. Looking [Fis07, TNT17]. Looks [CLKS19]. Lookup [MAST16]. Looming [RFSP19]. Loop [BKA^{+}11, KH19, LCP^{+}13, LRZM11, TGSP09]. Loops [FT13, XHF12]. Lossless [FM12a]. Low [ASDW14, CVS^{+}19, DRHK07, FSTG16, FKS16, GGZ^{+}18, HHM14, KM10, KWDG11, LLR18, OMD^{+}12, PSSC17, PS12, RZP^{+}07, WLHD17, WHX^{+}19, XYC^{+}18, YBZW14]. Low-Cost [DRHK07, KM10, RZP^{+}07]. Low-Dimensional [HHM14, XYC^{+}18]. Low-Pass [ASDW14]. Low-Rank [LLR18, WHX^{+}19]. Low-Resolution [PS12, YBZW14]. Lower [AS11]. LoyalTracker [SWL^{+}14a]. Loyalty [SWL^{+}14a]. LSTMVis [SGPR18]. Lumigraph [EHH^{+}19]. Luminaires [LPG15]. Lyapunov [BGT12, GT17, GHP^{+}16].

M [Ano13e]. Ma [Ano13f]. Machine [FWG09, KCM18, SKKC19, TKC17, ZWM^{+}19]. Machine- [TKC17]. Macro [MRSS^{+}13]. Magic [BHST17]. Magnetic [GPP^{+}16, SCT^{+}10, TYL^{+}18]. MagnetViz [SD12]. Magnification [PBA10]. Magnifier [ZZG^{+}12]. Magnitude [BDJ14, GUO00, KRHH11, ZBG^{+}17]. Magnostics [BBH^{+}17]. Maintaining [FSHH12]. Maintenance [HF11, SLMA06]. Majorization [WWS^{+}18]. Make [LKH^{+}16, OSSK12, YAE07]. Makes [BVB^{+}13]. Making [CDW^{+}16, IIS14, KDX^{+}12, PSTW^{+}17]. Management [BTC13, CMRS03, ET08, FWL17, GZL^{+}14, IG19, IYS13, KWS^{+}14, LSJ^{+}15, MBH^{+}12, PFK07, TLS17, WP18]. Manga [YHL^{+}17]. Manhattan [VAB12]. Manhattan-World [VAB12]. Manifold [GTLH01, HHQH17, KW11, KBH13, LPG12, PYW^{+}16, SJW07, TFO09, ZCFL15, ZWM^{+}19]. Manifolds [DMR04, ZWR14]. Manipulate [JAAL18]. Manipulated [SBS16]. Manipulating [GGL^{+}14a, HBM^{+}13, MJ09, PIS15]. Manipulation [BBK07, CK05b, CSC06, KC14, KYT^{+}18, MCG12, OKB^{+}19, SLMA06, TMDO15, YHL^{+}17, YJL^{+}15, ZWS^{+}17]. Manipulations [KISE14]. ManiWordle [KLKS10]. Manoeuvres [JPD^{+}18]. Manometry [KHZR18]. Mantle [SPO^{+}12]. Manufacturing [ACR^{+}19, JHP^{+}14, WAWS18]. Many [ASMP17, HBC12, KPBG13, OBS^{+}15, SHC^{+}09, Wal12, WC13, YDGM17, YCHZ12, ZK14a]. Many-Core [HBC12, SHC^{+}09, ZK14a]. Many-Light [WC13]. Many-to-Many [YDGM17]. ManyEyes [VWvH^{+}07]. ManyVis [RSD^{+}13]. Map [AABS^{+}14, APP11, APW16, BT13, BDM^{+}17, BDY06, BDD^{+}16, BSV11, CLG16, CDBR14, CM16, DWS10, EWWL98, FYWY16, GXH^{+}13, HGWW18, HR96, HSL19, Jen12, JhR10, KCH11, KLG^{+}16, KSBE18, NB12, SRMOW11, WS06a, WP16a, YYR17, YKL^{+}08, YSS^{+}12, ZMZM15, MTS07]. Mapped [FYWY16, PS12]. Mapping [AIS18, BKA^{+}11, CXM19, DT10, EWWL98, FST^{+}14, Guo09, GZ14, HAT^{+}00, HUPS14, HWL^{+}11, HZH14, JKRY12, JY17, KH19, KLS^{+}18, KKW^{+}17, LYY08, LBG^{+}08, LGQ09, LGY19, MZX15, MN07, MDB18, NWI17, PNML08, PSN10, PWG17, SGS^{+}19, SKP07, SJB10, SLS^{+}17, SHC^{+}09, SJ09, SCL08, THV^{+}14, XYGL13, YLY^{+}12, ZMG^{+}10, ZDM13, ZSG^{+}13, ZKK02, vHWV09]. Mappings [HSK17, NGK18, PMH18, RGE19]. Maps [AMJ^{+}12, AYRW09, APV^{+}15, BMJK09, BJB^{+}12, BSR^{+}14, CH17, JDL12, JVDF19, KMM^{+}13, LLHL14, MSME14, MV06, MKH12, MMK^{+}17, MDL^{+}19, NE04, NHB^{+}17, NW11, OBS^{+}15, RCSJ18, RASS17, RM13, SKB^{+}18, SWvdW^{+}11, SST^{+}17, SLQW17, TIW^{+}19, TGS11, VT08, WC11, WZC^{+}15, War09, XHL18, YDJ^{+}19, YGFX19, ZFL17, ZM17, vW14, KFS^{+}17]. MAQUI [LLMB19]. Marbling [AB06]. Marching [DSC^{+}08, DSS^{+}09, LB03, Nie03]. Marginal [XZM17]. Mark [MSM^{+}11]. Marker [KSNY17, LZD13, LJH^{+}18, NWI17]. Marker-Based [LJH^{+}18]. Markerless [CMPC06, KH19, LH09, MGL07, SM09, SLS^{+}17, WLT^{+}18b, XLC^{+}18]. Markers [AIS18, BDJ14]. Markov [PBL10, RS12]. Martian [YQK^{+}17]. Mashup [WDSC07]. Masking [GO15, LLPP19]. Masks [KM16]. Mass [AAFW17, CM14, DT10, EGG^{+}12, FT09, RB18, SVAC12, ZSG^{+}13, vLBR^{+}16]. Mass-Conserving [CM14]. Mass-Spring [SVAC12]. Masses [VAB12]. Massive [AK02, ADWK^{+}17, AA11, CL18, CCL^{+}16, KSY14, PSN10, SG05, vdEHBvW14]. Massively [LLB^{+}12]. Massless [SLNB11]. Master [TAK^{+}05]. Master-Slave [TAK^{+}05]. Match [LDW^{+}15, PYHZ14]. Matches [BNTM16]. Matching [CCM^{+}13b, FKLT10, FSHH12, HEWK03, KSBE18, LRP97, LWZQ17, LFP07, LB17, MGMP18, SSE15, TWSS16, WSW16, XLND11]. Material [AGDJ10, BDS^{+}03, HIH^{+}18, HKG07, IZM18, LLL^{+}10, POD^{+}13, RYKL13, SVAC12, SFA^{+}15, SRKL19, UMW^{+}12, WLLC15]. MaterialCloning [YL16]. Materials [BSS^{+}13, BHTF07, GKL^{+}16, HWW18, LB15, MBT^{+}18, RBDG15, Vis15].

Mathematical [GHK97, KLMA10, ZFSL19]. Mathematics [HP04]. Matrices [DWvW12, LBK^{+}18, OJK19, PDF14, ZCL09]. Matrix [BBH^{+}17, CMP09, CLRP13, EDF08, IML13, LLR18, TIW^{+}19, VMCJ10, WLLC15, YRWG13]. Matrix/Tree [YRWG13]. MatrixExplorer [HF06]. Matter [Ano08b, Ano09e, EBB^{+}15, JFTW07, KHA12, MSE^{+}06, SAM^{+}05, ZCL08]. Maxent [GHN13, MNS18]. Maxent-Stress [GHN13, MNS18]. Maximizing [SLF^{+}12]. May [CG16]. Maze [LYY^{+}16a]. Mazes [WLWL10]. MDS [IMO09, JFSK16]. Me [CLKS19, Ger17, MHS07, UKW19]. Mean [BSL^{+}14, CL09, CG14, FYWY16, PA06]. Mean-Shift [BSL^{+}14, FYWY16]. Meaning [BAW16, SGS^{+}19]. Means [RMW09]. Measure [GMS^{+}07, JAAL18, RKG^{+}11, RvWT08, RMCW19]. Measurement [LA11, ZHZ15]. Measurement-Based [LA11]. Measurements [PMT^{+}19]. Measures [CVG13, Far12, MJL^{+}13, PH11, ZLB^{+}05]. Measuring [CFEC17, CWRY06, FS14, KDM^{+}16, WTB^{+}19]. Mechanical [ATK16, LSZ^{+}18, WHFL14]. Mechanics [EGS03]. Mechanism [FWR00, YYD^{+}19]. Media [Bro07, CYW^{+}16, DK13, HWS17, LSB^{+}16, LWLM18, MS18a, MSW^{+}08, MYM16, SWL^{+}14b, SKLU^{+}11, TKE16, WH18, WLY^{+}14, XWW^{+}13, ZC03, ZFS^{+}19, ZCW^{+}14]. Medial [LYY^{+}16b, SAR96, SPB96, SCYW16]. Median [SB04]. Median-Based [SB04]. MEDIATE [PMvWC05]. Medical [DMR04, HJLH19, KCK^{+}19b, LLPY07, LRF^{+}11, PPM^{+}11, Sat13, SK13, SKYS14, SHC^{+}09, SLK^{+}17b, SEA09, UK12, WKB^{+}13, YL16, vLBB16, MTRP10]. Meeting [SBB^{+}18]. MeetingVis [SBB^{+}18]. MegaMol [GKM^{+}15]. MegaParallax [BCR19]. Melange [ERHRF10]. Melting [LIGF06]. MemAxes [GGJ^{+}18]. Members [Ano15h, Ano15k]. Membrane [DTT^{+}17, LHH^{+}12]. Membranes [Del08]. Mementos [TBHC16]. Memorability [BBK^{+}16, MPWG12]. Memorable [BVB^{+}13]. Memoriam [Ano05c, Bar05, EDK10, Han95]. Memories [TBHC16]. Memory [BSO^{+}12, CMCL06, CMRS03, GJ10, GGJ^{+}18, GJR^{+}14, GBP19, HBJP12, HE12, HSZ^{+}11, KSDA16, KYK11, KKCS98, LSJ^{+}15, PSSC17, PDBG18, RCW^{+}18, TSD09, VPF15, WX13, Zha14, ZD18]. Memory-Efficient [GJ10, GJR^{+}14, KSDA16, KYK11, KKCS98, WX13, ZD18]. Memory-Friendly [VPF15]. Memory-Hazard-Aware [Zha14]. Memory-Scalable [HSZ^{+}11]. Memoryless [LT99]. Mental [APP11, LLHL14, LS10]. Menus [GPL^{+}13]. Merge [SSEW19]. Merger [CWC^{+}06]. Merging [LH03]. Mesh [ALMF19, AZC^{+}12, CS18, CC08, CKLL09, DSC^{+}08, DSS^{+}09, FYP10, GABJ08, Gué01, GSPJ08, KSH03, KL03, KL07, LHH^{+}12, LRN96, LH03, LHBF19, LG15, LLT04, LLWQ13, LXB17, LSY^{+}18, LY06, LJWH08, LDC16, LGLR14, MZX15, ME11b, NWHWD16, OMD^{+}12, PA06, PGL^{+}12, PTC10, SRML07, TC09a, TDR10, WZW^{+}05, Wan08, WTW^{+}08, Wan11, WZC^{+}15, WWW^{+}19, WYP^{+}15, WHX^{+}19, XC19, YRP18, YRP19, YSZ04, YYSZ06, YL06, ZTP05, ZPS04, ZZC11, ZZL^{+}15, ZWZD15, ZDH^{+}19, ZFAT11, ZTA12, ZZBW08]. Mesh-Driven [PGL^{+}12]. Mesh-Free [OMD^{+}12]. Mesh-Guided [GSPJ08]. Meshes [AGL06, AFRS05, ATLF06, Bon98, CMRS03, CDM^{+}04, CHM11, CCS12, GHJ^{+}98, GWBO12, HHQH17, KTCG17, KWH19, LKT13, MW99, MKW07, MEB^{+}14, MCA^{+}10, PR00a, PS12, PCG15, RSOW18, Ros99, SE18, SG05, SZ11, SCYW16, TCL^{+}13, TNB11, THJ99, USM97, VP04a, VP04b, VCP08, VS11, VCL^{+}07, Wan06, XWL^{+}15, YPI13, YT02, YSS^{+}12, eYL07, ZC06, ZCW19, Zhu05, TGSP09, PR00b]. Meshing [BLW14, GCZL14, LLW15, MWK^{+}08, PD04, YWW14]. Meshless [BNPB13a, GLB^{+}06, PPL^{+}10]. Meshsweeper [Gué01]. Mesostructures [MNC14]. Message [Ano13n, Ano13o, Ano14n, Ano14o, Ano15i, Ano15j, Ano15l, Ano15m, Ano16n, Ano16o, Ano16p, Ano16q, BSS18, BRS18, CGGR18, CFK12, CLS13b, CKSB14, De 15d, DS16a, DS17b, De 18a, DS18b, De 19b, DLM^{+}12, Ert07a, Ert10d, De 16a, Flo17, GGHZ19, IKLW14, Lin11d, Lin11c, Lin12e, Lin12a, Lin13e, Lin14d, MB19b, Mue19, vHMM^{+}11]. Messages [BTH^{+}13]. Meta [ZGB^{+}17]. Meta-Analysis [ZGB^{+}17]. Metadata [IHK^{+}17]. Metal [AHR^{+}11, DBTH07, GBM^{+}12]. Metal-Artifact [AHR^{+}11]. Metamorphosis [BW01, LWCS96, LH03, LL05]. Metaphor [AABS^{+}14, KISE14, SGAS16, WBP07]. Metaphors [CVC^{+}12, MF11, ZK08]. MetaTracts [BWW^{+}17]. Meteorology [RBS^{+}18]. Metering [YWV^{+}19]. Method [AEM09, AWB11, BBBM18, BGB15, BTB10, CB15, Dru08, FYZ^{+}17, FA15, IHK05, IDAK15, KOF08, LLL06, LLLF08, LSY^{+}18, LCL^{+}19, MWCE09, ORC07, RGK^{+}13, Tay02, TS08, VS11, WH09, WHK15, WLMK04, WMK13, XTY^{+}11, XCZ^{+}19a, YSS^{+}12, YLY^{+}12, Zag96, ZCL08, JQD^{+}08]. Methodological [DBD18]. Methodology [AABW12, RHR16, SND05, SMM12, VJN^{+}15]. Methods [AL06, AJDL08, AMM^{+}08, AHH^{+}14, BS08, BNTM16, BN12, CA00, CWQ^{+}07, CMK15, CF10, CHM11, DLW^{+}17, DSC^{+}08, DLR09, EMdSP^{+}15, FCL09, GIMS18, HG01, KMLM16, LKJ^{+}05, LH11, LXL^{+}18, LD11b, SZB^{+}09, SBM^{+}06, SO17, TAE^{+}11, TWHS05, VF13, WHZ^{+}18, ZK12, HSKIH07]. Metric [CZN^{+}11, GSZ^{+}13, LJX^{+}10, LMZ^{+}14, MK13b, OJ12, PZLZ17, VCP08, WGS07a, WOO17, YYSZ06, CS18]. Metric-Dependent [VCP08]. Metric-Driven [LJX^{+}10]. Metrically [PMT^{+}19]. Metrics [DK10, GO15, GGZL16, JJ09, MDHB^{+}07, NHEM17, Vas16]. Metro [NW11, SRMOW11, WC11, WP16a]. Metrology [HKG07]. Metropolis [ZYM^{+}14]. MGV [AK02]. MIC [Wal12, BWW^{+}12]. Micro [LPF^{+}19, ZHLR14]. Micro-Robots [LPF^{+}19]. Microblog [BTH^{+}13, LLZ^{+}16]. Microfacets [MBT^{+}18]. Micrographs [DBTH07]. Micropolar [BKKW19]. Microscopy [BKA^{+}11, BJA^{+}19, HBJP12, MCS^{+}08, dLVvL06, WOCH09]. Microseconds [LBS^{+}16]. Microstructure [BZGV14]. Microvascular [GWE^{+}19, MAK08]. Midair [HQS18]. Middleware [BFE15]. Migration [KSG^{+}16]. Millennium [FSW09]. Mind [NBM19]. Minerals [GFG^{+}14]. Minimal [HYB^{+}17, LBW19, LBS^{+}16, PSF09]. Minimally [ES01]. Minimization [RB07, SSW18, WLT08]. Minimize [LT18]. Minimizing [OKSK16, YYSZ06]. Minimum [CXDR19, KMH11, VC17]. Mining [BISM14, GWP^{+}16, Kei02, KS02, LLMB19, LWC^{+}19, MOC^{+}14, VN19, dOL03]. MIP [EWWL98]. MIP-Map [EWWL98]. Mipmap [LKC09a]. Mirage [ZHF^{+}07]. Mirror [CSPN11, MSM^{+}11]. Mirror-Symmetric [CSPN11]. Mirrors [DTT^{+}17]. Mismatches [BNTM16]. Missing [SS19]. Mitigating [DBBF19]. Mix [AW14]. Mixed [Bro07, CLB13, CGB^{+}13, DH08, GKR14, GJK15, HKR^{+}08, JAM^{+}14, JLS15, KOJC12, KHSB11, KTWW13, KLD^{+}09, KLL12, LRM^{+}13, LBKD09, LABS10, NQX^{+}05, NW11, PLE^{+}18, PWG17, RPAC17, WDC^{+}18, XST^{+}18, HWA15]. Mixed-Initiative [HKR^{+}08, WDC^{+}18].

Mixed-Integer [NW11]. Mixed-Reality [KLL12, NQX^{+}05]. MixedFusion [ZX18]. Mixing [BJEYLW01, LBM^{+}06]. Mixture [LZLS16, WCZ^{+}11]. MizBee [MMP09]. MLS [LGM^{+}08]. Mo [XCZ^{+}19b]. MOBA [LXC^{+}17]. Mobile [BBBM18, BMWW18, BLIC19, BTJ^{+}13, CPW^{+}18, DH08, HBESB11, HAGS16, KPR^{+}15, KM16, KHSW17, LS07a, LHD18, LODI16, LSV^{+}18, NJ16, OKI15, PKMR15, RBDG15, RJG17, SOS^{+}17, VARS14, WRM^{+}10, XLZ^{+}19a, XCZ^{+}19b, XLZ^{+}19b, YC14, YLC^{+}19]. MobileFusion [OKI15]. Mobility [AAFW17, DSC^{+}16, WXZ^{+}16, ZFA^{+}14, vLBR^{+}16]. MobilityGraphs [vLBR^{+}16]. MObjects [RGK^{+}13]. Mobs [DGAB16]. Mock [VB18]. Mock-Ups [VB18]. Modal [CK05b, GMD13, MDS16, SLM18, WRHR19, DDBR^{+}19]. Modality [BGOJ16, DH08]. Modality-Driven [BGOJ16]. Mode [GPK14, PH07, vdCvW14]. Model [AB06, BMR^{+}19, BLL19, BSL^{+}12, BAF^{+}13, BMST97, BC12, BIRW19, CBL07, CZZ^{+}19, CK05a, DPW^{+}15, DK13, EASD^{+}19, EFN12, FG99, FBLS05, GHN13, GCL^{+}15, GPR^{+}01, GBP^{+}13, Guo95, HLRC^{+}12, HZL^{+}19, HLYL18, IZM18, JKMG07, JD13, JA18, JVDF19, KPBG13, KSI^{+}96, KEJK19, KPH^{+}03, KBVH17, KSDD14, LKL^{+}15, LKH^{+}16, LKT13, MD12, MOF09, MOF10, MGJ^{+}10, MHD^{+}18, ME09, MM11, MWC^{+}12, MTB17, MTB18, Mun09, MT01, NZ06, ORRL10, PY09, RKK16, RLA^{+}13, SSS^{+}14, SVAC12, SZD^{+}10, SASS16, SJH^{+}07, ŠTPV12, SO17, SRKL19, SSL^{+}12, TGH12, TL07, TLC^{+}10, THV^{+}14, TFH11, VSS08, WZW^{+}05, WCZ^{+}11, WPS^{+}16, WHM14, XFZ^{+}19, YLX^{+}12, YML^{+}17, YLSL11, YS17, ZPS04, ZWW^{+}12, ZGW^{+}19, ZWM^{+}19, ZFSL19, ZT09, vWN04, vLBB16]. Model-Agnostic [ZWM^{+}19]. Model-Based [BC12, CBL07, FBLS05, RKK16, SO17, WZW^{+}05, vLBB16]. Model-Driven [KSI^{+}96, SSL^{+}12].

Model-Guided [XFZ^{+}19]. Modelers [GDJ^{+}13]. Modeling [BTB^{+}04, BGK11, CVC10, CLRP13, CCB^{+}18, DC17, Del08, DCKY02, DQ07, EASS^{+}18, EHH^{+}19, FWZQ13, FGBB09, FM04, GRS95, GLX17, HWW18, HR96, HFG^{+}12, HQQ12, HQ04, KV03, KCA16, KPH^{+}03, KPB14, KGPS13, KKM^{+}09, LA11, LPS^{+}13, LvL12, LPQF14, MQV00, MAST16, MM11, NMGK17, NDM^{+}97, Ney98, NRS15, PDRK19, PK08, PW12, PDW^{+}14, Qin09, RBK^{+}19, RNE^{+}17, SPEB18, SKW^{+}11, SJM14, Sza18, TMH^{+}10, WLL^{+}12, WXJD17, WFW^{+}17, WX17, WBK^{+}07, WTB^{+}19, WZF^{+}04, WTP^{+}19, WKW06, XYS^{+}16, XSZ^{+}17, XA09, XA10, YJL^{+}15, ZZSS10, ZLD^{+}14]. Models [AG16b, ATK16, ADDG12, ASE16, BW04, BDK98, BW01, Bru17, CGD97, CC08, DvVH^{+}19, FBL^{+}18, GJG^{+}15, GNDV^{+}18, GMD^{+}17, HOT98, HC05, HKYM17, HG01, JK16, KAKC18, KTC^{+}19, KL14a, KMT14, KSY14, KML96, LH03, LL05, LJWH08, LR11, LS10, LZLS16, LSC^{+}18, LLL^{+}19b, Max95a, MJK06, MGM14, MF11, MWN^{+}19, MP13, NT03, OBLN17, PML97, RE01b, RLK19, STB18, SF19, SPK^{+}07, SGB^{+}19, TLQ^{+}08, TCYM09, TWBBM17, USKD12, Vas16, WKCB07, WH09, WQZ^{+}18, WSLL12, WSW^{+}18, XESV97, YFC^{+}19, YSGM05, ZK07, ZWM^{+}19, ZHC18, ZLDM16, ZJH07, ZSTR07]. Modification [LSJ^{+}15, MDS^{+}18]. Modified [GHA^{+}08, WZK12]. Modifying [AMA11, JAO^{+}14]. Modular [AHSS14, DiV15, OTKS15]. Modulating [UBH19]. Modulation [MDM10, RLL^{+}13, TIS16]. Modulator [IHS17]. Mohr [CRB^{+}05]. Molecular [BPG12, BLG^{+}16, CG07, DHR^{+}19, GRVE07, GBM^{+}12, GBCG^{+}14, HEG^{+}17, KBE09, LBH11, LBH14, PB13, SVGR16, TCM06, KFS^{+}17]. MoleView [HTE11]. Moment [BHSH15, CRT04, SHM^{+}07]. Monitor [TGW^{+}95]. Monitor-Based [TGW^{+}95]. Monitoring [BTH^{+}13, CLZ^{+}18, MKN^{+}07]. Monocular [HDBC15, VARS14, ZHQ^{+}07]. Monte [HKL17, LSPW12, RH19, Sbe97]. MOOC [FZCQ17]. Moore [TC13]. Morph [WR11]. Morphable [CLC^{+}15]. Morphing [CE01, RM15, SLGM09, WLL^{+}05]. Morse [CML^{+}07, CMLZ08, CDS^{+}12, GNPH07, GBHP08, GBP12, GKK^{+}12, GGL^{+}14b, GBP19, LLT04, SMN12, SZ12, Szy13, SS13b]. Mosaic [HCP^{+}16]. Moseying [AZL^{+}19]. Motifs [MRSS^{+}13]. Motion [AW03, BWK^{+}13, BCR19, BTV14, BWS^{+}19, BB19, BSWL12, CK16, CLL08, CLAL12, CBL07, FXG12, GXW^{+}18a, GXW^{+}18b, HOT98, HCMTH15, HZM13, JER16, KERC09, KCPS08, LPG^{+}18, LMD12, LZD13, LHC16, LYL19, LL19, LGYG12, LPLT11, LBS^{+}16, LC10, LSM03, LXRY18, LJH^{+}18, LKM^{+}18, MCP^{+}06, MK13a, PLW12, RK17, RFSP19, SKK^{+}14, SKC^{+}19, SYK^{+}18, SLG^{+}17, SZ11, SBE^{+}15, TAL^{+}07, WHK15, WLT^{+}18b, WSTH07, XLC^{+}18, XCZ^{+}19b, YN03, YAE07, ZJH^{+}11, ZZH19, FRG^{+}19, XLZ^{+}19a]. Motion-Blur [PLW12]. Motion-Sensitive [SZ11]. Motion-to-Pose [BWS^{+}19]. MotionExplorer [BWK^{+}13]. MotionFlow [JER16]. MotionRugs [BJC^{+}19]. Motions [ASvdP14, BvL06, HK09, HK16, yKL12, TH13, TMM^{+}13, XA10, XLZ^{+}19b]. Motivated [JKM06]. Mountains [CCB^{+}18]. Mounted [BSEN18, CDK^{+}17, GIMS18, HIH^{+}18, HRISI15, IK15, IDAK15, IAIK16, KBB^{+}12, Kra16, LCR16, LH16, LHC10, MIO^{+}15, dJOBNM17, OTKS15, PIN^{+}15, PGI^{+}17, QPNK18, SBK^{+}11, WRHR19, XST^{+}18, XLZ^{+}19a, XLZ^{+}19b, XCZ^{+}19b]. Mouse [LIRC12, RHR^{+}09]. Movement [ARH^{+}15, AA11, AABW12, AAH^{+}13, AAB^{+}13, AAFW17, BPS13, CYW^{+}16, DNN13, KTE15, RBK^{+}19, SMP17, SJL^{+}18b, ZMT^{+}19]. Movements [AdLH13, LKD19, PZ07, ySKK07]. Moves [FRG^{+}19, SSV18, ZWJZ12]. MovExp [PBO^{+}14]. Movie [BJEYLW01, LLKN17]. Moving [KGS98, LG12, MGM14, MEB^{+}14, PLK12, SP96, WBD14, ZLZY18]. MPI [CGC^{+}11, GPC^{+}17]. MPI-Hybrid [CGC^{+}11]. MPML3D [PUNI11]. MR [LBS13, MNKT01, ZDL03]. MR360 [RPAC17]. MRI [AS19, BPM^{+}13, JFTW07, KGP^{+}13, KGG^{+}12, MN07, STS07, STS10, vPBB^{+}10, vPBB^{+}11]. MRTouch [XST^{+}18]. Multi [APW16, BHHM19, BSM06, BJM07, BBP08, BL07, BM13, BDW^{+}08, CLG16, CWK^{+}07, CCM^{+}14, CLW18, CPG09, DVCD07, DDW14, EBB^{+}15, FT07, GSL^{+}17, GLG^{+}13, HA06a, HKG07, HBC12, HZM13, IWR^{+}18, IZM18, KKP^{+}17, KCK^{+}19a, KLS^{+}18, LLRR08, Lin16b, LPQF14, LD11a, LPK^{+}16, LRF^{+}11, MS18a, MS18b, MDS16, MiS^{+}18, ME11b, NSW^{+}17, PSTW^{+}17, PLS^{+}14, PTMB09, PLE^{+}18, PWIG18, PHF07, PBK^{+}12, PBC17, RLM10, SLGM09, SLM18, SKMH14, SLS^{+}17, SJH^{+}07, SGAS16, SKH^{+}19, TAK^{+}05, TIK15, TS08, VBK17, Wah14, WSW16, WLSL17, WCC^{+}18, WAG^{+}12, WRHR19, WS06b, WXC^{+}08, WCS^{+}18, YML^{+}17, YYD^{+}19, YBW^{+}19, YXG^{+}10, ZFL17, ZCL^{+}19, ZWC^{+}18, ZLC^{+}19, WOCH09, WBD14]. Multi- [HBC12]. Multi-Attribute [GLG^{+}13, SGAS16, WCC^{+}18]. multi-channel [WOCH09, WBD14]. Multi-Character [VBK17]. Multi-Charts [DDW14]. Multi-class [CCM^{+}14]. Multi-Client [SKH^{+}19]. Multi-Criteria [HZM13, PSTW^{+}17]. Multi-D.O.F. [TAK^{+}05]. Multi-Depth-Map [APW16]. Multi-Destination [ZFL17]. Multi-Dimensional [WXC^{+}08, ZWC^{+}18, ZLC^{+}19, LLRR08, LPK^{+}16]. Multi-field [LLRR08, WSW16]. Multi-Focus [TS08]. Multi-Focused [BDW^{+}08]. Multi-GPU [WAG^{+}12]. Multi-Granular [GSL^{+}17]. Multi-Instance [KCK^{+}19a]. Multi-Interaction [YML^{+}17].

Multi-Label [CLG16]. Multi-layer [YBW^{+}19]. Multi-Level [BM13, FT07, KKP^{+}17, ME11b, SJH^{+}07]. Multi-Material [HKG07, IZM18]. Multi-Modal [MDS16, SLM18, WRHR19]. Multi-Normal [ZCL^{+}19]. Multi-parameter [NSW^{+}17]. Multi-Phase [YYD^{+}19]. Multi-Pipeline [MiS^{+}18]. Multi-Projection [KLS^{+}18, PWIG18, SLS^{+}17, TIK15]. Multi-Projector [BSM06, BJM07, DVCD07, SLGM09]. Multi-Relational [PLS^{+}14]. Multi-Relaxation [LPQF14]. Multi-Resolution [CWK^{+}07, IWR^{+}18, PHF07, SKMH14, WLSL17]. Multi-Run [Lin16b]. Multi-Scale [BL07, CLW18, CPG09, EBB^{+}15, HA06a, KCK^{+}19a, PLE^{+}18]. Multi-Server [Wah14]. Multi-Stage [WCS^{+}18]. Multi-Threading [PTMB09]. Multi-Timepoint [BBP08]. Multi-Touch [LRF^{+}11]. Multi-Triangulation [MS18b]. Multi-User [BHHM19, PBK^{+}12, PBC17, RLM10]. Multi-Variate [LD11a, WS06b, YXG^{+}10]. Multi-View [MS18a]. Multi-Volume [PHF07]. Multiagent [SAC^{+}08]. Multibody [XZB14]. Multichannel [MNKT01, tCMR07, tCMR08]. Multicharacter [YLL^{+}12]. Multiclass [GCNF13, JVDF19, RAL^{+}17, WCG^{+}19]. Multicomputers [LRN96]. Multicore [CGC^{+}11, WJ08]. Multicriteria [SRMOW11]. Multidigraphs [AK02]. Multidimensional [CGSQ11, CZC^{+}15, DBD18, EDF08, EMdSP^{+}15, GK95, IML13, JCC^{+}11, KKH02, LAK^{+}11, LMZ^{+}14, LSPS10, LXR19, MMDP10, NA19, PNML08, SW13, STH02, VMCJ10, WZK12, WFW^{+}17, Wea10, XXM^{+}19b, ZKK02]. Multidomain [KJ12, NK06, YXG^{+}13]. Multifaceted [CSL^{+}10, KH13]. Multifield [BRT12, BBS^{+}08, DCK^{+}12, JWSK07, NN11a, NN11b, STS06, WKI^{+}17]. Multifield-Graphs [STS06]. Multifocal [BE06]. Multifocus [FWSL12, JE13]. Multiform [LSS^{+}11]. Multifragment [VF13, WX13]. Multiframe [ZFSL19]. Multigrid [CMF12, DGW11, DRW16]. Multilayer [ZPS04, ZFSL19]. Multilayer-Multiframe [ZFSL19]. Multilevel [AMA07, CK11, IMO09, IVJ12, LWS97, MNS18, YS03]. Multilinear [WL16]. Multimaterial [BLW14, MWK^{+}08]. Multimodal [BHWB07, BBB^{+}12, FWL17, GKL^{+}13, HBG11, RHR^{+}09, SS18, Vis15]. Multiperspective [CRPH10, WP16b, WP18]. Multiphase [MEB^{+}14, Wu16]. Multiplanar [HN13]. Multiple [AWHS16, AT16, BMGK08, BSKR19, BMR01, BBG^{+}18, BLLS17, BMST97, BNTM16, DZMQ16, GNCM^{+}16, GDST16, GK07, GGL^{+}14a, JME10, KERC09, KPBG13, KGS98, KBK11, LBD13, LMK07, LKD19, LD11b, MGJ^{+}10, MAF11, MTB17, MTB18, PBPP11, PLW11, QH18, RH19, RSD^{+}13, SM09, SGQ16, SGM08, SXX^{+}19, SLG^{+}17, SKY12, SSGM19, WCC^{+}19, XLC^{+}18, YDK^{+}18]. Multiple-View [BNTM16, KERC09, MTB17]. Multiples [APP11, LBK^{+}18, MDS^{+}17]. Multiplexing [DMAM19]. Multipole [Hau97]. Multiprojector [BE06, BMY05, JWS04, SM12]. Multiresolution [AB01, BRSP18, BG04, Bon98, CYB08, CL09, CMPS97, CSWP18, CWRY06, DL03, HG01, NW10, OR98, OR99, PD04, SPM^{+}13, SLNB11, SG05, SGM^{+}11, VP04a, WS06a, WD10]. Multiscale [AB06, CL06, GLRH13, Guo95, LHLW10, MMP09, MDS^{+}18, Mur95, RvWT08, SKU^{+}12, STH03, SGM^{+}11, TN14, WS09]. Multisensory [KJH^{+}18, PMvWC05]. Multistate [TMM^{+}13]. MultiStream [CSWP18]. Multisurface [CML^{+}12]. Multithreaded [LH09]. Multitouch [CML^{+}12, MCG12, SA19]. Multivalued [WKZL04]. Multivariate [BDSW13, CvW18, CXM19, CvW11, EG09, GHGM06, GHA^{+}08, GGA^{+}11, GCML06, Guo09, GXY12, HDSC19, HLCB18, HHG14, IML13, JFSK16, JBS08, KKSS13, KMM^{+}13, LS09, LT18, LWCC18, LS16, LDA12, NSL19, NGCL19, PHJ^{+}10, PV06, RML12, RASS17, RSA^{+}19, SWvdW^{+}11, SK16b, SSGM19, TIW^{+}19, TSH^{+}14, WFM^{+}12, ZW18, vdEvW14, HSKIH07]. Multiverse [SWR^{+}13]. Multiview [CMF^{+}18, LDX10, RHZN11, TCM^{+}12, WLDW11]. Multiviewpoint [Sim07]. Multiwavelength [LFH08]. Munin [BFE15]. Mural [JS98]. Muscle [AS11, CYZ^{+}09, DCKY02, TSB^{+}05]. Muscles [WGF08]. Museum [HSC08, MLMF12, RGFLL14]. Music [CQM10, FXG12]. Music-Driven [FXG12]. Musicians [JFS16]. Mutual [BBB^{+}12, TAK^{+}05]. My [BSB^{+}18, SS19, SHV^{+}18]. MyBrush [KPV^{+}18]. Myocardial [TBB^{+}08].

Name [SWY^{+}17, WBDS11]. NameClarifier [SWY^{+}17]. Nanocubes [LKS13b]. Nanoscale [AABS^{+}14, MAAB^{+}18]. Nanosphere [GKL^{+}16]. Nanostructures [MDS^{+}18]. Nanotechnology [QMK^{+}06]. Narrative [BMW17, FYF^{+}18, HD11, HDR^{+}13, RCW^{+}18, SH10, WLB^{+}14, WLF^{+}19]. Narratives [KBI^{+}18, SBB^{+}18]. Narrow [AS11, LKHW04]. Narrow-Band [LKHW04]. Narvis [WLF^{+}19]. Nasal [ZMH^{+}09]. Native [FYTL19, STYC12]. Natural [BPB14, BIPS12, FTB^{+}13, LCC^{+}17, LLL^{+}19b, MF11, NW15, NSN14, RJG17, SBK^{+}11, SLF^{+}12, TFJ12, XZX^{+}17, YN03, YON06]. Naturalness [KOF08]. Naturalness-Preserving [KOF08].

Navigable [WKW06]. Navigating [CVC^{+}12, CML^{+}12, CLB^{+}16, FWR00, JS98, SJMS19, SPK^{+}07, WS06a]. Navigation [AGN^{+}19, AN13, BFP14, BL07, BJNN98, BIPS12, CRI06, EDF08, GSA^{+}09, HD12, HMM00, HR11, IV11, KBKG07, LSJ^{+}15, LYY^{+}16a, LODI16, PvdBC^{+}11, PGI^{+}17, SW12, SM06, SAC^{+}08, TYSN06, TMM^{+}13, WHA07, WP18, ZHLR14, vWN04, JSV^{+}08]. ND [CXM19]. Near [ACR^{+}19, Ano96b, BBG^{+}18, DTT^{+}17, LH16, LHH16, LSJ96, RWBF18, RBDG15, XGS^{+}19]. Near-Eye [ACR^{+}19, DTT^{+}17, RWBF18, XGS^{+}19]. Near-Field [BBG^{+}18, LH16, LHH16, RBDG15]. Nearest [CK10, XLND11]. Nebulae [MKHD05, WAG^{+}12]. Needle [HJLH19]. Neighbor [CK10]. Neighborhood [DSF^{+}14]. Nematic [JKM06, MJK06, SPCJL06]. Nested [Mun09, WLSL17]. Nets [CD14, ST09, vHWV09]. NETSPEAK [RGP^{+}12]. Network [wAPS14, ABC^{+}19, BRH^{+}17, BF01, BEW95, BBH^{+}17, BKW03, CDK^{+}17, GKL^{+}13, HZM^{+}16, KAKC18, KDMW16, LLB^{+}12, MV06, MKN^{+}07, MPWG12, MGPH06, MJ09, NOB16, PGU^{+}13, SNG^{+}17, SFMB12, SHH11, SWW^{+}15, SSG12, SA06, SPEB18, VMCJ10, WLS^{+}19, WPZ^{+}16, YSD^{+}17, vdEHBvW14, vdEvW14, vdEHBV16]. Networked [DK11b]. Networks [AHSS14, AHK^{+}17, BFP14, BG04, BSKR19, BJY^{+}18, BN11, CGH^{+}19, CCM12, DVP^{+}18, DWvW12, FBL^{+}18, GRS^{+}19, GdBG12, GZ11, GWE^{+}19, HSS11, HSCW13, HS11, HF06, HFM07, HBF08, HKYM17, JHGH08, KCK^{+}19b, LPP^{+}06, LYL19, LZZ^{+}19, LLHL14, LSL^{+}17, MB19a, MK09, PS06, PHV^{+}18, RFFT17, RFLL18, SG09, SMER06, SS18, SGPR18, UMW^{+}12, VRW13, WGSY19, WKW06, XLZ19c, YDG^{+}16, YDC^{+}14, ZEC08, DMS^{+}08]. Neural [BF01, BJY^{+}18, JBH^{+}09, JSR^{+}19, JDL12, KAKC18, KCK^{+}19b, LYL19, LSL^{+}17, LBT^{+}18, PHV^{+}18, RFFT17, SLGM19, SGPR18, XLZ19c]. Neurobiology [BŠG^{+}09, WOCH09]. NeuroBlocks [AABH^{+}16]. Neurodegenerative [OIR^{+}17]. NeuroLines [AABS^{+}14]. Neuron [LHH^{+}12, UKF^{+}18]. Neuronal [AABS^{+}14, BJA^{+}19]. Neuroscience [BAAK^{+}13]. Neurosurgery [KOCC14]. Neurosurgical [BHWB07, JSV^{+}08]. Never [DBN06]. News [MBL^{+}06]. Next [HKPC19, Kas12, LSV^{+}18]. Nielson [Ano13e]. NLIZE [LLL^{+}19b]. Nmap [DSF^{+}14]. Node [HBF08, JRHT14, NBW14, OJK19, RP12, SSKB14]. Node-Link [JRHT14, NBW14, OJK19, SSKB14]. Node-Link-Group [SSKB14]. Nodes [WRF^{+}11]. NodeTrix [HFM07, YSD^{+}17]. Nodular [WKB^{+}13]. Noise [AGY^{+}17, CYC^{+}12, CCS12, FHHJ08, JEH02, KKSS13, LLD11, LLW06, YPI13]. Noise-Based [KKSS13]. Noise-Resistant [LLW06]. Nomograms [MMB^{+}19]. Non [AERA14, GAMD10, GXW^{+}18a, GXW^{+}18b, HUPS14, IAS19, JBS^{+}18, KW05, LYLG19, MWCE09, Mao96, MMH^{+}13, MYM16, MTB17, NWI17, NHPN14, OSS^{+}17, RJG17, SST^{+}17, SLS^{+}17, TvET14, YW16, YL18, ZBO13, FGS19, JQD^{+}08]. Non- [JBS^{+}18]. Non-Constant [HUPS14]. Non-Contact [IAS19]. Non-Euclidean [KW05]. Non-Experts [TvET14]. Non-Linear [GAMD10, MYM16, NHPN14, SST^{+}17, FGS19]. Non-Mobile [RJG17]. Non-Obtuse [YW16]. Non-Parametric [MWCE09, JQD^{+}08]. Non-Planar [MTB17]. Non-Premixed [YL18]. Non-Rigid [GXW^{+}18a, GXW^{+}18b, LYLG19, NWI17, SLS^{+}17]. Non-Skinned [ZBO13]. Non-Spatial [MMH^{+}13, OSS^{+}17]. Non-Visual [AERA14]. Nonlinear [AB01, BN12, Del08, HSK17, KBI^{+}18, KKW^{+}17, RW18, SVAC12, SKMR98]. Nonmanifold [BHS12, HG01].

Nonnegative [CLRP13]. Nonparametric [ASE16]. Nonphotorealistic [HCS^{+}07, RE01b]. Nonplanar [IYS13]. Nonregular [WPZ^{+}11]. Nonrigid [BBK07, HQ12, TCL^{+}13]. Nonsmooth [WB05]. Nonuniform [PRA^{+}10]. Nonuniformity [MS04]. Noodles [SZD^{+}10]. Normal [AWHS16, GQGP17, IWR^{+}18, JWC05, SJM14, WYP^{+}15, WTP^{+}19, WHX^{+}19, YRP18, YPI13, ZCL^{+}19, ZFAT11]. Normalization [BHSH15]. Normally [SKS12]. Not-so-Staggering [CDF14]. Note [Ano05a, Ano05b, De 15b, De 15a, De 15c, De 16b, De 17b, De 18c, Ebe03a, Ebe03b, Ebe04a, Ebe04b, Ebe06a, Ebe06b, Ert07b, Ert07c, Ert08, Ert09a, Ert09b, Ert10a, Ert10b, Ert11b, Flo16, De 17a, Hag99, Hag00, HE02, IKT15, Lin11a, Lin11b, Lin12b, Lin12c, Lin12d, Lin13a, Lin13c, Lin13d, Lin14c, Lin14a, Lin14b, Lin16a, Lin13b]. Notice [SGQ16]. Noticeability [KOK^{+}19]. Noticeable [NWHWD16]. Notifications [GWP^{+}18]. NotifiVR [GWP^{+}18]. Novel [ARRC11, AS98, CDZ^{+}09, CVC^{+}12, DSF^{+}14, INCB18, JAAL18, JSV^{+}08, KHA12, KK19, LHC10, NZ06, OHH06, PM08, RGK^{+}13, TCM10, WSM^{+}09, ZBZ^{+}13]. Novice [LKH^{+}16]. Novices [GTS10, GTS11, KPR^{+}14, LBW19, YEB18]. NPR [LCC^{+}17]. NPU [PMD^{+}07]. NPU-Based [PMD^{+}07]. Nuclear [DCK^{+}12]. Number [BDJ14, KWDG11, OKSK16]. Numerical [GBP^{+}13, SZD^{+}10, ZMZM15]. NURBS [SF14, KKM^{+}09, KML96, QT96, SF19]. NURBS-Based [SF14].

O [QK04]. O-Buffer [QK04]. Oasis [SGJM18]. Obesity [JAM^{+}14]. Object [DZL^{+}14, FG99, FST^{+}14, GHE19, KPBL16, KYT^{+}18, LLY^{+}13, LvL12, LSH07, LXT18, ORC07, PSG04, PLW11, QCH^{+}14, SJL^{+}18a, SPP^{+}14, Sil95, TWC^{+}18, TWMZ19, ZK17, ZGW^{+}19]. Object-Preserving [LLY^{+}13]. Object-Space [PSG04]. Objective [GO15, GT19, LLG17]. Objects [DGW11, DDKA06, EGS03, GT19, KTWW13, KDM^{+}16, LAM10, LKC09b, LZW^{+}13, NDS10, PK08, RNK^{+}15, SSI99, SPM^{+}13, SH12, SK99, SLMA06, WB05, YHJ^{+}17, YS03, YFM01, ZTP05, ZCJH12, ZWC^{+}16, ZX18, ZT99, ZCFL15]. OBL [Ros11]. Obliq [NB95]. Obliq- [NB95]. Oblivious [GSCO07]. obscuring [LSC08]. Observable [MOC^{+}14]. Observation [KCPE16]. Observation-Level [KCPE16]. Observational [YXG^{+}10]. Observer [HMTR19]. Observers [KNKH19]. obstruction [THT19]. obstruction-free [THT19]. Obtuse [YW16]. Occluded [MSHC99, QPNK18]. Occlusion [CM09, ET08, HDBC15, HLY10, IHS17, LGV^{+}16, PMP10, RWF19, SMG^{+}13, SVGR16, SG99, TYSN06, TCM06, TLS17, WWYP19, WP18]. Occlusion-Capable [RWF19]. Occlusion-Free [TYSN06, LGV^{+}16]. Occlusions [CRPH10, JNC^{+}15]. Occurrence [LCZ^{+}19, WXZ^{+}16]. Occurs [KHSS14]. Ocean [HMZ^{+}14, KSDD14, WPS^{+}16]. Octree [LCDP13, SJ06, WC09, WC10, WSC^{+}95, YS03, WSC^{+}95]. Octree-R [WSC^{+}95]. Octrees [LK11, ZGHG11]. Odor [MYI13]. Off [ALBR16, JK16, HKB^{+}19, PVF13]. Officers [GS16]. OLAP [LBGV13]. Old [OJK19]. Older [JBS^{+}18]. Olfactation [PBE19]. Olfactory [DDBR^{+}19, MYI13]. Omnidirectional [KNR17]. Omnistereo [TT05]. On-Demand [VHBS16]. On-Site [RNK^{+}15]. On-the-Fly [ZD18]. Oncological [SLK^{+}17a]. Online [CSL^{+}16, CCL^{+}16, EASB19, FT08, GCL^{+}18, KCH11, KKL^{+}16, LG12, LYW^{+}16]. Only [CPW^{+}18]. OnSet [SMDS14]. onto [FST^{+}14, KH19, NWI17, SSI99]. Ontology [SKKC19]. Opacity [MDM10]. Opaque [MBT^{+}18]. Open [BPB14, BSG18, CCL^{+}16, PAB^{+}08, SK13, WWB^{+}13]. Open-Boundary [BPB14]. Open-Box [SK13]. Opening [MPG^{+}14]. OpenMSI [RB18]. Operations [DL03, FGF^{+}05, KKM^{+}09, KKCS98, ME09, SKL^{+}14, Wan11, WM13b]. Operative [BSKR19]. Operator [GWK12, LRP97, PMCS11]. Operators [DZMQ16, OR98, OR99]. Opinion [DR08, WLY^{+}14]. OpinionFlow [WLY^{+}14]. OpinionSeer [WWL^{+}10]. Opportunistic [HF10]. Opportunities [JHKH13, KGD^{+}19]. Optical [BKA^{+}11, CWC^{+}06, DFG^{+}14, FM04, GIMS18, HIH^{+}18, HRISI15, IK15, IDAK15, IAIK16, IHS17, KSG^{+}16, LCR16, LH16, LHH16, LHC10, Max95a, MCS^{+}08, MIO^{+}15, PIN^{+}15, QPNK18, RWF19, SJK^{+}07, YLX^{+}12]. Optically [PK08]. Optics [RWF19, TT05]. Optimal [AEM09, Ano96b, BHW06, BH07, CGL^{+}17, DT10, DNP07, HKC^{+}12, HDJ05, LT16, LBG^{+}08, LSJ96, Max95b, MLMP18, PS12, RK17, SPP^{+}14, YKL^{+}08, ZSG^{+}13]. Optimality [AHRG10]. Optimisation [GCL^{+}18, LDM^{+}18]. Optimise [DSC^{+}16]. Optimization [AdLH13, CWM^{+}09a, CWQ^{+}07, DFG^{+}14, DMC^{+}12, EASD^{+}19, EDF11, FWD^{+}17, GSS^{+}15, HHQH17, HBW06, HLB^{+}18, KZW12, LZD13, LSS13, LL19, LGY19, LLL^{+}19a, LSPW12, MNS18, MPOW17, NXSL13, NZS^{+}17, PW12, SLM18, SJH^{+}07, SRMOW11, WL08, WG16, WCQ^{+}09, WDW16, WAWS18, YRP18, YCZ^{+}16, ZDZ18, vLFR17]. Optimize [CG16, CXR18]. Optimized [AZD17, GSPJ08, SJB10, WLHL13, WWY14, YC14]. Optimizers [TNT17]. Optimizing [TM12, VC17, WCG^{+}19, YYSZ06, ZWBH13]. Orbit [CML^{+}07]. Orchestra [PFC18]. Order [BYA15, BDJ14, BBG^{+}09, BSL^{+}12, GRT17, GUO00, HLL97, HVSW11, LVRL06, LFA^{+}16, MCK12, MNKW07, NKH11, NLKH12, NKH14, SCT06, SBM^{+}06, SK10, SW13, WXKP14, Zha14, ZG06]. Order-Independent [Zha14]. Ordered [MB18b, WD08, WBDS11, Ros11]. Ordering [CMSW04, PFC18, ZWM13]. Organic [OM09, YHJ^{+}17]. Organization [AN13, vHR08]. Organizational [NJJ11, RESC16]. Organizations [Ano11g]. Organized [FWR00, NSW^{+}17]. Organizing [NB12, SKB^{+}18]. Organs [HHCL01, Sel15]. Orientation [AGN^{+}19, JHBK19, LS02, RSBB17, REB^{+}16, ZQS11]. Orientation-Enhanced [REB^{+}16]. Orientation-Preserving [ZQS11]. Orientations [Nie04]. Oriented [AMM^{+}08, CS08, Kei00, RLA^{+}13]. origamic [LLLN^{+}14]. Origamizing [Tac10]. Origin [AAFW17, GZ14, YDJ^{+}19, ZMT^{+}19]. Origin-Destination [AAFW17, GZ14, YDJ^{+}19, ZMT^{+}19]. Orko [SS18]. Orthogonal [BW08a, BKH^{+}11, KDMW16, YYY16]. Orthographic [LT13]. Orthopedic [LRF^{+}11]. Orthopedics [DGBW09]. OSPRay [WJA^{+}17]. Other [RCL^{+}15, ZOC^{+}13, JHGH08, LSC08]. Our [BSSB10, EDK10, RCL^{+}15]. Out-of-Core [AGL06, HKC^{+}12, LP02, LCDP13, USM97, YSGM05]. Outcome [KNKH19]. Outcomes [WG12]. Outdoor [APV^{+}15, KGAM18, LBD13, LG12, SRK^{+}11, VGKS12, WLT^{+}18b]. Outdoors [BBBM18]. Outflow [WG12]. Outlier [NH06]. Outlier-Preserving [NH06]. Outliers [Wil18]. Outline [GUFM15]. Outlines [vGMSW15]. Output [DN12, FW08, HPC^{+}13, KSDD14]. Output-Coherent [HPC^{+}13]. Output-Sensitive [DN12, FW08]. Over-Plotting [DWA10]. Overcoming [MG13]. Overdraw [MG13]. Overestimation [HUPS14]. Overlaid [BCS11]. Overlapped [GPC^{+}17]. Overlapping [AAMH13, SLF^{+}12, VRW13]. Overlay [QWC^{+}09, SFMB12]. Overlays [KA12, MOC^{+}14, SA19]. Overview [ADG11, BISM14, CMP09, Chr03, EF10, JF16, LMK07, LBS^{+}19, NM13, TC09b, VGKS12, vdEvW14, DMS^{+}08]. Ovis [HMZ^{+}14]. Ownership [SSS13, WGR^{+}18]. Oxides [GBM^{+}12].

Pace [FPH19, vdCvW14]. Pacific [BHTY15, CCH14, DFQ12, HKQ13, LST^{+}16, NSvW11, HVY16]. PacificVis [BKL18, DW17, MSW19]. Package [IG19]. Packet [BWW^{+}12]. Packet-Ray [BWW^{+}12]. Packing [IYIK04, KW06, KXW^{+}18, YLG^{+}14]. Page [Ano12p, OAH14, Ano11i, Ano14j]. Pages [Ano08d]. Paint [Ger17]. Painted [YCLL08]. Painter [Ger17]. Painterly [CL06, OH12]. Painting [Ano14h, Bro07, DKMI13, DiV15, FYY^{+}19, Ger17, KBD^{+}11, KLYE13, KWL14, KISE14, LCC^{+}17, WWYS04, YLK12]. Paintings [TDM^{+}18]. Pair [ZCL^{+}19]. Paired [GHL15, HKF16]. Pairs [YHH^{+}19]. Pairwise [BMLC19, LBW19, WAG06, YML^{+}17, ZTZX13]. Palettes [PFC18, GLS17]. Palpation [UK12]. Panning [MOC^{+}14, RN19]. Panoramas [BCR19, PST^{+}15]. Panoramic [RPAC17]. PanoramicData [ZZD14]. Pants [LGQ09]. Paper [Ano12j, Ano13p, Ano14o, Ano16p, Ano16q, Ano16u, Ano17m, Ano18i, Ano19h, Ano19j, Ano19l, CKSB14, IIS^{+}17, MFS^{+}09, PDRK19, ZCL09, LLLN^{+}14, Ano13o, CFK12, CLS13b, DLM^{+}12, vHMM^{+}11]. Paper-Based [PDRK19]. Paper-Reference [ZCL09]. PaperCraft3D [PDRK19]. Papers [Ano16r, Ano17n, Ano18f, Ano18g, CR08, EDF08, WBDS11]. Paradigm [DVCD07, RLM10, SKBE17]. Paradox [AW14]. ParaGlide [BSM^{+}13]. Parallax [BCR19, BC18a, LXRY18, MSM^{+}11, PKS^{+}08, SKC^{+}19]. Parallax-Free [MSM^{+}11]. Parallax360 [LXRY18]. Parallel [AD12, BGM^{+}07, BGM^{+}08, BSO^{+}12, BBP08, BVB^{+}11, CL18, CW11, DK10, DK11a, EMP09, ED06, FPH^{+}08, GH00, Gor02, GXY12, GBP19, HKC^{+}12, HW09, HSH10, IBJ^{+}14, JFTW07, JF16, JCWD14, KCS^{+}16, KBH06, Lac96, LLB^{+}12, LT11, LDC96, LTP^{+}05, MB18b, MAWM11, MGM09, MFS^{+}09, NR18, NC07, NLS11, NH06, PZ12, REB^{+}16, RLS^{+}19, SMN12, SKLU^{+}11, VP09, VMCJ10, VHBS16, WZC^{+}15, WLSL17, YXSH13, YXM^{+}15, YGX^{+}09, ZWZ^{+}13, ZGH^{+}18, ZJX^{+}15, ZGHG11, ZW18, tCMR07]. Parallel-Hardware [JFTW07]. Parallelism [CGC^{+}11, GPC^{+}17, HBC12, SSIF09]. Parallelized [MS08]. Parameter [AAM^{+}12, BSM^{+}13, BVPtHR09, BM10, MHG10, OKB^{+}19, PBCR11, SHB^{+}14, TWSM^{+}11, WLSL17, NSW^{+}17]. Parameterization [BF01, GLB^{+}06, HAT^{+}00, KZW12, MGL07, NSZ^{+}17, NPPZ12, PTC10, WPZ^{+}11, YYSZ06, YKL^{+}08, ZHGH11]. Parameterizations [HFL18, SAS05]. Parameterized [VABW09]. Parameters [BTJ^{+}13, DGAB16, EASS^{+}18, FRG^{+}19, JBS^{+}18, KBD^{+}11, LSH07, TKTN09, YL16]. Parametric [ADDG12, CTT^{+}16, DQ07, Elb98, FGF^{+}05, IDAK15, MWCE09, JQD^{+}08]. Parametrization [AG17]. Parcels [VABW09]. Pareto [HHG14, MLMP18]. Pareto-Optimal [MLMP18]. Pargnostics [DK10]. Parity [JCWD14]. Part [Hu16b, MBH^{+}12, VB13]. Part-Level [Hu16b]. Partial [HQ12, TWSS16, Wan11]. Partially [KLC08]. Participating [WH18, ZC03]. Participation [Ano08c, Ano12e]. Participatory [VWF09]. Particle [BGB15, BC12, COJ15, CMK15, EGS03, GIK^{+}07, GKM^{+}15, GHS^{+}19, HL02, HWM95, Har16, IWR^{+}18, KL96, KKKW05, LSY^{+}18, LLRR08, LTKF08, LSC08, MSW^{+}08, MNKW07, MWK^{+}08, NJB07, RGC^{+}14, SYM14, SM17, SFBP09, STM08, WSTH07, YEII16, ZGH^{+}18, ZD18, vPVvdW10]. Particle-Based [COJ15, GIK^{+}07, GKM^{+}15, NJB07, LSC08, MWK^{+}08]. Particle-Particle [LSY^{+}18]. Particle/Flow [STM08]. Particle/Volume [SYM14, SM17]. Particles [ATT12, GT14, KSSW09, MKW07, ySKK07, YS03]. Partition [CGL^{+}17, LXR19, MP13, ZK14a]. Partition-Based [MP13]. Partitioned [FWT^{+}04]. Partitioning [AT16, BSM^{+}13, MW99, SN97, TDR10, WQZ^{+}18]. Partners [JOR^{+}19]. Parts [AAFG18, XYS^{+}16]. Partwise [WPZ^{+}11]. Pass [ASDW14, BTHD11, MPT03, WX13]. Passengers [Chi16]. Passing [BRNB19]. Passive [HSR18, NMN^{+}18, STH13, ZK17]. Past [LMW^{+}17, NJ16]. Patch [GGZ^{+}18, LSS^{+}15, WWC^{+}14, WHX^{+}19, XLND11, ZWS^{+}17]. Patch-Based [GGZ^{+}18, LSS^{+}15, XLND11, ZWS^{+}17]. Patches [Gor02, HTF97]. Patent [KBGE11]. Patents [FHKM17]. Path [AMA11, BMWW18, HWM95, HHCL01, HL09, KM16, SAC^{+}08, TWHS05, WBK^{+}08, WG16, ZLB^{+}05]. Path-Preserving [AMA11, HL09]. Pathline [COJ15]. Pathlines [MWSJ14]. Paths [AMA11, EDK10, LLBS17, LBH11, LWD^{+}17, WOO17, WKW06]. Pathways [JYC^{+}10, LPK^{+}13]. Patient [BSKR19, HHO^{+}17, KGPS13]. Patient-Centered [HHO^{+}17]. Patient-Specific [KGPS13]. Patients [LPK^{+}16]. Pattern [HEWK03, IFM14, LLMB19, LXR19, VN19, WSW16, YNM15]. Pattern-Based [YNM15]. Pattern-Growth [VN19]. Patterns [AAFW17, ASW12, BSH^{+}16, BW08a, BEJK12, CYW^{+}16, FPH19, GHA^{+}08, GZ11, GCML06, HA06b, HHKE16, JER16, KDA^{+}09, LCZ^{+}19, LWD^{+}17, MOJB^{+}19, PW13, SMO^{+}13, SCT^{+}10, TTS10, WWY14]. PC [Ano15j, GUO00, KGP^{+}13, KGG^{+}12]. PC-MRI [KGP^{+}13, KGG^{+}12]. PCs [NHN07, TBR^{+}12]. PDE [BF01]. PDEs [DQ07, RL08]. PDF [SKMH14]. Peacocks [AMA06]. Peak [KHW^{+}09]. Peaks [CCL^{+}16]. PeakVizor [CCL^{+}16]. Pearl [NDM^{+}97]. Pearl-Quality [NDM^{+}97]. Pedestrian [KO12, KGAM18]. Pedigree [TNS10]. PedVis [TNS10]. Peer [BGC^{+}11, BFE15]. Peer-to-Peer [BFE15]. PELs [XHT^{+}07]. Pelvic [SLK^{+}17a]. PelVis [SLK^{+}17a]. Pen [WLJ^{+}12, ZZD14]. Penalized [BKS01]. Penalized-Distance [BKS01]. Penalty [Dru08, XZB14]. Penalty-Based [XZB14]. Pendulum [TLC^{+}10]. Penetration [KLM04]. Penumbral [DF96]. People [HHH16, LKH^{+}16, OBKP18, RCL^{+}15, TAL^{+}07]. Per-Pixel [MDM10]. Perceived [BEDF16, GMS^{+}07, JSB13, KHSS14, RBK^{+}19]. Perception [BH07, BCS11, BI12, BG07, BSWL12, CWM^{+}09a, EMdSP^{+}15, FIB^{+}14, GLM^{+}17, GCNF13, GTPB19, HTP19, HKR^{+}08, HOT98, HV13, JAO^{+}14, JME10, JHBK19, KSTE06, LH16, LHH16, LML^{+}18, LBLE^{+}14, MRT00, PW12, PLE^{+}18, PBK^{+}12, PIS15, RYKL13, RFSP19, TSA14, TvET14, TYL^{+}18, UBH19, VS11, WK13, WFC^{+}18, WCG^{+}19, YBW^{+}19, ZK17, ZNZX16, ZHF^{+}19]. Perception-Based [CWM^{+}09a, EMdSP^{+}15, MRT00]. Perception-Driven [WFC^{+}18, ZNZX16, YBW^{+}19]. Perception-True [HV13]. Perceptions [RGFLL14, RCL^{+}15]. Perceptual [AL06, ACS^{+}18, BSG18, BAW16, DBH14, HBW06, KOCC14, KHA10, LKR^{+}18, MPOW17, PGRS13, ŠTPV12, SBE^{+}15, SBW17, WB08, YN06, ZK10, vHR08]. Perceptual-Statistics [ŠTPV12]. Perceptually [BTV14, GH00, LSS13, NSN14, QM08, ZWM13]. Perceptually-Based [ZWM13]. Perceptually-Driven [GH00]. Performance [ADDG12, BAP^{+}17, BBG^{+}18, BBB^{+}19, BC18a, BTB10, CTGH13, DK11b, GMS19, GGJ^{+}18, HBT14, KGAM18, KLL12, KOK^{+}19, LLB^{+}12, MZC^{+}16, PBO^{+}14, RAL^{+}17, SSMG13, TKAM06, WXY17, XXM19a, XMRC17, vdCvW14, GSL14]. Performing [KKM^{+}09]. Perfusion [ODH^{+}07, POM^{+}09, TBB^{+}08]. Peridynamics [HWW18]. Periodic [CML^{+}07]. Peripheral [JSB13]. Periphery [LBHW18]. Persistence [BEK10, FFST19, RM17, RKG^{+}11, RML12, RFLL18]. Persistence-Based [RML12]. Persistent [CLMO17, SJ06, WC09, WC10]. Person [TMM^{+}13]. Personal [FDPH17, HTA^{+}15, JhR10, KPV^{+}18, LLL^{+}19a, TBHC16, VBC^{+}16, WBA^{+}14, ZGW^{+}14]. Personality [Per95, SPW07, ZKM18, ZOC^{+}13, UKW19]. Personalization [WGR^{+}18]. Personalized [CTM^{+}13, YWV^{+}19]. Personified [TMM^{+}13]. Perspective [BHST17, DHM13a, HCS^{+}07, LS10, SH00a, SBK^{+}11, WB08, WCA^{+}17]. Perspectively [HN13]. Persuading [KV08]. Persuasive [PMN^{+}14]. Perturbation [CA00, LLL^{+}19b]. Perturbation-Driven [LLL^{+}19b]. Perturbations [CRB^{+}05]. Pervasive [GLZR17]. PET [RHR^{+}09]. PET/CT [RHR^{+}09]. Petascale [HBJP12]. PETMiner [HEFR18]. Petrophysical [HEFR18]. Pets [JAM^{+}14]. Phase [GPR^{+}01, GLX17, KBE^{+}18, PSN10, YYD^{+}19, ZC06]. PhenoBlocks [GHC^{+}16]. PhenoLines [GNDV^{+}18]. Phenomena [CMK15, SBHW11, WLMK04]. Phenomenological [MM17]. PhenoStacks [GGC^{+}17]. Phenotype [GHC^{+}16, GGC^{+}17, GNDV^{+}18]. Phone [VARS14]. Phones [BLIC19, OKI15, PKMR15, WRM^{+}10].

Photic [XHT^{+}07]. Photo [CORLS96, LWZ^{+}18, LZJ^{+}18, YHJ^{+}17, Yon14, YLG^{+}14, ZGW^{+}14]. Photo-Realistic [CORLS96]. Photograph [ZGW^{+}14]. Photographic [ASG15, EMRY02]. Photographing [HWZ^{+}19]. Photographs [CBLD11, CH03, HCS^{+}07, JDA^{+}11, XZX^{+}17]. Photography [CMF^{+}18, LDN11, MBW^{+}07, TKTN09]. Photometric [PWG17, WLDW11]. Photon [HUPS14, JKRY12, JY17, SJ09, SCL08, ZDM13]. Photon-Mapping [SCL08]. Photons [LBS^{+}16]. Photorealistic [PY09, RBDG15, ZLG^{+}06]. Photoreceptor [RD05]. PhotoRecomposer [LWZ^{+}18]. Photos [CdOKRV09, KCA16]. Phrase [vHWV09]. Phrasing [OPH^{+}16]. Physical [AN13, BGK11, DBP14, HSR18, JH16, LB17, MZS^{+}19, MLS18, Qin09, RNE^{+}17, STS^{+}14, TJW^{+}17, VBK17, YQK^{+}17, ZPS04]. Physicalization [LPF^{+}19]. Physically [BPS^{+}11, CMN13, GHK97, IZM18, JWL05, LXB17, Vis15, WB08]. Physically-Based [BPS^{+}11, CMN13, GHK97, IZM18, WB08]. Physician [OIR^{+}17]. Physics [HEG^{+}17, KJ12, LXW^{+}18, QT96, SAM^{+}07, TLC^{+}10, YGV^{+}13]. Physics-Based [HEG^{+}17, KJ12, LXW^{+}18, QT96, SAM^{+}07, TLC^{+}10, YGV^{+}13]. Physiologically [MOF09, MOF10]. Physiologically-Based [MOF10, MOF09]. Physiology [RD05, XSZ^{+}17]. Pick [WVFH12]. Pictorials [Chi16]. Picture [PTM^{+}18, WLL^{+}16]. Pictures [ZLZY18]. Pictus [IHD^{+}18]. Pie [GPL^{+}13]. PieceStack [WWS^{+}16]. Piecewise [HTF97, LVRL06, SM11, SZ12, XJF^{+}08, ZWZD15]. Piercings [SZHR11]. Pigment [TDLG19]. Pigment-Based [TDLG19]. Pigmento [TDLG19]. Pinhole [PRA^{+}10]. Pipeline [BGM^{+}07, BGM^{+}08, CQC^{+}08, GW06, HQK06, KRW19, LCP^{+}13, LZH^{+}13, LWP^{+}06, MiS^{+}18, SvW98, vLFR17]. Pipelines [DWBR06, KZX^{+}14, Koo08, Mor13, SRM19]. Pitch [SJL^{+}18b]. Pituitary [NWF^{+}05]. Pivoting [BMR^{+}99]. PivotPaths [DRRD12]. PIWI [YLZ^{+}13]. Pixel [FH16, Kei00, KHD02, MDM10, PSKN06, RGE19, RMCW19, WHL16, BPC^{+}10]. Pixel-Accurate [FH16]. Pixel-Based [RGE19, BPC^{+}10]. Pixel-Oriented [Kei00]. Pixels [KIS17, XDN11]. PixMix [HB14]. PizzaText [YFZ^{+}18]. Place [AAM^{+}12, Nie95, NSN14, PBA10]. Placegram [JhR10]. Placement [BMW17, CCK07, GWFI14, LAM10, LMG06, OJ15, PSKN06, PM08, RK17, WBK^{+}08, WLZM10, ZWZ^{+}13]. Placements [HK16]. Placenta [MMK^{+}17]. Placental [MMK^{+}17]. Places [AAH^{+}13, CDW^{+}16]. Planar [BJM07, KBH^{+}10, KBH13, LLW15, LYY^{+}16b, MK16, MTB17, MTB18, WGC^{+}08, WS01, ZWC^{+}16, ZCFL15]. Planar-Faced [ZCFL15]. Plane [DHL09, HSCS11, KH19, OJ15, SS13a, SBW17, SYR11]. Planetary [BAB^{+}18, HDJ05, KLJ^{+}09, MKHD05]. Planetary-Scale [KLJ^{+}09]. Planning [BHWB07, DGBW09, DBW11, EHH^{+}19, HK16, INCB18, KKMS11, KSI^{+}96, LAM10, LRF^{+}11, MKN^{+}07, MTRP10, OSS^{+}17, SLK^{+}17a, SAC^{+}08, WHK15, JSV^{+}08]. Plans [LAP19, SST^{+}17]. Plant [LZH^{+}13]. Plasma [RGC^{+}14]. Plasma-Based [RGC^{+}14]. Platforms [BTJ^{+}13]. Plausibility [ASMP17, BAP^{+}17, SNB^{+}17, HWA15]. Play [BBC15, KBS13, ZKM18]. PlenoPatch [ZWS^{+}17]. Plenoptic [ZWS^{+}17]. Plot [FHSW13, IML13, WHZ^{+}18]. Plots [ASW12, ED06, FKLT10, FBW16, HV13, HTE11, JFSK16, JM10, KNKH19, LT10, MMB^{+}19, MG13, REB^{+}16, RGE19, RW18, RSS14, SFP^{+}19, WLSL17, WH11]. Plotting [DWA10, TFO09]. Podium [WDC^{+}18].

Poem [MLCM16]. Poemage [MLCM16]. Point [ABCO^{+}03, APS^{+}14, BHS12, CK10, FFST19, FM12a, GR04, Gué01, HLM10, HPJG08, LI06, Lin14f, LDX10, LPG12, LZH^{+}13, LDW^{+}15, LWC^{+}18, Mal05, MSE^{+}06, MOG11, OHJ^{+}11, OHWS13, PSG04, PSKN06, PM08, RGFLL14, RL08, SHS11b, SRW^{+}16, TCL^{+}13, WLSW08, WSS09, WH18, WAG06, WG16, YNM15, YEII12, ZK06, ZG06, ZCW19, LSC08]. Point-Based [GR04, HLM10, LPG12, SHS11b, WSS09, WH18, ZG06, RL08, ZK06]. Point-Cloud [APS^{+}14]. Point-Cloud-Based [LDX10]. point-feature [LSC08]. Point-to-Polygonal-Mesh [Gué01]. Point-Wise [ZCW19]. Pointing [FKS16]. Points [BF01, CTT^{+}16, HGWW18, KV03, LBH18, RKG^{+}11, SXX^{+}19, Ste98, TSW^{+}07, YGX^{+}09, ZG12, ZW18, vdEHBV16]. Poisson [DRW16, LLLF08, LH13, YXSH13]. Poisson-Disk [LLLF08]. Polaris [STH02]. Polarity [GUFM15]. Policing [MMT^{+}14]. Policy [BKW03]. Polluted [MPBM^{+}18]. Pollution [QCX^{+}07]. Polycube [GXH^{+}13, WLL^{+}12]. Polycubes [LLWQ13]. Polygon [LRN96, PLK12, QM08, SvW98]. Polygonal [AGL06, ESV98, Gué01, NT03, PML97, XESV97, YT02]. Polygonizer [YOS13]. Polygons [DKMI13, ER97, GTLH01]. Polyhedral [LH03, LL05, MHDG11, PCG15, SPB96, Tac10, Wan11]. Polyhedrons [WLSW08]. Polymorphic [NK06]. Polynomial [EG09, HB03]. Polynomials [HWM95]. Polyp [HQK06, ZBB^{+}06]. Polytopes [CS08, KLM04]. Pondering [VI18]. pop [LLLN^{+}14]. pop-ups [LLLN^{+}14]. Population [KLG^{+}16]. Populations [CLB13, FLF^{+}11]. Popup [SFP^{+}19]. Popup-Plots [SFP^{+}19]. Pore [UMW^{+}12]. Porosity [MDG00]. Porous [PC13, WAWS18, ZFS^{+}19]. Portable [KCS^{+}16]. Portals [Ano05d]. Portfolio [BNTM16]. Portfolios [MEV^{+}14]. Portraiture [Bro07]. Pose [BWS^{+}19, CZN^{+}11, FCSF17, GSCO07, GCL^{+}18, HBESB11, HV00, IFM14, LIM^{+}12, MUS16, YHH^{+}19]. Pose-Inspired [FCSF17]. Pose-Oblivious [GSCO07]. Pose-Pairs [YHH^{+}19]. PoseShop [CTM^{+}13]. Position [JNC^{+}15, LYLG19, LPG15, RP12, WLD^{+}19]. Position-Based [WLD^{+}19]. Position-Dependent [LPG15]. Positional [PH11]. Positioning [BMST97, KIS17, TLH10]. Positions [AHRG10, VP09]. Possibilities [LPCC17]. Possible [LG13, LG15]. Post [BSKR19, Ano11c]. Post-Operative [BSKR19]. Postprocess [MDG00]. Posture [LZLS16, VSS08]. Postures [WHK15]. POT [SJ06]. Potential [BHHM19]. Potentials [JR07]. Powell [EM06]. Power [CBL07, HFMC12, LS13a, MGMP18, PMN^{+}14, WSM^{+}09]. Powered [JPD^{+}18]. PowerSet [AR17]. Practical [BMTD05, CVG13, EVM08, GBHP08, JKJTM06, MBW^{+}07]. Practice [DPW^{+}15, IIC^{+}13]. Practices [AZL^{+}19, HFM16]. Practicing [CLB13]. Pragmatics [HSTD18]. Pre [Ano08d, GAMD10]. Pre-Integrated [GAMD10]. Pre-Pages [Ano08d]. Preattentive [KK19]. Precepts [AS05]. Precise [EPS^{+}15]. Precision [HBKS09, HKB^{+}19, MMB^{+}19, PNML08]. Precomputed [BES12, XJF^{+}08, ZDM13]. Predicate [zBBKN14]. Predicate-Based [zBBKN14]. Predicates [BPM^{+}13, KGP^{+}13, SS06a]. Predict [AERA14, UKW19]. Predictability [BS02]. Predicting [KDM^{+}16, TWBBM17]. Prediction [BSR^{+}14, CLL08, FM12a, HZL^{+}19, HKKS18, LYL19, LLG17, LBR^{+}17, MTB18, PFC18, VB13]. Prediction-Based [FM12a]. Predictions [KBL19, LPCRH19]. Predictive [FM06, GBP^{+}13, KPB14, MHR^{+}11, MMT^{+}14, VBV^{+}18]. Predictor [BS95, SFBP09]. Predictor-Corrector [BS95, SFBP09]. Predominance [RCSJ18]. Preface [ALS^{+}17, Ano17o, CDF^{+}19, DEF^{+}18, FvHM^{+}10, Hag98, KSTW18]. preferable [GLS17]. Preferences [LLL^{+}19a]. Prefiltered [Csé08, Cse10]. Prefiltering [BN12, RHZN11]. Prefix [CDM^{+}06]. Preintegrated [LS13a]. Premixed [BWP^{+}10, YL18]. Preoperative [BHWB07]. Preparation [WZW^{+}05]. Prepare [RKA^{+}13]. Prerendered [CYB08]. Prescribed [ONL^{+}12, PT17]. Prescription [XGS^{+}19]. Presence [BC18b, SFL^{+}16, WGR^{+}18]. Present [NJ16]. Presentation [BGCS17, GJZ^{+}12, KK19, MHS07, TJW^{+}17, WBK^{+}08, vdEvW14, vdCvW14]. Presenting [HKF16, MYI13]. Preservation [APP11, CJTM05, DSF^{+}14, ZZG^{+}12, ZSG^{+}13]. Preserved [Wan08]. Preserving [ATT12, AMA11, ALMF19, BDD^{+}16, BGKG06, CWM09b, DK11a, FYP10, GNSP^{+}14, GLX17, HZH14, HYB^{+}17, HYZ^{+}12, HL09, IG19, KSW06, KOF08, KGZ^{+}12, LLL06, LLY^{+}13, LZJ^{+}18, LFR03, LDC16, LWC^{+}18, MSSH14, NGK18, NH06, RSB96, SFH06, SWS^{+}11, SRML07, TGS11, WWLM11, WXJD17, WCB^{+}18, WCC^{+}19, WJE01, YNBH11, ZQS11, CLB11, LLLN^{+}14, ML19]. Pressure [CMF12, DRW16, SB14]. Pretend [BBC15]. Preventing [GD01]. Prewriting [LFW^{+}19]. Priming [VZS18]. Primitive [WD09]. Primitives [BAW16, FGF^{+}05, Rot13]. Principle [CXDR19, VC17]. Principled [BRBF14, TWSM^{+}11]. Principles [BCB10, HSTD18]. Printable [SB17]. Printed [ZTZ17]. Printing [WQZ^{+}18]. Prior [SHM10]. Prioritized [KS00a, KS00b, KS01]. Prioritized-Layered [KS00a, KS00b, KS01]. Prism [ANR^{+}18].

Privacy [DK11a, WCC^{+}19].

Privacy-Preserving [DK11a]. PROACT [HHO^{+}17]. Proactive [MMT^{+}14]. Probabilistic [AHH^{+}14, AJ19, BMA^{+}19, CLG16, GR04, LLPY07, PH11, SNG^{+}17, SRKL19, YYR17]. Probabilities [LY12]. Probes [BDW^{+}08]. Probing [DVC18, SDMT16, vPBB^{+}11]. Problem [Mar18, QCX^{+}07, RM17]. Problem-Driven [Mar18]. Problems [ZYM^{+}14]. Procedural [DKMI13, HQQ12, HKYM17, MM11, VT08, VBV^{+}18, YHL^{+}17]. Procedurally [SGJM18]. Procedure [AGN^{+}19, DKM06b]. Procedures [TGW^{+}95]. Proceedings [Ano11i]. Process [CLS^{+}12, DC17, KS14b, LZLS16, MGM09, Mun09, NXW^{+}16, SASS16, TWBBM17]. Processes [CG16, GdBG12, JBH^{+}09, KSDD14, LSC^{+}18, LFA^{+}16, TKBH17, WYM12]. Processing [ALMF19, BYA15, BGM^{+}07, BGM^{+}08, BVW^{+}07, CCQ^{+}14, HKC^{+}12, HQ13, RGF^{+}04, XLND11, vLFR17]. Processors [WJ08, ZK14a]. Produced [AFRS05]. Product [MZS^{+}19, WPC^{+}13, WL16, WH11]. Production [EGH^{+}06]. Products [JHH^{+}10]. Professional [AZL^{+}19]. Profile [NWHWD16]. Profiles [LB19, OHWS13]. Profiling [GMD^{+}17, JFS16]. Program [Ano12i, Ano13l, Ano13m, Ano14m, Ano14t, Ano15m, Ano16m, Ano16w, Ano17l, Ano18f, Ano18h, Ano19g, Ano19i, BSS18, BRS18, CGGR18, GGHZ19]. Programmable [SvLF10]. Programming [DKM06a, GMD^{+}17, NW11, WJR^{+}13]. Progress [SPG14]. Progression [GJG^{+}19]. Progressive [CBPS06, EASS^{+}18, FSME14, FE17, GGLQ19, HPJG08, KL03, LL05, LYY^{+}16b, LXR19, PR00b, PR00a, PLvdM^{+}17, PHV^{+}18, SPG14, TKBH17, VP04b, YS03, ZGC^{+}17, ZNZX16, ZCFL15]. Project [PMvWC05]. Projected [IYS13, IFM14, Lin16b, SCT06, SFC^{+}07]. Projecting [PSN10]. Projection [AHR^{+}11, AIS18, BE06, CMSW04, DMC^{+}12, EMdSP^{+}15, FST^{+}14, GHE19, GXY12, HIH^{+}18, HK99, IMS15, IAS19, JCC^{+}11, KH19, KIS17, KS00a, KS00b, KS01, KLS^{+}18, MS04, NWI17, NA19, OR98, OR99, PNML08, PWG17, PWIG18, RLM10, SLS^{+}17, Sim07, TIS16, TIK15, WKME03, YYY16, YRWG13, ZXM10]. Projection-Based [AHR^{+}11, FST^{+}14, HK99, MS04, Sim07, ZXM10]. Projections [Gle13, Jen12, LT16, LT18, RKK16, SDMT16, SBK^{+}11, XZM17, FGS19]. Projective [DHL09, HWW18, HJLH19, KLCK17, LL19, OBLN17, YON05]. Projector [BSM06, BJM07, DVCD07, IMS15, KIS17, LAM10, RNK^{+}15, SLGM09, WG16]. Projector-Camera [RNK^{+}15]. Projectors [SM09, SM11, YHLJ08]. Projects [AABH^{+}16]. Promoting [PFG08, RM15]. Prone [NMGK17, ZMG^{+}10]. Proofreading [AABH^{+}16, HKBR^{+}14]. Propagation [AM13, CSG^{+}19, CLT^{+}08, CDM^{+}06, IK95, MAKM14, MRG^{+}15, MYM16, NZS^{+}17, RNL09, RSM^{+}16, RSR^{+}18, STB18, SRK^{+}11, XNT11, XWL^{+}15]. Propel [PSBS12]. Properties [BLG^{+}16, CSM07, CCS12, HEFR18, ORC07, RYKL13, RZP12, SFA^{+}15]. Property [KISE14]. Proportion [LJZ12]. Proportional [RSFH14, Wil12]. Prostate [BSKR19, HHO^{+}17, LDM^{+}18]. Protection [ZTP05]. Protein [BLG^{+}16, JYC^{+}10, LVRH07]. Proteome [LLB^{+}06]. Proteomic [JYC^{+}10]. Prototype [LHC10]. Prototyping [GKM^{+}15, KMSB14, MGJH08, SE18, SZK15]. Protovis [BH09]. Provenance [BYB^{+}13, CCSK19, GS06, NXW^{+}16, RESC16, SKB^{+}18, SGP^{+}19, WSD^{+}13, WGS^{+}13]. Provenance-Based [SGP^{+}19]. Provide [LKT13]. Providing [EASB19, KLKS10]. Proxemics [JHKH13, LBHW18]. Proxies [AM13]. Proximal [BYA15]. Proximity [KLL^{+}13, ZCL08]. Proximity-Based [ZCL08]. Proxy [NHEM17, TCM10, ZK17]. Prune [CL18]. Pruning [LFP07]. Pseudo [JAO^{+}14, UBH19]. Pseudo-Haptic [JAO^{+}14, UBH19]. Pseudonormal [BA05]. Pseudophysical [CK05a]. Psychological [DGAB16]. Psychologically [MZH^{+}08]. Psychophysical [JH16, LLPP19, SNB^{+}17, WKSS05]. PTOT [WC10, WC09]. Public [DLA^{+}09, LLKN17, DSC^{+}16, TKE16, ZFA^{+}14, ZYM^{+}14]. Publication [LB19]. Publications [IHK^{+}17]. Pull [AGY^{+}17]. Pulse [MDL^{+}17]. Purpose [AZM12, GUO00, YNM15]. Purposes [RESC16]. Pursuit [HQC^{+}19, LvL12]. Push [AGY^{+}17]. Putting [ZCJH12]. Puzzle [PSM12]. PVsolve [VP09].

Q [WGSY19]. Q-Networks [WGSY19]. Quad [KEJK19, PPT^{+}11, ZHLB13]. Quad-Based [PPT^{+}11]. Quadrangulation [TDN^{+}12]. Quadratic [DKM06a, HTF97, LXW^{+}18, RZNS04]. Quadrilinear [HWM95]. Quadtree [CE01, GSG96]. Quadtrees [LFR03]. Qualitative [BBIF12, Bru17, DJK^{+}06, GLvP^{+}12]. Quality [Ano09b, ASMP17, BMR01, BHWB07, CF10, Cse13, CWRY06, DSC^{+}08, DSS^{+}09, EGG^{+}12, FWL17, FAW10, GMS^{+}07, GO15, GQM^{+}18, GLX^{+}18, HWHK16, HE06, HB14, HAM11, JJ09, LAM10, LYS^{+}10, LKL^{+}15, LJL04, LLG17, LCDP13, MD12, MKW07, MSHC99, NDM^{+}97, NHEM17, NW11, RZHB^{+}08, RPHI08, SSS06, SJK^{+}12, TTR10, Vas16, WGS07a, WM08, WSS09, YDG^{+}16, vLBB16, ZK06]. Quantification [AL06, AE13, FHG^{+}09, HLRS^{+}08, WPB^{+}11]. Quantified [BJB^{+}12]. Quantifying [FPH19, ZM17]. Quantitative [Bru17, DJK^{+}06, JYC^{+}10, LSPS10, MMDP10, War09, ZBG^{+}17, HSKIH07]. Quantities [JDSR^{+}18]. Quantization [CM02, RG95]. Quantum [JV09]. Quartet [BAP^{+}17]. Quartic [Kim13, RZHB^{+}08]. Quasi [AMA06, Cse10, ME11a, TC09a, Wan08, ZMG^{+}10]. Quasi-Conformal [ZMG^{+}10]. Quasi-Developable [TC09a, Wan08]. Quasi-Interpolation [Cse10]. Quasi-Trees [AMA06]. Quaternion [HM95, Nie04]. Queries [BŠG^{+}09, CDM^{+}04, HTC09, KLL^{+}13, KPS16, KMH11, LLMB19, SAM^{+}05]. Query [BAAK^{+}13, DR08, GABJ07, GABJ08, GGA^{+}11, PHE^{+}18, SW13, SCL08, STH02, SA19]. Query-Driven [GABJ07, GABJ08, GGA^{+}11]. Query-Guided [BAAK^{+}13]. Query2Question [NW15]. Querying [ADP02, SVK^{+}07]. Question [OJK19]. Questionnaire [YAE07]. Questions [SDW09]. Quick [YSGM05]. Quick-VDR [YSGM05]. Quilts [BW11].

R [WSC^{+}95]. RACBVHs [KMKY10]. Radar [ALBR16, HLNW11]. Radial [ALBR16, AAMH13, Bac07, BKH^{+}11, CVS^{+}19, DBB10, DLA^{+}09, DLR09, FS19, LHLW10, SG09, WKB^{+}13]. Radian [CDDS18]. Radiance [ACTM12, GBP07, KGPB05, SSI99, SHR^{+}11, XJF^{+}08, YFM01]. Radiative [SZN^{+}18, SKLU^{+}11]. Radio [SRK^{+}11]. Radiofrequency [RKSH11]. Radiograph [SCT06]. Radiometric [GB08b, LCR16, TIK15]. Radiosity [GH00, HKL17, MPT03, NSS14, Sbe97]. RadViz [RSRDS16, SGM08]. Rail [MGJH08]. Random [CKLL09, KMKY10, MS18b, PBL10, Sbe97, eYL07, ZZC11, ZWLC19]. Random-Accessible [KMKY10, eYL07]. Randomized [GSCI15]. Range [GK95, LRP97, LCNG14, NSN14, RD05, SLW^{+}10, VAB12, YFM01, YNCP06, ZBG^{+}17]. Ranges [BLIC19]. Rank [GGZ^{+}18, LLR18, MLKS18, SS06b, WHX^{+}19]. Rank-Aware [MLKS18]. Rank-by-Feature [SS06b]. Ranked [KLC08]. RankExplorer [SCL^{+}12]. Ranking [HYFC14, KH16, SCL^{+}12, WDC^{+}18, WCD^{+}19]. Rankings [GLG^{+}13]. Rapid [MGJH08, MM08, PMT^{+}19, YEB18]. Rapidly [KHSS14]. Rare [LGG^{+}18]. Raster [KHPS07]. Rasterization [LCDP13, SRK^{+}11]. Rate [KPR^{+}15, Lin14f, SSB^{+}17]. Ratio [FHSW13, TGH12, WWF^{+}19]. Rational [AG17, HTF97, JSG03, PCS^{+}12]. Rational-Quadratic [HTF97]. Rationalization [ZK14b]. Ratios [WWZ^{+}18]. Raw [FH06, LSZ^{+}18, WX17]. Ray [BWW^{+}12, BPL^{+}19, CRPH10, GYK^{+}16, HH10, HBAB14, HK99, KHW^{+}09, KKCS98, LGM^{+}08, LYS^{+}10, MYM16, MGW10, NKP^{+}15, NO97, NK06, PPL^{+}99, SMP11, SF14, SYK^{+}18, SM97, SN10, SK00, Stü98, TCM^{+}12, WFM^{+}05, WFKH07, WJA^{+}17, WJ08, WWW^{+}19, WLMP19, WSC^{+}95, WKI^{+}17, vAPP^{+}11, AHR^{+}11, ERL^{+}13]. Ray-Casting [BPL^{+}19, HK99, SMP11, vAPP^{+}11]. Ray-Scene [WLMP19]. Ray-Space [GYK^{+}16]. Ray-Tracing [NK06, SK00, Stü98]. Ray-Triangle [HH10]. Raycasting [MHDH07]. Rays [KKMS11]. Raytracing [APW16, GN12, HL09]. RBF [LL14, YLY^{+}12]. RBF-Based [YLY^{+}12]. RCLens [LGG^{+}18]. Reach [HSR18]. Reaching [CK16, SSE15]. Reactive [SRHH16]. Read [OSSK12]. Readability [DFG^{+}14, GUFM15, HBF08, OSSK12]. Reading [BLE19, KA12, WHP^{+}18]. Real [AL11, BTB^{+}04, BTB10, BTH^{+}13, CLS^{+}12, CDAF18, CMF^{+}18, CK05b, CVC^{+}12, ČB19, CORLS96, CMPC06, CDA99, DHL09, EHH^{+}19, FS19, GSCI15, GMD13, GWBO12, GB08b, HB14, HLRC^{+}12, HRISI15, HDJ05, IIS14, JKM06, JHBK19, KKSM19, KLL12, KMDZ10, LCR16, LAM10, LDSM17, LKC09a, LKC09b, LKR^{+}18, LKS13b, LDR00, LXT18, MB03, MLKS18, MT01, NQX^{+}05, OKI15, PSSC17, PD04, PMP10, PFW12, Per95, PKMR15, QYH^{+}18, RGF^{+}04, RJD^{+}07, RHJ^{+}16, RYL^{+}18, RKSH11, RHD^{+}06, RASS17, RBLW07, SSI99, SLM18, SOS^{+}17, SKP07, SN10, SVGR16, SGJM18, SKH^{+}19, SAC^{+}08, SFR^{+}10, SHR^{+}11, SSE15, TCM06, TLC^{+}10, THV^{+}14, VMT06, WRM^{+}10, WTW^{+}08, WFW^{+}17, WOO17, WSE07, WXY17, WY19a, XESV97, XYS^{+}16, XCZ^{+}19b, YML^{+}17, YS03, ZQS11, ZHX^{+}11, ZWW^{+}12, ZDM13, ZX18, dRBS^{+}12, vdZCT16, vFWTS08]. Real-Time [AL11, BTB^{+}04, BTB10, BTH^{+}13, CMF^{+}18, CK05b, ČB19, CORLS96, CMPC06, CDA99, EHH^{+}19, FS19, GSCI15, GMD13, GWBO12, GB08b, HB14, HLRC^{+}12, HRISI15, HDJ05, KKSM19, KLL12, KMDZ10, LCR16, LDSM17, LKC09a, LKC09b, LKR^{+}18, LKS13b, LXT18, MB03, MLKS18, NQX^{+}05, OKI15, PSSC17, PD04, PMP10, PKMR15, QYH^{+}18, RGF^{+}04, RHJ^{+}16, RYL^{+}18, RKSH11, RHD^{+}06, RASS17, RBLW07, SOS^{+}17, SKP07, SN10, SVGR16, SKH^{+}19, SAC^{+}08, SHR^{+}11, TLC^{+}10, THV^{+}14, VMT06, WRM^{+}10, WTW^{+}08, WFW^{+}17, WSE07, WXY17, WY19a, XESV97, YML^{+}17, ZQS11, ZHX^{+}11, ZWW^{+}12, ZDM13, ZX18, dRBS^{+}12, vdZCT16, XCZ^{+}19b]. Real-Walking [PFW12]. Real-World [LAM10, SLM18, vFWTS08]. Realism [JDA^{+}11, KSS09, KFN06, LRM^{+}13]. Realistic [CDR^{+}18, CLZ^{+}03, CORLS96, DDKA06, JWL05, ZKM18]. Reality [Ano12e, Ano14j, Ano14d, Ano14e, Ano15c, Ano16d, Ano16e, AS11, BSB^{+}18, BBC15, Bai13, BBBM18, BDB^{+}16, BSE^{+}17, BZS^{+}13, BAP^{+}17, Bil13, BDH^{+}18, BLO^{+}05, BSWL12, BLRW05, BC18b, CDAF18, CLB13, CMPC06, CLS13a, CGB^{+}13, DS17a, DS18a, DJK^{+}06, DH08, DRHK07, EPS^{+}15, FBI07, Fuc13, GS08, GLM^{+}17, GKR14, GJK15, Ger17, GWP^{+}18, GLZR17, HK10, HBESB11, HIH^{+}18, HSR18, HCP^{+}15, HBKS09, HAGS16, HJLH19, HF10, HF11, HV00, IHD^{+}18, ILMH12, JAM^{+}14, JLS15, KMS09, KJH^{+}18, KSY16, KSNY17, KOJC12, KBS13, KGAM18, KBB^{+}18, KKL11, KM10, KHSB11, KTWW13, KLD^{+}09, KLL12, KYT^{+}18, KOK^{+}19, KV98, LKS^{+}19, LH09, LRM^{+}13, LSCN09, LBS^{+}16, LvL12, LG12, LBKD09, LABS10, LDFZ14, MTM^{+}16, MNZ^{+}15, MUS16, MBZB12, Meh17, MK13c, MKT^{+}18, NMN^{+}18, NQX^{+}05, NJJ11, NTS11, NSS03]. Reality [NTT^{+}19, OIR^{+}17, OSB^{+}15, PLW12, PLE^{+}18, PTM^{+}18, PIS15, RBK^{+}15, RSBB17, RHJ^{+}16, RWBF18, RWF19, RPAC17, RBDG15, RJG17, RFSP19, Ros13, STB18, Sat13, Sch13, SH00a, Seq12, SYYC11, SvLF10, SSL08, SL11, SFL^{+}16, SBHW11, SRKL19, SB14, SJK^{+}07, SSE15, TMM^{+}13, TL11, TGW^{+}95, UKF^{+}18, VGKS12, VBV^{+}18, WRM^{+}10, WLW17, WRHR19, XST^{+}18, YC14, YFZ^{+}18, YON05, YON06, ZK17, ZHF^{+}19, ZHLR14, HWA15]. Realizing [BLO^{+}05]. Realtime [Dan16, HZL^{+}19, TWMZ19]. Rear [RLM10]. Rearrangements [ORRL10]. Reasoning [CCM12, EFN12, LS10, MDF12, OPH^{+}16, VH16, WM16]. Reasons [LXC^{+}17]. Recalibration [ANR^{+}18, KHSS14]. Recall [BBK^{+}16, HKKS18]. Recalling [SBB^{+}18]. Recirculation [WRT19]. Recognition [BBK^{+}16, HBESB11]. Recoloring [KOF08]. Recombination [FYF^{+}18]. Recommendation [HWZ^{+}19]. Recommendations [WMA^{+}16]. Recomposition [LWZ^{+}18]. Reconciliation [PDW^{+}14]. Reconfigurable [BSM06, LYY^{+}16a, LPG^{+}18]. Reconstruct [LLR18]. Reconstructed [RCW^{+}18].

Reconstructing [LGS12, SvdBLM11, WWB^{+}13, WCB^{+}12, YOS13]. Reconstruction [AGDJ10, APS^{+}14, APW16, BF01, BMR^{+}99, BMR01, BK12, BDS^{+}03, CTT^{+}16, Csé08, Cse10, Cse13, Dan16, ERL^{+}13, EM06, EVM08, EDvW19, GCL^{+}18, GXW^{+}18a, GXW^{+}18b, HDBC15, JR07, KSG^{+}16, KEP08, Kim13, KKSM19, KY06, LJHY14, LZLS16, MY96, MKHD05, MES^{+}11, MGM14, MCS^{+}08, OKI15, PBL10, PY09, PKMR15, RZNS04, SCT06, SZN^{+}18, SRKL19, SKH^{+}19, SWF^{+}16, SGM^{+}11, WBK^{+}08, WSS09, Wan11, WWC^{+}14, WCW^{+}16, WXY17, WLDW11, XZM17, YYFX18, YHJ^{+}17, YON05, ZG12, ZX18, ZGHG11, ZCFL15]. Recorder [MSM^{+}11]. Records [KCK^{+}19b, LB19]. Recovering [AFRS05, CC08, LSZ^{+}18, Wan06]. Recovery [WHX^{+}19]. Rectangle [IYIK04]. Rectangle-Packing [IYIK04]. Rectangular [KHA10, RNL09]. Rectilinear [Mao96, MSHC99, SM06, SK00]. Recurrence [ASW12]. Recurrent [CZZ^{+}19, KCK^{+}19b, SCT^{+}10, SGPR18]. Recursive [LLMB19]. Redirected [BHHM19, BSSL19, BLS15, HB13, HBT14, NSE^{+}12, SHV^{+}18, SBJ^{+}10, ZWBH13]. Redirecting [BIPS12]. Reduce [JAM^{+}14]. Reduced [ATK16, KLLR07, KCT^{+}17, TGSP09]. Reducing [HKB^{+}19, Ros11, SCL08, vdEHBV16]. Reduction [AHR^{+}11, ED06, ED07, JJ09, KC04, QM08, RD05, SZS^{+}17, SMT13, VT08, WWLM11, WFC^{+}18, WCR^{+}18]. Reductions [DWF^{+}19, SDMT16]. Redundancy [WGS^{+}13]. Reeb [DN12, DN13, PSF09, TGSP09, TDN^{+}12, TC17]. Reference [JK16, KSDD14, NB12, SM17, ZCL09]. References [LYK^{+}12]. Referential [LDN11]. Refilming [ZDJ^{+}09]. Refinement [BT13, CDS^{+}12, CDM^{+}04, Fau99, GABJ08, KSH03, WWW^{+}19]. Reflectance [FBLS05, LF97, TIS16, TLQ^{+}08, TAL^{+}07, WPG05]. Reflecting [TBHC16]. Reflection [IZM18, NSS14, PSM06, TIS16]. Reflections [CA00, HQ07, SMM12, WXKP14, WDSC07, WBD14]. Reflective [KTWW13]. Reformation [AMB^{+}13, KBH^{+}10, LCMH09, WGC^{+}08, KST^{+}14]. Reformatting [STYC12]. Refraction [MB18a, dRBS^{+}12]. Refractions [HQ07]. Refractive [KTWW13, LSK^{+}18, WH18]. Regard [RKSB13]. Regarding [SNB^{+}17]. Region [YCLJ12]. Region-Based [YCLJ12]. Regions [FHS^{+}12, HS11, KRHH11]. Registered [SPK^{+}07]. Registration [AIS18, BJM07, FYZ^{+}17, HBKS09, HQ12, HZH14, LG13, LYLG19, MDS16, RNK^{+}15, RMW09, SM09, TCL^{+}13, XF04, YON05, YON06, ZMG^{+}10]. Registrations [RLNN11]. Regression [DvVH^{+}19, HSK17, KLG^{+}16, MP13]. RegressionExplorer [DvVH^{+}19]. Regular [GCT17, HAM11, MES^{+}11, SSS06, vW14]. Regularization [GXW^{+}18b, LDW^{+}15, WFS^{+}16, Won16, GXW^{+}18a]. Regulatory [DWvW12, YDC^{+}14]. Reidemeister [ZWJZ12]. Reinforced [BWW^{+}17, FHG^{+}09]. Reinforcement [YWV^{+}19]. Reinforcing [BW14]. Reinventing [AAMG12]. Relates [ZOC^{+}13]. Relation [CQC^{+}08, NN11a, NN11b, YHW^{+}07, GSL14]. Relation-Aware [CQC^{+}08, NN11a, NN11b]. Relational [KGS^{+}08, PLS^{+}14, STH02]. Relations [BMLC19, BN11, CXDR19, CPC09, CdOKRV09, Hol06, ZCCB13]. Relationship [LDM^{+}18]. Relationships [BCH^{+}13, CWK^{+}07, CC07, GW11, LS09, LPK^{+}13, LBT^{+}18, MAK08, NR18, SW13, TWC^{+}18]. Relative [MDS16, ZLB^{+}05]. Relativistic [MGW10]. Relativity [WBE^{+}06]. Relaxation [LPQF14, YLSL11]. Relaxing [RGE19]. Relevant [AAFG18, WLL^{+}16]. RelEx [SFMB12]. Reliable [GLM06, HHCL01]. Relief [FYY^{+}19, LWYM12, SRML09, SJM14, WTP^{+}19, YHJ^{+}17, ZZL^{+}15]. Relighting [CBLD11, LDR00, WPC^{+}13]. Relocalization [GSCI15, KCH11]. Remapped [HSR18]. Remeshing [AGL06, HYB^{+}17, LJX^{+}10, VCP08, WYL^{+}19, YBZW14, YW16, YFC^{+}19, ZHLB13, Zhu05]. Remodeling [CAH^{+}13]. Remote [CMP14, GLB16, LS07a, OIR^{+}17, PAB^{+}08, SJH^{+}07, YLC^{+}19]. Removal [JWS04, WWYP19]. Rendered [MDG00, PFK^{+}08, WQ07]. Renderer [FSTG16, WMS98]. Renderers [YESK95]. Rendering [ABCO^{+}03, Ano05d, Ano09b, BES12, BSS^{+}13, BGT12, BLM96, BLL19, BMTD05, BHWB07, BTB10, BTHD11, CR08, CICS05, CBPS06, CWM^{+}09a, CLZ^{+}03, CCB11, CMF^{+}18, CL06, Chr03, CH03, CMSW04, CHM11, Csé08, DH02, DDKA06, EMRY02, EHS13, EMP09, Elb95, EJR^{+}14, EBRI09, FM07, FM12b, FSW09, FAW10, GW06, GMS^{+}07, GUO00, GAMD10, Guo95, GQGP17, HR07a, HAAB^{+}18, HCS^{+}07, HIH^{+}18, HWHK16, HQS18, HMBG01, HSR13a, HSR13b, HA04, HLY10, HHM14, HK99, HBC12, HQ07, HC05, HGH^{+}10, IWR^{+}18, IZM18, JV09, JVDF19, JWC05, JM10, JKRY12, KSH03, KV03, KRW19, KKKT18, KKSS13, KWP01, KWH00, KKH02, KM16, KJL^{+}12, KKPS08, KMLM16, Lac96, LA11, LRN96, LKC09a, LS13a, LDC96, LS07b, LRZM11, LSR^{+}13, LY06, LR11, LHZ^{+}04, LCDP13, LLR18, LWP^{+}06, LMT^{+}03, LMC02, LLY06, LLPY07, MDM10, Max95a, MBT^{+}18]. Rendering [MAWM11, MTS07, MPBM^{+}18, MRT00, NK06, Ney98, NJ16, NvdVS00, NTS11, OWS15, PCY08, PLW12, PJ03, PAB^{+}08, PH08, PSM06, QCH^{+}14, RBG07, RE01a, RZHB^{+}08, RPAC17, RE01b, RDB^{+}12, SE17, SWB^{+}00, SLM18, SSB^{+}17, SF19, Sel15, SD11, SKMH14, SM97, SHC^{+}09, SEH08, SLW^{+}10, SHR^{+}11, TK14, TLQ^{+}08, TWBBM17, VF13, VPB^{+}11, Vis15, WZQK04, WW07, WSY07, WTW^{+}08, WK13, WX13, WXKP14, WH18, WLMP19, WBK^{+}07, WDC^{+}07, WKZL04, WPG05, WMGE12, WII^{+}12, WC13, WLLC15, WY19a, XESV97, XTY^{+}11, YHLJ08, YLX^{+}12, YCLL08, YL95, YSGM05, ZK06, ZM13, ZDM13, Zha14, ZEC08, ZWM13, ZG06, Zhu05, ZBDS12, ZD18, dRBS^{+}12, vRKEE17, DMAM19, LCM07]. Renderings [LB17]. Renderserver [PAB^{+}08]. Reordering [SCL08]. Reorganization [LBGV13]. Reorientation [BL15, PFW09]. REP [RF11]. Repair [BDK98, HF11, NT03, PPM^{+}11, ZJH07]. Repaired [CC08]. Reparameterization [YLY^{+}12]. Repeated [HJW99a, HJW99b, LPLT11]. Repetitive [XA10]. Replays [PKG12]. Replication [DJ18]. Replications [HKKS18]. Reports [BKM13, LTM18]. Repositioning [BL15]. Repositories [GCL^{+}15, KLK^{+}09]. Repository [PFG08]. Representation [AW03, CMPS97, CLB11, DVC18, FGF^{+}05, GIS03, HF06, Hu16b, KLCK17, KHPS07, KDA^{+}09, LF97, LHH^{+}12, LHFY12, LPLT11, LKT13, LXRY18, MZFM98, MB01, Mur95, NDS10, PB13, PBA10, SXM17, SG05, SLGM19, SP96, SCYW16, TvET14, WLSW08, WD10, YSD^{+}17, YS03, ZBZ^{+}13, ZKG07]. Representation-Independent [PBA10]. Representations [GK07, HHH16, HJC14, JV09, JEG12, JDL09, LZZ^{+}19, MCK12, WJD17]. Representative [LBR^{+}17, LPCRH19, TLLH12]. Representativity [PBPP11]. Representing [GHL15, LVRL06, PW13, RSFH14]. Reproducible [RB18]. Reproduction [CK16, IDAK15, LRP97, MK13c]. Request [SSMG13]. Request-Flow [SSMG13]. Requirements [GNCM^{+}16, TGW^{+}95, vLFR17]. Res [ETO^{+}10]. Resampling [CMF^{+}18, Mao96]. Research [BŠG^{+}09, EGG^{+}12, IIS^{+}17, JF16, KLM^{+}08, KBB^{+}18, Nie96, SDW09, TM04, WOCH09]. Resetting [BHHM19]. Reshaping [ZLDM16]. Residential [RB11]. Residue [ZT09]. Resistant [LLW06, YPRI17]. Resizing [DZL^{+}14, WLLM13]. Resolution [AL11, AJ19, CWK^{+}07, CAN14, DMAM19, FYTL19, HKB^{+}19, IWR^{+}18, KGS^{+}08, KHZR18, LMK07, LWW^{+}07, PS12, PHF07, SSIF09, SKMH14, TLQ^{+}08, WJ08, WLSL17, WPS^{+}16, WDW16, YBZW14, vAPP^{+}11, BPC^{+}10]. Resolution-Dependent [TLQ^{+}08]. Resolutions [BLLS17]. Resolving [FWT^{+}04, Kra16]. Resource [BLS15, DDBR^{+}19, MMT^{+}14, MKN^{+}07]. Resources [KSY14]. Respond [BZS^{+}13]. Response [MZH^{+}08, TWSM^{+}11, WGR^{+}18]. Responsive [Per95]. Restoration [LWYM12]. Restoring [QPNK18]. Restricted [CVC^{+}12, WLL^{+}12, YBZW14]. Result [BM10]. Result-Driven [BM10]. ResultMaps [CDF09]. Results [GNSP^{+}14, HOG^{+}12, NZ06, NB12, PHE^{+}18]. RetainVis [KCK^{+}19b]. Retargeting [DWK^{+}16, KPBL16, KEJK19, LLG17, LLY^{+}13, PPZ^{+}12, PMH18]. Retention [CB15]. Retexturing [GSPJ08, MTB17]. Rethinking [DWS10]. Reticulum [MQF06]. Retinal [FZC^{+}07]. Retrieval [ADP02, CLAL12, EHBA11, HKBE12, JSR^{+}19, LLZ^{+}16, LJH^{+}18, SGP^{+}19, SSW18, TL07]. Retrieving [HK09]. Retro [WCB^{+}12]. Retro-Deformation [WCB^{+}12]. Reusable [HA18]. Reusing [ZHF12]. Reveal [DW14, HRD^{+}19]. Revealing [AAFW17, ASW12, CC07, CPC09, GZ11]. Reverberation [AM13]. Reverse [SZK15]. Review [FIBK17, IIC^{+}13, KBB^{+}18].

Reviewer

[Ano05e, De 18b, Lin14c, Ano04b, Ano15n]. Reviewers [Ano00a, Ano01b, Ano03a, Ano06, Ano07, Ano08a, Ano09c, Ano10b, Ano11g, Ano11b, Ano12j, Ano12k, Ano12b, Ano13q, Ano13b, Ano15b, Ano15p, Ano16r, Ano16u, Ano17m, Ano17n, Ano17a, Ano17b, Ano18g, Ano19h, Ano19b, Ano19c, Ano19j, Ano19l, Ano13p, Ano14b, Ano14u, Ano18i]. Revisited [AS19, BLB^{+}17, HSH04, PWIG18, SCT06, SGQ16]. Revisiting [KBB^{+}18, PDF14, SSD^{+}08, WWS^{+}18]. Revitalized [AMC10]. Revolution [AG16a, AG17, Bai13]. Revolving [SFB^{+}12]. Reweighted [LYLG19]. Rewriting [NSS03]. ReX [KSY14]. Reynolds [KWDG11]. RFID [WCQ^{+}09]. RGB [GSCI15, IAS19, PZLZ17, PP09, Zho16]. RGB-D [GSCI15, PZLZ17, Zho16]. RGBD [LSWZ17, SKC^{+}19, WXY17]. Rheoscopic [BB07, HMT10]. Rhetoric [HD11]. Rhythm [MDL^{+}17]. Ribbon [WKW06]. Ricci [JKLG08]. Rich [CSL^{+}10, GBWI17, LBD13, MGPH06, SS13b, WBD14, YEB16]. Ridge [SP07]. Riemannian [GSZ^{+}13]. Rig [HSK17]. Right [PBK^{+}12, SHV^{+}18]. Rigid [BHTF07, BTT09, Dru08, ELF13, GSM^{+}14, GXW^{+}18a, GXW^{+}18b, KL14a, KMT14, LG15, LYLG19, LDW^{+}15, NWI17, ORC07, SM04, SL08, SLS^{+}17, TCL^{+}13, TF06]. Rigid-Fluid [BTT09]. Rigs [MAF11]. Ring [BMJK09, WLHD17]. RingText [XLZ^{+}19a, XLZ^{+}19b]. Risk [DBP14, HHO^{+}17, MMB^{+}19]. Risky [JOR^{+}19]. RIVA [LDC96]. RNA [LBLH19]. Road [BGC^{+}11, HS11, HR96, LLHL14, WSLL12]. Road-Crossing [BGC^{+}11]. Roadmaps [GSA^{+}09]. Roads [JOR^{+}19]. Roaming [CMCL06]. Robots [LPF^{+}19]. Robust [AZM12, CRT04, CL11, FYP10, GWE^{+}19, GXW^{+}18a, GXW^{+}18b, HQ12, JCWD14, KHS^{+}18, KCH11, KC04, LZH^{+}07, LYLG19, LYY^{+}16b, LDC16, NHYY18, PBPP11, RKZZ19, SSIF09, SPP^{+}14, SH12, SZ12, WWZ^{+}18, YRP19, YT02, ZTP05]. Robustness [BEK10, LB03, SWCR15, SRW^{+}16]. Robustness-Based [SWCR15]. Rock [Ano14p, JFBB10]. Rod [ZQS11]. Role [BDH^{+}18, GMD^{+}17, GPK14, SSK^{+}16, YaKSJ07, ZKM18, cKJG^{+}12]. Rolling [BDF16, EDF08]. Roman [CDW^{+}16]. Room [LLBS17]. Room-Scale [LLBS17]. Ropes [WRF^{+}11]. Rosenbrock [CAP18]. RoSy [LJX^{+}10]. Rotating [CDL^{+}16, RM17, TT05]. Rotation [FWK16, GST16, LH11, LG15, RSBB17, SHV^{+}18, WP19, ZLK^{+}18]. Rotational [CK05b, PZ11, ZBMY14]. Roth [Ano05c]. RotoTexture [FH06]. Rough [MBT^{+}18, dRBS^{+}12]. Roughness [UBH19]. Round [SHV^{+}18]. Route [KCJ^{+}10, QWC^{+}09, SLQW17]. Route-Zooming [SLQW17]. Routes [DSSK08, TYSN06]. Routing [LMZ06]. RSVP [BDK98]. RTcams [HCS^{+}07]. Rubbing [LWYM12]. Rule [FBL^{+}18, NT99]. Rule-Based [FBL^{+}18, NT99]. RuleMatrix [MQB19]. Rules [CvW18, MQB19]. Run [KWS^{+}14, Lin16b, SJH^{+}07, FSW09]. Run-Time [SJH^{+}07]. Running [STS^{+}14]. Runs [MGJ^{+}10]. Runtime [IPD^{+}07].

s [BWS^{+}19]. Saccadic [BL15]. Sacrificial [UKW19]. Safe [JOR^{+}19, NMN^{+}18]. Safely [CVC^{+}12]. Safety [BES12, BC18b, CB15, Chi16, LLQ^{+}17, TKE16]. SAH [Wal12]. Saliency [FKLT10, IV11, KV06, MHD^{+}18, WSYM17]. Saliency-Assisted [IV11]. Saliency-guided [KV06]. Salient [WM18]. Sample [BKW10, HEFR18, PSM06, QK04]. Sample-Based [BKW10, PSM06, QK04]. Sampled [ATT12, CDK04, CYW^{+}16, CL11, Wan06].

Sampler [VML97]. Samples [BT13, FHHJ08, LPCRH19, YCLL08]. Sampling [AT05, AGY^{+}17, AEM09, AFRS05, BMWM06, CSG^{+}19, CDBR14, CYC^{+}12, CCS12, DWB^{+}06, FYWY16, KSSW09, KHW^{+}09, KBH13, LL04, LBR^{+}17, LPG15, MY96, Max95b, MES^{+}11, MWK^{+}08, NHEM17, PBPP11, PRA^{+}10, QLLM13, WMWL11, WSW16, WLLC15, WCA^{+}17, WM19, YWW14, YHH^{+}19, YXSH13]. Samplings [LYS^{+}10]. Sand [ZLZY18]. SaNDVis [PGU^{+}13]. Sanity [CLKS19]. Satellite [LAP19, YXG^{+}10]. Satellite-based [YXG^{+}10]. Save [TKE16]. SCA [BvdP12, BDC17, KS14a, KL14b, SK15]. SCAAT [HR11]. Scaffolds [CG08]. Scagnostics [DAW13, DW14, MTL18]. Scalability [GBHP08, YN06, LKS^{+}19]. Scalable [ADG11, AAMG12, AAH^{+}13, BSM06, BPL^{+}19, BVB^{+}11, DVCD07, DK11b, EMP09, FSW09, FH07, GW06, GHGM06, GXY12, HSZ^{+}11, JBH^{+}09, KHPS07, KBGE11, LFH06, LMC02, MHDH07, MCA^{+}10, NSW^{+}17, NR18, NHN07, PAB^{+}08, RLM10, SCM^{+}06, WLD^{+}19, WLLC15, YEB16, YC14, YXM^{+}15, YXG^{+}10, ZK14a]. Scalar [CLB11, GBPW10, GSS^{+}19, GJR^{+}14, GNP^{+}06, GNPH07, HM10, LJHY14, PQMCR17, PMW13, RE01a, RKG^{+}11, STS06, SWC^{+}08, TN11, TN13, TN14, TP12, YL95]. Scale [AvHK06, APS^{+}14, BMA^{+}19, BL07, BWT^{+}11, CXDR19, CMK15, CLW18, CPG09, Dic14, DGWC10, DB07, EDF11, EBB^{+}15, FSW09, GWFI14, GBWI17, GCL^{+}18, GDST16, GHS^{+}19, HAAB^{+}18, HDSC19, HA06a, JH13, JST^{+}10, JHGH08, JSB13, KAKC18, KHSS14, KKP^{+}17, KSSW09, KLJ^{+}09, KCK^{+}19a, LLBS17, LFLH07, LBL19, LZH^{+}13, LXR19, LBS^{+}19, MDS16, MDS^{+}18, MDL^{+}19, MGM09, MOC^{+}14, NLS11, PFW12, PLE^{+}18, PY09, RKSB13, RKG^{+}11, SXM17, SZY^{+}18, SKH^{+}19, TKE16, VWvH^{+}07, WSLL12, YXM^{+}15, ZMT^{+}19, ZHL^{+}09, LFH06, LWL^{+}17, SMDS14]. Scale-Adaptive [MDS^{+}18]. Scale-Free [JHGH08]. Scale-Space [KSSW09]. Scale-Varying [DB07]. Scaled [CWC^{+}06, KCT^{+}17]. Scales [CVG13, SSGM19]. Scaling [LMZ^{+}14, LT18, ZKK02]. Scamming [KFS^{+}19]. Scan [AHRG10, RLNN11, YR95]. Scan-Based [RLNN11]. Scan-Conversion [YR95]. Scanned [SGJM18]. Scanning [PWG17, RBK^{+}15, TZL^{+}12, WG16, ZZM06]. Scans [BMR01, LSZ^{+}18, RHR^{+}09, VAB12]. Scatter [FHSW13, HOG^{+}12, LT10, MG13, RGE19, WHZ^{+}18]. Scatter/Gather [HOG^{+}12]. ScatterBlogs2 [BTH^{+}13]. Scattered [BF01, KH01, LWCS96, LWS97]. Scattering [ASW13, AWB11, KKSE17, NSS14, Sil95, YGX^{+}09, ZW07]. Scatterplot [CCM13a, EDF08, HKF16, KCPE16, SW12, SMT13, VMCJ10]. Scatterplots [BW08b, BGR06, CCM^{+}14, GCNF13, LWCC18, MPOW17, WCG^{+}19]. Scenario [BC18b]. Scenarios [GHE19, LBI^{+}12, TNT17]. Scene [APW16, FYF^{+}18, FM04, LSJ^{+}15, LXRY18, MB01, NHYY18, PDRK19, SSI99, SG99, WLMP19, ZX18, ZHC18, ZZM06, ZGM18, ZKG07]. Scenes [BGCS17, GN12, LBD13, LRP97, LWW^{+}07, LDR00, MFZ^{+}17, NKP^{+}15, Ney98, PZLZ17, SLM18, XA09, YYT16, LCM07]. Scenic [Hu16a]. Scent [PBE19]. Scented [WHA07]. Schedule [JHP^{+}14]. Schedules [PGSF16]. Scheduling [KLL^{+}13, TK14]. Schematization [vGMSW15]. Scheme [BPL^{+}19, Cse13, HPJG08, LH03, LYY08, LY06, LDC16, PB13, RF11, VP04b, ZC06, Zhu05]. SchemeLens [CLB^{+}16]. Schemes [Csé08, LRN96, SFBP09, TdJ14, TdJ15, ZTP05]. Schlieren [BPS^{+}11]. Science [Ano12g, BSS18, BRS18, CGGR18, DJ18, Ert10c, GGHZ19, MW13, UMW^{+}12, vW11]. Scientific [AL06, AB01, BKW16, BW17, FHKM17, HE99, KAM^{+}08, KMDH11, KH13, KCS^{+}16, LDC96, LS16, Mar18, MGKH09, Nie95, NTT^{+}19, OR98, OR99, OSBM14, RBGH14, Rob98, SKK^{+}14, WJA^{+}17, WBP07, WKZL04, YSI^{+}10]. Scientist [DLW^{+}17]. Scientists [GKL^{+}13]. Scission [DCK^{+}12]. Scissor [ZWC^{+}16, ZTA12]. SciVis [Ano19i, Ano19j]. Sclerosis [SGQ16]. Scream [BS11]. Screen [AJ17, CCAL12, DK10, GQM^{+}18, IWR^{+}18, MYI13, TIS16]. Screen-Space [DK10, IWR^{+}18]. Screenit [DSG^{+}17]. Screens [BGR06, DSG^{+}17]. Screentone [YHL^{+}17]. ScreenX [LLKN17]. Scripting [PUNI11]. Sculpting [CCB^{+}18, GM05, HRD^{+}19, WM13a]. Sculptures [STS^{+}14]. SD [HRN^{+}03]. Seamless [BMY05, FPB17, PWIG18, TS08]. Seamlessness [SLGM09]. Seams [PST^{+}15]. Search [BBH^{+}17, BWK^{+}13, CWT^{+}08, CDF09, GNSP^{+}14, GMD^{+}17, IDA^{+}14, JWC05, KBGE11, KOBH19, KOK^{+}19, LRM^{+}13, LCP^{+}13, LDFZ14, NHB^{+}17, NZ06, NB12, RSBB17, RH19, RGP^{+}12, SWL^{+}14a, SCM^{+}06, ZWBH13, vHP09]. Search-Based [CWT^{+}08]. Searching [SPP^{+}14, WPS^{+}09]. Seated [KKKT18]. Second [GRT17, HLL97, HVSW11, KH16, SK10, WXKP14]. Second-Order [GRT17, HLL97, SK10, WXKP14]. Section [AD12, Ano12g, BvdP12, BDC17, BHTY15, BGK11, BKL18, CCH14, CLS07, CW11, DFQ12, DW17, Ert10c, FBI07, GKR14, GJK15, GW13, GMM05, HK10, HVY16, HKQ13, HP04, HLM10, ILMH12, JLS15, KMN04, KRTvW06, KS14a, KKL11, KHSB11, KPGL12, KL14b, LS06, LSCN09, LST^{+}16, LBKD09, LABS10, MSW19, MH10, MW13, Moo03, MYM08, NSvW11, OA11, Ota17, PZ12, Pur09, Qin09, RvWT05, SK16a, SK15, SGR06, SSL08, SL11, SW17, TL11, VW12, WLW17, WM05, WY19b, vWMT04, vW11]. Sectional [GGC^{+}17, SKYS14]. sectioned [DBTH07]. Sections [LB17, SFH06]. Security [ABC^{+}19, MKN^{+}07, SSG12]. Sediment [UDSL18]. Sedimentation [HVF13]. See [BC18a, DFG^{+}14, DTT^{+}17, GMD^{+}17, GIMS18, GMY11, GLB16, HIH^{+}18, HRISI15, IK15, IDAK15, IAIK16, IHS17, LCR16, LH16, LHH16, LHC10, MIO^{+}15, PIN^{+}15, QPNK18, RHJ^{+}16, RWF19, SJK^{+}07, WVFH12]. See-Through [BC18a, DFG^{+}14, DTT^{+}17, GIMS18, HIH^{+}18, HRISI15, IK15, IDAK15, IAIK16, IHS17, LCR16, LH16, LHH16, LHC10, MIO^{+}15, PIN^{+}15, QPNK18, RHJ^{+}16, SJK^{+}07, RWF19]. Seed [HA04]. Seeding [FA15, MJL^{+}13]. Seeing [TAL^{+}07]. Seeking [BEDF16]. Seen [DBN06, IIS^{+}17, HMTR19]. Segmentation [AABH^{+}16, AZC^{+}12, BMA^{+}19, BWT^{+}11, DMR04, FZC^{+}07, HRN^{+}03, HNR^{+}06, IVJ12, JBH^{+}09, KSI^{+}96, KW11, LCS06, LY12, MBS^{+}04, MW99, MK09, NHYY18, PRH10, PPM^{+}11, SHM10, SF04, TWSM^{+}11, Wan08, XTY^{+}11, XLZ19c, ZTA12]. Segmentations [GJG^{+}15, vLBB16]. Segmented [CL06, RSOW18]. Seifert [vWC06]. Seismic [PGT^{+}08, YXG^{+}10, PGT^{+}08]. SeiVis [HFG^{+}12]. Selecting [FHSW13, LWL^{+}17, MJ09, WWZ^{+}18]. Selection [AT16, BAF^{+}13, BTJ^{+}13, CWL12, DMC15, EWWL98, JS06, JSR^{+}19, KPB14, LFA^{+}16, MLMP18, PLC^{+}11b, PBC17, SHVV16, TMWS13, THV^{+}14, WGS07a, WSS09, WHZ^{+}18, WWF^{+}19, YEII12, YEII16]. Selections [CvW18, vdEvW14]. Selective [CDM^{+}04]. Self [AL11, BPS13, BDH^{+}18, BSWL12, GD01, HDBC15, JNC^{+}15, JAO^{+}14, SKB^{+}18, SG09, SLF^{+}12]. Self-Avatar [BPS13, JAO^{+}14]. Self-Contact [BDH^{+}18]. Self-Generating [SG09]. Self-Intersection [AL11, GD01]. Self-Motion [BSWL12]. Self-Occlusion [HDBC15]. Self-Occlusions [JNC^{+}15].

Self-Organizing [SKB^{+}18].

Self-Overlapping [SLF^{+}12]. SellTrend [LSS09]. Semantic [AW03, BGR06, CQM10, EFN12, FWL17, HTE11, KTE15, LHJ^{+}18, LBT^{+}18, RBG07, SMER06, SA06, SMNR16, TWC^{+}18, WOO17, XCZ^{+}19a, ZGW^{+}19]. Semantic-Based [XCZ^{+}19a]. Semantic-Driven [LHJ^{+}18]. Semantic-Level [WOO17]. Semantics [HBM^{+}13, JSR^{+}19, KIL07, SKK06]. SemanticTraj [ADWK^{+}17]. Semen [DTW^{+}15]. Semi [Ano09b, BHST17, IDAK15, KGP^{+}13, KBVH17, MIO^{+}15, RNK^{+}15, TW18, Wu16, YYD^{+}19]. Semi-Analytical [Ano09b]. Semi-Automatic [KGP^{+}13, MIO^{+}15, RNK^{+}15, TW18]. Semi-Dense [BHST17]. Semi-Explicit [YYD^{+}19]. Semi-Lagrangian [Wu16]. Semi-Parametric [IDAK15]. Semi-Spatial [KBVH17]. Semiautomatic [SG09]. Semiotics [MRO^{+}12, VFR13]. Semiregular [DHM13b, PA06]. Send [GPC^{+}17]. Sensation [KKKT18]. Sense [CDW^{+}16, JAAL18, LKH^{+}16]. Sensemaking [AN13, EFN12, GLK^{+}13, aKS12, LKH^{+}16, MT14, NXW^{+}16, SMNR16, SZW^{+}19, ZGI^{+}18]. SensePath [NXW^{+}16]. Sensing [SD11, WAG^{+}12]. Sensitive [DN12, FW08, KGG^{+}12, SZ11]. Sensitivities [KRRW19]. Sensitivity [BVPtHR09, CCM13a, CCM12, LLBS17, MLMP18]. Sensitivity-Aware [MLMP18]. Sensor [BMJK09, FM04]. SentenTree [HWS17]. Separability [LWS^{+}17, MCG12, WCG^{+}19]. Separating [CMF12, FBTW10]. Separation [DKM06b, KHL99, RS12]. September [RPHI08]. Seq2seq [SGB^{+}19]. Seq2seq-Vis [SGB^{+}19]. Sequence [ADG11, CXR18, FNM13, GS14, GXZ^{+}18, GJG^{+}19, HDR^{+}13, LLMB19, MLL^{+}13, NJBJ09, SGB^{+}19, UDSL18, VN19, YYFX18, ZCD19, vdEHBvW14].

Sequence-to-Sequence [SGB^{+}19]. Sequences [CvW18, DSP^{+}17, HSF^{+}06, LWD^{+}17, MRT00, NTA^{+}19, RG95, UDSL18, VBW16, VJC09, WMWL11, War09, WG12]. Sequential [BB12, HKL17, JER16, MDL07, TTS10, WBD14]. Serial [DBTH07]. Serial-sectioned [DBTH07]. Seriation [SZW^{+}19]. Series [ASMP17, BAF^{+}13, CSWP18, DAW13, GSL^{+}17, GTPB19, HKF16, JME10, KLK^{+}09, Kin10, KBK11, LS13a, LCZ^{+}19, MHG10, MGMP18, MMMY97, SCL^{+}12, SKU^{+}12, SGAS16, SSGM19, WBJ16, WHZ^{+}18, XCH^{+}14, ZCPB11]. Serious [CB15, Chi16]. Server [PJ03, Wah14]. Servers [GHGM06]. Service [ZYM^{+}14]. Servoing [CMPC06]. Set [ABCO^{+}03, ARRC11, AR17, AJ19, BGB15, BW01, CDS^{+}12, CPC09, FMH08, HKC^{+}12, KS00a, KS00b, LTKF08, LWC^{+}18, MRS^{+}13, SMDS14, TSB^{+}05, WT10b, YEB16]. Set-Typed [FMH08]. Sets [AAMH13, BDSW13, BBP08, BMPB08, CML^{+}12, CPC09, CJTM05, CMP14, GWBO12, GTLH01, HHG14, HPNT18, JvdLR13, KSH03, KRW19, KHDL07, KWP01, KBH06, KGJ09, LKHW04, LT16, LGS^{+}14, LY12, LS16, LKT13, MS08, Mal05, MVN^{+}19, PSKN06, PSN10, PHJ^{+}10, PFK^{+}08, RSFH14, RP12, SBSG06, WKZL04, WMK13, YHW^{+}07]. Seven [LBI^{+}12]. Severe [PMvWC05]. SGaze [HZL^{+}19]. Shaded [Gor02, NWHWD16]. Shader [GHE19, PHF07, RKK16]. Shaders [VT08]. Shading [AD16, BHB04, BN12, JFZ^{+}18, LDSM17, SMP11, ŠTPV12, Ste98, WEE03]. Shadow [MDL^{+}19, OBS^{+}15, SFC^{+}07, WLMP19, ZH07]. Shadow-Driven [ZH07]. Shadows [ASDW14, BBBM18, IHK05, IIS14, JWS04, KSDA16, MDL^{+}19, WHL16, ZC03]. Shaft [KKSE17]. Shape [BH07, BBiA09, CVC10, CZZ17b, CZZ^{+}19, CL09, DVC18, Far12, FSHH12, FCSF17, FDC^{+}18, GSCO07, GIS03, GK95, GGLQ19, HBESB11, HHZH17, HLD^{+}08, HNR^{+}06, HTZ^{+}11, HKYM17, IFP97, JZLG09, KHSI04, LZZ^{+}19, LL14, LFP07, LBZ^{+}11, LWZ^{+}16, LSWZ17, LSM03, MGMP18, ML19, NGK18, PSKN06, PPM^{+}11, Qin09, SHM10, SAR96, SPK^{+}07, SPL^{+}13, SCOIT05, SO17, TH13, TAL^{+}07, VPB^{+}11, WX17, WKB^{+}13, YSZ04, YHMY08, ZHX^{+}11, ZZL^{+}15, ZZG^{+}12, ZDZ18, ZTZX13, ZHL^{+}09, vLBB16, SKW^{+}11]. Shape-Based [HNR^{+}06]. Shape-Encoding [WKB^{+}13]. Shape-Preserving [NGK18, ML19]. Shapes [ASvdP14, BES12, BSG18, CSPN11, HGWW18, HQ12, JSG03, KW14, KBS13, LYY^{+}16b, LGS12, MTB17, RvWT08, YOS13, vGRSW14]. Shaping [ZK08]. Shareable [RB18]. Shared [BSO^{+}12, CMCL06, GBP19, NBM19]. Shared-Memory [GBP19]. Sharp [AFRS05, CC08, KY06, NDS10, Wan06]. Sharpen [AFRS05]. Sharpening [DSP^{+}17]. Sharpness [CC08]. Sharpness-Dependent [CC08]. Shear [GLH^{+}14, Lac96, SBV^{+}11]. Shear-Warp [Lac96]. Sheet [RHR16]. Sheets [ATT12]. Shell [ELF13, GLB^{+}06, ZQS11]. Shelves [MSSH14]. Shift [BSL^{+}14, CL09, FYWY16]. Shifting [ZK17]. Shifty [ZK17]. Shimmering [ZHF^{+}07]. Shneiderman [Ano13c]. Short [DSSK08]. Shorthand [vGRSW14]. Shot [MI13]. Should [RZP12]. Show [GR04, LBS^{+}19, MHS07, vHP09]. Shrinking [BSSL19]. Shutter [BDF16]. SI [LLC15]. SI-Cut [LLC15]. Sibson [PLK^{+}06]. Sick [FRG^{+}19, YAE07]. Sickness [FRG^{+}19, LYL19, YAE07]. Side [AH11]. Side-by-Side [AH11]. Sided [YSL^{+}13, YHJ^{+}17]. SIGGRAPH [BvdP12, BDC17, GW13, KS14a, KL14b, MY14, SK15]. SIGGRAPH/Eurographics [BvdP12, BDC17, KS14a, KL14b, SK15]. Sigma [PBA10]. Signal [GS16].

SignalLens [Kin10]. Signature [GSCO07]. Signatures [CBH^{+}06, RML12, TL07, TSH^{+}14, Vis15, WFC^{+}06]. Signed [BA05, KL14a, KDBB17]. Significant [AAH^{+}13, MZFM98]. Significantly [KLLR07]. Silhouette [WTW^{+}08]. Silva [Ano14f]. Silvered [MSM^{+}11]. SIMD [SvW98]. Similar [KDM^{+}16, LWS^{+}17, LG13]. Similarity [CCK07, FSE12, FIB^{+}14, GTPB19, GNSP^{+}14, HBG11, Lin16b, MJL^{+}13, PFP^{+}11, PDW^{+}14, TIW^{+}19, VJC09, ZK10]. Similarity-Guided [CCK07]. Simple [AGY^{+}17, ER97, NMN^{+}18, PSSC17, SAS16, TNT17, YHL^{+}17, ZK12]. Simplex [YHMY08]. Simplicial [CMS06, WT10b]. Simplification [CMRS03, ESV98, GHK97, Gué99, GNP^{+}06, HHVW96, HHG14, HPNT18, LT99, LP02, MLL^{+}13, NE04, NT03, PJ03, RP12, SWCR15, TP12, THJ99, VCL^{+}07, VJN^{+}15, YSZ04]. Simplified [DBTH07, LP02]. SimpliFly [VJN^{+}15]. Simplifying [TNB11]. Simpson [AW14]. Simulate [CLB13, JAO^{+}14]. Simulated [ASvdP14, KH01, VABW09, WOO17, WBA^{+}14]. Simulating [BBBM18, DGW11, KO12, KM10, SKY12, ySKK07, TSB^{+}05, WLMK04]. Simulation [BKL^{+}11, BTB^{+}04, BM10, CZZ17a, CAP18, CM14, CMK15, CDL^{+}16, CK05b, CLMO17, CLEK13, CDA99, Dic14, Dru08, DBP14, DGAB16, ELF13, GLRH13, GH00, GMD13, GIK^{+}07, GLB^{+}06, GPP^{+}16, GLX17, HB13, KSS09, KZX^{+}14, KNO15, KWS^{+}14, KWC^{+}10, LBS13, LRM^{+}13, LXB17, LBL19, LCNG14, LLCD11, LPQF14, LSH07, LTKF08, MOF09, MOF10, MGJ^{+}10, MGS^{+}14, MZS^{+}19, MT01, OMD^{+}12, OBLN17, PQF^{+}09, QMK^{+}06, RYL^{+}18, RE14, RBLW07, Sat13, SM04, SSH14, SOL^{+}16, SSF13, TF06, UK12, USKD12, WZW^{+}05, WBK^{+}07, WMK13, WSLL12, YYY16, ZZSS10, ZQS11, ZLZY18, ZHF^{+}07, ZCF^{+}19]. Simulation-Based [KWS^{+}14, SOL^{+}16]. Simulations [BSM^{+}13, BBG^{+}18, BWT^{+}11, BvL06, CL18, DCH^{+}17, FSHH12, KHA12, LLB^{+}12, LHH^{+}12, Lin16b, LBS^{+}19, OJCJP16, PvdBC^{+}11, RWG^{+}12, RWF^{+}13, SLMA06, VBC^{+}16, WRF^{+}11, XSZ^{+}17]. Simulator [FRG^{+}19, NDM^{+}97, RGF^{+}04, STH13]. Simulators [SSH14]. Simultaneous [KIS17, Zho16]. Single [BWW^{+}12, BTHD11, Dan16, HPvU^{+}18, IHS17, KKSE17, LWYM12, LPS^{+}13, LSWZ17, OAH14, WX13, WJR^{+}13, ZHC18, ZTA12]. Single-Cell [HPvU^{+}18]. Single-Click [ZTA12]. Single-Image [WJR^{+}13]. Single-Layer [IHS17]. Single-Page [OAH14]. Single-Pass [BTHD11, WX13]. Singular [BHS12, SSC^{+}16]. Singularities [LVRL06]. Singularity [JHW^{+}14]. Sinogram [YOS13]. Sinus [KKPS08]. SIRIUS [DWF^{+}19]. Site [RNK^{+}15, VWvH^{+}07]. Sitting [LIM^{+}12]. Situ [DCH^{+}17, HSS11, WPS^{+}16, WKSS05, GRS^{+}19]. Situation [UKW19]. Situational [WKCB07]. Situations [SB14]. Six [CMHL11, JWC05, ORC07]. Size [ACS^{+}18, CK16, CCAL12, CM08, JH13, JH16, ZHF^{+}19, GSL14]. Size-based [CM08]. Sized [BW17, BI12, GT19]. Skeletal [SRCP02, SRCP03, TSB^{+}05, WQZ^{+}18]. Skeleton [BKS01, CSM07, EHP^{+}11, MTL18, SLG^{+}17, WL08, XTY^{+}11, YHMY08]. Skeleton-Based [EHP^{+}11, MTL18]. Skeletonization [SJL^{+}18a, ZT99]. Skeletons [LGS12, RvWT08, SWTH07, ZJH07]. Sketch [CLAL12, EHBA11, GXH^{+}13, LCP^{+}13, MSA17, SLGM19]. Sketch-Based [EHBA11, GXH^{+}13, LCP^{+}13]. Sketches [HKYM17, OK11]. Sketchiness [BBIF12]. Sketching [CSPN11, JK16, KAM^{+}08, LKS13a, LIM^{+}12, NTT^{+}19, RWG^{+}12, RHR16, SK16b, SYYC11, TW18, WFM^{+}06, XFZ^{+}19].

SketchPadN [WM13a]. SketchPadN-D [WM13a]. SketchStory [LKS13a]. Sketchy [WII^{+}12]. Skewed [JE13]. Skewed-Aspect [JE13]. Skills [ASvdP14, CLB13]. Skinned [LHBF19, ZBO13]. Skinning [BGB15, CZZ17a, KJ12]. Skipping [HAAB^{+}18]. Sky [Bru17, SJMS19]. SkyLens [ZWC^{+}18]. Skyline [ZWC^{+}18]. SLAM [APV^{+}15, VARS14, YYR17, ZGW^{+}19]. SLAMCast [SKH^{+}19]. Slave [TAK^{+}05]. Sleep [BDB^{+}16]. Slice [CML^{+}12]. Slices [SBW17]. Slicing [OJ15]. Slider [YNYH06]. Slideshows [WLF^{+}19]. Sliding [CLMO17, MBT07]. Slope [TGH12]. Smale [GNPH07, GBHP08, GBP12, GKK^{+}12, GGL^{+}14b, GBP19, SMN12]. Small [AT16, APP11, BBD06, BGR06, JH13, KO12, KPBG13, KBH^{+}10, LBK^{+}18, MDS^{+}17, NOB16, RKSB13, STM08, WWC^{+}14, WYL^{+}19, WFM^{+}12]. Small-Multiple [KPBG13]. Small-Scale [RKSB13]. Small-World [NOB16, WFM^{+}12]. Smart [CFM^{+}13, RLS^{+}19, XMRC17]. SmartAdP [LWL^{+}17]. SmartColor [HRISI15]. SmartCues [SA19]. Smartwatches [BBB^{+}19]. Smashing [AMA06]. SMCC [LH03]. Smelling [MYI13]. Smoke [LSY^{+}18, YYY16, vFWTS08]. Smooth [AGDJ10, BBG^{+}09, LG15, LBH18, Nie04, RN19, RL08, SM11, SEH08, vWN04]. Smoothed [SFBP09]. Smoothing [GZ14, SDHH12, SB04, WTW^{+}08, WJE01, ZHZ15]. Smoothly [IFP97]. Smoothness [SCKR08]. Smoothness-Increasing [SCKR08]. Snake [LG13]. SnapShot [PSBS12]. Snapshots [SJK^{+}12, vdEHBV16]. Snippet [GNSP^{+}14]. Snippet-Based [GNSP^{+}14]. Snooker [PLC^{+}11b]. Snowballing [LXC^{+}17]. Soccer [PVF13, WXW^{+}19]. SoccerStories [PVF13]. Social [BGC^{+}11, BN11, BRNB19, CYW^{+}16, CCM12, CdOKRV09, GKL^{+}13, HF06, HFM07, HBF08, HWS17, JHGH08, LKS^{+}19, PS06, PGU^{+}13, SMER06, SGJM18, SWL^{+}14b, TKE16, WK06, WLY^{+}14, WCS^{+}18, XWW^{+}13, YLL^{+}12, ZCW^{+}14]. Social-Event-Driven [YLL^{+}12]. Societal [KZX^{+}14]. Soft [CDA99, HCP^{+}15, IHK05, TMDO15, ZC03]. SoftAR [PIS15]. Softer [GLM^{+}17]. Softness [PIS15]. Software [CZ11, CAN14, FWL17, GUO00, HK10, HA06b, OM09, RLA^{+}13, TC09b, TL11, WLW17]. SolarView [CVS^{+}19]. Solid [BDK98, BW01, BGK11, CZZ17a, CMF12, DHM13b, Fau99, HR07a, HR96, HQ04, OHH06, PKS^{+}08, Qin09, SP96, VHLL14, YL18, ZJH07]. Solid-Fluid [VHLL14]. Solid-State [PKS^{+}08]. Solids [LIGF06, PC13, RSB96, SPB96, Wan11, WM13b]. Solution [BTC13, EG09, KFS^{+}19, LS07a, LMC02]. Solutions [ALBR16, JDSR^{+}18, LSV^{+}18, MS04, NLKH12, PLK12, XYGL13]. Solver [CMF12, JFTW07]. Solving [DRW16, LM96]. some [AL06]. SOMFlow [SKB^{+}18]. Sonic [MLCM16]. SonifEye [RNE^{+}17]. Sonification [RNE^{+}17]. Sort [LRN96, MAWM11]. Sort-First [MAWM11]. Sort-Last [LRN96]. Sorted [IK95]. Sorting [CICS05]. Sound [AM13, CSG^{+}19, CLT^{+}08, DRHK07, MAKM14, MRG^{+}15, MYM16, NTS11, OMD^{+}12, RNL09, RNE^{+}17, RSM^{+}16, RSR^{+}18, YL18, LCM07]. Sounding [YL18]. Sounds [NTS11]. Soups [PLK12]. Source [MYI13, MAKM14]. Sources [MBL^{+}06, MAF11, MTB18, SNM16]. Space [AJ17, AAB^{+}13, AMA08, BB12, BSM^{+}13, BTV14, BS02, BEJK12, BLB^{+}17, BJC^{+}19, CMSW04, DK10, DSF^{+}14, ERHRF10, FFB18, FSME14, FWT^{+}04, GCML06, GYK^{+}16, GLB16, HAAB^{+}18, HQ13, HPC^{+}13, IWR^{+}18, JH13, JA18, JZLG09, KBD^{+}11, KL03, KLM04, KLCK17, KSSW09, KDA^{+}09, KW13, LvWJH04, LMZ^{+}14, LSJ96, MWCE09, MJW^{+}13, ME09, MM08, NM13, OKB^{+}19, PSG04, PSTW^{+}17, PZ07, PBCR11, RBN^{+}19, RKG^{+}11, SFA^{+}15, SHS11a, SNHS13, SHB^{+}14, SKP07, STYC12, SPK^{+}07, STH13, TH13, TWMZ19, TC17, TNS10, VB18, WSY07, WZK12, WM13a, WFG^{+}19, Wen14, WFM^{+}12, ZWZD15, ZHL^{+}09, Ano96b]. Space-Based [LvWJH04]. Space-Efficient [TNS10]. Space-Filling [BB12, Wen14, WFM^{+}12, DSF^{+}14, WFG^{+}19]. Space-Frequency [HQ13]. Spaced [LMG06, WLZM10, ZG12]. Spaces [BFE15, BM10, DNL^{+}06, DRRD12, FWR00, IWSK07, JE13, JS98, MB01, dJOBNM17, PDW^{+}14, PCG15, RKG^{+}18, SGJM18, SLF^{+}12, VABW09, YSI^{+}10, vWN04]. Span [Ano96b, LSJ96]. SparkClouds [LRKC10]. Sparse [BT13, CTT^{+}16, DVC18, DdL14, GZL^{+}14, HYZ^{+}12, JvdLR13, KSH03, LBG^{+}16, LK11, SKMH14, WYL^{+}14, WSW16, WLT^{+}18b, YYFX18]. Sparse-Sequence [YYFX18]. SparseLeap [HAAB^{+}18]. Sparsely [CYW^{+}16, CL11]. Sparsity [LYLG19]. Spatia [BMLC19]. Spatia-temporal [BMLC19]. Spatial [AMJ^{+}12, AN13, AA11, AAFW17, BJB^{+}12, BIRW19, CJTM05, FTES13, Guo09, HIH^{+}18, HSZ^{+}11, IK15, IHS17, JSB13, KBVH17, LCZ^{+}19, LLL^{+}10, MWCR06, MMH^{+}13, MBL^{+}06, MK13c, NO97, dJOBNM17, OSS^{+}17, OPH^{+}16, PBN^{+}13, PLE^{+}18, PIS15, RKSB13, RSBB17, RCW^{+}18, SNM16, SR17, SJ06, Sim07, SPW02, TC13, VH16, VBV^{+}18, WBK^{+}08, WCD^{+}19, YPI13, BDM^{+}17, IPJT19, PSKN06]. Spatial-Temporal [SPW02]. Spatialization [TSW^{+}07]. Spatializations [HBM^{+}13, TSD09]. Spatialized [JWC05]. Spatially [LLL^{+}12, NSW^{+}17, SYYC11, WLHL13, WD08]. Spatio [BZGV14, BAB^{+}18, CHW^{+}18, FPV^{+}13, Lin16b, LXR19, SM17, SLQW17, TIS16, WG16, WDSC07, WXW^{+}19, vLBR^{+}16, DMAM19]. Spatio-Angular [BZGV14].

Spatio-Temporal

[BAB^{+}18, CHW^{+}18, FPV^{+}13, Lin16b, LXR19, SM17, SLQW17, TIS16, WG16, WXW^{+}19, vLBR^{+}16, WDSC07, DMAM19]. Spatioangular [RHZN11]. Spatiotemporal [BMJK09, BTHD11, CLS^{+}12, CLZ^{+}18, KJW^{+}18, KDA^{+}09, LKS13b, MRH^{+}10, MMT^{+}14, MLKS18, NXSL13, SvdBLM11, WMWL11, WLS^{+}19]. Speaker [Ano13t, Ano13u, Bai13, Bla12, Cze12, Fra12, Kas12, Min13, Sat13, Seq12]. Special [AD12, Ano12g, BDC17, BHTY15, BGK11, BKL18, BLRW05, CCH14, CLS07, CW11, DS17a, DS18a, DW17, Ert10c, FBI07, GKR14, GJK15, GW13, GMM05, HK10, Hag98, HVY16, HKQ13, HP04, HLM10, Joy02, JLS15, KMN04, KRTvW06, KS14a, KKL11, KHSB11, KPGL12, KL14b, LS06, LSCN09, LST^{+}16, LBKD09, LABS10, MSW19, MH10, MY14, MW13, Moo03, MGW10, MYM08, NSvW11, OA11, Ota17, PZ12, Pur09, Qin09, RvWT05, SK16a, SK15, SGR06, SSL08, SL11, SW17, TL11, Var01, VW12, WLW17, WM05, WBE^{+}06, WY19b, vWMT04, vW11, BvdP12, DFQ12, ILMH12]. Specific [CCQ^{+}14, KCS^{+}16, KGPS13, RBGH14]. Specification [AHSS14, DV95, LBW19, Sel15, SPEB18]. Specifications [KWP01, SFMB12]. Specifying [SKL^{+}14]. SPECT [MSSD^{+}08]. Spectral [ALMF19, BMTD05, BMWM06, NK06, NvdVS00, SK13]. Spectral/ [NK06]. Spectrometry [EGG^{+}12, RB18]. Spectrum [CM09, HHZH17, YNBH11]. Specular [CA00, WXKP14]. Specularities [MTB17]. Specularity [MTB18, STB18]. Speculative [EASD^{+}19, HFM16]. Speech [DNL^{+}06, KP05, LMD12, MCP^{+}06, MD12, PZ07]. Speech-Synchronized [KP05]. Speed [FC95]. Speeding [CMM^{+}97]. Speeds [NSN14]. SPH [ICS^{+}14, BK17, BKKW19, FAW10, LTKF08, PT17, YML^{+}17]. Sphere [CL06, FS04].

Spheres [KGS98, KBKG07, Mal05]. Spherical [BZGV14, CLCQ12, HFL18, LLW06, LHLW10, MBT07, NSZ^{+}17, XJF^{+}08, ZHF^{+}19, vAPP^{+}11]. SpicyNodes [DLA^{+}09]. Spin [SHV^{+}18]. Spine [AS11]. Spinel [GFG^{+}14, GFG^{+}14]. Spines [CLB11]. Spiral [BSV11]. Splatterplots [MG13]. Splatting [BVB^{+}11, Mao96, MMS^{+}98, MSHC99, NW10, PSG04, ZC03, ZPvBG02]. Spleen [GDKB17]. Splicing [SAB^{+}16]. Spline [BDHJ04, CLCQ12, Cse10, Cse13, EM06, HJW99a, HJW99b, KEP08, Kim13, KH01, LLWQ13, LQLX14, ONL^{+}12, RE01a, RZHB^{+}08, WBH04]. Splines [EVM08, LWS97, ME11a, Nie04, RZNS04, WLL^{+}12, XF04, HHQH17]. Split [SSEW19]. Splitting [ISC07]. SplitVectors [ZBG^{+}17]. Sport [SJL^{+}18b]. Sprayed [FHG^{+}09]. Spreading [HTC09, ZCW^{+}14]. Spreadsheet [JKM01, SPB08, WS09]. Spreadsheet-Like [JKM01]. Spring [DBD13, KW05, LSH07, LKT13, SVAC12]. Springs [Del08]. Sprites [MSE^{+}06]. SQ [MTS07]. SQ-Map [MTS07]. Square [PNML08, PA06]. Squares [MDS16, MGM14, PLK12, SZ11, WL08, RAL^{+}17]. Squash [yKL12]. Squash-and-Stretch [yKL12]. Squish [SM06]. SRVis [WCD^{+}19]. SSE4 [HH10]. St. [RLK19]. Stability [AHRG10, TC13, ZZH19]. Stabilization [WLHL13, ZZH19]. Stabilized [Ano14h, YLK12]. Stable [Dru08, LB15, PLW11, SSV18, Szy13, WTVP11]. Stack [JE13, XXM19a]. Stacked [BW08c, SXX^{+}19, WWS^{+}16]. Stacking [DWA10, TSAA12]. Stacking-Based [TSAA12]. Stackless [HL09]. Stacks [JST^{+}10, SMM12]. Stage [GXZ^{+}18, GGZ^{+}18, WCS^{+}18]. Staggered [CDF14]. Staggering [CDF14]. Stain [WLT18a]. Stained [Bro06]. Stains [ZCF^{+}19]. Stall [CDL^{+}16]. STAMP [GCML06]. Star [ER97, FIB^{+}14, LT13, ML19, RSS14, RSRDS16]. Stars [Ano14p]. Stasko [Ano13d]. State [AW03, BBG^{+}09, De 17c, FWL17, Flo18, KCWI13, Lin14e, LWP^{+}06, MB01, PKS^{+}08, PV06, SGPR18, WFS^{+}19, vHvdWvW02]. State-Based [AW03]. State-of-the-art [LWP^{+}06]. States [UDSL18]. Static [CZ11, FTES13, HLNW11, KISE14, KW19, LSM03, MFZ^{+}17, ZDH^{+}19, vRKEE17]. Static/Dynamic [ZDH^{+}19]. Statistical [AW14, BJEYLW01, BKL^{+}11, CK05a, GGA^{+}11, HR07b, JWSK07, KJW^{+}18, LS13b, LJWF12, MD12, MGB^{+}19, NASK18, SE17, WM08, ZCW19, vLBB16]. Statistics [CBB06, DCM13, FTB^{+}13, SSD^{+}08, ŠTPV12]. Steady [LJL04, SWCR15, WGS07b]. Steer [HOG^{+}12]. Steerable [AMA08, EASD^{+}19, KFL^{+}15, PLvdM^{+}17]. Steering [Ano12l, Ano13l, Ano14v, Ano15k, Ano16x, Ano18j, BSO^{+}12, EFN12, EGG^{+}12, FKS16, KCPE16, KZX^{+}14, MGJH08, MGS^{+}14, RWF^{+}13, WRF^{+}11, Ano13r]. Steganalysis [YPRI17]. Steganalysis-Resistant [YPRI17]. Steganalytic [YPRI17]. Steganography [ChLYL09]. StemView [GGPPS13]. Stenomaps [vGRSW14]. Stent [BSR^{+}14]. Steps [FSHH12, GSS^{+}15, XHF12]. Stereo [AHKMF11, BHST17, CK16, CCAL12, KWL14, LKL^{+}15, LDX10, RKSB13, WSYM17, WLDW11]. Stereolization [LGYG12]. StereoPasting [TZC13]. Stereoscopic [AHKMF11, DHM13a, GY02, LYL19, LODI16, LSS^{+}15, LXRY18, PBK^{+}12, SSB^{+}17, SC15, TZC13, WZQK04, WSYM17, WHR02, ZHQ^{+}07]. stereotactic [JSV^{+}08]. Stereotype [PDBG18]. Steve [Ano14d, Ano05c]. Stickies [KH19]. Sticky [RKK16]. Stiffness [GLM^{+}17, TYL^{+}18]. Stills [LWW^{+}07]. Stimulating [Ger17]. Stimulation [BTJ^{+}13, JSB13]. Stimuli [KW13]. Stimulus [MKT^{+}18]. Stimulus-driven [MKT^{+}18]. Stipple [LMT^{+}03]. Stippling [GSS^{+}19]. Stitching [GTLH01, GYK^{+}16]. STIX [SKL^{+}14]. Stochastic [ESSL11, GHK97, GHS^{+}19, Mao96, MN07, ZPG19, vRKEE17]. Stone [LWYM12]. Stories [LKS13a, LWW^{+}13, SH10]. Story [KHE09, KBI^{+}18, SPW07]. Storyboards [PLC^{+}11b]. StoryFlow [LWW^{+}13]. Storyline [TM12, THM15]. Storylines [TRL^{+}19]. Storytelling [BLB^{+}17]. Straight [IHR01]. Straightening [AH11]. Strain [SWTH07]. Strategies [AGN^{+}19, BVV^{+}19, DSP^{+}17, GR15, KWH00, MTM^{+}16, MPG^{+}14]. Strategy [MS08, PM08, PPM^{+}11, SKK^{+}14]. Stratified [QLLM13]. Streak [BFTW09, FBTW10, KGJ09, USE13, WT10a, WTS^{+}07]. Stream [BWF^{+}10, KHS^{+}18, KM96, TW18, TWHS05]. Streamable [LGLR14]. StreamExplorer [WCS^{+}18]. Streamgraph [CSWP18]. Streaming [CLZ^{+}18, HKC^{+}12, LS07a, LWG05, LKHW04, LBH18, LSV^{+}18, SRHH16, SKH^{+}19, THM15, VCL^{+}07, ZK07]. Streaming-Based [LS07a]. Streamline [CGC^{+}11, CCK07, FBW16, HM95, LS07b, LMG06, MJL^{+}13, NLS11, RS12, RT12, SS06a, SCKR08, Sun03, TMWS13, TWSS16, TEC^{+}16, USM96, USM97, WLZM10, YWSC12, ZWZ^{+}13]. Streamlines [MWSJ14, MCHM10, PW13, Wen14, WS01]. Streamlining [SLA^{+}09]. StreamMap [LBH18]. Streamribbon [USM96]. Streams [BSSB10, HBJP12, LYW^{+}16, WCS^{+}18]. StreamStory [SSGM19]. Streamsurfaces [AS19, ZDL03]. Streamtube [CCAL12, USM96]. Streamtubes [WB08, ZDL03]. Street [NZS^{+}17, SZY^{+}18]. Streets [VABW09]. StreetVizor [SZY^{+}18]. Strengths [DBB10]. Stress [CZZ17b, CRB^{+}05, DGBW09, GHN13, GLH^{+}14, MNS18, WWS^{+}18]. Stressful [RKA^{+}13]. Stretch [yKL12, SM06]. String [BAP^{+}17]. Stripes [KHH^{+}16]. Strips [MSE^{+}06]. Stroke [CRH05, LPK^{+}16, MSA17, OK11]. Strokes [GJZ^{+}12]. Strolling [DRRD12]. Strong [XA09]. Structural [GQM^{+}18, Hu16b, HRD^{+}19, LLC15, NRS15, SMER06, TKW08, WFS^{+}16]. Structure [BW14, BRT12, BL07, CQM10, CGH^{+}19, CLB11, DNP07, EBB^{+}15, FWR00, Guo95, HNR^{+}06, HTE11, KPBG13, LBG^{+}16, LBM^{+}06, LMZ^{+}14, LGLR14, MZFM98, SR17, STM08, SSW18, TLD^{+}12, Tay02, WWZ^{+}19, WKB^{+}13, XNT11, XC19, YLZ^{+}13, YEII12, ZHZ15, ZCL09, BDM^{+}17]. Structure-Aware [BL07, TLD^{+}12, YEII12, WWZ^{+}19]. Structure-Based [CGH^{+}19, FWR00, Guo95, HTE11, KPBG13, LMZ^{+}14]. Structure-preserving [CLB11]. Structure-Significant [MZFM98]. Structured [BKM13, BS16, CYB08, DF96, FPB17, GMD^{+}16, MZFM98, MHDH07, SZS^{+}17, TdJ14, TdJ15, TNS10, WHFL14, WZC^{+}15, XA10]. Structures [BJA^{+}19, BFL06, BWP^{+}10, GGTH07, GZ11, GSG96, GHP^{+}16, HVSW11, LWS^{+}17, LBLH19, MK16, MQF06, MWC^{+}12, MCS^{+}08, RHD^{+}06, RC06, SP07, SWB^{+}00, SFB^{+}12, TNB11, TGS11, WKZL04, WKB^{+}13, WJR^{+}13, WAWS18, XYC^{+}18, ZWC^{+}16, JQD^{+}08]. Structuring [MWCE09, VB18]. Student [TRd12]. Studies [AABW12, BBB^{+}19, FWD^{+}17, FIBK17, GS08, aKS12, LBI^{+}12, MHG10, PBO^{+}14, SS06b]. Study [ARRC11, BDJ14, BHZ^{+}18, BSKR19, BARM^{+}12, BOP15, BvL06, BKH^{+}11, DPW^{+}15, DLW^{+}17, DJ18, FPV^{+}13, FCL09, GKL^{+}13, GBWI17, GBFM16, GGZL16, HBAB14, HKB^{+}19, IHD^{+}18, JCRS09, KPHH12, KGAM18, KOJL^{+}14, KLG^{+}16, KMLM16, LKJ^{+}05, LTM18, LCS^{+}12, LD11b, MRO^{+}12, MRSS^{+}12, MGMP18, NBW14, OM09, PLC^{+}11b, PLE^{+}18, PFK^{+}08, SZB^{+}09, SNLD06, SMM12, TKE16, WBJ16, XRP^{+}12, YHR^{+}19, dLVvL06, RSRDS16]. Studying [BGC^{+}11, HZM^{+}16, RKSB13]. Style [ARB07, CH03, DFG^{+}14, HA18, KBD^{+}11, LTP^{+}05, MSSH14, MTW^{+}12, TSLR07, ZJX^{+}15, CDK^{+}17]. Style-Aware [ZJX^{+}15]. Style-Preserving [MSSH14]. Styles [CCM11, SJM14]. Styling [KZW^{+}16, WBK^{+}07]. Stylistic [vPBB^{+}10]. Stylization [HLYL18, KISE14, KKSE17, yKL12, KCWI13, LXT18]. Stylized [CL06, YCLL08]. Sub [WHL16]. Sub-Pixel [WHL16]. Subdivision [BDHJ04, CMS06, CM10, CWQ^{+}07, GGLQ19, KMDZ10, LRZM11, LHFY12, MS08, MQV00, NO97, PPT^{+}11, PP09, QMV98, WQS07, ZC06]. Subdivision-Based [LHFY12]. Subdivision-Surface [BDHJ04]. Subgraphs [MJ09]. Subgroup [DvVH^{+}19]. Subjective [AL06, MIO^{+}15]. Subliminal [BL15]. Subneighborhoods [FYP10]. Subset [YS17]. Subsets [GGL^{+}14a, LPK^{+}13]. Subspace [CZZ17b, KJ12, WX17, YXG^{+}13, YRWG13, WM18]. Subspaces [WM18]. Substitope [BLS04]. Substrates [SA06]. Substructures [Hu16b]. Subsurface [AWB11, CCB^{+}18, HFG^{+}12, NSS14, SAM^{+}07]. Subtle [LDFZ14]. Subtyping [GNDV^{+}18]. Subway [AABS^{+}14]. Success [KHE09]. Suggested [BEDF16]. Suggesting [ZAM11]. Suggestions [DC17, Koo08]. Suggestive [CGH^{+}19]. Sugiyama [Bac07]. Suite [SM04]. Summaries [FFB18, TSH^{+}14, WPS^{+}09]. Summarization [DZL^{+}14, GXZ^{+}18, MI13, PKG12]. Summarization-Based [DZL^{+}14]. Summarizing [BFL06]. Summary [BMW17, CXR18]. Summed [XHL18]. Super [CMF^{+}18, JNK19, RZNS04]. Super-Multiview [CMF^{+}18]. Super-Voxels [JNK19]. Supercluster [MQF06]. Supercomputing [EGH^{+}06].

Superconductor [GPP^{+}16]. Supercubes [WD09]. Superellipsoid [JKM06]. Superellipsoid-based [JKM06]. Superfluid [GLX^{+}18]. Superimpose [SSI99]. Superman [PLE^{+}18]. SuperMatching [CCM^{+}13b]. Superpixels [PZLZ17]. Superquadric [SK10]. Supersymmetric [CCM^{+}13b]. Supervised [KW11, WFC^{+}18]. Supervision [KEV^{+}18]. Supine [NMGK17, ZMG^{+}10]. Support [DBD18, DFD^{+}14, HJLH19, Hu16b, HTC09, KWS^{+}14, MTRP10, NHB^{+}17, PMCS11, PTMB09, STM17, SRCP02, SRCP03, SOL^{+}16, TAE^{+}11, TFJ12, WKCB07, WLW^{+}18, WQZ^{+}18, XGS^{+}19, YDC^{+}14, ZCD19]. Support-Free [WLW^{+}18, WQZ^{+}18]. Support-Induced [Hu16b]. Supported [NJJ11, ZHLR14]. Supporters [Ano12m]. Supporting [Ano11g, HSCW13, HTL13, HMSA08, KSL^{+}17, KPS16, LFW^{+}19, LBS^{+}19, MT14, MGPH06, PSPM15, RAL^{+}17, SVGR16, WPS^{+}09, ZGI^{+}18, vHP09]. Supports [BMST97, IDA^{+}14]. Suppression [BL15, SFC^{+}07]. Surface [AG16a, AG17, APS^{+}14, AMB^{+}13, BH07, BRT12, BMR^{+}99, BDHJ04, BK12, BAB^{+}18, Bon98, BS08, BB19, BN12, BFTW09, BKW10, CGL^{+}17, CWQ^{+}07, CG07, CG08, CPG09, DSKA19, EDvW19, GWBO12, GR04, GSG96, GCZL14, GH95, Gué99, GXW^{+}18a, GXW^{+}18b, HBG11, HAT^{+}00, HKG07, HCP^{+}16, HYB^{+}17, HPC^{+}13, IZM18, JR07, JKLG08, KBE09, KFS^{+}17, KY06, LLLN^{+}14, LBG^{+}08, LGQ09, LLWQ13, LPS^{+}13, LQLX14, LB17, LBH14, LLRR08, MS08, MW99, MDHB^{+}07, MDB18, NWI17, NKH11, OJ12, OMD^{+}12, OKI15, ORC07, PB13, PBL10, QLLM13, RYL^{+}18, RvWT08, RE14, RL08, SFH06, SSH14, SAR96, SB04, SO17, SF04, SSG16, TC09a, TWSM^{+}11, UBH19, VP04a, WWC^{+}14, WYL^{+}19, WPG05, YKL^{+}08, YYFX18, YYD^{+}19, YSS^{+}12, ZG12, ZWW^{+}12, ZGHG11, ZKK02].

Surface-Based [BRT12]. Surfaces [AT05, ABCO^{+}03, BHW06, BHS12, BWF^{+}10, BEHP04, BBK07, BS16, CGD97, CLCQ12, CFM^{+}13, CRT04, CRH05, DPR00, DBTH07, FT13, FBTW10, GKT^{+}08, Gor02, GR04, GTLH01, HB03, HDBC15, HGH^{+}10, IFP97, IYS13, JSG03, KCOY03, KLMA10, KSS09, KTCG17, KMDZ10, KGPS13, KGJ09, KLS^{+}18, LvWJH04, LVRL06, LLLF08, LTWH08, LBG^{+}08, LRZM11, LBZ^{+}11, LHZ^{+}04, MQV00, MGKH09, MS18b, MDG00, MWK^{+}08, MTB18, NWHWD16, NPPZ12, NHPN14, PZ11, PYW^{+}16, PLC^{+}11a, PPT^{+}11, PMCS11, QMV98, RBN^{+}19, RZHB^{+}08, RASS17, RSA^{+}19, SM09, SM11, STS10, SN10, Stü98, Tac10, TW18, TP12, TC17, USE13, VAW^{+}17, WWB^{+}13, WTW^{+}08, WSS09, WRT19, WB05, WDC08, WT10b, WKI^{+}17, YXSH13, YT02, ZHT07, ZTZ17, ZDW^{+}05, ZK14b, ZHGH11, vWC06, vFWTS08]. Surfel [YYR17]. Surgery [BTB^{+}04, CDA99, HCP^{+}15, KOJC12, LRF^{+}11, NWF^{+}05, RGF^{+}04, SSH14, SLK^{+}17a, WZW^{+}05, TGSP09]. Surgical [KSS09, KSI^{+}96, MTRP10, SLK^{+}17a, STH13]. Surprise [CH17]. Surveillance [MI13]. Survey [BNPB13b, BN12, CZ11, CAN14, COCSD03, DLR09, FHKM17, GIMS18, Han16, HMM00, HKPC19, HQC^{+}19, JBS06, KH13, LWC^{+}19, MUS16, Mor13, POM^{+}09, RBS^{+}18, SHS11a, SS06b, SSG12, TCL^{+}13, WHLS19, WBK^{+}07, dOL03]. Surveyor [ADG11]. SurVis [BKW16]. Survival [MMB^{+}19]. Suspicious [GRS^{+}19]. Suturing [BTB^{+}04]. Swap [PDBG18]. Sweep [CL18, IMS15, KGPS13, SM97, SYR11]. Sweeping [GMD^{+}16, SP96]. Sweeps [vRKEE17]. Swirling [WSTH07]. Switchable [XGS^{+}19]. Switching [GMS19]. Symmetric [CSPN11, DWF^{+}19, HLL97, HVSW11, JKM06, PYW^{+}16, RKZZ19, SK10].

Symmetry

[LJZ12, PZ11, TN11, TN13, TN14].

Symposium

[AD12, BvdP12, BDC17, BHTY15, CW11, Ert10c, GKR14, GJK15, GW13, HK10, HVY16, HKQ13, JLS15, KS14a, KHSB11, KL14b, LST^{+}16, LBKD09, LABS10, MH10, MY14, MFS^{+}09, OA11, Ota17, PZ12, SK16a, SK15, SW17, TL11, VW12, WLW17, vW11, CCH14, DFQ12, NSvW11]. Synchronization [HBKS09, LLL06]. Synchronized [KP05]. Synchronous [BE09]. SynCoPation [RSM^{+}16]. Synopsis [CXR18, NXSL13]. Synteny [MMP09, BGM^{+}17]. Synthesis [AS98, BJEYLW01, BB19, CDR^{+}18, CTM^{+}13, DNL^{+}06, DWK^{+}16, FvdPT97, FXG12, FR13, FCSF17, GIS03, HCP^{+}16, HKYM17, KCPS08, LPKK12, LHC16, LFP07, LBZ^{+}11, LHZ^{+}04, LCS^{+}12, LJWF12, LSS^{+}15, MCP^{+}06, MK13a, MM11, NTS11, RB07, RNE^{+}17, RSM^{+}16, SLGM19, WWYS04, WLD^{+}19, WWY14, XZX^{+}17, YL18, ZLG^{+}06, ZHQ^{+}07, ZSTR07, ZKG07]. Synthesis-Coupled [RSM^{+}16]. Synthesized [YYSZ06]. Synthesizing [MSSH14, VML97]. Synthetic [ALM11, ZLD^{+}14]. System [AK02, AvHK06, BDF^{+}10, CL06, CDW^{+}16, DRHK07, DH02, EGS03, FCZ15, GCML06, HF06, IG19, INCB18, aKS12, KNR17, KKKW05, Lac96, LBS14, LDC96, LRF^{+}11, MGJH08, MYI13, MFS^{+}09, NMN^{+}18, NB95, NLKH12, NQX^{+}05, NT99, PKS^{+}08, PWG17, PMD^{+}07, RHJ^{+}16, RBGH14, RNK^{+}15, SST^{+}17, SW13, SWY^{+}17, SvLF10, SYS^{+}06, STH02, TKTN09, TMH^{+}10, TIK15, TGW^{+}95, TLM05, WK13, WTL^{+}09, WDW16, WCS^{+}18, ZFA^{+}14, vLFR17]. Systematic [AAGS19, FIBK17, IIC^{+}13, PS06, WCR^{+}18]. Systems [ATK16, BSS^{+}19, BKDE00, BS02, CSL^{+}16, CCQ^{+}14, CLB^{+}16, CFEC17, DRHK07, DNN13, EG09, FWL17, GPK14, GS06, GBCG^{+}14, HKC^{+}12, HBC12, MGS^{+}14, MNKW07, NKHC08, NSS03, OWS15, PJ03, PSM12, PFK^{+}08, RK17, SSMG13, SSG12, Sim07, SBHW11, SLW^{+}10, TYSN06, YNYH06, YFZ^{+}18, YON05, YON06, ZLK^{+}18, vHvdWvW02, vPVvdW10].

T [Ano14f, KSY14]. T-ReX [KSY14]. Table [Ano10f, Ano11k, Ano12n, Ano12o, Ano13s, Ano14q, Ano14r, Ano15o, Ano16s, Ano16t, Ano19k, LRF^{+}11, RBLW07, WLS^{+}18, CLS13a, NSL19]. Tables [BLE19, NSH^{+}18, XHL18]. Tabletop [EASB19, IFP^{+}12, LH11, MBW^{+}07]. Tabletops [IIS14]. Tabular [GGL^{+}14a, HTL13, PDF14, WCC^{+}18, YEB18, YS17]. TACO [NSH^{+}18]. Tactile [BIAI17, YNYH06]. Tactile/Tangible [BIAI17]. Tag [LRKC10, RCSJ18]. tagged [CYW^{+}16]. Tailored [BLO^{+}05, HCMTH15]. Tails [SSS13]. Takes [Nie95]. Taking [FFB18, GS16, Seq12, YAE07]. Talk [SCB^{+}19]. Tangent [NJ99, ONL^{+}12, WT10a]. Tangential [VB13]. TanGeoMS [TMH^{+}10]. Tangible [BSB^{+}18, BIAI17, HF10, HJC14, LPG^{+}18, TMH^{+}10]. Tank [ZHF^{+}19]. Targets [SSE15]. TargetVue [CSL^{+}16]. Task [wAPS14, AG16b, BLIC19, DBP14, GHL18, GPC^{+}17, GLB16, JA18, KKC15, LFP07, LCS^{+}12, LWC^{+}19, LDA12, RKSB13, RBK^{+}15, SED19, TKAM06, BPC^{+}10]. Task-Based [BLIC19, LFP07, SED19]. Task-Driven [AG16b, LWC^{+}19]. Task-Overlapped [GPC^{+}17]. Tasks [BM13, CPG^{+}15, ERLW18, FM06, FKS16, GJC^{+}17, aKS12, KPR^{+}14, LTM18, LRM^{+}13, LB17, MCG12, MKT^{+}18, MLS18, NA19, OIR^{+}17, PTM^{+}18, RBS^{+}18, RFSP19, SNHS13, VBV^{+}18, WBK^{+}08]. Taste [KJH^{+}18]. Taxi [ADWK^{+}17, FPV^{+}13, HZM^{+}16, LWL^{+}17]. Taxonomy [wAPS14, ED07, ET08, KKC15, KOJC12, KCWI13, MRSS^{+}12, Rot13]. Taxonomy-Based [MRSS^{+}12]. Taylor [MMMY97]. Teaching [BGM^{+}17, KSI^{+}96, RRJH18]. Team [SJL^{+}18b, WXW^{+}19]. Teaming [RCL^{+}15]. Teammates [RCL^{+}15]. Techinal [Ano10d]. Technical [Ano11d, Ano12h, Ano13j, Ano13k, Ano13d, Ano13f, Ano14k, Ano14l, Ano14e, Ano14g, Ano14f, Ano16k, Ano16l, Ano16c, Ano16e, Bil13, Ebe17, Hee18, Sch13, Sil17a, Sil17b, Sil18a, Sil18b, Sil19, Ynn19]. Technique [ARRC11, Ano05d, BS95, BE06, CLB^{+}16, CM08, HBAB14, HHH16, JS98, KMM^{+}13, KK19, KSW06, MRS^{+}13, MIO^{+}15, PNML08, SMDS14, SMT13, SKP07, SFR^{+}10, TNS10, WWZ^{+}18, WSM^{+}09, WFM^{+}12, YON05, ZLB^{+}05, ZZG^{+}12, vFWTS08]. Techniques [AB01, BW03, BF01, BJM07, BL07, BMY05, BPS^{+}11, CFEC17, DDGL07, DRMM13, EF10, FWL17, JBS06, JCRS09, Kas12, KZL07, Kei00, KCWI13, MCG12, MJ09, NA19, NT03, PJ03, PFW09, RBS^{+}18, SPCJL06, SDMT16, SBJ^{+}10, SKL^{+}14, SOK^{+}16a, SOK^{+}16b, VGKS12, WBJ16, WEE03, YEII12, ZWBH13, JSV^{+}08]. Technology [Ano12g, BSS18, BRS18, CGGR18, Ert10c, GGHZ19, KJH^{+}18, MW13, TL11, WLW17, vW11, HK10, Min13]. Teens [ANR^{+}18]. Teichmüller [JZLG09]. Telco [WXZ^{+}16]. TelCoVis [WXZ^{+}16]. Tele [Ohl18]. Tele-Immersion [Ohl18]. Teleoperation [CWC^{+}06]. Telepresence [BKKF13, KNR17, LWG05, OWS15, PLE^{+}18, RZP^{+}07, SKH^{+}19, YLC^{+}19, ZLK^{+}18]. Telepresentation [YNYH06]. Telexistence [TAK^{+}05]. Tell [KDX^{+}12]. Telling [LKS13a, SH10]. Temperature [KLySK12]. Template [LG13, YLSL11]. Template-Based [YLSL11]. Templates [HA18]. Temporal [AAFW17, BFP14, BSH^{+}16, BJB^{+}12, BAB^{+}18, BMW17, CXR18, CHW^{+}18, DMAM19, DSP^{+}17, FPV^{+}13, FSE12, GJG^{+}15, GBP07, GHA^{+}08, GS14, GBFM16, GDKB17, HSCW13, HHB16, HSR13a, JFSK16, KKC15, KSS09, KW19, KPS16, KRRW19, LS09, Lin16b, LSS09, LXR19, LPK^{+}16, LYK^{+}12, MTM^{+}16, MLL^{+}13, SM17, SFP^{+}19, SBE^{+}15, SLQW17, SPW02, TIS16, TNT17, WPS^{+}09, WG16, WG12, WS09, WXW^{+}19, YGFX19, ZCD19, vLBR^{+}16, BMLC19, WDSC07]. Temporally [CRH05, MVN^{+}19, WLHL13]. Tennis [PYHZ14, RBLW07, WLS^{+}18]. TenniVis [PYHZ14]. Tensile [SVAC12]. Tensor [AWHS16, BRSP18, BRP19, CYZ^{+}09, DGBW09, FA15, HLL97, HVSW11, JKM06, KWH00, KW06, PYW^{+}16, PLC^{+}11a, RBN^{+}19, RKZZ19, RC06, SS08, STS10, SGM^{+}11, TKW08, WL16, WKB^{+}13, WXC^{+}08, YRP18, ZDL03, ZHT07, ZYLL09, ZSL^{+}16, ZPP05]. TensorFlow [WSW^{+}18]. Tensors [AS98, GRT17, OJ12, SK10]. Terascale [TBR^{+}12]. Term [LD11b]. Terrain [BKRE19, BS16, CE01, CCB^{+}18, GY02, KLJ^{+}09, LP02, Ste98, WSD^{+}13, WBP07, WCB^{+}12, ZSTR07]. Terrain-like [BS16]. Territorial [vGMSW15]. Tessellated [SPM^{+}13]. Tessellation [LSPW12, RLW^{+}11, YW16]. Test [JCWD14, KWP01, KYK11, LKK17, SVAC12]. Testbed [KMSB14]. Testing [WM19]. Tests [HFMC12, WLSW08]. Tetrahedra [SCT06]. Tetrahedral [BLW14, CDM^{+}04, GW06, GH95, KL96, KTCG17, LHBF19, NJ99, THJ99, VCL^{+}07, WFKH07, ZG06, ZCW19]. Tetrahedralization [Mal05]. Text [AMJ^{+}12, BLE19, CSL^{+}10, CLC^{+}15, CG08, CLT^{+}11, CLWW14, DFG^{+}14, DYW^{+}13, FPB17, FBM16, GUFM15, GBWI17, HKBE12, KJW^{+}14, LB19, LYW^{+}16, LWC^{+}19, LYK^{+}12, PKL^{+}18, SOK^{+}16a, SOK^{+}16b, SSBC19, XLZ^{+}19a, XLZ^{+}19b, YFZ^{+}18, vHWV09]. TextFlow [CLT^{+}11]. Textile [ZCF^{+}19]. Texton [War09]. Texts [BW17]. TextTile [FPB17]. Textural [HLYL18]. Texture [BW08a, BJEYLW01, BMR01, BBK07, CJR07, CR08, CDR^{+}18, DT10, DWK^{+}16, EWWL98, FST^{+}14, GIS03, GO15, GK95, HAT^{+}00, HPC^{+}13, HDJ05, IFP97, JWD^{+}14, KSNY17, KHSI04, KBH13, LLPP19, LYY08, LGY19, LHZ^{+}04, LJWF12, MZX15, PS12, RB07, SLW^{+}10, WWYS04, WEE03, WSE07, WDC08, YNBH11, YLY^{+}12, ZDW^{+}05, ZKK02]. Texture-Aware [DWK^{+}16]. Texture-Based [HPC^{+}13, WEE03, WSE07, CJR07, CR08]. Texture-Mapped [PS12]. Textured [APS^{+}14, IYS13, UBH19]. Textureless [SPP^{+}14]. Textures [BS16, HE99, HHM14, JEH02, KXW^{+}18, LA11, Ney98, QY07, Sel15, TDN^{+}12, WWY14]. Texturing [BHW06, BH07, BMPB08, CDK04, DHM13b, FH06, KAK^{+}07, MNZ^{+}15, RDB^{+}12]. Thaumatrope [SC15]. Theatres [LLKN17]. Their [HQ13, MSvG^{+}11, OHJ^{+}11, YLZ^{+}13, BOZ^{+}14, BLG^{+}16, FNM13, PFP^{+}11]. Thematic [CH17, HHWN02, VJN^{+}15]. Theme [Nie95, Nie96]. ThemeDelta [GJG^{+}15]. ThemeRiver [HHWN02]. Theoretic [CGJM19, XLS10, CJ10]. Theoretical [BTS^{+}18, CFEC17, LNS08]. Theory [BEJK12, DPW^{+}15, Kei00, LLT04, MY96, NBM19, SJL^{+}18a, STS10, SZHR11, TTR10, VFR13, ZWJZ12]. Therapy [MMB^{+}19, ZZSS10]. There [WWZ^{+}18]. Thermal [KJH^{+}18, SGAS16]. ThermalPlot [SGAS16]. Thermo [IAS19]. Thermo-Visual [IAS19]. Thermoforming [ZTZ17]. Thickness [GLH^{+}14, LGV^{+}16]. Thin [ELF13, GLB^{+}06, GQGP17, HNR^{+}06, SH12, VAW^{+}17, WKZL04, YNYH06, ZQS11]. Thin-Shell [GLB^{+}06, ZQS11]. Thinker [RKKF19]. Thinking [WCR^{+}11]. Thinning [IYK01]. Third [Kas12]. Thomas [EDK10]. Those [RJD^{+}07]. Thousands [WLMP19].

Thread [MMYK06, WKZL04]. Thread-like [MMYK06]. Threading [PTMB09]. Threat [PDBG18]. Threats [MKN^{+}07]. Three [ADP02, BSL^{+}14, CCAL12, EHS13, FZC^{+}07, GNP^{+}06, GNPH07, HEWK03, HNR^{+}06, HRD^{+}19, MBS^{+}04, NE04, RSFH14, SLB04]. Three-Dimensional [ADP02, BSL^{+}14, CCAL12, EHS13, GNP^{+}06, HEWK03, HNR^{+}06, MBS^{+}04, NE04, SLB04, FZC^{+}07, GNPH07]. Threshold [SHV^{+}18, TWBBM17, WTB^{+}19]. Thresholds [SBJ^{+}10, WP19, ZLK^{+}18]. Thumbnail [WSYM17]. Thumbnails [YDK^{+}18]. Thumbsticks [YFZ^{+}18]. tially [BJM07]. Tick [TLH10]. Tight [KZW^{+}16]. Tight-Fitting [KZW^{+}16]. Tile [HDJ05, NC07]. Tile-based [NC07]. Tileable [LPF^{+}07]. Tiled [BI12, ETO^{+}10, LSJ^{+}15, NHN07, RLM10, SM11, tCMR07, ETO^{+}10]. Tiles [HCP^{+}16]. Tiling [LLLF08]. Time [AL11, AMM^{+}08, Ano14h, ASMP17, BSH^{+}16, BGT12, BRT12, BTB^{+}04, BGM^{+}07, BGM^{+}08, BLLS17, BAF^{+}13, BTB10, BTH^{+}13, BLIC19, BJC^{+}19, CLS^{+}12, CKW^{+}12, CMF^{+}18, CLC^{+}15, CK05b, ČB19, CORLS96, CMPC06, CDA99, CSWP18, CMP14, DAW13, DS16b, EHH^{+}19, FWSL12, FKRW17, FSHH12, FS19, GKT^{+}08, GSS^{+}15, GSCI15, GMD13, GSL^{+}17, GTPB19, GABJ08, GSDJ04, GWBO12, GB08b, GW11, GWP^{+}16, GGPPS13, GT17, GHP^{+}16, HMTR19, HKF16, Har16, HE06, HB14, HLRC^{+}12, HRISI15, HDJ05, IBJ^{+}14, JEG12, JME10, JS06, JY17, KRHH11, KL96, KLK^{+}09, Kin10, KC14, KKSM19, KLL12, KMDZ10, KGJ09, KGG^{+}12, KDA^{+}09, KBK11, KG06, LBG^{+}16, LCR16, LDSM17, LS02, LKC09a, LKC09b, LS09, LKR^{+}18, LCZ^{+}19, LKS13b, LXT18, LMC02, MB03, MGMP18, MLKS18, MDL^{+}19, MT01, NSW^{+}17, NQX^{+}05, NSH^{+}18, OBJ16, OKI15, PSSC17, PD04, PMP10]. Time [PSR17, Per95, PKMR15, QYH^{+}18, RGF^{+}04, RHJ^{+}16, RS12, RYL^{+}18, RKSH11, RLA^{+}13, RHD^{+}06, RASS17, RBLW07, SBV^{+}11, SBS16, SOS^{+}17, SK16b, SKP07, SK98, SCL^{+}12, SYK^{+}18, SN10, SKU^{+}12, SJH^{+}07, SVGR16, SB06, SGAS16, SSGM19, SKH^{+}19, SAC^{+}08, SSR^{+}07, SHR^{+}11, SH00b, TIW^{+}19, TCM06, TWHS05, TWBBM17, TLC^{+}10, THV^{+}14, USE13, VMT06, WRM^{+}10, WFKH07, WBJ16, WYM08, WTW^{+}08, WC09, WC10, WFW^{+}17, WHZ^{+}18, WCJ06, WSE07, WXY17, WS06b, WS09, WY19a, XESV97, XCH^{+}14, YML^{+}17, YLK12, ZQS11, ZHX^{+}11, ZWW^{+}12, ZDM13, ZX18, ZCPB11, cKJG^{+}12, dRBS^{+}12, vdZCT16, CJR07, FSME14, GCML06, KBD^{+}11, KW13, SFA^{+}15, WSY07, XCZ^{+}19b]. Time-Dependent [GKT^{+}08, HMTR19, KRHH11, KL96, KGG^{+}12, SBV^{+}11, TWHS05, USE13]. Time-Discrete [RS12]. Time-Domain [LS02]. Time-Hierarchical [FKRW17]. Time-Octree [WC09, WC10]. Time-Oriented [AMM^{+}08, RLA^{+}13]. Time-Series [GSL^{+}17, KLK^{+}09, KBK11, SGAS16, WBJ16, XCH^{+}14, ZCPB11]. Time-Varying [Ano14h, BRT12, BLLS17, CKW^{+}12, CLC^{+}15, CMP14, FWSL12, GABJ08, GSDJ04, GW11, GWP^{+}16, Har16, HE06, JEG12, JS06, JY17, KGJ09, KG06, LS09, LMC02, OBJ16, PSR17, SK16b, SK98, SB06, SSR^{+}07, SH00b, TIW^{+}19, WFKH07, WYM08, WC09, WC10, WCJ06, YLK12, DS16b, KC14, NSW^{+}17, KBD^{+}11, CJR07]. TimeBench [RLA^{+}13]. Timeline [LZW^{+}13]. TimeLineCurator [FBM16]. Timelines [BLB^{+}17, FBM16]. TimeNotes [WBJ16]. Timepoint [BBP08]. Times [CDW^{+}16]. TimeSeer [DAW13]. TimeSpan [LPK^{+}16]. Tissue [HCP^{+}15, SWB^{+}00, SAM^{+}07, TYL^{+}18]. Tissues [CDA99]. Title [Ano12p, Ano11i, Ano14j]. TOD [GPC^{+}17]. TOD-Tree [GPC^{+}17]. Together [BRNB19, ZCJH12]. Tokens [HJC14]. Toleranced [Gué99]. Tomography [AHR^{+}11, WAG^{+}12]. Tone [BKA^{+}11, FYWY16, LRP97, SJB10]. Tone-Mapped [FYWY16]. Tone-Mapping [BKA^{+}11]. Tongue [YGV^{+}13]. Tool [BISM14, CXM19, CRI06, CGB^{+}13, FPB17, HEFR18, IDA^{+}14, KGS^{+}08, LLL^{+}19b, MLMF12, MHD^{+}18, MAAB^{+}18, NHYY18, PBO^{+}14, SZD^{+}10, SD12, SKW^{+}11, SGPR18, SGB^{+}19, UKF^{+}18, YYT16, WOCH09]. Tool-Hand [CGB^{+}13]. Toolkit [BDK98, BH09, CPW^{+}15, HA17, SLC^{+}19, TFL^{+}18, MTRP10]. Tools [BCB10, BDFM17, BYB^{+}13, FH06, HKBR^{+}14, NJ99, RBS^{+}18, SPS06, SH12]. Tooth [XLZ19c]. Top [LS10, LFR03, SOR^{+}09]. Top-Down [LFR03, LS10]. Topic [AG16b, CLRP13, DYW^{+}13, EASS^{+}18, EASD^{+}19, GJG^{+}15, GNDV^{+}18, KKP^{+}17, LWLM18, SWL^{+}14b, XWW^{+}13, ZGW^{+}19]. TopicLens [KKP^{+}17]. TopicPanorama [WLL^{+}16]. Topics [CLT^{+}11, CLWW14, WLL^{+}16]. TopKube [MLKS18]. TopoAngler [BDSS18]. Topographic [CDM^{+}06]. TopoLayout [AMA07]. Topological [AMA07, AJ19, BEHP04, CLB11, DFD^{+}14, DCK^{+}12, GKN05, GJR^{+}14, GNP^{+}06, KCA16, MBS^{+}04, OHWS13, RML12, RFLL18, SJJ^{+}17, STS07, SDHH12, SPCJL06, TL07, TWHS05, TP12, TKW08, TGS11, UMW^{+}12, WBP07, ZPP05]. Topologically [FA15, GDN^{+}07, SS13b]. Topologically-Informed [FA15]. Topologies [RE14]. Topology [AHK^{+}17, BDSS18, BDD^{+}16, BWT^{+}11, CL11, ESV98, ENS^{+}12, HHVW96, HLL97, LLT04, LBG^{+}08, LJWH08, LFR03, MLCM16, MKW07, OHJ^{+}11, PBC17, RLH11, RKZZ19, SKMR98, SPN^{+}16, SB06, TN11, USE13, WXJD17, WDC^{+}07, WJE01, WLZM10, WDW16, ZC06, ZJH07, TFL^{+}18].

Topology-Aware [WLZM10].

Topology-Based [BDSS18, BWT^{+}11]. Topology-Controlled [WDC^{+}07]. Topology-Preserving [BDD^{+}16, WXJD17, WJE01]. TORNADO [TT05]. Toroidal [SCT^{+}10]. Torso [SMTT^{+}17]. Total [ZWZD15]. Touch [EASB19, LODI16, LRF^{+}11, WLJ^{+}12, XST^{+}18, YSI^{+}10, ZZD14]. Tourism [NM13]. Tourism-Inspired [NM13]. Tournament [TSLR07]. Tournament-Style [TSLR07]. TPFlow [LXR19]. Traced [WHL16]. Traceless [JKM06]. Traceroutes [CDDS18]. Traces [AAM^{+}12, IBJ^{+}14]. Tracing [BWW^{+}12, CLS^{+}12, CLT^{+}08, COJ15, GWE^{+}19, GHS^{+}19, IHR01, KL96, MAST16, MYM16, MGW10, NKP^{+}15, NO97, NK06, PPL^{+}99, QCH^{+}14, SYK^{+}18, SN10, SK00, Stü98, TCM^{+}12, UKF^{+}18, WFM^{+}05, WFKH07, WJA^{+}17, WWW^{+}19, WSC^{+}95, ZGH^{+}18, LCM07]. Track [DNN13, NTT^{+}19, OKB^{+}19]. Trackballs [HSH04]. Tracked [HWHK16, OIR^{+}17]. Tracker [KDX^{+}12]. Tracking [AABH^{+}16, AJ17, BDF16, BWS^{+}19, BVPtHR09, BWP^{+}10, BC12, BKH^{+}11, CDF14, CMN13, CMPC06, DH08, DS16b, GHE19, GL17, GPP^{+}16, GXW^{+}18a, GXW^{+}18b, GLB16, IDA^{+}14, JER16, JA18, KW13, KHH^{+}16, KHSW17, LH09, LKC09b, LG12, LWW^{+}13, MWC^{+}12, MVN^{+}19, NBW14, OKI15, PLW11, PLW12, PSR17, PKMR15, RKSB13, RKK16, SGQ16, SYM14, SPP^{+}14, SW97, TNT17, THV^{+}14, VMN^{+}19, WRM^{+}10, Wu16, YNM15, YYD^{+}19, CJR07]. Tracks [LPCRH19]. Tract [EBB^{+}15]. Tractography [PFK07]. Tracts [JDL09, MSE^{+}06]. Trade [HKB^{+}19, RPHI08, WLS^{+}19]. Trade-off [HKB^{+}19]. Traditional [BKH^{+}11, CB15, Chi16]. Traffic [AAFG18, CDR^{+}18, LLB^{+}12, MV06, MKN^{+}07, SHVV16, SvdBLM11, WZvdW13, WYL^{+}14, WSLL12]. Trail [HEF^{+}14, HPNT18]. Trainable [OK11]. Training [AGN^{+}19, BBG^{+}18, BC18b, CPG^{+}15, DRHK07, FTES13, GS16, HKBE12, JPD^{+}18, LLQ^{+}17, LSC^{+}18, MTRP10, MLS18, RBK^{+}15, RSBB17, SSH14, SB14]. Traits [FDC^{+}18, UKW19]. Trajectories [ADWK^{+}17, AAFG18, COMP13, DHR^{+}19, HTC09, HRD^{+}19, JM10, LWL^{+}17, NBW14, SWvdW^{+}11, VJN^{+}15]. Trajectory [AdLH13, HZM^{+}16, KJW^{+}18, LLL06, MI13, SYM14, TSAA12, WZvdW13, WYL^{+}14]. Trajectory-Based [SYM14]. Trajectory-Preserving [LLL06]. TrajGraph [HZM^{+}16]. Transaction [KHDL07, LSS09]. Transactions [Ano14c, Ano14j, Ano15a, Ano16f, Ano16a, Ano17e, Ano18a, Ano18b, Ano19a, Ano19f, KHE09, XCH^{+}14]. TransCAIP [TKTN09]. Transcatheter [BSR^{+}14]. TransCut [LSR^{+}13]. Transfer [ACTM12, ADDG12, CR08, CPG^{+}15, CM08, CM11, EMRY02, KKH02, LWS^{+}17, LLL^{+}12, LLL^{+}10, LLY06, ME18, MWCE09, MJW^{+}13, MLS18, NHPN14, RSBB17, SKK06, SG09, Sel15, SHR^{+}11, WZK12, WQ07, XJF^{+}08, ZGI^{+}18, ZT09]. Transferring [PZ07]. Transfers [KHSS14]. Transfinite [SFA^{+}15]. Transform [FM07, HCMTH15, LS13b, SPB96, SCYW16, ES05]. Transformation [AAB^{+}13, FE17, HLCB18, LCP^{+}13, LYLG19, LDW^{+}15, PSKN06, WZ08]. Transformations [DSS^{+}09, LL05, MPK^{+}13, YHMY08]. Transforming [DW14, WSLL12]. Transforms [GGLQ19, SK00]. TransGraph [GW11]. Transition [AW03, GW11, MS18a, PV06, vHvdWvW02]. Transitional [JCG08]. Transitions [BFP14, BBG^{+}09, CDF14, HR07b, WASQ18]. Transitive [KL03]. Translating [NW15]. Translation [ZLK^{+}18]. Translucent [KSTE06, LSR^{+}13, NSS14, WLLC15]. Transmission [OHH06, YS03]. Transmissive [IHS17]. Transonic [CDL^{+}16, DCH^{+}17]. Transparency [BCS11, BS11, CFM^{+}13, CWM^{+}09a, ESSL11, HC05, MM17, SC15, Zha14]. Transparent [DK13, GQGP17, IFP97]. Transparently [FH07]. Transport [BB09, DT10, FT09, HOGJ13, DSC^{+}16, SPS06, SZN^{+}18, SKLU^{+}11, ZSG^{+}13]. Transportation [PGSF16, ZFA^{+}14]. Transpost [OHH06]. Travel [SFR^{+}10, ZLB^{+}05]. Traversal [GIK^{+}07, HL09]. Traversing [ZZM06]. Treatment [JDSR^{+}18, ZCD19]. Tree [BHP^{+}12, BKH^{+}11, CBLD11, GZ11, GPC^{+}17, HSF^{+}06, IPD^{+}07, IC07, LPP^{+}06, LXL^{+}18, QYH^{+}18, RH19, TSLR07, TdJ14, TdJ15, WXJD17, WV08, XYS^{+}16, YRWG13, ZGH^{+}18, ZT09, NSL19, TdJ14]. Tree-Parts [XYS^{+}16]. Tree-Structured [TdJ14, TdJ15]. Treelike [MK16]. Treemaps [BL07, GSWD18, KHA10, KW19, SSV18, TC13, TS07, TS08, VvWvdL06, WD08]. TreeNetViz [GZ11]. TreePlus [LPP^{+}06]. TreePOD [MLMP18]. Trees [AMA06, BSV11, CMM^{+}97, DN13, EG09, FDC^{+}18, GK07, HSCS11, KW19, MSvG^{+}11, MLMP18, PFP^{+}11, SN97, TFO09, WFM^{+}05, XXM19a, ZLD^{+}14, TGSP09]. TreeVersity2 [GGPPS13]. Trenches [SMM12]. Trend [GSL^{+}17, LS09, LD11a, RFF^{+}08, SKK^{+}14]. Trend-Centric [SKK^{+}14]. Trending [WOO17]. Trends [AAFW17, BJC^{+}19, KNKH19, KBB^{+}18, LRKC10, MOC^{+}14, OBJ16, WHZ^{+}18]. Tri [DDBR^{+}19]. Tri-modal [DDBR^{+}19]. Triangle [AFRS05, GHJ^{+}98, HH10, HHQH17, KWH19, LXB17, LGLR14, MSE^{+}06, Ros99, VP04b, WZW^{+}05, YPI13, eYL07]. Triangles [Far12, ZK12]. Triangular [CCS12, Del08, GSG96, HTF97, LFR03, Stü98, VCP08]. Triangulated [BG04, BEHP04, GH95, WDC08]. Triangulation [FHSW13, GCT17, MS18b, QCT13, YSS^{+}12]. Triangulation-Invariant [YSS^{+}12]. Triangulations [HB03]. Tricubic [KCOY03]. Trilinear [AS98, CM10, HTF97]. Trimmed [Stü98]. Trimming [HR07a, SF19]. TripAdvisor [NM13]. Trips [FPV^{+}13]. Trivariate [RE01a, WLL^{+}12]. Tropical [LPCRH19]. True [HV13]. Trumps [SOR^{+}09]. Truncated [NDS10]. Trust [DLW^{+}17, SSK^{+}16]. Truth [KDX^{+}12]. Truthful [MK13c]. Truths [LB17]. Try [HSR13b]. Try-On [HSR13b]. tSNE [PLvdM^{+}17]. Tubes [STM08]. Tuboids [PFK07]. Tubular [AH11, SMG^{+}13, WKB^{+}13]. Tugging [AMA11]. Tumor [KKMS11]. tunable [RWF19]. Tuner [TWSM^{+}11]. Tuning [BSWL12, LLL^{+}12, MGJ^{+}10]. Tunnel [ZZM06]. Tunnels [BLG^{+}16]. Turbine [SOL^{+}13]. Turbulence [KLySK12, LPQF14, TBR^{+}12]. Turbulent [BKKW19, HPAW07, LBM^{+}06, LSY^{+}18, SW97]. Turnaround [NB95]. turned [SBS16]. Turning [WLHD17]. TV [TKTN09]. TVCG [De 18b, Ano09a, Ano09c, Ano09e, Ano09d, Ano11l, Ano13a, Ano13b, CGGR18, GGHZ19]. Twitter [TKE16]. Two [CF10, GLX17, GGZ^{+}18, HMBG01, JDL12, KBE^{+}18, KRHH11, KBH13, KCPS08, LVRH07, LTKF08, MPT03, PSN10, VHLL14, WHK15, ZC06, HSKIH07]. Two-Character [KCPS08]. Two-Dimensional [JDL12, KRHH11]. Two-Level [HMBG01, LVRH07]. Two-Manifold [KBH13]. Two-Pass [MPT03]. Two-Phase [GLX17, KBE^{+}18, PSN10, ZC06]. Two-Stage [GGZ^{+}18]. Two-Way [LTKF08, VHLL14]. Type [LBH14, MS18a, ZCD19]. Typed [FMH08]. Types [BC18b, RESC16]. Typographic [AMJ^{+}12]. Typology [BM13, ERLW18].

Ubiquitous [BFE15, BMR^{+}19]. Ugly [BTS^{+}18]. Ultra [FSTG16, KWDG11, YDG^{+}16]. Ultra-Compact [YDG^{+}16]. Ultrasonic [PH07]. Ultrasound [HQS18, WKSS05, zBBKN14]. Unambiguous [BRH^{+}17]. Unaugmented [LBHW18]. Unbiased [YWW14]. Uncalibrated [SM09, WLDW11]. Uncertain [AJ19, FKLT10, GHP^{+}16, GHS^{+}19, LCP^{+}13, LPCRH19, PMW13, SE17, SNG^{+}17]. Uncertainty [AE13, ASE16, BBIF12, CZC^{+}15, FBW16, GSWD18, GBP^{+}13, GR04, GBFM16, GHL15, HLNW11, HKKS18, HQC^{+}19, LFLH07, LLZ^{+}16, LBR^{+}17, LLPY07, MRO^{+}12, PH11, PRH10, RWG^{+}12, SHM10, SSK^{+}16, SZB^{+}09, SZD^{+}10, SKS12, SDW11, SPB08, TK14, TvET14, WMK13, WPL96, WYM12]. Uncertainty-Aware [CZC^{+}15, LLZ^{+}16, PRH10]. Uncluttered [WJR^{+}13]. Uncluttering [FT09]. Unconstrained [THV^{+}14]. Uncover [GdBG12]. Uncovering [DBB10]. Underestimation [HUPS14]. Underexposed [ZNZX16]. Understand [GGZL16, LLB^{+}12, LWD^{+}17, WGYS18, WGSY19]. Understanding [AAB^{+}13, BCH^{+}13, BZS^{+}13, CLT^{+}11, DSC^{+}08, FLF^{+}11, GWP^{+}16, HDR^{+}13, KTC^{+}19, LBM^{+}06, LXC^{+}17, LDM^{+}18, LBLE^{+}14, LODI16, MRH^{+}10, MGJ^{+}10, MQB19, NXW^{+}16, NTA^{+}19, UDSL18, VFR13, WLJ^{+}12, WKCB07, WWS^{+}16, YaKSJ07, ZLC^{+}19]. Unfamiliar [LKH^{+}16]. Unicube [HWL^{+}11]. Unified [CM16, DVCD07, GLX17, KYK11, MFS^{+}09, RDB^{+}12, Sil95, TMWS13, WWS^{+}18, ZFSL19]. Uniform [BTV14, LLKN17, NO97, ONL^{+}12, Zhu05]. Union [DN13]. Unit [MGJ^{+}10, PDFE18]. Units [MB19a, tCMR08]. Universal [vdZCT16]. Universe [FH07, GB08a, MOC^{+}14]. Unobstructed [CGB^{+}13]. Unsteady [BS95, BMLC19, FC95, FPH^{+}08, GZL^{+}14, GHP^{+}16, GHS^{+}19, HLNW11, HLG^{+}14, JEH02, LM05, NTT^{+}19, SWCR15, WSTH07, WSE07, WGS07b]. Unstructured [BPL^{+}19, CICS05, CBPS06, CDM^{+}04, CMSW04, CHM11, EHH^{+}19, FPB17, FBM16, GJ10, Har16, KM96, MJK06, MEB^{+}14, MHDH07, RL08, USM96, USM97, WMS98]. Unsupervised [KEV^{+}18]. UnTangle [CLG16]. Untangling [DWvW12, RD10]. Untapped [HFM16]. Untrained [KNKH19]. Unusual [NTA^{+}19]. Unwrappable [FWZQ13]. Upper [MZH^{+}08]. Ups [VB18, LLLN^{+}14]. Upsampling [YGFX19]. UpSet [LGS^{+}14]. Urban [CWK^{+}07, CHW^{+}18, DZMQ16, DFD^{+}14, FPV^{+}13, HZM^{+}16, MDL^{+}17, MPBM^{+}18, OSS^{+}17, PY09, QWC^{+}09, SZY^{+}18, SWF^{+}16, VABW09, WXZ^{+}16]. Usability [GS08, KWL14, KLD^{+}09, Sim07]. Usage [SOK^{+}16a, SOK^{+}16b]. Use [BVW^{+}07, CFEC17, DPW^{+}15, HJC14, aKS12, LMK07, PGU^{+}13, RB11, TJW^{+}17, WZ08, WKB^{+}13]. Used [RZP12, WCR^{+}11]. Usefulness [KPR^{+}14]. User [BHHM19, BHST17, BTH^{+}13, BLO^{+}05, BGR06, CSL^{+}16, CLRP13, DC17, DLF^{+}09, EASD^{+}19, FCL09, FM06, GS08, GDJ^{+}13, GBFM16, HF10, HB13, HOG^{+}12, JJ09, KKKT18, KGAM18, KKW^{+}17, LKJ^{+}05, LKD19, LKS^{+}19, LWZQ17, LAK^{+}11, LGYG12, LCS^{+}12, MFZ^{+}17, MPG^{+}14, NBM19, NTA^{+}19, PLvdM^{+}17, PSM12, PBK^{+}12, PBC17, RGFLL14, RLM10, SKK06, SZB^{+}09, SS06b, SMP17, SPG14, XRP^{+}12, YEII16, YCHZ12, ZGB^{+}17, vHR08, GSL14]. User-Assisted [LAK^{+}11]. User-Authored [ZGB^{+}17]. User-Based [GS08]. User-Centered [GDJ^{+}13].

User-defined [JJ09]. User-Driven [CLRP13, SPG14]. User-Generated [DLF^{+}09, vHR08]. User-Guided [BTH^{+}13, MFZ^{+}17]. User-Perspective [BHST17]. User-Steerable [EASD^{+}19]. Users [BOZ^{+}14, FTES13, GMD^{+}17, HGWW18, LKD19, PPP12, SGQ16]. Uses [Sat13]. Using [AJDL08, AMJ^{+}12, AZM12, AHK^{+}17, Ano96b, ASMP17, AWB11, AERA14, AHL^{+}13, BHHM19, BA05, BSS^{+}13, BBC15, BB09, BDF16, BJEYLW01, BT13, BBD^{+}11, BSE^{+}17, BSKR19, BWC04, BJM07, BJB^{+}12, BWW^{+}17, BSO^{+}12, BARM^{+}12, BPM^{+}13, BWT^{+}11, BMPB08, BC12, BDW^{+}08, CGC^{+}11, CGD97, CvW18, CPW^{+}18, CDR^{+}18, CML^{+}07, CDS^{+}12, CCM^{+}13b, CVG13, CL06, CKLL09, CMM^{+}97, CML^{+}12, CCM12, CRB^{+}05, DK11a, DHL09, DRW16, DdL14, DWB^{+}06, DFD^{+}14, DYW^{+}13, DTT^{+}17, DKM06a, EG09, EJR^{+}14, FSHH12, FC95, FGF^{+}05, FT09, FG99, FMST96, GSA^{+}09, GMD13, GdBG12, GLM06, GUO00, GO15, GK95, GSG96, GGPPS13, GH95, GN12, GGZL16, GXW^{+}18a, GXW^{+}18b, GWK12, GSZ^{+}13, GKK^{+}12, HBJP12, HBG11, HTF97, HJW99a, HJW99b, HYFC14, HAGS16, HH10, HEWK03, HA04, HHM14, HRN^{+}03, Hu16a, HZM^{+}16, HR11, IHK05, IVJ12, IG19]. Using [IK95, IYK01, IYIK04, IHS17, IFM14, JK16, JNK19, JEG12, JWSK07, JBMS09, JFTW07, JWC05, JDA^{+}11, JFBB10, JM10, JFY16, JCRS09, KIL07, KSH03, KCJ^{+}10, KL96, KLM04, KL07, KJ12, KISE14, KHM^{+}98, KS01, KM96, KA12, KSDD14, LPF^{+}19, LH03, LYY08, LKC09a, LKL^{+}15, LKR^{+}18, LFP07, LGQ09, LBZ^{+}11, LDN11, LL05, LLY^{+}13, LB17, LLL^{+}10, LLRR08, Lin16b, LCS^{+}12, LGS12, LSJ96, LFR03, LPK^{+}16, LXT18, LSM03, LLPY07, LSS^{+}15, LXW^{+}18, LSK^{+}18, Mal05, MW99, MGJ^{+}10, MGM19, MDS16, MKW07, MMMY97, MM08, MZC^{+}16, NMGK17, NMN^{+}18, NWI17, NSW^{+}17, Ney98, NW10, NHN07, NT03, OHJ^{+}11, OK11, PBN^{+}13, PPT^{+}11, PKL^{+}18, PvdBC^{+}11, PPZ^{+}12, PW13, PGU12, QCT13, RNL09, RBK^{+}19, RS12, RvWT08, RNK^{+}15, RML12, Rob98, RHR16, RRJH18, RNE^{+}17, RL08, SCT06, SLGM09, SM09, SMO^{+}13, SE18, SSIF09, SG09]. Using [SBSG06, SW13, SXX^{+}19, SLG^{+}17, SR00, SK00, SLMA06, SRKL19, SB14, SGAS16, STH03, SRMOW11, SC15, SAC^{+}08, SRML09, TKTN09, TCYM09, TIW^{+}19, TTR10, TN13, TWMZ19, TZL^{+}12, TWSM^{+}11, TRd12, TLC^{+}10, VP09, VFR13, VJC09, WFM^{+}05, WDC^{+}18, WLL^{+}05, WGS07a, WL08, WC09, WC10, WCZ^{+}11, WFS^{+}16, WLHD17, WCJ06, Wen14, WBH04, WJR^{+}13, WHM14, WM19, XNT11, XF04, XA10, XYGL13, XLC^{+}18, XLZ^{+}19a, XLZ19c, YHMY08, YBZW14, YKL^{+}08, YSD^{+}17, YWV^{+}19, YCLJ12, YLSL11, YCLL08, YLG^{+}14, YFZ^{+}18, YON06, YCHZ12, YGFX19, ZMG^{+}10, ZC03, ZDL03, ZPS04, ZK07, ZCL08, ZZC11, ZHX^{+}11, ZDM13, ZGW^{+}14, ZK14a, ZWZD15, ZFSL19, ZDZ18, ZGI^{+}18, ZC06, ZJH07, ZT09, ZKK02, ZWBH13, ZHGH11, vPVvdW10, vRKEE17, AMB^{+}13, BMJK09, BL15, DMS^{+}08, EDF08, GHA^{+}08, HKG07, HBF08, HHG14, IBJ^{+}14, JR07, JKRY12, JQD^{+}08, KTC^{+}19, KGP^{+}13, LCM07, LYL19, LPK^{+}13, LPCRH19, LWP^{+}06]. using [MSE^{+}06, MGB^{+}19, NBM19, NTT^{+}19, RBN^{+}19, SHM10, SJ06, SFL^{+}16, TS07, UBH19, XLZ^{+}19b, YEB16, YLC^{+}19, ZSG^{+}13, vPBB^{+}10]. Usual [NTA^{+}19]. Utero [MMK^{+}17]. Utility [WCC^{+}18, ZYM^{+}14]. Utility-Aware [WCC^{+}18]. Utilizing [HFM16, ME09, NO97]. UTOPIAN [CLRP13].

VA [BJK^{+}16]. VAET [XCH^{+}14]. VAICo [SGB13]. VAiRoma [CDW^{+}16]. Validating [MP13]. Validation [JPD^{+}18, LBS13, Mun09, USKD12, YDC^{+}14, ZBG^{+}17]. Validations [MF11, QH18]. Valley [OMD^{+}12]. Value [GCNF13, KHDL07, WAM^{+}19, YHW^{+}07]. Value-Cell [KHDL07]. Value-Driven [WAM^{+}19]. Valued [GK95, GPP^{+}16]. Values [Zag96]. Valve [BSR^{+}14]. Variability [AAGS19, FFST19, FBW16, HSSK16, JBMS09, KBL19, PMW13]. Variable [BBIF12, FC95, GABJ07, JH16, MCP^{+}06]. Variable-Speed [FC95]. Variables [AT16, BI12, GDST16, GHL15]. Variance [BGOJ16]. Variant [FNM13]. Variants [FNM13]. Variate [LD11a, WS06b, YXG^{+}10]. Variation [GQM^{+}18, HHZH17, ZWZD15]. Variational [BS08, CYC^{+}12, ZWZD15]. Variations [LG12, RKC^{+}16]. Variety [DSP^{+}17]. Varifocal [DTT^{+}17, RWF19]. VarifocalReader [KJW^{+}14]. Varying [Ano14h, AM13, BRT12, BLLS17, CKW^{+}12, CLC^{+}15, CMP14, DB07, FWSL12, GABJ08, GSDJ04, GW11, GWP^{+}16, Har16, HE06, JEG12, JS06, JY17, KBD^{+}11, KLySK12, KGJ09, KG06, LLL^{+}12, LS09, LMC02, MVN^{+}19, NJ99, OBJ16, PSR17, RBDG15, SK16b, SK98, SB06, SSR^{+}07, SH00b, TIW^{+}19, WFKH07, WYM08, WC09, WC10, WCJ06, WS06b, YLK12, CJR07, DS16b, KC14, NSW^{+}17]. VASA [KZX^{+}14]. Vascular [KOCC14, KGPS13, RHD^{+}06, WJR^{+}13, JQD^{+}08]. VAST [Ano12g, Ert10c, MW13, vW11, Ano16u, Ano18h, Ano18i, Ano18j, Ano19l]. VAUD [CHW^{+}18]. VDR [YSGM05]. VDVR [ZXM10]. Vector [AJDL08, BW08a, CML^{+}07, CMLZ08, CKW^{+}12, CM02, CLB^{+}16, DANS10, DPR00, DNP07, ES05, FBW16, FCL09, GPR^{+}01, GKT^{+}08, Hau97, HEWK03, HA04, HE06, HLD^{+}08, LKJ^{+}05, LJL04, LVRL06, LCS06, LTWH08, LHFY12, LFR03, MBS^{+}04, MCHM10, MK13b, NJ99, NLS11, PGL^{+}12, RLH11, RLK19, RT12, SKMR98, SB04, SLB04, SWCR15, SRW^{+}16, SBW17, SZ12, SS13b, TWHS05, USE13, WASQ18, WT10a, WJE01, WGS07b, WPL96, ZBG^{+}17]. Vector-Based [CLB^{+}16]. Vectorial [IHR01]. Vectorization [YCZ^{+}16, YHL^{+}17]. Vectorized [HXF^{+}15, SGM08]. Vectorizing [ZCZ^{+}09]. VectorLens [DMC15]. Vectors [BHB04, FPH^{+}08, JCWD14, ONL^{+}12, VP09]. VEEVVIE [PGK16]. Vega [SRHH16, SMWH17]. Vega-Lite [SMWH17]. Vehicle [ZHLR14]. Vehicular [FCZ15]. velle [SB17]. Velocity [CS08, NSE^{+}12, PT17, YNBH11]. Velocity-Aligned [CS08]. Velocity-Dependent [NSE^{+}12]. Venant [RLK19]. Venn [Wil12]. Ventricular [HTP^{+}08]. VERAM [CZZ^{+}19]. Vergence [Kra16, XGS^{+}19]. Vergence-Accommodation [Kra16]. Verifiable [ESN^{+}09, ZXM10]. Verification [EGG^{+}12, ENS^{+}12, GRVE07, HAGS16]. Verifying [EJR^{+}14]. Verisimilitude [Yon14]. Versatile [HWW18, LDC96, PBO^{+}14, SR17, YLX^{+}12]. Version [CMFL16]. Versus [Chi16, HSKIH07, DFG^{+}14, GLM^{+}17, KYT^{+}18, PTM^{+}18]. Vertex [CM02, KXW^{+}18, LY06]. Vertically [SM11]. Vertices [FPH^{+}08]. Very [IV11, KPR^{+}15]. Vessel [AMB^{+}13, LLL^{+}12, MVB^{+}17, RHR^{+}09, JQD^{+}08]. VGTC [Ano12h, Ano13j, Ano13k, Ano13c, Ano13e, Ano13d, Ano13f, Ano14l, Ano15c, Ano16k, Ano16l, Ano16d, Ano16e, Meh17, Sil17a, Sil17b, Sil18a, Sil18b, Sil19]. Via [GQM^{+}18, ADDG12, BYA15, BMS17, BDD^{+}16, BMW17, BSV11, BHSH15, CGL^{+}17, CWT^{+}08, CYB08, DZMQ16, DT10, Elb95, GMD^{+}16, GSCI15, GNDV^{+}18, GGZ^{+}18, HKYM17, IFP97, JKM01, KCPE16, LYY^{+}16a, LS13b, LL14, LGY19, LLL^{+}19a, LSY^{+}18, LFW^{+}19, MSA17, NXSL13, OKSK16, PSKN06, PKL^{+}18, PW12, RB07, RB18, SSV18, SSW18, TC09a, TW18, WWB^{+}13, WSYM17, WGZ^{+}19, WHX^{+}19, WAG^{+}12, WMA^{+}16, WS09, WPZ^{+}11, YHLJ08, YYFX18, ZHZ15, ZWC^{+}16, ZNZX16, ZCL^{+}19, ZWLC19, ZCW19, ZKK02, vdEvW14, vLBR^{+}16]. Vials [SAB^{+}16]. ViBr [CXDR19]. Vibrotactile [LBHW18, dJOBNM17, UBH19]. Video [Ano14h, BGCS17, BC18a, BBS^{+}08, CBH^{+}06, CCL^{+}16, CRH05, DFG^{+}14, DTW^{+}15, FH06, FR13, GSPJ08, HB14, HKH^{+}12, KH19, KBD^{+}11, KCWI13, LHJ^{+}18, LCP^{+}13, LPS^{+}13, LGYG12, LLY^{+}13, LDX10, LZW^{+}13, LXT18, MI13, MFZ^{+}17, NXSL13, OH12, PLC^{+}11b, PZ07, PWG17, Rob98, RSSA08, RG95, SJL^{+}18b, SBE^{+}15, SJK^{+}12, TT05, TUG17, WBK^{+}08, WLHL13, WLT^{+}18b, WLH^{+}12, YLK12, ZHQ^{+}07, ZNZX16, ZLK^{+}18, ZZH19, SKC^{+}19]. Video-Based [FR13, ZLK^{+}18]. VideoPlus [Tay02]. Videos [ARL^{+}17, LYL19, RPAC17, SZN^{+}18, WKCB07, WBK^{+}08, WCW^{+}16, ZDJ^{+}09]. ViDX [XMRC17]. View [AS98, BH07, BSEN18, BNTM16, CZZ^{+}19, CLW18, DTT^{+}17, EHH^{+}19, FNM13, GLRH13, GR15, GY02, HSH10, IYS13, JS06, KLMA10, KERC09, KNO15, KKSM19, KOK^{+}19, LZZ^{+}19, LP02, MS18a, MCHM10, MGJ^{+}10, MTB17, MSM^{+}11, PD04, PBK^{+}12, RBK^{+}15, SM09, SOS^{+}17, TEC^{+}16, TLS17, Wah14, WC09, WC10, WKME03, YSGM05, ZEC08, Zhu05]. View-Dependent [GLRH13, GY02, HSH10, KNO15, KKSM19, LP02, MCHM10, PD04, TEC^{+}16, TLS17, WC09, WC10, YSGM05, ZEC08, Zhu05]. View-Enhanced [CZZ^{+}19]. View-Independent [SM09, WKME03]. Viewable [OHH06]. Viewing [BHW06, IDW^{+}13, KBKG07, LH11, LLKN17, LDN11, LODI16, LSK^{+}18, TKTN09, TUG17, vWN04]. Viewings [RCW^{+}18]. Viewpoint [HWZ^{+}19, IAIK16, JNC^{+}15, LPS^{+}13, LDX10, TMWS13]. Viewpoint-Dependent [IAIK16]. Views [BBH^{+}17, BBD06, BG06, CMP09, ČB19, GKN05, HTC09, LBD13, LKD19, Lin16b, QH18, SZY^{+}18, WWZ^{+}19, Wea10, WAG06, ZAM11, vW06, vdEHBvW14]. VIGOR [PHE^{+}18]. Virtual [ANR^{+}18, AGN^{+}19, Ano12e, Ano14j, Ano14d, Ano14e, Ano15c, Ano16d, Ano16e, AM13, AS11, BGC^{+}11, Bai13, BB07, BZS^{+}13, BAP^{+}17, BTB^{+}04, Bil13, BL15, BPS13, BDH^{+}18, BIPS12, BSWL12, BRNB19, BLRW05, BC18b, CPG^{+}15, CWL12, CDAF18, CLL08, CGJM19, CPK^{+}05, CRI06, CVC^{+}12, COMP13, CWC^{+}06, CMPC06, CLS13a, DS17a, DS18a, DV95, DJK^{+}06, DL12, DRHK07, DDBR^{+}19, DDKA06, ERL^{+}13, ESV98, FWK16, FS14, FBI07, Fuc13, GLM^{+}17, GPL^{+}13, Ger17, GWP^{+}18, HK10, HBJP12, HSR18, HWHK16, HSR13b, HHCL01, HMT10, HSH04, HLRC^{+}12, HBT14, HK16, HR11, ILMH12, INCB18, JBS^{+}18, JAAL18, JOR^{+}19, JPD^{+}18, JAM^{+}14, JSB13, JHBK19, JWL05, KJH^{+}18, KHSS14, KBS13, KPBL16, KCT^{+}17, KKL11, KYT^{+}18, KKPS08, LPG^{+}18, LKS^{+}19, LKC09b, LYS^{+}10, LBHW18, LLQ^{+}17, LLL^{+}19a, LSCN09, LIM^{+}12, LBS^{+}16, LvL12, LWP^{+}06]. Virtual [LDR00, MWCR06, MYI13, MBZB12, MRG^{+}15, Meh17, MZS^{+}19, MGO^{+}19, MF11, MLS18, MT01, NMN^{+}18, NBM19, NWF^{+}05, NSN14, NTS11, NSS03, NTT^{+}19, OBKP18, OBS^{+}15, OIR^{+}17, PFW09, PWHK16, PBK^{+}12, PKG12, PB16, PMT^{+}19, PGI^{+}17, RBK^{+}15, RSBB17, RJD^{+}07, RBK^{+}19, RCW^{+}18, RZP^{+}07, RGFLL14, RKA^{+}13, RCL^{+}15, RFSP19, Ros13, SGQ16, Sat13, SSI99, Sch13, Seq12, SL08, SMP17, SvLF10, SGJM18, SLMA06, SSL08, SFL^{+}16, SBHW11, SB14, SEB19, SAC^{+}08, SFR^{+}10, SLF^{+}12, TMDO15, TMM^{+}13, TL11, TWMZ19, TYL^{+}18, UK12, UKF^{+}18, VSS08, VBC^{+}16, Wah14, WGR^{+}18, WXJD17, WLW17, WKW06, WGC^{+}08, WRHR19, WBA^{+}14, YQK^{+}17, YAE07, YFZ^{+}18, ZLB^{+}05, ZK17, ZBB^{+}06, ZHF^{+}19, ZKM18, vPBB^{+}11, vTRvdM97, SL11]. Virtual-Reality [DJK^{+}06]. Virtuality [RRD^{+}13]. Virtualized [SvdBLM11, SRKL19]. VIS [GCML06, IHK^{+}17, Ano09e, Ano13v, Ano13t, Ano13u, Ano14o, Ano14s, Ano14t, Ano14u, Ano14v, Ano16p, Ano16v, Ano16w, Ano16x, Ano17p, Ano18k, Ano19m, LB19, Ano09d, OSS^{+}17, RSSA08, SGB^{+}19]. Vis-A-Ware [OSS^{+}17]. VIS-STAMP [GCML06]. Vis4Heritage [KLYE13]. VIS4ML [SKKC19]. VisBricks [LSS^{+}11]. VisComplete [Koo08]. Viscous [BK17, LBS^{+}19, PT17]. VisDock [CPW^{+}15]. VisFlow [YS17]. VisGets [DCCW08]. Visibility [CICS05, COCSD03, CF10, CMSW04, CM11, Dac11, FWT^{+}04, HXF^{+}15, KS01, KJL^{+}12, KKCS98, LRP97, LSS13, ME18, MAST16, OSS^{+}17, PPZ^{+}12, Ste98, WC13, ZZBW08, vRKEE17]. Visibility-Aware [OSS^{+}17]. Visibility-Driven [CM11, ZZBW08]. VisibilityCluster [WC13]. Visible [IIS14, KS00a, KS00b, MCP^{+}06, TSB^{+}05]. Vision [CK16, CWC^{+}06, DH08, EIKS18, MOF09, MOF10, OTKS15, RSSA08, DMAM19, SS95]. Vision-Based [DH08]. Visiting [GB08a]. Visitor [LWD^{+}17]. Visitors [MLMF12]. ViSizer [WLLM13]. ViSlang [RBGH14]. VisLink [CC07]. VisMashup [SLA^{+}09]. VisMatchmaker [LWZQ17]. VisOHC [KKL^{+}16]. Vispedia [CWT^{+}08]. Visplause [ASMP17]. Vispubdata.org [IHK^{+}17]. VisTiles [LHD18]. Vistrates [BMR^{+}19]. Visual [AHSS14, AJDL08, AABH^{+}16, AMM^{+}08, AAMG12, AAMH13, AHH^{+}14, AHRG10, AABW12, ABC^{+}19, Ano12g, ASMP17, AERA14, AIS18, BW14, BDF16, BBD^{+}11, BTC13, BKW16, BSS^{+}19, BMLC19, BHZ^{+}18, BAAK^{+}13, BI12, BBG^{+}09, BJK^{+}16, BAF^{+}13, BARM^{+}12, BSEN18, BPM^{+}13, BBIF12, BDD^{+}16, BISM14, BC12, BŠG^{+}09, BM10, BSWL12, CDDS18, CGSQ11, CWZ^{+}14, CSL^{+}16, CLG16, CLZ^{+}18, CD19, CGM^{+}17, CWT^{+}08, CBH^{+}06, CCM^{+}14, CYW^{+}16, CCL^{+}16, CXR18, CHW^{+}18, CVG13, CDF14, CDW^{+}16, CORLS96, CMPC06, CLB11, CCM12, CFEC17, DVP^{+}18, DLW^{+}17, DSG^{+}17, DGWC10, DDBR^{+}19, DR08, ERLW18, Eic00, EASS^{+}18, EASD^{+}19, EDF08, ERHRF10, EDF11, EGG^{+}12, Ert10c, EBB^{+}15, FHKM17, FKLT10, FPV^{+}13, FMH08, FHG^{+}09, FZCQ17, FWG09, FBM16, GFG^{+}14, GBPW10, GKL^{+}13, GJZ^{+}12, GS14, GHL18, GLG^{+}13, GS06, GRVE07, GBFM16, GHL15, GXZ^{+}18, GJG^{+}19, GDKB17, HSCW13, HWHK16]. Visual [HEFR18, HSR13a, HKR^{+}08, HE12, HKBE12, HHKE16, HEG^{+}17, HLRC^{+}12, HBW14, HKPC19, HFG^{+}12, HMZ^{+}14, HPvU^{+}18, HSTD18, HTA^{+}15, HZM^{+}16, HAS11, HOGJ13, HVF13, HJC14, IDA^{+}14, IFP^{+}12, IAS19, JER16, JBMS09, JFS16, JAO^{+}14, JE13, JSR^{+}19, JYC^{+}10, KAKC18, KTC^{+}19, KGS^{+}08, aKGS11, aKS12, KLYE13, KBE^{+}18, KNR17, KLM^{+}08, KMDH11, KH13, Kei02, KRTvW06, KHS^{+}19, KV08, KLK^{+}09, KKP^{+}17, KOJL^{+}14, KBGE11, KJW^{+}14, KDM^{+}16, KMG^{+}06, KBH06, KSDD14, KFS^{+}19, KPS16, KS02, KTE15, KSBE18, KG06, KRRW19, KW13, KHSW17, KKL^{+}16, KEV^{+}18, KCK^{+}19b, LSSB12, LBS13, LMK07, LHD18, LRM^{+}13, LCP^{+}13, LGM^{+}18, LBK^{+}18, LXC^{+}17, LWCC18, LLB^{+}06, Lin16b, LSS09, LS10, LH14, LS16, LYW^{+}16, LWL^{+}17, LDM^{+}18, LXL^{+}18, LBT^{+}18, LLL^{+}19b, LGS12, DSC^{+}16, LFA^{+}16, LDFZ14, LWLM18, LFW^{+}19, MWSJ14, MRO^{+}12, MRH^{+}10, MPK^{+}13, MHS07, MEV^{+}14, MRSS^{+}13, MMT^{+}14, MWCR06, MKN^{+}07]. Visual [MMH^{+}13, MPWG12, MGJH08, MGKH09, MGJ^{+}10, MGS^{+}14, MES^{+}11, MGB^{+}19, MBH^{+}12, MPOW17, MW13, MAAB^{+}18, MGM09, MZC^{+}16, MTRP10, NHB^{+}17, NHEM17, NSN14, NA19, OJ15, OBJ16, OSSK12, ODH^{+}07, OJEF19, PBN^{+}13, PSSC17, PFP^{+}11, PSPM15, PSTW^{+}17, PGK16, PKL^{+}18, PGU^{+}13, PVF13, PQF^{+}09, PLvdM^{+}17, PHV^{+}18, PHE^{+}18, PTMB09, PDW^{+}14, PSM12, PB16, POM^{+}09, PV06, QPNK18, QCX^{+}07, RBK^{+}15, RWF^{+}13, RFLL18, RLA^{+}13, RNE^{+}17, RM15, SSS^{+}14, SSK^{+}16, SZS^{+}17, SKKC19, SKBE17, SNLD06, SSEW19, SGB13, SZK15, SHB^{+}14, SW13, SMER06, SWY^{+}17, SZY^{+}18, SBB^{+}18, SPN^{+}16, SKU^{+}12, SYS^{+}06, SLK^{+}17b, SLMA06, SS18, SWS^{+}11, SPG14, SSL^{+}12, SGPR18, SGB^{+}19, SWL^{+}14b, TAE^{+}11, TKE16, TBHC16, TvET14, TFJ12, TWSM^{+}11, TSD09, TFH11, TLLH12, TSH^{+}14, USKD12, UDSL18, Vas16, VJC09, VN19, WSH^{+}19, WSD^{+}13, WDC^{+}18, WCR^{+}11, WS06a, WK13, WZvdW13]. Visual [WYL^{+}14, WM16, WFW^{+}17, WGYS18, WCC^{+}18, WGZ^{+}19, WGSY19, WCC^{+}19, WLS^{+}19, WHLS19, WV08, Wea09, Wea10, WFG^{+}19, WCR^{+}18, WAG06, WFM^{+}06, WFC^{+}06, WDSC07, WXC^{+}08, WCQ^{+}09, WLY^{+}14, WPZ^{+}16, WXZ^{+}16, XCH^{+}14, XZM17, XXM19a, XWW^{+}13, XMRC17, XXM^{+}19b, YYR17, YHW^{+}07, YSD^{+}17, YDC^{+}14, YRWG13, ZMH^{+}09, ZYM^{+}14, ZMZM15, ZM17, ZHF^{+}07, ZCW^{+}14, ZSCC18, ZWC^{+}18, ZWLC19, ZMT^{+}19, ZK08, dOL03, vW11, vdEHBV16, vLBR^{+}16, LSC08]. Visual-Interactive [BHZ^{+}18]. Visualisation [ED07, GMD^{+}17, KKC15, YDGM17]. Visualization [AvHK06, AL06, AJDL08, AMJ^{+}12, AZM12, AD12, AJ17, ADG11, ALBR16, AHK^{+}17, ARB07, ARRC11, AHKMF11, AR17, AWC10, AH11, Ano10c, Ano10d, Ano11i, Ano11c, Ano11d, Ano12h, Ano13j, Ano13k, Ano13c, Ano13e, Ano13d, Ano13f, Ano14c, Ano14j, Ano14k, Ano14l, Ano14g, Ano14f, Ano15a, Ano16f, Ano16a, Ano16k, Ano16l, Ano16b, Ano16c, Ano17e, Ano18a, Ano18b, Ano19a, Ano19f, APS^{+}14, AB01, ASE16, AMB^{+}13, BRH^{+}17, BW08a, BFE15, BRT12, BB07, BCH^{+}13, BE18, BKDE00, BGOJ16, BMWM06, BI12, BJB^{+}12, BWW^{+}17, BZGV14, BLLS17, BVW^{+}07, BAB^{+}18, BBB^{+}19, BSL^{+}12, BJA^{+}19, BARM^{+}12, BYB^{+}13, BVB^{+}13, BBK^{+}16, BSR^{+}14, BH09, BBS^{+}08, BSL^{+}14, BEJK12, BRBF14, BEDF16, BHTY15, BVPtHR09, BM13, BE09, BPS^{+}11, BIRW19, BGM^{+}17, BMW17, BFTW09, BvL06, BVB^{+}11, BTJ^{+}13, BAW16]. Visualization [CR08, CDC^{+}07, CSL^{+}10, CFM^{+}13, CCH14, Car19, CZ11, COJ15, CWK^{+}07, CBH^{+}06, CYB08, CWDH09, CJ10, CCAL12, CZC^{+}15, CG16, CGJM19, CCJ^{+}19, CDL^{+}16, CLC^{+}15, CRI06, CCQ^{+}14, CPW^{+}15, CFHH09, CMPS97, CDM^{+}04, CvW11, CDF09, CW11, CSC06, CM09, CDM^{+}06, CRB^{+}05, CZQ^{+}08, CRPH10, CMP14, DK11a, DPW^{+}15, DSSK08, DBM^{+}06, DDW14, DBH14, DZMQ16, DFQ12, DDGL07, DGBW09, DBW11, DPR00, Dil17, DBD17, DK11b, DLR09, DTW^{+}15, DWBR06, DHR^{+}19, DCH^{+}17, DRMM13, DB07, DWS10, ES01, Ebe17, EGH^{+}06, EDF08, ET08, EF10, EHP^{+}11, EMdSP^{+}15, ESN^{+}09, FYTL19, FWD^{+}17, FPB17, FKRW17, FCL09, FC95, FSE12, FSME14, FTES13, FH07, FH16, FM04, GGTH07, GJ10, GNBP11, GLvP^{+}12, GGJ^{+}18, GLH^{+}14, GHGM06, GNSP^{+}14, GJC^{+}17, GDJ^{+}13, GLK^{+}13, GSWD18, GABJ07, GABJ08, GGA^{+}11, GWE^{+}19, GHL18, GTS10, GTS11, GRS95, GMM05]. Visualization [GK95, GBM^{+}12, GKM^{+}15, GY02, GST16, GCML06, Guo09, GMY11, GXY12, GGZL16, GLX^{+}18, GHS^{+}19, GKK^{+}12, HBJP12, Hag98, HHO^{+}17, HVY16, Han16, Han18, HM95, HHB16, HW12, HKQ13, HE99, HKR^{+}08, HE12, HD12, HA06b, HMSA08, HB10, Hee18, HP04, HSF^{+}06, HA04, HPAW07, HFM07, HMM00, HSSK16, HBAB14, HSC08, HLNW11, HKB^{+}19, HKH^{+}12, Hol06, HZH14, HNR^{+}06, HPC^{+}13, HTA^{+}15, HD11, HDR^{+}13, HQC^{+}19, HEF^{+}14, IPD^{+}07, IG19, IC07, IDW^{+}13, IIC^{+}13, IHK^{+}17, IIS^{+}17, IYIK04, INCB18, JHKH13, JEG12, JWSK07, JBS08, JHH^{+}10, JKM01, JKM06, JKJTM06, JKMG07, JFTW07, JBH^{+}09, JPLŠ16, JHGH08, JEH02, JCG08, JH09, KMS09, KPHH12, KSG^{+}16, KCJ^{+}10, KHE09, KAM^{+}08, KERC09, KH13, Kei00, Kei02, KHS^{+}18, KOJC12, KOCC14, KGD^{+}19, KKSS13, KV06, KZD^{+}10, KJW^{+}18, KRH18, KW06, KS14b, KCS^{+}16]. Visualization [KAK^{+}18, KWDG11, KBH^{+}10, KBVH17, Koo08, KW19, KLL12, KK19, KGJ09, KBE09, KKKW05, KSW06, KHH^{+}16, KMLM16, KCM18, LBS14, LKJ^{+}05, Lam08, LBI^{+}12, LS07a, LVRH07, LCMH09, LvWJH04, LHH^{+}12, LLL^{+}12, LB19, LGV^{+}16, LTPH17, LS09, LIRC12, LSS13, LKH^{+}16, LKK17, LKHW04, LLT04, LSS^{+}11, LGS^{+}14, LDC96, LLL06, LTWH08, LBH18, LDN11, LBH11, LBLH19, LP02, LLRR08, LLCD11, LNS08, LS10, LCS^{+}12, LST^{+}16, LBR^{+}17, LWC^{+}19, LDA12, LPK^{+}16, LTP^{+}05, LWL14, LSM03, LLPY07, LRF^{+}11, LLCM12, MLMF12, ME19, MRO^{+}12, MKHD05, MB03, MB19a, MHG10, MDL07, MGL07, Mar18, MK16, MCK12, MHD^{+}18, MAK08, ME09, MMAM14, McK09, MGM19, MI13, MJK06, MSW^{+}08, MMYK06, MSE^{+}06, MRS^{+}13, MVB^{+}17, MSSD^{+}08, MNKW07, MDS^{+}18, MQF06, MiS^{+}18, MTW^{+}12, Moo03, MMCE09, ME11b, Mor13, MWN^{+}19, MHDG11, MGW10, Mun09, NMGK17, NGK18, NW15, NDM^{+}97]. Visualization [NJB07, NLKH12, NKH14, NSW^{+}17, NKHC08, NZ06, Nie95, Nie96, NSL19, NOB16, NSvW11, NH06, NASK18, ORRL10, OJ12, OHJ^{+}11, OM09, OR98, OR99, OSS^{+}17, OPH^{+}16, OSBM14, PW95, PZ12, PZ11, PLC^{+}11a, PBO^{+}14, PMN^{+}14, PSKN06, PK13, PGK16, PPL^{+}99, PLC^{+}11b, PGL^{+}12, PKS^{+}08, PPZ^{+}12, PK16, PHJ^{+}10, PSBS12, PW12, PYHZ14, PRA^{+}10, PSM07, PPvA^{+}11, PBCR11, PMD^{+}07, QMK^{+}06, QH18, RESC16, RGF^{+}04, RBN^{+}19, RVG06, RBGH14, RBS^{+}18, RPSC99, RGK^{+}13, RGP^{+}12, Rob98, RRJH18, RFF^{+}08, RASS17, RSA^{+}19, RB11, RHR^{+}09, RPHI08, RP12, RKG^{+}18, RB18, RM15, RSD^{+}13, RvWT05, SHM10, SMDS14, SPS06, SP07, SKBE17, SSRE18, SE17, SSC^{+}16, SLA^{+}09, SZB^{+}09, SZD^{+}10, SRHH16, SHVV16, SFBP09, SKS12, SFP^{+}19, SWC^{+}08, SPO^{+}12, SKK^{+}14, SK16b, STS07, SHS11a, SHS11b, SNHS13, SASS16, SNG^{+}17, SFMB12]. Visualization [SH10, SG09, SKYS14, SBSG06, SS16, SOL^{+}13, SAM^{+}07, SLB04, SCL^{+}12, SWW^{+}15, SRM19, SSG12, SA06, Sil17a, Sil17b, Sil18a, Sil18b, Sil19, SJH^{+}07, SVGR16, SPCJL06, SCM^{+}06, SOL^{+}16, SD12, SPEB18, SGP^{+}19, STH02, STH03, SB17, SEB19, SPB08, SOR^{+}09, SSBC19, SGM^{+}11, SEA09, SSG16, Sza18, SS13b, TH13, TD95, TSLR07, TLD^{+}12, TMWS13, TWSK14, TWSS16, TCM06, TDR10, TBB^{+}07, TBB^{+}08, TC09b, TCM10, TIC09, TSAA12, TLS17, TM04, TKAM06, TBR^{+}12, TGS11, TNS10, USM97, UKF^{+}18, VZS18, VABW09, VP04c, VFR13, VWvH^{+}07, VWF09, VKG05, VI18, VB18, WJA^{+}17, WBJ16, WS06a, WGS07a, WGM^{+}08, WLT08, WYM08, WWLM11, WSA^{+}16, WFC^{+}18, WWS^{+}18, WLF^{+}19, WHLS19, WCJ06, WTL^{+}09, WEE03, WBE^{+}06, WSE07, WZ08, WCD^{+}19, WAG^{+}12, WTS^{+}07, WKB^{+}13, WRT19, WFS^{+}19, WS01, WJR^{+}13, WWFT03, WSM^{+}09, WFM^{+}12, WG12, WMA^{+}16]. Visualization [WII^{+}12, WKD19, WS06b, WS09, WKSS05, WWL^{+}10, WLLM13, WP16b, WCA^{+}17, WLS^{+}18, XHT^{+}07, XSZ^{+}17, XLS10, XRP^{+}12, XC19, YESK95, YGV^{+}13, YHR^{+}19, YaKSJ07, Ynn19, YN06, YSI^{+}10, YS17, YNCP06, YSZ^{+}19, ZHT07, ZH07, ZYLL09, ZWA^{+}13, ZSL^{+}16, ZCD19, ZFS^{+}19, ZCCB12, ZGB^{+}17, ZBG^{+}17, ZXM10, ZAM11, ZZBW08, ZW18, ZOC^{+}13, dLVvL06, tCMR07, tCMR08, vHvdWvW02, vPBB^{+}10, vPVvdW10, vWPSP96, vWC06, vW06, vW14, vdEHBvW14, vFWTS08, vLFR17, zBBKN14, BPC^{+}10, CJR07, FHL10, GSL14, GLS17, JQD^{+}08, THT19, WOCH09, WBD14, Joy02, MFS^{+}09, SGR06, Var01, vWMT04]. Visualization-by-Sketching [SK16b]. Visualization-Driven [HBJP12]. Visualization-Opportunities [KGD^{+}19]. Visualizations [AS05, ARH^{+}15, BSB^{+}18, BLE19, BBD^{+}11, BDFM17, BVV^{+}19, BGP^{+}11, CC07, CPC09, CLKS19, CWRY06, DBB10, DBD18, DBBF19, DCCW08, DMC15, FPH19, FDFR10, GPL^{+}11, GHC^{+}16, GGC^{+}17, GNDV^{+}18, GWFI14, GTPB19, GTS10, GTS11, HYFC14, HV13, HBW06, HKKS18, JH13, JD13, KH16, KDX^{+}12, KC14, KKW^{+}17, LBGV13, LAM10, LKH^{+}16, LD11a, LPCC17, MRSS^{+}13, MK13b, MHD^{+}18, MJ09, MDF12, NBW14, NTT^{+}19, PDFE18, PDF14, PFG08, PMH18, RHY14, RRJH18, RM15, SED19, SND05, SVK^{+}07, SS13a, SGS^{+}19, SK16b, SLC^{+}19, SS19, SDES19, STS^{+}14, SNR14, TM12, THM15, TTR10, VC17, VMCJ10, WLB^{+}14, WAM^{+}19, WCR^{+}11, WHP^{+}18, WHA07, YEB16, YEII12, ZLB^{+}05, ZGC^{+}17, ZLC^{+}19]. Visualize [BSH^{+}16, BSKR19]. Visualizer [vTRvdM97]. Visualizing [AWHS16, AK02, AABS^{+}14, AW14, Bac07, BHS12, BS95, BPP^{+}16, BMGK08, BSSB10, BEW95, BHP^{+}12, BLIC19, BJC^{+}19, CQM10, CXDR19, CdOKRV09, DCK^{+}12, FTB^{+}13, FLF^{+}11, FNM13, FS04, GKN05, GHA^{+}08, Gle18, GDST16, GBP^{+}13, GIK^{+}07, GGPPS13, GZL^{+}14, GPP^{+}16, HHWN02, HHZH17, HWS17, IML13, IBJ^{+}14, IWSK07, JHP^{+}14, JM10, JCRS09, KIL07, KHA12, KHDL07, KLC08, KRH18, KBI^{+}18, KSSW09, KM96, KTB^{+}18, LMZ06, LLB^{+}12, LAK^{+}11, LRKC10, LPK^{+}13, LFLH07, LFH08, LMW^{+}17, LPCRH19, MRSS^{+}12, MKH12, MLCM16, MQB19, NSH^{+}18, NJBJ09, NGCL19, OHWS13, PFK07, PMW13, RFFT17, RSOW18, RKKF19, RSSA08, RM13, SSMG13, STM17, STS06, SBV^{+}11, SKMR98, SBM^{+}06, SK98, SWL^{+}14a, SW97, STM08, SGAS16, SAB^{+}16, Sun03, TRd12, TTS10, TS07, VRW13, VBW16, VvWvdL06, WHZ^{+}18, Wil18, WPL96, WSW^{+}18, WYM12, WXW^{+}19, XCZ^{+}19a, ZFA^{+}14, ZDL03]. Visualizing [ZCL09, ZWJZ12, ZWR14, vLdL03]. Visually [DYW^{+}13, GdBG12, LKC09b, LYK^{+}12, MOC^{+}14, PGSF16, PIS15, QM16, WCS^{+}18, YLZ^{+}13]. Visuo [CGB^{+}13, EPS^{+}15]. Visuo-Haptic [CGB^{+}13, EPS^{+}15]. VisWeek [Ano12q, Cox11, Cze12, Fra12, Heg10, Sza10, Tha11]. Vivaldi [CCQ^{+}14]. Viz [RSSA08]. Viz-A-Vis [RSSA08]. VizItCards [HA17]. VLAT [LKK17]. vLOD [CPK^{+}05]. Vocabulary [IHD^{+}18, TWSS16]. Voila [CLZ^{+}18]. Vol [Ano14c, Ano15a, Ano16f, Ano16a, Ano18a, Ano19a, SB17]. Volume [AWC10, ASW13, AD16, AZD17, AGDJ10, Ano09b, BRSP18, BLM96, BLL19, BMTD05, BMWM06, BHWB07, BPL^{+}19, BJNN98, BG06, BGKG06, BKW08, CR08, CICS05, CBPS06, CQC^{+}08, CWM^{+}09a, CYZ^{+}09, CCB11, CCJ^{+}19, CCQ^{+}14, CMPS97, CDM^{+}04, CHF95, CHF96, CML^{+}12, CMSW04, CSC06, CM08, CM09, CHM11, Csé08, DSP^{+}17, EMRY02, EJR^{+}14, FM07, FM12a, FM12b, FAW10, FE17, FH16, FMST96, GMD^{+}16, GMS^{+}07, GHGM06, GAMD10, Guo95, GMY11, GXY12, GWK12, HLRS^{+}08, HBJP12, HAAB^{+}18, HBG11, HWHK16, HMBG01, HLM10, HA04, HLY10, HHM14, HBC12, HZM13, HZH14, IZM18, ISC07, IYK01, JNK19, JV09, JH09, JKRY12, KSH03, KOCC14, KKSS13, KWP01, KV06, KMKY10, KWH00, KISE14, KHM^{+}98, KKH02, KPH^{+}03, KHPS07, KHW^{+}09, KJL^{+}12, KWH19, KKPS08, LBG^{+}16, Lac96, LSSB12, LBS13, LBS14, LCMH09, LGM^{+}08, LYS^{+}10, LS13a]. Volume [LR11, LLRR08, LCDP13, LWP^{+}06, LTP^{+}05, LMC02, LLY06, LLPY07, MB18a, MDM10, Max95a, MB18b, MAWM11, MHDG11, Mur95, NKH14, NJ16, NvdVS00, PMP10, PPL^{+}99, PK16, PHF07, PFK^{+}08, PRH10, RSB96, RBG07, RPSC99, RLNN11, RE01b, RDB^{+}12, RL08, RZNS04, SE17, SWB^{+}00, SYM14, SM17, SMP11, SE18, Sel15, SZN^{+}18, SRM19, SKMH14, SHC^{+}09, SK00, SB17, SYR11, SGM^{+}11, TLD^{+}12, TIW^{+}19, TCM10, TLM05, VKG05, WS06a, WGS07a, WM08, WCZ^{+}11, WWLM11, WK13, WDC^{+}07, WEE03, WFG^{+}19, WKZL04, WMS98, WQ07, WKI^{+}17, XTY^{+}11, YESK95, YLX^{+}12, YFC^{+}19, YL95, YNCP06, ZM13, ZDM13, ZXM10, ZAM11, ZWM13, vPVvdW10]. Volume-Accurate [AGDJ10]. Volume-Based [PMP10]. Volume-Enhanced [YFC^{+}19]. Volume-Preserving [HZH14, RSB96]. Volume-to-Volume [FE17]. Volumes [BMA^{+}19, CMCL06, EMRY02, FG99, HBAB14, HK99, IVJ12, JS06, KCOY03, KSTE06, Mao96, MWK^{+}08, SMG^{+}13, SKMH14, Sil95, WFKH07, WD10, WFG^{+}19, vRKEE17]. Volumetric [ASDW14, BKRE19, BAAK^{+}13, BKS01, BS02, BG07, EM06, FMST96, GWBO12, GW11, GWP^{+}16, HRN^{+}03, JFTW07, JFY16, JY17, KPR^{+}15, KKSS13, LLWQ13, ME18, MWCE09, MK09, MMCE09, MNC14, Ney98, NT03, OKI15, PCY08, QXF^{+}07, RWBF18, SFA^{+}15, SNM16, SE18, SBSG06, SLK^{+}17b, SH12, SEA09, WZQK04, WLL^{+}12, WSE07, WGC^{+}08, WMGE12, XYGL13, ZZSS10, ZT99, TGSP09, THT19].

VolVis [YNCP06]. Voronoi [BTC10, LBH11, LSPW12, MOG11, ME11a, RLW^{+}11, VCP08, YBZW14, YW16]. Voronoi-Based [LBH11, MOG11]. Vortex [GT14, GT19, JKJTM06, KRHH11, KWDG11, KGP^{+}13, RC06, SWTH07, SOL^{+}13]. Vortex-Turbine [SOL^{+}13]. Vortical [OJCJP16, PW95, VHLL14]. Vortices [GST16, GPP^{+}16, GLX^{+}18, HPAW07, SBV^{+}11, WTS^{+}07, WPB^{+}11]. Vorticity [HPAW07, PT17, SPS06]. Votes [DR08]. Voting [YRP18, ZCL^{+}19]. Voxel [KRW19, LK11]. Voxel-Based [KRW19]. Voxelization [DdL14, NDS10, SK99, ZD18]. Voxelized [DSKA19, KSDA16]. Voxels [JNK19, MSHC99]. Voyager [WM18, WMA^{+}16]. VPModel [MZS^{+}19]. VR [DBP14, ILMH12, KKL11, LSCN09, SL11, Ano14w, Ano15m, Ano15p, Ano16q, BTB10, CML^{+}12, DMAM19, FWL17, FSTG16, GLM^{+}17, GS16, JK16, KAM^{+}08, LBS14, LLBS17, LSV^{+}18, PGK16, PPZ^{+}12, WWYP19, WP18, XGS^{+}19]. VR-Based [DBP14]. VR-Systems [FWL17]. VRST [CLS07, TL11, KPGL12, LS06, Pur09]. vs [BSG18, CB15, DLW^{+}17, JOR^{+}19, PLE^{+}18]. VU [CRI06]. VU-Flow [CRI06]. Vulnerability [ABC^{+}19]. Vulnerable [CLB13]. Vulnus [ABC^{+}19].

Waldo [BOZ^{+}14]. Walk [Sbe97]. Walkers [LKM^{+}18]. Walking [AS11, BHHM19, BSSL19, BIPS12, BLS15, CVC^{+}12, COMP13, HB13, HBT14, JBS^{+}18, KKKT18, NSE^{+}12, NSN14, OBKP18, PFW09, PFW12, SHV^{+}18, SBJ^{+}10, SLF^{+}12, ZWBH13]. Walking-In-Place [NSN14]. Walks [ZZC11]. Walkthrough [CPK^{+}05, COCSD03, MRT00]. Wall [BI12, ETO^{+}10, GLH^{+}14, HPAW07, KLYE13, LGV^{+}16, LSY^{+}18, NHN07, PBC17]. Wall-Bounded [HPAW07, LSY^{+}18]. Wall-Displays [PBC17]. Wall-Sized [BI12]. Walls [LSJ^{+}15, RLM10]. Want [GMD^{+}17]. Ware [OSS^{+}17]. Warning [KGAM18]. Warp [GM05, Lac96]. Warped [JM10]. Warping [CK05b, IHK05, LYY08, LLY^{+}13, MZX15, SFP^{+}19, SSB^{+}17, STH13]. Waste [KHLM17]. Watchers [KWS^{+}14]. Water [LHBF19, LPS^{+}13, ZWW^{+}12]. Watercolor [DKMI13]. Watercolorization [Yon14]. Watermarking [DNP07, WH09, YPRI17, ZTP05]. Watershed [MW99]. Wave [DBM^{+}06, Kas12, MAKM14, MRG^{+}15, SRK^{+}11, MRG^{+}15]. Wave-Based [MAKM14, DBM^{+}06, MRG^{+}15]. Wavefront [XWL^{+}15]. Wavelet [CJTM05, DVP^{+}18, DL03, GGLQ19, JBMS09, LF97, LS13b, Mur95, PA06, VP04a, VP04b]. Wavelet-Based [DVP^{+}18, PA06, VP04a, VP04b]. Wavelets [BDHJ04, GK95, GSG96, HQ13, LWZ^{+}16, WQS07, WL16]. Wavemesh [VP04b]. Waves [KLCK17]. Way [BRBF14, KBS13, LTKF08, VHLL14]. Wayfinding [HLB^{+}18]. Weaknesses [DBB10]. Wearable [KNR17, YLC^{+}19]. Weather [BLLS17, FKRW17, KTB^{+}18, ME19, QM16, SZD^{+}10]. Weathering [JFBB10, MPBM^{+}18]. Weaving [HSKIH07]. Web [BKDE00, CCSK19, DDGL07, DCCW08, FPH19, FDFR10, GNSP^{+}14, KLK^{+}09, MEV^{+}14, McK09, NZ06, YDC^{+}14, YS17]. Web-Authored [FDFR10]. Web-Based [BKDE00, CCSK19, DCCW08, KLK^{+}09, MEV^{+}14, YDC^{+}14, YS17]. WebCharts [FDFR10]. Weber [HYFC14, KH16]. Weibull [HRN^{+}03]. Weight [JAO^{+}14, PSTW^{+}17, ZK17]. Weight-Shifting [ZK17]. Weighted [BA05, Cse13, DB07, PZLZ17, ZSCC18, HLD^{+}08]. Weighting [CH17, LDW^{+}15, LXW^{+}18]. WeightLifter [PSTW^{+}17]. Welcome [Ano15j, Min13]. Wellformedness [RZP12]. Wet [KSS09]. Wetting [PC13]. Wheel [AAMG12]. Wheelchair [JPD^{+}18].

Wheeled [LPF^{+}19]. Where [AJ17, SS19]. Which [RZP12]. While [Ger17, KDX^{+}12]. Whisper [BS11, CLS^{+}12]. White [EBB^{+}15, JFTW07, MSE^{+}06, SAM^{+}05, ZCL08]. White-Matter [ZCL08]. Whiteboards [WCR^{+}11, WLJ^{+}12]. Who [DR08, SBS16]. Whole [PFK07, RE14, WWFT03]. Whole-Brain [PFK07]. Wide [BJA^{+}19, DTT^{+}17, KOK^{+}19, MSM^{+}11, NZS^{+}17]. Wide-Baseline [NZS^{+}17]. Wide-Field [BJA^{+}19]. Wide-View [MSM^{+}11]. Widening [LSK^{+}18]. Widgets [WHA07]. Wijk [Ano13t]. Wiki [McK09]. Wiki-based [McK09]. Wikipedia [CWT^{+}08]. Wild [SFL^{+}16]. Wildcard [RGP^{+}12]. Wilhelms [Bar05]. Wilkinson [TLH10]. Will [YN03]. WIM [CML^{+}12, LFH06]. Wind [KFL^{+}15, SOL^{+}13]. Windowed [SJB10]. Wires [SLNB11]. Wise [SWF^{+}16, ZCW19]. Wisp [CK05a]. within [DMC15, KBL19, MBT07, OKB^{+}19, WKZL04, XHF12]. without [CMF^{+}18, Gor02, KJW^{+}18, LSC08, WYL^{+}19]. Wolfgang [Han95]. Word [BW17, BMS17, CLC^{+}15, FFB18, GWFI14, GBWI17, LBT^{+}18, PKL^{+}18, WCB^{+}18, WV08]. Word-Scale [GWFI14, GBWI17]. Word-Sized [BW17]. WORDGRAPH [RGP^{+}12]. Wordle [KLKS10, VWF09]. Workflow [IG19, MRSS^{+}13]. Workflows [BNTM16, MRSS^{+}12, PWHK16]. Working [PDBG18]. Works [CQM10]. Workshops [KGD^{+}19]. Workspace [LLL^{+}19a]. Workspaces [CVC^{+}12]. World [CDAF18, DBN06, HBT14, LAM10, NOB16, SLM18, VAB12, WFR^{+}10, WFM^{+}12, vFWTS08, RPHI08]. Worldmapper [DBN06]. Worlds [Bai13]. Worn [DFG^{+}14, CPW^{+}18]. Would [KCM18]. Woven [SZK15, ZBZ^{+}13]. Wrinkle [JFZ^{+}18]. Wrinkle-Aware [JFZ^{+}18]. Wrinkles [ZBO13]. Writings [OSSK12]. Wrong [PBK^{+}12]. Ws [ZWA^{+}13]. WYDIWYG [WM13a]. WYSIWYG [GMY11, KWL14]. WYSIWYP [WVFH12].

X [AHR^{+}11, ERL^{+}13]. X-ray

[AHR^{+}11, ERL^{+}13]. XCT [RGK^{+}13]. XPLORE [BHGK14]. xR [IPJT19].

Yarn [CLMO17]. Yarn-Level [CLMO17]. York [FPV^{+}13]. Younger [JBS^{+}18]. You’ve [DBN06].

Zeitgebers [SBS16]. Zero [KHLM17]. Zero-Waste [KHLM17]. Zometool [ZK14b]. Zone [RKSH11]. Zoom [BGR06]. Zoomable [BL07]. Zooming [JE13, MOC^{+}14, QWC^{+}09, RN19, SLQW17]. Zwart [EM06].

References

Andrienko:2011:SGA

[AA11] Natalia Andrienko and Gennady Andrienko. Spatial generalization and aggregation of massive movement data. IEEE Transactions on Visualization and Computer Graphics, 17(2):205–219, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrienko:2013:STU

[AAB^{+}13] Natalia Andrienko, Gennady Andrienko, Louise Barrett, Marcus Dostie, and Peter Henzi. Space transformation for understanding group movement. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2169–2178, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ai-Awami:2016:NVT

[AABH^{+}16] A. K. Ai-Awami, J. Beyer, D. Haehn, N. Kasthuri, J. W. Lichtman, H. Pfister, and M. Hadwiger. NeuroBlocks

— visual tracking of segmentation and proofreading for large connectomics projects. IEEE Transactions on Visualization and Computer Graphics, 22(1):738–746, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Al-Awami:2014:NSM

[AABS^{+}14] Ali K. Al-Awami, Johanna Beyer, Hendrik Strobelt, Narayanan Kasthuri, Jeff W. Lichtman, Hanspeter Pfister, and Markus Hadwiger. NeuroLines: A subway map metaphor for visualizing nanoscale neuronal connectivity. IEEE Transactions on Visualization and Computer Graphics, 20(12):2369– 2378, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875935-
abs.html.

Andrienko:2012:VAM

[AABW12] Gennady Andrienko, Natalia Andrienko, Michael Burch, and Daniel Weiskopf. Visual analytics methodology for eye movement studies. IEEE Transactions on Visualization and Computer Graphics, 18(12):2889–2898, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrienko:2018:CTR

[AAFG18] G. Andrienko, N. Andrienko, G. Fuchs, and J. M. C. Garcia. Clustering trajectories by relevant parts for air traffic analysis. IEEE Transactions on Visualization and Computer Graphics, 24(1):34–44, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrienko:2017:RPT

[AAFW17] Gennady Andrienko, Natalia Andrienko, Georg Fuchs, and Jo Wood. Revealing patterns and trends of mass mobility through spatial and temporal abstraction of origin-destination movement data. IEEE Transactions on Visualization and Computer Graphics, 23(9):2120– 2136, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/09/07587808-
                                        abs.html.

Andrienko:2019:AFV

[AAGS19] Natalia Andrienko, Gennady Andrienko, Jose Manuel Cordero Garcia, and David Scarlatti.

Analysis of flight variability: a systematic approach. IEEE Transactions on Visualization and Computer Graphics, 25 (1):54–64, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440121-
abs.html.

Andrienko:2013:SAM

[AAH^{+}13] Gennady Andrienko, Natalia Andrienko, Christophe Hurter, Salvatore Rinzivillo, and Stefan Wrobel. Scalable analysis of movement data for extracting and exploring significant places. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1078–1094, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrienko:2012:IPH

[AAM^{+}12] Gennady Andrienko, Natalia Andrienko, Martin Mladenov, Michael Mock, and Christian Politz. Identifying place histories from activity traces with an eye to parameter impact. IEEE Transactions on Visualization and Computer Graphics, 18(5):675–688, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alsallakh:2012:RCW

[AAMG12] Bilal Alsallakh, Wolfgang Aigner, Silvia Miksch, and M. Eduard Gröller. Reinventing the contingency wheel: Scalable visual analytics of large categorical data. IEEE Transactions on Visualization and Computer Graphics, 18(12):2849–2858, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alsallakh:2013:RSI

[AAMH13] Bilal Alsallakh, Wolfgang Aigner, Silvia Miksch, and Helwig Hauser. Radial sets: Interactive visual analysis of large overlapping sets. IEEE Transactions on Visualization and Computer Graphics, 19(12):2496–2505, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Asghar:2001:NMT

[AB01] M. W. Asghar and K. E. Barner. Nonlinear multiresolution techniques with applications to scientific visualization in a haptic environment. IEEE Transactions on Visualization and Computer Graphics, 7(1):76–??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0076.
pdf;   http://www.computer.
org/tvcg/tg2001/v0076abs.
htm.

Acar:2006:DMM

[AB06] Ruyam Acar and Pierre Boulanger. Digital marbling: a multiscale fluid model. IEEE Transactions on Visualization and Computer Graphics, 12(4):600– 614, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Angelini:2019:VVV

[ABC^{+}19] Marco Angelini, Graziano Blasilli, Tiziana Catarci, Simone Lenti, and Giuseppe Santucci. Vulnus: Visual vulnerability analysis for network security. IEEE Transactions on Visualization and Computer Graphics, 25(1): 183–192, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08443131-
abs.html.

Alexa:2003:CRP

[ABCO^{+}03] Marc Alexa, Johannes Behr, Daniel Cohen-Or, Shachar Fleishman, David Levin, and Claudio T. Silva. Computing and rendering point set surfaces. IEEE Transactions on Visualization and Computer Graphics, 9(1):3–15, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

                                        comp/trans/tg/2003/01/v0003abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/01/
                                        v0003.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/01/v0003.pdf.

Aksit:2019:MAD

[ACR^{+}19] Kaan Akşit, Praneeth Chakravarthula, Kishore Rathinavel, Youngmo Jeong, Rachel Albert, Henry Fuchs, and David Luebke. Manufacturing applicationdriven foveated near-eye displays. IEEE Transactions on Visualization and Computer Graphics, 25(5):1928– 1939, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642529/.

Alexander:2018:PBF

[ACS^{+}18] Eric Carlson Alexander, Chih-Ching Chang, Mariana Shimabukuro, Steven Franconeri, Christopher Collins, and Michael Gleicher. Perceptual biases in font size as a data encoding. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2397–2410, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/08/07968319-
                                        abs.html.

Antani:2012:DIA

[ACTM12] Lakulish Antani, Anish Chandak, Micah Taylor, and Dinesh Manocha. Direct-to-indirect acoustic radiance transfer. IEEE Transactions on Visualization and Computer Graphics, 18(2):261– 269, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ahrens:2012:GEI

[AD12] James Ahrens and Kurt Debattista. Guest Editor’s introduction: Special section on the Eurographics Symposium on Parallel Graphics and Visualization (EGPGV). IEEE Transactions on Visualization and Computer Graphics, 18 (1):3–4, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ament:2016:AAV

[AD16] M. Ament and C. Dachsbacher. Anisotropic ambient volume shading. IEEE Transactions on Visualization and Computer Graphics, 22 (1):1015–1024, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Asthana:2012:FPT

[ADDG12] Akshay Asthana, Miles Delahunty, Abhinav Dhall, and Roland Goecke. Facial performance transfer via deformable models and parametric correspondence. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1511–1519, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Albers:2011:SSL

[ADG11] Danielle Albers, Colin Dewey, and Michael Gleicher. Sequence surveyor: Leveraging overview for scalable genomic alignment visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2392–2401, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

AlBorno:2013:TOF

[AdLH13] Mazen Al Borno, Martin de Lasa, and Aaron Hertzmann. Trajectory optimization for full-body movements with complex contacts. IEEE Transactions on Visualization and Computer Graphics, 19 (8):1405–1414, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Assfalg:2002:TDI

[ADP02] Jürgen Assfalg, Alberto Del Bimbo, and Pietro Pala. Three-dimensional interfaces for querying by example in content-based image retrieval. IEEE Transactions on Visualization and Computer Graphics, 8(4):305–318, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0305abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0305.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0305.pdf.

Al-Dohuki:2017:SNA

[ADWK^{+}17] Shamal Al-Dohuki, Yingyu Wu, Farah Kamw, Jing Yang, Xin Li, Ye Zhao, Xinyue Ye, Wei Chen, Chao Ma, and Fei Wang. SemanticTraj: A new approach to interacting with massive taxi trajectories. IEEE Transactions on Visualization and Computer Graphics, 23(1):11–20, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Athawale:2013:UQL

[AE13] Tushar Athawale and Alireza Entezari. Uncertainty quantification in linear interpolation for isosurface extraction. IEEE Transactions on Visualization and Computer Graphics, 19(12):2723– 2732, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alim:2009:LBM

[AEM09] Usman R. Alim, Alireza Entezari, and Torsten Möller. The lattice-Boltzmann method on optimal sampling lattices. IEEE Transactions on Visualization and Computer Graphics, 15(4):630– 641, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Arietta:2014:CFU

[AERA14] Sean M. Arietta, Alexei A. Efros, Ravi Ramamoorthi, and Maneesh Agrawala. City forensics: Using visual elements to predict non-visual city attributes. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2624–2633, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875954-
abs.html.

Attene:2005:SBR

[AFRS05] Marco Attene, Bianca Falcidieno, Jarek Rossignac, and Michela Spagnuolo. Sharpen & bend: Recovering curved sharp edges in triangle meshes produced by featureinsensitive sampling. IEEE Transactions on Visualization and Computer Graphics, 11(2):181–192, March/April 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alcazar:2016:FAR

[AG16a] Juan G. Alcázar and Ron Goldman. Finding the axis of revolution of an algebraic surface of revolution. IEEE Transactions on Visualization and Computer Graphics, 22 (9):2082–2093, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alexander:2016:TDC

[AG16b] E. Alexander and M. Gleicher. Task-driven comparison of topic models. IEEE Transactions on Visualization and Computer Graphics, 22 (1):320–329, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alcazar:2017:DWI

[AG17] Juan Gerardo Alcazar and Ron Goldman. Detecting when an implicit equation or a rational parametrization defines a conical or cylindrical surface, or a surface of revolution. IEEE Transactions on Visualization and Computer Graphics, 23(12):2550– 2559, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07737012-
abs.html.

Anderson:2010:SVA

[AGDJ10] John C. Anderson, Christoph Garth, Mark A. Duchaineau, and Kenneth I. Joy. Smooth, volume-accurate material interface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 16(5):802–814, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ahn:2006:CRL

[AGL06] Minsu Ahn, Igor Guskov, and Seungyong Lee. Outof-core remeshing of large polygonal meshes. IEEE Transactions on Visualization and Computer Graphics, 12(5):1221–1228, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andreasen:2019:AFN

[AGN^{+}19] Anastassia Andreasen, Michele Geronazzo, Niels Christian Nilsson, Jelizaveta Zovnercuka, Kristian Konovalov, and Stefania Serafin. Auditory feedback for navigation with echoes in virtual environments: Training procedure and orientation strategies. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1876–1886, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8643846/.

Ahmed:2017:SPP

[AGY^{+}17] Abdalla G. M. Ahmed, Jianwei Guo, Dong-Ming Yan, Jean-Yves Franceschia, Xiaopeng Zhang, and Oliver Deussen. A simple push– pull algorithm for blue-noise sampling. IEEE Transactions on Visualization and Computer Graphics, 23(12):2496– 2508, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07790842-
abs.html.

Angelelli:2011:STF

[AH11] Paolo Angelelli and Helwig Hauser. Straightening tubular flow for side-by-side visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2063– 2070, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alsallakh:2014:VMA

[AHH^{+}14] Bilal Alsallakh, Allan Hanbury, Helwig Hauser, Silvia Miksch, and Andreas Rauber. Visual methods for analyzing probabilistic classification data. IEEE Transactions on Visualization and Computer Graphics, 20(12):1703– 1712, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875957-
abs.html.

Aldrich:2017:AVD

[AHK^{+}17] Garrett Aldrich, Jeffrey D. Hyman, Satish Karra, Carl W. Gable, Nataliia Makedonska, Hari Viswanathan, Jonathan Woodring, and Bernd Hamann. Analysis and visualization of discrete fracture networks using a flow topology graph. IEEE Transactions on Visualization and Computer Graphics, 23(8):1896– 1909, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07494624-
                                        abs.html.

Alper:2011:SHG

[AHKMF11] Basak Alper, Tobias Hollerer, JoAnn Kuchera-Morin, and Angus Forbes. Stereoscopic highlighting: 2D graph visualization on stereo displays. IEEE Transactions on Visualization and Computer Graphics, 17(12):2325– 2333, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Auber:2013:GUG

[AHL^{+}13] David Auber, Charles Huet, Antoine Lambert, Benjamin Renoust, Arnaud Sallaberry, and Agnes Saulnier. GosperMap: Using a Gosper curve for laying out hierarchical data. IEEE Transactions on Visualization and Computer Graphics, 19(11):1820–1832, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Amirkhanov:2011:PBM

[AHR^{+}11] Artem Amirkhanov, Christoph Heinzl, Michael Reiter, Johann Kastner, and Eduard Gröller. Projection-based metal-artifact reduction for industrial 3D X-ray computed tomography. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2193–2202, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Amirkhanov:2010:VOS

[AHRG10] Artem Amirkhanov, Christoph Heinzl, Michael Reiter, and Eduard Gröller. Visual optimality and stability analysis of 3DCT scan positions. IEEE Transactions on Visualization and Computer Graphics, 16(6):1477–1486, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Abello:2014:MDI

[AHSS14] James Abello, Steffen Hadlak, Heidrun Schumann, and Hans-Jörg Schulz. A modular degree-of-interest specification for the visual analysis of large dynamic networks. IEEE Transactions on Visualization and Computer Graphics, 20(3):337– 350, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Asayama:2018:FDV

[AIS18] H. Asayama, D. Iwai, and K. Sato. Fabricating diminishable visual markers for geometric registration in projection mapping. IEEE Transactions on Visualization and Computer Graphics, 24(2): 1091–1102, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andres:1997:DAH

[AJ97] E. Andres and M.-Andrée Jacob. The discrete analytical hyperspheres. IEEE Transactions on Visualization and Computer Graphics, 3(1):75– 86, January/March 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0075.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0075abs.
                                        htm.

Alam:2017:AET

[AJ17] Sayeed Safayet Alam and Radu Jianu. Analyzing eyetracking information in visualization and data space: From where on the screen to what on the screen. IEEE Transactions on Visualization and Computer Graphics, 23(5):1492–1505, May 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07420740-
abs.html.

Athawale:2019:PAD

[AJ19] Tushar Athawale and Chris R.

Johnson. Probabilistic asymptotic decider for topological ambiguity resolution in levelset extraction for uncertain 2D data. IEEE Transactions on Visualization and Computer Graphics, 25(1):1163– 1172, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440034-
abs.html.

Acevedo:2008:UVD

[AJDL08] Daniel Acevedo, Cullen D.

Jackson, Fritz Drury, and David H. Laidlaw. Using visual design experts in critique-based evaluation of 2D vector visualization methods. IEEE Transactions on Visualization and Computer Graphics, 14(4):877– 884, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Abello:2002:MSV

[AK02] James Abello and Jeffrey Korn. MGV: a system for visualizing massive multidigraphs. IEEE Transactions on Visualization and Computer Graphics, 8(1):21– 38, January 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0021.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0021abs.
                                        htm.

Kang:2011:HCV

[aKGS11] Youn ah Kang, Carsten Gorg, and John Stasko. How can visual analytics assist investigative analysis? design implications from an evaluation. IEEE Transactions on Visualization and Computer Graphics, 17(5):570–583, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kang:2012:EUV

[aKS12] Youn ah Kang and John Stasko. Examining the use of a visual analytics system for sensemaking tasks: Case studies with domain experts. IEEE Transactions on Visualization and Computer Graphics, 18(12):2869– 2878, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Acevedo:2006:SQP

[AL06] Daniel Acevedo and David Laidlaw. Subjective quantification of perceptual interactions among some 2D scientific visualization methods. IEEE Transactions on Visualization and Computer Graphics, 12(5):1133–1140, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ahn:2011:RTR

[AL11] Woojin Ahn and Doo Yong Lee. Real-time resolution of self-intersection in dynamic cylindrical free-form deformation. IEEE Transactions on Visualization and Computer Graphics, 17(4): 515–526, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Albo:2016:RCE

[ALBR16] Y. Albo, J. Lanir, P. Bak, and S. Rafaeli. Off the radar: Comparative evaluation of radial visualization solutions for composite indicators. IEEE Transactions on Visualization and Computer Graphics, 22(1):569–578, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Albuquerque:2011:SGH

[ALM11] Georgia Albuquerque, Thomas Lowe, and Marcus Magnor. Synthetic generation of highdimensional datasets. IEEE Transactions on Visualization and Computer Graphics, 17(12):2317–2324, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Arvanitis:2019:FPM

[ALMF19] G. Arvanitis, A. S. Lalos, K. Moustakas, and N. Fakotakis. Feature preserving mesh denoising based on graph spectral processing. IEEE Transactions on Visualization and Computer Graphics, 25(3):1513–1527, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrienko:2017:P

[ALS^{+}17] Gennady Andrienko, Shixia Liu, John Stasko, Niklas Elmqvist, Bongshin Lee, Kwan-Liu Ma, James Ahrens, Robert M. Kirby, and Jos Roerdink. Preface. IEEE Transactions on Visualization and Computer Graphics, 23 (1):xi–xv, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Antani:2013:APD

[AM13] Lakulish Antani and Dinesh Manocha. Aural proxies and directionally-varying reverberation for interactive sound propagation in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 19(4):

567–575, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Archambault:2006:SPF

[AMA06] Daniel Archambault, Tamara Munzner, and David Auber. Smashing peacocks further: Drawing quasi-trees from biconnected components. IEEE Transactions on Visualization and Computer Graphics, 12(5):813–820, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Archambault:2007:TMG

[AMA07] Daniel Archambault, Tamara Munzner, and David Auber. TopoLayout: Multilevel graph layout by topological features. IEEE Transactions on Visualization and Computer Graphics, 13(2):305– 317, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Archambault:2008:GSE

[AMA08] Daniel Archambault, Tamara Munzner, and David Auber. GrouseFlocks: Steerable exploration of graph hierarchy space. IEEE Transactions on Visualization and Computer Graphics, 14(4):900– 913, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Archambault:2011:TGF

[AMA11] Daniel Archambault, Tamara Munzner, and David Auber. Tugging graphs faster: Efficiently modifying pathpreserving hierarchies for browsing paths. IEEE Transactions on Visualization and Computer Graphics, 17(3): 276–289, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Auzinger:2013:VVU

[AMB^{+}13] Thomas Auzinger, Gabriel Mistelbauer, Ivan Baclija, Rudiger Schernthaner, Arnold Kochl, Michael Wimmer, M. Eduard Groller, and Stefan Bruckner. Vessel visualization using curved surface reformation. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2858–2867, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alim:2010:GER

[AMC10] Usman Alim, Torsten Moller, and Laurent Condat. Gradient estimation revitalized. IEEE Transactions on Visualization and Computer Graphics, 16(6):1495–1504, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Afzal:2012:STV

[AMJ^{+}12] Shehzad Afzal, Ross Maciejewski, Yun Jang, Niklas Elmqvist, and David S. Ebert. Spatial text visualization using automatic typographic maps. IEEE Transactions on Visualization and Computer Graphics, 18(12):2556– 2564, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Aigner:2008:VMA

[AMM^{+}08] Wolfgang Aigner, Silvia Miksch, Wolfgang Müller, Heidrun Schumann, and Christian Tominski. Visual methods for analyzing timeoriented data. IEEE Transactions on Visualization and Computer Graphics, 14(1):47– 60, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrews:2013:IPN

[AN13] Christopher Andrews and Chris North. The impact of physical navigation on spatial organization for sensemaking. IEEE Transactions on Visualization and Computer Graphics, 19(12):2207– 2216, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:1996:C

[Ano96a]  Anonymous.  Correction.

IEEE Transactions on Visualization and Computer Graphics, 2(3):280–??, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://dlib.computer.
                                        org/tg/books/tg1996/pdf/
                                        v0280.pdf.

Anonymous:1996:CNO

[Ano96b] Anonymous. Correction to “A Near Optimal Isosurface Extraction Algorithm Using the Span Space”. IEEE Transactions on Visualization and Computer Graphics, 2 (2):184, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0184.

pdf. See [LSJ96].

Anonymous:1997:I

[Ano97a]  Anonymous.  1997 index.

IEEE Transactions on Visualization and Computer Graphics, 3(4):381–384, October/December 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0381.
                                        pdf.

Anonymous:1997:E

[Ano97b] Anonymous. Editorial. IEEE Transactions on Visualization and Computer Graphics, 3(4):289–??, October/ December 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0289.
pdf.

Anonymous:1998:I

[Ano98] Anonymous. Index. IEEE Transactions on Visualization and Computer Graphics, 4(4):379–382, October/ December 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0379.
pdf.

Anonymous:1999:I

[Ano99] Anonymous. Index. IEEE Transactions on Visualization and Computer Graphics, 5(4):373–374, October/ December 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0373.
pdf.

Anonymous:2000:RL

[Ano00a] Anonymous. 1999 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 6(1):94–95, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0094.
                                        pdf.

Anonymous:2000:I

[Ano00b] Anonymous. Index. IEEE Transactions on Visualization and Computer Graphics, 6(4):381–??, October/ December 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0381.
                                        pdf.

Anonymous:2001:AIA

[Ano01a]  Anonymous.  Annual in-

dex. IEEE Transactions on Visualization and Computer Graphics, 7(4):380– 384, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2001/pdf/v0380.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0380abs.
                                        htm.

Anonymous:2001:RL

[Ano01b] Anonymous. Reviewers list 2000. IEEE Transactions on Visualization and Computer Graphics, 7(1):94–??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0094.
pdf.

Anonymous:2002:I

[Ano02]  Anonymous.  2002 index.

IEEE Transactions on Visualization and Computer Graphics, 8(4):395–399, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0395.
pdf;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0395.htm.

Anonymous:2003:RL

[Ano03a] Anonymous. 2002 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 9(1):111–112, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0111.
pdf;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0111.htm.

Anonymous:2003:AI

[Ano03b]  Anonymous.  Annual in-

dex. IEEE Transactions on Visualization and Computer Graphics, 9(4):587–592, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/04/v0587.
                                        pdf.

Anonymous:2004:AI

[Ano04a] Anonymous. Annual index. IEEE Transactions on Visualization and Computer Graphics, 10(6):730–736, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/06/v0730.
                                        pdf;   http://csdl.computer.
                                        org/dl/trans/tg/2004/06/
                                        v0730.htm.

Anonymous:2004:RL

[Ano04b] Anonymous. Reviewer’s list. IEEE Transactions on Visualization and Computer Graphics, 10(1):111–112, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/01/v0111.
                                        pdf.

Anonymous:2005:ENa

[Ano05a] Anonymous. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 11(1):1–??, January/February 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2005/01/v0001.

pdf;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0001.htm.

Anonymous:2005:ENb

[Ano05b] Anonymous. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 11(3):241–242, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2005/
03/v0241.pdf.

Anonymous:2005:LFS

[Ano05c] Anonymous. From the lab to the field: Steve Roth

— a memoriam. IEEE Transactions on Visualization and Computer Graphics, 11(6):609–610, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/06/v0609.
pdf.

Anonymous:2005:NTR

[Ano05d] Anonymous. A new technique for rendering complex portals. IEEE Transactions on Visualization and Computer Graphics, 11(1):81–90, January/February 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0081abs.

                                        htm;   http://csdl.computer.
                                        org/comp/trans/tg/2005/
                                        01/v0081s.zip;   http://
                                        csdl.computer.org/dl/trans/
                                        tg/2005/01/v0081.htm;
                                        http://csdl.computer.org/
                                        dl/trans/tg/2005/01/v0081.
                                        pdf.

Anonymous:2005:RL

[Ano05e] Anonymous. Reviewer’s list. IEEE Transactions on Visualization and Computer Graphics, 11(1):110–111, January/February 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2005/01/v0110.
                                        pdf.

Anonymous:2006:RL

[Ano06] Anonymous. 2005 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 12(1):126–128, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2006/01/v0126.
                                        pdf.

Anonymous:2007:RL

[Ano07] Anonymous. 2006 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 13(1):190–192, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/01/v0190.
pdf.

Anonymous:2008:RL

[Ano08a] Anonymous. 2007 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 14(1):242–244, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/01/ttg2008010242.
pdf.

Anonymous:2008:BMB

[Ano08b]  Anonymous.  Back mat-

ter: Back matter. IEEE Transactions on Visualization and Computer Graphics, 14(6):xxvii–xxviii, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2008:CP

[Ano08c] Anonymous. Call for participation. IEEE Transactions on Visualization and Computer Graphics, 14(2): 481, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/;   http://csdl.
computer.org/comp/trans/
tg/2008/02/ttg2008020481.
pdf.

Anonymous:2008:PP

[Ano08d] Anonymous. Pre-pages. IEEE Transactions on Visualization and Computer Graphics, 14(6):i–xxv, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2009:AIT

[Ano09a]  Anonymous.  Annual in-

dex: 2008 TVCG annual index. IEEE Transactions on Visualization and Computer Graphics, 15(1): INDEX:1–INDEX:23, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2009:HQS

[Ano09b] Anonymous. High-quality, semi-analytical volume rendering for AMR data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1611–1618, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2009:RLT

[Ano09c] Anonymous. Reviewers list: 2008 TVCG reviewers list. IEEE Transactions on Visualization and Computer Graphics, 15(1):174–176, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2009:TVIb

[Ano09d]  Anonymous.  TVCG Vis

and InfoVis 2009 author index. IEEE Transactions on Visualization and Computer Graphics, 15(6):xxviv–xxvv, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2009:TVIa

[Ano09e] Anonymous. TVCG Vis and InfoVis 2009 front matter. IEEE Transactions on Visualization and Computer Graphics, 15(6):i–XXVIII, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:AI

[Ano10a] Anonymous. 2009 annual index. IEEE Transactions on Visualization and Computer Graphics, 16(1): INDEX:1–INDEX:23, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:RL

[Ano10b] Anonymous. 2009 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 16(1):174–176, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:VCA

[Ano10c] Anonymous. The 2010 Visualization Career Award. IEEE Transactions on Visualization and Computer Graphics, 16 (6):xxii, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:VTA

[Ano10d] Anonymous. The 2010 Visualization Techinal Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 16(6):xxiii, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:AIC

[Ano10e]  Anonymous.  Author in-

dex and cover image credits. IEEE Transactions on Visualization and Computer Graphics, 16(6):xxvii–xxviii, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2010:TC

[Ano10f] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 16(6):i–x, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:AIa

[Ano11a] Anonymous. 2010 annual index. IEEE Transactions on Visualization and Computer Graphics, 17(1):[online only], January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:RL

[Ano11b] Anonymous. 2010 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 17(1):125–128, January 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:VCA

[Ano11c] Anonymous. The 2011 Visualization Career Award: Frits Post. IEEE Transactions on Visualization and Computer Graphics, 17(12):xxi, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:VTA

[Ano11d] Anonymous. The 2011 Visualization Technical Achievement Award: Daniel Keim. IEEE Transactions on Visualization and Computer Graphics, 17(12):xxii, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:AIb

[Ano11e]  Anonymous.  Author in-

dex. IEEE Transactions on Visualization and Computer Graphics, 17(12):xxv, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:BC

[Ano11f] Anonymous. [back cover]. IEEE Transactions on Visualization and Computer Graphics, 17(7):c4, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:CRS

[Ano11g] Anonymous. Committees, reviewers, and supporting organizations. IEEE Transactions on Visualization and Computer Graphics, 17(12): xii, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:FC

[Ano11h] Anonymous. [front cover]. IEEE Transactions on Visualization and Computer Graphics, 17(7):c1, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:IVC

[Ano11i] Anonymous. IEEE Visualization Conference and IEEE Information Visualization Conference Proceedings 2011 title page. IEEE Transactions on Visualization and Computer Graphics, 17(12):i–ii, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:IFC

[Ano11j] Anonymous. [inside front cover]. IEEE Transactions on Visualization and Computer Graphics, 17(7):c2, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:TC

[Ano11k] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 17(12):iii– viii, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2011:TIA

[Ano11l] Anonymous. TVCG information for authors. IEEE Transactions on Visualization and Computer Graphics, 17(7):c3, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:AIa

[Ano12a] Anonymous. 2011 annual index. IEEE Transactions on Visualization and Computer Graphics, 18(1):not in print, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:RL

[Ano12b] Anonymous. 2011 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 18(1):173–175, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:AIb

[Ano12c] Anonymous. Author index. IEEE Transactions on Visualization and Computer Graphics, 18(4):xv, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:AIc

[Ano12d]  Anonymous.  Author in-

dex. IEEE Transactions on Visualization and Computer Graphics, 18(12):xxv, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:CPI

[Ano12e] Anonymous. Call for participation: IEEE Virtual Reality 2012. IEEE Transactions on Visualization and Computer Graphics, 18(1):176, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:C

[Ano12f]  Anonymous.  Committees.

IEEE Transactions on Visualization and Computer Graphics, 18(4):vii–ix, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:GEI

[Ano12g] Anonymous. Guest Editors’ introduction: Special section on the IEEE Conference on Visual Analytics Science and Technology (VAST). IEEE Transactions on Visualization and Computer Graphics, 18 (5):660–661, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:IVG

[Ano12h] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 18 (12):xiii, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:IPC

[Ano12i] Anonymous. International program committees. IEEE Transactions on Visualization and Computer Graphics, 18 (12):xv–xvi, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:PR

[Ano12j] Anonymous. Paper reviewers.

IEEE Transactions on Visualization and Computer Graphics, 18(4):x, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:R

[Ano12k]  Anonymous.  Reviewers.

IEEE Transactions on Visualization and Computer Graphics, 18(12):xviii–xx, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:SC

[Ano12l] Anonymous. Steering committees. IEEE Transactions on Visualization and Computer Graphics, 18(12): xvii, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:S

[Ano12m]  Anonymous.  Supporters.

IEEE Transactions on Visualization and Computer Graphics, 18(12):xxiii, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:TCa

[Ano12n] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 18(4):iii–iv, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:TCb

[Ano12o] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 18(12):iiii– viii, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:TP

[Ano12p] Anonymous. Title page. IEEE Transactions on Visualization and Computer Graphics, 18 (4):i–ii, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2012:VCC

[Ano12q] Anonymous. VisWeek Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 18 (12):xiv, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:TAI

[Ano13a] Anonymous. 2012 TVCG annual index. IEEE Transactions on Visualization and Computer Graphics, 19(1):not in print, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:TRL

[Ano13b] Anonymous. 2012 TVCG reviewers list. IEEE Transactions on Visualization and Computer Graphics, 19(1): 173–175, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VVCa

[Ano13c] Anonymous. The 2012 VGTC Visualization Career Award: Ben Shneiderman. IEEE Transactions on Visualization and Computer Graphics, 19(12):xxiii, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VVTa

[Ano13d]  Anonymous.  The 2012

VGTC Visualization Technical Achievement Award: John Stasko. IEEE Transactions on Visualization and Computer Graphics, 19(12): xxiv, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VVCb

[Ano13e] Anonymous. The 2013 VGTC Visualization Career Award: Gregory M. Nielson. IEEE Transactions on Visualization and Computer Graphics, 19(12):xxv, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VVTb

[Ano13f]  Anonymous.  The 2013

VGTC Visualization Technical Achievement Award: Kwan-Liu Ma. IEEE Transactions on Visualization and Computer Graphics, 19(12): xxvi, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:AIa

[Ano13g] Anonymous. Author index.

IEEE Transactions on Visualization and Computer Graphics, 19(4):xix, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:AIb

[Ano13h] Anonymous. Author index.

IEEE Transactions on Visualization and Computer Graphics, 19(12):xxix–xxx, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:CC

[Ano13i]  Anonymous.  Conference

Committee. IEEE Transactions on Visualization and Computer Graphics, 19(4): viii, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:IVGa

[Ano13j] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 19 (4):vii, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:IVGb

[Ano13k] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 19 (12):xv, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:IPCa

[Ano13l] Anonymous. International Program Committee and Steering Committee. IEEE Transactions on Visualization and Computer Graphics, 19(4):ix, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:IPCb

[Ano13m] Anonymous. International Program Committees. IEEE Transactions on Visualization and Computer Graphics, 19(12):xvii–xviii, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:MEC

[Ano13n] Anonymous. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 19 (12):x, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:MPC

[Ano13o] Anonymous. Message from the Paper Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 19(12): xi–xiv, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:PR

[Ano13p] Anonymous. Paper reviewers.

IEEE Transactions on Visualization and Computer Graphics, 19(4):x–xi, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:R

[Ano13q]  Anonymous.  Reviewers.

IEEE Transactions on Visualization and Computer Graphics, 19(12):xx–xxii, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:SC

[Ano13r] Anonymous. Steering Committees. IEEE Transactions on Visualization and Computer Graphics, 19(12): xix, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:TC

[Ano13s] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 19(12):iii– ix, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VCS

[Ano13t] Anonymous. VIS 2013 capstone speaker: Jarke van Wijk. IEEE Transactions on Visualization and Computer Graphics, 19(12):xxviii, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VKS

[Ano13u]  Anonymous.  VIS 2013

keynote speaker: Erez Lieberman Aiden. IEEE Transactions on Visualization and Computer Graphics, 19(12): xxvii, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2013:VCC

[Ano13v] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 19(12): xvi, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:AI

[Ano14a] Anonymous. 2013 annual index. IEEE Transactions on Visualization and Computer Graphics, 20(1):not in print, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:RL

[Ano14b] Anonymous. 2013 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 20(1):155–158, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:IIT

[Ano14c] Anonymous. 2014 index IEEE Transactions on Visualization and Computer Graphics vol. 20. IEEE Transactions on Visualization and Computer Graphics, 20(12): xxvi–xxvii, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06935060.
pdf.

Anonymous:2014:VRC

[Ano14d] Anonymous. The 2014 virtual reality career award: Steve feiner. IEEE Transactions on Visualization and Computer Graphics, 20(4):xiii, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:VRT

[Ano14e] Anonymous. The 2014 Virtual Reality Technical Achievement Award: Doug A. Bowman. IEEE Transactions on Visualization and Computer Graphics, 20(4):xiv, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:VTAb

[Ano14f] Anonymous. The 2014 Visualization Technical Achievement Award: Claudio T. Silva. IEEE Transactions on Visualization and Computer Graphics, 20(12):xxv, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935102.
                                        pdf.

Anonymous:2014:VTAa

[Ano14g] Anonymous. The 2014 Visualization Technical Achievement Award: Ken Joy. IEEE Transactions on Visualization and Computer Graphics, 20(12):xxiv, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935101.
                                        pdf.

Anonymous:2014:EVP

[Ano14h]  Anonymous.  Erratum to

“Video Painting Based on a Stabilized Time-Varying Flow Field”. IEEE Transactions on Visualization and Computer Graphics, 20(2):316, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See [YLK12].

Anonymous:2014:FC

[Ano14i] Anonymous. [front cover].

IEEE Transactions on Visualization and Computer Graphics, 20(12):i–Bii, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06935065.
pdf.

Anonymous:2014:ITV

[Ano14j] Anonymous. IEEE Transactions on Visualization and Computer Graphics — 2014 IEEE Virtual Reality Conference [title page]. IEEE Transactions on Visualization and Computer Graphics, 20 (4):i–ii, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:IVGa

[Ano14k] Anonymous. IEEE Visualization and Graphics Technical Committee. IEEE Transactions on Visualization and Computer Graphics, 20(4):ix, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:IVGb

[Ano14l] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 20 (12):xv, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935062.
                                        pdf.

Anonymous:2014:IPC

[Ano14m] Anonymous. International program committee. IEEE Transactions on Visualization and Computer Graphics, 20 (4):xi, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:MEC

[Ano14n] Anonymous. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 20 (12):x, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935055.
                                        pdf.

Anonymous:2014:MVP

[Ano14o] Anonymous. Message from the VIS paper chairs Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 20(12):xi– xiv, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06935059.
pdf.

Anonymous:2014:RSC

[Ano14p] Anonymous. Rock stars of cybersecurity conference. IEEE Transactions on Visualization and Computer Graphics, 20 (7):1083, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:TCa

[Ano14q] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 20(4):iii–iv, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2014:TCb

[Ano14r] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 20(12):iii– xi, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06935069.
pdf.

Anonymous:2014:VCCb

[Ano14s] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 20(12): xvi, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935061.
                                        pdf.

Anonymous:2014:VIP

[Ano14t]  Anonymous.  VIS Inter-

national Program Committees. IEEE Transactions on Visualization and Computer Graphics, 20(12):xvii– xviii, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935057.
                                        pdf.

Anonymous:2014:VR

[Ano14u] Anonymous. VIS reviewers. IEEE Transactions on Visualization and Computer Graphics, 20(12):xx– xxiii, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06935066.
                                        pdf.

Anonymous:2014:VSE

[Ano14v] Anonymous. VIS Steering and Executive Committees. IEEE Transactions on Visualization and Computer Graphics, 20 (12):xix, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06935063.
pdf.

Anonymous:2014:VCCa

[Ano14w] Anonymous. VR Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 20(4):x, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:IIT

[Ano15a] Anonymous. 2014 index IEEE Transactions on Visualization and Computer Graphics vol. 20. IEEE Transactions on Visualization and Computer Graphics, 21(1): 140–171, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06966853.
pdf.

Anonymous:2015:RLa

[Ano15b] Anonymous. 2014 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 21(1):137– 139, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/01/06966856.
                                        pdf.

Anonymous:2015:VVR

[Ano15c] Anonymous. 2015 VGTC Virtual Reality Career Award. IEEE Transactions on Visualization and Computer Graphics, 21(4):xii, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:C

[Ano15d]  Anonymous.  Committees.

IEEE Transactions on Visualization and Computer Graphics, 21(4):vii–x, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:CAI

[Ano15e] Anonymous. Conference author index. IEEE Transactions on Visualization and Computer Graphics, 21(4): xiii, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:FCa

[Ano15f] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 21(4):i–ii, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:FCb

[Ano15g] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 21(11):i–Bii, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/csdl/trans/tg/2015/
11/07283728.pdf.

Anonymous:2015:ICC

[Ano15h] Anonymous. ISMAR 2015 Conference Committee members. IEEE Transactions on Visualization and Computer Graphics, 21(11):ix– x, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07283729.
pdf.

Anonymous:2015:IEM

[Ano15i] Anonymous. ISMAR 2015 EiC message. IEEE Transactions on Visualization and Computer Graphics, 21(11): v, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07283726.
pdf.

Anonymous:2015:IPW

[Ano15j] Anonymous. ISMAR 2015 PC welcome message. IEEE Transactions on Visualization and Computer Graphics, 21 (11):vi–viii, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/11/07283689.
                                        pdf.

Anonymous:2015:ISC

[Ano15k] Anonymous. ISMAR 2015 Steering Committee members. IEEE Transactions on Visualization and Computer Graphics, 21(11):xi, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/11/07283724.
                                        pdf.

Anonymous:2015:MEC

[Ano15l] Anonymous. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 21 (4):v, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:MVP

[Ano15m] Anonymous. Message from the VR Program Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 21(4):vi, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:RLb

[Ano15n]  Anonymous.  Reviewer’s

list. IEEE Transactions on Visualization and Computer Graphics, 21(11):xii, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07283722.
pdf.

Anonymous:2015:TC

[Ano15o] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 21(4):iii–iv, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2015:VR

[Ano15p] Anonymous. VR reviewers.

IEEE Transactions on Visualization and Computer Graphics, 21(4):xi, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:IIT

[Ano16a] Anonymous. 2015 index IEEE Transactions on Visualization and Computer Graphics vol.

21. IEEE Transactions on Visualization and Computer Graphics, 22(2):1–16, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VCA

[Ano16b] Anonymous. The 2015 Visualization Career Award. IEEE Transactions on Visualization and Computer Graphics, 22 (1):xxv, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VTA

[Ano16c] Anonymous. The 2015 Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 22(1): xxvi, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VVRa

[Ano16d] Anonymous. The 2016 VGTC Virtual Reality Career Award. IEEE Transactions on Visualization and Computer Graphics, 22(4):xi, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VVRb

[Ano16e] Anonymous. The 2016 VGTC Virtual Reality Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 22 (4):xii, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:AII

[Ano16f] Anonymous. Author index IEEE Transactions on Visualization and Computer Graphics vol. 20. IEEE Transactions on Visualization and Computer Graphics, 22(1):xxvii– xxviii, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:CAI

[Ano16g] Anonymous. Conference author index. IEEE Transactions on Visualization and Computer Graphics, 22(4): xiii, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:CC

[Ano16h] Anonymous. Conference committee. IEEE Transactions on Visualization and Computer Graphics, 22(4):viii, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:FCa

[Ano16i] Anonymous. Front cover.

IEEE Transactions on Visualization and Computer Graphics, 22(1):i–ii, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:FCb

[Ano16j] Anonymous. Front cover.

IEEE Transactions on Visualization and Computer Graphics, 22(4):i–Bii, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:IVGa

[Ano16k] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 22 (1):xvi, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:IVGb

[Ano16l] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 22 (4):vii, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:IPC

[Ano16m] Anonymous. International program committees. IEEE Transactions on Visualization and Computer Graphics, 22 (4):ix, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:MECa

[Ano16n] Anonymous. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 22 (1):x, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:MECb

[Ano16o] Anonymous. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 22 (4):v, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:MVPa

[Ano16p] Anonymous. Message from the VIS paper chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 22(1):xi– xv, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:MVPb

[Ano16q] Anonymous. Message from the VR paper chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 22(4): vi, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:PR

[Ano16r] Anonymous. Papers reviewers. IEEE Transactions on Visualization and Computer Graphics, 22(4):x, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:TCa

[Ano16s] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 22(1):iii–ix, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:TCb

[Ano16t] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 22(4):iii–iv, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VPR

[Ano16u] Anonymous. VAST paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 22(1): xxi–xxiv, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VCC

[Ano16v] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 22(1): xvii, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VIP

[Ano16w] Anonymous. VIS International Program Committees.

IEEE Transactions on Visualization and Computer Graphics, 22(1):xviii–xix, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2016:VSE

[Ano16x] Anonymous. VIS Steering and Executive Committees. IEEE Transactions on Visualization and Computer Graphics, 22 (1):xx, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:RLa

[Ano17a] Anonymous. 2016 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 23(3):1269– 1274, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07835768.
pdf.

Anonymous:2017:RLb

[Ano17b] Anonymous. 2017 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 23(12):2663– 2666, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/08089788.
pdf.

Anonymous:2017:AIa

[Ano17c] Anonymous. Author index. IEEE Transactions on Visualization and Computer Graphics, 23(1):xxvi–xxvii, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:AIb

[Ano17d] Anonymous. Author index. IEEE Transactions on Visualization and Computer Graphics, 23(4):xiii, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/04/07876889.
                                        pdf.

Anonymous:2017:ITV

[Ano17e] Anonymous. IEEE Transactions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 23 (1):C2, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:Cb

[Ano17f]  Anonymous.  Committee.

IEEE Transactions on Visualization and Computer Graphics, 23(1):xviii–xix, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:Cc

[Ano17g]  Anonymous.  Committee.

IEEE Transactions on Visualization and Computer Graphics, 23(1):xx, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:CCa

[Ano17h]  Anonymous.  Conference

Committee. IEEE Transactions on Visualization and Computer Graphics, 23(4): viii, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876896.
pdf.

Anonymous:2017:Ca

[Ano17i] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 23 (1):iii–ix, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:Cd

[Ano17j] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 23 (4):iii–iv, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876895.
pdf.

Anonymous:2017:I

[Ano17k] Anonymous. IEEE. IEEE Transactions on Visualization and Computer Graphics, 23 (4):i–ii, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07878333.
pdf.

Anonymous:2017:IPC

[Ano17l] Anonymous. International Program Committee. IEEE Transactions on Visualization and Computer Graphics, 23(4):ix, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876899.
pdf.

Anonymous:2017:PRa

[Ano17m] Anonymous. Paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 23(1):xxi–xxiii, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2017:PRb

[Ano17n]  Anonymous.  Papers re-

viewers. IEEE Transactions on Visualization and Computer Graphics, 23(4): x–xi, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876897.
pdf.

Anonymous:2017:P

[Ano17o] Anonymous. Preface. IEEE Transactions on Visualization and Computer Graphics, 23(4):vi, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876888.
pdf.

Anonymous:2017:VCC

[Ano17p] Anonymous. VIS conference committee. IEEE Transactions on Visualization and Computer Graphics, 23(1): xvii, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:IIT

[Ano18a]  Anonymous.  2017 index

IEEE Transactions on Visualization and Computer Graphics vol. 23. IEEE Transactions on Visualization and Computer Graphics, 24(2):1223– 1251, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:ITV

[Ano18b] Anonymous. IEEE Transactions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 24 (4):i–ii, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315165.
pdf.

Anonymous:2018:CC

[Ano18c]  Anonymous.  Conference

committee. IEEE Transactions on Visualization and Computer Graphics, 24(4): ix, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315168.
pdf.

Anonymous:2018:Ca

[Ano18d] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 24 (1):iii–ix, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:Cb

[Ano18e] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 24 (4):iii–iv, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315162.
pdf.

Anonymous:2018:IPC

[Ano18f] Anonymous. International program committee for journal papers. IEEE Transactions on Visualization and Computer Graphics, 24(4): x, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315169.
pdf.

Anonymous:2018:PR

[Ano18g]  Anonymous.  Papers re-

viewers. IEEE Transactions on Visualization and Computer Graphics, 24(4): xi–xii, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315157.
pdf.

Anonymous:2018:VIP

[Ano18h] Anonymous. VAST International Program Committee. IEEE Transactions on Visualization and Computer Graphics, 24(1):xviii–xix, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:VPR

[Ano18i] Anonymous. VAST Paper Reviewers. IEEE Transactions on Visualization and Computer Graphics, 24(1): xxi–xxv, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:VSC

[Ano18j] Anonymous. VAST Steering Committee. IEEE Transactions on Visualization and Computer Graphics, 24(1): xx, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2018:VCC

[Ano18k] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 24(1): xvii, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2019:IIT

[Ano19a] Anonymous. 2018 Index IEEE Transactions on Visualization and Computer Graphics vol. 24. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1–37, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/02/08594687.
                                        pdf.

Anonymous:2019:RLa

[Ano19b] Anonymous. 2018 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2331–2335, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2019:RLb

[Ano19c] Anonymous. 2019 reviewers list. IEEE Transactions on Visualization and Computer Graphics, 25(12):3258– 3261, December 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2019:Ca

[Ano19d]  Anonymous.  Committee.

IEEE Transactions on Visualization and Computer Graphics, 25(1):xx, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576486.
pdf.

Anonymous:2019:Cb

[Ano19e]  Anonymous.  Committee.

IEEE Transactions on Visualization and Computer Graphics, 25(1):xxiii, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08575978.
pdf.

Anonymous:2019:ITV

[Ano19f] Anonymous. IEEE Transactions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 25 (1):i, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576485.
pdf.

Anonymous:2019:IIP

[Ano19g] Anonymous. InfoVis International Program Committee. IEEE Transactions on Visualization and Computer Graphics, 25(1):xxi, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576492.
pdf.

Anonymous:2019:IPR

[Ano19h] Anonymous. InfoVis paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 25(1):xxv– xxvi, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576489.
pdf.

Anonymous:2019:SIP

[Ano19i] Anonymous. SciVis International Program Committee.

IEEE Transactions on Visualization and Computer Graphics, 25(1):xxii, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576494.
pdf.

Anonymous:2019:SPR

[Ano19j] Anonymous. SciVis paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 25(1): xxvii, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576493.
pdf.

Anonymous:2019:TC

[Ano19k] Anonymous. Table of content.

IEEE Transactions on Visualization and Computer Graphics, 25(1):ii–x, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08576487.
pdf.

Anonymous:2019:VPR

[Ano19l] Anonymous. VAST paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 25(1): xxiv, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08576488.
                                        pdf.

Anonymous:2019:VCC

[Ano19m] Anonymous. VIS conference committee. IEEE Transactions on Visualization and Computer Graphics, 25(1): xix, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08576490.
                                        pdf.

Adams:2018:LRP

[ANR^{+}18] Haley Adams, Gayathri Narasimham, John Rieser, Sarah Creem-Regehr, Jeanine Stefanucci, and Bobby Bodenheimer. Locomotive recalibration and prism adaptation of children and teens in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1408–1417, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/04/08267487-
                                        abs.html.

Archambault:2011:ASM

[APP11] Daniel Archambault, Helen C. Purchase, and Bruno Pinaud. Animation, small multiples, and the effect of mental map preservation in dynamic graphs. IEEE Transactions on Visualization and Computer Graphics, 17(4): 539–552, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Arikan:2014:LSP

[APS^{+}14] Murat Arikan, Reinhold Preiner, Claus Scheiblauer, Stefan Jeschke, and Michael Wimmer. Large-scale pointcloud visualization through localized textured surface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 20(9): 1280–1292, September 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Arth:2015:IOL

[APV^{+}15] Clemens Arth, Christian Pirchheim, Jonathan Ventura, Dieter Schmalstieg, and Vincent Lepetit. Instant outdoor localization and SLAM initialization from 2.5D maps. IEEE Transactions on Visualization and Computer Graphics, 21(11): 1309–1318, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07164332-
abs.html.

Arikan:2016:MDM

[APW16] M. Arikan, R. Preiner, and M. Wimmer. Multi-depth-map raytracing for efficient large-scene reconstruction. IEEE Transactions on Visualization and Computer Graphics, 22(2):1127–1137, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alsallakh:2017:PCV

[AR17] Bilal Alsallakh and Liu Ren. PowerSet: A comprehensive visualization of set intersections. IEEE Transactions on Visualization and Computer Graphics, 23(1):361–370, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Aliaga:2007:SGI

[ARB07] Daniel G. Aliaga, Paul A. Rosen, and Daniel R. Bekins. Style grammars for interactive visualization of architecture. IEEE Transactions on Visualization and Computer Graphics, 13(4):786– 797, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Amini:2015:IIC

[ARH^{+}15] Fereshteh Amini, Sebastien Rufiange, Zahid Hossain, Quentin Ventura, Pourang Irani, and Michael J. McGuffin. The impact of interactivity on comprehending 2D and 3D visualizations of movement data. IEEE Transactions on Visualization and Computer Graphics, 21(1): 122–135, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06826569-
abs.html.

Amini:2017:ADD

[ARL^{+}17] Fereshteh Amini, Nathalie Henry Riche, Bongshin Lee, Andres Monroy-Hernandez, and Pourang Irani. Authoring data-driven videos with Data-Clips. IEEE Transactions on Visualization and Computer Graphics, 23(1):501–510, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Alper:2011:DSL

[ARRC11] Basak Alper, Nathalie Riche, Gonzalo Ramos, and Mary Czerwinski. Design study of LineSets, a novel set visualization technique. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2259–2267, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Avidan:1998:NVS

[AS98] S. Avidan and A. Shashua.

Novel view synthesis by cascading trilinear tensors. IEEE Transactions on Visualization and Computer Graphics, 4(4):293–306, October/December 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0293.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0293abs.
                                        htm.

Amar:2005:KPD

[AS05] Robert A. Amar and John T. Stasko. Knowledge precepts for design and evaluation of information visualizations. IEEE Transactions on Visualization and Computer Graphics, 11(4):432– 442, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Antley:2011:ELS

[AS11] Angus Antley and Mel Slater. The effect on lower spine muscle activation of walking on a narrow beam in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 17(2): 255–259, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ankele:2019:DMS

[AS19] Michael Ankele and Thomas Schultz. DT-MRI streamsurfaces revisited. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1112–1121, January 2019.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440056-
abs.html.

Ament:2014:LPF

[ASDW14] Marco Ament, Filip Sadlo, Carsten Dachsbacher, and Daniel Weiskopf. Lowpass filtered volumetric shadows. IEEE Transactions on Visualization and Computer Graphics, 20(12):2437– 2446, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875905-
abs.html.

Athawale:2016:IVD

[ASE16] T. Athawale, E. Sakhaee, and A. Entezari. Isosurface visualization of data with nonparametric models for uncertainty. IEEE Transactions on Visualization and Computer Graphics, 22(1):777–786, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Aydin:2015:AAA

[ASG15] Tunc Ozan Aydin, Aljoscha Smolic, and Markus Gross. Automated aesthetic analysis of photographic images. IEEE Transactions on Visualization and Computer Graphics, 21 (1):31–42, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/01/06819054-
                                        abs.html.

Arbesser:2017:VVD

[ASMP17] Clemens Arbesser, Florian Spechtenhauser, Thomas Mühlbacher, and Harald Piringer. Visplause: Visual data quality assessment of many time series using plausibility checks. IEEE Transactions on Visualization and Computer Graphics, 23(1):641–650, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Agrawal:2014:DMC

[ASvdP14] Shailen Agrawal, Shuo Shen, and Michiel van de Panne. Diverse motions and character shapes for simulated skills. IEEE Transactions on Visualization and Computer Graphics, 20(10):1345– 1355, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/10/06781622-
                                        abs.html.

Angus:2012:CRP

[ASW12] Daniel Angus, Andrew Smith, and Janet Wiles. Conceptual recurrence plots: Revealing patterns in human discourse. IEEE Transactions on Visualization and Computer Graphics, 18(6):988–997, June 2012.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ament:2013:AVS

[ASW13] Marco Ament, Filip Sadlo, and Daniel Weiskopf. Ambient volume scattering. IEEE Transactions on Visualization and Computer Graphics, 19(12):2936–2945, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Agam:2005:SFA

[AT05] Gady Agam and Xiaojing Tang. A sampling framework for accurate curvature estimation in discrete surfaces. IEEE Transactions on Visualization and Computer Graphics, 11(5):573–583, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anand:2016:ASP

[AT16] A. Anand and J. Talbot. Automatic selection of partitioning variables for small multiple displays. IEEE Transactions on Visualization and Computer Graphics, 22(1): 669–677, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andrews:2016:BLM

[ATK16] S. Andrews, M. Teichmann, and P. G. Kry. Blended linear models for reduced compliant mechanical systems. IEEE Transactions on Visualization and Computer Graphics, 22 (3):1209–1222, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Au:2006:DLE

[ATLF06] Oscar Kin-Chung Au, Chiew-Lan Tai, Ligang Liu, and Hongbo Fu. Dual Laplacian editing for meshes. IEEE Transactions on Visualization and Computer Graphics, 12 (3):386–395, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ando:2012:PFS

[ATT12] Ryoichi Ando, Nils Thurey, and Reiji Tsuruno. Preserving fluid sheets with adaptively sampled anisotropic particles. IEEE Transactions on Visualization and Computer Graphics, 18(8):1202–1214, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Abello:2006:AGL

[AvHK06] James Abello, Frank van Ham, and Neeraj Krishnan. ASK-GraphView: a large scale graph visualization system. IEEE Transactions on Visualization and Computer Graphics, 12(5): 669–676, September/October 2006. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Andersen:2019:AHG

[AVP19] D. Andersen, P. Villano, and V. Popescu. AR HMD guidance for controlled hand-held 3D acquisition. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3073–3082, November 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ashraf:2003:SRC

[AW03] Golam Ashraf and Kok Cheong Wong. Semantic representation and correspondence for state-based motion transition. IEEE Transactions on Visualization and Computer Graphics, 9(4):481–499, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0481abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0481.pdf.

Armstrong:2014:VSM

[AW14] Zan Armstrong and Martin Wattenberg. Visualizing statistical mix effects and Simpson’s paradox. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2132–2141, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875927-
                                        abs.html.

Arbree:2011:HSS

[AWB11] A. Arbree, B. Walter, and K. Bala. Heterogeneous subsurface scattering using the finite element method. IEEE Transactions on Visualization and Computer Graphics, 17 (7):956–969, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ament:2010:DIV

[AWC10] Marco Ament, Daniel Weiskopf, and Hamish Carr. Direct interval volume visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1505–1514, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Abbasloo:2016:VTN

[AWHS16] A. Abbasloo, V. Wiens, M. Hermann, and T. Schultz. Visualizing tensor normal distributions at multiple levels of detail. IEEE Transactions on Visualization and Computer Graphics, 22(1):975–984, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ali:2009:CFD

[AYRW09] Saif Ali, Jieping Ye, Anshuman Razdan, and Peter Wonka. Compressed facade displacement maps. IEEE Transactions on Visualization and Computer Graphics, 15(2):262–273, March/April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Au:2012:MSC

[AZC^{+}12] Oscar Kin-Chung Au, Youyi Zheng, Menglin Chen, Pengfei Xu, and Chiew-Lan Tai. Mesh segmentation with concavityaware fields. IEEE Transactions on Visualization and Computer Graphics, 18(7): 1125–1134, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ament:2017:EOV

[AZD17] Marco Ament, Tobias Zirr, and Carsten Dachsbacher. Extinction-optimized volume illumination. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1767–1781, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07470264-
abs.html.

Alspaugh:2019:FMI

[AZL^{+}19] Sara Alspaugh, Nava Zokaei, Andrea Liu, Cindy Jin, and Marti A. Hearst. Futzing and moseying: Interviews with professional data analysts on exploration practices. IEEE Transactions on Visualization and Computer Graphics, 25 (1):22–31, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440815-
                                        abs.html.

Ahmed:2012:HCV

[AZM12] Nafees Ahmed, Ziyi Zheng, and Klaus Mueller. Human computation in visualization: Using purpose driven games for robust evaluation of visualization algorithms. IEEE Transactions on Visualization and Computer Graphics, 18(12):2104–2113, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baerentzen:2005:SDC

[BA05] J. Andreas Baerentzen and Henrik Aanaes. Signed distance computation using the angle weighted pseudonormal. IEEE Transactions on Visualization and Computer Graphics, 11(3):243–253, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beyer:2013:CQG

[BAAK^{+}13] Johanna Beyer, Ali Al-Awami, Narayanan Kasthuri, Jeff W. Lichtman, Hanspeter Pfister, and Markus Hadwiger. ConnectomeExplorer:

Query-guided visual analysis of large volumetric neuroscience data. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2868–2877, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bladin:2018:GBC

[BAB^{+}18] K. Bladin, E. Axelsson, E. Broberg, C. Emmart, P. Ljung, A. Bock, and A. Ynnerman. Globe browsing: Contextualized spatiotemporal planetary surface visualization. IEEE Transactions on Visualization and Computer Graphics, 24(1): 802–811, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bachmaier:2007:RAS

[Bac07] Christian Bachmaier. A radial adaptation of the sugiyama framework for visualizing hierarchical information. IEEE Transactions on Visualization and Computer Graphics, 13 (3):583–594, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bogl:2013:VAM

[BAF^{+}13] Markus Bogl, Wolfgang Aigner, Peter Filzmoser, Tim Lammarsch, Silvia Miksch, and Alexander Rind. Visual analytics for model selection in time series analysis. IEEE Transactions on Visualization and Computer Graphics, 19(12):2237–2246, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bailenson:2013:KSI

[Bai13] Jeremy Bailenson. Keynote speaker: Infinite reality: Avatars, eternal life, new worlds, and the dawn of the virtual revolution. IEEE Transactions on Visualization and Computer Graphics, 19 (4):xiv, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bergstrom:2017:PSQ

[BAP^{+}17] Ilias Bergstrom, Sergio Azevedo, Panos Papiotis, Nuno Saldanha, and Mel Slater. The plausibility of a string quartet performance in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 23(4):1352– 1359, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/04/07835714-
                                        abs.html.

Barsky:2005:MJW

[Bar05] Brian A. Barsky. In memoriam: Jane Wilhelms. IEEE Transactions on Visualization and Computer Graphics, 11(5):481–482, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/05/v0481.
pdf.

Borgo:2012:ESU

[BARM^{+}12] Rita Borgo, Alfie Abdul-Rahman, Farhan Mohamed, Philip W. Grant, Irene Reppa, Luciano Floridi, and Min Chen. An empirical study on using visual embellishments in visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2759–2768, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Byrne:2016:ACM

[BAW16] L. Byrne, D. Angus, and J. Wiles. Acquired codes of meaning in data visualization and infographics: Beyond perceptual primitives. IEEE Transactions on Visualization and Computer Graphics, 22(1):509–518, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Barth:2007:VRF

[BB07] William Barth and Christopher Burns. Virtual rheoscopic fluids for flow visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6):1751–1758, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1751s.
                                        mov.

Banks:2009:DII

[BB09] David C. Banks and Kevin Beason. Decoupling illumination from isosurface generation using 4D light transport. IEEE Transactions on Visualization and Computer Graphics, 15(6):1595–1602, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baudel:2012:CDS

[BB12] Thomas Baudel and Bertjan Broeskema. Capturing the design space of sequential spacefilling layouts. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2593–2602, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boukhayma:2019:SMC

[BB19] Adnane Boukhayma and Edmond Boyer. Surface motion capture animation synthesis. IEEE Transactions on Visualization and Computer Graphics, 25(6):2270– 2283, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8352750/.

Bramon:2012:MDF

[BBB^{+}12] Roger Bramon, Imma Boada, Anton Bardera, Joaquim Rodriguez, Miquel Feixas, Josep Puig, and Mateu Sbert. Multimodal data fusion based on mutual information. IEEE Transactions on Visualization and Computer Graphics, 18(9):1574–1587, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blascheck:2019:GVS

[BBB^{+}19] Tanja Blascheck, Lonni Besancon, Anastasia Bezerianos, Bongshin Lee, and Petra Isenberg. Glanceable visualization: Studies of data comparison performance on smartwatches. IEEE Transactions on Visualization and Computer Graphics, 25(1):630– 640, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08443125-
abs.html.

Barreira:2018:CAM

[BBBM18] João Barreira, Maximino Bessa, Luís Barbosa, and Luís Magalhães. A context-aware method for authentically simulating outdoors shadows for mobile augmented reality. IEEE Transactions on Visualization and Computer Graphics, 24(3):1223–1231, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bai:2015:UAR

[BBC15] Zhen Bai, A. F. Blackwell, and G. Coulouris. Using augmented reality to elicit pretend play for children with autism. IEEE Transactions on Visualization and Computer Graphics, 21(5):598– 610, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bottger:2006:CLV

[BBD06] Joachim Böttger, Michael Balzer, and Oliver Deussen. Complex logarithmic views for small details in large contexts. IEEE Transactions on Visualization and Computer Graphics, 12(5): 845–852, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Batagelj:2011:VAL

[BBD^{+}11] Vladimir Batagelj, Franz J. Brandenburg, Walter Didimo, Giuseppe Liotta, Pietro Palladino, and Maurizio Patrignani. Visual analysis of large graphs using (X, Y)- clustering and hybrid visualizations. IEEE Transactions on Visualization and Computer Graphics, 17(11):1587– 1598, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blaas:2009:SGV

[BBG^{+}09] Jorik Blaas, Charl Botha, Edward Grundy, Mark Jones, Robert Laramee, and Frits Post. Smooth graphs for visual exploration of higherorder state transitions. IEEE Transactions on Visualization and Computer Graphics, 15(6):969–976, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhargava:2018:EML

[BBG^{+}18] Ayush Bhargava, Jeffrey W.

Bertrand, Anand K. Gramopadhye, Kapil C. Madathil, and Sabarish V. Babu. Evaluating multiple levels of an interaction fidelity continuum on performance and learning in near-field training simulations. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1418–1427, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260967-
abs.html.

Behrisch:2017:MIB

[BBH^{+}17] Michael Behrisch, Benjamin Bach, Michael Hund, Michael Delz, Laura Von Rüden, Jean-Daniel Fekete, and Tobias Schreck. Magnostics: Imagebased search of interesting matrix views for guided network exploration. IEEE Transactions on Visualization and Computer Graphics, 23 (1):31–40, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baxter:2009:CES

[BBiA09] William V. Baxter III, Pascal Barla, and Ken ichi Anjyo. Compatible embedding for 2D shape animation. IEEE Transactions on Visualization and Computer Graphics, 15(5):867–879, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boukhelifa:2012:ESV

[BBIF12] Nadia Boukhelifa, Anastasia Bezerianos, Tobias Isenberg, and Jean-Daniel Fekete. Evaluating sketchiness as a visual variable for the depiction of qualitative uncertainty. IEEE Transactions on Visualization and Computer Graphics, 18(12):2769– 2778, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bronstein:2007:CNS

[BBK07] Alexander Bronstein, Michael Bronstein, and Ron Kimmel. Calculus of nonrigid surfaces for geometry and texture manipulation. IEEE Transactions on Visualization and Computer Graphics, 13(5): 902–913, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borkin:2016:BMV

[BBK^{+}16] M. A. Borkin, Z. Bylinskii, N. W. Kim, C. M. Bainbridge, C. S. Yeh, D. Borkin, H. Pfister, and A. Oliva. Beyond memorability: Visualization recognition and recall. IEEE Transactions on Visualization and Computer Graphics, 22(1):519–528, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blaas:2008:EPC

[BBP08] Jorik Blaas, Charl Botha, and Frits Post. Extensions of parallel coordinates for interactive exploration of large multitimepoint data sets. IEEE Transactions on Visualization and Computer Graphics, 14(6):1436–1451, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Botchen:2008:ABM

[BBS^{+}08] Ralf P. Botchen, Sven Bachthaler, Fabian Schick, Min Chen, Greg Mori, Daniel Weiskopf, and Thomas Ertl. Actionbased multifield video visualization. IEEE Transactions on Visualization and Computer Graphics, 14(4):885– 899, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040885s.
zip.

Brown:2012:FMB

[BC12] J. Anthony Brown and David W. Capson. A framework for 3D model-based visual tracking using a GPUaccelerated particle filter. IEEE Transactions on Visualization and Computer Graphics, 18(1):68–80, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borsoi:2018:PIP

[BC18a] Ricardo Augusto Borsoi and Guilherme Holsbach Costa. On the performance and implementation of parallax free video see-through displays. IEEE Transactions on Visualization and Computer Graphics, 24(6):2011– 2022, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07930426-
abs.html.

Buttussi:2018:EDT

[BC18b] F. Buttussi and L. Chittaro. Effects of different types of virtual reality display on presence and learning in a safety training scenario. IEEE Transactions on Visualization and Computer Graphics, 24(2):1063–1076, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bier:2010:PTC

[BCB10] Eric A. Bier, Stuart K. Card, and John W. Bodnar. Principles and tools for collaborative entity-based intelligence analysis. IEEE Transactions on Visualization and Computer Graphics, 16(2):178– 191, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Basole:2013:UIR

[BCH^{+}13] Rahul C. Basole, Trustin Clear, Mengdie Hu, Harshit Mehrotra, and John Stasko. Understanding interfirm relationships in business ecosystems with interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 19(12):2526– 2535, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bertel:2019:MCP

[BCR19] Tobias Bertel, Neill D. F. Campbell, and Christian Richardt. MegaParallax: Casual 360 panoramas with motion parallax. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1828–1835, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8651483/.

Bartram:2011:ECT

[BCS11] Lyn Bartram, Billy Cheung, and Maureen Stone. The effect of colour and transparency on the perception of overlaid grids. IEEE Transactions on Visualization and Computer Graphics, 17(12): 1942–1948, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baumeister:2016:ARC

[BDB^{+}16] J. Baumeister, J. Dorrlan, S. Banks, A. Chatburn, R. T. Smith, M. A. Carskadon, K. Lushington, and B. H. Thomas. Augmented reality as a countermeasure for sleep deprivation. IEEE Transactions on Visualization and Computer Graphics, 22(4): 1396–1405, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bertails-Descoubes:2017:GEI

[BDC17] Florence Bertails-Descoubes and Stelian Coros. Guest Editor’s introduction: Special section on the ACM SIG-GRAPH/Eurographics Symposium on Computer Animation (SCA). IEEE Transactions on Visualization and Computer Graphics, 23(3): 1165–1166, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07835799.
pdf.

Bouts:2016:VED

[BDD^{+}16] Quirijn W. Bouts, Tim Dwyer, Jason Dykes, Bettina Speckmann, Sarah Goodwin, Nathalie Henry Riche, Sheelagh Carpendale, and Ariel Liebman. Visual encoding of dissimilarity data via topology-preserving map deformation. IEEE Transactions on Visualization and Computer Graphics, 22(9): 2200–2213, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bezerianos:2010:GSE

[BDF^{+}10] Anastasia Bezerianos, Pierre Dragicevic, Jean-Daniel Fekete, Juhee Bae, and Ben Watson. GeneaQuilts: a system for exploring large genealogies. IEEE Transactions on Visualization and Computer Graphics, 16(6):1073–1081, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bapat:2016:TKH

[BDF16] Akash Bapat, Enrique Dunn, and Jan-Michael Frahm. Towards kilo-Hertz 6-DoF visual tracking using an egocentric cluster of rolling shutter cameras. IEEE Transactions on Visualization and Computer Graphics, 22(11): 2358–2367, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bigelow:2017:IBT

[BDFM17] Alex Bigelow, Steven Drucker, Danyel Fisher, and Miriah Meyer. Iterating between tools to create and edit visualizations. IEEE Transactions on Visualization and Computer Graphics, 23(1): 481–490, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bovet:2018:CRS

[BDH^{+}18] Sidney Bovet, Henrique Galvan Debarba, Bruno Herbelin, Eray Molla, and Ronan Boulic. The critical role of self-contact for embodiment in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1428–1436, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08283639.
pdf.

Bertram:2004:GBS

[BDHJ04] Martin Bertram, Mark A.

Duchaineau, Bernd Hamann, and Kenneth I. Joy. Generalized B-spline subdivisionsurface wavelets for geometry compression. IEEE Transactions on Visualization and Computer Graphics, 10 (3):326–338, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0326abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0326.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0326.pdf.

Bargo:2014:OMM

[BDJ14] Rita Bargo, Joel Dearden, and Mark W. Jones. Order of magnitude markers: An empirical study on large magnitude number detection. IEEE Transactions on Visualization and Computer Graphics, 20(12):2261– 2270, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

                                        csdl/trans/tg/2014/12/06875940-
                                        abs.html.

Barequet:1998:RGT

[BDK98] G. Barequet, C. A. Duncan, and S. Kumar. RSVP: a geometric toolkit for controlled repair of solid models. IEEE Transactions on Visualization and Computer Graphics, 4 (2):162–177, April/June 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0162.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0162abs.
                                        htm.

Beecham:2017:MLE

[BDM^{+}17] Roger Beecham, Jason Dykes, Wouter Meulemans, Aidan Slingsby, Cagatay Turkay, and Jo Wood. Map LineUps: Effects of spatial structure on graphical inference. IEEE Transactions on Visualization and Computer Graphics, 23 (1):391–400, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bonnell:2003:MIR

[BDS^{+}03] Kathleen S. Bonnell, Mark A. Duchaineau, Daniel R. Schikore, Bernd Hamann, and Kenneth I. Joy. Material interface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 9(4):500– 511, October/December 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0500abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0500.pdf.

Bock:2018:TIT

[BDSS18] A. Bock, H. Doraiswamy, A. Summers, and C. Silva. TopoAngler: Interactive topology-based extraction of fishes. IEEE Transactions on Visualization and Computer Graphics, 24(1):812–821, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Biswas:2013:IAF

[BDSW13] Ayan Biswas, Soumya Dutta, Han-Wei Shen, and Jonathan Woodring. An informationaware framework for exploring multivariate data sets. IEEE Transactions on Visualization and Computer Graphics, 19(12):2683–2692, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Butkiewicz:2008:MFG

[BDW^{+}08] Thomas Butkiewicz, Wenwen Dou, Zachary Wartell, William Ribarsky, and Remco Chang. Multi-focused geospatial analysis using probes. IEEE Transactions on Visualization and Computer Graphics, 14(6):1165–1172, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Been:2006:DML

[BDY06] Ken Been, Eli Daiches, and Chee Yap. Dynamic map labeling. IEEE Transactions on Visualization and Computer Graphics, 12(5): 773–780, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bimber:2006:MPM

[BE06] Oliver Bimber and Andreas Emmerling. Multifocal projection: a multiprojector technique for increasing focal depth. IEEE Transactions on Visualization and Computer Graphics, 12(4):658– 667, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/;   http://csdl.
                                        computer.org/comp/trans/
                                        tg/2006/04/v0658s.zip.

Bresciani:2009:BSC

[BE09] Sabrina Bresciani and Martin J. Eppler. The benefits of synchronous collaborative information visualization: Evidence from an experimental evaluation. IEEE Transactions on Visualization and Computer Graphics, 15(6):1073–1080, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Batch:2018:IVG

[BE18] A. Batch and N. Elmqvist.

The interactive visualization gap in initial exploratory data analysis. IEEE Transactions on Visualization and Computer Graphics, 24(1):278– 287, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boy:2016:SIS

[BEDF16] J. Boy, L. Eveillard, F. Detienne, and J. Fekete. Suggested interactivity: Seeking perceived affordances for information visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (1):639–648, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bremer:2004:THF

[BEHP04] Peer-Timo Bremer, Herbert Edelsbrunner, Bernd Hamann, and Valerio Pascucci. A topological hierarchy for functions on triangulated surfaces. IEEE Transactions on Visualization and Computer Graphics, 10(4): 385–396, July/August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

                                        dl/trans/tg/2004/04/v0385.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/04/
                                        v0385.pdf.

Bowman:2012:TVG

[BEJK12] Brian Bowman, Niklas Elmqvist, and T. J. Jankun-Kelly. Toward visualization for games: Theory, design space, and patterns. IEEE Transactions on Visualization and Computer Graphics, 18(11):1956– 1968, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bendich:2010:CRP

[BEK10] Paul Bendich, Herbert Edelsbrunner, and Michael Kerber. Computing robustness and persistence for images. IEEE Transactions on Visualization and Computer Graphics, 16(6):1251–1260, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baboud:2012:PSS

[BES12] Lionel Baboud, Elmar Eisemann, and Hans-Peter Seidel. Precomputed safety shapes for efficient and accurate height-field rendering. IEEE Transactions on Visualization and Computer Graphics, 18(11):1811–1823, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Becker:1995:VND

[BEW95] Richard A. Becker, Stephen G.

Eick, and Allan R. Wilks. Visualizing network data. IEEE Transactions on Visualization and Computer Graphics, 1 (1):16–28, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0016.
pdf;   http://www.computer.
org/tvcg/tg1995/v0016abs.
htm.

Barhak:2001:PRS

[BF01] J. Barhak and A. Fischer.

Parameterization and reconstruction from 3D scattered points based on neural network and PDE techniques. IEEE Transactions on Visualization and Computer Graphics, 7(1):1–??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0001.
pdf;   http://www.computer.
org/tvcg/tg2001/v0001abs.
htm.

Badam:2015:MPP

[BFE15] Sriram Karthik Badam, Eli Fisher, and Niklas Elmqvist. Munin: A peer-to-peer middleware for ubiquitous analytics and visualization spaces. IEEE Transactions on Visualization and Computer Graphics, 21(2):215– 228, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/02/06851203-
                                        abs.html.

Brandes:2006:SDB

[BFL06] Ulrik Brandes, Daniel Fleischer, and Jürgen Lerner. Summarizing dynamic bipolar conflict structures. IEEE Transactions on Visualization and Computer Graphics, 12(6):1486–1499, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bach:2014:GAT

[BFP14] Benjamin Bach, Jean-Daniel Fekete, and Emmanuel Pietriga. GraphDiaries: Animated transitions and temporal navigation for dynamic networks. IEEE Transactions on Visualization and Computer Graphics, 20(5):740–754, May 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Buerger:2009:ISS

[BFTW09] Kai Buerger, Florian Ferstl, Holger Theisel, and Rudiger Westermann. Interactive streak surface visualization on the GPU. IEEE Transactions on Visualization and Computer Graphics, 15(6):1259– 1266, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bartholdi:2004:MIT

[BG04] John J. Bartholdi and Paul Goldsman. Multiresolution indexing of triangulated irregular networks. IEEE Transactions on Visualization and Computer Graphics, 10(4): 484–495, July/August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/04/v0484.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0484.pdf.

Bruckner:2006:EVV

[BG06] Stefan Bruckner and M. Eduard Gröller. Exploded views for volume data. IEEE Transactions on Visualization and Computer Graphics, 12(5):1077–1084, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruckner:2007:EDP

[BG07] Stefan Bruckner and Eduard Gröller. Enhancing depth-perception with flexible volumetric halos. IEEE Transactions on Visualization and Computer Graphics, 13(6):1344–1351, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1344s.
                                        zip.

Bhattacharya:2015:LSM

[BGB15] Haimasree Bhattacharya, Yue Gao, and Adam W. Bargteil. A level-set method for skinning animated particle data. IEEE Transactions on Visualization and Computer Graphics, 21(3):315–327, March 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/csdl/trans/tg/2015/
                                        03/06919322-abs.html.

Babu:2011:IVP

[BGC^{+}11] Sabarish V. Babu, Timofey Y. Grechkin, Benjamin Chihak, Christine Ziemer, Joseph K. Kearney, James F. Cremer, and Jodie M. Plumert. An immersive virtual peer for studying social influences on child cyclists’ road-crossing behavior. IEEE Transactions on Visualization and Computer Graphics, 17(1):14–25, January 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baldacci:2017:PST

[BGCS17] Andrea Baldacci, Fabio Ganovelli, Massimiliano Corsini, and Roberto Scopigno. Presentation of 3D scenes through video example. IEEE Transactions on Visualization and Computer Graphics, 23(9): 2096–2107, September 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07565535-
abs.html.

Bronsvoort:2011:GEI

[BGK11] Willem F. Bronsvoort, Jens Gravesen, and John Keyser. Guest Editor’s introduction: Special section on the Joint Conference on Geometric Design and Solid and Physical Modeling (GDSPM). IEEE Transactions on Visualization and Computer Graphics, 17 (6):713–714, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruckner:2006:ICP

[BGKG06] Stefan Bruckner, Soren Grimm, Armin Kanitsar, and M. Eduard Gröller. Illustrative context-preserving exploration of volume data. IEEE Transactions on Visualization and Computer Graphics, 12(6):1559–1569, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Biddiscombe:2007:TDP

[BGM^{+}07] John Biddiscombe, Berk Geveci, Ken Martin, Kenneth Moreland, and David Thompson. Time dependent processing in a parallel pipeline architecture. IEEE Transactions on Visualization and Computer Graphics, 13(6):1376–1383, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1376s.

avi. See corrections [BGM^{+}08].

Biddiscombe:2008:CTD

[BGM^{+}08] John Biddiscombe, Berk Geveci, Ken Martin, Kenneth Moreland, and David Thompson. Corrections to “Time Dependent Processing in a Parallel Pipeline Architecture”. IEEE Transactions on Visualization and Computer Graphics, 14(1): 241, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/01/ttg2008010241.

pdf. See [BGM^{+}07].

Bryan:2017:SEI

[BGM^{+}17] Chris Bryan, Gregory Guterman, Kwan-Liu Ma, Harris Lewin, Denis Larkin, Jaebum Kim, Jian Ma, and Marta Farré. Synteny Explorer: An interactive visualization application for teaching genome evolution. IEEE Transactions on Visualization and Computer Graphics, 23 (1):711–720, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bensema:2016:MDC

[BGOJ16] Kevin Bensema, Luke Gosink, Harald Obermaier, and Kenneth I. Joy. Modality-driven classification and visualization of ensemble variance. IEEE Transactions on Visualization and Computer Graphics, 22(10):2289–2299, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borkin:2011:EAV

[BGP^{+}11] Michelle Borkin, Krzysztof Gajos, Amanda Peters, Dimitrios Mitsouras, Simone Melchionna, Frank Rybicki, Charles Feldman, and Hanspeter Pfister. Evaluation of artery visualizations for heart disease diagnosis. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2479–2488, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Buering:2006:UIS

[BGR06] Thorsten Buering, Jens Gerken, and Harald Reiterer. User interaction with scatterplots on small screens — a comparative evaluation of geometricsemantic zoom and fisheye distortion. IEEE Transactions on Visualization and Computer Graphics, 12(5): 829–836, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Barakat:2012:ICR

[BGT12] Samer Barakat, Christoph Garth, and Xavier Tricoche. Interactive computation and rendering of finite-time Lyapunov exponent fields. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1368–1380, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bair:2007:GVO

[BH07] Alethea Bair and Donald House. Grid with a view: Optimal texturing for perception of layered surface shape. IEEE Transactions on Visualization and Computer Graphics, 13(6):1656–1663, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bostock:2009:PGT

[BH09] Michael Bostock and Jeffrey Heer. Protovis: a graphical toolkit for visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1121–1128, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Barrera:2004:FSE

[BHB04] Tony Barrera, Anders Hast, and Ewert Bengtsson. Faster shading by equal angle interpolation of vectors. IEEE Transactions on Visualization and Computer Graphics, 10(2):217–223, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2004/
02/v0217abs.htm;   http:
//csdl.computer.org/dl/
trans/tg/2004/02/v0217.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0217.pdf.

Beham:2014:IX

[BHGK14] Michael Beham, Wolfgang Herzner, M. Eduard Groller, and Johannes Kehrer. IEEE XPLORE. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1693–1702, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875958-
abs.html.

Bachmann:2019:MUR

[BHHM19] Eric R. Bachmann, Eric Hodgson, Cole Hoffbauer, and Justin Messinger. Multi-user redirected walking and resetting using artificial potential fields. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2022–2031, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8645818/.

Block:2012:DEV

[BHP^{+}12] Florian Block, Michael S. Horn, Brenda Caldwell Phillips, Judy Diamond, E. Margaret Evans, and Chia Shen. The DeepTree Exhibit: Visualizing the tree of life to facilitate informal learning. IEEE Transactions on Visualization and Computer Graphics, 18(12):2789–2798, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Balsys:2012:VNS

[BHS12] Ron J. Balsys, Dirk J. Harbinson, and Kevin G. Suffern. Visualizing nonmanifold and singular implicit surfaces with point clouds. IEEE Transactions on Visualization and Computer Graphics, 18(2): 188–201, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bujack:2015:MIF

[BHSH15] Roxana Bujack, Ingrid Hotz, Gerik Scheuermann, and Eckhard Hitzer. Moment invariants for 2D flow fields via normalization in detail. IEEE Transactions on Visualization and Computer Graphics, 21(8):916–929, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/06951493-abs.html.

Baricevic:2017:UPA

[BHST17] Domagoj Baricevic, Tobias Hollerer, Pradeep Sen, and Matthew Turk. Userperspective AR magic lens from gradient-based IBR and semi-dense stereo. IEEE Transactions on Visualization and Computer Graphics, 23(7):1838–1851, July 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07460953-
abs.html.

Bao:2007:FRM

[BHTF07] Zhaosheng Bao, Jeong-Mo Hong, Joseph Teran, and Ronald Fedkiw. Fracturing rigid materials. IEEE Transactions on Visualization and Computer Graphics, 13(2): 370–378, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brandes:2015:GEI

[BHTY15] Ulrik Brandes, Hans Hagen, Shigeo Takahashi, and Xiaoru Yuan. Guest Editors’ introduction: Special section on the IEEE Pacific Visualization Symposium 2014. IEEE Transactions on Visualization and Computer Graphics, 21(8):887–888, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/07138667.pdf.

Bair:2006:TLS

[BHW06] Alethea Bair, Donald H. House, and Colin Ware. Texturing of layered surfaces for optimal viewing. IEEE Transactions on Visualization and Computer Graphics, 12(5):1125–1132, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beyer:2007:HQM

[BHWB07] Johanna Beyer, Markus Hadwiger, Stefan Wolfsberger, and Katja Bühler. High-quality multimodal volume rendering for preoperative planning of neurosurgical interventions. IEEE Transactions on Visualization and Computer Graphics, 13(6):1696–1703, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1696s.
zip.

Bernard:2018:CVI

[BHZ^{+}18] J. Bernard, M. Hutter, M. Zeppelzauer, D. Fellner, and M. Sedlmair. Comparing visual-interactive labeling with active learning: An experimental study. IEEE Transactions on Visualization and Computer Graphics, 24 (1):298–308, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bezerianos:2012:PVV

[BI12] Anastasia Bezerianos and Petra Isenberg. Perception of visual variables on tiled wall-sized displays for information visualization applications. IEEE Transactions on Visualization and Computer Graphics, 18(12):2516– 2525, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Besancon:2017:HTT

[BIAI17] Lonni Besançon, Paul Issartel, Mehdi Ammi, and Tobias Isenberg. Hybrid tactile/tangible interaction for 3D data exploration. IEEE Transactions on Visualization and Computer Graphics, 23 (1):881–890, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Billinghurst:2013:VRT

[Bil13]  Mark Billinghurst.  The

2013 Virtual Reality Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 19(4):xviii, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruder:2012:RWD

[BIPS12] Gerd Bruder, Victoria Interrante, Lane Phillips, and Frank Steinicke. Redirecting walking and driving for natural navigation in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18 (4):538–545, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruckner:2019:MSD

[BIRW19] S. Bruckner, T. Isenberg, T. Ropinski, and A. Wiebel. A model of spatial directness in interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 25(8):2514–2528, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brehmer:2014:ODA

[BISM14] Matthew Brehmer, Stephen Ingram, Jonathan Stray, and Tamara Munzner. Overview: The design, adoption, and analysis of a visual document mining tool for investigative journalists. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2271–2280, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875900-
abs.html.

Boorboor:2019:VNS

[BJA^{+}19] Saeed Boorboor, Shreeraj Jadhav, Mala Ananth, David Talmage, Lorna Role, and Arie Kaufman. Visualization of neuronal structures in wide-field microscopy brain images. IEEE Transactions on Visualization and Computer Graphics, 25(1):1018– 1028, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440805-
abs.html.

Bhatia:2012:FVQ

[BJB^{+}12] Harsh Bhatia, Shreeraj Jadhav, Peer-Timo Bremer, Guoning Chen, Joshua A. Levine, Luis Gustavo Nonato, and Valerio Pascucci. Flow visualization with quantified spatial and temporal errors using edge maps. IEEE Transactions on Visualization and Computer Graphics, 18(9):1383–1396, September 2012. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Buchmuller:2019:MVC

[BJC^{+}19] Juri Buchmuller, Dominik Jackle, Eren Cakmak, Ulrik Brandes, and Daniel A. Keim. MotionRugs: Visualizing collective trends in space and time. IEEE Transactions on Visualization and Computer Graphics, 25(1): 76–86, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440823-
                                        abs.html.

Bar-Joseph:2001:TMT

[BJEYLW01] Z. Bar-Joseph, R. El-Yaniv, D. Lischinski, and M. Werman. Texture mixing and texture movie synthesis using statistical learning. IEEE Transactions on Visualization and Computer Graphics, 7 (2):120–??, April/June 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2001/pdf/v0120.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0120abs.
                                        htm.

Blascheck:2016:VVA

[BJK^{+}16] T. Blascheck, M. John, K. Kurzhals, S. Koch, and T. Ertl. VA^{2}: A visual analytics approach for || evaluating visual analytics applications. IEEE Transactions on Visualization and Computer Graphics, 22(1):61–70, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhasker:2007:RTU

[BJM07] Ezekiel Bhasker, Ray Juang, and Aditi Majumder. Registration techniques for using imperfect and par tially calibrated devices in planar multi-projector displays. IEEE Transactions on Visualization and Computer Graphics, 13(6):1368–1375, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brady:1998:IVN

[BJNN98] M. L. Brady, K. K. Jung, H. T. Nguyen, and Thinh P. Q. Nguyen. Interactive volume navigation. IEEE Transactions on Visualization and Computer Graphics, 4(3):243– 256, July/September 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0243.
pdf;   http://www.computer.
org/tvcg/tg1998/v0243abs.
htm.

Bilal:2018:DCN

[BJY^{+}18] A. Bilal, A. Jourabloo, M. Ye, X. Liu, and L. Ren. Do convolutional neural networks learn class hierarchy? IEEE Transactions on Visualization and Computer Graphics, 24 (1):152–162, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Biggers:2012:IBS

[BK12] Keith Biggers and John Keyser. Inference-based surface reconstruction of cluttered environments. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1255–1267, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bender:2017:DFS

[BK17] Jan Bender and Dan Koschier. Divergence-free SPH for incompressible and viscous fluids. IEEE Transactions on Visualization and Computer Graphics, 23(3):1193– 1206, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07487018-
abs.html.

Bimber:2011:CLF

[BKA^{+}11] Oliver Bimber, Daniel Klock, Toshiyuki Amano, Anselm Grundhofer, and Daniel Kurz. Closed-loop feedback illumination for optical inverse tone-mapping in light microscopy. IEEE Transactions on Visualization and Computer Graphics, 17(6): 857–870, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bender:2000:FFW

[BKDE00] M. Bender, R. Klein, A. Disch, and A. Ebert. A functional framework for Webbased information visualization systems. IEEE Transactions on Visualization and Computer Graphics, 6(1):8– 23, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0008.
pdf;   http://www.computer.
org/tvcg/tg2000/v0008abs.
htm.

Burch:2011:ETO

[BKH^{+}11] Michael Burch, Natalia Konevtsova, Julian Heinrich, Markus Hoeferlin, and Daniel Weiskopf. Evaluation of traditional, orthogonal, and radial tree diagrams by an eye tracking study. IEEE Transactions on Visualization and Computer Graphics, 17(12):2440– 2448, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beck:2013:IGG

[BKKF13] Stephan Beck, Andre Kunert, Alexander Kulik, and Bernd Froehlich. Immersive groupto-group telepresence. IEEE Transactions on Visualization and Computer Graphics, 19 (4):616–625, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bender:2019:TMS

[BKKW19] Jan Bender, Dan Koschier, Tassilo Kugelstadt, and Marcel Weiler. Turbulent micropolar SPH fluids with foam. IEEE Transactions on Visualization and Computer Graphics, 25(6):2284– 2295, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8353115/.

Bennett:2011:FBS

[BKL^{+}11] Janine C. Bennett, Vaidyanathan Krishnamoorthy, Shusen Liu, Ray W. Grout, Evatt R. Hawkes, Jacqueline H. Chen, Jason Shepherd, Valerio Pascucci, and Peer-Timo Bremer. Feature-based statistical analysis of combustion simulation data. IEEE Transactions on Visualization and Computer Graphics, 17(12):1822– 1831, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruckner:2018:GEI

[BKL18] Stefan Bruckner, Koji Koyamada, and Bongshin Lee. Guest Editors’ introduction:

Special section on IEEE PacificVis 2018. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1879–1880, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/08352605.
pdf.

Bakke:2013:ALS

[BKM13] Eirik Bakke, David R. Karger, and Robert C. Miller. Automatic layout of structured hierarchical reports. IEEE Transactions on Visualization and Computer Graphics, 19(12):2586–2595, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Becher:2019:FBV

[BKRE19] Michael Becher, Michael Krone, Guido Reina, and Thomas Ertl. Feature-based volumetric terrain generation and decoration. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1283–1296, February 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08066313-
abs.html.

Bitter:2001:PDV

[BKS01] Ingmar Bitter, Arie E.

Kaufman, and Mie Sato.

Penalized-distance volumetric skeleton algorithm. IEEE Transactions on Visualization and Computer Graphics, 7 (3):195–206, July 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0195.
pdf;   http://www.computer.
org/tvcg/tg2001/v0195abs.
htm.

Brandes:2003:CCP

[BKW03] Ulrik Brandes, Patrick Kenis, and Dorothea Wagner. Communicating centrality in policy network drawings. IEEE Transactions on Visualization and Computer Graphics, 9 (2):241–253, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0241abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0241.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0241.pdf.

Burger:2008:DVE

[BKW08] Kai Bürger, Jens Krüger, and Rüdiger Westermann. Direct volume editing. IEEE Transactions on Visualization and Computer Graphics, 14(6):1388–1395, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Burger:2010:SBS

[BKW10] Kai Burger, Jens Kruger, and Rudiger Westermann. Sample-based surface coloring. IEEE Transactions on Visualization and Computer Graphics, 16(5): 763–776, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beck:2016:VAD

[BKW16] F. Beck, S. Koch, and D. Weiskopf. Visual analysis and dissemination of scientific literature collections with SurVis. IEEE Transactions on Visualization and Computer Graphics, 22(1): 180–189, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blanch:2007:BZT

[BL07] Renaud Blanch and Éric Lecolinet. Browsing zoomable treemaps: Structure-aware multi-scale navigation techniques. IEEE Transactions on Visualization and Computer Graphics, 13(6):1248– 1253, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bolte:2015:SRR

[BL15] B. Bolte and M. Lappe.

Subliminal reorientation and repositioning in immersive virtual environments using saccadic suppression. IEEE Transactions on Visualization and Computer Graphics, 21 (4):545–552, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blascovich:2012:CS

[Bla12] James J. Blascovich. Capstone speaker. IEEE Transactions on Visualization and Computer Graphics, 18(4): xii, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brehmer:2017:TRD

[BLB^{+}17] Matthew Brehmer, Bongshin Lee, Benjamin Bach, Nathalie Henry Riche, and Tamara Munzner. Timelines revisited: A design space and considerations for expressive storytelling. IEEE Transactions on Visualization and Computer Graphics, 23(9): 2151–2164, September 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07581076-
abs.html.

Badam:2019:EDC

[BLE19] Sriram Karthik Badam, Zhicheng Liu, and Niklas Elmqvist. Elastic documents: Coupling text and tables through contextual visualizations for enhanced document reading. IEEE Transactions on Visualization and Computer Graphics, 25(1):661– 671, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440810-
abs.html.

Byska:2016:AEP

[BLG^{+}16] J. Byska, M. Le Muzic, M. E.

Groller, I. Viola, and B. Kozlikova. AnimoAminoMiner: Exploration of protein tunnels and their properties in molecular dynamics. IEEE Transactions on Visualization and Computer Graphics, 22 (1):747–756, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brehmer:2019:VRT

[BLIC19] Matthew Brehmer, Bongshin Lee, Petra Isenberg, and Eun Kyoung Choe. Visualizing ranges over time on mobile phones: a task-based crowdsourced evaluation. IEEE Transactions on Visualization and Computer Graphics, 25 (1):619–629, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440812-
abs.html.

Berger:2019:GMV

[BLL19] Matthew Berger, Jixian Li, and Joshua A. Levine. A generative model for volume rendering. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1636–1650, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8316963/.

Biswas:2017:VTV

[BLLS17] Ayan Biswas, Guang Lin, Xiaotong Liu, and Han-Wei Shen. Visualization of timevarying weather ensembles across multiple resolutions. IEEE Transactions on Visualization and Computer Graphics, 23(1):841–850, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bentum:1996:FAG

[BLM96] Mark J. Bentum, Barthold B. A. Lichtenbelt, and Tom Malzbender. Frequency analysis of gradient estimators in volume rendering. IEEE Transactions on Visualization and Computer Graphics, 2 (3):242–254, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0242.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0242abs.
                                        htm.

Broll:2005:IRC

[BLO^{+}05] Wolfgang Broll, Irma Lindt, Jan Ohlenburg, Iris Herbst, Michael Wittkamper, and Thomas Novotny. An infrastructure for realizing customtailored augmented reality user interfaces. IEEE Transactions on Visualization and Computer Graphics, 11(6):722–733, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Burdea:2005:GES

[BLRW05] Grigore C. Burdea, Ming C.

Lin, William Ribarsky, and Benjamin Watson. Guest Editorial: Special issue on haptics, virtual, and augmented reality. IEEE Transactions on Visualization and Computer Graphics, 11(6):611–613, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/06/v0611.
pdf.

Banks:2004:CCS

[BLS04] David C. Banks, Stephen A.

Linton, and Paul K. Stockmeyer. Counting cases in substitope algorithms. IEEE Transactions on Visualization and Computer Graphics, 10(4):371–384, July/ August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        dl/trans/tg/2004/04/v0371.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/04/
                                        v0371.pdf.

Bodin:2012:CF

[BLS12] Kenneth Bodin, Claude Lacoursiere, and Martin Servin. Constraint fluids. IEEE Transactions on Visualization and Computer Graphics, 18 (3):516–526, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruder:2015:CRD

[BLS15] G. Bruder, P. Lubas, and F. Steinicke. Cognitive resource demands of redirected walking. IEEE Transactions on Visualization and Computer Graphics, 21(4):539– 544, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bronson:2014:LCM

[BLW14] Jonathan Bronson, Joshua A. Levine, and Ross Whitaker. Lattice cleaving: A multimaterial tetrahedral meshing algorithm with guarantees. IEEE Transactions on Visualization and Computer Graphics, 20(2):223–237, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruckner:2010:RDE

[BM10] Stefan Bruckner and Torsten Moller. Result-driven exploration of simulation parameter spaces for visual effects design. IEEE Transactions on Visualization and Computer Graphics, 16(6):1468– 1476, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brehmer:2013:MLT

[BM13] Matthew Brehmer and Tamara Munzner. A multi-level typology of abstract visualization tasks. IEEE Transactions on Visualization and Computer Graphics, 19(12):2376– 2385, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beyer:2019:CES

[BMA^{+}19] Johanna Beyer, Haneen Mohammed, Marco Agus, Ali K. Al-Awami, Hanspeter Pfister, and Markus Hadwiger. Culling for extreme-scale segmentation volumes: A hybrid deterministic and probabilistic approach. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1132–1141, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08444102-
abs.html.

Barsky:2008:CVM

[BMGK08] Aaron Barsky, Tamara Munzner, Jennifer Gardy, and Robert Kincaid. Cerebral: Visualizing multiple experimental conditions on a graph with biological context. IEEE Transactions on Visualization and Computer Graphics, 14(6):1253–1260, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bak:2009:SAS

[BMJK09] Peter Bak, Florian Mansmann, Halldor Janetzko, and Daniel Keim. Spatiotemporal analysis of sensor logs using growth ring maps. IEEE Transactions on Visualization and Computer Graphics, 15(6):913–920, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Berenjkoub:2019:VAS

[BMLC19] Marzieh Berenjkoub, Rodolfo Ostilla Monico, Robert S. Laramee, and Guoning Chen. Visual analysis of spatiatemporal relations of pairwise attributes in unsteady flow. IEEE Transactions on Visualization and Computer Graphics, 25(1):1246– 1256, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440118-
abs.html.

Brodersen:2008:GTU

[BMPB08] Anders Brodersen, Ken Museth, Serban Porumbescu, and Brian Budge. Geometric texturing using level sets. IEEE Transactions on Visualization and Computer Graphics, 14(2):277–288, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bernardini:1999:BPA

[BMR^{+}99] F. Bernardini, J. Mittleman, H. Rushmeier, C. Silva, and G. Taubin. The ball-pivoting algorithm for surface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 5(4):349–359, October/December 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0349.
pdf;   http://www.computer.
org/tvcg/tg1999/v0349abs.
htm.

Bernardini:2001:HQT

[BMR01] Fausto Bernardini, Ioana M.

Martin, and Holly Rushmeier. High-quality texture reconstruction from multiple scans. IEEE Transactions on Visualization and Computer Graphics, 7(4):318– 332, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2001/pdf/v0318.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0318abs.
                                        htm.

Badam:2019:VCM

[BMR^{+}19] Sriram Karthik Badam, Andreas Mathisen, Roman Radle, Clemens N. Klokmose, and Niklas Elmqvist. Vistrates: A component model for ubiquitous analytics. IEEE Transactions on Visualization and Computer Graphics, 25(1):586– 596, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440803-
                                        abs.html.

Berger:2017:CCD

[BMS17] Matthew Berger, Katherine McDonough, and Lee M. Seversky. cite2vec: Citationdriven document exploration via word embeddings. IEEE Transactions on Visualization and Computer Graphics, 23 (1):691–700, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boulic:1997:CCP

[BMST97] R. Boulic, R. Mas-Sanso, and D. Thalmann. Complex character positioning based on a compatible flow model of multiple supports. IEEE Transactions on Visualization and Computer Graphics, 3(3):245– 261, July/September 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0245.
pdf;   http://www.computer.
org/tvcg/tg1997/v0245abs.
htm.

Bergner:2005:PAS

[BMTD05] Steven Bergner, Torsten Möller, Melanie Tory, and Mark S. Drew. A practical approach to spectral volume rendering. IEEE Transactions on Visualization and Computer Graphics, 11(2):207– 216, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bryan:2017:TSI

[BMW17] Chris Bryan, Kwan-Liu Ma, and Jonathan Woodring. Temporal summary images: An approach to narrative visualization via interactive annotation generation and placement. IEEE Transactions on Visualization and Computer Graphics, 23(1): 511–520, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bergner:2006:SAF

[BMWM06] Steven Bergner, Torsten Möller, Daniel Weiskopf, and David J. Muraki. A spectral analysis of function composition and its implications for sampling in direct volume visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1353–1360, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Birsak:2018:DPE

[BMWW18] Michael Birsak, Przemyslaw Musialski, Peter Wonka, and Michael Wimmer. Dynamic path exploration on mobile devices. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1784–1798, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07891018-
abs.html.

Brown:2005:CBC

[BMY05] Michael Brown, Aditi Majumder, and Ruigang Yang. Camera-based calibration techniques for seamless multiprojector displays. IEEE Transactions on Visualization and Computer Graphics, 11(2):193–206, March/April 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brandes:2011:ARL

[BN11] Ulrik Brandes and Bobo Nick.

Asymmetric relations in longitudinal social networks. IEEE Transactions on Visualization and Computer Graphics, 17(12):2283–2290, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruneton:2012:SNP

[BN12] Eric Bruneton and Fabrice Neyret. A survey of nonlinear prefiltering methods for efficient and accurate surface shading. IEEE Transactions on Visualization and Computer Graphics, 18(2):242– 260, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhatia:2013:CMH

[BNPB13a] H. Bhatia, G. Norgard, V. Pascucci, and P. Bremer. Comments on the “Meshless Helmholtz–Hodge Decomposition”. IEEE Transactions on Visualization and Computer Graphics, 19(3): 527–528, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhatia:2013:HHD

[BNPB13b] Harsh Bhatia, Gregory Norgard, Valerio Pascucci, and Peer-Timo Bremer. The Helmholtz–Hodge decomposition — a survey. IEEE Transactions on Visualization and Computer Graphics, 19 (8):1386–1404, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brehmer:2016:MMM

[BNTM16] M. Brehmer, J. Ng, K. Tate, and T. Munzner. Matches, mismatches, and methods: Multiple-view workflows for energy portfolio analysis. IEEE Transactions on Visualization and Computer Graphics, 22(1):449–458, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bonneau:1998:MAI

[Bon98] G.-P. Bonneau. Multiresolution analysis on irregular surface meshes. IEEE Transactions on Visualization and Computer Graphics, 4(4):365– 378, October/December 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0365.
pdf;   http://www.computer.
org/tvcg/tg1998/v0365abs.
htm.

Bruneau:2015:GTG

[BOP15] J. Bruneau, A.-H. Olivier, and J. Pettre. Going through, going around: A study on individual avoidance of groups. IEEE Transactions on Visualization and Computer Graphics, 21(4):520–528, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brown:2014:FWL

[BOZ^{+}14] Eli T. Brown, Alvitta Ottley, Helen Zhao, Quan Lin, Richard Souvenir, Alex Endert, and Remco Chang. Finding Waldo: Learning about users from their interactions. IEEE Transactions on Visualization and Computer Graphics, 20(12):1663– 1672, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875913-
abs.html.

Bhatia:2014:NHH

[BPB14] Harsh Bhatia, Valerio Pascucci, and Peer-Timo Bremer. The natural Helmholtz– Hodge decomposition for open-boundary flow analysis. IEEE Transactions on Visualization and Computer Graphics, 20(11):1566– 1578, November 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06774477-
abs.html.

Borgo:2010:EIT

[BPC^{+}10] Rita Borgo, Karl Proctor, Min Chen, Heike Janicke, Tavi Murray, and Ian Thornton. Evaluating the impact of task demands and block resolution on the effectiveness of pixel-based visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):963–972, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bryden:2012:AIM

[BPG12] Aaron Bryden, George N. Phillips, Jr., and Michael Gleicher. Automated illustration of molecular flexibility. IEEE Transactions on Visualization and Computer Graphics, 18(1):132–145, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Binyahib:2019:SHS

[BPL^{+}19] R. Binyahib, T. Peterka, M. Larsen, K. Ma, and H. Childs. A scalable hybrid scheme for ray-casting of unstructured volume data. IEEE Transactions on Visualization and Computer Graphics, 25(7):2349–2361, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Born:2013:VAC

[BPM^{+}13] Silvia Born, Matthias Pfeifle, Michael Markl, Matthias Gutberlet, and Gerik Scheuermann. Visual analysis of cardiac 4D MRI blood flow using line predicates. IEEE Transactions on Visualization and Computer Graphics, 19 (6):900–912, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Barbosa:2016:VIK

[BPP^{+}16] A. Barbosa, F. V. Paulovich, A. Paiva, S. Goldenstein, F. Petronetto, and L. G. Nonato. Visualizing and interacting with kernelized data. IEEE Transactions on Visualization and Computer Graphics, 22(3):1314–1325, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brownlee:2011:PBI

[BPS^{+}11] Carson Brownlee, Vincent Pegoraro, Siddharth Shankar, Patrick S. McCormick, and Charles D. Hansen. Physicallybased interactive flow visualization based on Schlieren and interferometry experimental techniques. IEEE Transactions on Visualization and Computer Graphics, 17(11): 1574–1586, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borland:2013:ESA

[BPS13] David Borland, Tabitha Peck, and Mel Slater. An evaluation of self-avatar eye movement for virtual embodiment. IEEE Transactions on Visualization and Computer Graphics, 19 (4):591–596, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boy:2014:PWA

[BRBF14] Jeremy Boy, Ronald A. Rensink, Enrico Bertini, and Jean-Daniel Fekete. A principled way of assessing visualization literacy. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1963–1972, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875906-
abs.html.

Bach:2017:TUE

[BRH^{+}17] Benjamin Bach, Nathalie Henry Riche, Christophe Hurter, Kim Marriott, and Tim Dwyer. Towards unambiguous edge bundling: Investigating confluent drawings for network visualization. IEEE Transactions on Visualization and Computer Graphics, 23 (1):541–550, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Buck:2019:IAS

[BRNB19] Lauren E. Buck, John J. Rieser, Gayathri Narasimham, and Bobby Bodenheimer. Interpersonal affordances and social dynamics in collaborative immersive virtual environments: Passing together through apertures. IEEE Transactions on Visualization and Computer Graphics, 25(5):2123–2133, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643340/.

Brooks:2006:IBS

[Bro06]  Stephen Brooks.  Image-

based stained glass. IEEE Transactions on Visualization and Computer Graphics, 12(6):1547–1558, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/06/v1547s.
avi.

Brooks:2007:MMP

[Bro07] Stephen Brooks. Mixed media painting and portraiture. IEEE Transactions on Visualization and Computer Graphics, 13(5):1041–1054, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ballester-Ripoll:2019:TDI

[BRP19] Rafael Ballester-Ripoll and Renato Pajarola. Tensor decompositions for integral histogram compression and look-up. IEEE Transactions on Visualization and Computer Graphics, 25(2):1435– 1446, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08281540-
abs.html.

Broll:2018:MISb

[BRS18] Wolfgang Broll, Holger Regenbrecht, and J. Edward Swan. Message from the IS-MAR 2017 Science and Technology Program Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2266–2267, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/08365862.
pdf.

Ballester-Ripoll:2018:MVF

[BRSP18] Rafael Ballester-Ripoll, David Steiner, and Renato Pajarola. Multiresolution volume filtering in the tensor compressed domain. IEEE Transactions on Visualization and Computer Graphics, 24(10):2714– 2727, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08100972-
abs.html.

Barakat:2012:SBS

[BRT12] Samer S. Barakat, Markus Rutten, and Xavier Tricoche. Surface-based structure analysis and visualization for multifield time-varying datasets. IEEE Transactions on Visualization and Computer Graphics, 18(12):2392– 2401, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruneton:2017:QQE

[Bru17] Eric Bruneton. A qualitative and quantitative evaluation of 8 clear sky models. IEEE Transactions on Visualization and Computer Graphics, 23(12):2641– 2655, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07723935-
abs.html.

Banks:1995:PCT

[BS95] David C. Banks and Bart A.

Singer. A predictor-corrector technique for visualizing unsteady flow. IEEE Transactions on Visualization and Computer Graphics, 1(2): 151–163, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0151.
pdf;   http://www.computer.
org/tvcg/tg1995/v0151abs.
htm.

Blundell:2002:CVD

[BS02] Barry G. Blundell and Adam J. Schwarz. The classification of volumetric display systems: Characteristics and predictability of the image space. IEEE Transactions on Visualization and Computer Graphics, 8(1):66– 75, January 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0066.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0066abs.
                                        htm.

Botsch:2008:LVS

[BS08] Mario Botsch and Olga Sorkine. On linear variational surface deformation methods. IEEE Transactions on Visualization and Computer Graphics, 14(1):213–230, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bartram:2011:WDS

[BS11] Lyn Bartram and Maureen C. Stone. Whisper, don’t scream: Grids and transparency. IEEE Transactions on Visualization and Computer Graphics, 17(10):1444–1458, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Butkiewicz:2016:EST

[BS16] T. Butkiewicz and A. H.

Stevens. Effectiveness of structured textures on dynamically changing terrainlike surfaces. IEEE Transactions on Visualization and Computer Graphics, 22(1): 926–934, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bach:2018:HMH

[BSB^{+}18] B. Bach, R. Sicat, J. Beyer, M. Cordeil, and H. Pfister. The hologram in my hand: How effective is interactive exploration of 3D visualizations in immersive tangible augmented reality? IEEE Transactions on Visualization and Computer Graphics, 24 (1):457–467, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baumeister:2017:CCU

[BSE^{+}17] James Baumeister, Seung Youb Ssin, Neven A. M. ElSayed, Jillian Dorrian, David P. Webb, James A. Walsh, Timothy M. Simon, Andrew Irlitti, Ross T. Smith, Mark Kohler, and Bruce H. Thomas. Cognitive cost of using augmented reality displays. IEEE Transactions on Visualization and Computer Graphics, 23(11):2378– 2388, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/11/08007333-
                                        abs.html.

Bork:2018:TEV

[BSEN18] Felix Bork, Christian Schnelzer, Ulrich Eck, and Nassir Navab. Towards efficient visual guidance in limited fieldof-view head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 24(11):2983– 2992, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8456525/.

Bruckner:2009:BVQ

[BŠG^{+}09] Stefan Bruckner, Veronika Šoltészová, Eduard Gröller, Jiři Hladúvka, Katja Buhler, Jai Y. Yu, and Barry J. Dickson. BrainGazer — visual queries for neurobiology research. IEEE Transactions on Visualization and Computer Graphics, 15(6):1497– 1504, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Burlinson:2018:OVC

[BSG18] D. Burlinson, K. Subramanian, and P. Goolkasian. Open vs. closed shapes: New perceptual categories? IEEE Transactions on Visualization and Computer Graphics, 24 (1):574–583, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bach:2016:TCF

[BSH^{+}16] B. Bach, C. Shi, N. Heulot, T. Madhyastha, T. Grabowski, and P. Dragicevic. Time curves: Folding time to visualize patterns of temporal evolution in data. IEEE Transactions on Visualization and Computer Graphics, 22(1): 559–568, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bernard:2019:UDN

[BSKR19] J. Bernard, D. Sessler, J. Kohlhammer, and R. A. Ruddle. Using dashboard networks to visualize multiple patient histories: A design study on post-operative prostate cancer. IEEE Transactions on Visualization and Computer Graphics, 25(3): 1615–1628, March 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bock:2012:CBC

[BSL^{+}12] Alexander Bock, Erik Sunden, Bingchen Liu, Burkhard Wunsche, and Timo Ropinski. Coherency-based curve compression for high-order finite element model visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2315– 2324, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bottger:2014:TDM

[BSL^{+}14] Joachim Bottger, Alexander Schafer, Gabriele Lohmann, Arno Villringer, and Daniel S. Margulies. Three-dimensional mean-shift edge bundling for the visualization of functional connectivity in the brain. IEEE Transactions on Visualization and Computer Graphics, 20(3):471–480, March 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhasker:2006:ADC

[BSM06] Ezekiel S. Bhasker, Pinaki Sinha, and Aditi Majumder. Asynchronous distributed calibration for scalable and reconfigurable multi-projector displays. IEEE Transactions on Visualization and Computer Graphics, 12(5):1101– 1108, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bergner:2013:PIP

[BSM^{+}13] Steven Bergner, Michael Sedlmair, Torsten Moller, Sareh Nabi Abdolyousefi, and Ahmed Saad. ParaGlide: Interactive parameter space partitioning for computer simulations. IEEE Transactions on Visualization and Computer Graphics, 19(9):1499– 1512, September 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Biddiscombe:2012:PCS

[BSO^{+}12] John Biddiscombe, Jerome Soumagne, Guillaume Oger, David Guibert, and Jean-Guillaume Piccinali. Parallel computational steering for HPC applications using HDF5 files in distributed shared memory. IEEE Transactions on Visualization and Computer Graphics, 18(6): 852–864, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Born:2014:SMC

[BSR^{+}14] Silvia Born, Simon H. Sundermann, Christoph Russ, Raoul Hopf, Carlos E. Ruiz, Volkmar Falk, and Michael Gessat. Stent maps — comparative visualization for the prediction of adverse events of transcatheter aortic valve implantations. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2704–2713, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875945-
abs.html.

Bagher:2013:IRA

[BSS^{+}13] Mahdi Mohammad Bagher, Cyril Soler, Kartic Subr, Laurent Belcour, and Nicolas Holzschuch. Interactive rendering of acquired materials on dynamic geometry using frequency analysis. IEEE Transactions on Visualization and Computer Graphics, 19 (5):749–761, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Broll:2018:MISa

[BSS18] Wolfgang Broll, Hideo Saito, and J. Edward Swan II. Message from the ISMAR 2016 Science and Technology Program Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2265, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/08365873.
                                        pdf.

Behrisch:2019:CVA

[BSS^{+}19] M. Behrisch, D. Streeb, F. Stoffel, D. Seebacher, B. Matejek, S. H. Weber, S. Mittelstädt, H. Pfister, and D. Keim. Commercial visual analytics systems advances in the big data analytics field. IEEE Transactions on Visualization and Computer Graphics, 25(10):3011–3031, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baur:2010:SOL

[BSSB10] Dominikus Baur, Frederik Seiffert, Michael Sedlmair, and Sebastian Boring. The streams of our lives: Visualizing listening histories in context. IEEE Transactions on Visualization and Computer Graphics, 16(6):1119– 1128, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bolling:2019:SCA

[BSSL19] Luke Bölling, Niklas Stein, Frank Steinicke, and Markus Lappe. Shrinking circles: Adaptation to increased curvature gain in redirected walking. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2032–2039, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8645699/.

Buchin:2011:FML

[BSV11] Kevin Buchin, Bettina Speckmann, and Kevin Verbeek. Flow map layout via spiral trees. IEEE Transactions on Visualization and Computer Graphics, 17(12):2536– 2544, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bruder:2012:TSM

[BSWL12] Gerd Bruder, Frank Steinicke, Phil Wieland, and Markus Lappe. Tuning self-motion perception in virtual reality with visual illusions. IEEE Transactions on Visualization and Computer Graphics, 18(7):1068–1078, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Barakat:2013:ARF

[BT13] Samer S. Barakat and Xavier Tricoche. Adaptive refinement of the flow map using sparse samples. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2753–2762, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Berkley:2004:RTF

[BTB^{+}04] Jeffrey Berkley, George Turkiyyah, Daniel Berg, Mark Ganter, and Suzanne Weghorst. Realtime finite element modeling for surgery simulation: an application to virtual suturing. IEEE Transactions on Visualization and Computer Graphics, 10(3):314– 325, May/June 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/

comp/trans/tg/2004/03/v0314abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0314.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0314.pdf.

Borst:2010:RTR

[BTB10] Christoph W. Borst, Jan-Phillip Tiesel, and Christopher M. Best. Real-time rendering method and performance evaluation of composable 3D lenses for interactive VR. IEEE Transactions on Visualization and Computer Graphics, 16(3):394– 410, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brivio:2010:BLI

[BTC10] Paolo Brivio, Marco Tarini, and Paolo Cignoni. Browsing large image datasets through Voronoi diagrams. IEEE Transactions on Visualization and Computer Graphics, 16(6):1261–1270, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Baudel:2013:DEL

[BTC13] Thomas Baudel, Arthur G.

Telea, and Paolo Crisafulli. Decision exploration lab: A visual analytics solution for decision management. IEEE Transactions on Visualization and Computer Graphics, 19(12):1972–1981, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bosch:2013:SRT

[BTH^{+}13] Harald Bosch, Dennis Thom, Florian Heimerl, Edwin Puttmann, Steffen Koch, Robert Kruger, Michael Worner, and Thomas Ertl. ScatterBlogs2: Realtime monitoring of microblog messages through user-guided filtering. IEEE Transactions on Visualization and Computer Graphics, 19(12):2022– 2031, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borst:2011:SPC

[BTHD11] Christoph W. Borst, Jan-Phillip Tiesel, Emad Habib, and Kaushik Das. Single-pass composable 3D lens rendering and spatiotemporal 3D lenses. IEEE Transactions on Visualization and Computer Graphics, 17(9):1259–1272, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Butson:2013:EIV

[BTJ^{+}13] C. R. Butson, G. Tamm, S. Jain, T. Fogal, and J. Kruger. Evaluation of interactive visualization on mobile computing platforms for selection of deep brain stimulation parameters. IEEE Transactions on Visualization and Computer Graphics, 19(1): 108–117, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bujack:2018:GBU

[BTS^{+}18] R. Bujack, T. L. Turton, F. Samsel, C. Ware, D. H. Rogers, and J. Ahrens. The good, the bad, and the ugly: A theoretical framework for the assessment of continuous colormaps. IEEE Transactions on Visualization and Computer Graphics, 24(1): 923–933, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Becker:2009:DFL

[BTT09] Markus Becker, Hendrik Tessendorf, and Matthias Teschner. Direct forcing for Lagrangian rigid-fluid coupling. IEEE Transactions on Visualization and Computer Graphics, 15(3):493– 503, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Birkeland:2014:PUM

[BTV14] Asmund Birkeland, Cagatay Turkay, and Ivan Viola. Perceptually uniform motion space. IEEE Transactions on Visualization and Computer Graphics, 20(11):1542– 1554, November 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/11/06811168-
                                        abs.html.

Burch:2011:PES

[BVB^{+}11] Michael Burch, Corinna Vehlow, Fabian Beck, Stephan Diehl, and Daniel Weiskopf. Parallel edge splatting for scalable dynamic graph visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2344– 2353, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borkin:2013:WMV

[BVB^{+}13] Michelle A. Borkin, Azalea A. Vo, Zoya Bylinskii, Phillip Isola, Shashank Sunkavalli, Aude Oliva, and Hanspeter Pfister. What makes a visualization memorable? IEEE Transactions on Visualization and Computer Graphics, 19(12):2306–2315, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bargteil:2012:GEI

[BvdP12] Adam W. Bargteil and Michiel van de Panne. Guest Editor’s introduction: Special Section on the ACM SIG-GRAPH/Eurographics Symposium on Computer Animation (SCA). IEEE Transactions on Visualization and Computer Graphics, 18(8): 1189–1190, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Burakiewicz:2006:ACF

[BvL06] Wojciech Burakiewicz and Robert van Liere. Analyzing complex FTMS simulations: a case study in high-level visualization of ion motions. IEEE Transactions on Visualization and Computer Graphics, 12(5):1037–1044, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Brecheisen:2009:PSV

[BVPtHR09] Ralph Brecheisen, Anna Vilanova, Bram Platel, and Bart ter Haar Romeny. Parameter sensitivity visualization for DTI fiber tracking. IEEE Transactions on Visualization and Computer Graphics, 15(6):1441–1448, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blascheck:2019:ESD

[BVV^{+}19] Tanja Blascheck, Lindsay Mac-Donald Vermeulen, Jo Vermeulen, Charles Perin, Wesley Willett, Thomas Ertl, and Sheelagh Carpendale. Exploration strategies for discovery of interactivity in visualizations. IEEE Transactions on Visualization and Computer Graphics, 25(2):1407– 1420, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08281629-
abs.html.

Bitter:2007:CFF

[BVW^{+}07] Ingmar Bitter, Robert Van Uitert, Ivo Wolf, Luis Ibáñez, and Jan-Martin Kuhnigk. Comparison of four freely available frameworks for image processing and visualization that use ITK. IEEE Transactions on Visualization and Computer Graphics, 13 (3):483–493, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Braune:2000:FAE

[BW00] B. Braune and R. Wilhelm. Focusing in algorithm explanation. IEEE Transactions on Visualization and Computer Graphics, 6(1):1– 7, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0001.
pdf;   http://www.computer.
org/tvcg/tg2000/v0001abs.
htm.

Breen:2001:LSA

[BW01] D. E. Breen and R. T. Whitaker. A level-set approach for the metamorphosis of solid models. IEEE Transactions on Visualization and Computer Graphics, 7 (2):173–??, April/June 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0173.
pdf;   http://www.computer.
org/tvcg/tg2001/v0173abs.
htm.

Baciu:2003:IBT

[BW03] George Baciu and Wingo S. K. Wong. Image-based techniques in a hybrid collision detector. IEEE Transactions on Visualization and Computer Graphics, 9(2):254– 271, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0254abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0254.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0254.pdf.

Baciu:2004:IBC

[BW04] George Baciu and Wingo Sai-Keung Wong. Image-based collision detection for deformable cloth models. IEEE Transactions on Visualization and Computer Graphics, 10(6):649–663, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/06/v0649abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/06/
                                        v0649.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2004/06/v0649.pdf.

Bachthaler:2008:AOT

[BW08a] Sven Bachthaler and Daniel Weiskopf. Animation of orthogonal texture patterns for vector field visualization. IEEE Transactions on Visualization and Computer Graphics, 14(4):741– 755, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2008/04/ttg2008040741s.
                                        zip.

Bachthaler:2008:CS

[BW08b] Sven Bachthaler and Daniel Weiskopf. Continuous scatterplots. IEEE Transactions on Visualization and Computer Graphics, 14(6):1428– 1435, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Byron:2008:SGG

[BW08c] Lee Byron and Martin Wattenberg. Stacked graphs

— geometry & aesthetics. IEEE Transactions on Visualization and Computer Graphics, 14(6):1245–1252, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bae:2011:DEQ

[BW11] Juhee Bae and Benjamin Watson. Developing and evaluating quilts for the depiction of large layered graphs. IEEE Transactions on Visualization and Computer Graphics, 17(12):2268–2275, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bae:2014:RVG

[BW14] Juhee Bae and Benjamin Watson. Reinforcing visual grouping cues to communicate complex informational structure. IEEE Transactions on Visualization and Computer Graphics, 20(12):1973– 1982, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875960-
abs.html.

Beck:2017:WSG

[BW17] Fabian Beck and Daniel Weiskopf. Word-sized graphics for scientific texts. IEEE Transactions on Visualization and Computer Graphics, 23(6):1576–1587, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/06/07864462-
                                        abs.html.

Bhaniramka:2004:ICD

[BWC04] Praveen Bhaniramka, Rephael Wenger, and Roger Crawfis. Isosurface construction in any dimension using convex hulls. IEEE Transactions on Visualization and Computer Graphics, 10(2): 130–141, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/02/v0130abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/02/
                                        v0130.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2004/02/v0130.pdf.

Born:2010:ISS

[BWF^{+}10] Silvia Born, Alexander Wiebel, Jan Friedrich, Gerik Scheuermann, and Dirk Bartz. Illustrative stream surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1329–1338, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bernard:2013:MES

[BWK^{+}13] Jurgen Bernard, Nils Wilhelm, Bjorn Kruger, Thorsten May, Tobias Schreck, and Jorn Kohlhammer. Motion-Explorer: Exploratory search in human motion capture data based on hierarchical aggregation. IEEE Transactions on Visualization and Computer Graphics, 19(12):2257– 2266, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bremer:2010:ATB

[BWP^{+}10] Peer-Timo Bremer, Gunther H. Weber, Valerio Pascucci, Marc Day, and John B. Bell. Analyzing and tracking burning structures in lean premixed hydrogen flames. IEEE Transactions on Visualization and Computer Graphics, 16(2):248–260, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Blate:2019:IEK

[BWS^{+}19] Alex Blate, Mary Whitton, Montek Singh, Greg Welch, Andrei State, Turner Whitted, and Henry Fuchs. Implementation and evaluation of a 50 kHz, 28 μ s motionto-pose latency head tracking instrument. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1970–1980, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8658185/.

Bremer:2011:IEA

[BWT^{+}11] Peer-Timo Bremer, Gunther H. Weber, Julien Tierny, Valerio Pascucci, Marcus S. Day, and John B. Bell. Interactive exploration and analysis of large-scale simulations using topology-based data segmentation. IEEE Transactions on Visualization and Computer Graphics, 17(9): 1307–1324, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Benthin:2012:CSP

[BWW^{+}12] Carsten Benthin, Ingo Wald, Sven Woop, Manfred Ernst, and William R. Mark. Combining single and packet-ray tracing for arbitrary ray distributions on the Intel MIC architecture. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1438–1448, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bhattacharya:2017:IEV

[BWW^{+}17] Arindam Bhattacharya, Johannes Weissenbock, Rephael Wenger, Artem Amirkhanov, Johann Kastner, and Christoph Heinzl. Interactive exploration and visualization using MetaTracts extracted from carbon fiber reinforced composites. IEEE Transactions on Visualization and Computer Graphics, 23(8):1988– 2002, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/08/07493610-
abs.html.

Badri:2015:FEA

[BYA15] H. Badri, H. Yahia, and D. Aboutajdine. Fast edgeaware processing via first order proximal approximation. IEEE Transactions on Visualization and Computer Graphics, 21(6):743–755, June 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Borkin:2013:EFP

[BYB^{+}13] Michelle A. Borkin, Chelsea S.

Yeh, Madelaine Boyd, Peter Macko, Krzysztof Z. Gajos, Margo Seltzer, and Hanspeter Pfister. Evaluation of filesystem provenance visualization tools. IEEE Transactions on Visualization and Computer Graphics, 19(12):2476– 2485, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bista:2014:VBM

[BZGV14] Sujal Bista, Jiachen Zhuo, Rao P. Gullapalli, and Amitabh Varshney. Visualization of brain microstructure through spherical harmonics illumination of high fidelity spatio-angular fields. IEEE Transactions on Visualization and Computer Graphics, 20(12):2516– 2525, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875910-
abs.html.

Bekele:2013:UHA

[BZS^{+}13] Esubalew Bekele, Zhi Zheng, Amy Swanson, Julie Crittendon, Zachary Warren, and Nilanjan Sarkar. Understanding how adolescents with autism respond to facial expressions in virtual reality environments. IEEE Transactions on Visualization and Computer Graphics, 19(4): 711–720, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2000:PMI

[CA00] M. Chen and J. Arvo. Perturbation methods for interactive specular reflections. IEEE Transactions on Visualization and Computer Graphics, 6(3):253–264, July/ September 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0253.
pdf;   http://www.computer.
org/tvcg/tg2000/v0253abs.
htm.

Colburn:2013:IBR

[CAH^{+}13] A. Colburn, A. Agarwala, A. Hertzmann, B. Curless, and M. F. Cohen. Imagebased remodeling. IEEE Transactions on Visualization and Computer Graphics, 19 (1):56–66, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chung:2014:SSF

[CAN14] Haeyong Chung, Christopher Andrews, and Chris North. A survey of software frameworks for cluster-based large highresolution displays. IEEE Transactions on Visualization and Computer Graphics, 20 (8):1158–1177, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2018:ERE

[CAP18] Yu Ju Chen, Uri M. Ascher, and Dinesh K. Pai. Exponential Rosenbrock–Euler integrators for elastodynamic simulation. IEEE Transactions on Visualization and Computer Graphics, 24(10):2702– 2713, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08093708-
abs.html.

Carpendale:2019:VCA

[Car19] Sheelagh Carpendale. The 2018 Visualization Career Award. IEEE Transactions on Visualization and Computer Graphics, 25(1): xxviii, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08570932.
                                        pdf.

Chittaro:2015:AKR

[CB15] L. Chittaro and F. Buttussi. Assessing knowledge retention of an immersive serious game vs. a traditional education method in aviation safety. IEEE Transactions on Visualization and Computer Graphics, 21(4):529– 538, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cmolk:2019:RTE

[ČB19] L. Čmolík and J. Bittner. Real-time external labeling of ghosted views. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2458–2470, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carr:2006:HIS

[CBB06] Hamish Carr, Duffy Brian, and Denby Brian. On histograms and isosurface statistics. IEEE Transactions on Visualization and Computer Graphics, 12(5):1259– 1266, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2006:VSV

[CBH^{+}06] Min Chen, Ralf Botchen, Rudy Hashim, Daniel Weiskopf, Thomas Ertl, and Ian Thornton. Visual signatures in video visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1093–1100, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chattopadhyay:2007:HMC

[CBL07] Siddhartha Chattopadhyay, Suchendra M. Bhandarkar, and Kang Li. Human motion capture data compression by model-based indexing: a power aware approach. IEEE Transactions on Visualization and Computer Graphics, 13 (1):5–14, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cabral:2011:RPT

[CBLD11] Marcio Cabral, Nicolas Bonneel, Sylvain Lefebvre, and George Drettakis. Relighting photographs of tree canopies. IEEE Transactions on Visualization and Computer Graphics, 17(10):1459–1474, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Callahan:2006:PVR

[CBPS06] Steven P. Callahan, Louis Bavoil, Valerio Pascucci, and Claudio T. Silva. Progressive volume rendering of large unstructured grids. IEEE Transactions on Visualization and Computer Graphics, 12(5):1307–1314, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Collins:2007:VRR

[CC07] Christopher Collins and Sheelagh Carpendale. Vis-Link: Revealing relationships amongst visualizations. IEEE Transactions on Visualization and Computer Graphics, 13(6):1192–1199, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2008:SDF

[CC08] Chun-Yen Chen and Kuo-Young Cheng. A sharpnessdependent filter for recovering sharp features in repaired 3D mesh models. IEEE Transactions on Visualization and Computer Graphics, 14(1): 200–212, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Crouser:2012:ABF

[CC12] R. Jordon Crouser and Remco Chang. An affordance-based framework for human computation and human-computer collaboration. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2859–2868, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2012:ESS

[CCAL12] Jian Chen, Haipeng Cai, Alexander P. Auchus, and David H. Laidlaw. Effects of stereo and screen size on the legibility of threedimensional streamtube visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2130– 2139, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2011:EDV

[CCB11] Weifeng Chen, Wei Chen, and Hujun Bao. An efficient direct volume rendering approach for dichromats. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2144–2152, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cordonnier:2018:SMI

[CCB^{+}18] Guillaume Cordonnier, Marie-Paule Cani, Bedrich Benes, Jean Braun, and Eric Galin. Sculpting mountains: Interactive terrain modeling based on subsurface geology. IEEE Transactions on Visualization and Computer Graphics, 24(5):1756–1769, May 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07889056-
abs.html.

Carpendale:2014:GEI

[CCH14] Sheelagh Carpendale, Wei Chen, and Seok-Hee Hong. Guest Editors’ introduction: Special section on the IEEE Pacific Visualization Symposium. IEEE Transactions on Visualization and Computer Graphics, 20(8):1085– 1086, August 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2019:DLA

[CCJ^{+}19] Hsueh-Chien Cheng, Antonio Cardone, Somay Jain, Eric Krokos, Kedar Narayan, Sriram Subramaniam, and Amitabh Varshney. Deeplearning-assisted volume visualization. IEEE Transactions on Visualization and Computer Graphics, 25(2):1378– 1391, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2019/02/08265023-
abs.html.

Chen:2007:SGS

[CCK07] Yuan Chen, Jonathan Cohen, and Julian Krolik. Similarityguided streamline placement with error evaluation. IEEE Transactions on Visualization and Computer Graphics, 13(6):1448–1455, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2016:PVA

[CCL^{+}16] Qing Chen, Yuanzhe Chen, Dongyu Liu, Conglei Shi, Yingcai Wu, and Huamin Qu. PeakVizor: Visual analytics of peaks in video clickstreams from massive open online courses. IEEE Transactions on Visualization and Computer Graphics, 22(10): 2315–2330, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Clarke:2011:AGC

[CCM11] Lyndsey Clarke, Min Chen, and Benjamin Mora. Automatic generation of 3D caricatures based on artistic deformation styles. IEEE Transactions on Visualization and Computer Graphics, 17(6): 808–821, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2012:VRA

[CCM12] Carlos D. Correa, Tarik Crnovrsanin, and Kwan-Liu Ma. Visual reasoning about social networks using centrality sensitivity. IEEE Transactions on Visualization and Computer Graphics, 18(1): 106–120, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chan:2013:GSS

[CCM13a] Yu-Hsuan Chan, Carlos D. Correa, and Kwan-Liu Ma. The generalized sensitivity scatterplot. IEEE Transactions on Visualization and Computer Graphics, 19(10): 1768–1781, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2013:SFM

[CCM^{+}13b] Zhi-Quan Cheng, Yin Chen, Ralph R. Martin, Yu-Kun Lai, and Aiping Wang. SuperMatching: Feature matching using supersymmetric geometric constraints. IEEE Transactions on Visualization and Computer Graphics, 19(11):1885–1894, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2014:VAE

[CCM^{+}14] Haidong Chen, Wei Chen, Honghui Mei, Zhiqi Liu, Kun Zhou, Weifeng Chen, Wentao Gu, and Kwan-Liu Ma. Visual abstraction and exploration of multi-class scatterplots. IEEE Transactions on Visualization and Computer Graphics, 20(12):1683– 1692, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875982-
abs.html.

Choi:2014:VDS

[CCQ^{+}14] Hyungsuk Choi, Woohyuk Choi, Tran Minh Quan, David G. C. Hildebrand, Hanspeter Pfister, and Won-Ki Jeong. Vivaldi: A domain-specific language for volume processing and visualization on distributed heterogeneous systems. IEEE Transactions on Visualization and Computer Graphics, 20(12):2407– 2416, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875916-
abs.html.

Corsini:2012:EFS

[CCS12] Massimiliano Corsini, Paolo Cignoni, and Roberto Scopigno. Efficient and flexible sampling with blue noise properties of triangular meshes. IEEE Transactions on Visualization and Computer Graphics, 18 (6):914–924, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Camisetty:2019:EWB

[CCSK19] Akhilesh Camisetty, Chaitanya Chandurkar, Maoyuan Sun, and David Koop. Enhancing Web-based analytics applications through provenance. IEEE Transactions on Visualization and Computer Graphics, 25(1):131– 141, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08500765-
                                        abs.html.

Choi:2009:CCD

[CCW^{+}09] Yi-King Choi, Jung-Woo Chang, Wenping Wang, Myung-Soo Kim, and Gershon Elber. Continuous collision detection for ellipsoids. IEEE Transactions on Visualization and Computer Graphics, 15(2):311– 325, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carr:2014:JCN

[CD14] Hamish Carr and David Duke. Joint contour nets. IEEE Transactions on Visualization and Computer Graphics, 20 (8):1100–1113, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cavallo:2019:CGV

[CD19] Marco Cavallo and Cagatay Demiralp. Clustrophile 2: Guided visual clustering analysis. IEEE Transactions on Visualization and Computer Graphics, 25(1):267– 276, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440035-
abs.html.

Cotin:1999:RTE

[CDA99] Stéphane Cotin, Hervé Delingette, and Nicholas Ayache. Realtime elastic deformations of soft tissues for surgery simulation. IEEE Transactions on Visualization and Computer Graphics, 5(1):62–73, January/March 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0062.
pdf;   http://www.computer.
org/tvcg/tg1999/v0062abs.
htm.

Chakravarthula:2018:FAF

[CDAF18] Praneeth Chakravarthula, David Dunn, Kaan Akşit, and Henry Fuchs. FocusAR: Autofocus augmented reality eyeglasses for both real world and virtual imagery. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2906–2916, November 2018.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8458263/.

Chalmers:2014:IDE

[CDBR14] Alan Chalmers, Kurt Debattista, and Thomas Bashford-Rogers. Importance driven environment map sampling. IEEE Transactions on Visualization and Computer Graphics, 20(6):907–918, June 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cammarano:2007:VHD

[CDC^{+}07] Mike Cammarano, Xin (Luna) Dong, Bryan Chan, Jeff Klingner, Justin Talbot, Alon Halevey, and Pat Hanrahan. Visualization of heterogeneous data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1200–1207, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Candela:2018:RVE

[CDDS18] Massimo Candela, Marco Di Bartolomeo, Giuseppe Di Battista, and Claudio Squarcella. Radian: Visual exploration of traceroutes. IEEE Transactions on Visualization and Computer Graphics, 24(7):2194–2208, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07953527-
abs.html.

Clarkson:2009:RVS

[CDF09] Edward Clarkson, Krishna Desai, and James Foley. ResultMaps: Visualization for search interfaces. IEEE Transactions on Visualization and Computer Graphics, 15(6):1057–1064, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chevalier:2014:SES

[CDF14] Fanny Chevalier, Pierre Dragicevic, and Steven Franconeri. The not-so-staggering effect of staggered animated transitions on visual tracking. IEEE Transactions on Visualization and Computer Graphics, 20(12):2241– 2250, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876010-
abs.html.

Chang:2019:P

[CDF^{+}19] Remco Chang, Tim Dwyer, Issei Fujishiro, Petra Isenberg, Steve Franconeri, Huamin Qu, Tobias Schreck, Daniel Weiskopf, and Gunther H. Weber. Preface. IEEE Transactions on Visualization and Computer Graphics, 25(1): xii–xvii, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08570935.
pdf.

Chen:2004:FAS

[CDK04] Baoquan Chen, Frank Dachille, and Arie E. Kaufman. Footprint area sampled texturing. IEEE Transactions on Visualization and Computer Graphics, 10(2):230– 240, March/April 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0230abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0230.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/02/v0230.pdf.

Cordeil:2017:ICA

[CDK^{+}17] Maxime Cordeil, Tim Dwyer, Karsten Klein, Bireswar Laha, Kim Marriott, and Bruce H. Thomas. Immersive collaborative analysis of network connectivity: CAVEstyle or head-mounted display? IEEE Transactions on Visualization and Computer Graphics, 23(1):441–450, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cher:2016:VAR

[CDL^{+}16] C. Cher, S. Dutta, X. Liu, G. Heinlein, H. Shen, and J. Chen. Visualization and analysis of rotating stall for transonic jet engine simulation. IEEE Transactions on Visualization and Computer Graphics, 22(1):847–856, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cignoni:2004:SRQ

[CDM^{+}04] Paolo Cignoni, Leila De Floriani, Paola Magillo, Enrico Puppo, and Roberto Scopigno. Selective refinement queries for volume visualization of unstructured tetrahedral meshes. IEEE Transactions on Visualization and Computer Graphics, 10(1):29–45, January/ February 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0029abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/01/
v0029.pdf.

Cortese:2006:TVP

[CDM^{+}06] Pier Francesco Cortese, Giuseppe Di Battista, Antonello Moneta, Maurizio Patrignani, and Maurizio Pizzonia. Topographic visualization of prefix propagation in the Internet. IEEE Transactions on Visualization and Computer Graphics, 12(5):725–732, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Crampes:2009:VSP

[CdOKRV09] Michel Crampes, Jeremy de Oliveira-Kumar, Sylvie Ranwez, and Jean Villerd. Visualizing social photos on a Hasse diagram for eliciting relations and indexing new photos. IEEE Transactions on Visualization and Computer Graphics, 15(6):985–992, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chao:2018:RDD

[CDR^{+}18] Q. Chao, Z. Deng, J. Ren, Q. Ye, and X. Jin. Realistic data-driven traffic flow animation using texture synthesis. IEEE Transactions on Visualization and Computer Graphics, 24(2):1167–1178, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2012:MSC

[CDS^{+}12] Guoning Chen, Qingqing Deng, Andrzej Szymczak, Robert S. Laramee, and Eugene Zhang. Morse set classification and hierarchical refinement using Conley index. IEEE Transactions on Visualization and Computer Graphics, 18(5):767–782, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cho:2016:VVA

[CDW^{+}16] I. Cho, W. Dou, D. X. Wang, E. Sauda, and W. Ribarsky. VAiRoma: A visual analytics system for making sense of places, times, and events in Roman history. IEEE Transactions on Visualization and Computer Graphics, 22(1): 210–219, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2009:NII

[CDZ^{+}09] Wei Chen, Zi’ang Ding, Song Zhang, Anna MacKay-Brandt, Stephen Correia, Huamin Qu, John Allen Crow, David F. Tate, Zhicheng Yan, and Qunsheng Peng. A novel interface for interactive exploration of DTI fibers. IEEE Transactions on Visualization and Computer Graphics, 15(6):1433–1440, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cline:2001:TDT

[CE01] D. Cline and P. K. Egbert.

Terrain decimation through quadtree morphing. IEEE Transactions on Visualization and Computer Graphics, 7(1): 62–??, January/March 2001.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2001/pdf/v0062.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0062abs.
                                        htm.

Cole:2010:TFM

[CF10] Forrester Cole and Adam Finkelstein. Two fast methods for high-quality line visibility. IEEE Transactions on Visualization and Computer Graphics, 16(5):707–717, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Crouser:2017:TTT

[CFEC17] R. Jordan Crouser, Lyndsey Franklin, Alex Endert, and Kris Cook. Toward theoretical techniques for measuring the use of human effort in visual analytic systems. IEEE Transactions on Visualization and Computer Graphics, 23(1):121–130, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chu:2009:GGB

[CFHH09] Alan Chu, Chi-Wing Fu, Andrew Hanson, and Pheng-Ann Heng. GL4D: a GPUbased architecture for interactive 4D visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1587–1594, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coquillart:2012:MPC

[CFK12] Sabine Coquillart, Steven Feiner, and Kiyoshi Kiyokawa. Message from the Paper Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 18(4):vi, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carnecky:2013:STI

[CFM^{+}13] Robert Carnecky, Raphael Fuchs, Stephanie Mehl, Yun Jang, and Ronald Peikert. Smart transparency for illustrative visualization of complex flow surfaces. IEEE Transactions on Visualization and Computer Graphics, 19 (5):838–851, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cipriano:2007:MSA

[CG07] Gregory Cipriano and Michael Gleicher. Molecular surface abstraction. IEEE Transactions on Visualization and Computer Graphics, 13(6):1608–1615, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1608s.
                                        avi.

Cipriano:2008:TSE

[CG08] Gregory Cipriano and Michael Gleicher. Text scaffolds for effective surface labeling. IEEE Transactions on Visualization and Computer Graphics, 14(6):1675–1682, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correll:2014:EBC

[CG14] Michael Correll and Michael Gleicher. Error bars considered harmful: Exploring alternate encodings for mean and error. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2142–2151, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875915-
                                        abs.html.

Chen:2016:WMV

[CG16] Min Chen and Amos Golan. What may visualization processes optimize? IEEE Transactions on Visualization and Computer Graphics, 22(12): 2619–2632, December 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/

csdl/trans/tg/2016/12/07368928-
abs.html.

Cosco:2013:VHM

[CGB^{+}13] F. Cosco, C. Garre, F. Bruno, M. Muzzupappa, and M. A. Otaduy. Visuo-haptic mixed reality with unobstructed tool-hand integration. IEEE Transactions on Visualization and Computer Graphics, 19 (1):159–172, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Camp:2011:SIU

[CGC^{+}11] David Camp, Christoph Garth, Hank Childs, Dave Pugmire, and Kenneth I. Joy. Streamline integration using MPI-hybrid parallelism on a large multicore architecture. IEEE Transactions on Visualization and Computer Graphics, 17(11):1702– 1713, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cani-Gascuel:1997:ADM

[CGD97] M.-P. Cani-Gascuel and M. Desbrun. Animation of deformable models using implicit surfaces. IEEE Transactions on Visualization and Computer Graphics, 3(1):39– 50, January/March 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/

tg/books/tg1997/pdf/v0039.
pdf;   http://www.computer.
org/tvcg/tg1997/v0039abs.
htm.

Chu:2018:MIS

[CGGR18] David Chu, Joseph L. Gabbard, Jens Grubert, and Holger Regenbrecht. Message from the ISMAR 2018 Science and Technology Program Chairs and TVCG Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 24(11):2844– 2845, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8514064/.

Chen:2019:SBS

[CGH^{+}19] Wei Chen, Fangzhou Guo, Dongming Han, Jacheng Pan, Xiaotao Nie, Jiazhi Xia, and Xiaolong Zhang. Structurebased suggestive exploration: A new approach for effective exploration of large networks. IEEE Transactions on Visualization and Computer Graphics, 25(1):555– 565, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440813-
abs.html.

Chen:2019:ITA

[CGJM19] Min Chen, Kelly Gaither, Nigel W. John, and Brian Mccann. An informationtheoretic approach to the cost-benefit analysis of visualization in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 25(1): 32–42, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440821-
abs.html.

Cai:2017:SAA

[CGL^{+}17] Yiqi Cai, Xiaohu Guo, Yang Liu, Wenping Wang, Weihua Mao, and Zichun Zhong. Surface approximation via asymptotic optimal geometric partition. IEEE Transactions on Visualization and Computer Graphics, 23(12): 2613–2626, December 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07728112-
abs.html.

Ceneda:2017:CGV

[CGM^{+}17] Davide Ceneda, Theresia Gschwandtner, Thorsten May, Silvia Miksch, Hans-Jörg Schulz, Marc Streit, and Christian Tominski. Characterizing guidance in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 23(1):111–120, January 2017. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cao:2011:DIV

[CGSQ11] Nan Cao, David Gotz, Jimeng Sun, and Huamin Qu. DI-CON: Interactive visual analysis of multidimensional clusters. IEEE Transactions on Visualization and Computer Graphics, 17(12):2581– 2590, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Collomosse:2003:CSR

[CH03] John P. Collomosse and Peter M. Hall. Cubist style rendering from photographs. IEEE Transactions on Visualization and Computer Graphics, 9(4):443–453, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/04/v0443abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/04/
                                        v0443.pdf.

Correll:2017:SBW

[CH17] Michael Correll and Jeffrey Heer. Surprise! Bayesian weighting for de-biasing thematic maps. IEEE Transactions on Visualization and Computer Graphics, 23(1): 651–660, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cochran:1995:FVC

[CHF95] Wayne O. Cochran, John C.

Hart, and Patrick J. Flynn. Fractal volume compression. Internal report, Washington State University, School of EECS, Pullman, WA, USA, October 1995. See [CHF96].

Cochran:1996:FVC

[CHF96] Wayne O. Cochran, John C.

Hart, and Patrick J. Flynn. Fractal volume compression. IEEE Transactions on Visualization and Computer Graphics, 2(4):313– 322, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0313.
pdf;   http://www.computer.
org/tvcg/tg1996/v0313abs.
htm.

Chittaro:2016:DSG

[Chi16] L. Chittaro. Designing serious games for safety education: “learn to brace” versus traditional pictorials for aircraft passengers. IEEE Transactions on Visualization and Computer Graphics, 22 (5):1527–1539, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carmel:2004:CHE

[CHK04] Liran Carmel, David Harel, and Yehuda Koren. Combining hierarchy and energy drawing directed graphs. IEEE Transactions on Visualization and Computer Graphics, 10(1):46–57, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0046abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/01/
v0046.pdf.

Chao:2009:HCS

[ChLYL09] Min-Wen Chao, Chao hung Lin, Cheng-Wei Yu, and Tong-Yee Lee. A high capacity 3D steganography algorithm. IEEE Transactions on Visualization and Computer Graphics, 15(2):274– 284, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2011:CGE

[CHM11] Carlos D. Correa, Robert Hero, and Kwan-Liu Ma. A comparison of gradient estimation methods for volume rendering on unstructured meshes. IEEE Transactions on Visualization and Computer Graphics, 17(3): 305–319, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Christensen:2003:AIR

[Chr03] Per H. Christensen. Adjoints and importance in rendering: an overview. IEEE Transactions on Visualization and Computer Graphics, 9(3):329– 340, July/September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0329abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0329.pdf.

Chen:2018:VVA

[CHW^{+}18] Wei Chen, Zhaosong Huang, Feiran Wu, Minfeng Zhu, Huihua Guan, and Ross Maciejewski. VAUD: A visual analysis approach for exploring spatio-temporal urban data. IEEE Transactions on Visualization and Computer Graphics, 24(9):2636– 2648, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08054703-
abs.html.

Callahan:2005:HAV

[CICS05] Steven P. Callahan, Milan Ikits, Joao L. D. Comba, and Claudio T. Silva. Hardwareassisted visibility sorting for unstructured volume rendering. IEEE Transactions on Visualization and Computer Graphics, 11(3):285– 295, May/June 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2005/03/v0285s.
                                        pdf.

Chen:2010:ITF

[CJ10] Min Chen and Heike Jaenicke. An information-theoretic framework for visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1206–1215, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Caban:2007:TBF

[CJR07] Jesus Caban, Alark Joshi, and Penny Rheingans. Texturebased feature tracking for effective time-varying data visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6): 1472–1479, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2007/
                                        06/v1472s.avi.

Craciun:2005:SDW

[CJTM05] Gheorghe Craciun, Ming Jiang, David Thompson, and Raghu Machiraju. Spatial domain wavelet design for feature preservation in computational data sets. IEEE Transactions on Visualization and Computer Graphics, 11(2): 149–159, March/April 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Choe:2005:SWM

[CK05a] Byoungwon Choe and Hyeong-Seok Ko. A statistical wisp model and pseudophysical approaches for interactive hairstyle generation. IEEE Transactions on Visualization and Computer Graphics, 11(2):160–170, March/April 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Choi:2005:MWR

[CK05b] Min Gyu Choi and Hyeong-Seok Ko. Modal warping: Real-time simulation of large rotational deformation and manipulation. IEEE Transactions on Visualization and Computer Graphics, 11(1):91– 101, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0091abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0091.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0091.pdf.

Connor:2010:FCN

[CK10] Michael Connor and Piyush Kumar. Fast construction of k-nearest neighbor graphs for point clouds. IEEE Transactions on Visualization and Computer Graphics, 16(4): 599–608, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chiosa:2011:GBM

[CK11] Iurie Chiosa and Andreas Kolb. GPU-based multilevel clustering. IEEE Transactions on Visualization and Computer Graphics, 17(2): 132–145, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Camporesi:2016:EAS

[CK16] C. Camporesi and M. Kallmann. The effects of avatars, stereo vision and display size on reaching and motion reproduction. IEEE Transactions on Visualization and Computer Graphics, 22(5):1592– 1604, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

chulKwon:2012:ERT

[cKJG^{+}12] Bum chul Kwon, Waqas Javed, Sohaib Ghani, Niklas Elmqvist, Ji Soo Yi, and David S. Ebert. Evaluating the role of time in investigative analysis of document collections. IEEE Transactions on Visualization and Computer Graphics, 18(11):1992– 2004, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Choe:2009:RAM

[CKLL09] Sungyul Choe, Junho Kim, Haeyoung Lee, and Seungyong Lee. Random accessible mesh compression using mesh chartification. IEEE Transactions on Visualization and Computer Graphics, 15(1): 160–173, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coquillart:2014:MPC

[CKSB14] Sabine Coquillart, Kiyoshi Kiyokawav, J. Edward Swan, and Doug Bowman. Message from the paper chairs and guest editors. IEEE Transactions on Visualization and Computer Graphics, 20 (4):vi, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2012:DTV

[CKW^{+}12] Guoning Chen, Vivek Kwatra, Li-Yi Wei, Charles D. Hansen, and Eugene Zhang. Design of 2D time-varying vector fields. IEEE Transactions on Visualization and Computer Graphics, 18(10):

1717–1730, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chi:2006:SAP

[CL06] Ming-Te Chi and Tong-Yee Lee. Stylized and abstract painterly rendering system using a multiscale segmented sphere hierarchy. IEEE Transactions on Visualization and Computer Graphics, 12(1):61–72, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2006/01/v0061s.
                                        mp4.

Chu:2009:MMS

[CL09] Hung-Kuo Chu and Tong-Yee Lee. Multiresolution mean shift clustering algorithm for shape interpolation. IEEE Transactions on Visualization and Computer Graphics, 15(5):853–866, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2011:TRT

[CL11] Carlos Correa and Peter Lindstrom. Towards robust topology of sparsely sampled data. IEEE Transactions on Visualization and Computer Graphics, 17(12):1852– 1861, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Capannini:2018:ACC

[CL18] Gabriele Capannini and Thomas Larsson. Adaptive collision culling for massive simulations by a parallel and context-aware sweep and prune algorithm. IEEE Transactions on Visualization and Computer Graphics, 24(7):2064–2077, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07935404-
abs.html.

Chao:2012:HMR

[CLAL12] Min-Wen Chao, Chao-Hung Lin, Jackie Assa, and Tong-Yee Lee. Human motion retrieval from hand-drawn sketch. IEEE Transactions on Visualization and Computer Graphics, 18(5):729–740, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2011:TSS

[CLB11] Carlos Correa, Peter Lindstrom, and Peer-Timo Bremer. Topological spines: a structure-preserving visual representation of scalar fields. IEEE Transactions on Visualization and Computer Graphics, 17(12):1842– 1851, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chuah:2013:AMR

[CLB13] Joon Hao Chuah, Benjamin Lok, and Erik Black. Applying mixed reality to simulate vulnerable populations for practicing clinical communication skills. IEEE Transactions on Visualization and Computer Graphics, 19(4): 539–546, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cohe:2016:SCA

[CLB^{+}16] A. Cohe, B. Liutkus, G. Bailly, J. Eagan, and E. Lecolinet. SchemeLens: A content-aware vector-based fisheye technique for navigating large systems diagrams. IEEE Transactions on Visualization and Computer Graphics, 22(1): 330–338, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chi:2015:MWC

[CLC^{+}15] M. Chi, S. Lin, S. Chen, C. Lin, and T. Lee. Morphable word clouds for timevarying text data visualization. IEEE Transactions on Visualization and Computer Graphics, 21(12):1415– 1426, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cao:2012:SDS

[CLCQ12] Juan Cao, Xin Li, Zhonggui Chen, and Hong Qin. Spherical DCB-spline surfaces with hierarchical and adaptive knot insertion. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1290–1303, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coffey:2013:DDI

[CLEK13] Dane Coffey, Chi-Lun Lin, Arthur G. Erdman, and Daniel F. Keefe. Design by dragging: An interface for creative forward and inverse design with simulation ensembles. IEEE Transactions on Visualization and Computer Graphics, 19(12):2783– 2791, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cao:2016:UMV

[CLG16] Nan Cao, Yu-Ru Lin, and D. Gotz. UnTangle map: Visual analysis of probabilistic multi-label data. IEEE Transactions on Visualization and Computer Graphics, 22(2): 1149–1163, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correll:2019:LGM

[CLKS19] Michael Correll, Mingwei Li, Gordon Kindlmann, and Carlos Scheidegger. Looks good to me: Visualizations as sanity checks. IEEE Transactions on Visualization and Computer Graphics, 25(1): 830–839, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440818-
abs.html.

Chan:2008:HMP

[CLL08] Addison Chan, Rynson Lau, and Lewis Li. Hand motion prediction for distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 14(1):146–159, January/ February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cirio:2017:YLC

[CLMO17] Gabriel Cirio, Jorge Lopez-Moreno, and Miguel A. Otaduy. Yarn-level cloth simulation with sliding persistent contacts. IEEE Transactions on Visualization and Computer Graphics, 23(2):1152– 1162, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2017/02/07516643-
abs.html.

Choo:2013:UUD

[CLRP13] Jaegul Choo, Changhyun Lee, Chandan K. Reddy, and Haesun Park. UTOPIAN: Userdriven topic modeling based on interactive nonnegative matrix factorization. IEEE Transactions on Visualization and Computer Graphics, 19(12):1992–2001, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chrysanthou:2007:GEI

[CLS07] Yiorgos L. Chrysanthou, Rynson W. H. Lau, and Gurminder Singh. Guest Editors’ introduction: Special section on ACM VRST 2005. IEEE Transactions on Visualization and Computer Graphics, 13(1):3–4, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/01/v0003.
pdf.

Cao:2012:WTS

[CLS^{+}12] Nan Cao, Yu-Ru Lin, Xiaohua Sun, David Lazer, Shixia Liu, and Huamin Qu. Whisper: Tracing the spatiotemporal process of information diffusion in real time. IEEE Transactions on Visualization and Computer Graphics, 18(12):

2649–2658, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coquillart:2013:IVR

[CLS13a] Sabine Coquillart, Joseph J. LaViola, and Dieter Schmalstieg. IEEE Virtual Reality Conference 2013 [table of contents]. IEEE Transactions on Visualization and Computer Graphics, 19(4):i–iii, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coquillart:2013:MPC

[CLS13b] Sabine Coquillart, Joseph J. LaViola, and Dieter Schmalstieg. Message from the Paper Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 19(4):vi, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chandak:2008:AFA

[CLT^{+}08] Anish Chandak, Christian Lauterbach, Micah Taylor, Zhimin Ren, and Dinesh Manocha. AD-Frustum: Adaptive frustum tracing for interactive sound propagation. IEEE Transactions on Visualization and Computer Graphics, 14(6):1707– 1722, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2011:TTB

[CLT^{+}11] Weiwei Cui, Shixia Liu, Li Tan, Conglei Shi, Yangqiu Song, Zekai Gao, Huamin Qu, and Xin Tong. TextFlow: Towards better understanding of evolving topics in text. IEEE Transactions on Visualization and Computer Graphics, 17(12):2412–2421, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cho:2018:MSV

[CLW18] Isaac Cho, Jialei Li, and Zachary Wartell. Multi-scale 7DOF view adjustment. IEEE Transactions on Visualization and Computer Graphics, 24 (3):1331–1344, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2014:HHT

[CLWW14] Weiwei Cui, Shixia Liu, Zhuofeng Wu, and Hao Wei. How hierarchical topics evolve in large text corpora. IEEE Transactions on Visualization and Computer Graphics, 20(12):2281– 2290, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875938-
abs.html.

Chen:2003:RRA

[CLZ^{+}03] Yanyun Chen, Stephen Lin, Hua Zhong, Ying-Qing Xu, Baining Guo, and Heung-Yeung Shum. Realistic rendering and animation of knitwear. IEEE Transactions on Visualization and Computer Graphics, 9(1):43–55, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0043abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0043.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/01/v0043.pdf.

Cao:2018:VVA

[CLZ^{+}18] N. Cao, C. Lin, Q. Zhu, Y. R. Lin, X. Teng, and X. Wen. Voila: Visual anomaly detection and monitoring with streaming spatiotemporal data. IEEE Transactions on Visualization and Computer Graphics, 24 (1):23–33, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chou:2002:VDC

[CM02] Peter H. Chou and Teresa H. Meng. Vertex data compression through vector quantization. IEEE Transactions on Visualization and Computer Graphics, 8(4):373–382, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0373abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0373.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0373.pdf.

Correa:2008:SBT

[CM08] Carlos Correa and Kwan-Liu Ma. Size-based transfer functions: a new volume exploration technique. IEEE Transactions on Visualization and Computer Graphics, 14(6):1380–1387, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2009:OSV

[CM09] Carlos Correa and Kwan-Liu Ma. The occlusion spectrum for volume classification and visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1465–1472, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carr:2010:SAT

[CM10] Hamish Carr and Nelson Max.

Subdivision analysis of the trilinear interpolant. IEEE Transactions on Visualization and Computer Graphics, 16(4):533–547, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2011:VHV

[CM11] Carlos D. Correa and Kwan-Liu Ma. Visibility histograms and visibility-driven transfer functions. IEEE Transactions on Visualization and Computer Graphics, 17(2):192– 204, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chentanez:2014:MCE

[CM14] Nuttapong Chentanez and Matthias Muller. Massconserving Eulerian liquid simulation. IEEE Transactions on Visualization and Computer Graphics, 20(1): 17–29, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2016:DCM

[CM16] S. Cheng and K. Mueller. The data context map: Fusing data and attributes into a unified display. IEEE Transactions on Visualization and Computer Graphics, 22(1): 121–130, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Castanie:2006:DSM

[CMCL06] Laurent Castanie, Christophe Mion, Xavier Cavin, and Bruno Levy. Distributed shared memory for roaming large volumes. IEEE Transactions on Visualization and Computer Graphics, 12(5):1299–1306, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chentanez:2012:MFP

[CMF12] Nuttapong Chentanez and Matthias Muller-Fischer. A multigrid fluid pressure solver handling separating solid boundary conditions. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1191–1201, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2018:RTL

[CMF^{+}18] Guowen Chen, Cong Ma, Zhencheng Fan, Xiwen Cui, and Hongen Liao. Realtime lens based rendering algorithm for supermultiview integral photography without image resampling. IEEE Transactions on Visualization and Computer Graphics, 24(9):2600– 2609, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08052161-
abs.html.

Crassin:2016:AGB

[CMFL16] Cyril Crassin, Morgan McGuire, Kayvon Fatahalian, and Aaron Lefohn. Aggregate G-buffer anti-aliasing — extended version. IEEE Transactions on Visualization and Computer Graphics, 22(10): 2215–2228, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cirio:2011:SDF

[CMHL11] Gabriel Cirio, Maud Marchal, Sebastien Hillaire, and Anatole Lecuyer. Six degreesof-freedom haptic interaction with fluids. IEEE Transactions on Visualization and Computer Graphics, 17(11): 1714–1727, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chentanez:2015:CEH

[CMK15] Nuttapong Chentanez, Matthias Muller, and Tae-Yong Kim. Coupling 3D Eulerian, heightfield and particle methods for interactive simulation of large scale liquid phenomena. IEEE Transactions on Visualization and Computer Graphics, 21(10):1116– 1128, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07132780-
abs.html.

Chen:2007:VFE

[CML^{+}07] Guoning Chen, Konstantin Mischaikow, Robert S. Laramee, Pawel Pilarczyk, and Eugene Zhang. Vector field editing and periodic orbit extraction using Morse decomposition. IEEE Transactions on Visualization and Computer Graphics, 13(4):769– 785, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Coffey:2012:ISW

[CML^{+}12] Dane Coffey, Nicholas Malbraaten, Trung Bao Le, Iman Borazjani, Fotis Sotiropoulos, Arthur G. Erdman, and Daniel F. Keefe. Interactive slice WIM: Navigating and interrogating volume data sets using a multisurface, multitouch VR interface. IEEE Transactions on Visualization and Computer Graphics, 18(10):1614–1626, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2008:EMD

[CMLZ08] Guoning Chen, Konstantin Mischaikow, Robert S. Laramee, and Eugene Zhang. Efficient Morse decompositions of vector fields. IEEE Transactions on Visualization and Computer Graphics, 14(4):848– 862, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cignoni:1997:SIE

[CMM^{+}97] P. Cignoni, P. Marino, C. Montani, E. Puppo, and R. Scopigno. Speeding up isosurface extraction using interval trees. IEEE Transactions on Visualization and Computer Graphics, 3(2):158– 170, April/June 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0158.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0158abs.
                                        htm.

Clyne:2013:PBF

[CMN13] John Clyne, Pablo Mininni, and Alan Norton. Physicallybased feature tracking for CFD data. IEEE Transactions on Visualization and Computer Graphics, 19(6): 1020–1033, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2009:COD

[CMP09] Jin Chen, Alan M. MacEachren, and Donna J. Peuquet. Constructing overview + detail dendrogram-matrix views. IEEE Transactions on Visualization and Computer Graphics, 15(6):889–896, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2014:ADI

[CMP14] Jian Cui, Zhiqiang Ma, and Voicu Popescu. Animated depth images for interactive remote visualization of time-varying data sets. IEEE Transactions on Visualization and Computer Graphics, 20(11):1474– 1489, November 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06671915-
abs.html.

Comport:2006:RTM

[CMPC06] Andrew I. Comport, Eric Marchand, Muriel Pressigout, and François Chaumette. Real-time markerless tracking for augmented reality: The virtual visual servoing framework. IEEE Transactions on Visualization and Computer Graphics, 12(4):615– 628, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cignoni:1997:MRV

[CMPS97] P. Cignoni, C. Montani, E. Puppo, and R. Scopigno. Multiresolution representation and visualization of volume data. IEEE Transactions on Visualization and Computer Graphics, 3(4):352– 369, October/December 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0352.
pdf;   http://www.computer.
org/tvcg/tg1997/v0352abs.
htm.

Cignoni:2003:EMM

[CMRS03] Paolo Cignoni, Claudio Montani, Claudio Rocchini, and Roberto Scopigno. External memory management and simplification of huge meshes. IEEE Transactions on Visualization and Computer Graphics, 9(4):525–537, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0525abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0525.pdf.

Carr:2006:ACS

[CMS06] Hamish Carr, Torsten Moller, and Jack Snoeyink. Artifacts caused by simplicial subdivision. IEEE Transactions on Visualization and Computer Graphics, 12(2):231– 242, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cook:2004:ISV

[CMSW04] Richard Cook, Nelson Max, Cláudio T. Silva, and Peter L. Williams. Image-space visibility ordering for cell projection volume rendering of unstructured data. IEEE Transactions on Visualization and Computer Graphics, 10(6):695–707, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0695abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0695.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0695.pdf.

Cohen-Or:2003:SVW

[COCSD03] Daniel Cohen-Or, Yiorgos L.

Chrysanthou, Cláudio T. Silva, and Frédo Durand. A survey of visibility for walkthrough applications. IEEE Transactions on Visualization and Computer Graphics, 9(3):412–431, July/ September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0412abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0412.pdf.

Chandler:2015:IBP

[COJ15] Jennifer Chandler, Harald Obermaier, and Kenneth I. Joy. Interpolationbased pathline tracing in particle-based flow visualization. IEEE Transactions on Visualization and Computer Graphics, 21(1):68– 80, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06817592-
abs.html.

Cirio:2013:KEV

[COMP13] Gabriel Cirio, Anne-Helene Olivier, Maud Marchal, and Julien Pettre. Kinematic evaluation of virtual walking trajectories. IEEE Transactions on Visualization and Computer Graphics, 19(4): 671–680, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cohen-Or:1996:RTP

[CORLS96] Daniel Cohen-Or, Eran Rich, Uri Lerner, and Victor Shenkar. A real-time photorealistic visual flythrough. IEEE Transactions on Visualization and Computer Graphics, 2(3):255–265, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://dlib.computer.
org/tg/books/tg1996/pdf/
v0255.pdf;   http://www.
computer.org/tvcg/tg1996/
v0255abs.htm.

Cox:2011:VCA

[Cox11] Amanda Cox. VisWeek capstone address. IEEE Transactions on Visualization and Computer Graphics, 17(12): xxiv, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Collins:2009:BSR

[CPC09] Christopher Collins, Gerald Penn, and Sheelagh Carpendale. Bubble sets: Revealing set relations with isocontours over existing visualizations. IEEE Transactions on Visualization and Computer Graphics, 15(6):1009–1016, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cipriano:2009:MSS

[CPG09] Gregory Cipriano, George N.

Phillips, Jr., and Michael Gleicher. Multi-scale surface descriptors. IEEE Transactions on Visualization and Computer Graphics, 15(6):1201– 1208, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carlson:2015:VTL

[CPG^{+}15] P. Carlson, A. Peters, S. B.

Gilbert, J. M. Vance, and A. Luse. Virtual training: Learning transfer of assembly tasks. IEEE Transactions on Visualization and Computer Graphics, 21(6): 770–782, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chhugani:2005:VHF

[CPK^{+}05] Jatin Chhugani, Budirijanto Purnomo, Shankar Krishnan, Jonathan Cohen, Suresh Venkatasubramanian, David S. Johnson, and Subodh Kumar. vLOD: Highfidelity walkthrough of large virtual environments. IEEE Transactions on Visualization and Computer Graphics, 11(1):35–47, January/ February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0035abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0035.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0035.pdf.

Choi:2015:VTC

[CPW^{+}15] Jungu Choi, Deok Gun Park, Yuet Ling Wong, Eli Fisher, and Niklas Elmqvist. Vis-Dock: A toolkit for crosscutting interactions in visualization. IEEE Transactions on Visualization and Computer Graphics, 21(9): 1087–1100, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07063249-abs.html.

Cha:2018:TFM

[CPW^{+}18] Young-Woon Cha, True Price, Zhen Wei, Xinran Lu, Nicholas Rewkowski, Rohan Chabra, Zihe Qin, Hyounghun Kim, Zhaoqi Su, Yebin Liu, Adrian Ilie, Andrei State, Zhenlin Xu, Jan-Michael Frahm, and Henry Fuchs. Towards fully mobile 3D face, body, and environment capture using only headworn cameras. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2993–3004, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8458443/.

Chan:2008:RAV

[CQC^{+}08] Ming-Yuen Chan, Huamin Qu, Ka-Kei Chung, Wai-Ho Mak, and Yingcai Wu. Relation-aware volume exploration pipeline. IEEE Transactions on Visualization and Computer Graphics, 14(6):1683–1690, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chan:2010:VSS

[CQM10] Wing-Yi Chan, Huamin Qu, and Wai-Ho Mak. Visualizing the semantic structure in classical music works. IEEE Transactions on Visualization and Computer Graphics, 16(1):161–173, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Caban:2008:PIV

[CR08] Jesus J. Caban and Penny Rheingans. Papers of the IEEE Visualization Conference 2008: Texture-based transfer functions for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 14(6):1364–1371, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Crossno:2005:VGS

[CRB^{+}05] Patricia Crossno, David H. Rogers, Rebecca M. Brannon, David Coblentz, and Joanne T. Fredrich. Visualization of geologic stress perturbations using Mohr diagrams. IEEE Transactions on Visualization and Computer Graphics, 11(5):508–518, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Collomosse:2005:SST

[CRH05] John P. Collomosse, David Rowntree, and Peter M. Hall.

Stroke surfaces: Temporally coherent artistic animations from video. IEEE Transactions on Visualization and Computer Graphics, 11(5): 540–549, September/October 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chittaro:2006:VFV

[CRI06] Luca Chittaro, Roberto Ranon, and Lucio Ieronutti. VU-Flow: a visualization tool for analyzing navigation in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 12(6):1475–1485, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2010:CRC

[CRPH10] Jian Cui, Paul Rosen, Voicu Popescu, and Christoph Hoffmann. A curved ray camera for handling occlusions through continuous multiperspective visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1235–1242, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Clarenz:2004:RFD

[CRT04] Ulrich Clarenz, Martin Rumpf, and Alexandru Telea. Robust feature detection and local classification for surfaces based on moment analysis. IEEE Transactions on Visualization and Computer Graphics, 10(5):516–524, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0516.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0516.pdf.

Coming:2008:VAD

[CS08] Daniel S. Coming and Oliver G. Staadt. Velocity-aligned discrete oriented polytopes for dynamic collision detection. IEEE Transactions on Visualization and Computer Graphics, 14(1):1–12, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Centin:2018:MDG

[CS18] Marco Centin and Alberto Signoroni. Mesh denoising with (Geo)Metric fidelity. IEEE Transactions on Visualization and Computer Graphics, 24(8):2380– 2396, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07990536-
abs.html.

Correa:2006:FAV

[CSC06] Carlos Correa, Deborah Silver, and Min Chen. Feature aligned volume manipulation for illustration and visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1069– 1076, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Correa:2007:IDD

[CSC07] Carlos Correa, Debora Silver, and Mi Chen. Illustrative deformation for data exploration. IEEE Transactions on Visualization and Computer Graphics, 13(6):1320–1327, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1320s.
wmv.

Csebfalvi:2008:EPR

[Csé08] Balázs Csébfalvi. An evaluation of prefiltered reconstruction schemes for volume rendering. IEEE Transactions on Visualization and Computer Graphics, 14(2):289– 301, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Csebfalvi:2010:EPB

[Cse10] Balazs Csebfalvi. An evaluation of prefiltered B-spline reconstruction for quasiinterpolation on the bodycentered cubic lattice. IEEE Transactions on Visualization and Computer Graphics, 16 (3):499–512, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Csebfalvi:2013:CWB

[Cse13] Balazs Csebfalvi. Cosineweighted B-spline interpolation: A fast and high-quality reconstruction scheme for the body-centered cubic lattice. IEEE Transactions on Visualization and Computer Graphics, 19(9):1455–1466, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chaitanya:2019:ASS

[CSG^{+}19] Chakravarty R. Alla Chaitanya, John M. Snyder, Keith Godin, Derek Nowrouzezahrai, and Nikunj Raghuvanshi. Adaptive sampling for sound propagation. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1846–1854, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642450/.

Cao:2010:FMV

[CSL^{+}10] Nan Cao, Jimeng Sun, Yu-Ru Lin, David Gotz, Shixia Liu, and Huamin Qu. FacetAtlas: Multifaceted visualization for rich text corpora. IEEE Transactions on Visualization and Computer Graphics, 16(6):1172–1181, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cao:2016:TVA

[CSL^{+}16] N. Cao, C. Shi, S. Lin, J. Lu, Y. Lin, and C. Lin. TargetVue: Visual analysis of anomalous user behaviors in online communication systems. IEEE Transactions on Visualization and Computer Graphics, 22(1):280–289, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cornea:2007:CSP

[CSM07] Nicu D. Cornea, Deborah Silver, and Patrick Min. Curveskeleton properties, applications, and algorithms. IEEE Transactions on Visualization and Computer Graphics, 13 (3):530–548, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cordier:2011:SMS

[CSPN11] Frederic Cordier, Hyewon Seo, Jinho Park, and Junyong Noh. Sketching of mirrorsymmetric shapes. IEEE Transactions on Visualization and Computer Graphics, 17(11):1650–1662, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cuenca:2018:MMS

[CSWP18] Erick Cuenca, Arnaud Sallaberry, Florence Y. Wang, and Pascal Poncelet. MultiStream: A multiresolution streamgraph approach to explore hierarchical time series. IEEE Transactions on Visualization and Computer Graphics, 24(12):3160– 3173, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8267086/.

Casas:2013:IAP

[CTGH13] Dan Casas, Margara Tejera, Jean-Yves Guillemaut, and Adrian Hilton. Interactive animation of 4D performance capture. IEEE Transactions on Visualization and Computer Graphics, 19(5): 762–773, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2013:PHI

[CTM^{+}13] Tao Chen, Ping Tan, Li-Qian Ma, Ming-Ming Cheng, Ariel Shamir, and Shi-Min Hu. PoseShop: Human image database construction and personalized content synthesis. IEEE Transactions on Visualization and Computer Graphics, 19(5):824–837, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2016:PHB

[CTT^{+}16] Ke-Li Cheng, Ruo-Feng Tong, Min Tang, Jing-Ye Qian, and Michel Sarkis. Parametric human body reconstruction based on sparse key points. IEEE Transactions on Visualization and Computer Graphics, 22(11):2467–2479, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chang:2010:ASM

[CVC10] Herng-Hua Chang, Daniel J.

Valentino, and Woei-Chyn Chu. Active shape modeling with electric flows. IEEE Transactions on Visualization and Computer Graphics, 16(5):854–869, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cirio:2012:WCN

[CVC^{+}12] Gabriel Cirio, Peter Vangorp, Emmanuelle Chapoulie, Maud Marchal, Anatole Lecuyer, and George Drettakis. Walking in a cube: Novel metaphors for safely navigating large virtual environments in restricted real workspaces. IEEE Transactions on Visualization and Computer Graphics, 18(4):

546–554, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chevalier:2013:UCS

[CVG13] Fanny Chevalier, Romain Vuillemot, and Guia Gali. Using concrete scales: A practical framework for effective visual depiction of complex measures. IEEE Transactions on Visualization and Computer Graphics, 19(12):2426– 2435, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Castermans:2019:SLD

[CVS^{+}19] T. Castermans, K. Verbeek, B. Speckmann, M. A. Westenberg, R. Koopman, S. Wang, H. van den Berg, and A. Betti. SolarView: Low distortion radial embedding with a focus. IEEE Transactions on Visualization and Computer Graphics, 25(10):2969–2982, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Claessen:2011:FLA

[CvW11] Jarry H. T. Claessen and Jarke J. van Wijk. Flexible linked axes for multivariate data visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2310–2316, December 2011. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cappers:2018:EME

[CvW18] B. C. M. Cappers and J. J.

van Wijk. Exploring multivariate event sequences using rules, aggregations, and selections. IEEE Transactions on Visualization and Computer Graphics, 24(1):532–541, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Comba:2011:GEI

[CW11] Joao Comba and Daniel Weiskopf. Guest Editor’s introduction: Special section on the Eurographics Symposium on Parallel Graphics and Visualization (EGPGV). IEEE Transactions on Visualization and Computer Graphics, 17 (2):130–131, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Clanton:2006:OMD

[CWC^{+}06] Samuel T. Clanton, David C.

Wang, Vikram S. Chib, Yoky Matsuoka, and George D. Stetten. Optical merger of direct vision with virtual images for scaled teleoperation. IEEE Transactions on Visualization and Computer Graphics, 12(2):277– 285, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/.

Chen:2009:EBV

[CWDH09] Yanhua Chen, Lijun Wang, Ming Dong, and Jing Hua. Exemplar-based visualization of large document corpus (InfoVis2009-1115). IEEE Transactions on Visualization and Computer Graphics, 15(6):1161–1168, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chang:2007:LCF

[CWK^{+}07] Remco Chang, Ginette Wessel, Robert Kosara, Eric Sauda, and William Ribarsky. Legible cities: Focusdependent multi-resolution visualization of urban relationships. IEEE Transactions on Visualization and Computer Graphics, 13(6):1169– 1175, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2007/
06/v1169s.mov.

Cashion:2012:DDS

[CWL12] Jeffrey Cashion, Chadwick Wingrave, and Joseph J. LaViola, Jr. Dense and dynamic 3D selection for gamebased virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18(4):634–642, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chan:2009:PBT

[CWM^{+}09a] Ming-Yuen Chan, Yingcai Wu, Wai-Ho Mak, Wei Chen, and Huamin Qu. Perceptionbased transparency optimization for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 15(6):1283– 1290, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chuang:2009:HPC

[CWM09b] Johnson Chuang, Daniel Weiskopf, and Torsten Moller. Hue-preserving color blending. IEEE Transactions on Visualization and Computer Graphics, 15(6):1275–1282, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2007:DAO

[CWQ^{+}07] Kin-Shing Cheng, Wenping Wang, Hong Qin, Kwan-Yee Wong, Huaiping Yang, and Yang Liu. Design and analysis of optimization methods for subdivision surface fitting. IEEE Transactions on Visualization and Computer Graphics, 13(5):878–890, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2006:MDA

[CWRY06] Qingguang Cui, Matthew Ward, Elke Rundensteiner, and Jing Yang. Measuring data abstraction quality in multiresolution visualizations. IEEE Transactions on Visualization and Computer Graphics, 12(5): 709–716, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chan:2008:VIV

[CWT^{+}08] Bryan Chan, Leslie Wu, Justin Talbot, Mike Cammarano, and Pat Hanrahan. Vispedia: Interactive visual exploration of Wikipedia data via search-based integration. IEEE Transactions on Visualization and Computer Graphics, 14(6):1213–1220, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cao:2014:FFE

[CWZ^{+}14] Chen Cao, Yanlin Weng, Shun Zhou, Yiying Tong, and Kun Zhou. FaceWarehouse: A 3D facial expression database for visual computing. IEEE Transactions on Visualization and Computer Graphics, 20 (3):413–425, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chan:2019:VVB

[CXDR19] Gromit Yeuk-Yin Chan, Panpan Xu, Zeng Dai, and Liu Ren. ViBr: Visualizing bipartite relations at scale with the minimum description length principle. IEEE Transactions on Visualization and Computer Graphics, 25(1): 321–330, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440048-
abs.html.

Cheng:2019:CND

[CXM19] Shenghui Cheng, Wei Xu, and Klaus Mueller. ColorMap ND: A data-driven approach and tool for mapping multivariate data to color. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1361–1377, February 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08302605-
abs.html.

Chen:2018:SSO

[CXR18] Y. Chen, P. Xu, and L. Ren.

Sequence synopsis: Optimize visual summary of temporal event data. IEEE Transactions on Visualization and Computer Graphics, 24(1):45– 55, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2008:IID

[CYB08] Jerry Chen, Ilmi Yoon, and Wes Bethel. Interactive, Internet delivery of visualization via structured prerendered multiresolution imagery. IEEE Transactions on Visualization and Computer Graphics, 14(2):302– 312, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2012:VBN

[CYC^{+}12] Zhonggui Chen, Zhan Yuan, Yi-King Choi, Ligang Liu, and Wenping Wang. Variational blue noise sampling. IEEE Transactions on Visualization and Computer Graphics, 18(10):1784–1796, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2016:IVD

[CYW^{+}16] S. Chen, X. Yuan, Z. Wang, C. Guo, J. Liang, Z. Wang, X. L. Zhang, and J. Zhang. Interactive visual discovering of movement patterns from sparsely sampled geo-tagged social media data. IEEE Transactions on Visualization and Computer Graphics, 22 (1):270–279, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2009:VIM

[CYZ^{+}09] Wei Chen, Zhicheng Yan, Song Zhang, John Allen Crow, David S. Ebert, Ronald M. McLaughlin, Katie B. Mullins, Robert Cooper, Zi’ang Ding, and Jun Liao. Volume illustration of muscle from diffusion tensor images. IEEE Transactions on Visualization and Computer Graphics, 15(6):1425–1432, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Caserta:2011:VSA

[CZ11] P. Caserta and O. Zendra. Visualization of the static aspects of software: A survey. IEEE Transactions on Visualization and Computer Graphics, 17(7):913–933, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2015:UAM

[CZC^{+}15] Haidong Chen, Song Zhang, Wei Chen, Honghui Mei, Jiawei Zhang, Andrew Mercer, Ronghua Liang, and Huamin Qu. Uncertainty-aware multidimensional ensemble data visualization and exploration. IEEE Transactions on Visualization and Computer Graphics, 21(9):1072–1086, September 2015. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        09/07055260-abs.html.

Czerwinski:2012:VKS

[Cze12] Mary Czerwinski. VisWeek 2012 keynote speaker. IEEE Transactions on Visualization and Computer Graphics, 18 (12):xxi, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2011:LHP

[CZN^{+}11] Cheng Chen, Yueting Zhuang, Feiping Nie, Yi Yang, Fei Wu, and Jun Xiao. Learning a 3D human pose distance metric from geometric pose descriptor. IEEE Transactions on Visualization and Computer Graphics, 17(11):1676– 1689, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2008:GBE

[CZQ^{+}08] Weiwei Cui, Hong Zhou, Huamin Qu, Pak Chung Wong, and Xiaoming Li. Geometry-based edge clustering for graph visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1277–1284, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chai:2017:ASI

[CZZ17a] Menglei Chai, Changxi Zheng, and Kun Zhou. Adaptive skinning for interactive hairsolid simulation. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1725–1738, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07448467-
abs.html.

Chen:2017:EBS

[CZZ17b] Xiang Chen, Changxi Zheng, and Kun Zhou. Examplebased subspace stress analysis for interactive shape design. IEEE Transactions on Visualization and Computer Graphics, 24(10):2314– 2327, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07593384-
abs.html.

Chen:2019:VVE

[CZZ^{+}19] S. Chen, L. Zheng, Y. Zhang, Z. Sun, and K. Xu. VERAM: View-enhanced recurrent attention model for 3D shape classification. IEEE Transactions on Visualization and Computer Graphics, 25(12): 3244–3257, December 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dachsbacher:2011:AVC

[Dac11] Carsten Dachsbacher. Analyzing visibility configurations. IEEE Transactions on Visualization and Computer Graphics, 17(4):475– 486, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dang:2016:RRA

[Dan16] Gang Dang. Realtime reconstruction of an animating human body from a single depth camera. IEEE Transactions on Visualization and Computer Graphics, 22(8):2000– 2011, August 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Daniels:2010:IVF

[DANS10] Joel Daniels II, Erik W. Anderson, Luis Gustavo Nonato, and Claudio T. Silva. Interactive vector field feature identification. IEEE Transactions on Visualization and Computer Graphics, 16(6):1560– 1568, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dang:2013:TSH

[DAW13] Tuan Nhon Dang, Anushka Anand, and Leland Wilkinson. TimeSeer: Scagnostics for high-dimensional time series. IEEE Transactions on Visualization and Computer Graphics, 19(3):470– 483, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dykes:2007:GWV

[DB07] Jason Dykes and Chris Brunsdon. Geographically weighted visualization: Interactive graphics for scale-varying exploratory analysis. IEEE Transactions on Visualization and Computer Graphics, 13(6):1161–1168, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Diehl:2010:USW

[DBB10] Stephan Diehl, Fabian Beck, and Michael Burch. Uncovering strengths and weaknesses of radial visualizations—an empirical approach. IEEE Transactions on Visualization and Computer Graphics, 16(6):935–942, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dimara:2019:MAE

[DBBF19] Evanthia Dimara, Gilles Bailly, Anastasia Bezerianos, and Steven Franconeri. Mitigating the attraction effect with visualizations. IEEE Transactions on Visualization and Computer Graphics, 25 (1):850–860, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440809-
                                        abs.html.

Dogrusoz:2013:CCS

[DBD13] Ugur Dogrusoz, Mehmet E. Belviranli, and Alptug Dilek. CiSE: A circular spring embedder layout algorithm. IEEE Transactions on Visualization and Computer Graphics, 19(6):953–966, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dimara:2017:AEI

[DBD17] Evanthia Dimara, Anastasia Bezerianos, and Pierre Dragicevic. The attraction effect in information visualization. IEEE Transactions on Visualization and Computer Graphics, 23(1):471–480, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dimara:2018:CMI

[DBD18] E. Dimara, A. Bezerianos, and P. Dragicevic. Conceptual and methodological issues in evaluating multidimensional visualizations for decision support. IEEE Transactions on Visualization and Computer Graphics, 24 (1):749–759, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Demiralp:2014:LPK

[DBH14] Cagatay Demiralp Demiralp, Michael S. Bernstein, and Jeffrey Heer. Learning perceptual kernels for visualization design. IEEE Transactions on Visualization and Computer Graphics, 20(12):1933– 1942, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875950-
abs.html.

Deines:2006:CVW

[DBM^{+}06] Eduard Deines, Martin Bertram, Jan Mohring, Jevgenij Jegorovs, Frank Michel, Hans Hagen, and Gregory M. Nielson. Comparative visualization for wave-based and geometric acoustics. IEEE Transactions on Visualization and Computer Graphics, 12(5):1173–1180, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dorling:2006:WWY

[DBN06] Danny Dorling, Anna Barford, and Mark Newman. Worldmapper: The world as you’ve never seen it before. IEEE Transactions on Visualization and Computer Graphics, 12(5):757–764, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dumont:2014:AAV

[DBP14] Georges Dumont, Marwan Badawi, and Charles Pontonnier. Assessing the ability of a vr-based assembly task simulation to evaluate physical risk factors. IEEE Transactions on Visualization and Computer Graphics, 20 (5):664–674, May 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dillard:2007:CSB

[DBTH07] Scott Dillard, John Bingert, Dan Thoma, and Bernd Hamann. Construction of simplified boundary surfaces from serial-sectioned metal micrographs. IEEE Transactions on Visualization and Computer Graphics, 13(6):1528– 1535, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dick:2011:DVI

[DBW11] Christian Dick, Rainer Burgkart, and Rudiger Westermann. Distance visualization for interactive 3D implant planning. IEEE Transactions on Visualization and Computer Graphics, 17(12):2173– 2182, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dabek:2017:GBA

[DC17] Filip Dabek and Jesus J. Caban. A grammar-based approach for modeling user interactions and generating suggestions during the data exploration process. IEEE Transactions on Visualization and Computer Graphics, 23 (1):41–50, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dork:2008:VCV

[DCCW08] Marian Dörk, Sheelagh Carpendale, Christopher Collins, and Carey Williamson. VisGets: Coordinated visualizations for Web-based information exploration and discovery. IEEE Transactions on Visualization and Computer Graphics, 14(6):1205–1212, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dutta:2017:SDG

[DCH^{+}17] Soumya Dutta, Chun-Ming Chen, Gregory Heinlein, Han-Wei Shen, and Jen-Ping Chen. In situ distribution guided analysis and visualization of transonic jet engine simulations. IEEE Transactions on Visualization and Computer Graphics, 23(1):811–820, January 2017. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Duke:2012:VNS

[DCK^{+}12] David Duke, Hamish Carr, Aaron Knoll, Nicolas Schunck, Hai Ah Nam, and Andrzej Staszczak. Visualizing nuclear scission through a multifield extension of topological analysis. IEEE Transactions on Visualization and Computer Graphics, 18(12):2033– 2040, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dong:2002:ABA

[DCKY02] F. Dong, G. J. Clapworthy, M. A. Krokos, and J. Yao. An anatomy-based approach to human muscle modeling and deformation. IEEE Transactions on Visualization and Computer Graphics, 8(2): 154–170, April 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0154.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0154abs.
                                        htm.

Duffy:2013:IIS

[DCM13] B. Duffy, H. Carr, and T. Moller. Integrating isosurface statistics and histograms. IEEE Transactions on Visualization and Computer Graphics, 19(2):263–277, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doukakis:2019:AVO

[DDBR^{+}19] E. Doukakis, K. Debattista, T. Bashford-Rogers, A. Dhokia, A. Asadipour, A. Chalmers, and C. Harvey. Audio-visual-olfactory resource allocation for tri-modal virtual environments. IEEE Transactions on Visualization and Computer Graphics, 25(5):1865–1875, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642346/.

DiGiacomo:2007:GVT

[DDGL07] Emilio Di Giacomo, Walter Didimo, Luca Grilli, and Giuseppe Liotta. Graph visualization techniques for Web clustering engines. IEEE Transactions on Visualization and Computer Graphics, 13(2):294–304, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Duriez:2006:RHR

[DDKA06] Christian Duriez, Frédéric Dubois, Abderrahmane Kheddar, and Claude Andriot. Realistic haptic rendering of interacting deformable objects in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 12 (1):36–47, January/February 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dionne:2014:GBD

[DdL14] Olivier Dionne and Martin de Lasa. Geodesic binding for degenerate character geometry using sparse voxelization. IEEE Transactions on Visualization and Computer Graphics, 20(10):1367– 1378, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06809992-
abs.html.

Demir:2014:MCC

[DDW14] Ismail Demir, Christian Dick, and Rudiger Westermann. Multi-charts for comparative 3D ensemble visualization. IEEE Transactions on Visualization and Computer Graphics, 20(12):2694– 2703, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875990-
abs.html.

DeFloriani:2015:ENb

[De 15a] L. De Floriani. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 21(5):553–554, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2015:ENa

[De 15b] Leila De Floriani. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 21(2):172– 173, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/07000011.
pdf.

DeFloriani:2015:ENc

[De 15c]  Leila De Floriani.  Edi-

tor’s note. IEEE Transactions on Visualization and Computer Graphics, 21(9): 994–995, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07169659.pdf.

DeFloriani:2015:MNE

[De 15d] Leila De Floriani. A message from the new Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 21(1):3, January 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06966880.
pdf.

Floriani:2016:MEC

[De 16a] L. De Floriani. A message from the Editor-in-Chief.

IEEE Transactions on Visualization and Computer Graphics, 22(2):1035, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2016:EN

[De 16b]  Leila De Floriani.  Edi-

tor’s note. IEEE Transactions on Visualization and Computer Graphics, 22(12): 2507, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07707445.
pdf.

Floriani:2017:EN

[De 17a] Leila De Floriani. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 23(3):1163– 1164, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07835796.
pdf.

DeFloriani:2017:EN

[De 17b] Leila De Floriani. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 23(12):2494– 2495, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2017/12/08089787.
pdf.

DeFloriani:2017:SJ

[De 17c] Leila De Floriani. State of the journal. IEEE Transactions on Visualization and Computer Graphics, 23(2):1000– 1001, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07801956.
pdf.

DeFloriani:2018:MECa

[De 18a] L. De Floriani. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 24 (1):x, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2018:TBA

[De 18b] Leila De Floriani. 2017 TVCG Best Associate Editor Award and Best Reviewer Award. IEEE Transactions on Visualization and Computer Graphics, 24(8):2269, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/08400453.
pdf.

DeFloriani:2018:EN

[De 18c] Leila De Floriani. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 24(9):2470– 2472, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08423388.
pdf.

DeFloriani:2019:FNE

[De 19a] L. De Floriani. Farewell and new EIC introduction. IEEE Transactions on Visualization and Computer Graphics, 25 (3):1447–1448, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2019:MEC

[De 19b]  Leila De Floriani.  Mes-

sage from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 25(1):xi, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08570933.
pdf.

Dwyer:2018:P

[DEF^{+}18] T. Dwyer, N. Elmqvist, B. Fisher, S. Franconeri, I. Hotz, R. M. M. Kirby, S. Liu, T. Schreck, and X. Yuan. Preface. IEEE Transactions on Visualization and Computer Graphics, 24 (1):xi–xv, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Delingette:2008:TSM

[Del08] Hervé Delingette. Triangular springs for modeling nonlinear membranes. IEEE Transactions on Visualization and Computer Graphics, 14(2): 329–341, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Drettakis:1996:SPI

[DF96] George Drettakis and Eugene L. Fiume. Structured penumbral irradiance computation. IEEE Transactions on Visualization and Computer Graphics, 2(4):299– 312, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0299.
pdf;   http://www.computer.
org/tvcg/tg1996/v0299abs.
htm.

Doraiswamy:2014:UTA

[DFD^{+}14] Harish Doraiswamy, Nivan Ferreira, Theodoros Damoulas, Juliana Freire, and Claudio T. Silva. Using topological analysis to support event-guided exploration in urban data. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2634–2643, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876004-
abs.html.

Debernardis:2014:TRH

[DFG^{+}14] Saverio Debernardis, Michele Fiorentino, Michele Gattullo, Giuseppe Monno, and Antonio Emmanuele Uva. Text readability in head-worn displays: Color and style optimization in video versus optical see-through devices. IEEE Transactions on Visualization and Computer Graphics, 20 (1):125–139, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DiBattista:2012:GEI

[DFQ12] Giuseppe Di Battista, Jean-Daniel Fekete, and Huamin Qu. Guest Editor’s introduction: Special Section on the IEEE Pacific Visualization Symposium. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1381–1382, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Durupinar:2016:PPC

[DGAB16] Funda Durupınar, Uğur Güdükbay, Aytek Aman, and Norman I. Badler. Psychological parameters for crowd simulation: From audiences to mobs. IEEE Transactions on Visualization and Computer Graphics, 22(9): 2145–2159, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dick:2009:STF

[DGBW09] Christian Dick, Joachim Georgii, Rainer Burgkart, and Rudiger Westermann. Stress tensor field visualization for implant planning in orthopedics. IEEE Transactions on Visualization and Computer Graphics, 15(6):1399– 1406, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dick:2011:HMA

[DGW11] Christian Dick, Joachim Georgii, and Rudiger Westermann. A hexahedral multigrid approach for simulating cuts in deformable objects. IEEE Transactions on Visualization and Computer Graphics, 17(11):1663–1675, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dork:2010:VBL

[DGWC10] Marian Dork, Daniel Gruen, Carey Williamson, and Sheelagh Carpendale. A visual backchannel for large-scale events. IEEE Transactions on Visualization and Computer Graphics, 16(6):1129– 1138, November/December 2010. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dollner:2002:GRS

[DH02] J. Döllner and K. Hinrichs. A generic rendering system. IEEE Transactions on Visualization and Computer Graphics, 8(2):99–118, April 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0099.
pdf;   http://www.computer.
org/tvcg/tg2002/v0099abs.
htm.

DiVerdi:2008:HCV

[DH08] Stephen DiVerdi and Tobias Höllerer. Heads up and camera down: a vision-based tracking modality for mobile mixed reality. IEEE Transactions on Visualization and Computer Graphics, 14 (3):500–512, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Demiralp:2009:CLF

[DHL09]  Çağatay Demiralp, John F.

Hughes, and David H. Laidlaw. Coloring 3D line fields using Boy’s real projective plane immersion. IEEE Transactions on Visualization and Computer Graphics, 15(6):1457–1464, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Du:2013:CPS

[DHM13a] Song-Pei Du, Shi-Min Hu, and Ralph R. Martin. Changing perspective in stereoscopic images. IEEE Transactions on Visualization and Computer Graphics, 19(8):1288– 1297, August 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Du:2013:SST

[DHM13b] Song-Pei Du, Shi-Min Hu, and Ralph R. Martin. Semiregular solid texturing from 2D image exemplars. IEEE Transactions on Visualization and Computer Graphics, 19 (3):460–469, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Duran:2019:VLM

[DHR^{+}19] David Duran, Pedro Hermosilla, Timo Ropinski, Barbora Kozlikova, Alvar Vinacua, and Pere-Pau Vazquez. Visualization of large molecular trajectories. IEEE Transactions on Visualization and Computer Graphics, 25(1): 987–996, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08456856-
abs.html.

Dick:2014:LSL

[Dic14]  Christian Dick.  Large-

scale liquid simulation on adaptive hexahedral grids. IEEE Transactions on Visualization and Computer Graphics, 20(10):1405–1417, October 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/10/06747389-
                                        abs.html.

Dill:2017:VCA

[Dil17] John Dill. The 2016 Visualization Career Award. IEEE Transactions on Visualization and Computer Graphics, 23 (1):xxiv, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DiVerdi:2015:MFD

[DiV15] Stephen DiVerdi. A modular framework for digital painting. IEEE Transactions on Visualization and Computer Graphics, 21(7):783–793, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        07/07042343-abs.html.

Dragicevic:2018:BSI

[DJ18] P. Dragicevic and Y. Jansen. Blinded with science or informed by charts? A replication study. IEEE Transactions on Visualization and Computer Graphics, 24(1): 781–790, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Demiralp:2006:CFV

[DJK^{+}06] Cagatay Demiralp, Cullen D.

Jackson, David B. Karelitz, Song Zhang, and David H. Laidlaw. CAVE and fishtank virtual-reality displays: a qualitative and quantitative comparison. IEEE Transactions on Visualization and Computer Graphics, 12(3): 323–330, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Oliveira:2017:DVH

[dJOBNM17] Victor Adriel de Jesus Oliveira, Luca Brayda, Luciana Nedel, and Anderson Maciel. Designing a vibrotactile headmounted display for spatial awareness in 3D spaces. IEEE Transactions on Visualization and Computer Graphics, 23(4):1409–1417, April 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829406-
abs.html.

Dasgupta:2010:PSS

[DK10] Aritra Dasgupta and Robert Kosara. Pargnostics: Screenspace metrics for parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 16(6):1017– 1026, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dasgupta:2011:APP

[DK11a] Aritra Dasgupta and Robert Kosara. Adaptive privacypreserving visualization using parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 17(12):2241–2248, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doerr:2011:CSH

[DK11b] Kai-Uwe Doerr and Falko Kuester. CGLX: a scalable, high-performance visualization framework for networked display environments. IEEE Transactions on Visualization and Computer Graphics, 17 (3):320–332, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Debelov:2013:LML

[DK13] Victor A. Debelov and Dmitry S. Kozlov. A local model of light interaction with transparent crystalline media. IEEE Transactions on Visualization and Computer Graphics, 19(8):1274–1287, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dwyer:2006:DDG

[DKM06a] Tim Dwyer, Yehuda Koren, and Kim Marriott. Drawing directed graphs using quadratic programming. IEEE Transactions on Visualization and Computer Graphics, 12(4):536– 548, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dwyer:2006:ICI

[DKM06b] Tim Dwyer, Yehuda Koren, and Kim Marriott. IPSep-CoLa: an incremental procedure for separation constraint layout of graphs. IEEE Transactions on Visualization and Computer Graphics, 12(5): 821–828, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DiVerdi:2013:PPP

[DKMI13] Stephen DiVerdi, Aravind Krishnaswamy, Radomir Mach, and Daichi Ito. Painting with polygons: A procedural watercolor engine. IEEE Transactions on Visualization and Computer Graphics, 19 (5):723–735, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Drori:2003:FMI

[DL03] Iddo Drori and Dani Lischinski. Fast multiresolution image operations in the wavelet domain. IEEE Transactions on Visualization and Computer Graphics, 9(3):395–411, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0395abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0395.pdf.

Deng:2012:DAD

[DL12] Yunhua Deng and Rynson W. H. Lau. On delay adjustment for dynamic load balancing in distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18 (4):529–537, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Douma:2009:SRL

[DLA^{+}09] Michael Douma, Grzegorz Ligierko, Ovidiu Ancuta, Pavel Gritsai, and Sean Liu. SpicyNodes: Radial layout authoring for the general public. IEEE Transactions on Visualization and Computer Graphics, 15(6):1089–1096, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dwyer:2009:CUG

[DLF^{+}09] Tim Dwyer, Bongshin Lee, Danyel Fisher, Kori Inkpen Quinn, Petra Isenberg, George Robertson, and Chris North. A comparison of user-generated and automatic graph layouts. IEEE Transactions on Visualization and Computer Graphics, 15(6):961–968, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dykes:2012:MPC

[DLM^{+}12] Jason Dykes, David Laidlaw, Klaus Mueller, Giuseppe Santucci, Gerik Scheuermann, Matthew Ward, and Chris Weaver. Message from the Paper Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 18(12):x–xii, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Draper:2009:SRM

[DLR09] Geoffrey M. Draper, Yarden Livnat, and Richard F. Riesenfeld. A survey of radial methods for information visualization. IEEE Transactions on Visualization and Computer Graphics, 15(5): 759–776, September/October 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

deLeeuw:2006:VAL

[dLVvL06] Wim de Leeuw, Pernette Verschure, and Robert van Liere. Visualization and analysis of large data collections: a case study applied to confocal microscopy data. IEEE Transactions on Visualization and Computer Graphics, 12(5):1251–1258, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dasgupta:2017:FVT

[DLW^{+}17] Aritra Dasgupta, Joon-Yong Lee, Ryan Wilson, Robert A. Lafrance, Nick Cramer, Kristin Cook, and Samuel Payne. Familiarity vs trust: A comparative study of domain scientists’ trust in visual analytics and conventional analysis methods. IEEE Transactions on Visualization and Computer Graphics, 23(1): 271–280, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Denes:2019:TRM

[DMAM19] Gyorgy Denes, Kuba Maruszczyk, George Ash, and Rafał K. Mantiuk. Temporal resolution multiplexing: Exploiting the limitations of spatiotemporal vision for more efficient VR rendering. IEEE Transactions on Visualization and Computer Graphics, 25(5):2072–2082, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643566/.

Dellepiane:2012:FBL

[DMC^{+}12] Matteo Dellepiane, Ricardo Marroquim, Marco Callieri, Paolo Cignoni, and Roberto Scopigno. Flow-based local optimization for image-to-geometry projection. IEEE Transactions on Visualization and Computer Graphics, 18 (3):463–474, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dumas:2015:VAS

[DMC15] Maxime Dumas, Michael J.

McGuffin, and Patrick Chasse. VectorLens: Angular selection of curves within 2D dense visualizations. IEEE Transactions on Visualization and Computer Graphics, 21(3): 402–412, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06919281-
abs.html.

Deschamps:2004:FEI

[DMR04] Thomas Deschamps, Ravi Malladi, and Igor Ravve. Fast evolution of image manifolds and application to filtering and segmentation in 3D medical images. IEEE Transactions on Visualization and Computer Graphics, 10(5):525–535, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        dl/trans/tg/2004/05/v0525.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/05/
                                        v0525.pdf.

Dwyer:2008:ENU

[DMS^{+}08] Tim Dwyer, Kim Marriott, Falk Schreiber, Peter Stuckey, Michael Woodward, and Michael Wybrow. Exploration of networks using overview+detail with constraint-based cooperative layout. IEEE Transactions on Visualization and Computer Graphics, 14(6):1293– 1300, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doraiswamy:2012:OSC

[DN12] Harish Doraiswamy and Vijay Natarajan. Output-sensitive construction of Reeb graphs. IEEE Transactions on Visualization and Computer Graphics, 18(1):146–159, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doraiswamy:2013:CRG

[DN13] H. Doraiswamy and V. Natarajan. Computing Reeb graphs as a union of contour trees. IEEE Transactions on Visualization and Computer Graphics, 19(2):249–262, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Deng:2006:EFA

[DNL^{+}06] Zhigang Deng, Ulrich Neumann, J. P. Lewis, Tae-Yong Kim, Murtaza Bulut, and Shrikanth Narayanan. Expressive facial animation synthesis by learning speech coarticulation and expression spaces. IEEE Transactions on Visualization and Computer Graphics, 12(6):1523– 1534, November/December 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doraiswamy:2013:EFI

[DNN13] Harish Doraiswamy, Vijay Natarajan, and Ravi S. Nanjundiah. An exploration framework to identify and track movement of cloud systems. IEEE Transactions on Visualization and Computer Graphics, 19(12):2896– 2905, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doncel:2007:ODS

[DNP07] Victor Rodriguez Doncel, Nikos Nikolaidis, and Ioannis Pitas. An optimal detector structure for the Fourier descriptors domain watermarking of 2D vector graphics. IEEE Transactions on Visualization and Computer Graphics, 13(5):851–863, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

deOliveira:2003:VDE

[dOL03] Maria Cristina Ferreira de Oliveira and Haim Levkowitz. From visual data exploration to visual data mining: a survey. IEEE Transactions on Visualization and Computer Graphics, 9(3):378–394, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0378abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0378.pdf.

Diewald:2000:ADV

[DPR00] U. Diewald, T. Preußer, and M. Rumpf. Anisotropic diffusion in vector field visualization on Euclidean domains and surfaces. IEEE Transactions on Visualization and Computer Graphics, 6(2):139– 149, April/June 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0139.
pdf;   http://www.computer.
org/tvcg/tg2000/v0139abs.
htm.

Dasgupta:2015:BTP

[DPW^{+}15] Aritra Dasgupta, Jorge Poco, Yaxing Wei, Robert Cook, Enrico Bertini, and Claudio T. Silva. Bridging theory with practice: An exploratory study of visualization use and design for climate model comparison. IEEE Transactions on Visualization and Computer Graphics, 21(9):996–1014, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07061479-abs.html.

Du:2007:FFG

[DQ07] Haixia Du and Hong Qin.

Free-form geometric modeling by integrating parametric and implicit PDEs. IEEE Transactions on Visualization and Computer Graphics, 13 (3):549–561, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Draper:2008:WVW

[DR08] Geoffrey Draper and Richard Riesenfeld. Who votes for what? A visual query language for opinion data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1197–1204, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

deRousiers:2012:RTR

[dRBS^{+}12] Charles de Rousiers, Adrien Bousseau, Kartic Subr, Nicolas Holzschuch, and Ravi Ramamoorthi. Real-time rendering of rough refraction. IEEE Transactions on Visualization and Computer Graphics, 18(10):1591–1602, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Doerr:2007:ELC

[DRHK07] Kai-Uwe Doerr, Holger Rademacher, Silke Huesgen, and Wolfgang Kubbat. Evaluation of a lowcost 3D sound system for immersive virtual reality training systems. IEEE Transactions on Visualization and Computer Graphics, 13(2): 204–212, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dwyer:2013:ECT

[DRMM13] Tim Dwyer, Nathalie Henry Riche, Kim Marriott, and Christopher Mears. Edge compression techniques for visualization of dense directed graphs. IEEE Transactions on Visualization and Computer Graphics, 19(12):2596– 2605, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dork:2012:PST

[DRRD12] Marian Dork, Nathalie Henry Riche, Gonzalo Ramos, and Susan Dumais. PivotPaths: Strolling through faceted information spaces. IEEE Transactions on Visualization and Computer Graphics, 18(12):2709–2718, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Drumwright:2008:FSP

[Dru08] Evan Drumwright. A fast and stable penalty method for rigid body simulation. IEEE Transactions on Visualization and Computer Graphics, 14(1):231–240, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dick:2016:SFP

[DRW16] Christian Dick, Marcus Rogowsky, and Rüdiger Westermann. Solving the fluid pressure Poisson equation using multigrid — evaluation and improvements. IEEE Transactions on Visualization and Computer Graphics, 22(11): 2480–2492, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2016:MEC

[DS16a] Leila De Floriani and Dieter Schmalstieg. Message from the Editor-in-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 22(11):i, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dutta:2016:DDE

[DS16b] S. Dutta and H. Shen. Distribution driven extraction and tracking of features for timevarying data analysis. IEEE Transactions on Visualization and Computer Graphics, 22 (1):837–846, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DeFloriani:2017:IIV

[DS17a] Leila De Floriani and Dieter Schmalstieg. Introducing the IEEE Virtual Reality 2017 special issue. IEEE Transactions on Visualization and Computer Graphics, 23(4):v, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07878159-
abs.html.

DeFloriani:2017:MEC

[DS17b] Leila De Floriani and Dieter Schmalstieg. Message from the Editor-in-Chief and from the Associate Editorin-Chief. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2365, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08053887-
abs.html.

DeFloriani:2018:IIV

[DS18a] Leila De Floriani and Dieter Schmalstieg. Introducing the IEEE Virtual Reality 2018 special issue. IEEE Transactions on Visualization and Computer Graphics, 24(4):v, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08315163-
abs.html.

DeFloriani:2018:MECb

[DS18b] Leila De Floriani and Dieter Schmalstieg. Message from the Editor-in-Chief and from the Associate Editorin-Chief. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2843, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8514109/.

Dietrich:2008:EGA

[DSC^{+}08] Carlos Dietrich, Carlos Scheidegger, João Comba, Luciana Nedel, and Cláudio Silva. Edge groups: an approach to understanding the mesh quality of marching methods. IEEE Transactions on Visualization and Computer Graphics, 14(6):1651–1666, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lorenzo:2016:AVE

[DSC^{+}16] G. Di Lorenzo, M. Sbodio, F. Calabrese, M. Berlingerio, F. Pinelli, and R. Nair. AllAboard: Visual exploration of cellphone mobility data to optimise public transport. IEEE Transactions on Visualization and Computer Graphics, 22(2):1036–1050, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Duarte:2014:NNN

[DSF^{+}14] Felipe S. L. G. Duarte, Fabio Sikansi, Francisco M. Fatore, Samuel G. Fadel, and Fernando V. Paulovich. Nmap: A novel neighborhood preservation space-filling algorithm. IEEE Transactions on Visualization and Computer Graphics, 20(12):2063– 2071, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876012-
                                        abs.html.

Dinkla:2017:SVA

[DSG^{+}17] Kasper Dinkla, Hendrik Strobelt, Bryan Genest, Stephan Reiling, Mark Borowsky, and Hanspeter Pfister. Screenit: Visual analysis of cellular screens. IEEE Transactions on Visualization and Computer Graphics, 23(1):591– 600, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dolonius:2019:CCD

[DSKA19] Dan Dolonius, Erik Sintorn, Viktor Kampe, and Ulf Assarsson. Compressing color data for voxelized surface geometry. IEEE Transactions on Visualization and Computer Graphics, 25(2):1270– 1282, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08013133-
abs.html.

Du:2017:CVV

[DSP^{+}17] Fan Du, Ben Shneiderman, Catherine Plaisant, Sana Malik, and Adam Perer. Coping with volume and variety in temporal event sequences: Strategies for sharpening analytic focus. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1636–1649, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/06/07429778-
                                        abs.html.

Dietrich:2009:ETI

[DSS^{+}09] Carlos A. Dietrich, Carlos E. Scheidegger, John Schreiner, João L. D. Comba, Luciana P. Nedel, and Cláudio T. Silva. Edge transformations for improving mesh quality of marching cubes. IEEE Transactions on Visualization and Computer Graphics, 15(1): 150–159, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Degener:2008:EVS

[DSSK08] Patrick Degener, Ruwen Schnabel, Christopher Schwartz, and Reinhard Klein. Effective visualization of short routes. IEEE Transactions on Visualization and Computer Graphics, 14(6):1452–1458, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dominitz:2010:TMO

[DT10] Ayelet Dominitz and Allen Tannenbaum. Texture mapping via optimal mass transport. IEEE Transactions on Visualization and Computer Graphics, 16(3):419– 433, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dunn:2017:WFV

[DTT^{+}17] David Dunn, Cary Tippets, Kent Torell, Petr Kellnhofer, Kaan Aksit, Piotr Didyk, Karol Myszkowski, David Luebke, and Henry Fuchs. Wide field of view varifocal near-eye display using see-through deformable membrane mirrors. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1322–1331, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829412-
abs.html.

Duffy:2015:GBV

[DTW^{+}15] Brian Duffy, Jeyarajan Thiyagalingam, Simon Walton, David J. Smith, Anne Trefethen, Jackson C. Kirkman-Brown, Eamonn A. Gaffney, and Min Chen. Glyphbased video visualization for semen analysis. IEEE Transactions on Visualization and Computer Graphics, 21(8):980–993, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/06684146-abs.html.

DelBimbo:1995:SEV

[DV95] Alberto Del Bimbo and Enrico Vicario. Specification by-example of virtual agents behavior. IEEE Transactions on Visualization and Computer Graphics, 1(4):350– 360, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0350.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0350abs.
                                        htm.

Digne:2018:SGR

[DVC18] Julie Digne, Sebastien Valette, and Raphaelle Chaine. Sparse geometric representation through local shape probing. IEEE Transactions on Visualization and Computer Graphics, 24(7):2238–2250, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07956272-
                                        abs.html.

Damera-Venkata:2007:UPS

[DVCD07] Niranjan Damera-Venkata, Nelson Chang, and Jeffrey Dicarlo. A unified paradigm for scalable multi-projector displays. IEEE Transactions on Visualization and Computer Graphics, 13(6):1360– 1367, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

DalCol:2018:WBV

[DVP^{+}18] Alcebiades Dal Col, Paola Valdivia, Fabiano Petronetto, Fabio Dias, Claudio T. Silva, and L. Gustavo Nonato. Wavelet-based visual analysis of dynamic networks. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2456–2469, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/08017630-
abs.html.

Dingen:2019:RIE

[DvVH^{+}19] Dennis Dingen, Marcel van’t Veer, Patrick Houthuizen, Eveline H. J. Mestrom, Erik H. H. M. Korsten, Arthur R. A. Bouwman, and Jarke van Wijk. RegressionExplorer: Interactive exploration of logistic regression models with subgroup analysis. IEEE Transactions on Visualization and Computer Graphics, 25(1):246– 255, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08464305-
abs.html.

Dang:2014:TSR

[DW14] Tuan Nhon Dang and Leland Wilkinson. Transforming scagnostics to reveal hidden features. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1624–1632, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875999-
                                        abs.html.

Dwyer:2017:GEI

[DW17] Tim Dwyer and Yingcai Wu. Guest Editors’ introduction: Special section on IEEE PacificVis 2017. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1574–1575, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/06/07914611.
                                        html.

Dang:2010:SGE

[DWA10] Tuan Nhon Dang, Leland Wilkinson, and Anushka Anand. Stacking graphic elements to avoid over-plotting. IEEE Transactions on Visualization and Computer Graphics, 16(6):1044–1052, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Donikian:2006:ADI

[DWB^{+}06] Michael Donikian, Bruce Walter, Kavita Bala, Sebastian Fernandez, and Donald P. Greenberg. Accurate direct illumination using iterative adaptive sampling. IEEE Transactions on Visualization and Computer Graphics, 12 (3):353–364, May/June 2006.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Duke:2006:FGV

[DWBR06] David Duke, Malcolm Wallace, Rita Borgo, and Colin Runciman. Fine-grained visualization pipelines and lazy functional languages. IEEE Transactions on Visualization and Computer Graphics, 12(5):973–980, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dowling:2019:SDS

[DWF^{+}19] Michelle Dowling, John Wenskovitch, J. T. Fry, Scotland Leman, Leanna House, and Chris North. SIRIUS: Dual, symmetric, interactive dimension reductions. IEEE Transactions on Visualization and Computer Graphics, 25 (1):172–182, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440814-
abs.html.

Dong:2016:IRT

[DWK^{+}16] Weiming Dong, Fuzhang Wu, Yan Kong, Xing Mei, Tong-Yee Lee, and Xiaopeng Zhang. Image retargeting by textureaware synthesis. IEEE Transactions on Visualization and Computer Graphics, 22(2):

1088–1101, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dykes:2010:RML

[DWS10] Jason Dykes, Jo Wood, and Aidan Slingsby. Rethinking map legends with visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):890– 899, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dinkla:2012:CAM

[DWvW12] Kasper Dinkla, Michel A. Westenberg, and Jarke J. van Wijk. Compressed adjacency matrices: Untangling gene regulatory networks. IEEE Transactions on Visualization and Computer Graphics, 18(12):2457–2466, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dou:2013:HVE

[DYW^{+}13] Wenwen Dou, Li Yu, Xiaoyu Wang, Zhiqiang Ma, and William Ribarsky. HierarchicalTopics: Visually exploring large text collections using topic hierarchies. IEEE Transactions on Visualization and Computer Graphics, 19(12):2002–2011, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dong:2014:SBI

[DZL^{+}14] Weiming Dong, Ning Zhou, Tong-Yee Lee, Fuzhang Wu, Yan Kong, and Xiaopeng Zhang. Summarization-based image resizing by intelligent object carving. IEEE Transactions on Visualization and Computer Graphics, 20 (1):1, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Deng:2016:IUC

[DZMQ16] Hao Deng, Liqiang Zhang, Xiancheng Mao, and Huamin Qu. Interactive urban context-aware visualization via multiple disocclusion operators. IEEE Transactions on Visualization and Computer Graphics, 22(7):1862– 1874, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Emgin:2019:HIT

[EASB19] S. E. Emgin, A. Aghakhani, T. M. Sezgin, and C. Basdogan. HapTable: An interactive tabletop providing online haptic feedback for touch gestures. IEEE Transactions on Visualization and Computer Graphics, 25(9):2749– 2762, September 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

El-Assady:2019:VAT

[EASD^{+}19] Mennatallah El-Assady, Fabian Sperrle, Oliver Deussen, Daniel Keim, and Christopher Collins. Visual analytics for topic model optimization based on user-steerable speculative execution. IEEE Transactions on Visualization and Computer Graphics, 25 (1):374–384, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08467535-
abs.html.

El-Assady:2018:PLT

[EASS^{+}18] M. El-Assady, R. Sevastjanova, F. Sperrle, D. Keim, and C. Collins. Progressive learning of topic modeling parameters: A visual analytics framework. IEEE Transactions on Visualization and Computer Graphics, 24(1): 382–391, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Everts:2015:EBW

[EBB^{+}15] Maarten H. Everts, Eric Begue, Henk Bekker, Jos B. T. M. Roerdink, and Tobias Isenberg. Exploration of the brain’s white matter structure through visual abstraction and multiscale local fiber tract contraction. IEEE Transactions on Visualization and Computer Graphics, 21(7):808–821, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
07/07042344-abs.html.

Ebert:2000:GEI

[Ebe00] D. S. Ebert. Guest Editor’s introduction. IEEE Transactions on Visualization and Computer Graphics, 6(2): 97, April/June 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0097.
pdf.

Ebert:2003:ENa

[Ebe03a]  David Ebert.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 9(1):2, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0002.
pdf;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0002.htm.

Ebert:2003:ENb

[Ebe03b] David S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 9(2):188– 190, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/02/v0188.
                                        pdf;   http://csdl.computer.
                                        org/dl/trans/tg/2003/02/
                                        v0188.htm.

Ebert:2004:ENa

[Ebe04a] David S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 10(1):1, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/01/v0001.
                                        pdf.

Ebert:2004:ENb

[Ebe04b] David S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 10(5):545–547, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/05/v0545.
                                        pdf;   http://csdl.computer.
                                        org/dl/trans/tg/2004/05/
                                        v0545.htm.

Ebert:2006:ENa

[Ebe06a] David S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 12(1):1, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/01/v0001.
pdf.

Ebert:2006:ENb

[Ebe06b] David S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 12(4):417– 418, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/04/v0417.
pdf.

Ebert:2007:EEF

[Ebe07] David S. Ebert. Editorial:

EIC farewell and new EIC introduction. IEEE Transactions on Visualization and Computer Graphics, 13(1):1, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/01/v0001.
pdf.

Ebert:2017:VTA

[Ebe17] David Ebert. The 2016 Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 23(1): xxv, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Everts:2009:DDH

[EBRI09] Maarten H. Everts, Henk Bekker, Jos B. T. M. Roerdink, and Tobias Isenberg. Depth-dependent halos: Illustrative rendering of dense line data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1299– 1306, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ellis:2006:EAC

[ED06] Geoffrey Ellis and Alan Dix. Enabling automatic clutter reduction in parallel coordinate plots. IEEE Transactions on Visualization and Computer Graphics, 12(5): 717–724, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ellis:2007:TCR

[ED07] Geoffrey Ellis and Alan Dix. A taxonomy of clutter reduction for information visualisation. IEEE Transactions on Visualization and Computer Graphics, 13(6):1216– 1223, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elmqvist:2008:PII

[EDF08] Niklas Elmqvist, Pierre Dragicevic, and Jean-Daniel Fekete. Papers of the IEEE Information Visualization Conference 2008: Rolling the dice: Multidimensional visual exploration using scatterplot matrix navigation. IEEE Transactions on Visualization and Computer Graphics, 14(6):1141–1148, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elmqvist:2011:CLA

[EDF11] Niklas Elmqvist, Pierre Dragicevic, and Jean-Daniel Fekete. Color lens: Adaptive color scale optimization for visual exploration. IEEE Transactions on Visualization and Computer Graphics, 17(6): 795–807, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ebert:2010:MIO

[EDK10] David S. Ebert, John Dill, and David J. Kasik. In memoriam: Illuminating our paths — James (Jim) Joseph Thomas. IEEE Transactions on Visualization and Computer Graphics, 16(6):xxi, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Evrard:2019:SRD

[EDvW19] F. Evrard, F. Denner, and B. van Wachem. Surface reconstruction from discrete indicator functions. IEEE Transactions on Visualization and Computer Graphics, 25 (3):1629–1635, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elmqvist:2010:HAI

[EF10] Niklas Elmqvist and Jean-Daniel Fekete. Hierarchical aggregation for information visualization: Overview, techniques, and design guidelines. IEEE Transactions on Visualization and Computer Graphics, 16(3):439–454, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Endert:2012:SIS

[EFN12] Alex Endert, Patrick Fiaux, and Chris North. Semantic interaction for sensemaking: Inferring analytical reasoning for model steering. IEEE Transactions on Visualization and Computer Graphics, 18(12):2879–2888, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elber:2009:ESS

[EG09] Gershon Elber and Tom Grandine. An efficient solution to systems of multivariate polynomial using expression trees. IEEE Transactions on Visualization and Computer Graphics, 15(4):596– 604, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Engel:2012:VSV

[EGG^{+}12] Daniel Engel, Klaus Greff, Christoph Garth, Keith Bein, Anthony Wexler, Bernd Hamann, and Hans Hagen. Visual steering and verification of mass spectrometry data factorization in air quality research. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2275–2284, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ellsworth:2006:CVP

[EGH^{+}06] David Ellsworth, Bryan Green, Chris Henze, Patrick Moran, and Timothy Sandstrom. Concurrent visualization in a production supercomputing environment. IEEE Transactions on Visualization and Computer Graphics, 12(5):997–1004, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Etzmuss:2003:DPS

[EGS03] Olaf Etzmuss, Joachim Gross, and Wolfgang Strasser. Deriving a particle system from continuum mechanics for the animation of deformable objects. IEEE Transactions on Visualization and Computer Graphics, 9(4):538–550, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/04/v0538abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/04/
                                        v0538.pdf.

Eitz:2011:SBI

[EHBA11] Mathias Eitz, Kristian Hildebrand, Tamy Boubekeur, and Marc Alexa. Sketch-based image retrieval: Benchmark and bag-of-features descriptors. IEEE Transactions on Visualization and Computer Graphics, 17(11):1624– 1636, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Erat:2019:RTV

[EHH^{+}19] O. Erat, M. Hoell, K. Haubenwallner, C. Pirchheim, and D. Schmalstieg. Realtime view planning for unstructured lumigraph modeling. IEEE Transactions on Visualization and Computer Graphics, 25(11):3063– 3072, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ersoy:2011:SBE

[EHP^{+}11] Ozan Ersoy, Christophe Hurter, Fernando Paulovich, Gabriel Cantareiro, and Alex Telea. Skeleton-based edge bundling for graph visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2364– 2373, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Eichelbaum:2013:LIT

[EHS13] Sebastian Eichelbaum, Mario Hlawitschka, and Gerik Scheuermann. LineAO — improved three-dimensional line rendering. IEEE Transactions on Visualization and Computer Graphics, 19(3):433– 445, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Eick:2000:VDA

[Eic00] S. G. Eick. Visual discovery and analysis. IEEE Transactions on Visualization and Computer Graphics, 6(1):44– 58, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0044.
pdf;   http://www.computer.
org/tvcg/tg2000/v0044abs.
htm.

Erat:2018:DAH

[EIKS18] Okan Erat, Werner Alexander Isop, Denis Kalkofen, and Dieter Schmalstieg. Droneaugmented human vision: Exocentric control for drones exploring hidden areas. IEEE Transactions on Visualization and Computer Graphics, 24(4):1437–1446, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260942-
abs.html.

Etiene:2014:VVR

[EJR^{+}14] Tiago Etiene, Daniel Jonsson, Timo Ropinski, Carlos Scheidegger, Joao L. D. Comba, Luis Gustavo Nonato, Robert M. Kirby, Anders Ynnerman, and Claudio T. Silva. Verifying volume rendering using discretization error analysis. IEEE Transactions on Visualization and Computer Graphics, 20(1): 140–154, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elber:1995:LAR

[Elb95] Gershon Elber. Line art rendering via a coverage of isoparametric curves. IEEE Transactions on Visualization and Computer Graphics, 1 (3):231–239, September 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0231.
pdf;   http://www.computer.
org/tvcg/tg1995/v0231abs.
htm.

Elber:1998:LAI

[Elb98] G. Elber. Line art illustrations of parametric and implicit forms. IEEE Transactions on Visualization and Computer Graphics, 4(1):71– 81, January/March 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0071.
pdf;   http://www.computer.
org/tvcg/tg1998/v0071abs.
htm.

English:2013:IFS

[ELF13] R. Elliot English, Michael Lentine, and Ron Fedkiw. Interpenetration free simulation of thin shell rigid bodies. IEEE Transactions on Visualization and Computer Graphics, 19(6):991–1004, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Entezari:2006:EZP

[EM06] Alireza Entezari and Torsten Möller. Extensions of the Zwart–Powell box spline for volumetric data reconstruction on the Cartesian lattice. IEEE Transactions on Visualization and Computer Graphics, 12(5):1337–1344, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Etemadpour:2015:PBE

[EMdSP^{+}15] Ronak Etemadpour, Robson Motta, Jose Gustavo de Souza Paiva, Rosane Minghim, Maria Cristina Ferreira de Oliveira, and Lars Linsen. Perception-based evaluation of projection methods for multidimensional data visualization. IEEE Transactions on Visualization and Computer Graphics, 21(1): 81–94, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06832613-
abs.html.

Eilemann:2009:ESP

[EMP09] Stefan Eilemann, Maxim Makhinya, and Renato Pajarola. Equalizer: a scalable parallel rendering framework. IEEE Transactions on Visualization and Computer Graphics, 15(3):436–452, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ebert:2002:DET

[EMRY02] D. S. Ebert, C. J. Morris, P. Rheingans, and T. S. Yoo. Designing effective transfer functions for volume rendering from photographic volumes. IEEE Transactions on Visualization and Computer Graphics, 8(2): 183–197, April 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0183.
pdf;   http://www.computer.
org/tvcg/tg2002/v0183abs.
htm.

Etiene:2012:TVI

[ENS^{+}12] Tiago Etiene, L. Gustavo Nonato, Carlos Scheidegger, Julien Tierny, Thomas J. Peters, Valerio Pascucci, Robert M. Kirby, and Claudio T. Silva. Topology verification for isosurface extraction. IEEE Transactions on Visualization and Computer Graphics, 18(6): 952–965, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Eck:2015:PHD

[EPS^{+}15] U. Eck, F. Pankratz, C. Sandor, G. Klinker, and H. Laga. Precise haptic device colocation for visuo-haptic augmented reality. IEEE Transactions on Visualization and Computer Graphics, 21(12): 1427–1441, December 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Etzion:1997:CSD

[ER97] M. Etzion and A. Rappoport.

On compatible star decompositions of simple polygons. IEEE Transactions on Visualization and Computer Graphics, 3(1):87–95, January/March 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0087.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0087abs.
                                        htm.

Elmqvist:2010:MSF

[ERHRF10] Niklas Elmqvist, Yann Riche, Nathalie Henry-Riche, and Jean-Daniel Fekete. Melange: Space folding for visual exploration. IEEE Transactions on Visualization and Computer Graphics, 16(3):468– 483, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ehlke:2013:FGV

[ERL^{+}13] Moritz Ehlke, Heiko Ramm, Hans Lamecker, Hans-Christian Hege, and Stefan Zachow. Fast generation of virtual Xray images for reconstruction of 3D anatomy. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2673–2682, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Edge:2018:BTA

[ERLW18] D. Edge, N. H. Riche, J. Larson, and C. White. Beyond tasks: An activity typology for visual analytics.

IEEE Transactions on Visualization and Computer Graphics, 24(1):267–277, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2007:EMN

[Ert07a] Thomas Ertl. Editorial: a message from the new Editorin-Chief. IEEE Transactions on Visualization and Computer Graphics, 13(1):2, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/01/v0002.
pdf.

Ertl:2007:ENa

[Ert07b]  Thomas Ertl.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 13(3):417– 419, May/June 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/03/v0417.
pdf.

Ertl:2007:ENb

[Ert07c] Thomas Ertl. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 13(5):849–850, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/05/v0849.
                                        pdf.

Ertl:2008:EN

[Ert08]  Thomas Ertl.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 14(2):245, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2008/02/ttg2008020245.
                                        pdf.

Ertl:2009:EEN

[Ert09a] Thomas Ertl. Editorial: Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 15(1):1– 3, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2009:EN

[Ert09b] Thomas Ertl. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 15(5):707–708, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2010:ENa

[Ert10a]  Thomas Ertl.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 16(1):1, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2010:ENb

[Ert10b]  Thomas Ertl.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 16(4):529– 530, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2010:GEI

[Ert10c] Thomas Ertl. Guest Editor’s introduction: Special section on the IEEE Symposium on Visual Analytics Science and Technology (VAST). IEEE Transactions on Visualization and Computer Graphics, 16(2):177, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2010:MEC

[Ert10d] Thomas Ertl. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 16(6):x, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2011:EEF

[Ert11a] Thomas Ertl. Editorial: EIC farewell and new EIC introduction. IEEE Transactions on Visualization and Computer Graphics, 17(3):

261–262, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ertl:2011:EN

[Ert11b] Thomas Ertl. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 17(2):129, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ebert:2001:MIF

[ES01] David S. Ebert and Christopher D. Shaw. Minimally immersive flow visualization. IEEE Transactions on Visualization and Computer Graphics, 7(4):343– 350, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0343.
pdf;   http://www.computer.
org/tvcg/tg2001/v0343abs.
htm.

Ebling:2005:CFT

[ES05] Julia Ebling and Gerik Scheuermann. Clifford Fourier Transform on vector fields. IEEE Transactions on Visualization and Computer Graphics, 11(4):469–479, July/ August 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Etiene:2009:VVI

[ESN^{+}09] Tiago Etiene, Carlos Scheidegger, Luis Gustavo Nonato, Robert Mike Kirby, and Claudio Silva. Verifiable visualization for isosurface extraction. IEEE Transactions on Visualization and Computer Graphics, 15(6):1227–1234, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Enderton:2011:ST

[ESSL11] Eric Enderton, Erik Sintorn, Peter Shirley, and David Luebke. Stochastic transparency. IEEE Transactions on Visualization and Computer Graphics, 17(8):1036–1047, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

El-Sana:1998:TSP

[ESV98] J. El-Sana and A. Varshney. Topology simplification for polygonal virtual environments. IEEE Transactions on Visualization and Computer Graphics, 4(2):133– 144, April/June 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0133.
pdf;   http://www.computer.
org/tvcg/tg1998/v0133abs.
htm.

Elmqvist:2008:TOM

[ET08] Niklas Elmqvist and Philippas Tsigas. A taxonomy of 3D occlusion management for visualization. IEEE Transactions on Visualization and Computer Graphics, 14(5):1095–1109, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ebert:2010:TET

[ETO^{+}10] Achim Ebert, Sebastian Thelen, Peter-Scott Olech, Joerg Meyer, and Hans Hagen. Tiled++: an enhanced tiled hi-res display wall. IEEE Transactions on Visualization and Computer Graphics, 16(1):120–132, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Entezari:2008:PBS

[EVM08] Alireza Entezari, Dimitri Van De Ville, and Torsten Möller. Practical box splines for reconstruction on the body centered cubic lattice. IEEE Transactions on Visualization and Computer Graphics, 14(2):313–328, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ewins:1998:MML

[EWWL98] J. P. Ewins, M. D. Waller, M. White, and P. F. Lister. MIP-map level selection for texture mapping. IEEE Transactions on Visualization and Computer Graphics, 4(4):317–329, October/December 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0317.
pdf;   http://www.computer.
org/tvcg/tg1998/v0317abs.
htm.

Yoon:2007:RAC

[eYL07] Sung eui Yoon and Peter Lindstrom. Randomaccessible compressed triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 13(6):1536– 1543, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fu:2015:TIH

[FA15] Fred Fu and Nasser Mohieddin Abukhdeir. A topologically-informed hyperstreamline seeding method for alignment tensor fields. IEEE Transactions on Visualization and Computer Graphics, 21 (3):413–419, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06933942-
abs.html.

Farin:2012:SMT

[Far12] Gerald Farin. Shape measures for triangles. IEEE Transactions on Visualization and Computer Graphics, 18 (1):43–46, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Faure:1999:FIR

[Fau99] François Faure. Fast iterative refinement of articulated solid dynamics. IEEE Transactions on Visualization and Computer Graphics, 5(3):268– 276, July/September 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0268.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0268abs.
                                        htm.

Fraedrich:2010:EHQ

[FAW10] Roland Fraedrich, Stefan Auer, and Rudiger Westermann. Efficient high-quality volume rendering of SPH data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1533– 1540, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frohlich:2007:GEI

[FBI07] Bernd Fröhlich, Doug A. Bowman, and Hiroo Iwata. Guest Editors’ introduction: Special section on virtual reality. IEEE Transactions on Visualization and Computer Graphics, 13(3):420– 421, May/June 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/03/v0420.
pdf.

Forbes:2018:DIN

[FBL^{+}18] A. G. Forbes, A. Burks, K. Lee, X. Li, P. Boutillier, J. Krivine, and W. Fontana. Dynamic influence networks for rule-based models. IEEE Transactions on Visualization and Computer Graphics, 24 (1):184–194, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2005:RIM

[FBLS05] Martin Fuchs, Volker Blanz, Hendrik Lensch, and Hans-Peter Seidel. Reflectance from images: a model-based approach for human faces. IEEE Transactions on Visualization and Computer Graphics, 11 (3):296–305, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fulda:2016:TIA

[FBM16] J. Fulda, M. Brehmel, and T. Munzner. TimeLineCurator: Interactive authoring of visual timelines from unstructured text. IEEE Transactions on Visualization and Computer Graphics, 22(1): 300–309, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ferstl:2010:ISS

[FBTW10] Florian Ferstl, Kai Burger, Holger Theisel, and Rudiger Westermann. Interactive separating streak surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1569–1577, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ferstl:2016:SVP

[FBW16] F. Ferstl, K. Burger, and R. Westermann. Streamline variability plots for characterizing the uncertainty in vector field ensembles. IEEE Transactions on Visualization and Computer Graphics, 22 (1):767–776, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Forssell:1995:ULI

[FC95] Lisa K. Forssell and Scott D. Cohen. Using line integral convolution for flow visualization: Curvilinear grids, variable-speed animation, and unsteady flows. IEEE Transactions on Visualization and Computer Graphics, 1(2): 133–141, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0133.
pdf;   http://www.computer.
org/tvcg/tg1995/v0133abs.
htm.

Forsberg:2009:CVF

[FCL09] Andrew Forsberg, Jian Chen, and David Laidlaw. Comparing 3D vector field visualization methods: a user study. IEEE Transactions on Visualization and Computer Graphics, 15(6):1219–1226, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fu:2017:PIS

[FCSF17] Qiang Fu, Xiaowu Chen, Xiaoyu Su, and Hongbo Fu. Pose-inspired shape synthesis and functional hybrid. IEEE Transactions on Visualization and Computer Graphics, 23(12):2574– 2585, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/08089792-
abs.html.

Foxlin:2015:DEA

[FCZ15] E. Foxlin, T. Calloway, and H. Zhang. Design and error analysis of a vehicular AR system with autoharmonization. IEEE Transactions on Visualization and Computer Graphics, 21(12): 1323–1335, December 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fu:2018:HDA

[FDC^{+}18] S. Fu, H. Dong, W. Cui, J. Zhao, and H. Qu. How do ancestral traits shape family trees over generations? IEEE Transactions on Visualization and Computer Graphics, 24 (1):205–214, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fisher:2010:WEA

[FDFR10] Danyel Fisher, Steven Drucker, Roland Fernandez, and Scott Ruble. WebCharts: Extending applications with Webauthored, embeddable visualizations. IEEE Transactions on Visualization and Computer Graphics, 16(6):1157– 1163, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feng:2017:HEE

[FDPH17] Mi Feng, Cheng Deng, Evan M. Peck, and Lane Harrison. HindSight: Encouraging exploration through direct encoding of personal interaction history. IEEE Transactions on Visualization and Computer Graphics, 23(1): 351–360, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frey:2017:PDV

[FE17] Steffen Frey and Thomas Ertl.

Progressive direct volume-to-volume transformation. IEEE Transactions on Visualization and Computer Graphics, 23 (1):921–930, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Felix:2018:TWC

[FFB18] C. Felix, S. Franconeri, and E. Bertini. Taking word clouds apart: An empirical investigation of the design space for keyword summaries. IEEE Transactions on Visualization and Computer Graphics, 24(1):657–666, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Favelier:2019:PAC

[FFST19] Guillaume Favelier, Noura Faraj, Brian Summa, and Julien Tierny. Persistence atlas for critical point variability in ensembles. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1152–1162, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2019/01/08457259-
abs.html.

Frisken-Gibson:1999:ULV

[FG99] S. F. Frisken-Gibson. Using linked volumes to model object collisions, deformation, cutting, carving, and joining. IEEE Transactions on Visualization and Computer Graphics, 5(4):333–348, October/December 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0333.
pdf;   http://www.computer.
org/tvcg/tg1999/v0333abs.
htm.

Francois:2009:IBM

[FGBB09] Guillaume Francois, Pascal Gautron, Gaspard Breton, and Kadi Bouatouch. Imagebased modeling of the human eye. IEEE Transactions on Visualization and Computer Graphics, 15(5): 815–827, September/October 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fougerolle:2005:BOI

[FGF^{+}05] Yohan D. Fougerolle, Andrei Gribok, Sebti Foufou, Frederic Truchetet, and Mongi A. Abidi. Boolean operations with implicit and parametric representation of primitives using R-functions. IEEE Transactions on Visualization and Computer Graphics, 11(5):529–539, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Faust:2019:DAL

[FGS19] Rebecca Faust, David Glickenstein, and Carlos Scheidegger. DimReader: Axis lines that explain non-linear projections. IEEE Transactions on Visualization and Computer Graphics, 25(1):481– 490, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440820-
abs.html.

Fang:2006:RAT

[FH06] Hui Fang and John C. Hart.

RotoTexture: Automated tools for texturing raw video. IEEE Transactions on Visualization and Computer Graphics, 12(6):1580–1589, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/06/v1580s.
mpg.

Fu:2007:TSV

[FH07] Chi-Wing Fu and Andrew J.

Hanson. A transparently scalable visualization architecture for exploring the universe. IEEE Transactions on Visualization and Computer Graphics, 13(1):108–121, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2016:IIV

[FH16] F. G. Fuchs and J. M. Hjelmervik. Interactive isogeometric volume visualization with pixel-accurate geometry. IEEE Transactions on Visualization and Computer Graphics, 22(2):1102–1114, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fritz:2009:VAE

[FHG^{+}09] Laura Fritz, Markus Hadwiger, Georg Geier, Gerhard Pittino, and M. Eduard Gröller. A visual approach to efficient analysis and quantification of ductile iron and reinforced sprayed concrete. IEEE Transactions on Visualization and Computer Graphics, 15(6):1343–1350, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feng:2008:ANS

[FHHJ08] Louis Feng, Ingrid Hotz, Bernd Hamann, and Kenneth Joy. Anisotropic noise samples. IEEE Transactions on Visualization and Computer Graphics, 14(2):342– 354, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Federico:2017:SVA

[FHKM17] Paolo Federico, Florian Heimerl, Steffen Koch, and Silvia Miksch. A survey on visual approaches for analyzing scientific literature and patents. IEEE Transactions on Visualization and Computer Graphics, 23(9):2179– 2198, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07570239-
abs.html.

Forbes:2010:BFD

[FHL10] Angus Forbes, Tobias Hollerer, and George Legrady. “behaviorism”: a framework for dynamic data visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1164–1171, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fink:2012:ALF

[FHS^{+}12] Martin Fink, Jan-Henrik Haunert, Andre Schulz, Joachim Spoerhase, and Alexander Wolff. Algorithms for labeling focus regions. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2583–2592, December 2012.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fink:2013:SAR

[FHSW13] Martin Fink, Jan-Henrik Haunert, Joachim Spoerhase, and Alexander Wolff. Selecting the aspect ratio of a scatter plot based on its Delaunay triangulation. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2326–2335, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2014:ICS

[FIB^{+}14] Johannes Fuchs, Petra Isenberg, Anastasia Bezerianos, Fabian Fischer, and Enrico Bertini. The influence of contour on similarity perception of star glyphs. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2251–2260, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875973-
                                        abs.html.

Fuchs:2017:SRE

[FIBK17] Johannes Fuchs, Petra Isenberg, Anastasia Bezerianos, and Daniel Keim. A systematic review of experimental studies on data glyphs. IEEE Transactions on Visualization and Computer Graphics, 23(7):1863–1879, July 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07445239-
abs.html.

Fisher:2007:HLG

[Fis07]  Danyel Fisher.  Hotmap:

Looking at geographic attention. IEEE Transactions on Visualization and Computer Graphics, 13(6):1184– 1191, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feng:2010:MVS

[FKLT10] David Feng, Lester Kwock, Yueh Lee, and Russell Taylor. Matching visual saliency to confidence in plots of uncertain data. IEEE Transactions on Visualization and Computer Graphics, 16(6):980– 989, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ferstl:2017:THC

[FKRW17] Florian Ferstl, Mathias Kanzler, Marc Rautenhaus, and Rüdiger Westermann. Timehierarchical clustering and visualization of weather forecast ensembles. IEEE Transactions on Visualization and Computer Graphics, 23(1): 831–840, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Friston:2016:ELL

[FKS16] S. Friston, P. Karlström, and A. Steed. The effects of low latency on pointing and steering tasks. IEEE Transactions on Visualization and Computer Graphics, 22(5):1605– 1615, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ferreira:2011:BVU

[FLF^{+}11] Nivan Ferreira, Lauro Lins, Daniel Fink, Steve Kelling, Christopher Wood, Juliana Freire, and Claudio Silva. BirdVis: Visualizing and understanding bird populations. IEEE Transactions on Visualization and Computer Graphics, 17(12):2374– 2383, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Floriani:2016:EN

[Flo16]  L. D. Floriani.  Editor’s

note. IEEE Transactions on Visualization and Computer Graphics, 22(3):1179– 1180, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Floriani:2017:MEC

[Flo17]  Leila De Floriani.  Mes-

sage from the Editor-in-Chief.

IEEE Transactions on Visualization and Computer Graphics, 23(1):x, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Floriani:2018:SJ

[Flo18] L. D. Floriani. State of the journal. IEEE Transactions on Visualization and Computer Graphics, 24(2):1036– 1037, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fusiello:2004:ASM

[FM04] Andrea Fusiello and Vittorio Murino. Augmented scene modeling and visualization by optical and acoustic sensor integration. IEEE Transactions on Visualization and Computer Graphics, 10(6):625–636, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0625abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0625.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0625.pdf.

Forsyth:2006:PHG

[FM06] Benjamin A. C. Forsyth and Karon E. MacLean. Predictive haptic guidance: Intelligent user assistance for the control of dynamic tasks. IEEE Transactions on Visualization and Computer Graphics, 12(1):103–113, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fout:2007:TCH

[FM07] Nathaniel Fout and Kwan-Liu Ma. Transform coding for hardware-accelerated volume rendering. IEEE Transactions on Visualization and Computer Graphics, 13(6):1600– 1607, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fout:2012:APB

[FM12a] Nathaniel Fout and Kwan-Liu Ma. An adaptive predictionbased approach to lossless compression of floating-point volume data. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2295–2304, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fout:2012:FVR

[FM12b] Nathaniel Fout and Kwan-Liu Ma. Fuzzy volume rendering. IEEE Transactions on Visualization and Computer Graphics, 18(12):2335– 2344, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Freiler:2008:IVA

[FMH08] Wolfgang Freiler, Krešimir Matković, and Helwig Hauser. Interactive visual analysis of set-typed data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1340–1347, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fujishiro:1996:VDE

[FMST96] Issei Fujishiro, Yuji Maeda, Hiroshi Sato, and Yuriko Takeshima. Volumetric data exploration using interval volume. IEEE Transactions on Visualization and Computer Graphics, 2(2): 144–155, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0144.
pdf;   http://www.computer.
org/tvcg/tg1996/v0144abs.
htm.

Ferstay:2013:VVV

[FNM13] Joel A. Ferstay, Cydney B.

Nielsen, and Tamara Munzner. Variant view: Visualizing sequence variants in their gene context. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2546–2555, December 2013.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Felix:2017:TIV

[FPB17] Cristian Felix, Anshul Vikram Pandey, and Enrico Bertini. TextTile: An interactive visualization tool for seamless exploratory analysis of structured data and unstructured text. IEEE Transactions on Visualization and Computer Graphics, 23(1):161–170, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2008:PVC

[FPH^{+}08] Raphael Fuchs, Ronald Peikert, Helwig Hauser, Filip Sadlo, and Philipp Muigg. Parallel vectors criteria for unsteady flow vertices. IEEE Transactions on Visualization and Computer Graphics, 14 (3):615–626, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feng:2019:PPQ

[FPH19] Mi Feng, Evan Peck, and Lane Harrison. Patterns and pace: Quantifying diverse exploration behavior with visualizations on the Web. IEEE Transactions on Visualization and Computer Graphics, 25 (1):501–511, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08454489-
abs.html.

Ferreira:2013:VEB

[FPV^{+}13] Nivan Ferreira, Jorge Poco, Huy T. Vo, Juliana Freire, and Claudio T. Silva. Visual exploration of big spatiotemporal urban data: A study of New York City taxi trips. IEEE Transactions on Visualization and Computer Graphics, 19(12):2149– 2158, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Flagg:2013:VBC

[FR13] Matthew Flagg and James M.

Rehg. Video-based crowd synthesis. IEEE Transactions on Visualization and Computer Graphics, 19(11):1935– 1947, November 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frankel:2012:VCS

[Fra12]  Felice Frankel.  VisWeek

2012 capstone speaker. IEEE Transactions on Visualization and Computer Graphics, 18(12):xxii, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feigl:2019:SMM

[FRG^{+}19] T. Feigl, D. Roth, S. Gradl, M. Wirth, M. E. Latoschik, B. M. Eskofier, M. Philippsen, and C. Mutschler. Sick moves! Motion parameters as indicators of simulator sickness. IEEE Transactions on Visualization and Computer Graphics, 25(11):3146– 3157, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Francis:2004:VSE

[FS04] George Francis and John M. Sullivan. Visualizing a sphere eversion. IEEE Transactions on Visualization and Computer Graphics, 10(5):509–515, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0509.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0509.pdf.

Friston:2014:MLV

[FS14] Sebastian Friston and Anthony Steed. Measuring latency in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 20(4):616–625, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Friston:2019:RTC

[FS19] S. Friston and A. Steed. Realtime collision detection for deformable characters with radial fields. IEEE Transactions on Visualization and Computer Graphics, 25(8):2611– 2622, August 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frey:2012:VTS

[FSE12] Steffen Frey, Filip Sadlo, and Thomas Ertl. Visualization of temporal similarity in field data. IEEE Transactions on Visualization and Computer Graphics, 18(12):2023– 2032, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fierz:2012:MLT

[FSHH12] Basil Fierz, Jonas Spillmann, Iker Aguinaga Hoyos, and Matthias Harders. Maintaining large time steps in explicit finite element simulations using shape matching. IEEE Transactions on Visualization and Computer Graphics, 18 (5):717–728, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frey:2014:IPV

[FSME14] Steffen Frey, Filip Sadlo, Kwan-Liu Ma, and Thomas Ertl. Interactive progressive visualization with space–time error control. IEEE Transactions on Visualization and Computer Graphics, 20(12):

2397–2406, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875936-
                                        abs.html.

Fujimoto:2014:GCP

[FST^{+}14] Yuichiro Fujimoto, Ross T. Smith, Takafumi Taketomi, Goshiro Yamamoto, Jun Miyazaki, Hirokazu Kato, and Bruce H. Thomas. Geometrically-correct projectionbased texture mapping onto a deformable object. IEEE Transactions on Visualization and Computer Graphics, 20 (4):540–549, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Friston:2016:CEU

[FSTG16] S. Friston, A. Steed, S. Tilbury, and G. Gaydadjiev. Construction and evaluation of an ultra low latency frameless renderer for VR. IEEE Transactions on Visualization and Computer Graphics, 22(4): 1377–1386, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fraedrich:2009:EMR

[FSW09] Roland Fraedrich, Jens Schneider, and Rudiger Westermann. Exploring the Millennium Run — scalable rendering of large-scale cosmological datasets. IEEE Transactions on Visualization and Computer Graphics, 15(6):1251– 1258, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frishman:2007:MLG

[FT07] Yaniv Frishman and Ayellet Tal. Multi-level graph layout on the GPU. IEEE Transactions on Visualization and Computer Graphics, 13(6):1310–1319, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Frishman:2008:ODG

[FT08] Yaniv Frishman and Ayellet Tal. Online dynamic graph drawing. IEEE Transactions on Visualization and Computer Graphics, 14(4):727– 740, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040727s.
zip.

Frishman:2009:UGL

[FT09] Yaniv Frishman and Ayellet Tal. Uncluttering graph layouts using anisotropic diffusion and mass transport. IEEE Transactions on Visualization and Computer Graphics, 15(5):777–788, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feng:2013:CLS

[FT13] Xin Feng and Yiying Tong. Choking loops on surfaces. IEEE Transactions on Visualization and Computer Graphics, 19(8):1298–1306, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fang:2013:VNI

[FTB^{+}13] Hui Fang, Gary Kwok-Leung Tam, Rita Borgo, Andrew J. Aubrey, Philip W. Grant, Paul L. Rosin, Christian Wallraven, Douglas Cunningham, David Marshall, and Min Chen. Visualizing natural image statistics. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1228–1241, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Froese:2013:ESD

[FTES13] Maria-Elena Froese, Melanie Tory, Guy-Warwick Evans, and Kedar Shrikhande. Evaluation of static and dynamic visualization training approaches for users with different spatial abilities. IEEE Transactions on Visualization and Computer Graphics, 19(12):2810–2817, December 2013. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2013:VRC

[Fuc13] Henry Fuchs. The 2013 Virtual Reality Career Award. IEEE Transactions on Visualization and Computer Graphics, 19(4):xvii, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Faloutsos:1997:DFF

[FvdPT97] P. Faloutsos, M. van de Panne, and D. Terzopoulos. Dynamic free-form deformations for animation synthesis. IEEE Transactions on Visualization and Computer Graphics, 3(3):201–214, July/September 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0201.
pdf;   http://www.computer.
org/tvcg/tg1997/v0201abs.
htm.

Fekete:2010:P

[FvHM^{+}10] Jean-Daniel Fekete, Frank van Ham, Raghu Machiraju, Torsten Moller, and Hanspeter Pfister. Preface. IEEE Transactions on Visualization and Computer Graphics, 16(6):xi–xx, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Falk:2008:OSL

[FW08] Martin Falk and Daniel Weiskopf. Output-sensitive 3D line integral convolution. IEEE Transactions on Visualization and Computer Graphics, 14(4):820– 834, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2008/04/ttg2008040820s.
                                        zip.

Fang:2017:CCO

[FWD^{+}17] H. Fang, S. Walton, E. Delahaye, J. Harris, D. A. Storchak, and M. Chen. Categorical colormap optimization with visualization case studies. IEEE Transactions on Visualization and Computer Graphics, 23(1):871–880, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fuchs:2009:VHM

[FWG09] Raphael Fuchs, Jürgen Waser, and Meister Eduard Gröller. Visual human + machine learning. IEEE Transactions on Visualization and Computer Graphics, 15(6):1327– 1334, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Freitag:2016:ERG

[FWK16] S. Freitag, B. Weyers, and T. W. Kuhlen. Examining rotation gain in CAVE-like virtual environments. IEEE Transactions on Visualization and Computer Graphics, 22 (4):1462–1471, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fischbach:2017:SEC

[FWL17] Martin Fischbach, Dennis Wiebusch, and Marc Erich Latoschik. Semantic entitycomponent state management techniques to enhance software quality for multimodal VR-systems. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1342–1351, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829420-
abs.html.

Fua:2000:SBB

[FWR00] Ying-Huey Fua, M. O. Ward, and E. A. Rundensteiner. Structure-based brushes: a mechanism for navigating hierarchically organized data and information spaces. IEEE Transactions on Visualization and Computer Graphics, 6 (2):150–159, April/June 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0150.
pdf;   http://www.computer.

                                        org/tvcg/tg2000/v0150abs.
                                        htm.

Feng:2012:CTV

[FWSL12] Kun-Chuan Feng, Chaoli Wang, Han-Wei Shen, and Tong-Yee Lee. Coherent timevarying graph drawing with multifocus+context interaction. IEEE Transactions on Visualization and Computer Graphics, 18(8):1330– 1342, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fu:2004:BSP

[FWT^{+}04] Chi-Wing Fu, Tien-Tsin Wong, Wai-Shun Tong, Chi-Keung Tang, and Andrew J. Hanson. Binary-space-partitioned images for resolving image-based visibility. IEEE Transactions on Visualization and Computer Graphics, 10(1):58–71, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/01/v0058abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/01/
                                        v0058.pdf.

Fang:2013:IBM

[FWZQ13] Tian Fang, Zhexi Wang, Honghui Zhang, and Long Quan. Image-based modeling of unwrappable facades. IEEE Transactions on Visualization and Computer Graphics, 19(10):1720–1731, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fan:2012:EBA

[FXG12] Rukun Fan, Songhua Xu, and Weidong Geng. Examplebased automatic music-driven conventional dance motion synthesis. IEEE Transactions on Visualization and Computer Graphics, 18(3): 501–515, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fang:2018:NCI

[FYF^{+}18] Fei Fang, Miao Yi, Hui Feng, Shenghong Hu, and Chunxia Xiao. Narrative collage of image collections by scene graph recombination. IEEE Transactions on Visualization and Computer Graphics, 24(9): 2559–2572, September 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08057796-
abs.html.

Fan:2010:RFP

[FYP10] Hanqi Fan, Yizhou Yu, and Qunsheng Peng. Robust feature-preserving mesh denoising based on consistent subneighborhoods. IEEE Transactions on Visualization and Computer Graphics, 16(2):312–324, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Falk:2019:IVH

[FYTL19] Martin Falk, Anders Ynnerman, Darren Treanor, and Claes Lundstrom. Interactive visualization of 3D histopathology in native resolution. IEEE Transactions on Visualization and Computer Graphics, 25(1):1008– 1017, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440070-
abs.html.

Feng:2016:TMM

[FYWY16] Wei Feng, Ying Yang, Liang Wan, and Changguo Yu. Tone-mapped mean-shift based environment map sampling. IEEE Transactions on Visualization and Computer Graphics, 22(9):2187–2199, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fu:2019:HRB

[FYY^{+}19] Y. Fu, H. Yu, C. Yeh, J. Zhang, and T. Lee. High relief from brush painting. IEEE Transactions on Visualization and Computer Graphics, 25(9):2763–2776, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fan:2017:CHA

[FYZ^{+}17] Jingfan Fan, Jian Yang, Yitian Zhao, Danni Ai, Yonghuai Liu, Ge Wang, and Yongtian Wang. Convex hull aided registration method (CHARM). IEEE Transactions on Visualization and Computer Graphics, 23(9): 2042–2055, September 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07557083-
abs.html.

Fuller:2007:STD

[FZC^{+}07] Alfred Fuller, Robert Zawadzki, Stacey Choi, David Wiley, John Werner, and Bernd Hamann. Segmentation of three-dimensional retinal image data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1719–1726, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1719s.
mov.

Fu:2017:VAM

[FZCQ17] Siwei Fu, Jian Zhao, Weiwei Cui, and Huamin Qu. Visual analysis of MOOC forums with iForum. IEEE Transactions on Visualization and Computer Graphics, 23 (1):201–210, January 2017.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gosink:2007:VIQ

[GABJ07] Luke Gosink, John Anderson, Wes Bethel, and Kenneth Joy. Variable interactions in query-driven visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6):1400–1407, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gosink:2008:QDV

[GABJ08] Luke J. Gosink, John C. Anderson, E. Wes Bethel, and Kenneth I. Joy. Query-driven visualization of time-varying adaptive mesh refinement data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1715– 1722, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guetat:2010:PIV

[GAMD10] Amel Guetat, Alexandre Ancel, Stephane Marchesin, and Jean-Michel Dischler. Preintegrated volume rendering with non-linear gradient interpolation. IEEE Transactions on Visualization and Computer Graphics, 16(6):1487– 1494, November/December 2010. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grave:2008:VGU

[GB08a] Frank Grave and Michael Buser. Visiting the Gödel universe. IEEE Transactions on Visualization and Computer Graphics, 14(6):1563– 1570, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grundhofer:2008:RTA

[GB08b] Anselm Grundhöfer and Oliver Bimber. Real-time adaptive radiometric compensation. IEEE Transactions on Visualization and Computer Graphics, 14(1):97–108, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gunther:2014:CMI

[GBCG^{+}14] David Gunther, Roberto A.

Boto, Juila Contreras-Garcia, Jean-Philip Piquemal, and Julien Tierny. Characterizing molecular interactions in chemical systems. IEEE Transactions on Visualization and Computer Graphics, 20(12):2476–2485, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875922-
abs.html.

Gschwandtnei:2016:VET

[GBFM16] T. Gschwandtnei, M. Bogl, P. Federico, and S. Miksch. Visual encodings of temporal uncertainty: A comparative user study. IEEE Transactions on Visualization and Computer Graphics, 22(1): 539–548, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2008:PAM

[GBHP08] Attila Gyulassy, Peer-Timo Bremer, Bernd Hamann, and Valerio Pascucci. A practical approach to Morse–Smale complex computation: Scalability and generality. IEEE Transactions on Visualization and Computer Graphics, 14(6):1619–1626, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grottel:2012:VED

[GBM^{+}12] Sebastian Grottel, Philipp Beck, Christoph Muller, Guido Reina, Johannes Roth, Hans-Rainer Trebin, and Thomas Ertl. Visualization of electrostatic dipoles in molecular dynamics of metal oxides. IEEE Transactions on Visualization and Computer Graphics, 18(12):2061– 2068, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gautron:2007:TRC

[GBP07] Pascal Gautron, Kadi Bouatouch, and Sumanta Pattanaik. Temporal radiance caching. IEEE Transactions on Visualization and Computer Graphics, 13(5): 891–901, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2012:CMS

[GBP12] Attila Gyulassy, Peer-Timo Bremer, and Valerio Pascucci. Computing Morse– Smale complexes with accurate geometry. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2014–2022, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gosink:2013:CVP

[GBP^{+}13] Luke Gosink, Kevin Bensema, Trenton Pulsipher, Harald Obermaier, Michael Henry, Hank Childs, and Kenneth I. Joy. Characterizing and visualizing predictive uncertainty in numerical ensembles through Bayesian model averaging. IEEE Transactions on Visualization and Computer Graphics, 19(12):2703– 2712, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2019:SMP

[GBP19] Attila Gyulassy, Peer-Timo Bremer, and Valerio Pascucci. Shared-memory parallel computation of Morse– Smale complexes with improved accuracy. IEEE Transactions on Visualization and Computer Graphics, 25(1): 1183–1192, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440824-
abs.html.

Gerber:2010:VEH

[GBPW10] Samuel Gerber, Peer-Timo Bremer, Valerio Pascucci, and Ross Whitaker. Visual exploration of high dimensional scalar functions. IEEE Transactions on Visualization and Computer Graphics, 16(6):1271–1280, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Goffin:2017:ESW

[GBWI17] Pascal Goffin, Jeremy Boy, Wesley Willett, and Petra Isenberg. An exploratory study of word-scale graphics in data-rich text documents. IEEE Transactions on Visualization and Computer Graphics, 24(10):2275– 2287, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07593375-
abs.html.

Gao:2015:AEL

[GCL^{+}15] L. Gao, Y. Cao, Y. Lai, H. Huang, L. Kobbelt, and S. Hu. Active exploration of large 3D model repositories. IEEE Transactions on Visualization and Computer Graphics, 21(12):1390– 1402, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Golodetz:2018:CLS

[GCL^{+}18] Stuart Golodetz, Tommaso Cavallari, Nicholas A. Lord, Victor A. Prisacariu, David W. Murray, and Philip H. S. Torr. Collaborative largescale dense 3D reconstruction with online inter-agent pose optimisation. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2895–2905, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8492363/.

Guo:2006:VSS

[GCML06] Diansheng Guo, Jin Chen, Alan M. MacEachren, and Ke Liao. A visualization system for space–time and multivariate patterns (VIS-STAMP). IEEE Transactions on Visualization and Computer Graphics, 12(6):1461– 1474, November/December 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gleicher:2013:PAV

[GCNF13] Michael Gleicher, Michael Correll, Christine Nothelfer, and Steven Franconeri. Perception of average value in multiclass scatterplots. IEEE Transactions on Visualization and Computer Graphics, 19(12):2316–2325, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gao:2017:FRT

[GCT17] Mingcen Gao, Thanh-Tung Cao, and Tiow-Seng Tan. Flip to regular triangulation and convex hull. IEEE Transactions on Visualization and Computer Graphics, 23(2): 1056–1069, February 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/02/07406755-
                                        abs.html.

Gu:2014:SMC

[GCZL14] Xianfeng David Gu, Liming Chen, Wei Zeng, and Huibin Li. Surface meshing with curvature convergence. IEEE Transactions on Visualization and Computer Graphics, 20 (6):919–934, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gain:2001:PSI

[GD01] James E. Gain and Neil A.

Dodgson. Preventing selfintersection under free-form deformation. IEEE Transactions on Visualization and Computer Graphics, 7(4):289– 298, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0289.
pdf;   http://www.computer.
org/tvcg/tg2001/v0289abs.
htm.

Gorochowski:2012:UAV

[GdBG12] Thomas E. Gorochowski, Mario di Bernardo, and Claire S. Grierson. Using aging to visually uncover evolutionary processes on networks. IEEE Transactions on Visualization and Computer Graphics, 18(8):1343– 1352, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Goodwin:2013:CUC

[GDJ^{+}13] Sarah Goodwin, Jason Dykes, Sara Jones, Iain Dillingham, Graham Dove, Alison Duffy, Alexander Kachkaev, Aidan Slingsby, and Jo Wood. Creative user-centered visualization design for energy analysts and modelers. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2516–2525, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gutenko:2017:AVA

[GDKB17] Ievgeniia Gutenko, Konstantin Dmitriev, Arie E. Kaufman, and Matthew A. Barish. AnaFe: Visual analytics of image-derived temporal features — focusing on the spleen. IEEE Transactions on Visualization and Computer Graphics, 23(1):171–180, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2007:TCD

[GDN^{+}07] Attila Gyulassy, Mark Duchaineau, Vijay Natarajan, Valerio Pascucci, Eduardo Bringa, Andrew Higginbotham, and Bernd Hamann. Topologically clean distance fields. IEEE Transactions on Visualization and Computer Graphics, 13(6):1432–1439, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Goodwin:2016:VMV

[GDST16] S. Goodwin, J. Dykes, A. Slingsby, and C. Turkay. Visualizing multiple variables across scale and geography. IEEE Transactions on Visualization and Computer Graphics, 22(1):599–608, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gerry:2017:PMS

[Ger17] Lynda Joy Gerry. Paint with me: Stimulating creativity and empathy while painting with a painter in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1418–1426, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829415-
abs.html.

Ganuza:2014:SEI

[GFG^{+}14] Maria Lujan Ganuza, Gabriela Ferracutti, Maria Florencia Gargiulo, Silvia Mabel Castro, Ernesto Bjerg, Eduard Groller, and Kresimir Matkovic. The Spinel Explorer — interactive visual analysis of spinel group minerals. IEEE Transactions on Visualization and Computer Graphics, 20(12):1913– 1922, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875977-
abs.html.

Gosink:2011:AMS

[GGA^{+}11] Luke J. Gosink, Christoph Garth, John C. Anderson, E. Wes Bethel, and Kenneth I. Joy. An application of multivariate statistical analysis for query-driven visualization. IEEE Transactions on Visualization and Computer Graphics, 17(3): 264–275, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glueck:2017:PCS

[GGC^{+}17] Michael Glueck, Alina Gvozdik, Fanny Chevalier, Azam Khan, Michael Brudno, and Daniel Wigdor. PhenoStacks: Crosssectional cohort phenotype comparison visualizations. IEEE Transactions on Visualization and Computer Graphics, 23(1):191–200, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gabbard:2019:MIS

[GGHZ19] J. L. Gabbard, J. Grubert, S. Hu, and S. Zollmann. Message from the ISMAR 2019 Science and Technology Program Chairs and TVCG Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3050–3051, November 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gimenez:2018:MVA

[GGJ^{+}18] Alfredo Gimenez, Todd Gamblin, Ilir Jusufi, Abhinav Bhatele, Martin Schulz, Peer-Timo Bremer, and Bernd Hamann. MemAxes: Visualization and analytics for characterizing complex memory performance behaviors. IEEE Transactions on Visualization and Computer Graphics, 24(7):2180–2193, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07955092-
abs.html.

Gratzl:2014:DEC

[GGL^{+}14a] Samuel Gratzl, Nils Gehlenborg, Alexander Lex, Hanspeter Pfister, and Marc Streit. Domino: Extracting, comparing, and manipulating subsets across multiple tabular datasets. IEEE Transactions on Visualization and Computer Graphics, 20(12):2023– 2032, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875920-
abs.html.

Gyulassy:2014:CMS

[GGL^{+}14b] Attila Gyulassy, David Gunther, Joshua A. Levine, Julien Tierny, and Valerio Pascucci. Conforming Morse–Smale complexes. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2595–2603, December 2014.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875918-
                                        abs.html.

Guo:2019:BWT

[GGLQ19] H. Guo, Y. Guan, M. Liu, and K. Qin. Biorthogonal wavelet transforms and applications based on generalized progressive Catmull–Clark subdivision with shape control. IEEE Transactions on Visualization and Computer Graphics, 25(7):2392–2403, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guerra-Gomez:2013:VCT

[GGPPS13] JohnAlexis Guerra-Gomez, Michael L. Pack, Catherine Plaisant, and Ben Shneiderman. Visualizing change over time using dynamic hierarchies: TreeVersity2 and the StemView. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2566–2575, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Garth:2007:ECV

[GGTH07] Christoph Garth, Florian Gerhardt, Xavier Tricoche, and Hagen Hans. Efficient computation and visualization of coherent structures in fluid flow applications. IEEE Transactions on Visualization and Computer Graphics, 13(6):1464–1471, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2018:PBI

[GGZ^{+}18] Qiang Guo, Shanshan Gao, Xiaofeng Zhang, Yilong Yin, and Caiming Zhang. Patchbased image inpainting via two-stage low rank approximation. IEEE Transactions on Visualization and Computer Graphics, 24(6): 2023–2036, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07922581-
abs.html.

Guo:2016:CSU

[GGZL16] H. Guo, S. R. Gomez, C. Ziemkiewicz, and D. H. Laidlaw. A case study using visualization interaction logs and insight metrics to understand how analysts arrive at insights. IEEE Transactions on Visualization and Computer Graphics, 22(1): 51–60, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gueziec:1995:ETS

[GH95] André Guéziec and Robert Hummel. Exploiting triangulated surface extraction using tetrahedral decomposition. IEEE Transactions on Visualization and Computer Graphics, 1(4):328– 342, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0328.
pdf;   http://www.computer.
org/tvcg/tg1995/v0328abs.
htm.

Gibson:2000:PDP

[GH00] S. Gibson and R. J. Hubbold. A perceptually-driven parallel algorithm for efficient radiosity simulation. IEEE Transactions on Visualization and Computer Graphics, 6(3):220– 235, July/September 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0220.
pdf;   http://www.computer.
org/tvcg/tg2000/v0220abs.
htm.

Glatter:2008:VTP

[GHA^{+}08] Markus Glatter, Jian Huang, Sean Ahern, Jamison Daniel, and Aidong Lu. Visualizing temporal patterns in large multivariate data using modified globbing. IEEE Transactions on Visualization and Computer Graphics, 14(6):1467–1474, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glueck:2016:PPC

[GHC^{+}16] M. Glueck, P. Hamilton, F. Chevalier, S. Breslav, A. Khan, D. Wigdor, and M. Brudno. PhenoBlocks: Phenotype comparison visualizations. IEEE Transactions on Visualization and Computer Graphics, 22(1): 101–110, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gard:2019:PDB

[GHE19] N. Gard, A. Hilsmann, and P. Eisert. Projection distortion-based object tracking in shader lamp scenarios. IEEE Transactions on Visualization and Computer Graphics, 25(11):3105– 3113, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glatter:2006:SDS

[GHGM06] Markus Glatter, Jian Huang, Jinzhu Gao, and Colin Mollenhour. Scalable data servers for large multivariate volume visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1291–1298, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gieng:1998:CHT

[GHJ^{+}98] T. S. Gieng, B. Hamann, K. I. Joy, G. L. Schussman, and I. J. Trotts. Constructing hierarchies for triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 4(2):145– 161, April/June 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0145.
pdf;   http://www.computer.
org/tvcg/tg1998/v0145abs.
htm.

Grzeszczuk:1997:PBS

[GHK97] R. P. Grzeszczuk, M. Huang, and L. H. Kauffman. Physicallybased stochastic simplification of mathematical knots. IEEE Transactions on Visualization and Computer Graphics, 3(3):262–272, July/ September 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0262.
pdf;   http://www.computer.
org/tvcg/tg1997/v0262abs.
htm.

Guo:2015:RUG

[GHL15] Hua Guo, Jeff Huang, and David H. Laidlaw. Representing uncertainty in graph edges: An evaluation of paired visual variables. IEEE Transactions on Visualization and Computer Graphics, 21(10): 1173–1186, October 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07089294-
abs.html.

Gramazio:2018:AAV

[GHL18] Connor C. Gramazio, Jeff Huang, and David H. Laidlaw. An analysis of automated visual analysis classification: Interactive visualization task inference of cancer genomics domain experts. IEEE Transactions on Visualization and Computer Graphics, 24(8): 2270–2283, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07999244-
abs.html.

Gansner:2013:MSM

[GHN13] Emden R. Gansner, Yifan Hu, and Stephen North. A maxent-stress model for graph layout. IEEE Transactions on Visualization and Computer Graphics, 19(6): 927–940, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2016:FTL

[GHP^{+}16] Hanqi Guo, Wenbin He, Tom Peterka, Han-Wei Shen, Scott M. Collis, and Jonathan J.

Helmus. Finite-time Lyapunov exponents and Lagrangian coherent structures in uncertain unsteady flows. IEEE Transactions on Visualization and Computer Graphics, 22(6):1672–1682, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2019:ESS

[GHS^{+}19] H. Guo, W. He, S. Seo, H. Shen, E. M. Constantinescu, C. Liu, and T. Peterka. Extreme-scale stochastic particle tracing for uncertain unsteady flow visualization and analysis. IEEE Transactions on Visualization and Computer Graphics, 25(9):2710– 2724, September 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gribble:2007:CGT

[GIK^{+}07] Christiaan P. Gribble, Thiago Ize, Andrew Kensler, Ingo Wald, and Steven G. Parker. A coherent grid traversal approach to visualizing particle-based simulation data. IEEE Transactions on Visualization and Computer Graphics, 13(4):758– 768, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grubert:2018:SCM

[GIMS18] Jens Grubert, Yuta Itoh, Kenneth Moser, and J. Edward Swan. A survey of calibration methods for optical seethrough head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 24(9):2649– 2662, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08052554.
pdf.

Gorla:2003:TSS

[GIS03] Gabriele Gorla, Victoria Interrante, and Guillermo Sapiro. Texture synthesis for 3D shape representation. IEEE Transactions on Visualization and Computer Graphics, 9(4):512–524, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0512abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0512.pdf.

Garth:2010:FME

[GJ10] Christoph Garth and Kenneth I. Joy. Fast, memoryefficient cell location in unstructured grids for visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1541– 1550, November/December 2010. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gomez:2017:FCG

[GJC^{+}17] Steven R. Gomez, Radu Jianu, Ryan Cabeen, Hua Guo, and David H. Laidlaw. Fauxvea: Crowdsourcing gaze location estimates for visualization analysis tasks. IEEE Transactions on Visualization and Computer Graphics, 23(2):1042– 1055, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/02/07414495-
                                        abs.html.

Gad:2015:TDS

[GJG^{+}15] S. Gad, W. Javed, S. Ghani, N. Elmqvist, T. Ewing, K. N. Hampton, and N. Ramakrishnan. ThemeDelta: Dynamic segmentations over temporal topic models. IEEE Transactions on Visualization and Computer Graphics, 21(5): 672–685, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2019:VPA

[GJG^{+}19] Shunan Guo, Zhuochen Jin, David Gotz, Fan Du, Hongyuan Zha, and Nan Cao. Visual progression analysis of event sequence data. IEEE Transactions on Visualization and Computer Graphics, 25(1): 417–426, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440811-
abs.html.

Gandy:2015:GEI

[GJK15] M. Gandy, S. Julier, and K. Kiyokawa. Guest Editor’s introduction to the special section on the International Symposium on Mixed and Augmented Reality 2013. IEEE Transactions on Visualization and Computer Graphics, 21(5):555–556, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gunther:2014:FME

[GJR^{+}14] David Gunther, Alec Jacobson, Jan Reininghaus, Hans-Peter Seidel, Olga Sorkine-Hornung, and Tino Weinkauf. Fast and memory-efficient topological denoising of 2D and 3D scalar fields. IEEE Transactions on Visualization and Computer Graphics, 20(12):2585–2594, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875939-
abs.html.

Gomez:2012:DSD

[GJZ^{+}12] Steven R. Gomez, Radu Jianu, Caroline Ziemkiewicz, Hua Guo, and David Laidlaw. Different strokes for different folks: Visual presentation design between disciplines. IEEE Transactions on Visualization and Computer Graphics, 18(12):2411– 2420, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gross:1995:VMS

[GK95] Markus H. Gross and Rolf Koch. Visualization of multidimensional shape and texture features in laser range data using complex-valued Gabor wavelets. IEEE Transactions on Visualization and Computer Graphics, 1(1): 44–59, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0044.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0044abs.
                                        htm.

Graham:2007:EMT

[GK07] Martin Graham and Jessie Kennedy. Exploring multiple trees through DAG representations. IEEE Transactions on Visualization and Computer Graphics, 13(6):1294– 1301, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2012:DFV

[GKK^{+}12] Attila Gyulassy, Natallia Kotava, Mark Kim, Charles Hansen, Hans Hagen, and Valerio Pascucci. Direct feature visualization using Morse– Smale complexes. IEEE Transactions on Visualization and Computer Graphics, 18(9):1549–1562, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ghani:2013:VAM

[GKL^{+}13] Sohaib Ghani, Bum Chul Kwon, Seungyoon Lee, Ji Soo Yi, and Niklas Elmqvist. Visual analytics for multimodal social network analysis: A design study with social scientists. IEEE Transactions on Visualization and Computer Graphics, 19(12):2032– 2041, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2016:III

[GKL^{+}16] A. Gyulassy, A. Knoll, K. C.

Lau, B. Wang, P. Bremer, M. E. Papka, L. A. Curtiss, and V. Pascucci. Interstitial and interlayer ion diffusion geometry extraction in graphitic nanosphere battery materials. IEEE Transactions on Visualization and Computer Graphics, 22(1):916–925, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grottel:2015:MPF

[GKM^{+}15] Sebastian Grottel, Michael Krone, Christoph Muller, Guido Reina, and Thomas Ertl. MegaMol — a prototyping framework for particlebased visualization. IEEE Transactions on Visualization and Computer Graphics, 21 (2):201–214, February 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/06881728-
abs.html.

Gansner:2005:TFV

[GKN05] Emden R. Gansner, Yehuda Koren, and Stephen C. North. Topological fisheye views for visualizing large graphs. IEEE Transactions on Visualization and Computer Graphics, 11(4):457– 468, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gandy:2014:GEI

[GKR14] Maribeth Gandy, Kiyoshi Kiyokawa, and Gerhard Reitmayr. Guest Editor’s introduction: Special section on the international symposium on mixed and augmented reality 2012. IEEE Transactions on Visualization and Computer Graphics, 20(6): 823–824, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Garth:2008:GAI

[GKT^{+}08] Christoph Garth, Han Krishnan, Xavier Tricoche, Tom Tricoche, and Kenneth I. Joy. Generation of accurate integral surfaces in timedependent vector fields. IEEE Transactions on Visualization and Computer Graphics, 14(6):1404–1411, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Garon:2017:DDT

[GL17] Mathieu Garon and Jean-Francois Lalonde. Deep 6-DOF tracking. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2410–2418, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007334-
abs.html.

Guo:2006:MTS

[GLB^{+}06] Xiaohu Guo, Xin Li, Yunfan Bao, Xianfeng Gu, and Hong Qin. Meshless thin-shell simulation based on global conformal parameterization. IEEE Transactions on Visualization and Computer Graphics, 12 (3):375–385, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gupta:2016:DYS

[GLB16] Kunal Gupta, Gun A. Lee, and Mark Billinghurst. Do you see what I see? The effect of gaze tracking on task space remote collaboration. IEEE Transactions on Visualization and Computer Graphics, 22 (11):2413–2422, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gleicher:2013:EEE

[Gle13] Michael Gleicher. Explainers: Expert explorations with crafted projections. IEEE Transactions on Visualization and Computer Graphics, 19(12):2042–2051, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gleicher:2018:CVC

[Gle18]  M. Gleicher.  Considera-

tions for visualizing comparison. IEEE Transactions on Visualization and Computer Graphics, 24(1):413–423, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gratzl:2013:LVA

[GLG^{+}13] Samuel Gratzl, Alexander Lex, Nils Gehlenborg, Hanspeter Pfister, and Marc Streit. LineUp: Visual analysis of multi-attribute rankings. IEEE Transactions on Visualization and Computer Graphics, 19(12):2277–2286, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glaber:2014:CVW

[GLH^{+}14] Sylvia Glaber, Kai Lawonn, Thomas Hoffmann, Martin Skalej, and Bernhard Preim. Combined visualization of wall thickness and wall shear stress for the evaluation of aneurysms. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2506–2515, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06877722-
abs.html.

Gorg:2013:CCA

[GLK^{+}13] C. Gorg, Zhicheng Liu, Jaeyeon Kihm, Jaegul Choo, Haesun Park, and J. Stasko. Combining computational analyses and interactive visualization for document exploration and sensemaking in Jigsaw. IEEE Transactions on Visualization and Computer Graphics, 19(10):1646– 1663, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Govindaraju:2006:FRC

[GLM06] Naga K. Govindaraju, Ming C.

Lin, and Dinesh Manocha.

Fast and reliable collision culling using graphics hardware. IEEE Transactions on Visualization and Computer Graphics, 12(2):143– 154, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gaffary:2017:AFS

[GLM^{+}17] Yoren Gaffary, Benoit Le Gouis, Maud Marchal, Ferran Argelaguet, Bruno Arnaldi, and Anatole Lecuyer. AR feels “softer” than VR: Haptic perception of stiffness in augmented versus virtual reality. IEEE Transactions on Visualization and Computer Graphics, 23(11):2372– 2377, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007246-
abs.html.

Gao:2013:VDM

[GLRH13] Yue Gao, Chen-Feng Li, Bo Ren, and Shi-Min Hu. View-dependent multiscale fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 19(2): 178–188, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gramazio:2017:CCD

[GLS17] Connor C. Gramazio, David H. Laidlaw, and Karen B.

Schloss. Colorgorical: Creating discriminable and preferable color palettes for information visualization. IEEE Transactions on Visualization and Computer Graphics, 23 (1):521–530, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gasteiger:2012:ADV

[GLvP^{+}12] Rocco Gasteiger, Dirk J.

Lehmann, Roy van Pelt, Gabor Janiga, Oliver Beuing, Anna Vilanova, Holger Theisel, and Bernhard Preim. Automatic detection and visualization of qualitative hemodynamic characteristics in cerebral aneurysms. IEEE Transactions on Visualization and Computer Graphics, 18(12):2178–2187, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2017:UDP

[GLX17] Yulong Guo, Xiaopei Liu, and Xuemiao Xu. A unified detailpreserving liquid simulation by two-phase lattice Boltzmann modeling. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1479–1491, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07414526-
abs.html.

Guo:2018:THQ

[GLX^{+}18] Yulong Guo, Xiaopei Liu, Chi Xiong, Xuemiao Xu, and Chi-Wing Fu. Towards highquality visualization of superfluid vortices. IEEE Transactions on Visualization and Computer Graphics, 24(8): 2440–2455, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/08/07959130-
                                        abs.html.

Grubert:2017:TPA

[GLZR17] Jens Grubert, Tobias Langlotz, Stefanie Zollmann, and Holger Regenbrecht. Towards pervasive augmented reality: Context-awareness in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1706–1724, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/06/07435333-
                                        abs.html.

Gain:2005:WS

[GM05] James Gain and Patrick Marais. Warp sculpting. IEEE Transactions on Visualization and Computer Graphics, 11(2):217– 227, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glondu:2013:RTS

[GMD13] L. Glondu, M. Marchal, and G. Dumont. Real-time simulation of brittle fracture using modal analysis. IEEE Transactions on Visualization and Computer Graphics, 19 (2):201–209, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gao:2016:SVD

[GMD^{+}16] Xifeng Gao, Tobias Martin, Sai Deng, Elaine Cohen, Zhigang Deng, and Guoning Chen. Structured volume decomposition via generalized sweeping. IEEE Transactions on Visualization and Computer Graphics, 22(7):1899– 1911, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Goodwin:2017:WDC

[GMD^{+}17] Sarah Goodwin, Christopher Mears, Tim Dwyer, Maria Garcia de la Banda, Guido Tack, and Mark Wallace. What do constraint programming users want to see? Exploring the role of visualisation in profiling of models and search. IEEE Transactions on Visualization and Computer Graphics, 23(1): 281–290, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Groller:2005:GEI

[GMM05] Eduard Gröller, Kwan-Liu Ma, and Klaus Mueller. Guest Editors’ introduction: Special section on IEEE visualization applications. IEEE Transactions on Visualization and Computer Graphics, 11(5):483–484, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/05/v0483.
pdf.

Giesen:2007:CAM

[GMS^{+}07] Joachim Giesen, Klaus Mueller, Eva Schuberth, Lujin Wang, and Peter Zolliker. Conjoint analysis to measure the perceived quality in volume rendering. IEEE Transactions on Visualization and Computer Graphics, 13(6):1664– 1671, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gabbard:2019:EAD

[GMS19] Joseph L. Gabbard, Divya Gupta Mehra, and J. Edward Swan. Effects of AR display context switching and focal distance switching on human performance. IEEE Transactions on Visualization and Computer Graphics, 25(6):2228–2241, June 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8353823/.

Guo:2011:WWY

[GMY11] Hanqi Guo, Ningyu Mao, and Xiaoru Yuan. WYSI-WYG (what you see is what you get) volume visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2106– 2114, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guntury:2012:RDS

[GN12] Sashidhar Guntury and P. J.

Narayanan. Raytracing dynamic scenes on the GPU using Grids. IEEE Transactions on Visualization and Computer Graphics, 18(1):5– 16, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gasteiger:2011:FFC

[GNBP11] Rocco Gasteiger, Mathias Neugebauer, Oliver Beuing, and Bernhard Preim. The FLOWLENS: a focusand-context visualization approach for exploration of blood flow in cerebral aneurysms. IEEE Transactions on Visualization and Computer Graphics, 17(12):2183–2192, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gomez-Nieto:2016:DMR

[GNCM^{+}16] E. Gomez-Nieto, W. Casaca, D. Motta, I. Hartmann, G. Taubin, and L. G. Nonato. Dealing with multiple requirements in geometric arrangements. IEEE Transactions on Visualization and Computer Graphics, 22(3):1223– 1235, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glueck:2018:PPC

[GNDV^{+}18] M. Glueck, M. P. Naeini, F. Doshi-Velez, F. Chevalier, A. Khan, D. Wigdor, and M. Brudno. PhenoLines: Phenotype comparison visualizations for disease subtyping via topic models. IEEE Transactions on Visualization and Computer Graphics, 24(1): 371–381, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2006:TAS

[GNP^{+}06] Attila Gyulassy, Vijay Natarajan, Valerio Pascucci, Peer-Timo Bremer, and Bernd Hamann. A topological approach to simplification of three-dimensional scalar functions. IEEE Transactions on Visualization and Computer Graphics, 12(4):474– 484, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gyulassy:2007:ECM

[GNPH07] Attila Gyulassy, Vijay Natarajan, Valerio Pascucci, and Bernd Hamann. Efficient computation of Morse– Smale complexes for threedimensional scalar functions. IEEE Transactions on Visualization and Computer Graphics, 13(6):1440–1447, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gomez-Nieto:2014:SPS

[GNSP^{+}14] Erick Gomez-Nieto, Frizzi San Roman, Paulo Pagliosa, Wallace Casaca, Elias S. Helou, Maria Cristina F. de Oliveira, and Luis Gustavo Nonato. Similarity preserving snippetbased visualization of Web search results. IEEE Transactions on Visualization and Computer Graphics, 20(3): 457–470, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Griffin:2015:ETC

[GO15] Wesley Griffin and Marc Olano. Evaluating texture compression masking effects using objective image quality assessment metrics. IEEE Transactions on Visualization and Computer Graphics, 21(8):970–979, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/07101291-abs.html.

Gordon:2002:FCA

[Gor02] Dan Gordon. The floating column algorithm for shaded, parallel display of function surfaces without patches. IEEE Transactions on Visualization and Computer Graphics, 8(1):76–91, January 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0076.
pdf;   http://www.computer.
org/tvcg/tg2002/v0076abs.
htm.

Govyadinov:2018:E

[Gov18] Pavel A. Govyadinov. Erratum. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2264, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/08365872.
pdf.

Grosset:2017:TTT

[GPC^{+}17] A. V. Pascal Grosset, Manasa Prasad, Cameron Christensen, Aaron Knoll, and Charles Hansen. TOD-tree: Task-overlapped direct send tree image compositing for hybrid MPI parallelism and GPUs. IEEE Transactions on Visualization and Computer Graphics, 23(6):1677– 1690, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07433468-
abs.html.

Grechkin:2014:DAE

[GPK14] Timofey Y. Grechkin, Jodie M.

Plumert, and Joseph K. Kearney. Dynamic affordances in embodied interactive systems: The role of display and mode of locomotion. IEEE Transactions on Visualization and Computer Graphics, 20 (4):596–605, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Geng:2011:AHF

[GPL^{+}11] Zhao Geng, ZhenMin Peng, Robert S. Laramee, Jonathan C. Roberts, and Rick Walker. Angular histograms: Frequencybased visualizations for large, high dimensional data. IEEE Transactions on Visualization and Computer Graphics, 17(12):2572–2580, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gebhardt:2013:EPM

[GPL^{+}13] Sascha Gebhardt, Sebastian Pick, Franziska Leithold, Bernd Hentschel, and Torsten Kuhlen. Extended pie menus for immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 19(4): 644–651, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2016:ETV

[GPP^{+}16] H. Guo, C. L. Phillips, T. Peterka, D. Karpeyev, and A. Glatz. Extracting, tracking, and visualizing magnetic flux vortices in 3D complexvalued superconductor simulation data. IEEE Transactions on Visualization and Computer Graphics, 22(1): 827–836, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Garcke:2001:PFM

[GPR^{+}01] Harald Garcke, Tobias Preußer, Martin Rumpf, Alexandru C. Telea, Ulrich Weikard, and Jarke J. van Wijk. A phase field model for continuous clustering on vector fields. IEEE Transactions on Visualization and Computer Graphics, 7(3):230–241, July 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0230.
pdf;   http://www.computer.

org/tvcg/tg2001/v0230abs.
htm.

Guo:2017:RTT

[GQGP17] Jie Guo, Jinghui Qian, Yanwen Guo, and Jingui Pan. Rendering thin transparent layers with extended normal distribution functions. IEEE Transactions on Visualization and Computer Graphics, 23(9):2108– 2119, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07590146-
abs.html.

Gu:2018:EQS

[GQM^{+}18] Ke Gu, Junfei Qiao, Xiongkuo Min, Guanghui Yue, Weisi Lin, and Daniel Thalmann. Evaluating quality of screen content images via structural variation analysis. IEEE Transactions on Visualization and Computer Graphics, 24(10):2689–2701, October 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL https://www.computer.
org/csdl/trans/tg/2018/
10/08100977-abs.html.

Grigoryan:2004:PBP

[GR04] Gevorg Grigoryan and Penny Rheingans. Point-based probabilistic surfaces to show surface uncertainty. IEEE Transactions on Visualization and Computer Graphics, 10(5):564–573, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0564.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0564.pdf.

Griffin:2015:CCL

[GR15] Amy L. Griffin and Anthony C. Robinson. Comparing color and leader line highlighting strategies in coordinated view geovisualizations. IEEE Transactions on Visualization and Computer Graphics, 21(3):339– 349, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06965627-
abs.html.

Groller:1995:MVK

[GRS95] Eduard Gröller, René T. Rau, and Wolfgang Straßer. Modeling and visualization of knitwear. IEEE Transactions on Visualization and Computer Graphics, 1(4):302– 310, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0302.
pdf;   http://www.computer.
org/tvcg/tg1995/v0302abs.
htm.

Goodall:2019:SIE

[GRS^{+}19] John R. Goodall, Eric D. Ragan, Chad A. Steed, Joel W. Reed, G. David Richardson, Kelly M. T. Huffer, Robert A. Bridges, and Jason A. Laska. Situ: Identifying and explaining suspicious behavior in networks. IEEE Transactions on Visualization and Computer Graphics, 25(1):204– 214, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440825-
abs.html.

Gerrits:2017:GGS

[GRT17] Tim Gerrits, Christian Rössl, and Holger Theisel. Glyphs for general second-order 2D and 3D tensors. IEEE Transactions on Visualization and Computer Graphics, 23(1): 980–989, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grottel:2007:VVA

[GRVE07] Sebastian Grottel, Guido Reina, Jadran Vrabec, and Thomas Ertl. Visual verification and analysis of cluster detection for molecular dynamics. IEEE Transactions on Visualization and Computer Graphics, 13(6):1624– 1631, November/December 2007. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Groth:2006:PAV

[GS06] Dennis P. Groth and Kristy Streefkerk. Provenance and annotation for visual exploration systems. IEEE Transactions on Visualization and Computer Graphics, 12(6):1500–1510, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gabbard:2008:UEA

[GS08] Joe L. Gabbard and J. E. Swan II. Usability engineering for augmented reality: Employing user-based studies to inform design. IEEE Transactions on Visualization and Computer Graphics, 14 (3):513–525, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gotz:2014:DVA

[GS14] David Gotz and Harry Stavropoulos. DecisionFlow: Visual analytics for highdimensional temporal event sequence data. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1783–1792, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/

csdl/trans/tg/2014/12/06875996-
abs.html.

Greunke:2016:TIV

[GS16] L. Greunke and A. Sadagic.

Taking immersive VR leap in training of landing signal officers. IEEE Transactions on Visualization and Computer Graphics, 22(4):1482– 1491, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gayle:2009:INH

[GSA^{+}09] Russell Gayle, Avneesh Sud, Erik Andersen, Stephen J. Guy, Ming C. Lin, and Dinesh Manocha. Interactive navigation of heterogeneous agents using adaptive roadmaps. IEEE Transactions on Visualization and Computer Graphics, 15(1):34– 48, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glocker:2015:RTR

[GSCI15] B. Glocker, J. Shotton, A. Criminisi, and S. Izadi. Real-time RGB-D camera relocalization via randomized ferns for keyframe encoding. IEEE Transactions on Visualization and Computer Graphics, 21(5):571–583, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gal:2007:POS

[GSCO07] Ran Gal, Ariel Shamir, and Daniel Cohen-Or. Poseoblivious shape signature. IEEE Transactions on Visualization and Computer Graphics, 13(2):261–271, March/ April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gregorski:2004:AET

[GSDJ04] Benjamin Gregorski, Joshua Senecal, Mark A. Duchaineau, and Kenneth I. Joy. Adaptive extraction of timevarying isosurfaces. IEEE Transactions on Visualization and Computer Graphics, 10(6):683–694, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0683abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0683.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0683.pdf.

Gross:1996:ETS

[GSG96] Markus H. Gross, Oliver G. Staadt, and Roger Gatti. Efficient triangular surface approximations using wavelets and quadtree data structures. IEEE Transactions on Visualization and Computer Graphics, 2(2):130–143, June 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0130.
pdf;   http://www.computer.
org/tvcg/tg1996/v0130abs.
htm.

Gramazio:2014:RBV

[GSL14] Connor C. Gramazio, Karen B.

Schloss, and David H. Laidlaw. The relation between visualization size, grouping, and user performance. IEEE Transactions on Visualization and Computer Graphics, 20(12):1953–1962, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875989-
abs.html.

Goethem:2017:MGT

[GSL^{+}17] Arthur Van Goethem, Frank Staals, Maarten Löffler, Jason Dykes, and Bettina Speckmann. Multi-granular trend detection for time-series analysis. IEEE Transactions on Visualization and Computer Graphics, 23(1):661–670, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Glondu:2014:FCD

[GSM^{+}14] Loeiz Glondu, Sara C.

Schvartzman, Maud Marchal, Georges Dumont, and Miguel A. Otaduy. Fast collision detection for fracturing rigid bodies. IEEE Transactions on Visualization and Computer Graphics, 20(1): 30–41, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2008:MGO

[GSPJ08] Yanwen Guo, Hanqiu Sun, Qunsheng Peng, and Zhongding Jiang. Mesh-guided optimized retexturing for image and video. IEEE Transactions on Visualization and Computer Graphics, 14(2):426– 439, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gast:2015:OIL

[GSS^{+}15] Theodore F. Gast, Craig Schroeder, Alexey Stomakhin, Chenfanfu Jiang, and Joseph M. Teran. Optimization integrator for large time steps. IEEE Transactions on Visualization and Computer Graphics, 21(10):1103– 1115, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/10/07164346-
                                        abs.html.

Gortler:2019:SSF

[GSS^{+}19] Jochen Görtler, Marc Spicker, Christoph Schulz, Daniel Weiskopf, and Oliver Deussen. Stippling of 2D scalar fields.

IEEE Transactions on Visualization and Computer Graphics, 25(6):2193–2204, June 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL https://ieeexplore.
ieee.org/document/8667696/

.

Gunther:2016:RIV

[GST16] T. Gunther, M. Schulze, and H. Theisel. Rotation invariant vortices for flow visualization. IEEE Transactions on Visualization and Computer Graphics, 22(1):817–826, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gortler:2018:BTU

[GSWD18] J. Görtler, C. Schulz, D. Weiskopf, and O. Deussen. Bubble treemaps for uncertainty visualization. IEEE Transactions on Visualization and Computer Graphics, 24(1): 719–728, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gurijala:2013:CFU

[GSZ^{+}13] Krishna Chaitanya Gurijala, Rui Shi, Wei Zeng, Xianfeng Gu, and Arie Kaufman. Colon flattening using heat diffusion Riemannian metric. IEEE Transactions on Visualization and Computer Graphics, 19(12):2848– 2857, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gunther:2014:VCI

[GT14] Tobias Gunther and Holger Theisel. Vortex cores of inertial particles. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2535–2544, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875993-
abs.html.

Gunther:2017:BFT

[GT17] Tobias Günther and Holger Theisel. Backward finite-time Lyapunov exponents in inertial flows. IEEE Transactions on Visualization and Computer Graphics, 23(1): 970–979, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gunther:2019:OVC

[GT19] Tobias Gunther and Holger Theisel. Objective vortex corelines of finite-sized objects in fluid flows. IEEE Transactions on Visualization and Computer Graphics, 25 (1):956–966, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440096-
abs.html.

Gueziec:2001:CSC

[GTLH01] A. Guéziec, G. Taubin, F. Lazarus, and B. Horn. Cutting and stitching: Converting sets of polygons to manifold surfaces. IEEE Transactions on Visualization and Computer Graphics, 7(2):136– ??, April/June 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0136.
pdf;   http://www.computer.
org/tvcg/tg2001/v0136abs.
htm.

Gogolou:2019:CSP

[GTPB19] Anna Gogolou, Theophanis Tsandilas, Themis Palpanas, and Anastasia Bezerianos. Comparing similarity perception in time series visualizations. IEEE Transactions on Visualization and Computer Graphics, 25(1):523– 533, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440826-
abs.html.

Grammel:2010:HIV

[GTS10] Lars Grammel, Melanie Tory, and Margaret-Anne Storey. How information visualization novices construct visualizations. IEEE Transactions on Visualization and Computer Graphics, 16(6):943– 952, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Grammel:2011:EHI

[GTS11] Lars Grammel, Melanie Tory, and Margaret-Anne Storey. Erratum to “How Information Visualization Novices Construct Visualizations”. IEEE Transactions on Visualization and Computer Graphics, 17 (2):260, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gueziec:1999:LTS

[Gué99] A. Guéziec. Locally toleranced surface simplification. IEEE Transactions on Visualization and Computer Graphics, 5(2):168– 189, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0168.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0168abs.
                                        htm.

Gueziec:2001:MDP

[Gué01]  A. P. Guéziec.  “Mesh-

sweeper”: Dynamic point-to-polygonal-mesh distance and applications. IEEE Transactions on Visualization and Computer Graphics, 7(1):47– ??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0047.
pdf;   http://www.computer.
org/tvcg/tg2001/v0047abs.
htm.

Gattullo:2015:ETO

[GUFM15] M. Gattullo, A. E. Uva, M. Fiorentino, and G. Monno. Effect of text outline and contrast polarity on AR text readability in industrial lighting. IEEE Transactions on Visualization and Computer Graphics, 21(5):638–651, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:1995:MMS

[Guo95] Baining Guo. A multiscale model for structure-based volume rendering. IEEE Transactions on Visualization and Computer Graphics, 1(4):291– 301, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0291.
pdf;   http://www.computer.
org/tvcg/tg1995/v0291abs.
htm.

Grevera:2000:OMF

[GUO00] G. J. Grevera, J. K. Udupa, and D. Odhner. An order of magnitude faster isosurface rendering in software on a PC than using dedicated, general purpose rendering hardware. IEEE Transactions on Visualization and Computer Graphics, 6(4):335– ??, October/December 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0335.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0335abs.
                                        htm.

Guo:2009:FMM

[Guo09] Diansheng Guo. Flow mapping and multivariate visualization of large spatial interaction data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1041– 1048, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Georgii:2006:GSP

[GW06] Joachim Georgii and Rüdiger Westermann. A generic and scalable pipeline for GPU tetrahedral grid rendering. IEEE Transactions on Visualization and Computer Graphics, 12(5):1345–1352, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gu:2011:THE

[GW11] Yi Gu and Chaoli Wang. TransGraph: Hierarchical exploration of transition relationships in time-varying volumetric data. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2015–2024, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Garland:2013:GEI

[GW13] Michael Garland and Rui Wang. Guest Editors’ introduction: Special section on the ACM SIGGRAPH Symposium on Interactive 3D Graphics and Games (I3D 2012). IEEE Transactions on Visualization and Computer Graphics, 19(5):721–722, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Griffin:2012:RTG

[GWBO12] Wesley Griffin, Yu Wang, David Berrios, and Marc Olano. Real-time GPU surface curvature estimation on deforming meshes and volumetric data sets. IEEE Transactions on Visualization and Computer Graphics, 18(10):1603–1613, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Govyadinov:2019:RTV

[GWE^{+}19] Pavel A. Govyadinov, Tasha Womack, Jason L. Eriksen, Guoning Chen, and David Mayerich. Robust tracing and visualization of heterogeneous microvascular networks. IEEE Transactions on Visualization and Computer Graphics, 25(4):1760– 1773, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8326724/.

Goffin:2014:EPD

[GWFI14] Pascal Goffin, Wesley Willett, Jean-Daniel Fekete, and Petra Isenberg. Exploring the placement and design of word-scale visualizations. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2291–2300, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875917-
                                        abs.html.

Gurijala:2012:CHD

[GWK12] Krishna Chaitanya Gurijala, Lei Wang, and Arie Kaufman. Cumulative heat diffusion using volume gradient operator for volume analysis. IEEE Transactions on Visualization and Computer Graphics, 18(12):2069– 2077, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gu:2016:MGU

[GWP^{+}16] Y. Gu, C. Wang, T. Peterka, R. Jacob, and S. H. Kim. Mining graphs for understanding time-varying volumetric data. IEEE Transactions on Visualization and Computer Graphics, 22(1):965–974, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ghosh:2018:NEI

[GWP^{+}18] Sarthak Ghosh, Lauren Winston, Nishant Panchal, Philippe Kimura-Thollander, Jeff Hotnog, Douglas Cheong, Gabriel Reyes, and Gregory D. Abowd. NotifiVR: Exploring interruptions and notifications in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 24(4):1447–1456, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260856-
abs.html.

Garcia:2013:IAS

[GXH^{+}13] Ismael Garcia, Jiazhi Xia, Ying He, Shi-Qing Xin, and Gustavo Patow. Interactive applications for sketch-based editable polycube map. IEEE Transactions on Visualization and Computer Graphics, 19(7):1158–1171, July 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2018:ERN

[GXW^{+}18a] Kaiwen Guo, Feng Xu, Yangang Wang, Yebin Liu, and Qionghai Dai. Errata to “Robust Non-Rigid Motion Tracking and Surface Reconstruction Using L_{0} Regularization. IEEE Transactions on Visualization and Computer Graphics, 24(7):2268, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/08365924.

pdf. See [GXW^{+}18b].

Guo:2018:RNR

[GXW^{+}18b] Kaiwen Guo, Feng Xu, Yangang Wang, Yebin Liu, and Qionghai Dai. Robust nonrigid motion tracking and surface reconstruction using L_{0} regularization. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1770–1783, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07888591-

abs.html. See erratum [GXW^{+}18a].

Guo:2012:SMV

[GXY12] Hanqi Guo, He Xiao, and Xiaoru Yuan. Scalable multivariate volume visualization and analysis based on dimension projection and parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1397–1410, September 2012.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2018:EVS

[GXZ^{+}18] S. Guo, K. Xu, R. Zhao, D. Gotz, H. Zha, and N. Cao. EventThread: Visual summarization and stage analysis of event sequence data. IEEE Transactions on Visualization and Computer Graphics, 24 (1):56–65, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gudukbay:2002:SVD

[GY02] Ugur Güdükbay and Türker Yilmaz. Stereoscopic viewdependent visualization of terrain height fields. IEEE Transactions on Visualization and Computer Graphics, 8(4):330–345, October/ December 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0330abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0330.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0330.pdf.

Guo:2016:ELF

[GYK^{+}16] Xinqing Guo, Zhan Yu, Sing Bing Kang, Haiting Lin, and Jingyi Yu. Enhancing light fields through rayspace stitching. IEEE Transactions on Visualization and Computer Graphics, 22(7): 1852–1861, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gou:2011:TRP

[GZ11] Liang Gou and Xiaolong (Luke) Zhang. TreeNetViz: Revealing patterns of networks over tree structures. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2449–2458, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2014:ODF

[GZ14] Diansheng Guo and Xi Zhu. Origin-destination flow data smoothing and mapping. IEEE Transactions on Visualization and Computer Graphics, 20(12):2043–2052, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875983-
                                        abs.html.

Guo:2014:ABS

[GZL^{+}14] Hanqi Guo, Jiang Zhang, Richen Liu, Lu Liu, Xiaoru Yuan, Jian Huang, Xiangfei Meng, and Jingshan Pan. Advection-based sparse data management for visualizing unsteady flow. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2555–2564, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875971-
abs.html.

Helgeland:2004:VVF

[HA04] Anders Helgeland and Oyvind Andreassen. Visualization of vector fields using seed LIC and volume rendering. IEEE Transactions on Visualization and Computer Graphics, 10(6):673–682, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0673abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0673.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0673.pdf.

Heer:2006:MSB

[HA06a] Jeffrey Heer and Maneesh Agrawala. Multi-scale banking to 45 degrees. IEEE Transactions on Visualization and Computer Graphics, 12(5):701–708, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heer:2006:SDP

[HA06b] Jeffrey Heer and Maneesh Agrawala. Software design patterns for information visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5): 853–860, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

He:2017:VCB

[HA17] Shiqing He and Eytan Adar. VizItCards: A card-based toolkit for Infovis design education. IEEE Transactions on Visualization and Computer Graphics, 23(1):561–570, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Harper:2018:CBD

[HA18] Jonathan Harper and Maneesh Agrawala. Converting basic D3 charts into reusable style templates. IEEE Transactions on Visualization and Computer Graphics, 24(3): 1274–1286, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hadwiger:2018:SEE

[HAAB^{+}18] M. Hadwiger, A. K. Al-Awami, J. Beyer, M. Agus, and H. Pfister. SparseLeap: Efficient empty space skipping for large-scale volume rendering. IEEE Transactions on Visualization and Computer Graphics, 24(1):974–983, January 2018. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hagen:1998:PSI

[Hag98] H. Hagen. Preface: Special issue on visualization. IEEE Transactions on Visualization and Computer Graphics, 4(2): 97, April/June 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0097.
pdf.

Hagen:1999:EN

[Hag99] Hans Hagen. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 5(3):193–??, July/ September 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0193.
pdf.

Hagen:2000:EN

[Hag00] H. Hagen. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 6(3):193–??, July/ September 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0193.
pdf.

Hagen:2003:ECF

[Hag03] Hans Hagen. Editor-in-Chief’s farewell. IEEE Transactions on Visualization and Computer Graphics, 9(1):1, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/01/v0001.
                                        pdf;   http://csdl.computer.
                                        org/dl/trans/tg/2003/01/
                                        v0001.htm.

Hartl:2016:EVH

[HAGS16] Andreas Daniel Hartl, Clemens Arth, Jens Grubert, and Dieter Schmalstieg. Efficient verification of holograms using mobile augmented reality. IEEE Transactions on Visualization and Computer Graphics, 22(7):1843–1851, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hossain:2011:THQ

[HAM11] Zahid Hossain, Usman R. Alim, and Torsten Moller. Toward high-quality gradient estimation on regular lattices. IEEE Transactions on Visualization and Computer Graphics, 17(4): 426–439, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hanrahan:1995:MDW

[Han95] Pat Hanrahan. In memoriam: Dr. Wolfgang Krueger. IEEE Transactions on Visualization and Computer Graphics, 1(3):209, September 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0209.
pdf.

Hansen:2016:SCV

[Han16] Charles D. Hansen. A survey of colormaps in visualization. IEEE Transactions on Visualization and Computer Graphics, 22(8):2051–2069, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hansen:2018:VCA

[Han18] C. Hansen. The 2017 Visualization Career Award. IEEE Transactions on Visualization and Computer Graphics, 24 (1):xxvi, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hart:2016:FCP

[Har16] John C. Hart. Fast coherent particle advection through time-varying unstructured flow datasets. IEEE Transactions on Visualization and Computer Graphics, 22(8): 1959–1972, August 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hullman:2011:BIV

[HAS11] Jessica Hullman, Eytan Adar, and Priti Shah. Benefitting InfoVis with visual difficulties. IEEE Transactions on Visualization and Computer Graphics, 17(12):2213– 2222, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haker:2000:CSP

[HAT^{+}00] S. Haker, S. Angenent, A. Tannenbaum, R. Kikinis, G. Sapiro, and M. Halle. Conformal surface parameterization for texture mapping. IEEE Transactions on Visualization and Computer Graphics, 6(2):181– 189, April/June 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0181.
pdf;   http://www.computer.
org/tvcg/tg2000/v0181abs.
htm.

Hausner:1997:MEL

[Hau97] A. Hausner. Multipole expansion of the light vector. IEEE Transactions on Visualization and Computer Graphics, 3(1): 12–22, January/March 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0012.
pdf;   http://www.computer.

org/tvcg/tg1997/v0012abs.
htm.

Hahmann:2003:PSI

[HB03] Stefanie Hahmann and Georges-Pierre Bonneau. Polynomial surfaces interpolating arbitrary triangulations. IEEE Transactions on Visualization and Computer Graphics, 9(1): 99–109, January/March 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0099abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0099.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/01/v0099.pdf.

Heer:2010:DLD

[HB10] Jeffrey Heer and Michael Bostock. Declarative language design for interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1149– 1156, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hodgson:2013:CFA

[HB13] Eric Hodgson and Eric Bachmann. Comparing four approaches to generalized redirected walking: Simulation and live user data. IEEE Transactions on Visualization and Computer Graphics, 19 (4):634–643, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Herling:2014:HQR

[HB14] Jan Herling and Wolfgang Broll. High-quality real-time video inpainting with PixMix. IEEE Transactions on Visualization and Computer Graphics, 20(6):866–879, June 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Herrera:2014:SRC

[HBAB14] Imanol Herrera, Carlos Buchart, Iker Aguinaga, and Diego Borro. Study of a ray casting technique for the visualization of deformable volumes. IEEE Transactions on Visualization and Computer Graphics, 20(11):1555– 1565, November 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06851204-
abs.html.

Howison:2012:HPV

[HBC12] Mark Howison, E. Wes Bethel, and Hank Childs. Hybrid parallelism for volume rendering on large-, multi-, and many-core systems. IEEE Transactions on Visualization and Computer Graphics, 18 (1):17–29, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hagbi:2011:SRP

[HBESB11] Nate Hagbi, Oriel Bergig, Jihad El-Sana, and Mark Billinghurst. Shape recognition and pose estimation for mobile augmented reality. IEEE Transactions on Visualization and Computer Graphics, 17(10):1369–1379, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Henry:2008:IRC

[HBF08] Nathalie Henry, Anastasia Bezerianos, and Jean-Daniel Fekete. Improving the readability of clustered social networks using node duplication. IEEE Transactions on Visualization and Computer Graphics, 14(6):1317–1324, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haidacher:2011:VAU

[HBG11] Martin Haidacher, Stefan Bruckner, and Eduard Gröller. Volume analysis using multimodal surface similarity. IEEE Transactions on Visualization and Computer Graphics, 17(12):1969–1978, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hadwiger:2012:IVE

[HBJP12] Markus Hadwiger, Johanna Beyer, Won-Ki Jeong, and Hanspeter Pfister. Interactive volume exploration of petascale microscopy data streams using a visualizationdriven virtual memory approach. IEEE Transactions on Visualization and Computer Graphics, 18(12):2285– 2294, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Harders:2009:CRS

[HBKS09] Matthias Harders, Gérald Bianchi, Benjamin Knoerlein, and Gábor Székely. Calibration, registration, and synchronization for high precision augmented reality haptics. IEEE Transactions on Visualization and Computer Graphics, 15(1):138–149, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2013:SDM

[HBM^{+}13] Xinran Hu, Lauren Bradel, Dipayan Maiti, Leanna House, Chris North, and Scotland Leman. Semantics of directly manipulating spatializations. IEEE Transactions on Visualization and Computer Graphics, 19(12):2052– 2059, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hodgson:2014:PRW

[HBT14] Eric Hodgson, Eric Bachmann, and Tyler Thrash. Performance of redirected walking algorithms in a constrained virtual world. IEEE Transactions on Visualization and Computer Graphics, 20 (4):579–587, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

House:2006:APO

[HBW06] Donald H. House, Alethea S.

Bair, and Colin Ware. An approach to the perceptual optimization of complex visualizations. IEEE Transactions on Visualization and Computer Graphics, 12(4):509– 521, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hlawatsch:2014:VAL

[HBW14] Marcel Hlawatsch, Michael Burch, and Daniel Weiskopf. Visual adjacency lists for dynamic graphs. IEEE Transactions on Visualization and Computer Graphics, 20(11): 1590–1603, November 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06812198-
abs.html.

Huang:2005:ITR

[HC05] Jianbing Huang and Michael B. Carter. Interactive transparency rendering for large CAD models. IEEE Transactions on Visualization and Computer Graphics, 11(5): 584–595, September/October 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2015:HMC

[HCMTH15] Junhui Hou, Lap-Pui Chau, Nadia Magnenat-Thalmann, and Ying He. Human motion capture data tailored transform coding. IEEE Transactions on Visualization and Computer Graphics, 21(7):848–859, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        07/07042272-abs.html.

Haouchine:2015:IST

[HCP^{+}15] N. Haouchine, S. Cotin, I. Peterlik, J. Dequidt, M. S. Lopez, E. Kerrien, and M.-O. Berger. Impact of soft tissue heterogeneity on augmented reality for liver surgery. IEEE Transactions on Visualization and Computer Graphics, 21 (5):584–597, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2016:SMS

[HCP^{+}16] W. Hu, Z. Chen, H. Pan, Y. Yu, E. Grinspun, and W. Wang. Surface mosaic synthesis with irregular tiles. IEEE Transactions on Visualization and Computer Graphics, 22(3):1302– 1313, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hall:2007:RNP

[HCS^{+}07] Peter M. Hall, John P. Collomosse, Yi-Zhe Song, Peiyi Shen, and Chuan Li. RTcams: a new perspective on nonphotorealistic rendering from photographs. IEEE Transactions on Visualization and Computer Graphics, 13(5): 966–979, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hullman:2011:VRF

[HD11] Jessica Hullman and Nick Diakopoulos. Visualization rhetoric: Framing effects in narrative visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2231–2240, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Healey:2012:IDN

[HD12] Christopher G. Healey and Brent M. Dennis. Interest driven navigation in visualization. IEEE Transactions on Visualization and Computer Graphics, 18(10):1744– 1756, October 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haouchine:2015:MRA

[HDBC15] N. Haouchine, J. Dequidt, M. Berger, and S. Cotin. Monocular 3D reconstruction and augmentation of elastic surfaces with self-occlusion handling. IEEE Transactions on Visualization and Computer Graphics, 21(12):1363– 1376, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hwa:2005:RTO

[HDJ05] Lok M. Hwa, Mark A. Duchaineau, and Kenneth I. Joy. Real-time optimal adaptation for planetary geometry and texture: 4-8 tile hierarchies. IEEE Transactions on Visualization and Computer Graphics, 11(4):355– 368, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hullman:2013:DUS

[HDR^{+}13] Jessica Hullman, Steven Drucker, Nathalie Henry Riche, Bongshin Lee, Danyel Fisher, and Eytan Adar. A deeper understanding of sequence in narrative visualization. IEEE Transactions on Visualization and Computer Graphics, 19(12):2406– 2415, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hazarika:2019:CFC

[HDSC19] Subhashis Hazarika, Soumya Dutta, Han-Wei Shen, and Jen-Ping Chen. CoDDA: A flexible copula-based distribution driven analysis framework for large-scale multivariate data. IEEE Transactions on Visualization and Computer Graphics, 25(1):1214– 1224, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440043-
abs.html.

Healey:1999:LDG

[HE99] C. G. Healey and J. T.

Enns. Large datasets at a glance: Combining textures and colors in scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 5(2):145– 167, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0145.
pdf;   http://www.computer.
org/tvcg/tg1999/v0145abs.
htm.

Hagen:2002:EN

[HE02] H. Hagen and D. S. Ebert. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 8(2): 97–98, April 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0097.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0097abs.
                                        htm.

Helgeland:2006:HQI

[HE06] Anders Helgeland and Thomas Elboth. High-quality and interactive animations of 3D time-varying vector fields. IEEE Transactions on Visualization and Computer Graphics, 12(6):1535–1546, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Healey:2012:AVM

[HE12] Christopher G. Healey and James T. Enns. Attention and visual memory in visualization and computer graphics. IEEE Transactions on Visualization and Computer Graphics, 18(7): 1170–1188, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heer:2018:VTA

[Hee18] J. Heer. The 2017 Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 24(1):xxvii– xxviii, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hurter:2014:BVD

[HEF^{+}14] Christophe Hurter, Ozan Ersoy, Sara Irina Fabrikant, Tijmen R. Klein, and Alexandru C. Telea. Bundled visualization of dynamic graph and trail data. IEEE Transactions on Visualization and Computer Graphics, 20(8):1141– 1157, August 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Harrison:2018:PVA

[HEFR18] Dave Graham Harrison, Nick D.

Efford, Quentin J. Fisher, and Roy Alan Ruddle. PET-Miner — a visual analysis tool for petrophysical properties of core sample data. IEEE Transactions on Visualization and Computer Graphics, 24(5):1728–1741, May 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07879300-
abs.html.

Hegarty:2010:VKA

[Heg10]  Mary Hegarty.  VisWeek

keynote address. IEEE Transactions on Visualization and Computer Graphics, 16(6): xxiv, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hermosilla:2017:PBV

[HEG^{+}17] Pedro Hermosilla, Jorge Estrada, Victor Guallar, Timo Ropinski, Álvar Vinacua, and Pere-Pau Vázquez. Physicsbased visual characterization of molecular interaction forces. IEEE Transactions on Visualization and Computer Graphics, 23(1):731–740, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heiberg:2003:TDF

[HEWK03] Einar Heiberg, Tino Ebbers, Lars Wigström, and Matts Karlsson. Three-dimensional flow characterization using vector pattern matching. IEEE Transactions on Visualization and Computer Graphics, 9(3):313–319, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/03/v0313abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/03/
                                        v0313.pdf.

Henry:2006:MDR

[HF06] Nathalie Henry and Jean-Daniel Fekete. MatrixExplorer: a dual-representation system to explore social networks. IEEE Transactions on Visualization and Computer Graphics, 12(5): 677–684, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Henderson:2010:OTU

[HF10] Steven Henderson and Steven Feiner. Opportunistic tangible user interfaces for augmented reality. IEEE Transactions on Visualization and Computer Graphics, 16(1):4– 16, January/February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Henderson:2011:EBA

[HF11] Steven Henderson and Steven Feiner. Exploring the benefits of augmented reality documentation for maintenance and repair. IEEE Transactions on Visualization and Computer Graphics, 17(10): 1355–1368, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hollt:2012:SIV

[HFG^{+}12] Thomas Hollt, Wolfgang Freiler, Fritz-M. Gschwantner, Helmut Doleisch, Gabor Heinemann, and Markus Hadwiger. SeiVis: an interactive visual subsurface modeling application. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2226–2235, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2018:AHS

[HFL18] Xin Hu, Xiao-Ming Fu, and Ligang Liu. Advanced hierarchical spherical parameterizations. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1930–1941, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07927729-
abs.html.

Henry:2007:NHV

[HFM07] Nathalie Henry, Jean-Daniel Fekete, and Michael J. McGuffin. NodeTrix: a hybrid visualization of social networks. IEEE Transactions on Visualization and Computer Graphics, 13(6):1302–1309, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1302s.
mov.

Hinrichs:2016:SPU

[HFM16] U. Hinrichs, S. Forlini, and B. Moynihan. Speculative practices: Utilizing InfoVis to explore untapped literary collections. IEEE Transactions on Visualization and Computer Graphics, 22(1): 429–438, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hofmann:2012:GTP

[HFMC12] Heike Hofmann, Lendie Follett, Mahbubul Majumder, and Dianne Cook. Graphical tests for power comparison of competing designs. IEEE Transactions on Visualization and Computer Graphics, 18(12):2441–2448, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hubeli:2001:MMN

[HG01] Andreas Hubeli and Markus Gross. Multiresolution methods for nonmanifold models. IEEE Transactions on Visualization and Computer Graphics, 7(3):207–221, July 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0207.
pdf;   http://www.computer.
org/tvcg/tg2001/v0207abs.
htm.

Hummel:2010:IIR

[HGH^{+}10] Mathias Hummel, Christoph Garth, Bernd Hamann, Hans Hagen, and Kenneth I. Joy. IRIS: Illustrative rendering for integral surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1319–1328, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hecher:2018:HDU

[HGWW18] Michael Hecher, Paul Guerrero, Peter Wonka, and Michael Wimmer. How do users map points between dissimilar shapes? IEEE Transactions on Visualization and Computer Graphics, 24(8): 2327–2338, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/08/07990196-
                                        abs.html.

Havel:2010:YFR

[HH10] Jiři Havel and Adam Herout. Yet faster ray-triangle intersection (using SSE4). IEEE Transactions on Visualization and Computer Graphics, 16 (3):434–438, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hao:2016:EVT

[HHB16] L. Hao, C. G. Healey, and S. A. Bass. Effective visualization of temporal ensembles. IEEE Transactions on Visualization and Computer Graphics, 22(1):787–796, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

He:2001:RPV

[HHCL01] Taosong He, Lichan Hong, Dongqing Chen, and Zhengrong Liang. Reliable path for virtual endoscopy: Ensuring complete examination of human organs. IEEE Transactions on Visualization and Computer Graphics, 7(4):333– 342, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0333.
pdf;   http://www.computer.
org/tvcg/tg2001/v0333abs.
htm.

Huettenberger:2014:DSM

[HHG14] Lars Huettenberger, Christian Heine, and Christoph Garth. Decomposition and simplification of multivariate data using Pareto sets. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2684–2693, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875963-
abs.html.

Hogan:2016:EIT

[HHH16] Trevor Hogan, Uta Hinrichs, and Eva Hornecker. The elicitation interview technique: Capturing people’s experiences of data representations. IEEE Transactions on Visualization and Computer Graphics, 22(12):2579– 2593, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2016/12/07369991-
                                        abs.html.

Heimerl:2016:CVA

[HHKE16] F. Heimerl, Q. Han, S. Koch, and T. Ertl. CiteRivers: Visual analytics of citation patterns. IEEE Transactions on Visualization and Computer Graphics, 22(1):190–199, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hero:2014:VRC

[HHM14] Robert Hero, Chris Ho, and Kwan-Liu Ma. Volume rendering of curvilinear-grid data using low-dimensional deformation textures. IEEE Transactions on Visualization and Computer Graphics, 20(9): 1330–1343, September 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hakone:2017:PID

[HHO^{+}17] Anzu Hakone, Lane Harrison, Alvitta Ottley, Nathan Winters, Caitlin Gutheil, Paul K. J. Han, and Remco Chang. PROACT: Iterative design of a patient-centered visualization for effective prostate cancer health risk communication. IEEE Transactions on Visualization and Computer Graphics, 23(1):601–610, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2017:KOB

[HHQH17] Fei Hou, Ying He, Hong Qin, and Aimin Hao. Knot optimization for biharmonic Bsplines on manifold triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 23(9):2082– 2095, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07558244-
abs.html.

He:1996:CTS

[HHVW96] Taosong He, Lichan Hong, Amitabh Varshney, and Sidney W. Wang. Controlled topology simplification. IEEE Transactions on Visualization and Computer Graphics, 2(2): 171–184, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0171.
pdf;   http://www.computer.
org/tvcg/tg1996/v0171abs.
htm.

Havre:2002:TVT

[HHWN02] Susan Havre, Elizabeth Hetzler, Paul Whitney, and Lucy Nowell. ThemeRiver: Visualizing thematic changes in large document collections. IEEE Transactions on Visualization and Computer Graphics, 8(1):9–20, January 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0009.
pdf;   http://www.computer.
org/tvcg/tg2002/v0009abs.
htm.

Hu:2017:VSD

[HHZH17] Jiaxi Hu, Hajar Hamidian, Zichun Zhong, and Jing Hua. Visualizing shape deformations with variation of geometric spectrum. IEEE Transactions on Visualization and Computer Graphics, 23 (1):721–730, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hamasaki:2018:HHM

[HIH^{+}18] Takumi Hamasaki, Yuta Itoh, Yuichi Hiroi, Daisuke Iwai, and Maki Sugimoto. HySAR: Hybrid material rendering by an optical see-through headmounted display with spatial augmented reality projection. IEEE Transactions on Visualization and Computer Graphics, 24(4):1457– 1466, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260968-
abs.html.

Huron:2014:CVR

[HJC14] Samuel Huron, Yvonne Jansen, and Sheelagh Carpendale. Constructing visual representations: Investigating the use of tangible tokens. IEEE Transactions on Visualization and Computer Graphics, 20(12):2102–2111, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875946-
abs.html.

Heinrich:2019:CPA

[HJLH19] Florian Heinrich, Fabian Joeres, Kai Lawonn, and Christian Hansen. Comparison of projective augmented reality concepts to support medical needle insertion. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2157–2167, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8667734/.

Hamann:1999:CCH

[HJW99a] B. Hamann, B. W. Jordan, and D. F. Wiley. Corrections to “On a Construction of a Hierarchy of Best Linear Spline Approximations Using Repeated Bisections”. IEEE Transactions on Visualization and Computer Graphics, 5(2): 190, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0190.

pdf. See [HJW99b].

Hamann:1999:CHB

[HJW99b] Bernd Hamann, Benjamin W. Jordan, and David F. Wiley. On a construction of a hierarchy of best linear spline approximations using repeated bisection. IEEE Transactions on Visualization and Computer Graphics, 5(1):30–46, January/March 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0030.
pdf;   http://www.computer.
org/tvcg/tg1999/v0030abs.

htm. See corrections [HJW99a].

Hong:1999:FPB

[HK99] Lichan Hong and A. E. Kaufman. Fast projectionbased ray-casting algorithm for rendering curvilinear volumes. IEEE Transactions on Visualization and Computer Graphics, 5(4):322–332, October/December 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0322.
pdf;   http://www.computer.
org/tvcg/tg1999/v0322abs.
htm.

Ho:2009:IRM

[HK09] Edmond S. L. Ho and Taku Komura. Indexing and retrieving motions of characters in close contact. IEEE Transactions on Visualization and Computer Graphics, 15 (3):481–492, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hachet:2010:GEI

[HK10] Martin Hachet and Ernst Kruijff. Guest Editor’s introduction: Special section on the ACM Symposium on Virtual Reality Software and Technology. IEEE Transactions on Visualization and Computer Graphics, 16(1):2– 3, January/February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2016:PMP

[HK16] Y. Huang and M. Kallmann.

Planning motions and placements for virtual demonstrators. IEEE Transactions on Visualization and Computer Graphics, 22(5):1568– 1579, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hoang:2019:STB

[HKB^{+}19] Duong Hoang, Pavol Klacansky, Harsh Bhatia, Peer-Timo Bremer, Peter Lindstrom, and Valerio Pascucci. A study of the trade-off between reducing precision and reducing resolution for data analysis and visualization. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1193–1203, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440822-
abs.html.

Heimerl:2012:VCT

[HKBE12] Florian Heimerl, Steffen Koch, Harald Bosch, and Thomas Ertl. Visual classifier training for text document retrieval. IEEE Transactions on Visualization and Computer Graphics, 18(12):2839– 2848, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haehn:2014:DEI

[HKBR^{+}14] Daniel Haehn, Seymour Knowles-Barley, Mike Roberts, Johanna Beyer, Narayanan Kasthuri, Jeff W. Lichtman, and Hanspeter Pfister. Design and evaluation of interactive proofreading tools for connectomics. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2466–2475, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

csdl/trans/tg/2014/12/06875931-
abs.html.

Ha:2012:IOC

[HKC^{+}12] Linh Khanh Ha, Jens Kruger, Joao Luiz Dihl Comba, Claudio T. Silva, and Sarang Joshi. ISP: an optimal out-of-core image-set processing streaming architecture for parallel heterogeneous systems. IEEE Transactions on Visualization and Computer Graphics, 18 (6):838–851, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haroz:2016:CSP

[HKF16] Steve Haroz, Robert Kosara, and Steven L. Franconeri. The connected scatterplot for presenting paired time series. IEEE Transactions on Visualization and Computer Graphics, 22(9):2174– 2186, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heinzl:2007:SEM

[HKG07] Christoph Heinzl, Johann Kastner, and Eduard Gröller. Surface extraction from multimaterial components for metrology using dual energy CT. IEEE Transactions on Visualization and Computer Graphics, 13(6):1520–1527, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hoferlin:2012:EFF

[HKH^{+}12] Markus Hoferlin, Kuno Kurzhals, Benjamin Hoferlin, Gunther Heidemann, and Daniel Weiskopf. Evaluation of fast-forward video visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2095– 2103, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hullman:2018:IRG

[HKKS18] J. Hullman, M. Kay, Y. S. Kim, and S. Shrestha. Imagining replications: Graphical prediction — discrete visualizations improve recall — estimation of effect uncertainty. IEEE Transactions on Visualization and Computer Graphics, 24(1):446–456, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hedman:2017:SMC

[HKL17] Peter Hedman, Tero Karras, and Jaakko Lehtinen. Sequential Monte Carlo instant radiosity. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1442–1453, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/05/07867077-
                                        abs.html.

Hohman:2019:VAD

[HKPC19] F. Hohman, M. Kahng, R. Pienta, and D. H. Chau. Visual analytics in deep learning: An interrogative survey for the next frontiers. IEEE Transactions on Visualization and Computer Graphics, 25 (8):2674–2693, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hauser:2013:GEI

[HKQ13] Helwig Hauser, Stephen Kobourov, and Huamin Qu. Guest Editors’ introduction: Special section on the IEEE Pacific Visualization Symposium 2012. IEEE Transactions on Visualization and Computer Graphics, 19(6): 898–899, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Healey:2008:VPM

[HKR^{+}08] Christopher Healey, Sarat Kocherlakota, Vivek Rao, Reshma Mehta, and Robert St. Amant. Visual perception and mixed-initiative interaction for assisted visualization design. IEEE Transactions on Visualization and Computer Graphics, 14(2):396– 411, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2017:SSS

[HKYM17] Haibin Huang, Evangelos Kalogerakis, Ersin Yumer, and Radomir Mech. Shape synthesis from sketches via procedural models and convolutional networks. IEEE Transactions on Visualization and Computer Graphics, 23 (8):2003–2013, August 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/08/07530838-
abs.html.

Habibi:2002:DPC

[HL02] Arash Habibi and Annie Luciani. Dynamic particle coating. IEEE Transactions on Visualization and Computer Graphics, 8(4):383–394, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0383abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0383.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0383.pdf.

Hughes:2009:KJP

[HL09] David M. Hughes and Ik Soo Lim. Kd-Jump: a pathpreserving stackless traversal for faster isosurface raytracing on GPUs. IEEE Transactions on Visualization and Computer Graphics, 15(6):1555– 1562, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2018:AOW

[HLB^{+}18] Haikun Huang, Ni-Ching Lin, Lorenzo Barrett, Darian Springer, Hsueh-Cheng Wang, Marc Pomplun, and Lap-Fai Yu. Automatic optimization of wayfinding design. IEEE Transactions on Visualization and Computer Graphics, 24(9):2516– 2530, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08063899-
abs.html.

Hristova:2018:TMG

[HLCB18] Hristina Hristova, Olivier Le Meur, Remi Cozot, and Kadi Bouatouch. Transformation of the multivariate generalized Gaussian distribution for image editing. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2813–2826, October 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08094021-
abs.html.

Hua:2008:GDW

[HLD^{+}08] Jing Hua, Zhaoqiang Lai, Ming Dong, Xianfeng Gu, and Hong Qin. Geodesic distanceweighted shape vector image diffusion. IEEE Transactions on Visualization and Computer Graphics, 14(6):1643– 1650, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hong:2014:FLD

[HLG^{+}14] Fan Hong, Chufan Lai, Hanqi Guo, Enya Shen, Xiaoru Yuan, and Sikun Li. FLDA: Latent Dirichlet allocation based unsteady flow analysis. IEEE Transactions on Visualization and Computer Graphics, 20(12):2545– 2554, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875956-
abs.html.

Hesselink:1997:TSS

[HLL97] L. Hesselink, Yuval Levy, and Yingmei Lavin. The topology of symmetric, secondorder 3D tensor fields. IEEE Transactions on Visualization and Computer Graphics, 3(1): 1–11, January/March 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0001.
pdf;   http://www.computer.
org/tvcg/tg1997/v0001abs.
htm.

Hege:2010:GEI

[HLM10] Hans-Christian Hege, David H.

Laidlaw, and Raghu Machiraju. Guest Editors’ introduction: Special section on volume graphics and point-based graphics. IEEE Transactions on Visualization and Computer Graphics, 16(4):531– 532, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hlawatsch:2011:FRG

[HLNW11] Marcel Hlawatsch, Philipp Leube, Wolfgang Nowak, and Daniel Weiskopf. Flow radar glyphs — static visualization of unsteady flow with uncertainty. IEEE Transactions on Visualization and Computer Graphics, 17(12):1949– 1958, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hillaire:2012:DAR

[HLRC^{+}12] Sebastien Hillaire, Anatole Lecuyer, Tony Regia-Corte, Remi Cozot, Jerome Royan, and Gaspard Breton. Design and application of realtime visual attention model for the exploration of 3D virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18 (3):356–368, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hadwiger:2008:IVE

[HLRS^{+}08] Markus Hadwiger, Fritz Laura, Christof Rezk-Salama, Thomas Höllt, Georg Geier, and Thomas Pabel. Interactive volume exploration for feature detection and quantification in industrial CT data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1507–1514, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hernell:2010:LAO

[HLY10] Frida Hernell, Patric Ljung, and Anders Ynnerman. Local ambient occlusion in direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 16(4):548–559, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2018:GTB

[HLYL18] Y. J. Huang, W. C. Lin, I. C. Yeh, and T. Y. Lee. Geometric and textural blending for 3D model stylization. IEEE Transactions on Visualization and Computer Graphics, 24(2):1114–1126, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hanson:1995:QFA

[HM95] Andrew J. Hanson and Hui Ma. Quaternion frame approach to streamline visualization. IEEE Transactions on Visualization and Computer Graphics, 1(2): 164–174, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0164.
pdf;   http://www.computer.
org/tvcg/tg1995/v0164abs.
htm.

Hosssain:2010:EAA

[HM10] Zahid Hosssain and Torsten Moller. Edge aware anisotropic diffusion for 3D scalar data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1376–1385, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hauser:2001:TLV

[HMBG01] Helwig Hauser, Lukas Mroz, Gian Italo Bischi, and M. Eduard Gröller. Two-level volume rendering. IEEE Transactions on Visualization and Computer Graphics, 7(3): 242–252, July 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0242.
pdf;   http://www.computer.
org/tvcg/tg2001/v0242abs.
htm.

Herman:2000:GVN

[HMM00] Ivan Herman, Guy Melançon, and M. Scott Marshall. Graph visualization and navigation in information visualization: a survey. IEEE Transactions on Visualization and Computer Graphics, 6(1):24–43, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0024.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0024abs.
                                        htm.

Heer:2008:GHV

[HMSA08] Jeffrey Heer, Jock Mackinlay, Chris Stolte, and Maneesh Agrawala. Graphical histories for visualization: Supporting analysis, communication, and evaluation. IEEE Transactions on Visualization and Computer Graphics, 14(6):1189–1196, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hecht:2010:VRF

[HMT10] Florian Hecht, Peter J. Mucha, and Greg Turk. Virtual rheoscopic fluids. IEEE Transactions on Visualization and Computer Graphics, 16(1):147–160, January/ February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hadwiger:2019:TDF

[HMTR19] Markus Hadwiger, Matej Mlejnek, Thomas Theussl, and Peter Rautek. Timedependent flow seen through approximate observer Killing fields. IEEE Transactions on Visualization and Computer Graphics, 25(1):1257– 1266, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440037-
abs.html.

Hollt:2014:OFV

[HMZ^{+}14] Thomas Hollt, Ahmed Magdy, Peng Zhan, Guoning Chen, Ganesh Gopalakrishnan, Ibrahim Hoteit, Charles D. Hansen, and Markus Hadwiger. Ovis: A framework for visual analysis of ocean forecast ensembles. IEEE Transactions on Visualization and Computer Graphics, 20(8):1114– 1126, August 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Harish:2013:DPC

[HN13] Pawan Harish and P. J.

Narayanan. Designing perspectively correct multiplanar displays. IEEE Transactions on Visualization and Computer Graphics, 19(3):

407–419, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2006:TSS

[HNR^{+}06] Adam Huang, Gregory M. Nielson, Anshuman Razdan, Gerald E. Farin, D. Page Baluch, and David G. Capco. Thin structure segmentation and visualization in threedimensional biomedical images: a shape-based approach. IEEE Transactions on Visualization and Computer Graphics, 12(1):93–102, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hossain:2012:SGC

[HOG^{+}12] M. Shahriar Hossain, Praveen Kumar Reddy Ojili, Cindy Grimm, Rolf Muller, Layne T. Watson, and Naren Ramakrishnan. Scatter/gather clustering: Flexibly incorporating user feedback to steer clustering results. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2829–2838, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hummel:2013:CVA

[HOGJ13] Mathias Hummel, Harald Obermaier, Christoph Garth, and Kenneth I. Joy. Comparative visual analysis of Lagrangian transport in CFD ensembles. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2743–2752, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Holten:2006:HEB

[Hol06] Danny Holten. Hierarchical edge bundles: Visualization of adjacency relations in hierarchical data. IEEE Transactions on Visualization and Computer Graphics, 12(5): 741–748, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hodgins:1998:PHM

[HOT98] J. K. Hodgins, J. F. O’Brien, and J. Tumblin. Perception of human motion with different geometric models. IEEE Transactions on Visualization and Computer Graphics, 4(4):307–316, October/December 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0307.
pdf;   http://www.computer.
org/tvcg/tg1998/v0307abs.
htm.

Hege:2004:GEI

[HP04] Hans-Christian Hege and Konrad Polthier. Guest Editors’ introduction: Special section on mathematics and visualization. IEEE Transactions on Visualization and Computer Graphics, 10(5):497–498, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/05/v0497.
pdf;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0497.htm.

Helgeland:2007:VVV

[HPAW07] Anders Helgeland, B. Anders Pettersson, Reif Øyvind Andreassen, and Carl Erik Wasberg. Visualization of vorticity and vortices in wall-bounded turbulent flows. IEEE Transactions on Visualization and Computer Graphics, 13(5):1055–1067, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2013:IST

[HPC^{+}13] Jin Huang, Zherong Pan, Guoning Chen, Wei Chen, and Hujun Bao. Imagespace texture-based outputcoherent surface flow visualization. IEEE Transactions on Visualization and Computer Graphics, 19(9):1476– 1487, September 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2008:GSP

[HPJG08] Yan Huang, Jingliang Peng, C.-C. Jay, and Kuo M. Gopi. A generic scheme for progressive point cloud coding. IEEE Transactions on Visualization and Computer Graphics, 14(2):440–453, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hurter:2018:FDB

[HPNT18] C. Hurter, S. Puechmorel, F. Nicol, and A. Telea. Functional decomposition for bundled simplification of trail sets. IEEE Transactions on Visualization and Computer Graphics, 24(1):500–510, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hollt:2018:CVG

[HPvU^{+}18] T. Höllt, N. Pezzotti, V. van Unen, F. Koning, B. P. F. Lelieveldt, and A. Vilanova. CyteGuide: Visual guidance for hierarchical singlecell analysis. IEEE Transactions on Visualization and Computer Graphics, 24(1): 739–748, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hua:2004:HBD

[HQ04] Jing Hua and Hong Qin.

Haptics-based dynamic implicit solid modeling. IEEE Transactions on Visualization and Computer Graphics, 10(5):574–586, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0574.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0574.pdf.

Hu:2007:IAR

[HQ07] Wei Hu and Kaihuai Qin. Interactive approximate rendering of reflections, refractions, and caustics. IEEE Transactions on Visualization and Computer Graphics, 13(1):46– 57, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2012:RDR

[HQ12] Tingbo Hou and Hong Qin. Robust dense registration of partial nonrigid shapes. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1268–1280, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2013:ADW

[HQ13] Tingbo Hou and Hong Qin. Admissible diffusion wavelets and their applications in space-frequency processing. IEEE Transactions on Visualization and Computer Graphics, 19(1):3–15, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hullman:2019:PES

[HQC^{+}19] Jessica Hullman, Xiaoli Qiao, Michael Correll, Alex Kale, and Matthew Kay. In pursuit of error: A survey of uncertainty visualization evaluation. IEEE Transactions on Visualization and Computer Graphics, 25(1):903– 913, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08457476-
abs.html.

Hong:2006:PCA

[HQK06] Wei Hong, Feng Qiu, and Arie Kaufman. A pipeline for computer aided polyp detection. IEEE Transactions on Visualization and Computer Graphics, 12(5): 861–868, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2012:DBP

[HQQ12] Fei Hou, Yue Qi, and Hong Qin. Drawing-based procedural modeling of Chinese architectures. IEEE Transactions on Visualization and Computer Graphics, 18(1):30– 42, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hasegawa:2018:MUF

[HQS18] Keisuke Hasegawa, Liwei Qiu, and Hiroyuki Shinoda. Midair ultrasound fragrance rendering. IEEE Transactions on Visualization and Computer Graphics, 24(4):1477– 1485, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/04/08260917.
                                        pdf.

Hoffmann:1996:RMS

[HR96] Christoph M. Hoffmann and Jaroslaw R. Rossignac. A road map to solid modeling. IEEE Transactions on Visualization and Computer Graphics, 2(1):3–10, March 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0003.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0003abs.
                                        htm.

Hable:2007:CCS

[HR07a] John Hable and Jarek Rossignac. CST: Constructive solid trimming for rendering BReps and CSG. IEEE Transactions on Visualization and Computer Graphics, 13(5):1004– 1014, September/October 2007. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heer:2007:ATS

[HR07b] Jeffrey Heer and George Robertson. Animated transitions in statistical data graphics. IEEE Transactions on Visualization and Computer Graphics, 13(6):1240–1247, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1240s.
mov.

Hutson:2011:JAN

[HR11] Malcolm Hutson and Dirk Reiners. JanusVF: Accurate navigation using SCAAT and virtual fiducials. IEEE Transactions on Visualization and Computer Graphics, 17 (1):3–13, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hurter:2019:FST

[HRD^{+}19] Christophe Hurter, Nathalie Henry Riche, Steven M. Drucker, Maxime Cordeil, Richard Alligier, and Romain Vuillemot. FiberClay: Sculpting three dimensional trajectories to reveal structural insights. IEEE Transactions on Visualization and Computer Graphics, 25 (1):704–714, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440808-
abs.html.

Hincapie-Ramos:2015:SRT

[HRISI15] J. David Hincapie-Ramos, L. Ivanchuk, S. K. Sridharan, and P. P. Irani. Smart-Color: Real-time color and contrast correction for optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 21(12):1336– 1348, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2003:VSU

[HRN^{+}03] Jiuxiang Hu, Anshuman Razdan, Gregory M. Nielson, Gerald E. Farin, D. Page Baluch, and David G. Capco. Volumetric segmentation using Weibull E-SD fields. IEEE Transactions on Visualization and Computer Graphics, 9(3):320–328, July/ September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0320abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0320.pdf.

Haunert:2011:DRN

[HS11] Jan-Henrik Haunert and Leon Sering. Drawing road networks with focus regions.

IEEE Transactions on Visualization and Computer Graphics, 17(12):2555–2562, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hinrichs:2008:EBI

[HSC08] Uta Hinrichs, Holly Schmidt, and Sheelagh Carpendale. EMDialog: Bringing information visualization into the museum. IEEE Transactions on Visualization and Computer Graphics, 14(6):1181– 1188, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heine:2011:DCT

[HSCS11] Christian Heine, Dominic Schneider, Hamish Carr, and Gerik Scheuermann. Drawing contour trees in the plane. IEEE Transactions on Visualization and Computer Graphics, 17(11):1599– 1611, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hadlak:2013:SVA

[HSCW13] Steffen Hadlak, Heidrun Schumann, Clemens H. Cap, and Till Wollenberg. Supporting the visual analysis of dynamic networks by clustering associated temporal attributes. IEEE Transactions on Visualization and Computer Graphics, 19(12):2267– 2276, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heine:2006:VBT

[HSF^{+}06] Christian Heine, Gerik Scheuermann, Christoph Flamm, Ivo L. Hofacker, and Peter F. Stadler. Visualization of barrier tree sequences. IEEE Transactions on Visualization and Computer Graphics, 12(5):781–788, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Henriksen:2004:VTR

[HSH04] Knud Henriksen, Jon Sporring, and Kasper Hornbæk. Virtual trackballs revisited. IEEE Transactions on Visualization and Computer Graphics, 10(2):206–216, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2004/
                                        02/v0206abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        trans/tg/2004/02/v0206.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/02/
                                        v0206.pdf.

Hu:2010:PVD

[HSH10] Liang Hu, Pedro V. Sander, and Hugues Hoppe. Parallel view-dependent level-of-detail control. IEEE Transactions on Visualization and Computer Graphics, 16(5): 718–728, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Henry:2014:IFC

[HSK14] Joseph Henry, Hubert P. H.

Shum, and Taku Komura. Interactive formation control in complex environments. IEEE Transactions on Visualization and Computer Graphics, 20 (2):211–222, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Holden:2017:LIR

[HSK17] Daniel Holden, Jun Saito, and Taku Komura. Learning inverse rig mappings by nonlinear regression. IEEE Transactions on Visualization and Computer Graphics, 23 (3):1167–1178, March 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07742381-
abs.html.

Hagh-Shenas:2007:WVB

[HSKIH07] Haleh Hagh-Shenas, Sunghee Kim, Victoria Interrante, and Christopher Healey. Weaving versus blending: a quantitative assessment of the information carrying capacities of two alternative methods for conveying multivariate data with color. IEEE Transactions on Visualization and Computer Graphics, 13(6):1270–1277, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2019:CAD

[HSL19] Yifan Hu, Lei Shi, and Qingsong Liu. A coloring algorithm for disambiguating graph and map drawings. IEEE Transactions on Visualization and Computer Graphics, 25(2):1321– 1335, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/02/08269823-
                                        abs.html.

Hauswiesner:2013:TCI

[HSR13a] Stefan Hauswiesner, Matthias Straka, and Gerhard Reitmayr. Temporal coherence in image-based visual hull rendering. IEEE Transactions on Visualization and Computer Graphics, 19(10):1758– 1767, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hauswiesner:2013:VTT

[HSR13b] Stefan Hauswiesner, Matthias Straka, and Gerhard Reitmayr. Virtual try-on through image-based rendering. IEEE Transactions on Visualization and Computer Graphics, 19(9):1552–1565, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Han:2018:ERP

[HSR18] Dustin T. Han, Mohamed Suhail, and Eric D. Ragan. Evaluating remapped physical reach for hand interactions with passive haptics in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1467–1476, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260974-
abs.html.

Hadlak:2011:SEL

[HSS11] Steffen Hadlak, Hans-Jörg Schulz, and Heidrun Schumann. In situ exploration of large dynamic networks. IEEE Transactions on Visualization and Computer Graphics, 17(12):2334– 2343, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hermann:2016:AIV

[HSSK16] M. Hermann, A. C. Schunke, T. Schultz, and R. Klein. Accurate interactive visualization of large deformations and variability in biomedical image ensembles. IEEE Transactions on Visualization and Computer Graphics, 22(1): 708–717, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hoque:2018:APP

[HSTD18] E. Hoque, V. Setlur, M. Tory, and I. Dykeman. Applying pragmatics principles for interaction with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 24(1):309–318, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hlawatsch:2011:HLI

[HSW11] Marcel Hlawatsch, Filip Sadlo, and Daniel Weiskopf. Hierarchical line integration. IEEE Transactions on Visualization and Computer Graphics, 17(8):1148–1163, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hou:2011:MSG

[HSZ^{+}11] Qiming Hou, Xin Sun, Kun Zhou, Christian Lauterbach, and Dinesh Manocha. Memory-scalable GPU spatial hierarchy construction. IEEE Transactions on Visualization and Computer Graphics, 17 (4):466–474, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2015:PVP

[HTA^{+}15] Dandan Huang, Melanie Tory, Bon Adriel Aseniero, Lyn Bartram, Scott Bateman, Sheelagh Carpendale, Anthony Tang, and Robert Woodbury. Personal visualization and personal visual analytics. IEEE Transactions on Visualization and Computer Graphics, 21(3): 420–433, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06908006-
abs.html.

Hurter:2009:FSA

[HTC09] Christophe Hurter, Benjamin Tissoires, and Stephane Conversy. FromDaDy: Spreading aircraft trajectories across views to support iterative queries. IEEE Transactions on Visualization and Computer Graphics, 15(6):1017– 1024, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hurter:2011:MAS

[HTE11] Christophe Hurter, Alexandru Telea, and Ozan Ersoy. MoleView: an attribute and structure-based semantic lens for large element-based plots. IEEE Transactions on Visualization and Computer Graphics, 17(12):2600– 2609, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hamann:1997:ACP

[HTF97] B. Hamann, I. J. Trotts, and G. E. Farin. On approximating contours of the piecewise trilinear interpolant using triangular rational-quadratic Bézier patches. IEEE Transactions on Visualization and Computer Graphics, 3(3):215– 227, July/September 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0215.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0215abs.
                                        htm.

Hajizadeh:2013:SAT

[HTL13] Amir Hossein Hajizadeh, Melanie Tory, and Rock Leung. Supporting awareness through collaborative brushing and linking of tabular data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2189– 2197, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hentschel:2008:IBD

[HTP^{+}08] Bernd Hentschel, Irene Tedjo, Markus Probst, Marc Wolter, Marek Behr, Christian Bischof, and Torsten Kuhlen. Interactive blood damage analysis for ventricular assist devices. IEEE Transactions on Visualization and Computer Graphics, 14(6):1515–1522, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haehn:2019:EGP

[HTP19] Daniel Haehn, James Tompkin, and Hanspeter Pfister. Evaluating ‘graphical perception’ with CNNs. IEEE Transactions on Visualization and Computer Graphics, 25 (1):641–650, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440806-
abs.html.

Huang:2011:ISI

[HTZ^{+}11] J. Huang, Y. Tong, K. Zhou, H. Bao, and M. Desbrun. Interactive shape interpolation through controllable dynamic deformation. IEEE Transactions on Visualization and Computer Graphics, 17 (7):983–992, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2016:FCI

[Hu16a] Shi-Min Hu. Faithful completion of images of scenic landmarks using Internet images.

IEEE Transactions on Visualization and Computer Graphics, 22(8):1945–1958, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2016:SSS

[Hu16b] Shi-Min Hu. Support substructures: Support-induced part-level structural representation. IEEE Transactions on Visualization and Computer Graphics, 22(8):2024– 2036, August 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hubbard:1995:CDI

[Hub95] Philip M. Hubbard. Collision detection for interactive graphics applications. IEEE Transactions on Visualization and Computer Graphics, 1 (3):218–230, September 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0218.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0218abs.
                                        htm.

Hernandez:2014:OUB

[HUPS14] Ruben Jesus Garcia Hernandez, Carlos Ureña, Jordi Poch, and Mateu Sbert. Overestimation and underestimation biases in photon mapping with non-constant kernels. IEEE Transactions on Visualization and Computer Graphics, 20(10):1441– 1450, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06781612-
abs.html.

Hoff:2000:AHP

[HV00] W. Hoff and T. Vincent.

Analysis of head pose accuracy in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 6(4):319–??, October/December 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0319.
pdf;   http://www.computer.
org/tvcg/tg2000/v0319abs.
htm.

Hofmann:2013:CAP

[HV13] Heike Hofmann and Marie Vendettuoli. Common angle plots as perception-true visualizations of categorical associations. IEEE Transactions on Visualization and Computer Graphics, 19(12):2297– 2305, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huron:2013:VS

[HVF13] Samuel Huron, Romain Vuillemot, and Jean-Daniel Fekete. Visual sedimentation. IEEE Transactions on Visualization and Computer Graphics, 19(12):2446–2455, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hlawatsch:2011:CSC

[HVSW11] Marcel Hlawatsch, Joachim E. Vollrath, Filip Sadlo, and Daniel Weiskopf. Coherent structures of characteristic curves in symmetric second order tensor fields. IEEE Transactions on Visualization and Computer Graphics, 17 (6):781–794, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hansen:2016:GEI

[HVY16] Charles Hansen, Ivan Viola, and Xiaoru Yuan. Guest Editor’s introduction to the special section on the 2016 IEEE Pacific visualization symposium. IEEE Transactions on Visualization and Computer Graphics, 22(6):1638– 1639, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heinrich:2009:CPC

[HW09] Julian Heinrich and Daniel Weiskopf. Continuous parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 15(6):1531–1538, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Haroz:2012:HCL

[HW12] Steve Haroz and David Whitney. How capacity limits of attention influence information visualization effectiveness. IEEE Transactions on Visualization and Computer Graphics, 18(12):2402– 2410, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hough:2015:FPB

[HWA15] G. Hough, I. Williams, and C. Athwal. Fidelity and plausibility of bimanual interaction in mixed reality. IEEE Transactions on Visualization and Computer Graphics, 21(12):1377–1389, December 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hanel:2016:VQA

[HWHK16] C. Hänel, B. Weyers, B. Hentschel, and T. W. Kuhlen. Visual quality adjustment for volume rendering in a head-tracked virtual environment. IEEE Transactions on Visualization and Computer Graphics, 22 (4):1472–1481, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ho:2011:UDE

[HWL^{+}11] Tze-Yiu Ho, Liang Wan, Chi-Sing Leung, Ping-Man Lam, and Tien-Tsin Wong. Unicube for dynamic environment mapping. IEEE Transactions on Visualization and Computer Graphics, 17(1): 51–63, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hamann:1995:PPG

[HWM95] Bernd Hamann, Donghua Wu, and Robert J. Moorhead II. On particle path generation based on quadrilinear interpolation and Bernstein–Bézier polynomials. IEEE Transactions on Visualization and Computer Graphics, 1(3):210– 217, September 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0210.
pdf;   http://www.computer.
org/tvcg/tg1995/v0210abs.
htm.

Hu:2017:VSM

[HWS17] Mengdie Hu, Krist Wongsuphasawat, and John Stasko. Visualizing social media content with SentenTree. IEEE Transactions on Visualization and Computer Graphics, 23 (1):621–630, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

He:2018:PPM

[HWW18] Xiaowei He, Huamin Wang, and Enhua Wu. Projective peridynamics for modeling versatile elastoplastic materials. IEEE Transactions on Visualization and Computer Graphics, 24(9):2589– 2599, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08048582-
abs.html.

He:2019:VAR

[HWZ^{+}19] J. He, L. Wang, W. Zhou, H. Zhang, X. Cui, and Y. Guo. Viewpoint assessment and recommendation for photographing architectures. IEEE Transactions on Visualization and Computer Graphics, 25(8):2636–2649, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ho:2015:AFD

[HXF^{+}15] Tze-Yiu Ho, Yi Xiao, Rui-Bin Feng, Chi-Sing Leung, and Tien-Tsin Wong. Allfrequency direct illumination with vectorized visibility. IEEE Transactions on Visualization and Computer Graphics, 21(8):945–958, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.

org/csdl/trans/tg/2015/
08/07050311-abs.html.

Hu:2017:EBF

[HYB^{+}17] Kaimo Hu, Dong-Ming Yan, David Bommes, Pierre Alliez, and Bedrich Benes. Errorbounded and feature preserving surface remeshing with minimal angle improvement. IEEE Transactions on Visualization and Computer Graphics, 23(12):2560– 2573, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07756294-
abs.html.

Harrison:2014:RVC

[HYFC14] Lane Harrison, Fumeng Yang, Steven Franconeri, and Remco Chang. Ranking visualizations of correlation using Weber’s Law. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1943–1952, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875978-
abs.html.

Huang:2012:DPC

[HYZ^{+}12] Haoda Huang, KangKang Yin, Ling Zhao, Yue Qi, Yizhou Yu, and Xin Tong. Detail-preserving controllable deformation from sparse examples. IEEE Transactions on Visualization and Computer Graphics, 18(8):1215– 1227, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2013:CEF

[HZ13] Jian Huang and Xiaomin Zhu.

Coupled ensemble flow line advection and analysis. IEEE Transactions on Visualization and Computer Graphics, 19(12):2733–2742, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2014:VPM

[HZH14] Jiaxi Hu, Guangyu Jeff Zou, and Jing Hua. Volumepreserving mapping and registration for collective data visualization. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2664–2673, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875953-
abs.html.

Hu:2019:SDD

[HZL^{+}19] Zhiming Hu, Congyi Zhang, Sheng Li, Guoping Wang, and Dinesh Manocha. SGaze: A data-driven eye-head coordination model for realtime gaze prediction. IEEE Transactions on Visualization and Computer Graphics, 25(5):

2002–2010, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8643434/.

Hsu:2013:MCA

[HZM13] Wei-Hsien Hsu, Yubo Zhang, and Kwan-Liu Ma. A multicriteria approach to camera motion design for volume data animation. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2792–2801, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2016:TGB

[HZM^{+}16] X. Huang, Y. Zhao, C. Ma, J. Yang, X. Ye, and C. Zhang. TrajGraph: A graph-based visual analytics approach to studying urban network centralities using taxi trajectory data. IEEE Transactions on Visualization and Computer Graphics, 22(1):160–169, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Itoh:2016:GLF

[IAIK16] Yuta Itoh, Toshiyuki Amano, Daisuke Iwai, and Gudrun Klinker. Gaussian light field: Estimation of viewpointdependent blur for optical seethrough head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 22(11):2368– 2376, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Iwai:2019:NCT

[IAS19] Daisuke Iwai, Mei Aoki, and Kosuke Sato. Non-contact thermo-visual augmentation by IR-RGB projection. IEEE Transactions on Visualization and Computer Graphics, 25(4):1707–1716, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8327511/.

Isaacs:2014:CCH

[IBJ^{+}14] Katherine E. Isaacs, Peer-Timo Bremer, Ilir Jusufi, Todd Gamblin, Abhinav Bhatele, Martin Schulz, and Bernd Hamann. Combing the communication hairball: Visualizing parallel execution traces using logical time. IEEE Transactions on Visualization and Computer Graphics, 20(12):2349– 2358, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876005-
abs.html.

Isenberg:2007:ITC

[IC07] Petra Isenberg and Sheelagh Carpendale. Interactive tree comparison for colocated collaborative information visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6):1232–1239, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1232s.
avi.

Ihmsen:2014:IIS

[ICS^{+}14] Markus Ihmsen, Jens Cornelis, Barbara Solenthaler, Christopher Horvath, and Matthias Teschner. Implicit incompressible sph. IEEE Transactions on Visualization and Computer Graphics, 20 (3):426–435, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Isaacs:2014:FVS

[IDA^{+}14] Ellen Isaacs, Kelly Damico, Shane Ahern, Eugene Bart, and Mudita Singhal. Footprints: A visual search tool that supports discovery and coverage tracking. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1793–1802, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875947-
abs.html.

Itoh:2015:SPC

[IDAK15] Yuta Itoh, Maksym Dzitsiuk, Toshiyuki Amano, and Gudrun Klinker. Semiparametric color reproduction method for optical seethrough head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 21(11):1269– 1278, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165643-
abs.html.

Isenberg:2013:HIV

[IDW^{+}13] Petra Isenberg, Pierre Dragicevic, Wesley Willett, Anastasia Bezerianos, and Jean-Daniel Fekete. Hybrid-image visualization for large viewing environments. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2346–2355, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Izadi:2014:KPE

[IFM14] Shahram Izadi, Andrew Fitzgibbon, and Paul McIlroy. Kinectrack: 3D pose estimation using a projected dense dot pattern. IEEE Transactions on Visualization and Computer Graphics, 20 (6):839–851, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Interrante:1997:CSS

[IFP97] V. Interrante, H. Fuchs, and S. M. Pizer. Conveying the 3D shape of smoothly curving transparent surfaces via texture. IEEE Transactions on Visualization and Computer Graphics, 3(2):98– 117, April/June 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0098.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0098abs.
                                        htm.

Isenberg:2012:CLC

[IFP^{+}12] Petra Isenberg, Danyel Fisher, Sharoda A. Paul, Meredith Ringel Morris, Kori Inkpen, and Mary Czerwinski. Co-located collaborative visual analytics around a tabletop display. IEEE Transactions on Visualization and Computer Graphics, 18 (5):689–702, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Isaacs:2019:PCL

[IG19] K. E. Isaacs and T. Gamblin. Preserving command line workflow for a package management system using ASCII DAG visualization. IEEE Transactions on Visualization and Computer Graphics, 25(9):2804–2820, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ibrahim:2018:APS

[IHD^{+}18] Adam Ibrahim, Brandon Huynh, Jonathan Downey, Tobias Höllerer, Dorothy Chun, and John O’donovan. ARbis Pictus: A study of vocabulary learning with augmented reality. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2867–2874, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8457524/.

Im:2005:MGS

[IHK05] Yeon-Ho Im, Chang-Young Han, and Lee-Sup Kim. A method to generate soft shadows using a layered depth image and warping. IEEE Transactions on Visualization and Computer Graphics, 11 (3):265–272, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Isenberg:2017:VOM

[IHK^{+}17] Petra Isenberg, Florian Heimerl, Steffen Koch, Tobias Isenberg, Panpan Xu, Charles D. Stolper, Michael Sedlmair, Jian Chen, Torsten Moller, and John Stasko. Vispubdata.org: A metadata collection about IEEE visualization (VIS) publications. IEEE Transactions on Visualization and Computer Graphics, 23(9):2199– 2206, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07583708-
abs.html.

Ibanez:2001:VAT

[IHR01] L. Ibáñez, C. Hamitouche, and C. Roux. A vectorial algorithm for tracing discrete straight lines in N-dimensional generalized grids. IEEE Transactions on Visualization and Computer Graphics, 7(2):97– ??, April/June 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0097.
pdf;   http://www.computer.
org/tvcg/tg2001/v0097abs.
htm.

Itoh:2017:OLC

[IHS17] Yuta Itoh, Takumi Hamasaki, and Maki Sugimoto. Occlusion leak compensation for optical see-through displays using a single-layer transmissive spatial light modulator. IEEE Transactions on Visualization and Computer Graphics, 23(11):2463– 2473, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007218-
abs.html.

Isenberg:2013:SRP

[IIC^{+}13] Tobias Isenberg, Petra Isenberg, Jian Chen, Michael Sedlmair, and Torsten Moller. A systematic review on the practice of evaluating visualization. IEEE Transactions on Visualization and Computer Graphics, 19(12):2818– 2827, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Isogawa:2014:MGI

[IIS14] Mariko Isogawa, Daisuke Iwai, and Kosuke Sato. Making graphical information visible in real shadows on interactive tabletops. IEEE Transactions on Visualization and Computer Graphics, 20 (9):1, September 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Isenberg:2017:VST

[IIS^{+}17] Petra Isenberg, Tobias Isenberg, Michael Sedlmair, Jian Chen, and Torsten Möller. Visualization as seen through its research paper keywords. IEEE Transactions on Visualization and Computer Graphics, 23(1):771–780, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Itoh:1995:AIP

[IK95] Takayuki Itoh and Koji Koyamada. Automatic isosurface propagation using an extrema graph and sorted boundary cell lists. IEEE Transactions on Visualization and Computer Graphics, 1(4):319– 327, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0319.
pdf;   http://www.computer.
org/tvcg/tg1995/v0319abs.
htm.

Itoh:2015:LFC

[IK15] Y. Itoh and G. Klinker. Lightfield correction for spatial calibration of optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 21 (4):471–480, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Interrante:2014:MGC

[IKLW14] Victoria Interrante, Daniel F. Keefe, Benjamin Lok, and Greg Welch. Message from the general chairs. IEEE Transactions on Visualization and Computer Graphics, 20 (4):vii, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Igarashi:2015:EN

[IKT15] T. Igarashi, G. Klinker, and B. H. Thomas. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 21(12):1319– 1320, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Interrante:2012:GEI

[ILMH12] Victoria Interrante, Benjamin C. Lok, Aditi Majumder, and Michitaka Hirose. Guest Editors’ introduction: Special Section on the IEEE Virtual Reality Conference (VR). IEEE Transactions on Visualization and Computer Graphics, 18(7): 1013–1016, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Im:2013:GGP

[IML13] Jean-Francois Im, Michael J.

McGuffin, and Rock Leung. GPLOM: The generalized plot matrix for visualizing multidimensional multivariate data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2606– 2614, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ingram:2009:GMM

[IMO09] Stephen Ingram, Tamara Munzner, and Marc Olano.

Glimmer: Multilevel MDS on the GPU. IEEE Transactions on Visualization and Computer Graphics, 15(2):249– 261, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Iwai:2015:EDF

[IMS15] D. Iwai, S. Mihara, and K. Sato. Extended depthof-field projector by fast focal sweep projection. IEEE Transactions on Visualization and Computer Graphics, 21 (4):462–470, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ivson:2018:CNV

[INCB18] P. Ivson, D. Nascimento, W. Celes, and S. D. Barbosa. CasCADe: A novel 4D visualization system for virtual construction planning. IEEE Transactions on Visualization and Computer Graphics, 24 (1):687–697, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Insley:2007:RVH

[IPD^{+}07] Joseph A. Insley, Michael E. Papka, Suchuan Dong, George Karniadakis, and Nicholas T. Karonis. Runtime visualization of the human arterial tree. IEEE Transactions on Visualization and Computer Graphics, 13(4):810– 821, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Irlitti:2019:CSA

[IPJT19] A. Irlitti, T. Piumsomboon, D. Jackson, and B. H. Thomas. Conveying spatial awareness cues in xR collaborations. IEEE Transactions on Visualization and Computer Graphics, 25(11):3178– 3189, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Islam:2007:VSA

[ISC07] Shoukat Islam, Deborah Silver, and Min Chen. Volume splitting and its applications. IEEE Transactions on Visualization and Computer Graphics, 13(2):193– 203, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ip:2011:SAN

[IV11] Cheuk Yiu Ip and Amitabh Varshney. Saliency-assisted navigation of very large landscape images. IEEE Transactions on Visualization and Computer Graphics, 17(12): 1737–1746, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ip:2012:HEV

[IVJ12] Cheuk Yiu Ip, Amitabh Varshney, and Joseph JaJa. Hierarchical exploration of volumes using multilevel segmentation of the intensitygradient histograms. IEEE Transactions on Visualization and Computer Graphics, 18(12):2355–2363, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ibrahim:2018:SSN

[IWR^{+}18] M. Ibrahim, P. Wickenhäuser, P. Rautek, G. Reina, and M. Hadwiger. Screen-space normal distribution function caching for consistent multiresolution rendering of large particle data. IEEE Transactions on Visualization and Computer Graphics, 24(1): 944–953, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ivanov:2007:VHL

[IWSK07] Yuri Ivanov, Christopher Wren, Alexander Sorokin, and Ishwinder Kaur. Visualizing the history of living spaces. IEEE Transactions on Visualization and Computer Graphics, 13(6):1153–1160, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

comp/trans/tg/2007/06/v1153s.
zip.

Itoh:2004:HDV

[IYIK04] Takayuki Itoh, Yumi Yamaguchi, Yuko Ikehata, and Yasumasa Kajinaga. Hierarchical data visualization using a fast rectangle-packing algorithm. IEEE Transactions on Visualization and Computer Graphics, 10(3):302– 313, May/June 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0302abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0302.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0302.pdf.

Itoh:2001:FIG

[IYK01] T. Itoh, Y. Yamaguchi, and K. Koyamada. Fast isosurface generation using the volume thinning algorithm. IEEE Transactions on Visualization and Computer Graphics, 7(1): 32–??, January/March 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0032.
pdf;   http://www.computer.
org/tvcg/tg2001/v0032abs.
htm.

Iwai:2013:VMP

[IYS13] Daisuke Iwai, Tatsunori Yabiki, and Kosuke Sato.

View management of projected labels on nonplanar and textured surfaces. IEEE Transactions on Visualization and Computer Graphics, 19 (8):1415–1424, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Igouchkine:2018:MMV

[IZM18] Oleg Igouchkine, Yubo Zhang, and Kwan-Liu Ma. Multimaterial volume rendering with a physically-based surface reflection model. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3147–3159, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8226853/.

Jianu:2018:DMT

[JA18] Radu Jianu and Sayeed Safayet Alam. A data model and task space for data of interest (DOI) eye-tracking analyses. IEEE Transactions on Visualization and Computer Graphics, 24(3):1232–1245, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jeunet:2018:DYF

[JAAL18] Camille Jeunet, Louis Albert, Ferran Argelaguet, and Anatole Lecuyer. “Do you feel in control?”: Towards novel approaches to characterise, manipulate and measure the sense of agency in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 24(4):1486–1495, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260944-
abs.html.

Johnsen:2014:MRV

[JAM^{+}14] Kyle Johnsen, Sun Joo Ahn, James Moore, Scott Brown, Thomas P. Robertson, Amanda Marable, and Aryabrata Basu. Mixed reality virtual pets to reduce childhood obesity. IEEE Transactions on Visualization and Computer Graphics, 20 (4):523–530, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jauregui:2014:TPH

[JAO^{+}14] David Antonio Gomez Jauregui, Ferran Argelaguet, Anne-Helene Olivier, Maud Marchal, Franck Multon, and Anatole Lecuyer. Toward “pseudo-haptic avatars”: Modifying the visual animation of self-avatar can simulate the perception of weight lifting. IEEE Transactions on Visualization and Computer Graphics, 20(4):654–661, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jeong:2009:SIS

[JBH^{+}09] Won-Ki Jeong, Johanna Beyer, Markus Hadwiger, Amelio Vazquez, Hanspeter Pfister, and Ross T. Whitaker. Scalable and interactive segmentation and visualization of neural processes in EM datasets. IEEE Transactions on Visualization and Computer Graphics, 15(6):1505– 1514, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Janicke:2009:VEC

[JBMS09] Heike Janicke, Michael Bottinger, Uwe Mikolajewicz, and Gerik Scheuermann. Visual exploration of climate variability changes using wavelet analysis. IEEE Transactions on Visualization and Computer Graphics, 15(6):1375– 1382, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jones:2006:DFS

[JBS06] Mark W. Jones, J. Andreas Baerentzen, and Milos Sramek. 3D distance fields: a survey of techniques and applications. IEEE Transactions on Visualization and Computer Graphics, 12(4):581– 599, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Janicke:2008:BAC

[JBS08] Heike Jänicke, Michael Böttinger, and Gerik Scheuermann. Brushing of attribute clouds for the visualization of multivariate data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1459–1466, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Janeh:2018:AGP

[JBS^{+}18] Omar Janeh, Gerd Bruder, Frank Steinicke, Alessandro Gulberti, and Monika Poetter-Nerger. Analyses of gait parameters of younger and older adults during (non-)isometric virtual walking. IEEE Transactions on Visualization and Computer Graphics, 24(10):2663– 2674, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08103804-
abs.html.

Joia:2011:LAM

[JCC^{+}11] Paulo Joia, Danilo Coimbra, Jose A. Cuminato, Fernando V. Paulovich, and Luis G. Nonato. Local affine multidimensional projection. IEEE Transactions on Visualization and Computer Graphics, 17(12):2563– 2571, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Johnson:2008:IVA

[JCG08] Gregory P. Johnson, Victor M. Calo, and Kelly P. Gaither. Interactive visualization and analysis of transitional flow. IEEE Transactions on Visualization and Computer Graphics, 14(6):1420–1427, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Joshi:2009:CSV

[JCRS09] Alark Joshi, Jesus Caban, Penny Rheingans, and Lynn Sparling. Case study on visualizing hurricanes using illustration-inspired techniques. IEEE Transactions on Visualization and Computer Graphics, 15(5): 709–718, September/October 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ju:2014:RPT

[JCWD14] Tao Ju, Minxin Cheng, Xu Wang, and Ye Duan. A robust parity test for extracting parallel vectors in 3D. IEEE Transactions on Visualization and Computer Graphics, 20(12):2526– 2534, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875965-
abs.html.

Jansen:2013:IMV

[JD13] Yvonne Jansen and Pierre Dragicevic. An interaction model for visualizations beyond the desktop. IEEE Transactions on Visualization and Computer Graphics, 19(12):2396–2405, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Johnson:2011:CIR

[JDA^{+}11] Micah K. Johnson, Kevin Dale, Shai Avidan, Hanspeter Pfister, William T. Freeman, and Wojciech Matusik. CG2Real: Improving the realism of computer generated images using a large collection of photographs. IEEE Transactions on Visualization and Computer Graphics, 17(9):1273–1285, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jianu:2009:EDF

[JDL09] Radu Jianu, Cagatay Demiralp, and David Laidlaw. Exploring 3D DTI fiber tracts with linked 2D representations. IEEE Transactions on Visualization and Computer Graphics, 15(6):1449– 1456, November/December 2009. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jianu:2012:EBC

[JDL12] Radu Jianu, Çağatay Demiralp, and David H. Laidlaw. Exploring brain connectivity with two-dimensional neural maps. IEEE Transactions on Visualization and Computer Graphics, 18(6): 978–987, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jallepalli:2018:TFQ

[JDSR^{+}18] A. Jallepalli, J. Docampo-Sánchez, J. K. Ryan, R. Haimes, and R. M. Kirby. On the treatment of field quantities and elemental continuity in FEM solutions. IEEE Transactions on Visualization and Computer Graphics, 24(1): 903–912, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Javed:2013:SZM

[JE13] Waqas Javed and Niklas Elmqvist. Stack zooming for multifocus interaction in skewed-aspect visual spaces. IEEE Transactions on Visualization and Computer Graphics, 19(8):1362–1374, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jang:2012:TVD

[JEG12] Yun Jang, David S. Ebert, and Kelly Gaither. Timevarying data visualization using functional representations. IEEE Transactions on Visualization and Computer Graphics, 18(3):421– 433, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jobard:2002:LEA

[JEH02] Bruno Jobard, Gordon Erlebacher, and M. Yousuff Hussaini. Lagrangian–Eulerian advection of noise and dye textures for unsteady flow visualization. IEEE Transactions on Visualization and Computer Graphics, 8(3):211– 222, July/September 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/03/v0211abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/03/
v0211.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/03/v0211.pdf.

Jenny:2012:ACM

[Jen12]  Bernhard Jenny.  Adap-

tive composite map projections. IEEE Transactions on Visualization and Computer Graphics, 18(12):2575– 2582, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jang:2016:MVA

[JER16] S. Jang, N. Elmqvist, and K. Ramani. MotionFlow: Visual abstraction and aggregation of sequential patterns in human motion tracking data. IEEE Transactions on Visualization and Computer Graphics, 22(1):21–30, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Johansson:2016:EPC

[JF16] J. Johansson and C. Forsell. Evaluation of parallel coordinates: Overview, categorization and guidelines for future research. IEEE Transactions on Visualization and Computer Graphics, 22(1): 579–588, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jones:2010:DWC

[JFBB10] Michael D. Jones, McKay Farley, Joseph Butler, and Matthew Beardall. Directable weathering of concave rock using curvature estimation. IEEE Transactions on Visualization and Computer Graphics, 16(1):81–94, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Janicke:2016:IVP

[JFS16] S. Janicke, J. Focht, and G. Scheuermann. Interactive visual profiling of musicians. IEEE Transactions on Visualization and Computer Graphics, 22(1):200–209, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jackle:2016:TMP

[JFSK16] D. Jackle, F. Fischer, T. Schreck, and D. A. Keim. Temporal MDS plots for analysis of multivariate data. IEEE Transactions on Visualization and Computer Graphics, 22(1): 141–150, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jeong:2007:IVV

[JFTW07] Won-Ki Jeong, P. Thomas Fletcher, Ran Tao, and Ross Whitaker. Interactive visualization of volumetric white matter connectivity in DT-MRI using a parallelhardware Hamilton–Jacobi solver. IEEE Transactions on Visualization and Computer Graphics, 13(6):1480–1487, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1480s.
mov.

Jonsson:2016:IEV

[JFY16] D. Jonsson, M. Falk, and A. Ynnerman. Intuitive exploration of volumetric data using dynamic galleries. IEEE Transactions on Visualization and Computer Graphics, 22 (1):896–905, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jayaraman:2018:GCW

[JFZ^{+}18] Pradeep Kumar Jayaraman, Chi-Wing Fu, Jianmin Zheng, Xueting Liu, and Tien-Tsin Wong. Globally consistent wrinkle-aware shading of line drawings. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2103–2117, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07930445-
                                        abs.html.

Johnson:2009:DDV

[JH09] C. Ryan Johnson and Jian Huang. Distribution-driven visualization of volume data. IEEE Transactions on Visualization and Computer Graphics, 15(5):734–746, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jakobsen:2013:IVL

[JH13] Mikkel R. Jakobsen and Kasper Hornbæk. Interactive visualizations on large and small displays: The interrelation of display size, information space, and scale. IEEE Transactions on Visualization and Computer Graphics, 19(12):2336–2345, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jansen:2016:PIS

[JH16] Y. Jansen and K. Hornbak.

A psychophysical investigation of size as a physical variable. IEEE Transactions on Visualization and Computer Graphics, 22(1):479–488, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jones:2019:OPR

[JHBK19] J. Adam Jones, Jonathan E.

Hopper, Mark T. Bolas, and David M. Krum. Orientation perception in real and virtual environments. IEEE Transactions on Visualization and Computer Graphics, 25(5):2050–2060, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642384/.

Jia:2008:VSO

[JHGH08] Yuntao Jia, Jared Hoberock, Michael Garland, and John Hart. On the visualization of social and other scale-free networks. IEEE Transactions on Visualization and Computer Graphics, 14(6):1285– 1292, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Janicke:2010:VGP

[JHH^{+}10] Stefan Janicke, Christian Heine, Marc Hellmuth, Peter F. Stadler, and Gerik Scheuermann. Visualization of graph products. IEEE Transactions on Visualization and Computer Graphics, 16(6):1082–1089, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jakobsen:2013:IVP

[JHKH13] Mikkel R. Jakobsen, Yonas Sahlemariam Haile, Soren Knudsen, and Kasper Hornbæk. Information visualization and proxemics: Design opportunities and empirical findings. IEEE Transactions on Visualization and Computer Graphics, 19(12):2386– 2395, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jo:2014:LIV

[JHP^{+}14] Jaemin Jo, Jaeseok Huh, Jonghun Park, Bohyoung Kim, and Jinwook Seo. Live-Gantt: Interactively visualizing a large manufacturing schedule. IEEE Transactions on Visualization and Computer Graphics, 20(12):2329– 2338, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875942-
abs.html.

Jo:2010:PDM

[JhR10] Hyungeun Jo and Jung hee Ryu. Placegram: a diagrammatic map for personal geotagged data browsing. IEEE Transactions on Visualization and Computer Graphics, 16(2):221–234, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jiang:2014:FFS

[JHW^{+}14] Tengfei Jiang, Jin Huang, Yuanzhen Wang, Yiying Tong, and Hujun Bao. Frame field singularity correction for automatic hexahedralization. IEEE Transactions on Visualization and Computer Graphics, 20(8):1189–1199, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Johansson:2009:IDR

[JJ09] Sara Johansson and Jimmy Johansson. Interactive dimensionality reduction through user-defined combinations of quality metrics. IEEE Transactions on Visualization and Computer Graphics, 15(6):993–1000, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jackson:2016:LUR

[JK16] B. Jackson and D. F. Keefe. Lift-off: Using reference imagery and freehand sketching to create 3D models in VR. IEEE Transactions on Visualization and Computer Graphics, 22(4):1442– 1451, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jankun-Kelly:2006:VVP

[JKJTM06] Monika Jankun-Kelly, Ming Jiang, David Thompson, and Raghu Machiraju. Vortex visualization for practical engineering applications. IEEE Transactions on Visualization and Computer Graphics, 12(5):957–964, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jin:2008:DSR

[JKLG08] Miao Jin, Junho Kim, Feng Luo, and Xianfeng Gu. Discrete surface Ricci flow. IEEE Transactions on Visualization and Computer Graphics, 14(5):1030–1043, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jankun-Kelly:2001:VEE

[JKM01] T. J. Jankun-Kelly and Kwan-Liu Ma. Visualization exploration and encapsulation via a spreadsheetlike interface. IEEE Transactions on Visualization and Computer Graphics, 7(3): 275–287, July 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0275.
pdf;   http://www.computer.
org/tvcg/tg2001/v0275abs.
htm.

Jankun-Kelly:2006:SBR

[JKM06] T. J. Jankun-Kelly and Ketan Mehta. Superellipsoid-based, real symmetric traceless tensor glyphs motivated by nematic liquid crystal alignment visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1197–1204, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jankun-Kelly:2007:MFV

[JKMG07] T. J. Jankun-Kelly, Kwan-Liu Ma, and Michael Gertz. A model and framework for visualization exploration. IEEE Transactions on Visualization and Computer Graphics, 13(2):357–369, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jonsson:2012:HEI

[JKRY12] Daniel Jonsson, Joel Kronander, Timo Ropinski, and Anders Ynnerman. Historygrams: Enabling interactive global illumination in direct volume rendering using photon mapping. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2364–2371, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Julier:2015:GEI

[JLS15] S. Julier, R. Lindeman, and C. Sandor. Guest Editor’s introduction to the special section on the IEEE International Symposium on Mixed and Augmented Reality 2014. IEEE Transactions on Visualization and Computer Graphics, 21(12):1321– 1322, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jones:2010:VFT

[JM10] Chad Jones and Kwan-Liu Ma. Visualizing flow trajectories using locality-based rendering and warped curve plots. IEEE Transactions on Visualization and Computer Graphics, 16(6):1587– 1594, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Javed:2010:GPM

[JME10] Waqas Javed, Bryan McDonnel, and Niklas Elmqvist. Graphical perception of multiple time series. IEEE Transactions on Visualization and Computer Graphics, 16(6):927–934, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jang:2015:FCC

[JNC^{+}15] Youngkyoon Jang, Seung-Tak Noh, Hyung Jin Chang, Tae-Kyun Kim, and Woontack Woo. 3D finger CAPE: Clicking action and position estimation under self-occlusions in egocentric viewpoint. IEEE Transactions on Visualization and Computer Graphics, 21 (4):501–510, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jadhav:2019:FVE

[JNK19] S. Jadhav, S. Nadeem, and A. Kaufman. FeatureLego: Volume exploration using exhaustive clustering of supervoxels. IEEE Transactions on Visualization and Computer Graphics, 25(9):2725– 2737, September 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jiang:2019:JAV

[JOR^{+}19] Y. Jiang, E. E. O’Neal, P. Rahimian, J. P. Yon, J. M. Plumert, and J. K. Kearney. Joint action in a virtual environment: Crossing roads with risky vs. safe human and agent partners. IEEE Transactions on Visualization and Computer Graphics, 25(10):2886–2895, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Joy:2002:GEI

[Joy02] Kenneth I. Joy. Guest Editor’s introduction: Special issue on IEEE Visualization. IEEE Transactions on Visualization and Computer Graphics, 8(3):209–210, July/September 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2002/03/v0209.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/03/
v0209.pdf.

John:2018:IVV

[JPD^{+}18] Nigel W. John, Serban R. Pop, Thomas W. Day, Panagiotis D. Ritsos, and Christopher J. Headleand. The implementation and validation of a virtual environment for training powered wheelchair manoeuvres. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1867–1878, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07917262-
abs.html.

Jerman:2016:BEV

[JPLŠ16] Tim Jerman, Franjo Pernuš, Boštjan Likar, and Žiga Špiclin. Blob enhancement and visualization for improved intracranial aneurysm detection. IEEE Transactions on Visualization and Computer Graphics, 22(6):1705– 1717, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Joshi:2008:EVC

[JQD^{+}08] Alark Joshi, Xiaoning Qian, Donald Dione, Ketan Bulsara, Christopher Breuer, Albert Sinusas, and Xenophon Papademetris. Effective visualization of complex vascular structures using a nonparametric vessel detection method. IEEE Transactions on Visualization and Computer Graphics, 14(6):1603– 1610, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jalba:2007:ESR

[JR07] Andrei C. Jalba and Jos B. T. M. Roerdink. Efficient surface reconstruction using generalized Coulomb potentials. IEEE Transactions on Visualization and Computer Graphics, 13(6):1512–1519, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jianu:2014:HDG

[JRHT14] Radu Jianu, Adrian Rusu, Yifan Hu, and Douglas Taggart. How to display group information on node-link diagrams: An evaluation. IEEE Transactions on Visualization and Computer Graphics, 20(11): 1530–1541, November 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/11/06787045-
                                        abs.html.

Jerding:1998:IMT

[JS98] D. F. Jerding and J. T. Stasko. The information mural: a technique for displaying and navigating large information spaces. IEEE Transactions on Visualization and Computer Graphics, 4(3):257– 271, July/September 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0257.

pdf;   http://www.computer.
org/tvcg/tg1998/v0257abs.
htm.

Ji:2006:DVS

[JS06] Guangfeng Ji and Han-Wei Shen. Dynamic view selection for time-varying volumes. IEEE Transactions on Visualization and Computer Graphics, 12(5):1109–1116, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jones:2013:PSE

[JSB13] J. Adam Jones, J. Edward Swan II, and Mark Bolas. Peripheral stimulation and its effect on perceived spatial scale in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 19 (4):701–710, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jackowski:2003:ADS

[JSG03] Marcel Jackowski, Martin Satter, and Ardeshir Goshtasby. Approximating digital 3D shapes by rational Gaussian surfaces. IEEE Transactions on Visualization and Computer Graphics, 9(1):56– 69, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0056abs.

                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/01/
                                        v0056.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/01/v0056.pdf.

Ji:2019:VEN

[JSR^{+}19] Xiaonan Ji, Han-Wei Shen, Alan Ritter, Raghu Machiraju, and Po-Yin Yen. Visual exploration of neural document embedding in information retrieval: Semantics and feature selection. IEEE Transactions on Visualization and Computer Graphics, 25(6):2181–2192, June 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8667702/.

Jeong:2010:IHL

[JST^{+}10] Won-Ki Jeong, Jens Schneider, Stephen Turney, Beverly E. Faulkner-Jones, Dominik Meyer, Rudiger Westermann, R. Clay Reid, Jeff Lichtman, and Hanspeter Pfister. Interactive histology of large-scale biomedical image stacks. IEEE Transactions on Visualization and Computer Graphics, 16(6):1386–1395, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Joshi:2008:NIT

[JSV^{+}08] Alark Joshi, Dustin Scheinost, Kenneth Vives, Dennis Spencer, Lawrence Staib, and Xenophon Papademetris. Novel interaction techniques for neurosurgical planning and stereotactic navigation. IEEE Transactions on Visualization and Computer Graphics, 14(6):1587–1594, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jang:2009:IVR

[JV09] Yun Jang and Ugo Varetto.

Interactive volume rendering of functional representations in quantum chemistry. IEEE Transactions on Visualization and Computer Graphics, 15(6):1579–5186, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jo:2019:DRM

[JVDF19] Jaemin Jo, Frederic Vernier, Pierre Dragicevic, and Jean-Daniel Fekete. A declarative rendering model for multiclass density maps. IEEE Transactions on Visualization and Computer Graphics, 25 (1):470–480, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440817-
abs.html.

Jalba:2013:FSL

[JvdLR13] A. C. Jalba, Wladimir J. van der Laan, and J. B. T. M. Roerdink. Fast sparse level sets on graphics hardware. IEEE Transactions on Visualization and Computer Graphics, 19(1):30–44, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Johnson:2005:SDF

[JWC05] David E. Johnson, Peter Willemsen, and Elaine Cohen. Six degree-of-freedom haptic rendering using spatialized normal cone search. IEEE Transactions on Visualization and Computer Graphics, 11(6):661–670, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jarabo:2014:EAF

[JWD^{+}14] Adrian Jarabo, Hongzhi Wu, Julie Dorsey, Holly Rushmeier, and Diego Gutierrez. Effects of approximate filtering on the appearance of bidirectional texture functions. IEEE Transactions on Visualization and Computer Graphics, 20(6):880–892, June 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jorissen:2005:DIP

[JWL05] Pieter Jorissen, Maarten Wijnants, and Wim Lamotte. Dynamic interactions in physically realistic collaborative virtual environments. IEEE Transactions on Visualization and Computer Graphics, 11(6):649–660, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jaynes:2004:CBD

[JWS04] Christopher Jaynes, Stephen Webb, and R. Matt Steele. Camera-based detection and removal of shadows from interactive multiprojector displays. IEEE Transactions on Visualization and Computer Graphics, 10(3):290– 301, May/June 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0290abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0290.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0290.pdf.

Janicke:2007:MVU

[JWSK07] Heike Jänicke, Alexander Wiebel, Gerik Scheuermann, and Wolfgang Kollmann. Multifield visualization using local statistical complexity. IEEE Transactions on Visualization and Computer Graphics, 13(6):1384–1391, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1384s.
                                        zip.

Jonsson:2017:CPM

[JY17] Daniel Jönsson and Anders Ynnerman. Correlated photon mapping for interactive global illumination of time-varying volumetric data. IEEE Transactions on Visualization and Computer Graphics, 23(1):901–910, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jianu:2010:VIQ

[JYC^{+}10] Radu Jianu, Kebing Yu, Lulu Cao, Vinh Nguyen, Arthur R. Salomon, and David H. Laidlaw. Visual integration of quantitative proteomic data, pathways, and protein interactions. IEEE Transactions on Visualization and Computer Graphics, 16(4):609– 620, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jin:2009:CTS

[JZLG09] Miao Jin, Wei Zeng, Feng Luo, and Xianfeng Gu. Computing Teichmüller shape space. IEEE Transactions on Visualization and Computer Graphics, 15(3):504– 517, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kong:2012:GOU

[KA12] Nicholas Kong and Maneesh Agrawala. Graphical overlays: Using layered elements to aid chart reading. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2631–2638, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwatra:2007:TF

[KAK^{+}07] Vivek Kwatra, David Adalsteinsson, Theodore Kim, Nipun Kwatra, Mark Carlson, and Ming Lin. Texturing fluids. IEEE Transactions on Visualization and Computer Graphics, 13(5): 939–952, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2007/
05/v0939s.mpg.

Klein:2018:ICV

[KAK^{+}18] T. Klein, L. Autin, B. Kozlíková, D. S. Goodsell, A. Olson, M. E. Gröller, and I. Viola. Instant construction and visualization of crowded biological environments. IEEE Transactions on Visualization and Computer Graphics, 24(1): 862–872, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kahng:2018:AVE

[KAKC18] Minsuk Kahng, Pierre Y. Andrews, Aditya Kalro, and Duen Horng (Polo) Chau. ActiVis: Visual exploration of industry-scale deep neural network models. IEEE Transactions on Visualization and Computer Graphics, 24 (1):88–97, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keefe:2008:SSC

[KAM^{+}08] Daniel F. Keefe, Daniel Acevedo, Jadrian Miles, Fritz Drury, Sharon M. Swartz, and David H. Laidlaw. Scientific sketching for collaborative VR visualization design. IEEE Transactions on Visualization and Computer Graphics, 14(4):835–847, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kasik:2012:BSW

[Kas12] David J. Kasik. Banquet speaker: What’s next?: The third wave in computer graphics and interactive techniques. IEEE Transactions on Visualization and Computer Graphics, 18(4):xiii–xiv, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kaufman:1995:Eb

[Kau95]  Arie E. Kaufman.  Edi-

torial. IEEE Transactions on Visualization and Computer Graphics, 1(4):289– ??, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0289.
pdf.

Kaufman:1996:Ea

[Kau96a] Arie E. Kaufman. Editorial.

IEEE Transactions on Visualization and Computer Graphics, 2(1):1–??, March 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0001.
pdf.

Kaufman:1996:Eb

[Kau96b]  Arie E. Kaufman.  Edi-

torial. IEEE Transactions on Visualization and Computer Graphics, 2(3):184– ??, September 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0185.
pdf.

Kaufman:1996:Ec

[Kau96c]  Arie E. Kaufman.  Edi-

torial. IEEE Transactions on Visualization and Computer Graphics, 2(4):281– ??, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0281.
pdf.

Kaufman:1997:E

[Kau97] A. E. Kaufman. Editorial. IEEE Transactions on Visualization and Computer Graphics, 3(2):97, April/June 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0097.
pdf.

Kaufman:1998:ECE

[Kau98] Arie E. Kaufman. Editor-in-Chief’s editorial. IEEE Transactions on Visualization and Computer Graphics, 4(4):289– ??, October/December 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0289.
pdf.

Kellner:2012:GCH

[KBB^{+}12] Falko Kellner, Benjamin Bolte, Gerd Bruder, Ulrich Rautenberg, Frank Steinicke, Markus Lappe, and Reinhard Koch. Geometric calibration of head-mounted displays and its effects on distance estimation. IEEE Transactions on Visualization and Computer Graphics, 18(4): 589–596, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2018:RTA

[KBB^{+}18] Kangsoo Kim, Mark Billinghurst, Gerd Bruder, Henry Been-Lirn Duh, and Gregory F. Welch. Revisiting trends in augmented reality research: a review of the 2nd decade of ISMAR (2008– 2017). IEEE Transactions on Visualization and Computer Graphics, 24(11):2947– 2962, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8456568/.

Kagaya:2011:VPS

[KBD^{+}11] Mizuki Kagaya, William Brendel, Qingqing Deng, Todd Kesterson, Sinisa Todorovic, Patrick J. Neill, and Eugene Zhang. Video painting with space–time-varying style parameters. IEEE Transactions on Visualization and Computer Graphics, 17(1): 74–87, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krone:2009:IVM

[KBE09] Michael Krone, Katrin Bidmon, and Thomas Ertl. Interactive visualization of molecular surface dynamics. IEEE Transactions on Visualization and Computer Graphics, 15(6):1391–1398, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Karch:2018:VAI

[KBE^{+}18] Grzegorz Karol Karch, Fabian Beck, Moritz Ertl, Christian Meister, Kathrin Schulte, Bernhard Weigand, Thomas Ertl, and Filip Sadlo. Visual analysis of inclusion dynamics in two-phase flow. IEEE Transactions on Visualization and Computer Graphics, 24(5):1841–1855, May 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07895170-
abs.html.

Koch:2011:IIV

[KBGE11] Steffen Koch, Harald Bosch, Mark Giereth, and Thomas Ertl. Iterative integration of visual insights during scalable patent search and analysis. IEEE Transactions on Visualization and Computer Graphics, 17(5):557–569, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kosara:2006:PSI

[KBH06] Robert Kosara, Fabian Bendix, and Helwig Hauser. Parallel sets: Interactive exploration and visual analysis of categorical data. IEEE Transactions on Visualization and Computer Graphics, 12(4):558– 568, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kok:2010:APR

[KBH^{+}10] Peter Kok, Martin Baiker, Emile A. Hendriks, Frits H. Post, Jouke Dijkstra, Clemens W. G. M. Lowik, Boudewijn P. F. Lelieveldt, and Charl P. Botha. Articulated planar reformation for change visualization in small animal imaging. IEEE Transactions on Visualization and Computer Graphics, 16(6):1396–1404, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kratz:2013:ASP

[KBH13] A. Kratz, D. Baum, and I. Hotz. Anisotropic sampling of planar and twomanifold domains for texture generation and glyph distribution. IEEE Transactions on Visualization and Computer Graphics, 19(11):1782– 1794, November 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2018:VNN

[KBI^{+}18] N. W. Kim, B. Bach, H. Im, S. Schriber, M. Gross, and H. Pfister. Visualizing nonlinear narratives with story curves. IEEE Transactions on Visualization and Computer Graphics, 24(1):595–604, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krstajic:2011:CCD

[KBK11] Milos Krstajic, Enrico Bertini, and Daniel Keim. Cloud-Lines: Compact display of event episodes in multiple time-series. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2432–2439, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kohlmann:2007:LDV

[KBKG07] Peter Kohlmann, Stefan Bruckner, Armin Kanitsar, and Eduard Gröller. LiveSync: Deformed viewing spheres for knowledgebased navigation. IEEE Transactions on Visualization and Computer Graphics, 13(6):1544–1551, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1544s.
                                        zip.

Kappe:2019:EVW

[KBL19] C. P. Kappe, M. Böttinger, and H. Leitte. Exploring variability within ensembles of decadal climate predictions. IEEE Transactions on Visualization and Computer Graphics, 25(3):1499–1512, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kilteni:2013:DIV

[KBS13] Konstantina Kilteni, Ilias Bergstrom, and Mel Slater. Drumming in immersive virtual reality: The body shapes the way we play. IEEE Transactions on Visualization and Computer Graphics, 19(4): 597–605, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kolesar:2017:FCC

[KBVH17] Ivan Kolesár, Stefan Bruckner, Ivan Viola, and Helwig Hauser. A fractional Cartesian composition model for semi-spatial comparative visualization design. IEEE Transactions on Visualization and Computer Graphics, 23 (1):851–860, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koren:2004:RLD

[KC04] Yehuda Koren and Liran Carmel. Robust linear dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 10(4):459–470, July/ August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        dl/trans/tg/2004/04/v0459.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/04/
                                        v0459.pdf.

Kondo:2014:DET

[KC14] Brittany Kondo and Christopher Collins. DimpVis: Exploring time-varying information visualizations by direct manipulation. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2003–2012, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875985-
                                        abs.html.

Kim:2016:MPG

[KCA16] Young Min Kim, Junghyun Cho, and Sang Chul Ahn. 3D modeling from photos given topological information. IEEE Transactions on Visualization and Computer Graphics, 22(9):2070–2081, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kuo:2017:GAF

[KCC^{+}17] Ying-Miao Kuo, Hung-Kuo Chu, Ming-Te Chi, Ruen-Rone Lee, and Tong-Yee Lee.

Generating ambiguous figureground images. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1534–1545, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07420730-
abs.html.

Kim:2011:RRE

[KCH11] S. Kim, C. Coffin, and T. Hollerer. Robust relocalization and its evaluation for online environment map construction. IEEE Transactions on Visualization and Computer Graphics, 17(7): 875–887, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Karnick:2010:RVU

[KCJ^{+}10] Pushpak Karnick, David Cline, Stefan Jeschke, Anshuman Razdan, and Peter Wonka. Route visualization using detail lenses. IEEE Transactions on Visualization and Computer Graphics, 16(2):235–247, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kouril:2019:LLL

[KCK^{+}19a] David Kouril, Ladislav Cmolik, Barbora Kozlikova, Hslanc-Yun Wu, Graham Johnson, David S. Goodsell, Arthur Olson, M. Eduard Groller, and Ivan Viola. Labels on levels: Labeling of multi-scale multiinstance and crowded 3D biological environments. IEEE Transactions on Visualization and Computer Graphics, 25 (1):977–986, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440077.
pdf.

Kwon:2019:RVA

[KCK^{+}19b] Bum Chul Kwon, Min-Je Choi, Joanne Taery Kim, Edward Choi, Young Bin Kim, Soonwook Kwon, Jimeng Sun, and Jaegul Choo. RetainVis: Visual analytics with interpretable and interactive recurrent neural networks on electronic medical records. IEEE Transactions on Visualization and Computer Graphics, 25 (1):299–309, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440842-
abs.html.

Kwon:2018:WWG

[KCM18] O. H. Kwon, T. Crnovrsanin, and K. L. Ma. What would a graph look like in this layout? A machine learning approach to large graph visualization. IEEE Transactions on Visualization and Computer Graphics, 24(1):478–488, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kadosh:2003:TID

[KCOY03] Arie Kadosh, Daniel Cohen-Or, and Roni Yagel. Tricubic interpolation of discrete surfaces for binary volumes. IEEE Transactions on Visualization and Computer Graphics, 9(4):580–586, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0580abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0580.pdf.

Kim:2016:ISS

[KCPE16] H. Kim, J. Choo, H. Park, and A. Endert. InterAxis: Steering scatterplot axes via observation-level interaction. IEEE Transactions on Visualization and Computer Graphics, 22(1):131–140, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwon:2008:TCM

[KCPS08] Taesoo Kwon, Young-Sang Cho, Sang I. Park, and Sung Yong Shin. Twocharacter motion analysis and synthesis. IEEE Transactions on Visualization and Computer Graphics, 14(3):707– 720, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/;   http://csdl.
                                        computer.org/comp/trans/
                                        tg/2008/03/ttg2008030707s.
                                        zip.

Kindlmann:2016:DDS

[KCS^{+}16] G. Kindlmann, C. Chiw, N. Seltzer, L. Samuels, and J. Reppy. Diderot: a domainspecific language for portable parallel scientific visualization and image analysis. IEEE Transactions on Visualization and Computer Graphics, 22 (1):867–876, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2017:SJG

[KCT^{+}17] MyoungGon Kim, Sunglk Cho, Tanh Quang Tran, Seong-Pil Kim, Ohung Kwon, and JungHyun Han. Scaled jump in gravity-reduced virtual environments. IEEE Transactions on Visualization and Computer Graphics, 23(4):1360–1368, April 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/04/07829423-
                                        abs.html.

Kyprianidis:2013:SAT

[KCWI13] Jan Eric Kyprianidis, John Collomosse, Tinghuai Wang, and Tobias Isenberg. State of the “art”: A taxonomy of artistic stylization techniques for images and video. IEEE Transactions on Visualization and Computer Graphics, 19 (5):866–885, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kristensson:2009:EST

[KDA^{+}09] Per Ola Kristensson, Nils Dahlbäck, Daniel Anundi, Marius Björnstad, Hanna Gillberg, Jonas Haraldsson, Ingrid Mårtensson, Mathias Nordvall, and Josefine Ståhl. An evaluation of space time cube representation of spatiotemporal patterns. IEEE Transactions on Visualization and Computer Graphics, 15(4):696–702, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koschier:2017:ADA

[KDBB17] Dan Koschier, Crispin Deul, Magnus Brand, and Jan Bender. hp-adaptive discretization algorithm for signed distance field generation: An hp-adaptive discretization algorithm for signed distance field generation. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2208–2221, October 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/10/07987773-
                                        abs.html.

Kong:2016:MPV

[KDM^{+}16] Yan Kong, Weiming Dong, Xing Mei, Chongyang Ma, Tong-Yee Lee, Siwei Lyu, Feiyue Huang, and Xiaopeng Zhang. Measuring and predicting visual importance of similar objects. IEEE Transactions on Visualization and Computer Graphics, 22(12): 2564–2578, December 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2016/12/07374748-
                                        abs.html.

Kieffer:2016:HHL

[KDMW16] S. Kieffer, T. Dwyer, K. Marriott, and M. Wybrow. HOLA: Human-like orthogonal network layout. IEEE Transactions on Visualization and Computer Graphics, 22 (1):349–358, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2012:DET

[KDX^{+}12] Sung-Hee Kim, Zhihua Dong, Hanjun Xian, Benjavan Upatising, and Ji Soo Yi. Does an eye tracker tell the truth about visualizations?: Findings while investigating visualizations for decision making. IEEE Transactions on Visualization and Computer Graphics, 18(12):2421– 2430, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keim:2000:DPO

[Kei00] D. A. Keim. Designing pixeloriented visualization techniques: Theory and applications. IEEE Transactions on Visualization and Computer Graphics, 6(1):59–78, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0059.
pdf;   http://www.computer.
org/tvcg/tg2000/v0059abs.
htm.

Keim:2002:IVV

[Kei02] Daniel A. Keim. Information visualization and visual data mining. IEEE Transactions on Visualization and Computer Graphics, 8 (1):1–8, January 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0001.
pdf;   http://www.computer.
org/tvcg/tg2002/v0001abs.
htm.

Kim:2019:QEB

[KEJK19] Y. Kim, H. Eun, C. Jung, and C. Kim. A quad edgebased grid encoding model for content-aware image retargeting. IEEE Transactions on Visualization and Computer Graphics, 25(12):3202– 3215, December 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2008:BSR

[KEP08] Minho Kim, Alireza Entezari, and Jörg Peters. Box spline reconstruction on the facecentered cubic lattice. IEEE Transactions on Visualization and Computer Graphics, 14(6):1523–1530, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keefe:2009:ICM

[KERC09] Daniel Keefe, Marcus Ewert, William Ribarsky, and Remco Chang. Interactive coordinated multiple-view visualization of biomechanical motion data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1383– 1390, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwon:2018:CVS

[KEV^{+}18] B. C. Kwon, B. Eysenbach, J. Verma, K. Ng, C. De Filippi, W. F. Stewart, and A. Perer. Clustervision: Visual supervision of unsupervised clustering. IEEE Transactions on Visualization and Computer Graphics, 24(1): 142–151, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kulkarni:2015:FBS

[KFL^{+}15] Sandip D. Kulkarni, Charles J.

Fisher, Price Lefler, Aditya Desai, Shanthanu Chakravarthy, Eric R. Pardyjak, Mark A. Minor, and John M. Hollerbach. A full body steerable wind display for a locomotion interface. IEEE Transactions on Visualization and Computer Graphics, 21(10):1146– 1159, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07089286-
abs.html.

Kuchenbecker:2006:ICR

[KFN06] Katherine J. Kuchenbecker, Jonathan Fiene, and Gunter Niemeyer. Improving contact realism through event-based haptic feedback. IEEE Transactions on Visualization and Computer Graphics, 12(2): 219–230, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krone:2017:MSM

[KFS^{+}17] Michael Krone, Florian Frieß, Katrin Scharnowski, Guido Reina, Silvia Fademrecht, Tobias Kulschewski, Jürgen Pleiss, and Thomas Ertl. Molecular Surface Maps. IEEE Transactions on Visualization and Computer Graphics, 23(1):701–710, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koven:2019:LLD

[KFS^{+}19] Jay Koven, Cristian Felix, Hossein Siadati, Markus Jakobsson, and Enrico Bertini. Lessons learned developing a visual analytics solution for investigative analysis of scamming activities. IEEE Transactions on Visualization and Computer Graphics, 25(1): 225–234, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440841-
                                        abs.html.

Kumar:2006:VEC

[KG06] Gautam Kumar and Michael Garland. Visual exploration of complex timevarying graphs. IEEE Transactions on Visualization and Computer Graphics, 12(5): 805–812, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2018:DBP

[KGAM18] Hyungil Kim, Joseph L. Gabbard, Alexandre Miranda Anon, and Teruhisa Misu. Driver behavior and performance with augmented reality pedestrian collision warning: An outdoor user study. IEEE Transactions on Visualization and Computer Graphics, 24(4):1515–1524, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08302393-
abs.html.

Kerzner:2019:FCV

[KGD^{+}19] Ethan Kerzner, Sarah Goodwin, Jason Dykes, Sara Jones, and Miriah Meyer. A framework for creative visualization-opportunities workshops. IEEE Transactions on Visualization and Computer Graphics, 25(1):748– 758, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440830-
abs.html.

Krishnan:2012:ATD

[KGG^{+}12] Harinarayan Krishnan, Christoph Garth, Jens Guhring, M. Akif Gulsun, Andreas Greiser, and Kenneth I. Joy. Analysis of time-dependent flow-sensitive PC-MRI data. IEEE Transactions on Visualization and Computer Graphics, 18(6): 966–977, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krishnan:2009:TSS

[KGJ09] Hari Krishnan, Christoph Garth, and Kenneth Joy. Time and streak surfaces for flow visualization in large time-varying data sets. IEEE Transactions on Visualization and Computer Graphics, 15(6):1267–1274, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kohler:2013:SAV

[KGP^{+}13] Benjamin Kohler, Rocco Gasteiger, Uta Preim, Holger Theisel, Matthias Gutberlet, and Bernhard Preim. Semiautomatic vortex extraction in 4D PC-MRI cardiac blood flow data using line predicates. IEEE Transactions on Visualization and Computer Graphics, 19(12):2773– 2782, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krivanek:2005:RCE

[KGPB05] Jaroslav Krivanek, Pascal Gautron, Sumanta Pattanaik, and Kadi Bouatouch. Radiance caching for efficient global illumination computation. IEEE Transactions on Visualization and Computer Graphics, 11(5): 550–561, September/October 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kretschmer:2013:IPS

[KGPS13] Jan Kretschmer, Christian Godenschwager, Bernhard Preim, and Marc Stamminger. Interactive patientspecific vascular modeling with sweep surfaces. IEEE Transactions on Visualization and Computer Graphics, 19(12):2828–2837, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:1998:FCD

[KGS98] D.-J. Kim, L. J. Guibas, and S.-Y. Shin. Fast collision detection among multiple moving spheres. IEEE Transactions on Visualization and Computer Graphics, 4(3):230– 242, July/September 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0230.
pdf;   http://www.computer.
org/tvcg/tg1998/v0230abs.
htm.

Kang:2008:IER

[KGS^{+}08] Hyunmo Kang, Lise Getoor, Ben Shneiderman, Mustafa Bilgic, and Louis Licamele. Interactive entity resolution in relational data: a visual analytic tool and its evaluation. IEEE Transactions on Visualization and Computer Graphics, 14(5):999–1014, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kuhne:2012:DDA

[KGZ^{+}12] Lars Kuhne, Joachim Giesen, Zhiyuan Zhang, Sungsoo Ha, and Klaus Mueller. A data-driven approach to huepreserving color-blending. IEEE Transactions on Visualization and Computer Graphics, 18(12):2122–2129, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kreylos:2001:SAC

[KH01] O. Kreylos and B. Hamann. On simulated annealing and the construction of linear spline approximations for scattered data. IEEE Transactions on Visualization and Computer Graphics, 7(1):17– ??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0017.
pdf;   http://www.computer.
org/tvcg/tg2001/v0017abs.
htm.

Kehrer:2013:VVA

[KH13] Johannes Kehrer and Helwig Hauser. Visualization and visual analysis of multifaceted scientific data: A survey. IEEE Transactions on Visualization and Computer Graphics, 19(3):495– 513, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kay:2016:BWL

[KH16] M. Kay and J. Heer. Beyond Weber’s law: A second look at ranking visualizations of correlation. IEEE Transactions on Visualization and Computer Graphics, 22 (1):469–478, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kagami:2019:ASF

[KH19] S. Kagami and K. Hashimoto.

Animated stickies: Fast video projection mapping onto a markerless plane through a direct closed-loop alignment. IEEE Transactions on Visualization and Computer Graphics, 25(11):3094– 3104, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kong:2010:PGC

[KHA10] Nicholas Kong, Jeffrey Heer, and Maneesh Agrawala. Perceptual guidelines for creating rectangular treemaps. IEEE Transactions on Visualization and Computer Graphics, 16(6):990–998, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kaehler:2012:NAV

[KHA12] Ralf Kaehler, Oliver Hahn, and Tom Abel. A novel approach to visualizing dark matter simulations. IEEE Transactions on Visualization and Computer Graphics, 18(12):2078–2087, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keim:2002:HPB

[KHD02] Daniel A. Keim, Ming C. Hao, and Umeshwar Dayal. Hierarchical pixel bar charts. IEEE Transactions on Visualization and Computer Graphics, 8(3):255–269, July/ September 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2002/03/v0255abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2002/03/
                                        v0255.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2002/03/v0255.pdf.

Keim:2007:VCB

[KHDL07] Daniel A. Keim, Ming C. Hao, Umeshwar Dayal, and Martha Lyons. Value-cell bar charts for visualizing large transaction data sets. IEEE Transactions on Visualization and Computer Graphics, 13(4):

822–833, July/August 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kaufman:2009:AIT

[KHE09] Arie E. Kaufman, Hans Hagen, and David S. Ebert. The 15th anniversary of the IEEE Transactions on Visualization and Computer Graphics: Celebrating a success story. IEEE Transactions on Visualization and Computer Graphics, 15(5):705–706, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kurzhals:2016:GSI

[KHH^{+}16] K. Kurzhals, M. Hlawatsch, F. Heimerl, M. Burch, T. Ertl, and D. Weiskopf. Gaze stripes: Image-based visualization of eye tracking data. IEEE Transactions on Visualization and Computer Graphics, 22(1):1005–1014, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kenwright:1999:FES

[KHL99] D. N. Kenwright, C. Henze, and C. Levit. Feature extraction of separation and attachment lines. IEEE Transactions on Visualization and Computer Graphics, 5(2):135– 144, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0135.
pdf;   http://www.computer.
org/tvcg/tg1999/v0135abs.
htm.

Koo:2017:TZW

[KHLM17] Bongjin Koo, Jean Hergel, Sylvain Lefebvre, and Niloy J. Mitra. Towards zero-waste furniture design. IEEE Transactions on Visualization and Computer Graphics, 23(12): 2627–2640, December 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07762730-
abs.html.

Klosowski:1998:ECD

[KHM^{+}98] J. T. Klosowski, M. Held, J. S. B. Mitchell, H. Sowizral, and K. Zikan. Efficient collision detection using bounding volume hierarchies of k-DOPs. IEEE Transactions on Visualization and Computer Graphics, 4(1):21– 36, January/March 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0021.
pdf;   http://www.computer.
org/tvcg/tg1998/v0021abs.
htm.

Kniss:2007:IRR

[KHPS07] Joe Kniss, Warren Hunt, Kristin Potter, and Pradeep Sen. IStar: a raster representation for scalable image and volume data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1424–1431, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kern:2018:RDV

[KHS^{+}18] M. Kern, T. Hewson, F. Sadlo, R. Westermann, and M. Rautenhaus. Robust detection and visualization of jet-stream core lines in atmospheric flow. IEEE Transactions on Visualization and Computer Graphics, 24(1):893–902, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kern:2019:IVA

[KHS^{+}19] Michael Kern, Tim Hewson, Andreas Schatler, Rudiger Westermann, and Marc Rautenhaus. Interactive 3D visual analysis of atmospheric fronts. IEEE Transactions on Visualization and Computer Graphics, 25(1):1080– 1090, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440076-
abs.html.

Klinker:2011:GEI

[KHSB11] Gudrun Klinker, Tobias Hollerer, Hideo Saito, and Oliver Bimber. Guest Editors’ introduction: Special section on the IEEE International Symposium on Mixed and Augmented Reality (IS-MAR). IEEE Transactions on Visualization and Computer Graphics, 17(10):1353– 1354, October 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2004:CST

[KHSI04] Sunghee Kim, Haleh Hagh-Shenas, and Victoria Interrante. Conveying shape with texture: Experimental investigations of texture’s effects on shape categorization judgments. IEEE Transactions on Visualization and Computer Graphics, 10(4):471– 483, July/August 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/04/v0471.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0471.pdf.

Kelly:2014:RPD

[KHSS14] Jonathan W. Kelly, William W. Hammel, Zachary D. Siegel, and Lori A. Sjolund. Recalibration of perceived distance in virtual environments occurs rapidly and transfers asymmetrically across scale. IEEE Transactions on Visualization and Computer Graphics, 20 (4):588–595, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kurzhals:2017:VAM

[KHSW17] Kuno Kurzhals, Marcel Hlawatsch, Christof Seeger, and Daniel Weiskopf. Visual analytics for mobile eye tracking. IEEE Transactions on Visualization and Computer Graphics, 23 (1):301–310, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Knoll:2009:VRC

[KHW^{+}09] Aaron Knoll, Younis Hijazi, Rolf Westerteiger, Mathias Schott, Charles Hansen, and Hans Hagen. Volume ray casting with peak finding and differential sampling. IEEE Transactions on Visualization and Computer Graphics, 15(6):1571–1578, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kreiser:2018:DGE

[KHZR18] J. Kreiser, A. Hann, E. Zizer, and T. Ropinski. Decision graph embedding for highresolution manometry diagnosis. IEEE Transactions on Visualization and Computer Graphics, 24(1):873–882, January 2018. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kadaba:2007:VCS

[KIL07] Nivedita Kadaba, Pourang Irani, and Jason Leboe. Visualizing causal semantics using animations. IEEE Transactions on Visualization and Computer Graphics, 13(6):1254–1261, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2013:QBS

[Kim13] Minho Kim. Quartic boxspline reconstruction on the BCC lattice. IEEE Transactions on Visualization and Computer Graphics, 19(2): 319–330, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kincaid:2010:SFA

[Kin10]  Robert Kincaid.  Signal-

Lens: Focus+Context applied to electronic time series. IEEE Transactions on Visualization and Computer Graphics, 16(6):900–907, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kitajima:2017:SPP

[KIS17] Yuki Kitajima, Daisuke Iwai, and Kosuke Sato. Simultaneous projection and positioning of laser projector pixels. IEEE Transactions on Visualization and Computer Graphics, 23(11):2419– 2429, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007248-
abs.html.

Klehm:2014:PLM

[KISE14] Oliver Klehm, Ivo Ihrke, Hans-Peter Seidel, and Elmar Eisemann. Property and lighting manipulations for static volume stylization using a painting metaphor. IEEE Transactions on Visualization and Computer Graphics, 20 (7):983–995, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2012:PBC

[KJ12] Theodore Kim and Doug L.

James. Physics-based character skinning using multidomain subspace deformations. IEEE Transactions on Visualization and Computer Graphics, 18(8):1228–1240, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Karunanayaka:2018:NTT

[KJH^{+}18] Kasun Karunanayaka, Nurafiqah Johari, Surina Hariri, Hanis Camelia, Kevin Stanley Bielawski, and Adrian David Cheok. New thermal taste actuation technology for future multisensory virtual reality and Internet. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1496–1505, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260970-
abs.html.

Kronander:2012:EVE

[KJL^{+}12] Joel Kronander, Daniel Jonsson, Joakim Low, Patric Ljung, Anders Ynnerman, and Jonas Unger. Efficient visibility encoding for dynamic illumination in direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 18 (3):447–462, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koch:2014:VDV

[KJW^{+}14] Steffen Koch, Markus John, Michael Worner, Andreas Muller, and Thomas Ertl. VarifocalReader — in-depth visual analysis of large text documents. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1723–1732, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

csdl/trans/tg/2014/12/06875959-
abs.html.

Kim:2018:DFA

[KJW^{+}18] Seokyeon Kim, Seongmin Jeong, Insoo Woo, Yun Jang, Ross Maciejewski, and David S. Ebert. Data flow analysis and visualization for spatiotemporal statistical data without trajectory information. IEEE Transactions on Visualization and Computer Graphics, 24(3): 1287–1300, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krekhov:2019:DNP

[KK19] Andrey Krekhov and Jens Kruger. Deadeye: A novel preattentive visualization technique based on dichoptic presentation. IEEE Transactions on Visualization and Computer Graphics, 25 (1):936–945, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440097-
abs.html.

Kerracher:2015:TTT

[KKC15] Natalie Kerracher, Jessie Kennedy, and Kevin Chalmers. A task taxonomy for temporal graph visualisation. IEEE Transactions on Visualization and Computer Graphics, 21(10):1160–1172, October 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07091028-
abs.html.

Kwon:1998:MER

[KKCS98] Bomjun Kwon, Dae Seoung Kim, Kyung-Yong Chwa, and Sung Yong Shin. Memoryefficient ray classification for visibility operations. IEEE Transactions on Visualization and Computer Graphics, 4(3):193–201, July/ September 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0193.
pdf;   http://www.computer.
org/tvcg/tg1998/v0193abs.
htm.

Kniss:2002:MTF

[KKH02] Joe Kniss, Gordon Kindlmann, and Charles Hansen. Multidimensional transfer functions for interactive volume rendering. IEEE Transactions on Visualization and Computer Graphics, 8(3):270– 285, July/September 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/03/v0270abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/03/
v0270.htm;   http://csdl.

computer.org/dl/trans/tg/
2002/03/v0270.pdf.

Kato:2018:FRE

[KKKT18] Ginga Kato, Yoshihiro Kuroda, Kiyoshi Kiyokawa, and Haruo Takemura. Force rendering and its evaluation of a friction-based walking sensation display for a seated user. IEEE Transactions on Visualization and Computer Graphics, 24(4):1506– 1514, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08267106-
abs.html.

Kruger:2005:PSI

[KKKW05] Jens Kruger, Peter Kipfer, Polina Kondratieva, and Rudiger Westermann. A particle system for interactive visualization of 3D flows. IEEE Transactions on Visualization and Computer Graphics, 11(6):744–756, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kiyokawa:2011:GEI

[KKL11] Kiyoshi Kiyokawa, Gudrun Klinker, and Benjamin Lok. Guest Editors’ introduction: Special section on the IEEE Virtual Reality Conference (VR). IEEE Transactions on Visualization and Computer Graphics, 17(9):1193– 1194, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwon:2016:VDV

[KKL^{+}16] B. C. Kwon, S. Kim, S. Lee, J. Choo, J. Huh, and J. S. Yi. VisOHC: Designing visual analytics for online health communities. IEEE Transactions on Visualization and Computer Graphics, 22(1):71– 80, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krishnamurthy:2009:PEN

[KKM^{+}09] Adarsh Krishnamurthy, Rahul Khardekar, Sara McMains, Kirk Haller, and Gershon Elber. Performing efficient NURBS modeling operations on the GPU. IEEE Transactions on Visualization and Computer Graphics, 15(4): 530–543, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Khlebnikov:2011:CRT

[KKMS11] Rostislav Khlebnikov, Bernhard Kainz, Judith Muehl, and Dieter Schmalstieg. Crepuscular rays for tumor accessibility planning. IEEE Transactions on Visualization and Computer Graphics, 17(12):2163–2172, December 2011. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2017:TEM

[KKP^{+}17] Minjeong Kim, Kyeongpil Kang, Deokgun Park, Jaegul Choo, and Niklas Elmqvist. TopicLens: Efficient multilevel visual topic exploration of large-scale document collections. IEEE Transactions on Visualization and Computer Graphics, 23(1):151–160, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krueger:2008:SEA

[KKPS08] Arno Krueger, Christoph Kubisch, Bernhard Preim, and Gero Strauss. Sinus endoscopy — application of advanced GPU volume rendering for virtual endoscopy. IEEE Transactions on Visualization and Computer Graphics, 14(6):1491–1498, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kol:2017:ESS

[KKSE17] Timothy R. Kol, Oliver Klehm, Hans-Peter Seidel, and Elmar Eisemann. Expressive single scattering for light shaft stylization. IEEE Transactions on Visualization and Computer Graphics, 23(7):1753–1766, July 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/07/07452672-
                                        abs.html.

Koniaris:2019:CAL

[KKSM19] Charalampos Koniaris, Maggie Kosek, David Sinclair, and Kenny Mitchell. Compressed animated light fields with real-time view-dependent reconstruction. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1666–1680, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8322310/.

Khlebnikov:2013:NBV

[KKSS13] Rostislav Khlebnikov, Bernhard Kainz, Markus Steinberger, and Dieter Schmalstieg. Noise-based volume rendering for the visualization of multivariate volumetric data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2926– 2935, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwon:2017:AIN

[KKW^{+}17] Bum Chul Kwon, Hannah Kim, Emily Wall, Jaegul Choo, Haesun Park, and Alex Endert. AxiSketcher: Interactive nonlinear axis mapping of visualizations through user drawings. IEEE Transactions on Visualization and Computer Graphics, 23(1): 221–230, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kenwright:1996:ITD

[KL96] David N. Kenwright and David A. Lane. Interactive time-dependent particle tracing using tetrahedral decomposition. IEEE Transactions on Visualization and Computer Graphics, 2(2): 120–129, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0120.
pdf;   http://www.computer.
org/tvcg/tg1996/v0120abs.
htm.

Kim:2003:TMS

[KL03] Junho Kim and Seungyong Lee. Transitive mesh space of a progressive mesh. IEEE Transactions on Visualization and Computer Graphics, 9(4):463–480, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0463abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0463.pdf.

Kim:2007:FAL

[KL07] Theodore Kim and Ming C. Lin. Fast animation of lightning using an adaptive mesh. IEEE Transactions on Visualization and Computer Graphics, 13(2):390– 402, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2014:EAS

[KL14a] Young J. Kim and Fuchang Liu. Exact and adaptive signed distance fields computation for rigid and deformable models on GPUs. IEEE Transactions on Visualization and Computer Graphics, 20(5):714–725, May 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kry:2014:GEI

[KL14b] Paul G. Kry and Jehee Lee. Guest Editor’s introduction: Special section on the ACM SIGGRAPH/ Eurographics Symposium on Computer Animation (SCA). IEEE Transactions on Visualization and Computer Graphics, 20(1):2–3, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kidwell:2008:VIP

[KLC08] Paul Kidwell, Guy Lebanon, and William Cleveland. Visualizing incomplete and partially ranked data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1356–1363, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kang:2009:FBI

[KLC09] Henry Kang, Seungyong Lee, and Charles K. Chui. Flow-based image abstraction. IEEE Transactions on Visualization and Computer Graphics, 15(1):62–76, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2017:ERD

[KLCK17] Jong-Hyun Kim, Jung Lee, Sungdeok Cha, and Chang-Hun Kim. Efficient representation of detailed foam waves by incorporating projective space. IEEE Transactions on Visualization and Computer Graphics, 23(9):2056– 2068, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07567502-
abs.html.

Kotranza:2009:MRH

[KLD^{+}09] Aaron Kotranza, Benjamin Lok, Adeline Deladisma, Carla M. Pugh, and D. Scott Lind. Mixed reality humans: Evaluating behavior, usability, and acceptability. IEEE Transactions on Visualization and Computer Graphics, 15 (3):369–382, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Klemm:2016:RHM

[KLG^{+}16] P. Klemm, K. Lawonn, S. Glaser, U. Niemann, K. Hegenscheid, H. Volzke, and B. Preim. 3D regression heat map analysis of population study data. IEEE Transactions on Visualization and Computer Graphics, 22 (1):81–90, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kooima:2009:PST

[KLJ^{+}09] Robert Kooima, Jason Leigh, Andrew Johnson, Doug Roberts, Mark SubbaRao, and Thomas A. DeFanti. Planetary-scale terrain composition. IEEE Transactions on Visualization and Computer Graphics, 15(5):719–733, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2009:GWB

[KLK^{+}09] Bohyoung Kim, Bongshin Lee, Susan Knoblach, Eric Hoffman, and Jinwook Seo. GeneShelf: a Web-based visual interface for large gene expression time-series data repositories. IEEE Transactions on Visualization and Computer Graphics, 15(6):905–912, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koh:2010:MPF

[KLKS10] Kyle Koh, Bongshin Lee, Bohyoung Kim, and Jinwook Seo. ManiWordle: Providing flexible control over Wordle. IEEE Transactions on Visualization and Computer Graphics, 16(6):1190–1197, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kotranza:2012:RTE

[KLL12] Aaron Kotranza, D. Scott Lind, and Benjamin Lok. Real-time evaluation and visualization of learner performance in a mixed-reality environment for clinical breast examination. IEEE Transactions on Visualization and Computer Graphics, 18(7): 1101–1114, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2013:SHC

[KLL^{+}13] Duksu Kim, Jinkyu Lee, Junghwan Lee, Insik Shin, John Kim, and Sung-Eui Yoon. Scheduling in heterogeneous computing environments for proximity queries. IEEE Transactions on Visualization and Computer Graphics, 19(9):1513–1525, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2007:ASR

[KLLR07] ByungMoon Kim, Yingjie Liu, Ignacio Llamas, and Jarek Rossignac. Advections with significantly reduced dissipation and diffusion. IEEE Transactions on Visualization and Computer Graphics, 13(1):135–144, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2004:IPD

[KLM04] Young J. Kim, Ming C. Lin, and Dinesh Manocha. Incremental penetration depth estimation between convex polytopes using dual-space expansion. IEEE Transactions on Visualization and Computer Graphics, 10(2): 152–163, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0152abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0152.htm;   http://csdl.

computer.org/dl/trans/tg/
2004/02/v0152.pdf.

Kehrer:2008:HGC

[KLM^{+}08] Johannes Kehrer, Florian Ladstädter, Philipp Muigg, Helmut Doleisch, Andrea Steiner, and Helwig Hauser. Hypothesis generation in climate research with interactive visual data exploration. IEEE Transactions on Visualization and Computer Graphics, 14(6):1579–1586, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Karpenko:2010:EVD

[KLMA10] Olga Karpenko, Wilmot Li, Niloy Mitra, and Maneesh Agrawala. Exploded view diagrams of mathematical surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1311– 1318, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kurth:2018:ACD

[KLS^{+}18] Philipp Kurth, Vanessa Lange, Christian Siegl, Marc Stamminger, and Frank Bauer. Auto-calibration for dynamic multi-projection mapping on arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 24(11):2886–2894, November 2018. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL https://ieeexplore.
                                        ieee.org/document/8466021/

.

Kang:2013:VVA

[KLYE13] Kai Kang, Dajian Liu, Ye Yuan, and Yanli E. Vis4Heritage: Visual analytics approach on grotto wall painting degradations. IEEE Transactions on Visualization and Computer Graphics, 19(12):1982–1991, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2012:BTV

[KLySK12] Doyub Kim, Seung Woo Lee, Oh young Song, and Hyeong-Seok Ko. Baroclinic turbulence with varying density and temperature. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1488–1495, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Knight:1996:VUF

[KM96] David Knight and Gordon Mallinson. Visualizing unstructured flow data using dual stream functions. IEEE Transactions on Visualization and Computer Graphics, 2 (4):355–363, December 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0355.
pdf;   http://www.computer.
org/tvcg/tg1996/v0355abs.
htm.

Klein:2010:SLC

[KM10] Georg Klein and David W.

Murray. Simulating lowcost cameras for augmented reality compositing. IEEE Transactions on Visualization and Computer Graphics, 16 (3):369–380, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krajcevski:2016:CCM

[KM16] Pavel Krajcevski and Dinesh Manocha. Compressed coverage masks for path rendering on mobile GPUs. IEEE Transactions on Visualization and Computer Graphics, 22 (10):2229–2238, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kehrer:2011:IVA

[KMDH11] J. Kehrer, P. Muigg, H. Doleisch, and H. Hauser. Interactive visual analysis of heterogeneous scientific data across an interface. IEEE Transactions on Visualization and Computer Graphics, 17(7):934–946, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kovacs:2010:RTC

[KMDZ10] Denis Kovacs, Jason Mitchell, Shanon Drone, and Denis Zorin. Real-time creased approximate subdivision surfaces with displacements. IEEE Transactions on Visualization and Computer Graphics, 16(5):742–751, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Konyha:2006:IVA

[KMG^{+}06] Zoltan Konyha, Krešimir Matković, Denis Gracanin, Mario Jelovic, and Helwig Hauser. Interactive visual analysis of families of function graphs. IEEE Transactions on Visualization and Computer Graphics, 12(6):1373– 1385, November/December 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krishnamurthy:2011:GAM

[KMH11] Adarsh Krishnamurthy, Sara McMains, and Kirk Haller. GPU-accelerated minimum distance and clearance queries. IEEE Transactions on Visualization and Computer Graphics, 17(6):729–742, June 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2010:RRA

[KMKY10] Tae-Joon Kim, Bochang Moon, Duksu Kim, and Sung-Eui Yoon. RACB-VHs: Random-accessible compressed bounding volume hierarchies. IEEE Transactions on Visualization and Computer Graphics, 16(2): 273–286, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kumar:1996:IDL

[KML96] Subodh Kumar, Dinesh Manocha, and Anselmo Lastra. Interactive display of large NURBS models. IEEE Transactions on Visualization and Computer Graphics, 2(4):323– 336, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0323.
pdf;   http://www.computer.
org/tvcg/tg1996/v0323abs.
htm.

Kwon:2016:SLR

[KMLM16] Oh-Hyun Kwon, Chris Muelder, Kyungwon Lee, and Kwan-Liu Ma. A study of layout, rendering, and interaction methods for immersive graph visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (7):1802–1815, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2013:BMM

[KMM^{+}13] SungYe Kim, Ross Maciejewski, Abish Malik, Yun Jang, David S. Ebert, and Tobias Isenberg. Bristle maps: A multivariate abstraction technique for geovisualization. IEEE Transactions on Visualization and Computer Graphics, 19(9):1438–1454, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keim:2004:GEI

[KMN04] Daniel A. Keim, Tamara Munzner, and Stephen C. North. Guest Editors’ introduction: Special section on InfoVis. IEEE Transactions on Visualization and Computer Graphics, 10(4):446– ??, July/August 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/04/v0446.
                                        pdf;   http://csdl.computer.
                                        org/dl/trans/tg/2004/04/
                                        v0446.htm.

Kalkofen:2009:CVA

[KMS09] Denis Kalkofen, Erick Mendez, and Dieter Schmalstieg. Comprehensible visualization for augmented reality. IEEE Transactions on Visualization and Computer Graphics, 15(2):193–204, March/April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kapadia:2014:AAD

[KMSB14] Mubbasir Kapadia, Nathan Marshak, Alexander Shoulson, and Norman I. Badler. ADAPT: The agent development and prototyping testbed. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1035–1047, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2014:HCA

[KMT14] Young J. Kim, Dinesh Manocha, and Min Tang. Hierarchical and controlled advancement for continuous collision detection of rigid and articulated models. IEEE Transactions on Visualization and Computer Graphics, 20 (5):755–766, May 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kaufman:1995:Ea

[KN95] Arie E. Kaufman and Gregory M. Nielson. Editorials. IEEE Transactions on Visualization and Computer Graphics, 1(1):1–??, March 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0001.
pdf.

Kale:2019:HOP

[KNKH19] Alex Kale, Francis Nguyen, Matthew Kay, and Jessica Hullman. Hypothetical outcome plots help untrained observers judge trends in ambiguous data. IEEE Transactions on Visualization and Computer Graphics, 25(1): 892–902, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440816-
abs.html.

Koh:2015:VDA

[KNO15] Woojong Koh, Rahul Narain, and James F. O’Brien. Viewdependent adaptive cloth simulation with buckling compensation. IEEE Transactions on Visualization and Computer Graphics, 21(10):1138– 1145, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07127098-
abs.html.

Keim:2004:CFA

[KNP04] Daniel A. Keim, Stephen C. North, and Christian Panse. CartoDraw: a fast algorithm for generating contiguous cartograms. IEEE Transactions on Visualization and Computer Graphics, 10(1):95–110, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0095abs.
htm;   http://csdl.computer.

org/dl/trans/tg/2004/01/
v0095.pdf.

Kasahara:2017:JHI

[KNR17] Shunichi Kasahara, Shohei Nagai, and Jun Rekimoto. JackIn Head: Immersive visual telepresence system with omnidirectional wearable camera. IEEE Transactions on Visualization and Computer Graphics, 23(3): 1222–1234, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07792698-
abs.html.

Karamouzas:2012:SEL

[KO12] Ioannis Karamouzas and Mark Overmars. Simulating and evaluating the local behavior of small pedestrian groups. IEEE Transactions on Visualization and Computer Graphics, 18(3): 394–406, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kovacs:2019:CAA

[KOBH19] B. Kovacs, P. O’Donovan, K. Bala, and A. Hertzmann. Context-aware asset search for graphic design. IEEE Transactions on Visualization and Computer Graphics, 25(7):2419–2429, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kersten-Oertel:2014:EDE

[KOCC14] Marta Kersten-Oertel, Sean Jy-Shyang Chen, and D. Louis Collins. An evaluation of depth enhancing perceptual cues for vascular volume visualization in neurosurgery. IEEE Transactions on Visualization and Computer Graphics, 20(3):391–403, March 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kuhn:2008:ENP

[KOF08] Giovane R. Kuhn, Manuel M. Oliveira, and Leandro A. F. Fernandes. An efficient naturalness-preserving imagerecoloring method for dichromats. IEEE Transactions on Visualization and Computer Graphics, 14(6):1747– 1754, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kersten-Oertel:2012:DTM

[KOJC12] Marta Kersten-Oertel, Pierre Jannin, and D. Louis Collins. DVV: a taxonomy for mixed reality visualization in image guided surgery. IEEE Transactions on Visualization and Computer Graphics, 18 (2):332–352, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Klemm:2014:IVA

[KOJL^{+}14] Paul Klemm, Steffen Oeltze-Jafra, Kai Lawonn, Katrin Hegenscheid, Henry Volzke, and Bernhard Preim. Interactive visual analysis of image-centric cohort study data. IEEE Transactions on Visualization and Computer Graphics, 20(12):1673– 1682, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876009-
abs.html.

Kruijff:2019:ILD

[KOK^{+}19] E. Kruijff, J. Orlosky, N. Kishishita, C. Trepkowski, and K. Kiyokawa. The influence of label design on search performance and noticeability in wide field of view augmented reality displays. IEEE Transactions on Visualization and Computer Graphics, 25(9):2821– 2837, September 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koop:2008:VAS

[Koo08] David Koop. VisComplete:

Automating suggestions for visualization pipelines. IEEE Transactions on Visualization and Computer Graphics, 14(6):1691–1698, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

King:2005:CSS

[KP05] Scott A. King and Richard E. Parent. Creating speechsynchronized animation. IEEE Transactions on Visualization and Computer Graphics, 11 (3):341–352, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krause:2014:IIF

[KPB14] Josua Krause, Adam Perer, and Enrico Bertini. IN-FUSE: Interactive feature selection for predictive modeling of high dimensional data. IEEE Transactions on Visualization and Computer Graphics, 20(12):1614– 1623, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876047-
                                        abs.html.

Kehrer:2013:MSB

[KPBG13] Johannes Kehrer, Harald Piringer, Wolfgang Berger, and M. Eduard Groller. A model for structure-based comparison of many categories in small-multiple displays. IEEE Transactions on Visualization and Computer Graphics, 19(12):2287– 2296, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2016:RHO

[KPBL16] Yeonjoon Kim, Hangil Park, Seungbae Bang, and Sung-Hee Lee. Retargeting humanobject interaction to virtual avatars. IEEE Transactions on Visualization and Computer Graphics, 22(11):2405– 2412, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Komura:2012:GEI

[KPGL12] Taku Komura, Qunsheng Peng, George Gaciu, and Rynson W. H. Lau. Guest Editors’ introduction: Special section on ACM VRST. IEEE Transactions on Visualization and Computer Graphics, 18 (3):354–355, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kniss:2003:MVL

[KPH^{+}03] Joe Kniss, Simon Premoze, Charles Hansen, Peter Shirley, and Allen McPherson. A model for volume lighting and modeling. IEEE Transactions on Visualization and Computer Graphics, 9 (2):150–162, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0150abs.

                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/02/
                                        v0150.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/02/v0150.pdf.

Kandel:2012:EDA

[KPHH12] Sean Kandel, Andreas Paepcke, Joseph M. Hellerstein, and Jeffrey Heer. Enterprise data analysis and visualization: an interview study. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2917–2926, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krivanek:2014:TEU

[KPR^{+}14] Jaroslav Krivanek, Fabio Pellacini, Martin Ruzicka, Vaclav Gassenbauer, and Ondrej Karlik. Toward evaluating the usefulness of global illumination for novices in lighting design tasks. IEEE Transactions on Visualization and Computer Graphics, 20 (6):944–954, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kahler:2015:VHF

[KPR^{+}15] Olaf Kahler, Victor Adrian Prisacariu, Carl Yuheng Ren, Xin Sun, Philip Torr, and David Murray. Very high frame rate volumetric integration of depth images on mobile devices. IEEE Transactions on Visualization and Computer Graphics, 21(11): 1241–1250, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165673-
abs.html.

Krause:2016:SIC

[KPS16] J. Krause, A. Perer, and H. Stavropoulos. Supporting iterative cohort construction with visual temporal queries. IEEE Transactions on Visualization and Computer Graphics, 22(1):91–100, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koytek:2018:MBL

[KPV^{+}18] P. Koytek, C. Perin, J. Vermeulen, E. André, and S. Carpendale. MyBrush: Brushing and linking with personal agency. IEEE Transactions on Visualization and Computer Graphics, 24(1): 605–615, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kramida:2016:RVA

[Kra16] Gregory Kramida. Resolving the vergence-accommodation conflict in head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 22(7):1912– 1931, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2018:DTO

[KRH18] Y. S. Kim, K. Reinecke, and J. Hullman. Data through others’s eyes: The impact of visualizing others’s expectations on visualization interpretation. IEEE Transactions on Visualization and Computer Graphics, 24(1): 760–769, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kasten:2011:TDT

[KRHH11] Jens Kasten, Jan Reininghaus, Ingrid Hotz, and Hans-Christian Hege. Twodimensional time-dependent vortex regions based on the acceleration magnitude. IEEE Transactions on Visualization and Computer Graphics, 17(12):2080–2087, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kumpf:2019:VAT

[KRRW19] Alexander Kumpf, Marc Rautenhaus, Michael Riemer, and Rudiger Westermann. Visual analysis of the temporal evolution of ensemble forecast sensitivities. IEEE Transactions on Visualization and Computer Graphics, 25(1): 98–108, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440839-
abs.html.

Keim:2006:GES

[KRTvW06] Daniel A. Keim, George G.

Robertson, Jim J. Thomas, and Jarke J. van Wijk. Guest editorial: Special section on visual analytics. IEEE Transactions on Visualization and Computer Graphics, 12(6):1361–1362, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/06/v1361.
pdf.

Kanzler:2019:VBR

[KRW19] M. Kanzler, M. Rautenhaus, and R. Westermann. A voxelbased rendering pipeline for large 3D line sets. IEEE Transactions on Visualization and Computer Graphics, 25(7):2378–2391, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Klosowski:2000:CPL

[KS00a] J. T. Klosowski and C. T.

Silva. Corrections to The Prioritized-Layered Projection Algorithm for Visible Set Estimation. IEEE Transactions on Visualization and Computer Graphics, 6(4):380–??, October/December 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0380.
pdf;   http://www.computer.
org/tvcg/tg2000/v0380abs.

htm. See [KS00b].

Klosowski:2000:PLP

[KS00b] J. T. Klosowski and C. T. Silva. The prioritized-layered projection algorithm for visible set estimation. IEEE Transactions on Visualization and Computer Graphics, 6 (2):108–123, April/June 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0108.
pdf;   http://www.computer.
org/tvcg/tg2000/v0108abs.

htm. See corrections [KS00a].

Klosowski:2001:ECV

[KS01] James T. Klosowski and Cláudio T. Silva. Efficient conservative visibility culling using the prioritized-layered projection algorithm. IEEE Transactions on Visualization and Computer Graphics, 7(4): 365–379, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0365.
pdf;   http://www.computer.
org/tvcg/tg2001/v0365abs.
htm.

Kreuseler:2002:FAV

[KS02] Matthias Kreuseler and Heidrun Schumann. A flexible approach for visual data mining. IEEE Transactions on Visualization and Computer Graphics, 8(1):39–51, January 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0039.
pdf;   http://www.computer.
org/tvcg/tg2002/v0039abs.
htm.

Kim:2014:GEI

[KS14a] Theodore Kim and Robert Summer. Guest Editors’ introduction: Special section on the ACM SIG-GRAPH/Eurographics Symposium on Computer Animation (SCA). IEEE Transactions on Visualization and Computer Graphics, 20(10): 1344, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06881790.
pdf.

Kindlmann:2014:APV

[KS14b] Gordon Kindlmann and Carlos Scheidegger. An algebraic process for visualization design. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2181–2190, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875930-
                                        abs.html.

Kruger:2018:VIM

[KSBE18] Robert Kruger, Georgi Simeonov, Fabian Beck, and Thomas Ertl. Visual interactive map matching. IEEE Transactions on Visualization and Computer Graphics, 24(6):1881–1892, June 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/06/08320847-
                                        abs.html.

Kampe:2016:FME

[KSDA16] Viktor Kämpe, Erik Sintorn, Dan Dolonius, and Ulf Assarsson. Fast, memoryefficient construction of voxelized shadows. IEEE Transactions on Visualization and Computer Graphics, 22(10): 2239–2248, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kothur:2014:VAC

[KSDD14] Patrick Kothur, Mike Sips, Henryk Dobslaw, and Doris Dransch. Visual analytics for comparison of ocean model output with reference data: Detecting and analyzing geophysical processes using clustering ensembles. IEEE Transactions on Visualization and Computer Graphics, 20(12):1893– 1902, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876007-
abs.html.

Kappe:2016:RVC

[KSG^{+}16] C. P. Kappe, L. Schutz, S. Gunther, L. Hufnagel, S. Lemke, and H. Leitte. Reconstruction and visualization of coordinated 3D cell migration based on optical flow. IEEE Transactions on Visualization and Computer Graphics, 22(1):995–1004, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kahler:2003:IVR

[KSH03] Ralf Kähler, Mark Simon, and Hans-Christian Hege. Interactive volume rendering of large sparse data sets using adaptive mesh refinement hierarchies. IEEE Transactions on Visualization and Computer Graphics, 9(3):341–351, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0341abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0341.pdf.

Kikinis:1996:DBA

[KSI^{+}96] Ron Kikinis, Martha E. Shenton, Dan V. Iosifescu, Robert W. McCarley, Pairash Saiviroonporn, Hiroto H. Hokama, Andre Robatino, David Metcalf, Cynthia G. Wible, Chiara M. Portas, Robert M. Donnino, and Ferenc A. Jolesz. A digital brain atlas for surgical planning, model-driven segmentation, and teaching. IEEE Transactions on Visualization and Computer Graphics, 2 (3):232–241, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0232.
pdf;   http://www.computer.
org/tvcg/tg1996/v0232abs.

htm. Illustrates the use of Java for display and visualization of medical images.

Kim:2017:DDG

[KSL^{+}17] Nam Wook Kim, Eston Schweickart, Zhicheng Liu, Mira Dontcheva, Wilmot Li, Jovan Popovic, and Hanspeter Pfister. Datadriven guides: Supporting expressive design for information graphics. IEEE Transactions on Visualization and Computer Graphics, 23(1): 491–500, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kawai:2017:ARM

[KSNY17] Norihiko Kawai, Tomokazu Sato, Yuta Nakashima, and Naokazu Yokoya. Augmented reality marker hiding with texture deformation. IEEE Transactions on Visualization and Computer Graphics, 24(10):2288– 2300, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07593269-
abs.html.

Kerwin:2009:ERW

[KSS09] Thomas Kerwin, Han-Wei Shen, and Don Stredney. Enhancing realism of wet surfaces in temporal bone surgical simulation. IEEE Transactions on Visualization and Computer Graphics, 15(5): 747–758, September/October 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kindlmann:2009:SVC

[KSSW09] Gordon L. Kindlmann, Raul San Jose Estepar, Stephen M. Smith, and Carl-Fredrik Westin. Sampling and visualizing creases with scale-space particles. IEEE Transactions on Visualization and Computer Graphics, 15(6):1415– 1424, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kretschmer:2014:AAD

[KST^{+}14] Jan Kretschmer, Grzegorz Soza, Christian Tietjen, Michael Suehling, Bernhard Preim, and Marc Stamminger. ADR —- Anatomy-Driven Reformation. IEEE Transactions on Visualization and Computer Graphics, 20(12):2496–2505, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876018-
                                        abs.html.

Kersten:2006:EDP

[KSTE06] Marta Kersten, James Stewart, Niko Troje, and Randy Ellis. Enhancing depth perception in translucent volumes. IEEE Transactions on Visualization and Computer Graphics, 12(5):1117– 1124, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kiyokawa:2018:P

[KSTW18] Kiyoshi Kiyokawa, Frank Steinicke, Bruce Thomas, and Greg Welch. Preface. IEEE Transactions on Visualization and Computer Graphics, 24 (4):vi–vii, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/

csdl/trans/tg/2018/04/08315160-
abs.html.

Kruger:2006:CIC

[KSW06] Jens Krüger, Jens Schneider, and Rüdiger Westermann. ClearView: an interactive context preserving hotspot visualization technique. IEEE Transactions on Visualization and Computer Graphics, 12(5):941–948, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2014:RIG

[KSY14] Tae-Joon Kim, Xin Sun, and Sung-Eui Yoon. T-ReX: Interactive global illumination of massive models on heterogeneous computing resources. IEEE Transactions on Visualization and Computer Graphics, 20(3):481– 494, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kawai:2016:DRB

[KSY16] N. Kawai, T. Sato, and N. Yokoya. Diminished reality based on image inpainting considering background geometry. IEEE Transactions on Visualization and Computer Graphics, 22(3):1236– 1247, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kumpf:2018:VCC

[KTB^{+}18] A. Kumpf, B. Tost, M. Baumgart, M. Riemer, R. Westermann, and M. Rautenhaus. Visualizing confidence in cluster-based ensemble weather forecast analyses. IEEE Transactions on Visualization and Computer Graphics, 24(1):109–119, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kahng:2019:GLU

[KTC^{+}19] Minsuk Kahng, Nikhil Thorat, Duen Horng Polo Chau, Fernanda B. Viegas, and Martin Wattenberg. GAN Lab: Understanding complex deep generative models using interactive visual experimentation. IEEE Transactions on Visualization and Computer Graphics, 25(1):310– 320, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440049-
abs.html.

Klacansky:2017:FEF

[KTCG17] Pavol Klacansky, Julien Tierny, Hamish Carr, and Zhao Geng. Fast and exact fiber surfaces for tetrahedral meshes. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1782–1795, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07471499-
abs.html.

Krueger:2015:SEM

[KTE15] Robert Krueger, Dennis Thom, and Thomas Ertl. Semantic enrichment of movement behavior with Foursquare

— a visual analytics approach. IEEE Transactions on Visualization and Computer Graphics, 21(8):903– 915, August 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/06960909-abs.html.

Knecht:2013:RRO

[KTWW13] Martin Knecht, Christoph Traxler, Christoph Winklhofer, and Michael Wimmer. Reflective and refractive objects for mixed reality. IEEE Transactions on Visualization and Computer Graphics, 19 (4):576–582, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kutulakos:1998:CFA

[KV98] K. N. Kutulakos and J. R.

Vallino. Calibration-free augmented reality. IEEE Transactions on Visualization and Computer Graphics, 4(1):1– 20, January/March 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0001.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0001abs.
                                        htm.

Kalaiah:2003:MRP

[KV03] Aravind Kalaiah and Amitabh Varshney. Modeling and rendering of points with local geometry. IEEE Transactions on Visualization and Computer Graphics, 9(1):30–42, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/01/v0030abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/01/
                                        v0030.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/01/v0030.pdf.

Kim:2006:SGE

[KV06] Youngmin Kim and Amitabh Varshney. Saliency-guided enhancement for volume visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5): 925–932, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2008:PVA

[KV08] Youngmin Kim and Amitabh Varshney. Persuading visual attention through geometry. IEEE Transactions on Visualization and Computer Graphics, 14(4):772– 782, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kobourov:2005:NES

[KW05] Stephen G. Kobourov and Kevin Wampler. Non-Euclidean spring embedders. IEEE Transactions on Visualization and Computer Graphics, 11(6):757–767, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kindlmann:2006:DTV

[KW06] Gordon Kindlmann and Carl-Fredrik Westin. Diffusion tensor visualization with glyph packing. IEEE Transactions on Visualization and Computer Graphics, 12(5):1329– 1336, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Khoury:2010:FDI

[KW10] Marc Khoury and Rephael Wenger. On the fractal dimension of isosurfaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1198–1205, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kniss:2011:SMD

[KW11] Joe Kniss and Guanyu Wang. Supervised manifold distance segmentation. IEEE Transactions on Visualization and Computer Graphics, 17(11): 1637–1649, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kurzhals:2013:STV

[KW13] Kuno Kurzhals and Daniel Weiskopf. Space–time visual analytics of eye-tracking data for dynamic stimuli. IEEE Transactions on Visualization and Computer Graphics, 19(12):2129–2138, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kaufman:2014:IDA

[KW14] Arie E. Kaufman and Lei Wang. Importance-driven accessory lights design for enhancing local shapes. IEEE Transactions on Visualization and Computer Graphics, 20 (5):781–794, May 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kopp:2019:TTS

[KW19] Wiebke Kopp and Tino Weinkauf. Temporal treemaps: Static visualization of evolving trees. IEEE Transactions on Visualization and Computer Graphics, 25(1):

534–543, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08443124-
abs.html.

Kwatra:2010:FSA

[KWC^{+}10] Nipun Kwatra, Chris Wojtan, Mark Carlson, Irfan Essa, Peter J. Mucha, and Greg Turk. Fluid simulation with articulated bodies. IEEE Transactions on Visualization and Computer Graphics, 16(1):70– 80, January/February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Koehler:2011:VVU

[KWDG11] Christopher Koehler, Thomas Wischgoll, Haibo Dong, and Zachary Gaston. Vortex visualization in ultra low Reynolds number insect flight. IEEE Transactions on Visualization and Computer Graphics, 17(12):2071– 2079, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kindlmann:2000:SDV

[KWH00] G. Kindlmann, D. Weinstein, and D. Hart. Strategies for direct volume rendering of diffusion tensor fields. IEEE Transactions on Visualization and Computer Graphics, 6 (2):124–138, April/June 2000.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0124.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0124abs.
                                        htm.

Kronenberger:2019:ECC

[KWH19] M. Kronenberger, O. Wirjadi, and H. Hagen. Empirical comparison of curvature estimators on volume images and triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 25(10):3032–3041, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2014:WSP

[KWL14] Yongjin Kim, Holger Winnemoller, and Seungyong Lee. WYSIWYG stereo painting with usability enhancements. IEEE Transactions on Visualization and Computer Graphics, 20(7):957–969, July 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2001:EST

[KWP01] Kwansik Kim, Craig M. Wittenbrink, and Alex Pang. Extended specifications and test data sets for data level comparisons of direct volume rendering algorithms. IEEE Transactions on Visualization and Computer Graphics, 7(4):

299–317, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0299.
pdf;   http://www.computer.
org/tvcg/tg2001/v0299abs.
htm.

Konev:2014:RWA

[KWS^{+}14] Artem Konev, Jurgen Waser, Bernhard Sadransky, Daniel Cornel, Rui A. P. Perdigao, Zsolt Horvath, and M. Eduard Groller. Run watchers: Automatic simulation-based decision support in flood management. IEEE Transactions on Visualization and Computer Graphics, 20(12):1873– 1882, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875941-
abs.html.

Kwan:2018:PVD

[KXW^{+}18] Kin Chung Kwan, Xuemiao Xu, Liang Wan, Tien-Tsin Wong, and Wai-Man Pang. Packing vertex data into hardware-decompressible textures. IEEE Transactions on Visualization and Computer Graphics, 24(5):1705– 1716, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2018/05/07903744-
                                        abs.html.

Kuo:2006:NCA

[KY06] Chuan-Chu Kuo and Hong-Tzong Yau. A new combinatorial approach to surface reconstruction with sharp features. IEEE Transactions on Visualization and Computer Graphics, 12(1):73–82, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2011:MEU

[KYK11] Hong-Yun Kim, Chang-Hyo Yu, and Lee-Sup Kim. A memory-efficient unified early Z-test. IEEE Transactions on Visualization and Computer Graphics, 17(9):1286– 1294, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krichenbauer:2018:ARV

[KYT^{+}18] M. Krichenbauer, G. Yamamoto, T. Taketom, C. Sandor, and H. Kato. Augmented reality versus virtual reality for 3D object manipulation. IEEE Transactions on Visualization and Computer Graphics, 24(2):1038–1048, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2010:IVH

[KZD^{+}10] Seon Joo Kim, Shaojie Zhuo, Fanbo Deng, Chi-Wing Fu, and Michael Brown. Interactive visualization of hyperspectral images of historical documents. IEEE Transactions on Visualization and Computer Graphics, 16(6):1441–1448, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Keefe:2007:DAI

[KZL07] Daniel Keefe, Robert Zeleznik, and David Laidlaw. Drawing on air: Input techniques for controlled 3D line illustration. IEEE Transactions on Visualization and Computer Graphics, 13(5):1067–1081, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwok:2012:EOC

[KZW12] Tsz-Ho Kwok, Yunbo Zhang, and Charlie C. L. Wang. Efficient optimization of common base domains for cross parameterization. IEEE Transactions on Visualization and Computer Graphics, 18(10): 1678–1692, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwok:2016:SET

[KZW^{+}16] T. H. Kwok, Y. Q. Zhang, C. C. L. Wang, Y. J. Liu, and K. Tang. Styling evolution for tight-fitting garments. IEEE Transactions on Visualization and Computer Graphics, 22 (5):1580–1591, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ko:2014:VIC

[KZX^{+}14] Sungahn Ko, Jieqiong Zhao, Jing Xia, Shehzad Afzal, Xiaoyu Wang, Greg Abram, Niklas Elmqvist, Len Kne, David Van Riper, Kelly Gaither, Shaun Kennedy, William Tolone, William Ribarsky, and David S. Ebert. VASA: Interactive computational steering of large asynchronous simulation pipelines for societal infrastructure. IEEE Transactions on Visualization and Computer Graphics, 20(12):1853–1862, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875926-
abs.html.

Lang:2011:MBM

[LA11] Jochen Lang and Sheldon Andrews. Measurement-based modeling of contact forces and textures for haptic rendering. IEEE Transactions on Visualization and Computer Graphics, 17(3):380– 391, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livingston:2010:GEI

[LABS10] Mark A. Livingston, Ronald T.

Azuma, Oliver Bimber, and Hideo Saito. Guest Editors’ introduction: Special section on The International Symposium on Mixed and Augmented Reality (ISMAR). IEEE Transactions on Visualization and Computer Graphics, 16(3):353–354, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lacroute:1996:APV

[Lac96] Philippe Lacroute. Analysis of a parallel volume rendering system based on the shearwarp factorization. IEEE Transactions on Visualization and Computer Graphics, 2 (3):218–231, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0218.
pdf;   http://www.computer.
org/tvcg/tg1996/v0218abs.
htm.

Lawrence:2011:UAA

[LAK^{+}11] Jason Lawrence, Sean Arietta, Michael Kazhdan, Daniel Lepage, and Colleen O’Hagan. A user-assisted approach to visualizing multidimensional images. IEEE Transactions on Visualization and Computer Graphics, 17(10):1487– 1498, October 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lam:2008:FIC

[Lam08] Heidi Lam. A framework of interaction costs in information visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1149–1156, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Law:2010:PPP

[LAM10] Alvin J. Law, Daniel G. Aliaga, and Aditi Majumder. Projector placement planning for high quality visualizations on real-world colored objects. IEEE Transactions on Visualization and Computer Graphics, 16(6):1633–1641, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lobo:2019:APB

[LAP19] Maria-Jesus Lobo, Caroline Appert, and Emmanuel Pietriga. Animation plans for before-and-after satellite images. IEEE Transactions on Visualization and Computer Graphics, 25(2):1347– 1360, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/

csdl/trans/tg/2019/02/08267097-
abs.html.

Lopes:2003:IRA

[LB03] Adriano Lopes and Ken Brodlie. Improving the robustness and accuracy of the marching cubes algorithm for isosurfacing. IEEE Transactions on Visualization and Computer Graphics, 9(1):16– 29, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0016abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0016.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/01/v0016.pdf.

Li:2015:SAM

[LB15] Yijing Li and Jernej Barbic. Stable anisotropic materials. IEEE Transactions on Visualization and Computer Graphics, 21(10):1129– 1137, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07130660-
abs.html.

Lind:2017:CCS

[LB17] Andreas J. Lind and Stefan Bruckner. Comparing crosssections and 3D renderings for surface matching tasks using physical ground truths.

IEEE Transactions on Visualization and Computer Graphics, 23(1):781–790, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Latif:2019:VAP

[LB19] Shahid Latif and Fabian Beck. VIS author profiles: Interactive descriptions of publication records combining text and visualization. IEEE Transactions on Visualization and Computer Graphics, 25 (1):152–161, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440852-
abs.html.

Laffont:2013:RII

[LBD13] P. Laffont, A. Bousseau, and G. Drettakis. Rich intrinsic image decomposition of outdoor scenes from multiple views. IEEE Transactions on Visualization and Computer Graphics, 19(2):210– 224, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2008:GOS

[LBG^{+}08] Xin Li, Yunfan Bao, Xiaohu Guo, Miao Jin, Xianfeng Gu, and Hong Qin. Globally optimal surface mapping for surfaces with arbitrary topology. IEEE Transactions on Visualization and Computer Graphics, 14(4): 805–819, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040805s.
zip.

Labschutz:2016:JJT

[LBG^{+}16] M. Labschutz, S. Bruckner, M. E. Groller, M. Hadwiger, and P. Rautek. JiTTree: A just-in-time compiled sparse GPU volume data structure. IEEE Transactions on Visualization and Computer Graphics, 22(1):1025–1034, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lafon:2013:HRD

[LBGV13] Sebastien Lafon, Fatma Bouali, Christiane Guinot, and Gilles Venturini. Hierarchical reorganization of dimensions in OLAP visualizations. IEEE Transactions on Visualization and Computer Graphics, 19(11):1833–1845, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindow:2011:VBE

[LBH11] Norbert Lindow, Daniel Baum, and Hans-Christian Hege. Voronoi-based extraction and visualization of molecular paths. IEEE Transactions on Visualization and Computer Graphics, 17(12):

2025–2034, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindow:2014:LES

[LBH14] Norbert Lindow, Daniel Baum, and Hans-Christian Hege. Ligand excluded surface: A new type of molecular surface. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2486–2495, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876051-
                                        abs.html.

Li:2018:SSD

[LBH18] Chenhui Li, George Baciu, and Yu Han. StreamMap: Smooth dynamic visualization of high-density streaming points. IEEE Transactions on Visualization and Computer Graphics, 24(3):1381– 1393, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2018:EUP

[LBHW18] Myungho Lee, Gerd Bruder, Tobias Hollerer, and Greg Welch. Effects of unaugmented periphery and vibrotactile feedback on proxemics with virtual humans in AR. IEEE Transactions on Visualization and Computer Graphics, 24(4):1525– 1534, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08302409-
abs.html.

Lam:2012:ESI

[LBI^{+}12] Heidi Lam, Enrico Bertini, Petra Isenberg, Catherine Plaisant, and Sheelagh Carpendale. Empirical studies in information visualization: Seven scenarios. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1520–1536, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lekschas:2018:HVE

[LBK^{+}18] F. Lekschas, B. Bach, P. Kerpedjiev, N. Gehlenborg, and H. Pfister. HiPiler: Visual exploration of large genome interaction matrices with interactive small multiples. IEEE Transactions on Visualization and Computer Graphics, 24 (1):522–531, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livingston:2009:GEI

[LBKD09] Mark A. Livingston, Reinhold Behringer, Hirokazu Kato, and Tom Drummond. Guest Editors’ introduction: Special section on The International Symposium on Mixed and Augmented Reality (IS-MAR). IEEE Transactions on Visualization and Computer Graphics, 15(2):177– 178, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2019:CSK

[LBL19] W. Li, K. Bai, and X. Liu. Continuous-scale kinetic fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 25(9): 2694–2709, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lopez:2014:EAT

[LBLE^{+}14] Thomas Lopez, Rozenn Bouville, Emilie Loup-Escande, Florian Nouviale, Valerie Gouranton, and Bruno Arnaldi. Exchange of avatars: Toward a better perception and understanding. IEEE Transactions on Visualization and Computer Graphics, 20 (4):644–653, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindow:2019:IVR

[LBLH19] Norbert Lindow, Daniel Baum, Morgan Leborgne, and Hans-Christian Hege. Interactive visualization of RNA and DNA structures. IEEE Transactions on Visualization and Computer Graphics, 25 (1):967–976, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08494830-
abs.html.

Laney:2006:UST

[LBM^{+}06] D. Laney, P.-T. Bremer, A. Mascarenhas, P. Miller, and V. Pascucci. Understanding the structure of the turbulent mixing layer in hydrodynamic instabilities. IEEE Transactions on Visualization and Computer Graphics, 12(5):1053–1060, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2017:UVR

[LBR^{+}17] Le Liu, Alexander P. Boone, Ian T. Ruginski, Lace Padilla, Mary Hegarty, Sarah H. Creem-Regehr, William B. Thompson, Cem Yuksel, and Donald H. House. Uncertainty visualization by representative sampling from prediction ensembles. IEEE Transactions on Visualization and Computer Graphics, 23(9): 2165–2178, September 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/09/07563342-
abs.html.

Laha:2013:VMS

[LBS13] Bireswar Laha, Doug A. Bowman, and James D. Schiffbauer. Validation of the MR simulation approach for evaluating the effects of immersion on visual analysis of volume data. IEEE Transactions on Visualization and Computer Graphics, 19(4): 529–538, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Laha:2014:EVS

[LBS14] Bireswar Laha, Doug A. Bowman, and John J. Socha. Effects of VR system fidelity on analyzing isosurface visualization of volume datasets. IEEE Transactions on Visualization and Computer Graphics, 20 (4):513–522, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lincoln:2016:MPM

[LBS^{+}16] P. Lincoln, A. Blate, M. Singh, T. Whitted, A. State, A. Lastra, and H. Fuchs. From motion to photons in 80 microseconds: Towards minimal latency for virtual and augmented reality. IEEE Transactions on Visualization and Computer Graphics, 22(4): 1367–1376, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Luciani:2019:DFS

[LBS^{+}19] Timothy Luciani, Andrew Burks, Cassiano Sugiyama, Jonathan Komperda, and G. Elisabeta Marai. Detailsfirst, show context, overview last: Supporting exploration of viscous fingers in large-scale ensemble simulations. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1225–1235, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440850-
abs.html.

Liu:2018:VES

[LBT^{+}18] S. Liu, P. T. Bremer, J. J.

Thiagarajan, V. Srikumar, B. Wang, Y. Livnat, and V. Pascucci. Visual exploration of semantic relationships in neural word embeddings. IEEE Transactions on Visualization and Computer Graphics, 24(1):553–562, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Law:2019:DHD

[LBW19] Po-Ming Law, Rahul C. Basole, and Yanhong Wu. Duet: Helping data analysis novices conduct pairwise comparisons by minimal specification. IEEE Transactions on Visualization and Computer Graphics, 25(1):427– 437, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440115-
                                        abs.html.

Li:2011:GSS

[LBZ^{+}11] Yuanyuan Li, Fan Bao, Eugene Zhang, Yoshihiro Kobayashi, and Peter Wonka. Geometry synthesis on surfaces using field-guided shape grammars. IEEE Transactions on Visualization and Computer Graphics, 17(2): 231–243, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lou:2010:EBH

[LC10] Hui Lou and Jinxiang Chai. Example-based human motion denoising. IEEE Transactions on Visualization and Computer Graphics, 16(5): 870–879, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lai:2017:DDN

[LCC^{+}17] Yu-Chi Lai, Bo-An Chen, Kuo-Wei Chen, Wei-Lin Si, Chih-Yuan Yao, and Eugene Zhang. Data-driven NPR illustrations of natural flows in Chinese painting. IEEE Transactions on Visualization and Computer Graphics, 23(12):2535– 2549, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07726076-
abs.html.

Liu:2013:ORA

[LCDP13] Baoquan Liu, Gordon J. Clapworthy, Feng Dong, and Edmond C. Prakash. Octree rasterization: Accelerating highquality out-of-core GPU volume rendering. IEEE Transactions on Visualization and Computer Graphics, 19(10): 1732–1745, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2019:IMI

[LCL^{+}19] Shixia Liu, Changjian Chen, Yafeng Lu, Fangxin Ouyang, and Bin Wang. An interactive method to improve crowdsourced annotations. IEEE Transactions on Visualization and Computer Graphics, 25 (1):235–245, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440116-
abs.html.

Lauterbach:2007:ISR

[LCM07] Christian Lauterbach, Anish Chandak, and Dinesh Manocha. Interactive sound rendering in complex and dynamic scenes using frustum tracing. IEEE Transactions on Visualization and Computer Graphics, 13(6):1672– 1679, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2007/
                                        06/v1672s.mp4.

Lampe:2009:CCV

[LCMH09] Ove Daae Lampe, Carlos Correa, Kwan-Liu Ma, and Helwig Hauser. Curve-centric volume reformation for comparative visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1235–1242, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:HLR

[LCNG14] Ming C. Lin, Sean Curtis, Rahul Narain, and Abhinav Golas. Hybrid long-range collision avoidance for crowd simulation. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1022–1034, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Legg:2013:TUV

[LCP^{+}13] Philip A. Legg, David H. S. Chung, Matthew L. Parry, Rhodri Bown, Mark W. Jones, Iwan W. Griffiths, and Min Chen. Transformation of an uncertain video search pipeline to a sketch-based visual analytics loop. IEEE Transactions on Visualization and Computer Graphics, 19(12):2109–2118, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Langlotz:2016:RTR

[LCR16] Tobias Langlotz, Matthew Cook, and Holger Regenbrecht. Real-time radiometric compensation for optical seethrough head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 22(11):2385– 2394, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2006:SDV

[LCS06] Hongyu Li, Wenbin Chen, and I-Fan Shen. Segmentation of discrete vector fields. IEEE Transactions on Visualization and Computer Graphics, 12 (3):289–300, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2012:FVU

[LCS^{+}12] Zhanping Liu, Shangshu Cai, J. Edward Swan, Robert J. Moorhead, Joel P. Martin, and T. J. Jankun-Kelly. A 2D flow visualization user study using explicit flow synthesis and implicit task design.

IEEE Transactions on Visualization and Computer Graphics, 18(5):783–796, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2019:CIE

[LCZ^{+}19] J. Li, S. Chen, K. Zhang, G. Andrienko, and N. Andrienko. COPE: Interactive exploration of co-occurrence patterns in spatial time series. IEEE Transactions on Visualization and Computer Graphics, 25(8):2554–2567, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livingston:2011:ETL

[LD11a] Mark Livingston and Jonathan Decker. Evaluation of trend localization with multi-variate visualizations. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2053–2062, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lloyd:2011:HCA

[LD11b] David Lloyd and Jason Dykes. Human-centered approaches in geovisualization design: Investigating multiple methods through a long-term case study. IEEE Transactions on Visualization and Computer Graphics, 17(12):2498– 2507, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livingston:2012:EMV

[LDA12] Mark A. Livingston, Jonathan W.

Decker, and Zhuming Ai. Evaluation of multivariate visualization on a multivariate task. IEEE Transactions on Visualization and Computer Graphics, 18(12):2114– 2121, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:1996:RVP

[LDC96] P. Peggy Li, William H.

Duquette, and David W. Curkendall. RIVA: a versatile parallel rendering system for interactive scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 2(3):186– 201, September 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0186.
pdf;   http://www.computer.
org/tvcg/tg1996/v0186abs.
htm.

Lu:2016:RSF

[LDC16] X. Lu, Z. Deng, and W. Chen.

A robust scheme for featurepreserving mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 22(3):1181–1194, ???? 2016. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2014:ASC

[LDFZ14] Weiquan Lu, Henry Been-Lirn Duh, Steven Feiner, and Qi Zhao. Attributes of subtle cues for facilitating visual search in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 20(3):404– 412, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2009:DFI

[LDH09] Cha Lee, Stephen DiVerdi, and Tobias Höllerer. Depthfused 3D imagery on an immaterial display. IEEE Transactions on Visualization and Computer Graphics, 15(1):20– 33, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2018:URB

[LDM^{+}18] J. Liu, T. Dwyer, K. Marriott, J. Millar, and A. Haworth. Understanding the relationship between interactive optimisation and visual analytics in the context of prostate brachytherapy. IEEE Transactions on Visualization and Computer Graphics, 24 (1):319–329, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liao:2011:ADL

[LDN11] Hongen Liao, Takeyoshi Dohi, and Keisuke Nomura. Autostereoscopic 3D display with long visualization depth using referential viewing areabased integral photography. IEEE Transactions on Visualization and Computer Graphics, 17(11):1690–1701, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Loscos:2000:IVR

[LDR00] C. Loscos, G. Drettakis, and L. Robert. Interactive virtual relighting of real scenes. IEEE Transactions on Visualization and Computer Graphics, 6(4):289–??, October/December 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0289.
pdf;   http://www.computer.
org/tvcg/tg2000/v0289abs.
htm.

Lecocq:2017:AAR

[LDSM17] Pascal Lecocq, Arthur Dufay, Gael Sourimant, and Jean-Eudes Marvie. Analytic approximations for real-time area light shading. IEEE Transactions on Visualization and Computer Graphics, 23(5):1428–1441, May 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/05/07829413-
                                        abs.html.

Liu:2015:RBI

[LDW^{+}15] Yonghuai Liu, Luigi De Dominicis, Baogang Wei, Liang Chen, and Ralph R. Martin. Regularization based iterative point match weighting for accurate rigid transformation estimation. IEEE Transactions on Visualization and Computer Graphics, 21(9):1058–1071, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        09/07055263-abs.html.

Liu:2010:PCB

[LDX10] Yebin Liu, Qionghai Dai, and Wenli Xu. A pointcloud-based multiview stereo algorithm for free-viewpoint video. IEEE Transactions on Visualization and Computer Graphics, 16(3):407– 418, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lalonde:1997:WRR

[LF97] P. Lalonde and A. Fournier. A wavelet representation of reflectance functions. IEEE Transactions on Visualization and Computer Graphics, 3(4):329–336, October/ December 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0329.
pdf;   http://www.computer.
org/tvcg/tg1997/v0329abs.
htm.

Lowe:2016:VAD

[LFA^{+}16] T. Lowe, E. Forster, G. Albuquerque, J. Kreiss, and M. Magnor. Visual analytics for development and evaluation of order selection criteria for autoregressive processes. IEEE Transactions on Visualization and Computer Graphics, 22(1):151–159, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2006:SWE

[LFH06] Yinggang Li, Chi-Wing Fu, and Andrew Hanson. Scalable WIM: Effective exploration in large-scale astrophysical environments. IEEE Transactions on Visualization and Computer Graphics, 12(5):1005–1012, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2008:VMA

[LFH08] Hongwei Li, Chi-Wing Fu, and Andrew Hanson. Visualizing multiwavelength astrophysical data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1555–1562, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2007:VLS

[LFLH07] Hongwei Li, Chi-Wing Fu, Yinggang Li, and Andrew Hanson. Visualizing largescale uncertainty in astrophysical data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1640–1647, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1640s.
avi.

Li:2007:DDG

[LFP07] Ying Li, Jiaxin L. Fu, and Nancy S. Pollard. Datadriven grasp synthesis using shape matching and taskbased pruning. IEEE Transactions on Visualization and Computer Graphics, 13(4): 732–747, July/August 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lodha:2003:TPT

[LFR03] Suresh K. Lodha, Nikolai M. Faaland, and Jose C. Renteria. Topology preserving top-down compression of 2D vector fields using bintree and triangular quadtrees. IEEE Transactions on Visualization and Computer Graphics, 9(4):433– 442, October/December 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0433abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0433.pdf.

Lu:2019:ISP

[LFW^{+}19] Zhicong Lu, Mingming Fan, Yun Wang, Jian Zhao, Michelle Annett, and Daniel Wigdor. InkPlanner: Supporting prewriting via intelligent visual diagramming. IEEE Transactions on Visualization and Computer Graphics, 25(1):277– 287, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440848-
abs.html.

Liu:2012:OTO

[LG12] Yanli Liu and Xavier Granier.

Online tracking of outdoor lighting variations for augmented reality with moving cameras. IEEE Transactions on Visualization and Computer Graphics, 18(4): 573–580, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Levi:2013:DSI

[LG13] Z. Levi and C. Gotsman. D-Snake: Image registration by as-similar-as-possible template deformation. IEEE Transactions on Visualization and Computer Graphics, 19 (2):331–343, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Levi:2015:SRE

[LG15] Zohar Levi and Craig Gotsman. Smooth rotation enhanced as-rigid-as-possible mesh animation. IEEE Transactions on Visualization and Computer Graphics, 21(2): 264–277, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/02/06905844-
                                        abs.html.

Lin:2018:RIR

[LGG^{+}18] Hanfei Lin, Siyuan Gao, David Gotz, Fan Du, Jingrui He, and Nan Cao. RCLens: Interactive rare category exploration and identification. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2223–2237, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07939996-
                                        abs.html.

Luffel:2014:GCS

[LGLR14] Mark Luffel, Topraj Gurung, Peter Lindstrom, and Jarek Rossignac. Grouper: A compact, streamable triangle mesh data structure. IEEE Transactions on Visualization and Computer Graphics, 20 (1):84–98, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ledergerber:2008:VMR

[LGM^{+}08] Christian Ledergerber, Gaël Guennebaud, Miriah Meyer, Moritz Bächer, and Hanspeter Pfister. Volume MLS ray casting. IEEE Transactions on Visualization and Computer Graphics, 14(6):1372– 1379, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Leite:2018:EVA

[LGM^{+}18] R. A. Leite, T. Gschwandtner, S. Miksch, S. Kriglstein, M. Pohl, E. Gstrein, and J. Kuntner. EVA: Visual analytics to identify fraudulent events. IEEE Transactions on Visualization and Computer Graphics, 24(1):330–339, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2009:SMU

[LGQ09] Xin Li, Xianfeng Gu, and Hong Qin. Surface mapping using consistent pants decomposition. IEEE Transactions on Visualization and Computer Graphics, 15(4):558– 571, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lentine:2011:CCF

[LGS^{+}11] Michael Lentine, Jon Tomas Gretarsson, Craig Schroeder, Avi Robinson-Mosher, and Ronald Fedkiw. Creature control in a fluid environment. IEEE Transactions on Visualization and Computer Graphics, 17(5):682–693, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livesu:2012:RCS

[LGS12] Marco Livesu, Fabio Guggeri, and Riccardo Scateni. Reconstructing the curve-skeletons of 3D shapes using the visual hull. IEEE Transactions on Visualization and Computer Graphics, 18(11):1891– 1901, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lex:2014:UVI

[LGS^{+}14] Alexander Lex, Nils Gehlenborg, Hendrik Strobelt, Romain Vuillemot, and Hanspeter Pfister. UpSet: Visualization of intersecting sets. IEEE Transactions on Visualization and Computer Graphics, 20(12):1983–1992, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876017-
abs.html.

Lawonn:2016:OFB

[LGV^{+}16] K. Lawonn, S. Glaber, A. Vilanova, B. Preim, and T. Isenberg. Occlusion-free blood flow animation with wall thickness visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (1):728–737, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2019:FTM

[LGY19] Wei Li, Huajun Gong, and Ruigang Yang. Fast texture mapping adjustment via local/global optimization. IEEE Transactions on Visualization and Computer Graphics, 25(6):2296– 2303, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8352725/.

Liao:2012:VSC

[LGYG12] Miao Liao, Jizhou Gao, Ruigang Yang, and Minglun Gong. Video stereolization: Combining motion analysis with user interaction. IEEE Transactions on Visualization and Computer Graphics, 18(7):1079–1088, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2003:FIM

[LH03] Tong-Yee Lee and Po-Hua Huang. Fast and intuitive metamorphosis of 3D polyhedral models using SMCC mesh merging scheme. IEEE Transactions on Visualization and Computer Graphics, 9(1): 85–98, January/March 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/01/v0085abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/01/
                                        v0085.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/01/v0085.pdf.

Lee:2009:MHF

[LH09] Taehee Lee and Tobias Höllerer. Multithreaded hybrid feature tracking for markerless augmented reality. IEEE Transactions on Visualization and Computer Graphics, 15(3):355–368, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2011:EVC

[LH11] Sangyoon Lee and Hong Hua. Effects of viewing conditions and rotation methods in a collaborative tabletop AR environment. IEEE Transactions on Visualization and Computer Graphics, 17(9):1245– 1258, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2013:PC

[LH13] Xian-Ying Li and Shi-Min Hu.

Poisson coordinates. IEEE Transactions on Visualization and Computer Graphics, 19 (2):344–352, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2014:EIL

[LH14] Zhicheng Liu and Jeffrey Heer. The effects of interactive latency on exploratory visual analysis. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2122–2131, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876022-
abs.html.

Lee:2016:ECO

[LH16] S. Lee and H. Hua. Effects of configuration of optical combiner on near-field depth perception in optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 22 (4):1432–1441, ???? 2016.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2019:STM

[LHBF19] M. Lee, D. Hyde, M. Bao, and R. Fedkiw. A skinned tetrahedral mesh for hair animation and hair-water interaction. IEEE Transactions on Visualization and Computer Graphics, 25(3):1449– 1459, March 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2010:NPO

[LHC10] Sheng Liu, Hong Hua, and Dewen Cheng. A novel prototype for an optical seethrough head-mounted display with addressable focus cues. IEEE Transactions on Visualization and Computer Graphics, 16(3):381– 393, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2016:MES

[LHC16] Jaebong Lee, Bohyung Han, and Seungmoon Choi. Motion effects synthesis for 4D films. IEEE Transactions on Visualization and Computer Graphics, 22(10):2300–2314, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Langner:2018:VCC

[LHD18] R. Langner, T. Horak, and R. Dachselt. VisTiles: Coordinating and combining colocated mobile devices for visual data exploration. IEEE Transactions on Visualization and Computer Graphics, 24 (1):626–636, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liao:2012:SBR

[LHFY12] Zicheng Liao, Hugues Hoppe, David Forsyth, and Yizhou Yu. A subdivision-based representation for vector image editing. IEEE Transactions on Visualization and Computer Graphics, 18(11):1858– 1867, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lasserre:2012:NMM

[LHH^{+}12] Sebastien Lasserre, Juan Hernando, Sean Hill, Felix Schurmann, Pedro de Miguel Anasagasti, Georges Abou Jaoude, and Henry Markram. A neuron membrane mesh representation for visualization of electrophysiological simulations. IEEE Transactions on Visualization and Computer Graphics, 18(2):214–227, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2016:EOC

[LHH16] S. Lee, X. Hu, and H. Hua. Effects of optical combiner and IPD change for convergence on near-field depth perception in an optical see-through HMD. IEEE Transactions on Visualization and Computer Graphics, 22(5):1540– 1554, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lai:2018:SDG

[LHJ^{+}18] Wei-Sheng Lai, Yujia Huang, Neel Joshi, Christopher Buehler, Ming-Hsuan Yang, and Sing Bing Kang. Semantic-driven generation of hyperlapse from 360 degree video. IEEE Transactions on Visualization and Computer Graphics, 24(9): 2610–2621, September 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08031049-
abs.html.

Lam:2010:AFL

[LHLW10] Ping-Man Lam, Tze-Yiu Ho, Chi-Sing Leung, and Tien-Tsin Wong. All-frequency lighting with multiscale spherical radial basis functions. IEEE Transactions on Visualization and Computer Graphics, 16(1):43–56, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2006:GDL

[LHV06] Chang Ha Lee, Xuejun Hao, and Amitabh Varshney. Geometry-dependent lighting. IEEE Transactions on Visualization and Computer Graphics, 12(2):197– 207, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2004:SRB

[LHZ^{+}04] Xinguo Liu, Yaohua Hu, Jingdan Zhang, Xin Tong, Baining Guo, and Heung-Yeung Shum. Synthesis and rendering of bidirectional texture functions on arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 10(3):278– 289, May/June 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0278abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0278.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0278.pdf.

Lindstrom:2006:FEC

[LI06] Peter Lindstrom and Martin Isenburg. Fast and efficient compression of floating-point data. IEEE Transactions on Visualization and Computer Graphics, 12(5):1245– 1250, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Losasso:2006:MBS

[LIGF06] Frank Losasso, Geoffrey Irving, Eran Guendelman, and Ron Fedkiw. Melting and burning solids into liquids and gases. IEEE Transactions on Visualization and Computer Graphics, 12(3):343– 352, May/June 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:SIS

[LIM^{+}12] Juncong Lin, Takeo Igarashi, Jun Mitani, Minghong Liao, and Ying He. A sketching interface for sitting pose design in the virtual environment. IEEE Transactions on Visualization and Computer Graphics, 18(11):1979– 1991, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2011:ENa

[Lin11a] Ming C. Lin. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 17(5):553–554, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2011:ENb

[Lin11b] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 17(8):1033, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2011:MNE

[Lin11c] Ming C. Lin. A message from the new Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 17(3):263, March 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2011:MEC

[Lin11d] Ming Lin Lin. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 17 (12):ix, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:MECb

[Lin12a] Ming Lin. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 18(12): ix, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:ENa

[Lin12b] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 18(1):1–2, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:ENb

[Lin12c] Ming C. Lin. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 18(1):2, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:ENc

[Lin12d] Ming C. Lin. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 18(3):353, March 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2012:MECa

[Lin12e] Ming C. Lin. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 18 (4):v, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:ENa

[Lin13a] Ming C. Lin. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 19(1):1, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:ENb

[Lin13b] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 19(1):2, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:ENc

[Lin13c] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 19(6):897, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:ENd

[Lin13d] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 19(10):1605, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:MEC

[Lin13e] Ming C. Lin. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 19 (4):v, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:ENa

[Lin14a] Ming C. Lin. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 20(5):662–663, May 2014.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:ENc

[Lin14b] Ming C. Lin. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 20(8):1084, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:ENB

[Lin14c] Ming C. Lin. Editor’s note [2013 Best Associate Editor Award & 2013 Best Reviewer Award]. IEEE Transactions on Visualization and Computer Graphics, 20 (6):822, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:MEC

[Lin14d] Ming C. Lin. Message from the editor-in chief. IEEE Transactions on Visualization and Computer Graphics, 20 (4):v, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:SJ

[Lin14e] Ming C. Lin. State of the journal. IEEE Transactions on Visualization and Computer Graphics, 20(1):1, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindstrom:2014:FRC

[Lin14f] Peter Lindstrom. Fixed-rate compressed floating-point arrays. IEEE Transactions on Visualization and Computer Graphics, 20(12):2674– 2683, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876024-
abs.html.

Lin:2015:EFN

[Lin15] Ming C. Lin. EIC farewell and new EIC introduction. IEEE Transactions on Visualization and Computer Graphics, 21(1):1–2, January 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06966881-
abs.html.

Lindstrom:2016:EN

[Lin16a] Peter Lindstrom. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 22(8):1932, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2016/
08/07501981.pdf.

Linsen:2016:VAM

[Lin16b] Lars Linsen. Visual analysis of multi-run spatio-temporal simulations using isocontour similarity for projected views. IEEE Transactions on Visualization and Computer Graphics, 22(8):2037–2050, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2012:BMK

[LIRC12] Bongshin Lee, Petra Isenberg, Nathalie Henry Riche, and Sheelagh Carpendale. Beyond mouse and keyboard: Expanding design considerations for information visualization interactions. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2689–2698, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lv:2018:GCB

[LJH^{+}18] Na Lv, Zifei Jiang, Yan Huang, Xiangxu Meng, Gopi Meenakshisundaram, and Jingliang Peng. Generic content-based retrieval of marker-based motion capture data. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1969–1982, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07922586-
abs.html.

Lindholm:2014:BAR

[LJHY14] Stefan Lindholm, Daniel Jonsson, Charles Hansen, and Anders Ynnerman. Boundary aware reconstruction of scalar fields. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2447–2455, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876035-
abs.html.

Lefer:2004:HQA

[LJL04] Wilfrid Lefer, Bruno Jobard, and Claire Leduc. Highquality animation of 2D steady vector fields. IEEE Transactions on Visualization and Computer Graphics, 10(1):2–14, January/ February 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0002abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/01/
v0002.pdf.

Liu:2012:SIT

[LJWF12] Xiaopei Liu, Lei Jiang, Tien-Tsin Wong, and Chi-Wing Fu. Statistical invariance for texture synthesis. IEEE Transactions on Visualization and Computer Graphics, 18(11): 1836–1848, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2008:MCM

[LJWH08] Juncong Lin, Xiaogang Jin, Charlie Wang, and Kin-Chuen Hui. Mesh composition on models with arbitrary boundary topology. IEEE Transactions on Visualization and Computer Graphics, 14(3): 653–665, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2008/03/ttg2008030653s.
                                        zip.

Lai:2010:MDR

[LJX^{+}10] Yu-Kun Lai, Miao Jin, Xuexiang Xie, Ying He, Jonathan Palacios, Eugene Zhang, Shi-Min Hu, and Xianfeng Gu. Metric-driven RoSy field design and remeshing. IEEE Transactions on Visualization and Computer Graphics, 16(1):95–108, January/ February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liao:2012:ESP

[LJZ12] Qiqi Liao, Xiaogang Jin, and Wenting Zeng. Enhancing the symmetry and proportion of 3D face geometry. IEEE Transactions on Visualization and Computer Graphics, 18(10):1704–1716, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Laine:2011:ESV

[LK11] Samuli Laine and Tero Karras. Efficient sparse voxel octrees. IEEE Transactions on Visualization and Computer Graphics, 17(8):1048– 1059, August 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2009:RTD

[LKC09a] Sungkil Lee, Gerard Jounghyun Kim, and Seungmoon Choi. Real-time depth-of-field rendering using anisotropically filtered Mipmap interpolation. IEEE Transactions on Visualization and Computer Graphics, 15(3):453– 464, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2009:RTT

[LKC09b] Sungkil Lee, Gerard Jounghyun Kim, and Seungmoon Choi. Real-time tracking of visually attended objects in virtual environments and its application to LOD. IEEE Transactions on Visualization and Computer Graphics, 15(1):6– 19, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Langner:2019:MCV

[LKD19] Ricardo Langner, Ulrike Kister, and Raimund Dachselt.

Multiple coordinated views at large displays for multiple users: Empirical findings on user behavior, movements, and distances. IEEE Transactions on Visualization and Computer Graphics, 25 (1):608–618, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440846-
abs.html.

Lee:2016:HDP

[LKH^{+}16] S. Lee, S. Kim, Y. Hung, H. Lam, Y. Kang, and J. S. Yi. How do people make sense of unfamiliar visualizations?: a grounded model of novice’s information visualization sensemaking. IEEE Transactions on Visualization and Computer Graphics, 22 (1):499–508, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lefohn:2004:SNB

[LKHW04] Aaron E. Lefohn, Joe M. Kniss, Charles D. Hansen, and Ross T. Whitaker. A streaming narrow-band algorithm: Interactive computation and visualization of level sets. IEEE Transactions on Visualization and Computer Graphics, 10(4): 422–433, July/August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/04/v0422s.
pdf;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0422.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/04/v0422.pdf.

Laidlaw:2005:CVF

[LKJ^{+}05] David H. Laidlaw, Robert M.

Kirby, Cullen D. Jackson, J. Scott Davidson, Timothy S. Miller, Marco da Silva, William H. Warren, and Michael J. Tarr. Comparing 2D vector field visualization methods: a user study. IEEE Transactions on Visualization and Computer Graphics, 11(1):59–70, January/February 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0059abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0059.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0059.pdf.

Lee:2017:VDV

[LKK17] Sukwon Lee, Sung-Hee Kim, and Bum Chul Kwon. VLAT: Development of a visualization literacy assessment test. IEEE Transactions on Visualization and Computer Graphics, 23(1):551–560, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2015:HQD

[LKL^{+}15] Jungjin Lee, Younghui Kim, Sangwoo Lee, Bumki Kim, and Junyong Noh. Highquality depth estimation using an exemplar 3D model for stereo conversion. IEEE Transactions on Visualization and Computer Graphics, 21(7):835–847, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        07/07027857-abs.html.

Lynch:2018:CAB

[LKM^{+}18] Sean Dean Lynch, Richard Kulpa, Laurentius Antonius Meerhoff, Julien Pettre, Armel Cretual, and Anne-Helene Olivier. Collision avoidance behavior between walkers: Global and local motion cues. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2078–2088, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07955099-
                                        abs.html.

Leimkuhler:2018:PRT

[LKR^{+}18] Thomas Leimkuhler, Petr Kellnhofer, Tobias Ritschel, Karol Myszkowski, and Hans-Peter Seidel. Perceptual realtime 2D-to-3D conversion using cue fusion. IEEE Transactions on Visualization and Computer Graphics, 24(6): 2037–2050, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07926422-
abs.html.

Lee:2013:STM

[LKS13a] Bongshin Lee, Rubaiat Habib Kazi, and Greg Smith. SketchStory: Telling more engaging stories with data through freeform sketching. IEEE Transactions on Visualization and Computer Graphics, 19(12):2416– 2425, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lins:2013:NRT

[LKS13b] Lauro Lins, James T. Klosowski, and Carlos Scheidegger. Nanocubes for real-time exploration of spatiotemporal datasets. IEEE Transactions on Visualization and Computer Graphics, 19(12):2456–2465, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Latoschik:2019:AHS

[LKS^{+}19] Marc Erich Latoschik, Florian Kern, Jan-Philipp Stauffert, Andrea Bartl, Mario Botsch, and Jean-Luc Lugrin. Not alone here?! Scalability and user experience of embodied ambient crowds in distributed social virtual reality. IEEE Transactions on Visualization and Computer Graphics, 25(5):2134–2144, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643417/.

Lucas:2013:SLS

[LKT13] Blake C. Lucas, Michael Kazhdan, and Russell H. Taylor. Spring level sets: A deformable model representation to provide interoperability between meshes and level sets. IEEE Transactions on Visualization and Computer Graphics, 19(5):852–865, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2004:CSF

[LL04] Ping-Hsien Lin and Tong-Yee Lee. Camera-sampling field and its applications. IEEE Transactions on Visualization and Computer Graphics, 10 (3):241–251, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0241abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0241.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0241.pdf.

Lin:2005:MPM

[LL05] Chao-Hung Lin and Tong-Yee Lee. Metamorphosis of 3D polyhedral models using progressive connectivity transformations. IEEE Transactions on Visualization and Computer Graphics, 11(1):2– 12, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0002abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0002.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0002.pdf.

Levin:2014:SDI

[LL14] David Levin and Zohar Levi.

Shape deformation via interior rbf. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1062–1075, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2019:PMC

[LL19] Sukwon Lee and Sung-Hee Lee. Projective motion correction with contact optimization. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1746–1759, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8323196/.

Linsen:2006:VAG

[LLB^{+}06] Lars Linsen, Julia Locherbach, Matthias Berth, Dorte Becher, and Jorg Bernhardt. Visual analysis of gel-free proteome data. IEEE Transactions on Visualization and Computer Graphics, 12(4): 497–508, July/August 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2006/04/v0497s.
                                        zip.

Landge:2012:VNT

[LLB^{+}12] Aaditya G. Landge, Joshua A. Levine, Abhinav Bhatele, Katherine E. Isaacs, Todd Gamblin, Martin Schulz, Steve H. Langer, Peer-Timo Bremer, and Valerio Pascucci. Visualizing network traffic to understand the performance of massively parallel simulations. IEEE Transactions on Visualization and Computer Graphics, 18(12):2467– 2476, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Langbehn:2017:BCS

[LLBS17] Eike Langbehn, Paul Lubos, Gerd Bruder, and Frank Steinicke. Bending the curve: Sensitivity to bending of curved paths and application in room-scale VR. IEEE Transactions on Visualization and Computer Graphics, 23(4):1389–1398, April 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07833190-
abs.html.

Lin:2015:SCS

[LLC15] I-Chen Lin, Yu-Chien Lan, and Po-Wen Cheng. SI-Cut: Structural inconsistency analysis for image foreground extraction. IEEE Transactions on Visualization and Computer Graphics, 21(7):860–872, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
07/07018970-abs.html.

Lipsa:2011:FVF

[LLCD11] Dan Lipsa, Robert Laramee, Simon Cox, and Tudur Davies. FoamVis: Visualization of 2D foam simulation data. IEEE Transactions on Visualization and Computer Graphics, 17(12):2096– 2105, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Luo:2012:AFE

[LLCM12] Sheng-Jie Luo, Chun-Liang Liu, Bing-Yu Chen, and Kwan-Liu Ma. Ambiguityfree edge-bundling for interactive graph visualization. IEEE Transactions on Visualization and Computer Graphics, 18 (5):810–821, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lagae:2011:IGN

[LLD11] Ares Lagae, Sylvain Lefebvre, and Philip Dutre. Improving Gabor noise. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1096–1107, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liang:2017:OQP

[LLG17] Yun Liang, Yong-Jin Liu, and Diego Gutierrez. Objective quality prediction of image retargeting algorithms. IEEE Transactions on Visualization and Computer Graphics, 23(2):1099– 1110, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/02/07381692-
                                        abs.html.

Lin:2014:DRN

[LLHL14] Shih-Syun Lin, Chao-Hung Lin, Yan-Jhang Hu, and Tong-Yee Lee. Drawing road networks with mental maps. IEEE Transactions on Visualization and Computer Graphics, 20(9):1241–1252, September 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2017:SPI

[LLKN17] Jungjin Lee, Sangwoo Lee, Younghui Kim, and Junyong Noh. ScreenX: Public immersive theatres with uniform movie viewing experiences. IEEE Transactions on Visualization and Computer Graphics, 23(2):1124– 1138, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07414506-
abs.html.

Li:2006:TPS

[LLL06] Lewis W. F. Li, Frederick W. B. Li, and Rynson W. H. Lau. A trajectory-preserving synchronization method for collaborative visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):989–996, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindholm:2010:SCT

[LLL^{+}10] Stefan Lindholm, Patric Ljung, Claes Lundstrom, Anders Persson, and Anders Ynnerman. Spatial conditioning of transfer functions using local material distributions. IEEE Transactions on Visualization and Computer Graphics, 16(6):1301–1310, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lathen:2012:ATS

[LLL^{+}12] Gunnar Lathen, Stefan Lindholm, Reiner Lenz, Anders Persson, and Magnus Borga. Automatic tuning of spatially varying transfer functions for blood vessel visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2345– 2354, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liang:2019:FWO

[LLL^{+}19a] Wei Liang, Jingjing Liu, Yining Lang, Bing Ning, and Lap-Fai Yu. Functional workspace optimization via learning personal preferences from virtual experiences. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1836–1845, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642445/.

Liu:2019:NPD

[LLL^{+}19b] Shusen Liu, Zhimin Li, Tao Li, Vivek Srikumar, Valerio Pascucci, and Peer-Timo Bremer. NLIZE: A perturbationdriven visual interrogation tool for analyzing and interpreting natural language inference models. IEEE Transactions on Visualization and Computer Graphics, 25(1): 651–660, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08454904-
abs.html.

Li:2008:DPD

[LLLF08] Hongwei Li, Kui-Yip Lo, Man-Kang Leung, and Chi-Wing Fu. Dual Poisson-disk tiling: an efficient method for distributing features on arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 14(5): 982–998, September/October 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Le:2014:SCP

[LLLN^{+}14] Sang N. Le, Su-Jun Leow, Tuong-Vu Le-Nguyen, Conrado Ruiz, and Kok-Lim Low. Surface and contourpreserving origamic architecture paper pop-ups. IEEE Transactions on Visualization and Computer Graphics, 20 (2):276–288, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Law:2019:MIQ

[LLMB19] Po-Ming Law, Zhicheng Liu, Sana Malik, and Rahul C. Basole. MAQUI: Interweaving queries and pattern mining for recursive event sequence exploration. IEEE Transactions on Visualization and Computer Graphics, 25 (1):396–406, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440851-
                                        abs.html.

Lavoue:2019:PET

[LLPP19] Guillaume Lavoue, Michael Langer, Adrien Peytavie, and Pierre Poulin. A psychophysical evaluation of texture compression masking effects. IEEE Transactions on Visualization and Computer Graphics, 25(2):1336– 1346, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/02/08292880-
                                        abs.html.

Lundstrom:2007:UVM

[LLPY07] Claes Lundström, Patric Ljung, Anders Persson, and Anders Ynnerman. Uncertainty visualization in medical volume rendering using probabilistic animation. IEEE Transactions on Visualization and Computer Graphics, 13(6):1648–1655, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1648s.
avi.

Li:2017:EST

[LLQ^{+}17] Changyang Li, Wei Liang, Chris Quigley, Yibiao Zhao, and Lap-Fai Yu. Earthquake safety training through virtual drills. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1275–1284, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829397-
abs.html.

Liu:2018:LRM

[LLR18] Ping Liu, John Lewis, and Taehyun Rhee. Low-rank matrix completion to reconstruct incomplete rendering images. IEEE Transactions on Visualization and Computer Graphics, 24(8):2353– 2365, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07967887-
abs.html.

Linsen:2008:SEM

[LLRR08] Lars Linsen, Tran Van Long, Paul Rosenthal, and Stephan Rosswog. Surface extraction from multi-field particle volume data using multidimensional cluster visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1483– 1490, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lewiner:2004:AFD

[LLT04] Thomas Lewiner, Hélio Lopes, and Geovan Tavares. Applications of Forman’s discrete Morse theory to topology visualization and mesh compression. IEEE Transactions on Visualization and Computer Graphics, 10(5): 499–508, September/October 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/dl/trans/tg/2004/05/
v0499.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/05/v0499.pdf.

Lam:2006:NRF

[LLW06] Ping-Man Lam, Chi-Sing Leung, and Tien-Tsin Wong. Noise-resistant fitting for spherical harmonics. IEEE Transactions on Visualization and Computer Graphics, 12(2):254–265, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2015:PHM

[LLW15] Yufei Li, Yang Liu, and Wenping Wang. Planar hexagonal meshing for architecture. IEEE Transactions on Visualization and Computer Graphics, 21(1):95– 106, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06846311-
abs.html.

Li:2013:SMV

[LLWQ13] Bo Li, Xin Li, Kexiang Wang, and Hong Qin. Surface mesh to volumetric spline conversion with generalized polycubes. IEEE Transactions on Visualization and Computer Graphics, 19(9):1539– 1551, September 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lundstrom:2006:LHD

[LLY06] Claes Lundstrom, Patric Ljung, and Anders Ynnerman. Local histograms for design of transfer functions in direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 12(6):1570–1579, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2013:CAV

[LLY^{+}13] Shih-Syun Lin, Chao-Hung Lin, I-Cheng Yeh, Shu-Huai Chang, Chih-Kuo Yeh, and Tong-Yee Lee. Contentaware video retargeting using object-preserving warping. IEEE Transactions on Visualization and Computer Graphics, 19(10):1677– 1686, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2016:UAA

[LLZ^{+}16] M. Liu, S. Liu, X. Zhu, Q. Liao, F. Wei, and S. Pan. An uncertainty-aware approach for exploratory microblog retrieval. IEEE Transactions on Visualization and Computer Graphics, 22 (1):250–259, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lamure:1996:SGC

[LM96] Hervé Lamure and Dominique Michelucci. Solving geometric constraints by homotopy. IEEE Transactions on Visualization and Computer Graphics, 2(1):28–34, March 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0028.
pdf;   http://www.computer.
org/tvcg/tg1996/v0028abs.
htm.

Liu:2005:AUF

[LM05] Zhanping Liu and Robert J.

Moorhead. Accelerated unsteady flow line integral convolution. IEEE Transactions on Visualization and Computer Graphics, 11(2):113– 125, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lum:2002:HAS

[LMC02] Eric B. Lum, Kwan-Liu Ma, and John Clyne. A hardware-assisted scalable solution for interactive volume rendering of time-varying data. IEEE Transactions on Visualization and Computer Graphics, 8(3):286–301, July/September 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/03/v0286abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/03/
v0286.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/03/v0286.pdf.

Le:2012:LSD

[LMD12] Binh H. Le, Xiaohan Ma, and Zhigang Deng. Live speech driven head-and-eye motion generators. IEEE Transactions on Visualization and Computer Graphics, 18(11): 1902–1914, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2006:AES

[LMG06] Zhanping Liu, Robert Moorhead, and Joe Groner. An advanced evenly-spaced streamline placement algorithm. IEEE Transactions on Visualization and Computer Graphics, 12(5):965–972, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lam:2007:OUM

[LMK07] Heidi Lam, Tamara Munzner, and Robert Kincaid. Overview use in multiple visual information resolution interfaces. IEEE Transactions on Visualization and Computer Graphics, 13(6):1278– 1285, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2018:GEA

[LML^{+}18] S. S. Lin, C. C. Morace, C. H. Lin, L. F. Hsu, and T. Y. Lee. Generation of Escher arts with dual perception. IEEE Transactions on Visualization and Computer Graphics, 24(2):1103– 1113, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2003:IIS

[LMT^{+}03] Aidong Lu, Christopher J.

Morris, Joe Taylor, David S. Ebert, Charles Hansen, Penny Rheingans, and Mark Hartner. Illustrative interactive stipple rendering. IEEE Transactions on Visualization and Computer Graphics, 9 (2):127–138, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0127abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0127.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0127.pdf.

Liu:2017:VHD

[LMW^{+}17] Shusen Liu, Dan Maljovec, Bei Wang, Peer-Timo Bremer, and Valerio Pascucci. Visualizing high-dimensional data: Advances in the past decade. IEEE Transactions on Visualization and Computer Graphics, 23(3):1249– 1268, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07784854-
abs.html.

Lad:2006:VIR

[LMZ06] Mohit Lad, Dan Massey, and Lixia Zhang. Visualizing Internet routing changes. IEEE Transactions on Visualization and Computer Graphics, 12(6):1450–1460, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2014:SBD

[LMZ^{+}14] Jenny Hyunjung Lee, Kevin T. McDonnell, Alla Zelenyuk, Dan Imre, and Klaus Mueller. A structure-based distance metric for high-dimensional space exploration with multidimensional scaling. IEEE Transactions on Visualization and Computer Graphics, 20 (3):351–364, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liao:2016:NCC

[LNHS16] Jing Liao, Diego Nehab, Hugues Hoppe, and Pedro V. Sander. New controls for combining images in correspondence. IEEE Transactions on Visualization and Computer Graphics, 22(7): 1875–1885, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2008:DCT

[LNS08] Zhicheng Liu, Nancy Nersessian, and John Stasko. Distributed cognition as a theoretical framework for information visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1173–1180, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lopez:2016:TUM

[LODI16] D. López, L. Oehlberg, C. Doger, and T. Isenberg. Towards an understanding of mobile touch navigation in a stereoscopic viewing environment for 3D data exploration. IEEE Transactions on Visualization and Computer Graphics, 22(5):1616– 1629, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindstrom:2002:TSS

[LP02] Peter Lindstrom and Valerio Pascucci. Terrain simplification simplified: a general framework for viewdependent out-of-core visualization. IEEE Transactions on Visualization and Computer Graphics, 8(3):239–254, July/September 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/03/v0239abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/03/
v0239.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/03/v0239.pdf.

Loorak:2017:EPE

[LPCC17] Mona Hosseinkhani Loorak, Charles Perin, Christopher Collins, and Sheelagh Carpendale. Exploring the possibilities of embedding heterogeneous data attributes in familiar visualizations. IEEE Transactions on Visualization and Computer Graphics, 23 (1):581–590, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2019:VUT

[LPCRH19] Le Liu, Lace Padilla, Sarah H. Creem-Regehr, and Donald H. House. Visualizing uncertain tropical cyclone predictions using representative samples from ensembles of forecast tracks. IEEE Transactions on Visualization and Computer Graphics, 25(1): 882–891, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440837-
                                        abs.html.

Leung:2007:TB

[LPF^{+}07] Man-Kang Leung, Wai-Man Pang, Chi-Wing Fu, Tien-Tsin Wong, and Pheng-Ann Heng. Tileable BTF. IEEE Transactions on Visualization and Computer Graphics, 13(5):953–965, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/05/v0953s.
avi.

LeGoc:2019:DCD

[LPF^{+}19] Mathieu Le Goc, Charles Perin, Sean Follmer, Jean-Daniel Fekete, and Pierre Dragicevic. Dynamic composite data physicalization using wheeled micro-robots. IEEE Transactions on Visualization and Computer Graphics, 25 (1):737–747, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440836-
abs.html.

Liu:2012:PBM

[LPG12] Yang Liu, Balakrishnan Prabhakaran, and Xiaohu Guo. Point-based manifold harmonics. IEEE Transactions on Visualization and Computer Graphics, 18(10):1693– 1703, October 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2015:PDI

[LPG15] Heqi Lu, Romain Pacanowski, and Xavier Granier. Positiondependent importance sampling of light field luminaires. IEEE Transactions on Visualization and Computer Graphics, 21(2):241– 251, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/06905831-
abs.html.

Lamberti:2018:VCA

[LPG^{+}18] Fabrizio Lamberti, Gianluca Paravati, Valentina Gatteschi, Alberto Cannavo, and Paolo Montuschi. Virtual character animation based on affordable motion capture and reconfigurable tangible interfaces. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1742–1755, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07891591-
abs.html.

Lex:2013:EVR

[LPK^{+}13] Alexander Lex, Christian Partl, Denis Kalkofen, Marc Streit, Samuel Gratzl, Anne Mai Wassermann, Dieter Schmalstieg, and Hanspeter Pfister. Entourage: Visualizing relationships between biological pathways using contextual subsets. IEEE Transactions on Visualization and Computer Graphics, 19(12):2536– 2545, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Loorak:2016:TUV

[LPK^{+}16] M. H. Loorak, C. Perin, N. Kamal, M. Hill, and S. Carpendale. TimeSpan: Using visualization to explore temporal multi-dimensional data of stroke patients. IEEE Transactions on Visualization and Computer Graphics, 22 (1):409–418, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2012:ASD

[LPKK12] Sung-Ho Lee, Taejung Park, Jong-Hyeon Kim, and Chang-Hun Kim. Adaptive synthesis of distance fields. IEEE Transactions on Visualization and Computer Graphics, 18(7):1135–1145, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2011:AMD

[LPLT11] I-Chen Lin, Jen-Yu Peng, Chao-Chih Lin, and Ming-Han Tsai. Adaptive motion data representation with repeated motion analysis. IEEE Transactions on Visualization and Computer Graphics, 17 (4):527–538, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2006:TIE

[LPP^{+}06] Bongshin Lee, Cynthia S.

Parr, Catherine Plaisant, Benjamin B. Bederson, Vladislav D. Veksler, Wayne D. Gray, and Christopher Kotfila. TreePlus: Interactive exploration of networks with enhanced tree layouts. IEEE Transactions on Visualization and Computer Graphics, 12(6):1414–1426, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2014:TSA

[LPQF14] Xiaopei Liu, Wai-Man Pang, Jing Qin, and Chi-Wing Fu. Turbulence simulation by adaptive multi-relaxation lattice Boltzmann modeling. IEEE Transactions on Visualization and Computer Graphics, 20(2):289–302, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2013:WSM

[LPS^{+}13] Chuan Li, David Pickup, Thomas Saunders, Darren Cosker, David Marshall, Peter Hall, and Philip Willis. Water surface modeling from a single viewpoint video. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1242–1251, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2014:AAB

[LQLX14] Hongwei Lin, Yang Qin, Hongwei Liao, and Yunyang Xiong. Affine arithmeticbased B-spline surface intersection with GPU acceleration. IEEE Transactions on Visualization and Computer Graphics, 20(2):172– 181, February 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindemann:2011:AII

[LR11] Florian Lindemann and Timo Ropinski. About the influence of illumination models on image comprehension in direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 17(12):1922–1931, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lundstrom:2011:MTT

[LRF^{+}11] Claes Lundstrom, Thomas Rydell, Camilla Forsell, Anders Persson, and Anders Ynnerman. Multi-touch table system for medical visualization: Application to orthopedic surgery planning. IEEE Transactions on Visualization and Computer Graphics, 17(12):1775–1784, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2010:SVT

[LRKC10] Bongshin Lee, Nathalie Henry Riche, Amy K. Karlson, and Sheelash Carpendale. SparkClouds: Visualizing trends in tag clouds. IEEE Transactions on Visualization and Computer Graphics, 16(6):1182–1189, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2013:EVR

[LRM^{+}13] Cha Lee, Gustavo A. Rincon, Greg Meyer, Tobias Hollerer, and Doug A. Bowman. The effects of visual realism on search tasks in mixed reality simulation. IEEE Transactions on Visualization and Computer Graphics, 19(4): 547–556, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:1996:ICS

[LRN96] Tong-Yee Lee, C. S. Raghavendra, and John B. Nicholas. Image composition schemes for sort-last polygon rendering on 2D mesh multicomputers. IEEE Transactions on Visualization and Computer Graphics, 2(3):202– 217, September 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0202.
pdf;   http://www.computer.
org/tvcg/tg1996/v0202abs.
htm.

Larson:1997:VMT

[LRP97] G. W. Larson, H. Rushmeier, and C. Piatko. A visibility matching tone reproduction operator for high dynamic range scenes. IEEE Transactions on Visualization and Computer Graphics, 3(4):291– 306, October/December 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0291.
pdf;   http://www.computer.
org/tvcg/tg1997/v0291abs.
htm.

Li:2011:ALS

[LRZM11] Guiqing Li, Canjiang Ren, Jiahua Zhang, and Weiyin Ma. Approximation of loop subdivision surfaces for fast rendering. IEEE Transactions on Visualization and Computer Graphics, 17(4): 500–514, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2002:GCT

[LS02] J. Lee and S. Y. Shin. General construction of time-domain filters for orientation data. IEEE Transactions on Visualization and Computer Graphics, 8(2):119–128, April 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0119.

                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0119abs.
                                        htm.

Lau:2006:GEI

[LS06] Rynson W. H. Lau and Hans-Peter Seidel. Guest Editors’ introduction: Special section on ACM VRST. IEEE Transactions on Visualization and Computer Graphics, 12(2): 129–130, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2006/02/v0129.
                                        pdf.

Lamberti:2007:SBS

[LS07a] Fabrizio Lamberti and Andrea Sanna. A streamingbased solution for remote visualization of 3D graphics on mobile devices. IEEE Transactions on Visualization and Computer Graphics, 13(2): 247–260, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2007:IBS

[LS07b] Liya Li and Han-Wei Shen. Image-based streamline generation and rendering. IEEE Transactions on Visualization and Computer Graphics, 13 (3):630–640, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/.

Lee:2009:VET

[LS09] Teng-Yok Lee and Han-Wei Shen. Visualization and exploration of temporal trend relationships in multivariate time-varying data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1359–1366, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2010:MMV

[LS10] Zhicheng Liu and John Stasko. Mental models, visual reasoning and interaction in information visualization: a top-down perspective. IEEE Transactions on Visualization and Computer Graphics, 16(6):999–1008, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2013:AIP

[LS13a] Byeonghun Lee and Yeong-Gil Shin. Advanced interactive preintegrated volume rendering with a power series. IEEE Transactions on Visualization and Computer Graphics, 19(8):1264–1273, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2013:ELS

[LS13b] Teng-Yok Lee and Han-Wei Shen. Efficient local statistical analysis via integral histograms with discrete wavelet transform. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2693–2702, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2016:AAV

[LS16] X. Liu and H. Shen. Association analysis for visual exploration of multivariate scientific data sets. IEEE Transactions on Visualization and Computer Graphics, 22(1): 955–964, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2016:EEM

[LSB^{+}16] Y. Lu, M. Steptoe, S. Burke, H. Wang, J. Tsai, H. Davulcu, D. Montgomery, S. R. Corman, and R. Maciejewski. Exploring evolving media discourse through event cueing. IEEE Transactions on Visualization and Computer Graphics, 22(1):220–229, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Luboschik:2008:PBL

[LSC08] Martin Luboschik, Heidrun Schumann, and Hilko Cords. Particle-based labeling: Fast point-feature labeling without obscuring other visual features. IEEE Transactions on Visualization and Computer Graphics, 14(6):1237– 1244, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2018:ATP

[LSC^{+}18] M. Liu, J. Shi, K. Cao, J. Zhu, and S. Liu. Analyzing the training processes of deep generative models. IEEE Transactions on Visualization and Computer Graphics, 24 (1):77–87, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2009:GEI

[LSCN09] Ming C. Lin, Anthony Steed, and Carolina Cruz-Neira. Guest Editor’s introduction: Special section on the IEEE Virtual Reality Conference (VR). IEEE Transactions on Visualization and Computer Graphics, 15(3):353– 354, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lloyd:2007:ISP

[LSH07] Bryn Lloyd, Gábor Székely, and Matthias Harders. Identification of spring parameters for deformable object simulation. IEEE Transactions on Visualization and Computer Graphics, 13(5):1081– 1094, September/October 2007. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Livnat:1996:NOI

[LSJ96] Yarden Livnat, Han-Wei Shen, and Christopher R. Johnson. A near optimal isosurface extraction algorithm using the span space. IEEE Transactions on Visualization and Computer Graphics, 2 (1):73–84, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0073.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0073abs.

htm. See correction [Ano96b].

Lai:2015:DMH

[LSJ^{+}15] Duy-Quoc Lai, B. Sajadi, Shan Jiang, G. Meenakshisundaram, and A. Majumder. A distributed memory hierarchy and data management for interactive scene navigation and modification on tiled display walls. IEEE Transactions on Visualization and Computer Graphics, 21 (6):714–729, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lyu:2018:WVA

[LSK^{+}18] Geng Lyu, Xukun Shen, Taku Komura, Kartic Subr, and Lijun Teng. Widening viewing angles of automultiscopic displays using refractive inserts. IEEE Transactions on Visualization and Computer Graphics, 24(4):1554– 1563, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260976-
abs.html.

Liu:2017:TBA

[LSL^{+}17] Mengchen Liu, Jiaxin Shi, Zhen Li, Chongxuan Li, Jun Zhu, and Shixia Liu. Towards better analysis of deep convolutional neural networks. IEEE Transactions on Visualization and Computer Graphics, 23(1):91–100, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lum:2003:UMI

[LSM03] Eric B. Lum, Aleksander Stompel, and Kwan-Liu Ma. Using motion to illustrate static 3D shape—kinetic visualization. IEEE Transactions on Visualization and Computer Graphics, 9(2):115– 126, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0115abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0115.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0115.pdf.

Lex:2010:CAM

[LSPS10] Alexander Lex, Marc Streit, Christian Partl, and Dieter Schmalstieg. Comparative analysis of multidimensional, quantitative data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1027–1035, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2012:GOC

[LSPW12] Lin Lu, Feng Sun, Hao Pan, and Wenping Wang. Global optimization of centroidal Voronoi tessellation with Monte Carlo approach. IEEE Transactions on Visualization and Computer Graphics, 18(11):1880–1890, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2013:TIR

[LSR^{+}13] Dongping Li, Xin Sun, Zhong Ren, Stephen Lin, Yiying Tong, Baining Guo, and Kun Zhou. TransCut: Interactive rendering of translucent cutouts. IEEE Transactions on Visualization and Computer Graphics, 19(3): 484–494, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2009:SIA

[LSS09] Zhicheng Liu, John Stasko, and Timothy Sullivan. Sell-Trend: Inter-attribute visual analysis of temporal transaction data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1025– 1032, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lex:2011:VMV

[LSS^{+}11] Alexander Lex, Hans-Jörg Schulz, Marc Streit, Christian Partl, and Dieter Schmalstieg. VisBricks: Multiform visualization of large, inhomogeneous data. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2291–2300, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2013:PDV

[LSS13] Sungkil Lee, Mike Sips, and Hans-Peter Seidel. Perceptually driven visibility optimization for categorical data visualization. IEEE Transactions on Visualization and Computer Graphics, 19(10):1746– 1757, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Luo:2015:GCS

[LSS^{+}15] Sheng-Jie Luo, Ying-Tse Sun, I-Chao Shen, Bing-Yu Chen, and Yung-Yu Chuang. Geometrically consistent stereoscopic image editing using patch-based synthesis. IEEE Transactions on Visualization and Computer Graphics, 21 (1):56–67, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06824802-
abs.html.

Laha:2012:EIV

[LSSB12] Bireswar Laha, Kriti Sensharma, James D. Schiffbauer, and Doug A. Bowman. Effects of immersion on visual analysis of volume data. IEEE Transactions on Visualization and Computer Graphics, 18 (4):597–606, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2016:GEI

[LST^{+}16] Shixia Liu, Gerik Scheuermann, Shigeo Takahashi, Tim Dwyer, and Yingcai Wu. Guest Editors’ introduction: Special section on the IEEE Pacific Visualization Symposium 2015. IEEE Transactions on Visualization and Computer Graphics, 22(7): 1786–1787, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lungaro:2018:GAS

[LSV^{+}18] Pietro Lungaro, Rickard Sjoberg, Alfredo Jose Fanghella Valero, Ashutosh Mittal, and Konrad Tollmar. Gaze-aware streaming solutions for the next generation of mobile VR experiences. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1535–1544, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08269373-
abs.html.

Li:2017:SCS

[LSWZ17] Dongping Li, Tianjia Shao, Hongzhi Wu, and Kun Zhou. Shape completion from a single RGBD image. IEEE Transactions on Visualization and Computer Graphics, 23(7):1809–1822, July 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07451283-
abs.html.

Liao:2018:AWB

[LSY^{+}18] Xiangyun Liao, Weixin Si, Zhiyong Yuan, Hanqiu Sun, Jing Qin, Qiong Wang, and Pheng-Ann Heng. Animating wall-bounded turbulent smoke via filament-mesh particle-particle method. IEEE Transactions on Visualization and Computer Graphics, 24 (3):1260–1273, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lin:2018:RFM

[LSZ^{+}18] Minmin Lin, Tianjia Shao, Youyi Zheng, Niloy Jyoti Mitra, and Kun Zhou. Recovering functional mechanical assemblies from raw scans. IEEE Transactions on Visualization and Computer Graphics, 24(3):1354–1367, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lindstrom:1999:EMS

[LT99] P. Lindstrom and G. Turk. Evaluation of memoryless simplification. IEEE Transactions on Visualization and Computer Graphics, 5(2):98– 115, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0098.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0098abs.
                                        htm.

Lehmann:2010:DCS

[LT10] Dirk J. Lehmann and Holger Theisel. Discontinuities in continuous scatter plots. IEEE Transactions on Visualization and Computer Graphics, 16(6):1291–1300, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lehmann:2011:FCP

[LT11] Dirk J. Lehmann and Holger Theisel. Features in continuous parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 17(12):1912– 1921, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lehmann:2013:OSC

[LT13] Dirk J. Lehmann and Holger Theisel. Orthographic star coordinates. IEEE Transactions on Visualization and Computer Graphics, 19(12):2615– 2624, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lehmann:2016:OSP

[LT16] D. J. Lehmann and H. Theisel.

Optimal sets of projections of high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 22 (1):609–618, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lehmann:2018:LAM

[LT18] Dirk Joachim Lehmann and Holger Theisel. The LloydRelaxer: An approach to minimize scaling effects for multivariate projections.

IEEE Transactions on Visualization and Computer Graphics, 24(8):2424–2439, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07930525-
abs.html.

Losasso:2008:TWC

[LTKF08] Frank Losasso, Jerry Talton, Nipun Kwatra, and Ronald Fedkiw. Two-way coupled SPH and particle level set fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 14(4): 797–804, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040797s.
zip.

Lam:2018:BGT

[LTM18] H. Lam, M. Tory, and T. Munzner. Bridging from goals to tasks with design study analysis reports. IEEE Transactions on Visualization and Computer Graphics, 24 (1):435–445, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lowe:2005:PCS

[LTP^{+}05] Nick Lowe, Melanie Tory, Simeon Potts, Amitava Datta, and Torsten Möller. A parallel coordinates style interface for exploratory volume visualization. IEEE Transactions on Visualization and Computer Graphics, 11(1):71– 80, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0071abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0071.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0071.pdf.

Lawonn:2017:VEC

[LTPH17] Kai Lawonn, Erik Trostmann, Bernhard Preim, and Klaus Hildebrandt. Visualization and extraction of carvings for heritage conservation. IEEE Transactions on Visualization and Computer Graphics, 23 (1):801–810, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2008:FCV

[LTWH08] Guo-Shi Li, Xavier Tricoche, Daniel Weiskopf, and Charles D. Hansen. Flow Charts: Visualization of vector fields on arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 14(5):1067– 1080, September/October 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2012:MOP

[LvL12] Lei Liu and Robert van Liere. Modeling object pursuit for desktop virtual reality. IEEE Transactions on Visualization and Computer Graphics, 18(7):1017–1026, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lampe:2007:TLA

[LVRH07] Ove Daae Lampe, Ivan Viola, Nathalie Reuter, and Helwig Hauser. Two-level approach to efficient visualization of protein dynamics. IEEE Transactions on Visualization and Computer Graphics, 13(6):1616–1623, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1616s.
                                        avi.

Li:2006:RHO

[LVRL06] Wan-Chiu Li, Bruno Vallet, Nicolas Ray, and Bruno Lévy. Representing higherorder singularities in vector fields on piecewise linear surfaces. IEEE Transactions on Visualization and Computer Graphics, 12(5):1315– 1322, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Laramee:2004:III

[LvWJH04] Robert S. Laramee, Jarke J.

van Wijk, Bruno Jobard, and Helwig Hauser. ISA and IBFVS: Image space-based visualization of flow on surfaces. IEEE Transactions on Visualization and Computer Graphics, 10(6):637–648, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0637abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0637.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0637.pdf.

Lu:2018:GGI

[LWC^{+}18] Xuequan Lu, Shihao Wu, Honghua Chen, Sai-Kit Yeung, Wenzhi Chen, and Matthias Zwicker. GPF: GMM-inspired feature-preserving point set filtering. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2315–2326, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07974776-
abs.html.

Liu:2019:BTV

[LWC^{+}19] S. Liu, X. Wang, C. Collins, W. Dou, F. Ouyang, M. El-Assady, L. Jiang, and D. A. Keim. Bridging text visualization and mining: A taskdriven survey. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2482–2504, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liao:2018:CBV

[LWCC18] Hongsen Liao, Yingcai Wu, Li Chen, and Wei Chen. Cluster-based visual abstraction for multivariate scatterplots. IEEE Transactions on Visualization and Computer Graphics, 24(9):2531– 2545, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08047300-
abs.html.

Lee:1996:IMS

[LWCS96] Seungyong Lee, George Wolberg, Kyung-Yong Chwa, and Sung Yong Shin. Image metamorphosis with scattered feature constraints. IEEE Transactions on Visualization and Computer Graphics, 2(4):337– 354, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0337.
pdf;   http://www.computer.
org/tvcg/tg1996/v0337abs.
htm.

Liu:2017:PSI

[LWD^{+}17] Zhicheng Liu, Yang Wang, Mira Dontcheva, Matthew Hoffman, Seth Walker, and Alan Wilson. Patterns and sequences: Interactive exploration of clickstreams to understand common visitor paths. IEEE Transactions on Visualization and Computer Graphics, 23(1):321–330, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lamboray:2005:DST

[LWG05] Edouard Lamboray, Stephan Wurmlin, and Markus Gross. Data streaming in telepresence environments. IEEE Transactions on Visualization and Computer Graphics, 11(6):637–348, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2014:CEV

[LWL14] Wei-Li Lu, Yu-Shuen Wang, and Wen-Chieh Lin. Chess evolution visualization. IEEE Transactions on Visualization and Computer Graphics, 20 (5):702–713, May 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2017:SVA

[LWL^{+}17] Dongyu Liu, Di Weng, Yuhong Li, Jie Bao, Yu Zheng, Huamin Qu, and Yingcai Wu. SmartAdP: Visual analytics of large-scale taxi trajectories for selecting billboard locations. IEEE Transactions on Visualization and Computer Graphics, 23(1):1–10, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2018:VAF

[LWLM18] Yafeng Lu, Hong Wang, Steven Landis, and Ross Maciejewski. A visual analytics framework for identifying topic drivers in media events. IEEE Transactions on Visualization and Computer Graphics, 24(9):2501– 2515, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08037991-
abs.html.

Ljung:2006:FBV

[LWP^{+}06] Patric Ljung, Calle Winskog, Anders Persson, Claes Lundström, and Anders Ynnerman. Full body virtual autopsies using a stateof-the-art volume rendering pipeline. IEEE Transactions on Visualization and Computer Graphics, 12(5): 869–876, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:1997:SDI

[LWS97] S. Lee, G. Wolberg, and S. Y.

Shin. Scattered data interpolation with multilevel Bsplines. IEEE Transactions on Visualization and Computer Graphics, 3(3):228–244, July/September 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0228.
pdf;   http://www.computer.
org/tvcg/tg1997/v0228abs.
htm.

Lan:2017:ISS

[LWS^{+}17] Shouren Lan, Lisheng Wang, Yipeng Song, Yu ping Wang, Liping Yao, Kun Sun, Bin Xia, and Zongben Xu. Improving separability of structures with similar attributes in 2D transfer function design. IEEE Transactions on Visualization and Computer Graphics, 23(5):1546– 1560, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07423790.
pdf.

Lin:2007:HRA

[LWW^{+}07] Zhouchen Lin, Lifeng Wang, Yunbo Wang, Sing Bing Kang, and Tian Fang. High resolution animated scenes from stills. IEEE Transactions on Visualization and Computer Graphics, 13(3): 562–568, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2013:STE

[LWW^{+}13] Shixia Liu, Yingcai Wu, Enxun Wei, Mengchen Liu, and Yang Liu. StoryFlow: Tracking the evolution of stories. IEEE Transactions on Visualization and Computer Graphics, 19(12):2436– 2445, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2012:RBS

[LWYM12] Zhuwen Li, Song Wang, Jinhui Yu, and Kwan-Liu Ma. Restoration of brick and stone relief from single rubbing images. IEEE Transactions on Visualization and Computer Graphics, 18(2):177– 187, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2016:GLG

[LWZ^{+}16] Nannan Li, Shengfa Wang, Ming Zhong, Zhixun Su, and Hong Qin. Generalized localto-global shape feature detection based on graph wavelets. IEEE Transactions on Visualization and Computer Graphics, 22(9):2094–2106, ???? 2016. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liang:2018:PIP

[LWZ^{+}18] Yuan Liang, Xiting Wang, Song-Hai Zhang, Shi-Min Hu, and Shixia Liu. PhotoRecomposer: Interactive photo recomposition by cropping. IEEE Transactions on Visualization and Computer Graphics, 24(10):2728– 2742, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08078205-
abs.html.

Law:2017:VCU

[LWZQ17] Po-Ming Law, Wenchao Wu, Yixian Zheng, and Huamin Qu. VisMatchmaker: Cooperation of the user and the computer in centralized matching adjustment. IEEE Transactions on Visualization and Computer Graphics, 23(1): 231–240, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2017:ETM

[LXB17] Yijing Li, Hongyi Xu, and Jernej Barbic. Enriching triangle mesh animations with physically based simulation. IEEE Transactions on Visualization and Computer Graphics, 24(10):2301– 2313, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/10/07636982-
                                        abs.html.

Li:2017:VAA

[LXC^{+}17] Quan Li, Peng Xu, Yeuk Yin Chan, Yun Wang, Zhipeng Wang, Huamin Qu, and Xiaojuan Ma. A visual analytics approach for understanding reasons behind snowballing and comeback in MOBA games. IEEE Transactions on Visualization and Computer Graphics, 23(1):211–220, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2018:VDT

[LXL^{+}18] S. Liu, J. Xiao, J. Liu, X. Wang, J. Wu, and J. Zhu. Visual diagnosis of tree boosting methods. IEEE Transactions on Visualization and Computer Graphics, 24(1): 163–173, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2019:TPP

[LXR19] Dongyu Liu, Panpan Xu, and Liu Ren. TPFlow: Progressive partition and multidimensional pattern extraction for large-scale spatiotemporal data analysis. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1–11, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440840-
abs.html.

Luo:2018:PSS

[LXRY18] Bicheng Luo, Feng Xu, Christian Richardt, and Jun-Hai Yong. Parallax360: Stereoscopic 360^{∘} scene representation for head-motion parallax. IEEE Transactions on Visualization and Computer Graphics, 24(4):1545– 1553, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260916-
abs.html.

Lu:2018:RTV

[LXT18] Cewu Lu, Yao Xiao, and Chi-Keung Tang. Real-time video stylization using object flows. IEEE Transactions on Visualization and Computer Graphics, 24(6): 2051–2063, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/07920337-
abs.html.

Luo:2018:PBQ

[LXW^{+}18] Ran Luo, Weiwei Xu, Huamin Wang, Kun Zhou, and Yin Yang. Physics-based quadratic deformation using elastic weighting. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3188–3199, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8207608/.

Lin:2006:IVC

[LY06] Gang Lin and Thomas P.- Y. Yu. An improved vertex caching scheme for 3D mesh rendering. IEEE Transactions on Visualization and Computer Graphics, 12(4):640– 648, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2012:IIS

[LY12] Yugang Liu and Yizhou Yu. Interactive image segmentation based on level sets of probabilities. IEEE Transactions on Visualization and Computer Graphics, 18(2): 202–213, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Luo:2012:EVE

[LYK^{+}12] Dongning Luo, Jing Yang, Milos Krstajic, William Ribarsky, and Daniel A. Keim. EventRiver: Visually exploring text collections with temporal references. IEEE Transactions on Visualization and Computer Graphics, 18(1):93– 105, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2019:MSP

[LYL19] Tae Min Lee, Jong-Chul Yoon, and In-Kwon Lee. Motion sickness prediction in stereoscopic videos using 3D convolutional neural networks. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1919–1927, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642906/.

Li:2019:RNR

[LYLG19] Kun Li, Jingyu Yang, Yu-Kun Lai, and Daoliang Guo. Robust non-rigid registration with reweighted position and transformation sparsity. IEEE Transactions on Visualization and Computer Graphics, 25(6):2255–2269, June 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8353493/.

Lee:2010:FHQ

[LYS^{+}10] Byeonghun Lee, Jihye Yun, Jinwook Seo, Byonghyo Shim, Yeong-Gil Shin, and Bohyoung Kim. Fast highquality volume ray casting with virtual samplings. IEEE Transactions on Visualization and Computer Graphics, 16(6):1525–1532, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2016:OVA

[LYW^{+}16] Shixia Liu, Jialun Yin, Xiting Wang, Weiwei Cui, Kelei Cao, and Jian Pei. Online visual analytics of text streams. IEEE Transactions on Visualization and Computer Graphics, 22(11):2451–2466, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2008:TMH

[LYY08] Tong-Yee Lee, Shao-Wei Yen, and I-Cheng Yeh. Texture mapping with hard constraints using warping scheme. IEEE Transactions on Visualization and Computer Graphics, 14(2):382– 395, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lai:2016:NIF

[LYY^{+}16a] Chi-Fu William Lai, Sai-Kit Yeung, Xiaoqi Yan, Chi-Wing Fu, and Chi-Keung Tang. 3D navigation on impossible figures via dynamically reconfigurable maze. IEEE Transactions on Visualization and Computer Graphics, 22(10):

2275–2288, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2016:RDC

[LYY^{+}16b] Yong-Jin Liu, Cheng-Chi Yu, Min-Jing Yu, Kai Tang, and Deok-Soo Kim. A robust divide and conquer algorithm for progressive medial axes of planar shapes. IEEE Transactions on Visualization and Computer Graphics, 22(12): 2522–2536, December 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07364294-
abs.html.

Le:2013:MOF

[LZD13] Binh H. Le, Mingyang Zhu, and Zhigang Deng. Marker optimization for facial motion acquisition and deformation. IEEE Transactions on Visualization and Computer Graphics, 19(11):1859– 1871, November 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lai:2007:RFC

[LZH^{+}07] Yu-Kun Lai, Qian-Yi Zhou, Shi-Min Hu, Johannes Wallner, and Helmut Pottmann. Robust feature classification and editing. IEEE Transactions on Visualization and Computer Graphics, 13(1):34– 45, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2013:CDL

[LZH^{+}13] Yong-Jin Liu, Jun-Bin Zhang, Ji-Chun Hou, Ji-Cheng Ren, and Wei-Qing Tang. Cylinder detection in large-scale point cloud of pipeline plant. IEEE Transactions on Visualization and Computer Graphics, 19(10):1700–1707, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2018:CPP

[LZJ^{+}18] Lingjie Liu, Hongjie Zhang, Guangmei Jing, Yanwen Guo, Zhonggui Chen, and Wenping Wang. Correlation-preserving photo collage. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1956–1968, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/06/07927420-
                                        abs.html.

Liu:2016:KPR

[LZLS16] Zhiguang Liu, Liuyang Zhou, Howard Leung, and Hubert P. H. Shum. Kinect posture reconstruction based on a local mixture of Gaussian process models. IEEE Transactions on Visualization and Computer Graphics, 22(11):

2437–2450, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2013:TEO

[LZW^{+}13] Shao-Ping Lu, Song-Hai Zhang, Jin Wei, Shi-Min Hu, and Ralph R. Martin. Timeline editing of objects in video. IEEE Transactions on Visualization and Computer Graphics, 19(7):1218–1227, July 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Leng:2019:LDS

[LZZ^{+}19] B. Leng, C. Zhang, X. Zhou, C. Xu, and K. Xu. Learning discriminative 3D shape representations by view discerning networks. IEEE Transactions on Visualization and Computer Graphics, 25(10): 2896–2909, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mohammed:2018:AVT

[MAAB^{+}18] H. Mohammed, A. K. Al-Awami, J. Beyer, C. Cali, P. Magistretti, H. Pfister, and M. Hadwiger. Abstractocyte: A visual tool for exploring nanoscale astroglial cells. IEEE Transactions on Visualization and Computer Graphics, 24(1):853–861, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miller:2011:FBC

[MAF11] Christian Miller, Okan Arikan, and Don Fussell. Frankenrigs: Building character rigs from multiple sources. IEEE Transactions on Visualization and Computer Graphics, 17(8): 1060–1070, August 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mayerich:2008:VCM

[MAK08] David Mayerich, Louise Abbott, and John Keyser. Visualization of cellular and microvascular relationships. IEEE Transactions on Visualization and Computer Graphics, 14(6):1611–1618, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mehra:2014:SLD

[MAKM14] Ravish Mehra, Lakulish Antani, Sujeong Kim, and Dinesh Manocha. Source and listener directivity for interactive wave-based sound propagation. IEEE Transactions on Visualization and Computer Graphics, 20(4): 495–503, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maltz:2005:TPS

[Mal05] Martin S. Maltz. Tetrahedralization of point sets using expanding spheres. IEEE Transactions on Visualization and Computer Graphics, 11(1):102–109, January/ February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0102abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0102.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0102.pdf.

Mao:1996:SNR

[Mao96]  Xiaoyang Mao.  Splat-

ting of non rectilinear volumes through stochastic resampling. IEEE Transactions on Visualization and Computer Graphics, 2(2): 156–170, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0156.
pdf;   http://www.computer.
org/tvcg/tg1996/v0156abs.
htm.

Marai:2018:ACD

[Mar18]  G. E. Marai.  Activity-

centered domain characterization for problem-driven scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 24 (1):913–922, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Markovic:2016:BTB

[MAST16] Dejan Marković, Fabio Antonacci, Augusto Sarti, and Stefano Tubaro. 3D beam tracing based on visibility lookup for interactive acoustic modeling. IEEE Transactions on Visualization and Computer Graphics, 22(10): 2262–2274, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moloney:2011:SFP

[MAWM11] Brendan Moloney, Marco Ament, Daniel Weiskopf, and Torsten Moller. Sort-first parallel volume rendering. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1164–1177, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Max:1995:OMD

[Max95a] Nelson Max. Optical models for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 1(2):99–108, June 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0099.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0099abs.
                                        htm.

Max:1995:OSH

[Max95b] Nelson Max. Optimal sampling for hemicubes. IEEE Transactions on Visualization and Computer Graphics, 1 (1):60–76, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0060.
pdf;   http://www.computer.
org/tvcg/tg1995/v0060abs.
htm.

Mason:2001:GRS

[MB01] A. E. W. Mason and E. H.

Blake. A graphical representation of the state spaces of hierarchical level-of-detail scene descriptions. IEEE Transactions on Visualization and Computer Graphics, 7(1):70– ??, January/March 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0070.
pdf;   http://www.computer.
org/tvcg/tg2001/v0070abs.
htm.

Mahovsky:2003:AJB

[MB03] Jeffrey Mahovsky and Luigi Benedicenti. An architecture for Java-based realtime distributed visualization. IEEE Transactions on Visualization and Computer Graphics, 9(4):570–579, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/04/v0570abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/04/
                                        v0570.pdf.

Magnus:2018:IDV

[MB18a] J. G. Magnus and S. Bruckner. Interactive dynamic volume illumination with refraction and caustics. IEEE Transactions on Visualization and Computer Graphics, 24 (1):984–993, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meister:2018:PLO

[MB18b] Daniel Meister and Jiří Bittner. Parallel locally-ordered clustering for bounding volume hierarchy construction. IEEE Transactions on Visualization and Computer Graphics, 24(3):1345–1353, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Major:2019:GEU

[MB19a] Timothy Major and Rahul C. Basole. Graphicle: Exploring units, networks, and context in a blended visualization approach. IEEE Transactions on Visualization and Computer Graphics, 25(1): 576–585, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440854-
abs.html.

Mueller:2019:MEC

[MB19b] K. Mueller and D. Bowman.

Message from the Editor-in-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3049, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meyer:2012:VDA

[MBH^{+}12] Joerg Meyer, E. Wes Bethel, Jennifer L. Horsman, Susan S. Hubbard, Harinarayan Krishnan, Alexandru Romosan, Elizabeth H. Keating, Laura Monroe, Richard Strelitz, Phil Moore, Glenn Taylor, Ben Torkian, Timothy C. Johnson, and Ian Gorton. Visual data analysis as an integral part of environmental management. IEEE Transactions on Visualization and Computer Graphics, 18(12):2088– 2094, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mehler:2006:SAN

[MBL^{+}06] Andrew Mehler, Yunfan Bao, Xin Li, Yue Wang, and Steven Skiena. Spatial analysis of news sources. IEEE Transactions on Visualization and Computer Graphics, 12(5): 765–772, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mahrous:2004:TST

[MBS^{+}04] Karim Mahrous, Janine Bennett, Gerik Scheuermann, Bernd Hamann, and Kenneth I. Joy. Topological segmentation in threedimensional vector fields. IEEE Transactions on Visualization and Computer Graphics, 10(2):198–205, March/ April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2004/
                                        02/v0198abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        trans/tg/2004/02/v0198.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/02/
                                        v0198.pdf.

Maciel:2007:ECD

[MBT07] Anderson Maciel, Ronan Boulic, and Daniel Thalmann. Efficient collision detection within deforming spherical sliding contact. IEEE Transactions on Visualization and Computer Graphics, 13(3): 518–529, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meneveaux:2018:RRO

[MBT^{+}18] Daniel Meneveaux, Benjamin Bringier, Emmanuelle Tauzia, Mickaël Ribardière, and Lionel Simonot. Rendering rough opaque materials with interfaced Lambertian microfacets. IEEE Transactions on Visualization and Computer Graphics, 24(3):1368– 1380, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mohan:2007:TCL

[MBW^{+}07] Ankit Mohan, Reynold Bailey, Jonathan Waite, Jack Tumblin, Cindy Grimm, and Bobby Bodenheimer. Tabletop computed lighting for practical digital photography. IEEE Transactions on Visualization and Computer Graphics, 13(4):652– 662, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McMahan:2012:EDF

[MBZB12] Ryan P. McMahan, Doug A.

Bowman, David J. Zielinski, and Rachael B. Brady. Evaluating display fidelity and interaction fidelity in a virtual reality game. IEEE Transactions on Visualization and Computer Graphics, 18 (4):626–633, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Munteanu:2010:SIC

[MCA^{+}10] Adrian Munteanu, Dan C. Cernea, Alin Alecu, Jan Cornelis, and Peter Schelkens. Scalable L-infinite coding of meshes. IEEE Transactions on Visualization and Computer Graphics, 16(3):513– 528, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Martinet:2012:ISM

[MCG12] Anthony Martinet, Gery Casiez, and Laurent Grisoni. Integrality and separability of multitouch interaction techniques in 3D manipulation tasks. IEEE Transactions on Visualization and Computer Graphics, 18(3):369– 380, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Marchesin:2010:VDS

[MCHM10] Stephane Marchesin, Cheng-Kai Chen, Chris Ho, and Kwan-Liu Ma. Viewdependent streamlines for 3D vector fields. IEEE Transactions on Visualization and Computer Graphics, 16(6):1578–1586, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McKeon:2009:HWI

[McK09] Matt McKeon. Harnessing the Web information ecosystem with Wiki-based visualization dashboards. IEEE Transactions on Visualization and Computer Graphics, 15(6):1081–1088, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Martin:2012:DIV

[MCK12] Tobias Martin, Elaine Cohen, and Robert M. Kirby. Direct isosurface visualization of hex-based high-order geometry and attribute representations. IEEE Transactions on Visualization and Computer Graphics, 18(5):753–766, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2006:AVS

[MCP^{+}06] Jiyong Ma, Ron Cole, Bryan Pellom, Wayne Ward, and Barbara Wise. Accurate visible speech synthesis based on concatenating variable length motion capture data. IEEE Transactions on Visualization and Computer Graphics, 12(2):266–276, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mosaliganti:2008:RCB

[MCS^{+}08] Kishore Mosaliganti, Lee Cooper, Richard Sharp, Raghu Machiraju, Gustavo Leone, Kun Huang, and Joel Saltz. Reconstruction of cellular biological structures from optical microscopy data. IEEE Transactions on Visualization and Computer Graphics, 14(4):863– 876, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2012:SQM

[MD12] Xiaohan Ma and Zhigang Deng. A statistical quality model for data-driven speech animation. IEEE Transactions on Visualization and Computer Graphics, 18(11): 1915–1927, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Molla:2018:EMB

[MDB18] Eray Molla, Henrique Galvan Debarba, and Ronan Boulic. Egocentric mapping of body surface constraints. IEEE Transactions on Visualization and Computer Graphics, 24(7):2089–2102, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07942001.
pdf.

Micallef:2012:AEV

[MDF12] Luana Micallef, Pierre Dragicevic, and Jean-Daniel Fekete.

Assessing the effect of visualizations on Bayesian reasoning through crowdsourcing. IEEE Transactions on Visualization and Computer Graphics, 18(12):2536– 2545, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Merillou:2000:BPI

[MDG00] S. Mérillou, J.-M. Dischler, and D. Ghazanfarpour. A BRDF postprocess to integrate porosity on rendered surfaces. IEEE Transactions on Visualization and Computer Graphics, 6(4):306–??, October/December 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0306.
pdf;   http://www.computer.
org/tvcg/tg2000/v0306abs.
htm.

Michel:2007:LBA

[MDHB^{+}07] Frank Michel, Eduard Deines, Martin Hering-Bertram, Christoph Garth, and Hans Hagen. Listener-based analysis of surface importance for acoustic metrics. IEEE Transactions on Visualization and Computer Graphics, 13(6):1680– 1687, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mao:2007:SDV

[MDL07] Yi Mao, Joshua Dillon, and Guy Lebanon. Sequential document visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6):1208–1215, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miranda:2017:UPC

[MDL^{+}17] Fabio Miranda, Harish Doraiswamy, Marcos Lage, Kai Zhao, Bruno Gonçalves, Luc Wilson, Mondrian Hsieh, and Cláudio T. Silva. Urban pulse: Capturing the rhythm of cities. IEEE Transactions on Visualization and Computer Graphics, 23(1):791– 800, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miranda:2019:SAM

[MDL^{+}19] F. Miranda, H. Doraiswamy, M. Lage, L. Wilson, M. Hsieh, and C. T. Silva. Shadow accrual maps: Efficient accumulation of city-scale shadows over time. IEEE Transactions on Visualization and Computer Graphics, 25(3):1559– 1574, March 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Marchesin:2010:PPO

[MDM10] Stephane Marchesin, Jean-Michel Dischler, and Catherine Mongenet. Per-pixel opacity modulation for feature enhancement in volume rendering. IEEE Transactions on Visualization and Computer Graphics, 16(4):560– 570, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mellado:2016:RSE

[MDS16] Nicolas Mellado, Matteo Dellepiane, and Roberto Scopigno. Relative scale estimation and 3D registration of multi-modal geometry using growing least squares. IEEE Transactions on Visualization and Computer Graphics, 22 (9):2160–2173, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meulemans:2017:SMG

[MDS^{+}17] Wouter Meulemans, Jason Dykes, Aidan Slingsby, Cagatay Turkay, and Jo Wood. Small multiples with gaps. IEEE Transactions on Visualization and Computer Graphics, 23(1):381–390, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miao:2018:MVS

[MDS^{+}18] H. Miao, E. De Llano, J. Sorger, Y. Ahmadi, T. Kekic, T. Isenberg, M. E. Gröller, I. Bari i, and I. Viola. Multiscale visualization and scale-adaptive modification of DNA nanostructures. IEEE Transactions on Visualization and Computer Graphics, 24(1):1014–1024, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McDonnel:2009:TUG

[ME09] Bryan McDonnel and Niklas Elmqvist. Towards utilizing GPUs in information visualization: a model and implementation of image-space operations. IEEE Transactions on Visualization and Computer Graphics, 15(6):1105– 1112, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mirzargar:2011:QIV

[ME11a] Mahsa Mirzargar and Alireza Entezari. Quasi interpolation with Voronoi splines. IEEE Transactions on Visualization and Computer Graphics, 17(12):1832–1841, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moran:2011:VAD

[ME11b] Patrick Moran and David Ellsworth. Visualization of AMR data with multilevel dual-mesh interpolation. IEEE Transactions on Visualization and Computer Graphics, 17(12):1862– 1871, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2018:VFB

[ME18] Bo Ma and Alireza Entezari.

Volumetric feature-based classification and visibility analysis for transfer function design. IEEE Transactions on Visualization and Computer Graphics, 24(12):3253– 3267, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8119816/.

Ma:2019:IFV

[ME19] Bo Ma and Alireza Entezari.

An interactive framework for visualization of weather forecast ensembles. IEEE Transactions on Visualization and Computer Graphics, 25(1): 1091–1101, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440052-
abs.html.

Misztal:2014:MFI

[MEB^{+}14] Marek Krzysztof Misztal, Kenny Erleben, Adam Bargteil, Jens Fursund, Brian Bunch Christensen, Jakob Andreas Baerentzen, and Robert Bridson. Multiphase flow of immiscible fluids on unstructured moving meshes. IEEE Transactions on Visualization and Computer Graphics, 20 (1):4–16, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mehra:2017:VVR

[Meh17] Ravish Mehra. The 2016 VGTC Virtual Reality: Best Dissertation Award. IEEE Transactions on Visualization and Computer Graphics, 23(4):xii, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/04/07878158-
                                        abs.html.

Meng:2011:VCR

[MES^{+}11] Tai Meng, Alireza Entezari, Benjamin Smith, Torsten Moller, Daniel Weiskopf, and Arthur E. Kirkpatrick. Visual comparability of 3D regular sampling and reconstruction. IEEE Transactions on Visualization and Computer Graphics, 17(10):1420–1432, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Madhavan:2014:DWB

[MEV^{+}14] Krishna Madhavan, Niklas Elmqvist, Mihaela Vorvoreanu, Xin Chen, Yuetling Wong, Hanjun Xian, Zhihua Dong, and Aditya Johri. DIA2: Web-based cyberinfrastructure for visual analysis of funding portfolios. IEEE Transactions on Visualization and Computer Graphics, 20(12):1823–1832, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876046-
abs.html.

Moehring:2011:NIM

[MF11] Mathias Moehring and Bernd Froehlich. Natural interaction metaphors for functional validations of virtual car models. IEEE Transactions on Visualization and Computer Graphics, 17(9):1195–1208, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muller:2009:BPE

[MFS^{+}09] Christoph Müller, Steffen Frey, Magnus Strengert, Carsten Dachsbacher, and Thomas Ertl. Best paper of EGPGV — Eurographics Symposium on Parallel Graphics and Visualization, Guest Editor Jean Favre: a compute unified system architecture for graphics clusters incorporating data locality. IEEE Transactions on Visualization and Computer Graphics, 15(4):605– 617, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meka:2017:LUG

[MFZ^{+}17] Abhimitra Meka, Gereon Fox, Michael Zollhofer, Christian Richardt, and Christian Theobalt. Live userguided intrinsic video for static scenes. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2447–2454, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/11/08008808-
                                        abs.html.

Mayorga:2013:SOO

[MG13] Adrian Mayorga and Michael Gleicher. Splatterplots: Overcoming overdraw in scatter plots. IEEE Transactions on Visualization and Computer Graphics, 19(9):1526– 1538, September 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meuschke:2019:VAA

[MGB^{+}19] Monique Meuschke, Tobias Gunther, Philipp Berg, Ralph Wickenhofer, Bernhard Preim, and Kai Lawonn. Visual analysis of aneurysm data using statistical graphics. IEEE Transactions on Visualization and Computer Graphics, 25(1): 997–1007, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440110-
abs.html.

Matkovic:2010:IVA

[MGJ^{+}10] Krešimir Matković, Denis Gracanin, Mario Jelovic, Andreas Ammer, Alan Lez, and Helwig Hauser. Interactive visual analysis of multiple simulation runs using the simulation model view: Understanding and tuning of an electronic unit injector. IEEE Transactions on Visualization and Computer Graphics, 16(6):1449–1457, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Matkovic:2008:IVS

[MGJH08] Krešimir Matković, Denis Gracanin, Mario Jelovic, and Helwig Hauser. Interactive visual steering — rapid visual prototyping of a common rail injection system. IEEE Transactions on Visualization and Computer Graphics, 14(6):1699–1706, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Matkovic:2009:IVA

[MGKH09] Krešimir Matković, Denis Gracanin, Borislav Klarin, and Helwig Hauser. Interactive visual analysis of complex scientific data as families of data surfaces. IEEE Transactions on Visualization and Computer Graphics, 15(6):1351–1358, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Marai:2007:AJM

[MGL07] G. Elisabeta Marai, Cindy M. Grimm, and David H. Laidlaw. Arthrodial joint markerless cross-parameterization and biomechanical visualization. IEEE Transactions on Visualization and Computer Graphics, 13(5):1095–1104, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/.

Muelder:2009:VAI

[MGM09] Chris Muelder, Francois Gygi, and Kwan-Liu Ma. Visual analysis of inter-process communication for large-scale parallel computing. IEEE Transactions on Visualization and Computer Graphics, 15(6):1129–1136, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Merry:2014:MLS

[MGM14] Bruce Merry, James Gain, and Patrick Marais. Moving least-squares reconstruction of large models with GPUs. IEEE Transactions on Visualization and Computer Graphics, 20(2):249– 261, February 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mccurdy:2019:FEI

[MGM19] Nina Mccurdy, Julie Gerdes, and Miriah Meyer. A framework for externalizing implicit error using visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1):925– 935, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08449328-
abs.html.

Mendhurwar:2018:DPS

[MGMP18] Kaustubha Mendhurwar, Qing Gu, Sudhir Mudur, and Tiberiu Popa. The discriminative power of shape an empirical study in time series matching. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1799–1813, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2018/05/07892972-
                                        abs.html.

Mirhosseini:2019:IVC

[MGO^{+}19] Seyedkoosha Mirhosseini, Ievgeniia Gutenko, Sushant Ojal, Joseph Marino, and Arie Kaufman. Immersive virtual colonoscopy. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2011–2021, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642297/.

Marsh:2006:NAS

[MGPH06] James Marsh, Mashhuda Glencross, Steve Pettifer, and Roger Hubbold. A network architecture supporting consistent rich behavior in collaborative interactive applications. IEEE Transactions on Visualization and Computer Graphics, 12(3):405– 416, May/June 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/.

Matkovic:2014:VAC

[MGS^{+}14] Kresimir Matkovic, Denis Gracanin, Rainer Splechtna, Mario Jelovic, Benedikt Stehno, Helwig Hauser, and Werner Purgathofer. Visual analytics for complex engineering systems: Hybrid visual steering of simulation ensembles. IEEE Transactions on Visualization and Computer Graphics, 20(12):1803– 1812, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876045-
abs.html.

Muller:2010:SRV

[MGW10] Thomas Muller, Sebastian Grottel, and Daniel Weiskopf. Special relativistic visualization by local ray tracing. IEEE Transactions on Visualization and Computer Graphics, 16(6):1243–1250, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McGuire:2010:GEI

[MH10] Morgan McGuire and Eric Haines. Guest Editor’s introduction: Special section on the symposium on interactive 3D graphics and games (I3D). IEEE Transactions on Visualization and Computer Graphics, 16(5): 705–706, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Matzen:2018:DVS

[MHD^{+}18] L. E. Matzen, M. J. Haass, K. M. Divis, Z. Wang, and A. T. Wilson. Data visualization saliency model: A tool for evaluating abstract data visualizations. IEEE Transactions on Visualization and Computer Graphics, 24(1): 563–573, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muigg:2011:IVV

[MHDG11] Philipp Muigg, Markus Hadwiger, Helmut Doleisch, and Eduard Gröller. Interactive volume visualization of general polyhedral grids. IEEE Transactions on Visualization and Computer Graphics, 17(12):2115–2124, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muigg:2007:SHU

[MHDH07] Philipp Muigg, Markus Hadwiger, Helmut Doleisch, and Helwig Hauser. Scalable hybrid unstructured and structured grid raycasting. IEEE Transactions on Visualization and Computer Graphics, 13(6):1592–1599, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1592s.
                                        zip.

Malik:2010:CVP

[MHG10] Muhammad Muddassir Malik, Christoph Heinzl, and M. Eduard Gröller. Comparative visualization for parameter studies of dataset series. IEEE Transactions on Visualization and Computer Graphics, 16(5):829–840, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maciejewski:2011:FHP

[MHR^{+}11] Ross Maciejewski, Ryan Hafen, Stephen Rudolph, Stephen G. Larew, Michael A. Mitchell, William S. Cleveland, and David S. Ebert. Forecasting hotspots — a predictive analytics approach. IEEE Transactions on Visualization and Computer Graphics, 17(4):440–453, April 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mackinlay:2007:SMA

[MHS07] Jock Mackinlay, Pat Hanrahan, and Chris Stolte. Show me: Automatic presentation for visual analysis. IEEE Transactions on Visualization and Computer Graphics, 13(6):1137–1144, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meghdadi:2013:IES

[MI13] Amir H. Meghdadi and Pourang Irani. Interactive exploration of surveillance video through action shot summarization and trajectory visualization. IEEE Transactions on Visualization and Computer Graphics, 19(12):2119– 2128, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mine:2013:KSW

[Min13] Mark Mine. Keynote speaker: Welcome to the future! Technology and innovation at Disney. IEEE Transactions on Visualization and Computer Graphics, 19(4):xiii, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moser:2015:SES

[MIO^{+}15] K. Moser, Y. Itoh, K. Oshima, J. E. Swan, G. Klinker, and C. Sandor. Subjective evaluation of a semiautomatic optical see-through head-mounted display calibration technique. IEEE Transactions on Visualization and Computer Graphics, 21(4): 491–500, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mindek:2018:VMP

[MiS^{+}18] P. Mindek, D. Kou il, J. Sorger, D. Toloudis, B. Lyons, G. Johnson, M. E. Gröller, and I. Viola. Visualization multi-pipeline for communicating biology. IEEE Transactions on Visualization and Computer Graphics, 24 (1):883–892, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McGuffin:2009:ITS

[MJ09] Michael J. McGuffin and Igor Jurisica. Interaction techniques for selecting and manipulating subgraphs in network visualizations. IEEE Transactions on Visualization and Computer Graphics, 15(6):937–944, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mehta:2006:DVD

[MJK06] Ketan Mehta and T. J.

Jankun-Kelly. Detection and visualization of defects in 3D unstructured models of nematic liquid crystals. IEEE Transactions on Visualization and Computer Graphics, 12(5):1045–1052, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McLoughlin:2013:SME

[MJL^{+}13] Tony McLoughlin, Mark W.

Jones, Robert S. Laramee, Rami Malki, Ian Masters, and Charles D. Hansen. Similarity measures for enhancing interactive streamline seeding. IEEE Transactions on Visualization and Computer Graphics, 19(8):1342–1353, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maciejewski:2013:AAS

[MJW^{+}13] R. Maciejewski, Yun Jang, Insoo Woo, H. Jänicke, K. P. Gaither, and D. S. Ebert. Abstracting attribute space for transfer function exploration and design. IEEE Transactions on Visualization and Computer Graphics, 19(1):94– 107, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mayerich:2009:HAS

[MK09] David Mayerich and John Keyser. Hardware accelerated segmentation of complex volumetric filament networks. IEEE Transactions on Visualization and Computer Graphics, 15(4):670– 681, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mahmudi:2013:ALS

[MK13a] Mentar Mahmudi and Marcelo Kallmann. Analyzing locomotion synthesis with featurebased motion graphs. IEEE Transactions on Visualization and Computer Graphics, 19 (5):774–786, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Matvienko:2013:MED

[MK13b] Victor Matvienko and Jens Kruger. A metric for the evaluation of dense vector field visualizations. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1122–1132, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Menk:2013:TCR

[MK13c] C. Menk and R. Koch. Truthful color reproduction in spatial augmented reality applications. IEEE Transactions on Visualization and Computer Graphics, 19(2):236– 248, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Marino:2016:PVT

[MK16] J. Marino and A. Kaufman.

Planar visualization of treelike structures. IEEE Transactions on Visualization and Computer Graphics, 22(1): 906–915, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mashima:2012:VDD

[MKH12] Daisuke Mashima, Stephen G.

Kobourov, and Yifan Hu. Visualizing dynamic data with maps. IEEE Transactions on Visualization and Computer Graphics, 18(9):1424– 1437, September 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Magnor:2005:RVP

[MKHD05] Marcus Magnor, Gordon Kindlmann, Charles Hansen, and Neb Duric. Reconstruction and visualization of planetary nebulae. IEEE Transactions on Visualization and Computer Graphics, 11(5): 485–496, September/October 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mansmann:2007:VAN

[MKN^{+}07] Florian Mansmann, Daniel A. Keim, Stephen C. North, Brian Rexroad, and Daniel Sheleheda. Visual analysis of network traffic for resource planning, interactive monitoring, and interpretation of security threats. IEEE Transactions on Visualization and Computer Graphics, 13(6):1105–1112, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Merenda:2018:ARI

[MKT^{+}18] Coleman Merenda, Hyungil Kim, Kyle Tanous, Joseph L. Gabbard, Blake Feichtl, Teruhisa Misu, and Chihiro Suga. Augmented reality interface design approaches for goal-directed and stimulus-driven driving tasks. IEEE Transactions on Visualization and Computer Graphics, 24(11):2875– 2885, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8466859/.

Meyer:2007:TAQ

[MKW07] Miriah Meyer, Robert M.

Kirby, and Ross Whitaker. Topology, accuracy, and quality of isosurface meshes using dynamic particles. IEEE Transactions on Visualization and Computer Graphics, 13(6):1704–1711, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Molchanov:2019:SPS

[ML19] Vladimir Molchanov and Lars Linsen. Shape-preserving star coordinates. IEEE Transactions on Visualization and Computer Graphics, 25(1): 449–458, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440845-
abs.html.

McCurdy:2016:PVS

[MLCM16] N. McCurdy, J. Lein, K. Coles, and M. Meyer. Poemage: Visualizing the sonic topology of a poem. IEEE Transactions on Visualization and Computer Graphics, 22(1): 439–448, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miranda:2018:TRA

[MLKS18] Fabio Miranda, Lauro Lins, James T. Klosowski, and Claudio T. Silva. TopKube: A rank-aware data cube for realtime exploration of spatiotemporal data. IEEE Transactions on Visualization and Computer Graphics, 24(3): 1394–1407, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Monroe:2013:TES

[MLL^{+}13] Megan Monroe, Rongjian Lan, Hanseung Lee, Catherine Plaisant, and Ben Shneiderman. Temporal event sequence simplification. IEEE Transactions on Visualization and Computer Graphics, 19(12):2227–2236, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2012:LLD

[MLMF12] Joyce Ma, Isaac Liao, Kwan-Liu Ma, and Jennifer Frazier. Living liquid: Design and evaluation of an exploratory visualization tool for museum visitors. IEEE Transactions on Visualization and Computer Graphics, 18(12):2799– 2808, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muhlbacher:2018:TSA

[MLMP18] T. Mühlbacher, L. Linhardt, T. Möller, and H. Piringer. TreePOD: Sensitivity-aware selection of Pareto-optimal decision trees. IEEE Transactions on Visualization and Computer Graphics, 24(1): 174–183, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Murcia-Lopez:2018:CVP

[MLS18] Maria Murcia-Lopez and Anthony Steed. A comparison of virtual and physical training transfer of bimanual assembly tasks. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1574–1583, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260860-
abs.html.

Muelder:2008:RGL

[MM08] Chris Muelder and Kwan-Liu Ma. Rapid graph layout using space filling curves. IEEE Transactions on Visualization and Computer Graphics, 14(6):1301–1308, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Merrell:2011:MSG

[MM11] Paul Merrell and Dinesh Manocha. Model synthesis: a general procedural modeling algorithm. IEEE Transactions on Visualization and Computer Graphics, 17(6): 715–728, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McGuire:2017:PT

[MM17] Morgan McGuire and Michael Mara. Phenomenological transparency. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1465–1478, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/05/07828148-
                                        abs.html.

McKenna:2014:DAF

[MMAM14] Sean McKenna, Dominika Mazur, James Agutter, and Miriah Meyer. Design activity framework for visualization design. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2191–2200, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/

csdl/trans/tg/2014/12/06876000-
abs.html.

Marai:2019:PRA

[MMB^{+}19] G. Elisabeta Marai, Chihua Ma, Andrew Thomas Burks, Filippo Pellolio, Guadalupe Canahuate, David M. Vock, Abdallah S. R. Mohamed, and Clifton David Fuller. Precision risk analysis of cancer therapy with interactive nomograms and survival plots. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1732–1745, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8320386/.

Mora:2009:VCG

[MMCE09] Benjamin Mora, Ross Maciejewski, Min Chen, and David S. Ebert. Visualization and computer graphics on isotropically emissive volumetric displays. IEEE Transactions on Visualization and Computer Graphics, 15(2): 221–234, March/April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meyer:2010:CAM

[MMDP10] Miriah Meyer, Tamara Munzner, Angela DePace, and Hanspeter Pfister. Comparative analysis of multidimensional, quantitative data.

IEEE Transactions on Visualization and Computer Graphics, 16(6):908–917, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maries:2013:GVC

[MMH^{+}13] Adrian Maries, Nathan Mays, MeganOlson Hunt, Kim F. Wong, William Layton, Robert Boudreau, Caterina Rosano, and G. Elisabeta Marai. GRACE: A visual comparison framework for integrated spatial and non-spatial geriatric data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2916– 2925, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miao:2017:PMU

[MMK^{+}17] Haichao Miao, Gabriel Mistelbauer, Alexey Karimov, Amir Alansary, Alice Davidson, David F. A. Lloyd, Mellisa Damodaram, Lisa Story, Jana Hutter, Joseph V. Hajnal, Mary Rutherford, Bernhard Preim, Bernhard Kainz, and M. Eduard Groller. Placenta maps: In utero placental health assessment of the human fetus. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1612–1623, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07864470-
abs.html.

Moller:1997:EDF

[MMMY97] T. Möller, R. Machiraju, K. Mueller, and Roni Yagel. Evaluation and design of filters using a Taylor series expansion. IEEE Transactions on Visualization and Computer Graphics, 3(2):184– 199, April/June 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0184.
pdf;   http://www.computer.
org/tvcg/tg1997/v0184abs.
htm.

Meyer:2009:MMS

[MMP09] Miriah Meyer, Tamara Munzner, and Hanspeter Pfister. MizBee: a multiscale synteny browser. IEEE Transactions on Visualization and Computer Graphics, 15(6):897–904, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:1998:SEA

[MMS^{+}98] K. Mueller, T. Möller, J. E.

Swan II, R. Crawfis, N. Shareef, and R. Yagel. Splatting errors and antialiasing. IEEE Transactions on Visualization and Computer Graphics, 4 (2):178–191, April/June 1998.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0178.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0178abs.
                                        htm.

Malik:2014:PSR

[MMT^{+}14] Abish Malik, Ross Maciejewski, Sherry Towers, Sean Mc-Cullough, and David S. Ebert. Proactive spatiotemporal resource allocation and predictive visual analytics for community policing and law enforcement. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1863–1872, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875970-
                                        abs.html.

Melek:2006:VFT

[MMYK06] Zeki Melek, David Mayerich, Cem Yuksel, and John Keyser. Visualization of fibrous and thread-like data. IEEE Transactions on Visualization and Computer Graphics, 12(5):1165–1172, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

McGraw:2007:SDM

[MN07] Tim McGraw and Mariappan Nadar. Stochastic DT-MRI connectivity mapping on the GPU. IEEE Transactions on Visualization and Computer Graphics, 13(6):1504– 1511, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mourning:2014:IMV

[MNC14] Chad Mourning, Scott Nykl, and David Chelberg. Interactive mesostructures with volumetric collisions. IEEE Transactions on Visualization and Computer Graphics, 20 (7):970–982, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muraki:2001:ACM

[MNKT01] Shigeru Muraki, Toshiharu Nakai, Yasuyo Kita, and Koji Tsuda. An attempt for coloring multichannel MR imaging data. IEEE Transactions on Visualization and Computer Graphics, 7(3): 265–274, July 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0265.
pdf;   http://www.computer.
org/tvcg/tg2001/v0265abs.
htm.

Meyer:2007:PSE

[MNKW07] Miriah Meyer, Blake Nelson, Robert Kirby, and Ross Whitaker. Particle systems for efficient and accurate high-order finite element visualization. IEEE Transactions on Visualization and Computer Graphics, 13(5):1015–1026, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meyerhenke:2018:DLG

[MNS18] Henning Meyerhenke, Martin Nollenburg, and Christian Schulz. Drawing large graphs by multilevel maxentstress optimization. IEEE Transactions on Visualization and Computer Graphics, 24(5):1814–1827, May 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07889042-
abs.html.

Magnenat:2015:LTA

[MNZ^{+}15] Stephane Magnenat, Dat Tien Ngo, Fabio Zund, Mattia Ryffel, Gioacchino Noris, Gerhard Rothlin, Alessia Marra, Maurizio Nitti, Pascal Fua, Markus Gross, and Robert W. Sumner. Live texturing of augmented reality characters from colored drawings. IEEE Transactions on Visualization and Computer Graphics, 21(11):1201– 1210, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165658-
abs.html.

Myers:2014:LSO

[MOC^{+}14] Sean Myers, Daniel Oliphant, Brian Cherinka, Timothy Basil Luciani, W. Michael Wood-Vasey, Alexandros Labrinidis, and G. Elisabeta Marai. Large-scale overlays and trends: Visually mining, panning and zooming the observable universe. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1048–1061, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Machado:2009:PBM

[MOF09] Gustavo M. Machado, Manuel M.

Oliveira, and Leandro A. F. Fernandes. A physiologicallybased model for simulation of color vision deficiency. IEEE Transactions on Visualization and Computer Graphics, 15(6):1291–1298, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See corrections [MOF10].

Machado:2010:CPB

[MOF10] Gustavo M. Machado, Manuel M.

Oliveira, and Leandro A. F. Fernandes. Corrections to “A Physiologically-Based Model for Simulation of Color Vision Deficiency”. IEEE Transactions on Visualization and Computer Graphics, 16(2): 352, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See [MOF09].

Merigot:2011:VBC

[MOG11] Quentin Merigot, Maks Ovsjanikov, and Leonidas Guibas. Voronoi-based curvature and feature estimation from point clouds. IEEE Transactions on Visualization and Computer Graphics, 17(6): 743–756, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meuschke:2019:CBF

[MOJB^{+}19] M. Meuschke, S. Oeltze-Jafra, O. Beuing, B. Preim, and K. Lawonn. Classification of blood flow patterns in cerebral aneurysms. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2404–2418, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moorhead:2003:GEI

[Moo03] Robert J. Moorhead II. Guest Editor’s introduction: Special section on visualization. IEEE Transactions on Visualization and Computer Graphics, 9 (2):113–114, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0113.
pdf;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0113.htm.

Moreland:2013:SVP

[Mor13] Kenneth Moreland. A survey of visualization pipelines. IEEE Transactions on Visualization and Computer Graphics, 19(3):367–378, March 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muhlbacher:2013:PBF

[MP13] Thomas Mühlbacher and Harald Piringer. A partitionbased framework for building and validating regression models. IEEE Transactions on Visualization and Computer Graphics, 19(12):1962– 1971, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Munoz-Pandiella:2018:UWI

[MPBM^{+}18] Imanol Muñoz-Pandiella, Carles Bosch, Nicolas Mérillou, Gustavo Patow, Stéphane Mérillou, and Xavier Pueyo. Urban weathering: Interactive rendering of polluted cities. IEEE Transactions on Visualization and Computer Graphics, 24(12):3239– 3252, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8260954/.

Muhlbacher:2014:OBB

[MPG^{+}14] Thomas Mühlbacher, Harald Piringer, Samuel Gratzl, Michael Sedlmair, and Marc Streit. Opening the black box: Strategies for increased user involvement in existing algorithm implementations. IEEE Transactions on Visualization and Computer Graphics, 20(12):1643– 1652, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875995-
abs.html.

Maciejewski:2013:ABC

[MPK^{+}13] R. Maciejewski, A. Pattath, Sungahn Ko, R. Hafen, W. S. Cleveland, and D. S. Ebert. Automated Box–Cox transformations for improved visual encoding. IEEE Transactions on Visualization and Computer Graphics, 19(1): 130–140, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Micallef:2017:TPO

[MPOW17] Luana Micallef, Gregorio Palmas, Antti Oulasvirta, and Tino Weinkauf. Towards perceptual optimization of the visual design of scatterplots. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1588–1599, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07864468-
abs.html.

Martin:2003:FFC

[MPT03] Ignacio Martín, Xavier Pueyo, and Dani Tost. Frameto-frame coherent animation with two-pass radiosity. IEEE Transactions on Visualization and Computer Graphics, 9(1): 70–84, January/March 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0070abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/01/
v0070.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/01/v0070.pdf.

Marriott:2012:MVF

[MPWG12] Kim Marriott, Helen Purchase, Michael Wybrow, and Cagatay Goncu. Memorability of visual features in network diagrams. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2477–2485, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ming:2019:RVU

[MQB19] Yao Ming, Huamin Qu, and Enrico Bertini. RuleMatrix: Visualizing and understanding classifiers with rules. IEEE Transactions on Visualization and Computer Graphics, 25(1):342– 352, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440085-
abs.html.

Miller:2006:IVI

[MQF06] Jameson Miller, Cory Quammen, and Matthew Fleenor. Interactive visualization of intercluster galaxy structures in the horologiumreticulum supercluster. IEEE Transactions on Visualization and Computer Graphics, 12(5):1149–1156, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mandal:2000:DMB

[MQV00] C. Mandal, H. Qin, and B. C. Vemuri. Dynamic modeling of butterfly subdivision surfaces. IEEE Transactions on Visualization and Computer Graphics, 6(3):265–287, July/September 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0265.
pdf;   http://www.computer.
org/tvcg/tg2000/v0265abs.
htm.

Mehra:2015:WIW

[MRG^{+}15] R. Mehra, A. Rungta, A. Golas, Ming Lin, and D. Manocha. WAVE: Interactive wavebased sound propagation for virtual environments. IEEE Transactions on Visualization and Computer Graphics, 21 (4):434–442, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maciejewski:2010:VAA

[MRH^{+}10] Ross Maciejewski, Stephen Rudolph, Ryan Hafen, Ahmad M. Abusalah, Mohamed Yakout, Mourad Ouzzani, William S. Cleveland, Shaun J. Grannis, and David S. Ebert. A visual analytics approach to understanding spatiotemporal hotspots. IEEE Transactions on Visualization and Computer Graphics, 16(2):205– 220, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

MacEachren:2012:VSU

[MRO^{+}12] Alan M. MacEachren, Robert E.

Roth, James O’Brien, Bonan Li, Derek Swingley, and Mark Gahegan. Visual semiotics & uncertainty visualization: an empirical study. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2496–2505, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meulemans:2013:KHS

[MRS^{+}13] Wouter Meulemans, Nathalie Henry Riche, Bettina Speckmann, Basak Alper, and Tim Dwyer. KelpFusion: A hybrid set visualization technique. IEEE Transactions on Visualization and Computer Graphics, 19(11):1846–1858, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maguire:2012:TBG

[MRSS^{+}12] Eamonn Maguire, Philippe Rocca-Serra, Susanna-Assunta Sansone, Jim Davies, and Min Chen. Taxonomy-based glyph design — with a case study on visualizing workflows of biological experiments. IEEE Transactions on Visualization and Computer Graphics, 18(12):2603–2612, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maguire:2013:VCW

[MRSS^{+}13] Eamonn Maguire, Philippe Rocca-Serra, Susanna-Assunta Sansone, Jim Davies, and Min Chen. Visual compression of workflow visualizations with automated detection of macro motifs. IEEE Transactions on Visualization and Computer Graphics, 19(12):2576– 2585, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Myszkowski:2000:PBF

[MRT00] K. Myszkowski, P. Rokita, and T. Tawara. Perceptionbased fast rendering and antialiasing of walkthrough sequences. IEEE Transactions on Visualization and Computer Graphics, 6(4):360–??, October/December 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0360.
pdf;   http://www.computer.
org/tvcg/tg2000/v0360abs.
htm.

Majumder:2004:CNP

[MS04] Aditi Majumder and Rick Stevens. Color nonuniformity in projection-based displays: Analysis and solutions. IEEE Transactions on Visualization and Computer Graphics, 10(2):177– 188, March/April 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0177abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0177.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/02/v0177.pdf.

Ma:2008:PSE

[MS08] YingLiang Ma and Kurt Saetzler. A parallelized surface extraction algorithm for large binary image data sets based on an adaptive 3-D Delaunay subdivision strategy. IEEE Transactions on Visualization and Computer Graphics, 14(1):160–172, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

MacQuarrie:2018:ETT

[MS18a] Andrew MacQuarrie and Anthony Steed. The effect of transition type in multi-view 360^{∘} media. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1564–1573, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/04/08260946.
                                        pdf.

Medeiros:2018:GRM

[MS18b] Esdras Medeiros and Marcelo Siqueira. Good random multi-triangulation of surfaces. IEEE Transactions on Visualization and Computer Graphics, 24(6):1983– 1996, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/06/07927461-
                                        abs.html.

Matsui:2017:DDA

[MSA17] Yusuke Matsui, Takaaki Shiratori, and Kiyoharu Aizawa.

DrawFromDrawings: 2D drawing assistance via stroke interpolation with a sketch database. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1852–1862, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07452668.
pdf.

Merhof:2006:HVW

[MSE^{+}06] Dorit Merhof, Markus Sonntag, Frank Enders, Christopher Nimsky, Peter Hastreiter, and Guenther Greiner. Hybrid visualization for white matter tracts using triangle strips and point sprites. IEEE Transactions on Visualization and Computer Graphics, 12(5):1181–1188, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:1999:HQS

[MSHC99] K. Mueller, N. Shareef, J. Huang, and R. Crawfis. High-quality splatting on rectilinear grids with efficient culling of occluded voxels. IEEE Transactions on Visualization and Computer Graphics, 5(2):116– 134, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0116.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0116abs.
                                        htm.

Mori:2011:WVP

[MSM^{+}11] H. Mori, E. Sumiya, T. Mashita, K. Kiyokawa, and H. Takemura. A wide-view parallaxfree eye-mark recorder with a hyperboloidal half-silvered mirror and appearance-based gaze estimation. IEEE Transactions on Visualization and Computer Graphics, 17(7): 900–912, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Machado:2014:EM

[MSME14] Gustavo Machado, Filip Sadlo, Thomas Muller, and Thomas Ertl. Escape maps. IEEE Transactions on Visualization and Computer Graphics, 20(12):2604–2613, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875962-
                                        abs.html.

Meyer-Spradow:2008:GBS

[MSSD^{+}08] Jennis Meyer-Spradow, Lars Stegger, Christian Döring, Timo Ropinski, and Klaus Hinrichs. Glyph-based SPECT visualization for the diagnosis of coronary artery disease.

IEEE Transactions on Visualization and Computer Graphics, 14(6):1499–1506, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Majerowicz:2014:FYS

[MSSH14] Lucas Majerowicz, Ariel Shamir, Alla Sheffer, and Holger H. Hoos. Filling your shelves: Synthesizing diverse style-preserving artifact arrangements. IEEE Transactions on Visualization and Computer Graphics, 20(11): 1507–1518, November 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06636298-
abs.html.

Marriott:2011:HTT

[MSvG^{+}11] Kim Marriott, Peter Sbarski, Tim van Gelder, Daniel Prager, and Andy Bulka. Hitrees and their layout. IEEE Transactions on Visualization and Computer Graphics, 17 (3):290–304, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meier:2008:VPI

[MSW^{+}08] Holger A. Meier, Michael Schlemmer, Christian Wagner, Andreas Kerren, Hans Hagen, Ellen Kuhl, and Paul Steinmann. Visualization of particle interactions in granular media. IEEE Transactions on Visualization and Computer Graphics, 14(5):1110–1125, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maciejewski:2019:GEI

[MSW19] Ross Maciejewski, Jinwook Seo, and Rüdiger Westermann. Guest Editors’ introduction: Special section on IEEE PacificVis 2019. IEEE Transactions on Visualization and Computer Graphics, 25(6):2155–2156, June 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8703194/.

Musse:2001:HMR

[MT01] S. R. Musse and D. Thalmann. Hierarchical model for real time simulation of virtual human crowds. IEEE Transactions on Visualization and Computer Graphics, 7 (2):152–??, April/June 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0152.
pdf;   http://www.computer.
org/tvcg/tg2001/v0152abs.
htm.

Moller:2005:CIA

[MT05] Christian N. Moller and Adrian R. L. Travis. Correcting interperspective aliasing in autostereoscopic displays. IEEE Transactions on Visualization and Computer Graphics, 11(2):228– 236, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mahyar:2014:SCC

[MT14] Narges Mahyar and Melanie Tory. Supporting communication and coordination in collaborative sensemaking. IEEE Transactions on Visualization and Computer Graphics, 20(12):1633– 1642, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875986-
abs.html.

Morgand:2017:MVG

[MTB17] Alexandre Morgand, Mohamed Tamaazousti, and Adrien Bartoli. A multipleview geometric model of specularities on non-planar shapes with application to dynamic retexturing. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2485–2493, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007318-
abs.html.

Morgand:2018:GMS

[MTB18] Alexandre Morgand, Mohamed Tamaazousti, and Adrien Bartoli. A geometric model for specularity prediction on planar surfaces with multiple light sources. IEEE Transactions on Visualization and Computer Graphics, 24(5):1691–1704, May 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/05/07869421-
                                        abs.html.

Matute:2018:SBS

[MTL18] J. Matute, A. C. Telea, and L. Linsen. Skeleton-based scagnostics. IEEE Transactions on Visualization and Computer Graphics, 24(1): 542–552, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Madsen:2016:TCS

[MTM^{+}16] J. B. Madsen, M. Tatzqern, C. B. Madsen, D. Schmalstieg, and D. Kalkofen. Temporal coherence strategies for augmented reality labeling. IEEE Transactions on Visualization and Computer Graphics, 22(4):1415–1423, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muhler:2010:MET

[MTRP10] Konrad Muhler, Christian Tietjen, Felix Ritter, and Bernhard Preim. The Medical Exploration Toolkit: an efficient support for visual computing in surgical planning and training. IEEE Transactions on Visualization and Computer Graphics, 16(1):133–146, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moustakas:2007:SME

[MTS07] Konstantinos Moustakas, Dimitrios Tzovaras, and Michael Gerassimos Strintzis. SQ-Map: Efficient layered collision detection and haptic rendering. IEEE Transactions on Visualization and Computer Graphics, 13(1):80–93, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moere:2012:EES

[MTW^{+}12] Andrew Vande Moere, Martin Tomitsch, Christoph Wimmer, Boesch Christoph, and Thomas Grechenig. Evaluating the effect of style in information visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2739–2748, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:2019:MNE

[Mue19]  Klaus Mueller.  A mes-

sage from the new Editorin-Chief. IEEE Transactions on Visualization and Computer Graphics, 25(2):1267– 1268, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/02/08594692.
                                        pdf.

Munzner:2009:NPM

[Mun09] Tamara Munzner. A nested process model for visualization design and validation. IEEE Transactions on Visualization and Computer Graphics, 15(6):921–928, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muraki:1995:MVR

[Mur95] Shigeru Muraki. Multiscale volume representation by a DoG wavelet. IEEE Transactions on Visualization and Computer Graphics, 1(2): 109–116, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0109.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0109abs.
                                        htm.

Marchand:2016:PEA

[MUS16] Eric Marchand, Hideaki Uchiyama, and Fabien Spindler.

Pose estimation for augmented reality: A hands-on survey. IEEE Transactions on Visualization and Computer Graphics, 22(12):2633– 2651, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07368948-
abs.html.

Mansmann:2006:IED

[MV06] Florian Mansmann and Svetlana Vinnik. Interactive exploration of data traffic with hierarchical network maps. IEEE Transactions on Visualization and Computer Graphics, 12(6):1440–1449, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meuschke:2017:CVV

[MVB^{+}17] Monique Meuschke, Samuel Voss, Oliver Beuing, Bernhard Preim, and Kai Lawonn. Combined visualization of vessel deformation and hemodynamics in cerebral aneurysms. IEEE Transactions on Visualization and Computer Graphics, 23(1): 761–770, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muthumanickam:2019:ITV

[MVN^{+}19] Prithiviraj K. Muthumanickam, Katerina Vrotsou, Aida Nordman, Jimmy Johansson, and Matthew Cooper. Identification of temporally varying areas of interest in long-duration eye-tracking data sets. IEEE Transactions on Visualization and Computer Graphics, 25(1): 87–97, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08454276-
abs.html.

Mangan:1999:PSM

[MW99] A. P. Mangan and R. T. Whitaker. Partitioning 3D surface meshes using watershed segmentation. IEEE Transactions on Visualization and Computer Graphics, 5(4):308–321, October/ December 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0308.
pdf;   http://www.computer.
org/tvcg/tg1999/v0308abs.
htm.

Miksch:2013:GEI

[MW13] Silvia Miksch and Matthew Ward. Guest Editors’ introduction: Special section on the IEEE Conference on Visual Analytics Science and Technology (VAST). IEEE Transactions on Visualization and Computer Graphics, 19(7):1076–1077, July 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Miller:2012:IMA

[MWC^{+}12] Andrew Miller, Brandyn White, Emiko Charbonneau, Zach Kanzler, and Joseph J. LaViola, Jr. Interactive 3D model acquisition and tracking of building block structures. IEEE Transactions on Visualization and Computer Graphics, 18(4): 651–659, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Maciejewski:2009:SFS

[MWCE09] Ross Maciejewski, Insoo Woo, Wei Chen, and David Ebert. Structuring feature space: a non-parametric method for volumetric transfer function generation. IEEE Transactions on Visualization and Computer Graphics, 15(6):1473–1480, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mania:2006:EVI

[MWCR06] Katerina Mania, Dave Wooldridge, Matthew Coxon, and Andrew Robinson. The effect of visual and interaction fidelity on spatial cognition in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 12(3):396–404, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meyer:2008:PBS

[MWK^{+}08] Miriah Meyer, Ross Whitaker, Robert M. Kirby, Christian Ledergerber, and Hanspeter Pfister. Particle-based sampling and meshing of surfaces in multimaterial volumes. IEEE Transactions on Visualization and Computer Graphics, 14(6):1539– 1546, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mirzargar:2014:CBG

[MWK14] Mahsa Mirzargar, Ross T. Whitaker, and Robert M. Kirby. Curve boxplot: Generalization of boxplot for ensembles of curves. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2654–2663, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875964-
                                        abs.html.

Moritz:2019:FVD

[MWN^{+}19] Dominik Moritz, Chenglong Wang, Greg L. Nelson, Halden Lin, Adam M. Smith, Bill Howe, and Jeffrey Heer. Formalizing visualization design knowledge as constraints: Actionable and extensible models in Draco. IEEE Transactions on Visualization and Computer Graphics, 25(1): 438–448, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440847-
abs.html.

Ma:2014:GBI

[MWSJ14] Jun Ma, Chaoli Wang, Ching-Kuang Shene, and Jingfeng Jiang. A graph-based interface for visual analytics of 3D streamlines and pathlines. IEEE Transactions on Visualization and Computer Graphics, 20(8):1127–1140, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2013:CBI

[MXW^{+}13] Li-Qian Ma, Kun Xu, Tien-Tsin Wong, Bi-Ye Jiang, and Shi-Min Hu. Change blindness images. IEEE Transactions on Visualization and Computer Graphics, 19(11): 1808–1819, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Machiraju:1996:REC

[MY96] Raghu Machiraju and Roni Yagel. Reconstruction error characterization and control: a sampling theory approach. IEEE Transactions on Visualization and Computer Graphics, 2(4):364– 378, December 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0364.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0364abs.
                                        htm.

Meenakshisundaram:2014:GEI

[MY14] Gopi Meenakshisundaram and Sung-Eui Yoon. Guest editor’ introduction: Special issue on the ACM SIGGRAPH Symposium on Interactive 3D Graphics and Games 2013. IEEE Transactions on Visualization and Computer Graphics, 20(7):955–956, July 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Matsukura:2013:SSD

[MYI13] Haruka Matsukura, Tatsuhiro Yoneda, and Hiroshi Ishida. Smelling screen: Development and evaluation of an olfactory display system for presenting a virtual odor source. IEEE Transactions on Visualization and Computer Graphics, 19 (4):606–615, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Museth:2008:GEI

[MYM08] Ken Museth, Anders Ynnerman, and Torsten Möller.

Guest Editor’s introduction: Special section on Euro-Vis. IEEE Transactions on Visualization and Computer Graphics, 14(4):725– 726, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040725.
pdf.

Mo:2016:TAR

[MYM16] Qi Mo, Hengchin Yeh, and Dinesh Manocha. Tracing analytic ray curves for light and sound propagation in nonlinear media. IEEE Transactions on Visualization and Computer Graphics, 22(11): 2493–2506, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Muelder:2016:VAC

[MZC^{+}16] Chris Muelder, Biao Zhu, Wei Chen, Hongxin Zhang, and Kwan-Liu Ma. Visual analysis of cloud computing performance using behavioral lines. IEEE Transactions on Visualization and Computer Graphics, 22(6):1694– 1704, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Machiraju:1998:SSR

[MZFM98] R. Machiraju, Z. Zhu, B. Fry, and R. Moorhead. Structuresignificant representation of structured datasets. IEEE Transactions on Visualization and Computer Graphics, 4 (2):117–132, April/June 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0117.
pdf;   http://www.computer.
org/tvcg/tg1998/v0117abs.
htm.

Metoyer:2008:PIA

[MZH^{+}08] Ronald Metoyer, Victor Zordan, Benjamin Hermens, Chun-Chi Wu, and Marc Soriano. Psychologically inspired anticipation and dynamic response for impacts to the head and upper body. IEEE Transactions on Visualization and Computer Graphics, 14(1): 173–185, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Min:2019:VHF

[MZS^{+}19] X. Min, W. Zhang, S. Sun, N. Zhao, S. Tang, and Y. Zhuang. VPModel: Highfidelity product simulation in a virtual–physical environment. IEEE Transactions on Visualization and Computer Graphics, 25(11):3083– 3093, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2015:FFM

[MZX15] Yuewen Ma, Jianmin Zheng, and Jian Xie. Foldover-free mesh warping for constrained texture mapping. IEEE Transactions on Visualization and Computer Graphics, 21 (3):375–388, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06942245-
abs.html.

Nonato:2019:MPV

[NA19] L. G. Nonato and M. Aupetit. Multidimensional projection for visual analytics: Linking techniques with distortions, tasks, and layout enrichment. IEEE Transactions on Visualization and Computer Graphics, 25(8):2650– 2673, August 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nusrat:2018:ECE

[NAK18] S. Nusrat, M. J. Alam, and S. Kobourov. Evaluating cartogram effectiveness. IEEE Transactions on Visualization and Computer Graphics, 24(2):1077–1090, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nusrat:2018:CVB

[NASK18] Sabrina Nusrat, Muhammad Jawaherul Alam, Carlos Scheidegger, and Stephen Kobourov. Cartogram visualization for bivariate geostatistical data. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2675–2688, October 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08078198-
abs.html.

Najork:1995:OHL

[NB95] Marc A. Najork and Marc H. Brown. Obliq-3D: a highlevel, fast-turnaround 3D animation system. IEEE Transactions on Visualization and Computer Graphics, 1(2): 175–193, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0175.
pdf;   http://www.computer.
org/tvcg/tg1995/v0175abs.
htm.

Nocaj:2012:OSR

[NB12] Arlind Nocaj and Ulrik Brandes. Organizing search results with a reference map. IEEE Transactions on Visualization and Computer Graphics, 18(12):2546–2555, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Narang:2019:IUI

[NBM19] Sahil Narang, Andrew Best, and Dinesh Manocha. Inferring user intent using Bayesian theory of mind in shared avatar-agent virtual environments. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2113–2122, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642370/.

Netzel:2014:CET

[NBW14] Rudolf Netzel, Michel Burch, and Daniel Weiskopf. Comparative eye tracking study on node-link visualizations of trajectories. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2221–2230, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875968-
abs.html.

Niski:2007:TBL

[NC07] Krzysztof Niski and Jonathan D.

Cohen. Tile-based level of detail for the parallel age. IEEE Transactions on Visualization and Computer Graphics, 13(6):1352–1359, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1352s.
mov.

Nagata:1997:MVP

[NDM^{+}97] N. Nagata, T. Dobashi, Y. Manabe, T. Usami, and S. Inokuchi. Modeling and visualization for a pearlquality evaluation simulator. IEEE Transactions on Visualization and Computer Graphics, 3(4):307–315, October/December 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0307.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0307abs.
                                        htm.

Nimeroff:1996:IAI

[NDR96] Jeffry Nimeroff, Julie Dorsey, and Holly Rushmeier. Implementation and analysis of an image-based global illumination framework for animated environments. IEEE Transactions on Visualization and Computer Graphics, 2 (4):283–298, December 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0283.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0283abs.
                                        htm.

Novotny:2010:EVR

[NDS10] Pavol Novotny, Leonid I. Dimitrov, and Milos Sramek. Enhanced voxelization and representation of objects with sharp details in truncated distance fields. IEEE Transactions on Visualization and Computer Graphics, 16(3): 484–498, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Natarajan:2004:STD

[NE04] Vijay Natarajan and Herbert Edelsbrunner. Simplification of three-dimensional density maps. IEEE Transactions on Visualization and Computer Graphics, 10(5): 587–597, September/October 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/dl/trans/tg/2004/05/
v0587.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/05/v0587.pdf.

Neyret:1998:MAR

[Ney98] F. Neyret. Modeling, animating, and rendering complex scenes using volumetric textures. IEEE Transactions on Visualization and Computer Graphics, 4(1):55–70, January/March 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0055.
pdf;   http://www.computer.
org/tvcg/tg1998/v0055abs.
htm.

Nobre:2019:LVM

[NGCL19] C. Nobre, N. Gehlenborg, H. Coon, and A. Lex. Lineage: Visualizing multivariate clinical data in genealogy graphs. IEEE Transactions on Visualization and Computer Graphics, 25(3):1543–1558, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nadeem:2018:LSP

[NGK18] Saad Nadeem, Xianfeng Gu, and Arie E. Kaufman. LMap: Shape-preserving local mappings for biomedical visualization. IEEE Transactions on Visualization and Computer Graphics, 24(12):3111– 3122, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8106712/.

Novotny:2006:OPF

[NH06] Matej Novotny and Helwig Hauser. Outlier-preserving Focus + Context visualization in parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 12(5):893–900, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Netzel:2017:EVS

[NHB^{+}17] Rudolf Netzel, Marcel Hlawatsch, Michael Burch, Sanjeev Balakrishnan, Hansjörg Schmauder, and Daniel Weiskopf. An evaluation of visual search support in maps. IEEE Transactions on Visualization and Computer Graphics, 23(1): 421–430, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2017:PGV

[NHEM17] Quan Hoang Nguyen, Seok-Hee Hong, Peter Eades, and Amyra Meidiana. Proxy graph: Visual quality metrics of big graph sampling. IEEE Transactions on Visualization and Computer Graphics, 23(6):1600–1611, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07864456-
abs.html.

Nirnimesh:2007:GST

[NHN07] Nirnimesh, Pawan Harish, and P. J. Narayanan. Garuda: a scalable tiled display wall using commodity PCs. IEEE Transactions on Visualization and Computer Graphics, 13(5):864–877, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nowrouzezahrai:2014:FNL

[NHPN14] Derek Nowrouzezahrai, Eric Heitz, Pierre Poulin, and Fabrice Neyret. Filtering nonlinear transfer functions on surfaces. IEEE Transactions on Visualization and Computer Graphics, 20(7): 996–1008, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2018:RIT

[NHYY18] Duc Thanh Nguyen, Binh-Son Hua, Lap-Fai Yu, and Sai-Kit Yeung. A robust 3D–2D interactive tool for scene segmentation and annotation. IEEE Transactions on Visualization and Computer Graphics, 24(12):3005– 3018, December 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8115231/.

Nielson:1995:TII

[Nie95] Gregory M. Nielson. Theme issue introduction: Visualization takes its place in the scientific community. IEEE Transactions on Visualization and Computer Graphics, 1 (2):97–??, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0097.
                                        pdf.

Nielson:1996:TII

[Nie96] Gregory M. Nielson. Theme issue introduction: Challenges in visualization research. IEEE Transactions on Visualization and Computer Graphics, 2(2): 97–??, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0097.
pdf.

Nielson:2003:MC

[Nie03] Gregory M. Nielson. On marching cubes. IEEE Transactions on Visualization and Computer Graphics, 9(3):283– 297, July/September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0283abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0283.pdf.

Nielson:2004:QSS

[Nie04]  Gregory M. Nielson.  ν-

quaternion splines for the smooth interpolation of orientations. IEEE Transactions on Visualization and Computer Graphics, 10(2): 224–229, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0224abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/

                                        v0224.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2004/02/v0224.pdf.

Nielson:1999:TCT

[NJ99] G. M. Nielson and I.-H. Jung. Tools for computing tangent curves for linearly varying vector fields over tetrahedral domains. IEEE Transactions on Visualization and Computer Graphics, 5(4):360–372, October/December 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0360.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0360abs.
                                        htm.

Noguera:2016:MVR

[NJ16] J. M. Noguera and J. R. Jimenez. Mobile volume rendering: Past, present and future. IEEE Transactions on Visualization and Computer Graphics, 22(2):1164– 1178, February 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Navratil:2007:VCP

[NJB07] Paul Navratil, Jarrett Johnson, and Volker Bromm. Visualization of cosmological particle-based datasets. IEEE Transactions on Visualization and Computer Graphics, 13(6):1712–1718, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nielsen:2009:AEV

[NJBJ09] Cydney B. Nielsen, Shaun D.

Jackman, Inanc Birol, and Steven J. M. Jones. ABySS-Explorer: Visualizing genome sequence assemblies. IEEE Transactions on Visualization and Computer Graphics, 15(6):881–888, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nilsson:2011:COC

[NJJ11] Susanna Nilsson, Bjorn J. E.

Johansson, and Arne Jonsson. Cross-organizational collaboration supported by augmented reality. IEEE Transactions on Visualization and Computer Graphics, 17(10):1380–1392, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nelson:2006:RTP

[NK06] Blake Nelson and Robert M.

Kirby. Ray-tracing polymorphic multidomain spectral/hp elements for isosurface rendering. IEEE Transactions on Visualization and Computer Graphics, 12(1):114–125, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nelson:2011:GBI

[NKH11] Blake Nelson, Robert M. Kirby, and Robert Haimes. GPU-based interactive cutsurface extraction from highorder finite element fields. IEEE Transactions on Visualization and Computer Graphics, 17(12):1803–1811, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nelson:2014:GBV

[NKH14] Blake Nelson, Robert M. Kirby, and Robert Haimes. GPU-based volume visualization from high-order finite element fields. IEEE Transactions on Visualization and Computer Graphics, 20(1): 70–83, January 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

New:2008:DVC

[NKHC08] Joshua New, Wesley Kendall, Jian Huang, and Elissa Chesler. Dynamic visualization of coexpression in systems genetics data. IEEE Transactions on Visualization and Computer Graphics, 14(5):1081–1095, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/05/ttg2008051081s.
zip.

Nah:2015:HHA

[NKP^{+}15] Jae-Ho Nah, Jin-Woo Kim, Junho Park, Won-Jong Lee, Jeong-Soo Park, Seok-Yoon Jung, Woo-Chan Park, Dinesh Manocha, and Tack-Don Han. HART: A hybrid architecture for ray tracing animated scenes. IEEE Transactions on Visualization and Computer Graphics, 21(3): 389–401, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06960897-
abs.html.

Nelson:2012:ESA

[NLKH12] Blake Nelson, Eric Liu, Robert M. Kirby, and Robert Haimes. ElVis: a system for the accurate and interactive visualization of high-order finite element solutions. IEEE Transactions on Visualization and Computer Graphics, 18(12):2325–2334, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nouanesengsy:2011:LBP

[NLS11] Boonthanome Nouanesengsy, Teng-Yok Lee, and Han-Wei Shen. Load-balanced parallel streamline generation on large scale vector fields. IEEE Transactions on Visualization and Computer Graphics, 17(12):1785–1794, December 2011. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nam:2013:TTI

[NM13] J. E. Nam and K. Mueller. TripAdvisor^{N−D}: A tourisminspired high-dimensional space exploration framework with overview and detail. IEEE Transactions on Visualization and Computer Graphics, 19(2):291–305, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nadeem:2017:CSP

[NMGK17] Saad Nadeem, Joseph Marino, Xianfeng Gu, and Arie Kaufman. Corresponding supine and prone colon visualization using eigenfunction analysis and fold modeling. IEEE Transactions on Visualization and Computer Graphics, 23 (1):751–760, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nagao:2018:ADV

[NMN^{+}18] Ryohei Nagao, Keigo Matsumoto, Takuji Narumi, Tomohiro Tanikawa, and Michitaka Hirose. Ascending and descending in virtual reality: Simple and safe system using passive haptics. IEEE Transactions on Visualization and Computer Graphics, 24(4):1584–1593, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260962-
abs.html.

Nagaraj:2011:ERA

[NN11a] Suthambhara Nagaraj and Vijay Natarajan. Errata to “Relation-Aware Isosurface Extraction in Multifield Data”. IEEE Transactions on Visualization and Computer Graphics, 17(5):709–710, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See [NN11b].

Nagaraj:2011:RAI

[NN11b] Suthambhara Nagaraj and Vijay Natarajan. Relationaware isosurface extraction in multifield data. IEEE Transactions on Visualization and Computer Graphics, 17(2): 182–191, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See errata [NN11a].

Nakamaru:1997:BFR

[NO97] K. Nakamaru and Y. Ohno.

Breadth-first ray tracing utilizing uniform spatial subdivision. IEEE Transactions on Visualization and Computer Graphics, 3(4):316–328, October/December 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/

                                        tg/books/tg1997/pdf/v0316.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0316abs.
                                        htm.

Nocaj:2016:ADB

[NOB16] Arlind Nocaj, Mark Ortmann, and Ulrik Brandes. Adaptive disentanglement based on local clustering in small-world network visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (6):1662–1671, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nieser:2012:HGP

[NPPZ12] Matthias Nieser, Jonathan Palacios, Konrad Polthier, and Eugene Zhang. Hexagonal global parameterization of arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 18 (6):865–878, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2005:RTH

[NQX^{+}05] Ta Huynh Duy Nguyen, Tran Cong Thien Qui, Ke Xu, Adrian David Cheok, Sze Lee Teo, ZhiYing Zhou, Asitha Mallawaarachchi, Shang Ping Lee, Wei Liu, Hui Siang Teo, Le Nam Thang, Yu Li, and Hirokazu Kato. Realtime 3D human capture system for mixed-reality art and entertainment. IEEE Transactions on Visualization and Computer Graphics, 11(6):706–721, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/06/v0706s.
pdf.

Nguyen:2018:DDS

[NR18] Hoa Nguyen and Paul Rosen.

DSPCP: A data scalable approach for identifying relationships in parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 24(3):1301– 1315, March 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2015:SMD

[NRS15] Thanh Nguyen, Gerhard Reitmayr, and Dieter Schmalstieg. Structural modeling from depth images. IEEE Transactions on Visualization and Computer Graphics, 21(11): 1230–1240, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165661-
abs.html.

Neth:2012:VDD

[NSE^{+}12] Christian T. Neth, Jan L.

Souman, David Engel, Uwe Kloos, Heinrich H. Bülthoff, and Betty J. Mohler. Velocitydependent dynamic curvature gain for redirected walking. IEEE Transactions on Visualization and Computer Graphics, 18(7):1041–1052, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Niederer:2018:TVC

[NSH^{+}18] C. Niederer, H. Stitz, R. Hourieh, F. Grassinger, W. Aigner, and M. Streit. TACO: Visualizing changes in tables over time. IEEE Transactions on Visualization and Computer Graphics, 24(1):677–686, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nobre:2019:JTA

[NSL19] Carolina Nobre, Marc Streit, and Alexander Lex. Juniper: A Tree+Table approach to multivariate graph visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 544–554, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08454344-
                                        abs.html.

Nilsson:2014:ERP

[NSN14] Niels Christian Nilsson, Stefania Serafin, and Rolf Nordahl. Establishing the range of perceptually natural visual walking speeds for virtual walkingin-place locomotion. IEEE Transactions on Visualization and Computer Graphics, 20 (4):569–578, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Noser:2003:DVR

[NSS03] Hansrudi Noser, Christian Stern, and Peter Stucki. Distributed virtual reality environments based on rewriting systems. IEEE Transactions on Visualization and Computer Graphics, 9(2):213– 225, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0213abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0213.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0213.pdf.

Narasimhan:2014:TRE

[NSS14] Srinivasa G. Narasimhan, Yulong Shi, and Yu Sheng. Translucent radiosity: Efficiently combining diffuse inter-reflection and subsurface scattering. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1009–1021, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

North:2011:GEI

[NSvW11] Stephen North, Han-Wei Shen, and Jarke J. van Wijk. Guest Editor’s introduction: Special section on the IEEE Pacific Visualization Symposium. IEEE Transactions on Visualization and Computer Graphics, 17(11):1545– 1546, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Neuroth:2017:SVT

[NSW^{+}17] Tyson Neuroth, Franz Sauer, Weixing Wang, Stephane Ethier, Choong-Seock Chang, and Kwan-Liu Ma. Scalable visualization of time-varying multi-parameter distributions using spatially organized histograms. IEEE Transactions on Visualization and Computer Graphics, 23(12):2599– 2612, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/12/07792155-
                                        abs.html.

Nadeem:2017:SPB

[NSZ^{+}17] Saad Nadeem, Zhengyu Su, Wei Zeng, Arie Kaufman, and Xianfeng Gu. Spherical parameterization balancing angle and area distortions. IEEE Transactions on Visualization and Computer Graphics, 23(6):1663–1676, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07434077-
abs.html.

Noser:1999:RBI

[NT99] H. Noser and D. Thalmann.

A rule-based interactive behavioral animation system for humanoids. IEEE Transactions on Visualization and Computer Graphics, 5(4):281– 307, October/December 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0281.
pdf;   http://www.computer.
org/tvcg/tg1999/v0281abs.
htm.

Nooruddin:2003:SRP

[NT03] Fakir S. Nooruddin and Greg Turk. Simplification and repair of polygonal models using volumetric techniques. IEEE Transactions on Visualization and Computer Graphics, 9 (2):191–205, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0191abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0191.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0191.pdf.

Nguyen:2019:UUB

[NTA^{+}19] P. H. Nguyen, C. Turkay, G. Andrienko, N. Andrienko, O. Thonnard, and J. Zouaoui. Understanding user behaviour through action sequences: From the usual to the unusual. IEEE Transactions on Visualization and Computer Graphics, 25(9):2838– 2852, September 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nordahl:2011:SSE

[NTS11] Rolf Nordahl, Luca Turchet, and Stefania Serafin. Sound synthesis and evaluation of interactive footsteps and environmental sounds rendering for virtual reality applications. IEEE Transactions on Visualization and Computer Graphics, 17(9):1234– 1244, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Novotny:2019:DVR

[NTT^{+}19] Johannes Novotny, Joshua Tveite, Morgan L. Turner, Stephen Gatesy, Fritz Drury, Peter Falkingham, and David H. Laidlaw. Developing virtual reality visualizations for unsteady flow analysis of dinosaur track formation using scientific sketching. IEEE Transactions on Visualization and Computer Graphics, 25(5):2145–2154, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8672601/.

Noordmans:2000:SVR

[NvdVS00] H. J. Noordmans, H. T. M.

van der Voort, and A. W. M. Smeulders. Spectral volume rendering. IEEE Transactions on Visualization and Computer Graphics, 6(3):196–207, July/September 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0196.
pdf;   http://www.computer.
org/tvcg/tg2000/v0196abs.
htm.

Nichols:2010:III

[NW10] Greg Nichols and Chris Wyman. Interactive indirect illumination using adaptive multiresolution splatting. IEEE Transactions on Visualization and Computer Graphics, 16(5):729–741, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nollenburg:2011:DLH

[NW11] Martin Nollenburg and Alexander Wolff. Drawing and labeling high-quality metro maps by mixed-integer programming. IEEE Transactions on Visualization and Computer Graphics, 17(5):626–641, May 2011. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nafari:2015:QTV

[NW15] M. Nafari and C. Weaver. Query2Question: Translating visualization interaction into natural language. IEEE Transactions on Visualization and Computer Graphics, 21 (6):756–769, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Neubauer:2005:AVE

[NWF^{+}05] Andre Neubauer, Stefan Wolfsberger, Marie-Therese Forster, Lukas Mroz, Rainer Wegenkittl, and Katja Buhler. Advanced virtual endoscopic pituitary surgery. IEEE Transactions on Visualization and Computer Graphics, 11(5):497–507, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nader:2016:JND

[NWHWD16] Georges Nader, Kai Wang, Franck Hétroy-Wheeler, and Florent Dupont. Just noticeable distortion profile for flat-shaded 3D mesh surfaces. IEEE Transactions on Visualization and Computer Graphics, 22(11):2423–2436, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Narita:2017:DPM

[NWI17] Gaku Narita, Yoshihiro Watanabe, and Masatoshi Ishikawa. Dynamic projection mapping onto deforming non-rigid surface using deformable dot cluster marker. IEEE Transactions on Visualization and Computer Graphics, 23(3): 1235–1248, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07516689-
abs.html.

Nie:2013:CVS

[NXSL13] Yongwei Nie, Chunxia Xiao, Hanqiu Sun, and Ping Li. Compact video synopsis via global spatiotemporal optimization. IEEE Transactions on Visualization and Computer Graphics, 19(10):1664– 1676, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2016:SUS

[NXW^{+}16] P. H. Nguyen, K. Xu, A. Wheat, B. W. Wong, S. Attfield, and B. Fields. SensePath: Understanding the sensemaking process through analytic provenance. IEEE Transactions on Visualization and Computer Graphics, 22(1):41–50, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2006:NVM

[NZ06] Tien Nguyen and Jun Zhang. A novel visualization model for Web search results. IEEE Transactions on Visualization and Computer Graphics, 12(5):981–988, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nie:2017:HPO

[NZS^{+}17] Yongwei Nie, Zhensong Zhang, Hanqiu Sun, Tan Su, and Guiqing Li. Homography propagation and optimization for wide-baseline street image interpolation. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2328–2341, October 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/10/07593383-
                                        abs.html.

Oliveira:2011:GEI

[OA11] Manuel M. Oliveira and Daniel G. Aliaga. Guest Editor’s introduction: Special section on the symposium on interactive 3D graphics and games (I3D). IEEE Transactions on Visualization and Computer Graphics, 17(8): 1034–1035, August 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

ODonovan:2014:LLS

[OAH14] Peter O’Donovan, Aseem Agarwala, and Aaron Hertzmann. Learning layouts for single-page graphic designs. IEEE Transactions on Visualization and Computer Graphics, 20(8):1200–1213, August 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Obermaier:2016:VTA

[OBJ16] Harald Obermaier, Kevin Bensema, and Kenneth I. Joy. Visual trends analysis in timevarying ensembles. IEEE Transactions on Visualization and Computer Graphics, 22 (10):2331–2342, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Olivier:2018:WVP

[OBKP18] Anne-Helene Olivier, Julien Bruneau, Richard Kulpa, and Julien Pettre. Walking with virtual people: Evaluation of locomotion interfaces in dynamic environments. IEEE Transactions on Visualization and Computer Graphics, 24(7):2251–2263, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07946183-
abs.html.

Overby:2017:APD

[OBLN17] Matthew Overby, George E. Brown, Jie Li, and Rahul Narain. ADMM ⊇ projective dynamics: Fast simulation of hyperelastic models with dynamic constraints. IEEE Transactions on Visualization and Computer Graphics, 24(10):2222–2234, October 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07990052-
abs.html.

Olsson:2015:MEV

[OBS^{+}15] O. Olsson, M. Billeter, E. Sintorn, V. Kampe, and U. Assarsson. More efficient virtual shadow maps for many lights. IEEE Transactions on Visualization and Computer Graphics, 21(6):701–713, June 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Oeltze:2007:IVA

[ODH^{+}07] Steffen Oeltze, Helmut Doleisch, Helwig Hauser, Philipp Muigg, and Bernhard Preim. Interactive visual analysis of perfusion data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1392–1399, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1392s.
avi.

ODonovan:2012:AIP

[OH12] Peter O’Donovan and Aaron Hertzmann. AniPaint: Interactive painterly animation from video. IEEE Transactions on Visualization and Computer Graphics, 18(3): 475–487, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Otsuka:2006:TNA

[OHH06] Rieko Otsuka, Takeshi Hoshino, and Youichi Horry. Transpost: a novel approach to the display and transmission of 360 degrees-viewable 3D solid images. IEEE Transactions on Visualization and Computer Graphics, 12(2):178– 185, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Oesterling:2011:VHD

[OHJ^{+}11] Patrick Oesterling, Christian Heine, Heike Janicke, Gerik Scheuermann, and Gerhard Heyer. Visualization of highdimensional point clouds using their density distribution’s topology. IEEE Transactions on Visualization and Computer Graphics, 17(11):1547– 1559, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ohl:2018:TIC

[Ohl18] Stephan Ohl. Tele-immersion concepts. IEEE Transactions on Visualization and Computer Graphics, 24(10):2827– 2842, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/10/08093688-
                                        abs.html.

Oesterling:2013:VPC

[OHWS13] Patrick Oesterling, Christian Heine, Gunther H. Weber, and Gerik Scheuermann. Visualizing nD point clouds as topological landscape profiles to guide local data analysis. IEEE Transactions on Visualization and Computer Graphics, 19(3):514– 526, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Orlosky:2017:EPT

[OIR^{+}17] Jason Orlosky, Yuta Itoh, Maud Ranchet, Kiyoshi Kiyokawa, John Morgan, and Hannes Devos. Emulation of physician tasks in eye-tracked virtual reality for remote diagnosis of neurodegenerative disease. IEEE Transactions on Visualization and Computer Graphics, 23(4):1302– 1311, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829437-
abs.html.

Obermaier:2012:DMT

[OJ12] Harald Obermaier and Kenneth I. Joy. Derived metric tensors for flow surface visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2149– 2158, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Obermaier:2015:AAS

[OJ15] H. Obermaier and K. I. Joy.

An automated approach for slicing plane placement in visual data analysis. IEEE Transactions on Visualization and Computer Graphics, 21(12):1403–1414, December 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Oeltze-Jafra:2016:CAV

[OJCJP16] S. Oeltze-Jafra, J. R. Cebral, G. Janiga, and B. Preim. Cluster analysis of vortical flow in simulations of cerebral aneurysm hemodynamics. IEEE Transactions on Visualization and Computer Graphics, 22(1):757–766, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ondov:2019:FFE

[OJEF19] Brian Ondov, Nicole Jardine, Niklas Elmqvist, and Steven Franconeri. Face to face: Evaluating visual comparison. IEEE Transactions on Visualization and Computer Graphics, 25(1):861– 871, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440856-
abs.html.

Okoe:2019:NLA

[OJK19] M. Okoe, R. Jianu, and S. Kobourov. Node-link or adjacency matrices: Old question, new insights. IEEE Transactions on Visualization and Computer Graphics, 25(10):2940–2952, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Orbay:2011:BDS

[OK11] Gunay Orbay and Levent Burak Kara. Beautification of design sketches using trainable stroke clustering and curve fitting. IEEE Transactions on Visualization and Computer Graphics, 17(5): 694–708, May 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Orban:2019:DTD

[OKB^{+}19] Daniel Orban, Daniel F. Keefe, Ayan Biswas, James Ahrens, and David Rogers. Drag and track: A direct manipulation interface for contextualizing data instances within a continuous parameter space. IEEE Transactions on Visualization and Computer Graphics, 25(1): 256–266, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440838-
abs.html.

Ondruska:2015:MRT

[OKI15] Peter Ondruska, Pushmeet Kohli, and Shahram Izadi. MobileFusion: Real-time volumetric surface reconstruction and dense tracking on mobile phones. IEEE Transactions on Visualization and Computer Graphics, 21(11): 1251–1258, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165662-
abs.html.

Onoue:2016:MNE

[OKSK16] Yosuke Onoue, Nobuyuki Kukimoto, Naohisa Sakamoto, and Koji Koyamada. Minimizing the number of edges via edge concentration in dense layered graphs. IEEE Transactions on Visualization and Computer Graphics, 22 (6):1652–1661, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ogawa:2009:CDS

[OM09] Michael Ogawa and Kwan-Liu Ma. code swarm: a design study in organic software visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1097– 1104, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Obermaier:2012:MFV

[OMD^{+}12] Harald Obermaier, Jan Mohring, Eduard Deines, Martin Hering-Bertram, and Hans Hagen. On mesh-free valley surface extraction with application to low frequency sound simulation. IEEE Transactions on Visualization and Computer Graphics, 18(2):270– 282, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Okaniwa:2012:UBS

[ONL^{+}12] Shoichi Okaniwa, Ahmad Nasri, Hongwei Lin, Abdulwahed Abbas, Yuki Kineri, and Takashi Maekawa. Uniform B-spline curve interpolation with prescribed tangent and curvature vectors. IEEE Transactions on Visualization and Computer Graphics, 18(9):1474–1487, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Olano:2015:IGE

[OO15] M. Olano and M. Otaduy.

I3D 2014 Guest Editor’s introduction. IEEE Transactions on Visualization and Computer Graphics, 21(6): 686–687, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ottley:2016:IBR

[OPH^{+}16] A. Ottley, E. M. Peck, L. T. Harrison, D. Afergan, C. Ziemkiewicz, H. A. Taylor, P. K. J. Han, and R. Chang. Improving Bayesian reasoning: The effects of phrasing, visualization, and spatial ability. IEEE Transactions on Visualization and Computer Graphics, 22(1):529–538, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ohlberger:1998:APO

[OR98] M. Ohlberger and M. Rumpf.

Adaptive projection operators in multiresolution scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 4(4):344– 364, October/December 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0344.
pdf;   http://www.computer.
org/tvcg/tg1998/v0344abs.

htm. See corrected reprinting [OR99].

Ohlberger:1999:APO

[OR99] Mario Ohlberger and Martin Rumpf. Adaptive projection operators in multiresolution scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 5(1): 74–94, January/March 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0074.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0074abs.

htm. Corrected printing of [OR98].

Ortega:2007:SDF

[ORC07] Michael Ortega, Stephane Redon, and Sabine Coquillart. A six degree-of-freedom Godobject method for haptic display of rigid bodies with surface properties. IEEE Transactions on Visualization and Computer Graphics, 13 (3):458–469, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

OBrien:2010:GIV

[ORRL10] Trevor O’Brien, Anna Ritz, Benjamin Raphael, and David Laidlaw. Gremlin: an interactive visualization model for analyzing genomic rearrangements. IEEE Transactions on Visualization and Computer Graphics, 16(6):918– 926, November/December 2010. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Oskiper:2015:ARB

[OSB^{+}15] T. Oskiper, M. Sizintsev, V. Branzoi, S. Samarasekera, and R. Kumar. Augmented reality binoculars. IEEE Transactions on Visualization and Computer Graphics, 21 (5):611–623, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ozer:2014:ADS

[OSBM14] Sedat Ozer, Deborah Silver, Karen Bemis, and Pino Martin. Activity detection in scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 20 (3):377–390, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ortner:2017:VWI

[OSS^{+}17] Thomas Ortner, Johannes Sorger, Harald Steinlechner, Gerd Hesina, Harald Piringer, and Eduard Groller. Vis-A-Ware: Integrating spatial and non-spatial visualization for visibility-aware urban planning. IEEE Transactions on Visualization and Computer Graphics, 23(2):1139– 1151, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2017/02/07390069-
abs.html.

Oelke:2012:VRA

[OSSK12] Daniela Oelke, David Spretke, Andreas Stoffel, and Daniel A. Keim. Visual readability analysis: How to make your writings easier to read. IEEE Transactions on Visualization and Computer Graphics, 18 (5):662–674, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Otaduy:2017:ISS

[Ota17] Miguel Otaduy. Introduction to the special section on the ACM/Eurographics Symposium on Computer Animation 2016. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2207, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/08023920.
pdf.

Orlosky:2015:MEC

[OTKS15] Jason Orlosky, Takumi Toyama, Kiyoshi Kiyokawa, and Daniel Sonntag. ModulAR: Eyecontrolled vision augmentations for head mounted displays. IEEE Transactions on Visualization and Computer Graphics, 21(11):1259– 1268, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07164337-
abs.html.

Ohl:2015:LDA

[OWS15] S. Ohl, M. Willert, and O. Staadt. Latency in distributed acquisition and rendering for telepresence systems. IEEE Transactions on Visualization and Computer Graphics, 21(12):1442– 1448, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Payan:2006:MSE

[PA06] Frederic Payan and Marc Antonini. Mean square error approximation for waveletbased semiregular mesh compression. IEEE Transactions on Visualization and Computer Graphics, 12(4):649– 657, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paul:2008:CRS

[PAB^{+}08] Brian Paul, Sean Ahern, Wes Bethel, Eric Brugger, Rich Cook, Jamison Daniel, Ken Lewis, Jens Owen, and Dale Southard. Chromium renderserver: Scalable and open remote rendering infrastructure. IEEE Transactions on Visualization and Computer Graphics, 14(3):627–639, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Parulek:2013:FBS

[PB13] Julius Parulek and Andrea Brambilla. Fast blending scheme for molecular surface representation. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2653–2662, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prachyabrued:2016:DEV

[PB16] Mores Prachyabrued and Christoph W. Borst. Design and evaluation of visual interpenetration cues in virtual grasping. IEEE Transactions on Visualization and Computer Graphics, 22(6): 1718–1731, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pietriga:2010:RIP

[PBA10] Emmanuel Pietriga, Olivier Bau, and Caroline Appert. Representation-independent in-place magnification with sigma lenses. IEEE Transactions on Visualization and Computer Graphics, 16(3): 455–467, May/June 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prouzeau:2017:EMU

[PBC17] Arnaud Prouzeau, Anastasia Bezerianos, and Olivier Chapuis. Evaluating multiuser selection for exploring graph topology on walldisplays. IEEE Transactions on Visualization and Computer Graphics, 23(8):1936– 1951, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/08/07516722-
abs.html.

Pretorius:2011:VPS

[PBCR11] A. Johannes Pretorius, Mark-Anthony Bray, Anne E. Carpenter, and Roy A. Ruddle. Visualization of parameter space for image analysis. IEEE Transactions on Visualization and Computer Graphics, 17(12):2402– 2411, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patnaik:2019:IOH

[PBE19] Biswaksen Patnaik, Andrea Batch, and Niklas Elmqvist. Information olfactation: Harnessing scent to convey data. IEEE Transactions on Visualization and Computer Graphics, 25(1):726– 736, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08444077-
                                        abs.html.

Pollock:2012:RVW

[PBK^{+}12] Brice Pollock, Melissa Burton, Jonathan W. Kelly, Stephen Gilbert, and Eliot Winer. The right view from the wrong location: Depth perception in stereoscopic multi-user virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18 (4):581–588, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paulsen:2010:MRF

[PBL10] Rasmus R. Paulsen, Jakob Andreas Baerentzen, and Rasmus Larsen. Markov random field surface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 16(4):636– 646, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Packer:2013:VAS

[PBN^{+}13] Eli Packer, Peter Bak, Mikko Nikkila, Valentin Polishchuk, and Harold J. Ship. Visual analytics for spatial clustering: Using a heuristic approach for guided exploration. IEEE Transactions on Visualization and Computer Graphics, 19(12):2179– 2188, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Palmas:2014:MVV

[PBO^{+}14] Gregorio Palmas, Myroslav Bachynskyi, Antti Oulasvirta, Hans-Peter Seidel, and Tina Weinkauf. MovExp: A versatile visualization tool for human-computer interaction studies with 3D performance and biomechanical data. IEEE Transactions on Visualization and Computer Graphics, 20(12):2359– 2368, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876050-
abs.html.

Pajot:2011:RRA

[PBPP11] Anthony Pajot, Loic Barthe, Mathias Paulin, and Pierre Poulin. Representativity for robust and adaptive multiple importance sampling. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1108–1121, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patkar:2013:WPS

[PC13] Saket Patkar and Parag Chaudhuri. Wetting of porous solids. IEEE Transactions on Visualization and Computer Graphics, 19(9):1592– 1604, September 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Poranne:2015:LSP

[PCG15] R. Poranne, Renjie Chen, and C. Gotsman. On linear spaces of polyhedral meshes. IEEE Transactions on Visualization and Computer Graphics, 21 (5):652–662, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pacanowski:2012:RB

[PCS^{+}12] Romain Pacanowski, Oliver Salazar Celis, Christophe Schlick, Xavier Granier, Pierre Poulin, and Annie Cuyt. Rational BRDF. IEEE Transactions on Visualization and Computer Graphics, 18(11):1824– 1835, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Palmerius:2008:HRD

[PCY08] Karljohan Lundin Palmerius, Matthew Cooper, and Anders Ynnerman. Haptic rendering of dynamic volumetric data. IEEE Transactions on Visualization and Computer Graphics, 14(2):263– 276, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pajarola:2004:EIR

[PD04] Renato Pajarola and Christopher DeCoro. Efficient implementation of real-time view-dependent multiresolution meshing. IEEE Transactions on Visualization and Computer Graphics, 10(3): 353–368, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/V0353-
supp.pdf;   http://csdl.
computer.org/comp/trans/
tg/2004/03/v0353abs.htm;
http://csdl.computer.org/
dl/trans/tg/2004/03/v0353.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0353.pdf.

Peck:2018:EGB

[PDBG18] Tabitha C. Peck, My Doan, Kimberly A. Bourne, and Jessica J. Good. The effect of gender body-swap illusions on working memory and stereotype threat. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1604–1612, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260949-
abs.html.

Perin:2014:RBM

[PDF14] Charles Perin, Pierre Dragicevic, and Jean-Daniel Fekete.

Revisiting Bertin matrices: New interactions for crafting tabular visualizations. IEEE Transactions on Visualization and Computer Graphics, 20(12):2082–2091, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875988-
abs.html.

Park:2018:AGU

[PDFE18] Deokgun Park, Steven M. Drucker, Roland Fernandez, and Niklas Elmqvist. Atom: A grammar for unit visualizations. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3032–3043, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8233127/.

Paczkowski:2019:PPB

[PDRK19] Patrick Paczkowski, Julie Dorsey, Holly Rushmeier, and Min H. Kim. PaperCraft3D: Paper-based 3D modeling and scene fabrication. IEEE Transactions on Visualization and Computer Graphics, 25(4):1717–1731, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8327516/.

Poco:2014:VRA

[PDW^{+}14] Jorge Poco, Aritra Dasgupta, Yaxing Wei, William Hargrove, Christopher R. Schwalm, Deborah N. Huntzinger, Robert Cook, Enrico Bertini, and Claudio T. Silva. Visual reconciliation of alternative similarity spaces in climate modeling. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1923–1932, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876041-
abs.html.

Perlin:1995:RTR

[Per95] Ken Perlin. Real time responsive animation with personality. IEEE Transactions on Visualization and Computer Graphics, 1(1): 5–15, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0005.
pdf;   http://www.computer.
org/tvcg/tg1995/v0005abs.
htm.

Phan:2018:COO

[PFC18] Huy Q. Phan, Hongbo Fu, and Antoni B. Chan. Color Orchestra: Ordering color palettes for interpolation and prediction. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1942–1955, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/06/07911336-
                                        abs.html.

Plaisant:2008:PIB

[PFG08] Catherine Plaisant, Jean-Daniel Fekete, and Georges Grinstein. Promoting insightbased evaluation of visualizations: From contest to benchmark repository. IEEE Transactions on Visualization and Computer Graphics, 14(1): 120–134, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Petrovic:2007:VWB

[PFK07] Vid Petrovic, James Fallon, and Falko Kuester. Visualizing whole-brain DTI tractography with GPU-based tuboids and LoD management. IEEE Transactions on Visualization and Computer Graphics, 13(6):1488–1495, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1488s.
                                        mov.

Prabhat:2008:CSD

[PFK^{+}08] Prabhat, Andrew Forsberg, Michael Katzourin, Kristi Wharton, and Mel Slater. A comparative study of desktop, fishtank, and cave systems for the exploration of volume rendered confocal data sets. IEEE Transactions on Visualization and Computer Graphics, 14(3):551–563, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paiva:2011:IST

[PFP^{+}11] Jose Gustavo Paiva, Laura Florian, Helio Pedrini, Guilherme Telles, and Rosane Minghim. Improved similarity trees and their application to visual data classification. IEEE Transactions on Visualization and Computer Graphics, 17(12):2459– 2468, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peck:2009:ERT

[PFW09] Tabitha C. Peck, Henry Fuchs, and Mary C. Whitton. Evaluation of reorientation techniques and distractors for walking in large virtual environments. IEEE Transactions on Visualization and Computer Graphics, 15(3):383– 394, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peck:2012:DEL

[PFW12] Tabitha C. Peck, Henry Fuchs, and Mary C. Whitton.

The design and evaluation of a large-scale real-walking locomotion interface. IEEE Transactions on Visualization and Computer Graphics, 18(7):1053–1067, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Punpongsanon:2017:ELV

[PGI^{+}17] Parinya Punpongsanon, Emilie Guy, Daisuke Iwai, Kosuke Sato, and Tamy Boubekeur. Extended LazyNav: Virtual 3D ground navigation for large displays and headmounted displays. IEEE Transactions on Visualization and Computer Graphics, 23 (8):1952–1963, August 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07501805-
                                        abs.html.

Papadopoulos:2016:VVE

[PGK16] C. Papadopoulos, I. Gutenko, and A. E. Kaufman. VEEVVIE: Visual explorer for empirical visualization, VR and interaction experiments. IEEE Transactions on Visualization and Computer Graphics, 22 (1):111–120, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peng:2012:MDV

[PGL^{+}12] Zhenmin Peng, Edward Grundy, Robert S. Laramee, Guoning Chen, and Nick Croft. Meshdriven vector field clustering and visualization: An imagebased approach. IEEE Transactions on Visualization and Computer Graphics, 18(2): 283–298, February 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ponto:2013:PCI

[PGRS13] Kevin Ponto, Michael Gleicher, Robert G. Radwin, and Hyun Joon Shin. Perceptual calibration for immersive display environments. IEEE Transactions on Visualization and Computer Graphics, 19 (4):691–700, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Palomo:2016:VET

[PGSF16] C. Palomo, Z. Guo, C. T.

Silva, and J. Freire. Visually exploring transportation schedules. IEEE Transactions on Visualization and Computer Graphics, 22(1): 170–179, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patel:2008:SAI

[PGT^{+}08] Daniel Patel, Christopher Giertsen, John Thurmond, John Gjelberg, and Eduard Grøller. The Seismic Analyzer: Interpreting and illustrating 2D seismic data. IEEE Transactions on Visualization and Computer Graphics, 14(6):1571–1578, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pilhofer:2012:CCU

[PGU12] Alexander Pilhofer, Alexander Gribov, and Antony Unwin. Comparing clusterings using Bertin’s idea. IEEE Transactions on Visualization and Computer Graphics, 18(12):2506–2515, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Perer:2013:LUS

[PGU^{+}13] Adam Perer, Ido Guy, Erel Uziel, Inbal Ronen, and Michal Jacovi. The longitudinal use of SaNDVis: Visual social network analytics in the enterprise. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1095–1108, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Petersch:2007:BFC

[PH07] Bernhard Petersch and Dieter Hönigmann. Blood flow in its context: Combining 3D Bmode and color Doppler ultrasonic data. IEEE Transactions on Visualization and Computer Graphics, 13(4):

748–757, July/August 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ponchio:2008:IRD

[PH08] Federico Ponchio and Kai Hormann. Interactive rendering of dynamic geometry. IEEE Transactions on Visualization and Computer Graphics, 14(4):914– 925, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040914s.
zip.

Pothkow:2011:PUI

[PH11] Kai Pothkow and Hans-Christian Hege. Positional uncertainty of isocontours: Condition analysis and probabilistic measures. IEEE Transactions on Visualization and Computer Graphics, 17(10):1393–1406, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pienta:2018:VIV

[PHE^{+}18] R. Pienta, F. Hohman, A. Endert, A. Tamersoy, K. Roundy, C. Gates, S. Navathe, and D. H. Chau. VIGOR: Interactive visual exploration of graph query results. IEEE Transactions on Visualization and Computer Graphics, 24(1):215–225, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Plate:2007:FMV

[PHF07] John Plate, Thorsten Holtkaemper, and Bernd Froehlich. A flexible multi-volume shader framework for arbitrarily intersecting multi-resolution datasets. IEEE Transactions on Visualization and Computer Graphics, 13(6):1584– 1591, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2007/
                                        06/v1584s.avi.

Pham:2010:VDL

[PHJ^{+}10] Tuan Pham, Rob Hess, Crystal Ju, Eugene Zhang, and Ronald Metoyer. Visualization of diversity in large multivariate data sets. IEEE Transactions on Visualization and Computer Graphics, 16(6):1053–1062, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pezzotti:2018:DPV

[PHV^{+}18] N. Pezzotti, T. Höllt, J. Van Gemert, B. P. F. Lelieveldt, E. Eisemann, and A. Vilanova. DeepEyes: Progressive visual analytics for designing deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 24(1):98–108, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Plopski:2015:CIC

[PIN^{+}15] A. Plopski, Y. Itoh, C. Nitschke, K. Kiyokawa, G. Klinker, and H. Takemura. Cornealimaging calibration for optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 21(4):481– 490, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Punpongsanon:2015:SVM

[PIS15] Parinya Punpongsanon, Daisuke Iwai, and Kosuke Sato. SoftAR: Visually manipulating haptic softness perception in spatial augmented reality. IEEE Transactions on Visualization and Computer Graphics, 21(11):1279– 1288, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165660-
abs.html.

Pasman:2003:CSI

[PJ03] W. Pasman and F. W. Jansen.

Comparing simplification and image-based techniques for 3D client-server rendering systems. IEEE Transactions on Visualization and Computer Graphics, 9(2):226– 240, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/02/v0226abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/02/
                                        v0226.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/02/v0226.pdf.

Park:2008:MOC

[PK08] Johnny Park and Avinash Kak. 3D modeling of optically challenging objects. IEEE Transactions on Visualization and Computer Graphics, 14(2):246–262, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Papadopoulos:2013:ADG

[PK13] Charilaos Papadopoulos and Arie E. Kaufman. Acuitydriven gigapixel visualization. IEEE Transactions on Visualization and Computer Graphics, 19(12):2886– 2895, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Petkov:2016:FVV

[PK16] K. Petkov and A. E. Kaufman. Frameless volume visualization. IEEE Transactions on Visualization and Computer Graphics, 22(2):1076– 1087, February 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ponto:2012:ERS

[PKG12] Kevin Ponto, Joe Kohlmann, and Michael Gleicher. Effective replays and summarization of virtual experiences. IEEE Transactions on Visualization and Computer Graphics, 18(4):607–616, April 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Park:2018:CTV

[PKL^{+}18] D. Park, S. Kim, J. Lee, J. Choo, N. Diakopoulos, and N. Elmqvist. ConceptVector: Text visual analytics via interactive lexicon building using word embedding. IEEE Transactions on Visualization and Computer Graphics, 24(1):361–370, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prisacariu:2015:RTT

[PKMR15] V. A. Prisacariu, O. Kahler, D. W. Murray, and I. D. Reid. Real-time 3d tracking and reconstruction on mobile phones. IEEE Transactions on Visualization and Computer Graphics, 21(5): 557–570, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peterka:2008:ADS

[PKS^{+}08] Tomy Peterka, Robert L. Kooima, Daniel J. Sandin, Andrew Johnson, Jason Leigh, and Thomas A. De-Fanti. Advances in the dynallax solid-state dynamic parallax barrier autostereoscopic visualization display system. IEEE Transactions on Visualization and Computer Graphics, 14(3):487–499, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Palke:2011:ATF

[PLC^{+}11a] Darrel Palke, Zhongzang Lin, Guoning Chen, Harry Yeh, Paul Vincent, Robert Laramee, and Eugene Zhang. Asymmetric tensor field visualization for surfaces. IEEE Transactions on Visualization and Computer Graphics, 17(12):1979–1988, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Parry:2011:HES

[PLC^{+}11b] Matthew L. Parry, Philip A. Legg, David H. S. Chung, Iwan W. Griffiths, and Min Chen. Hierarchical event selection for video storyboards with a case study on snooker video visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):1747–1756, December 2011. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Piumsomboon:2018:SVG

[PLE^{+}18] Thammathip Piumsomboon, Gun A. Lee, Barrett Ens, Bruce H. Thomas, and Mark Billinghurst. Superman vs Giant: A study on spatial perception for a multi-scale mixed reality flying telepresence interface. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2974–2982, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8466636/.

Park:2006:DSI

[PLK^{+}06] Sung W. Park, Lars Linsen, Oliver Kreylos, John D. Owens, and Bernd Hamann. Discrete sibson interpolation. IEEE Transactions on Visualization and Computer Graphics, 12(2):243– 253, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Park:2012:ASI

[PLK12] Taejung Park, Sung-Ho Lee, and Chang-Hun Kim. Analytic solutions of integral moving least squares for polygon soups. IEEE Transactions on Visualization and Computer Graphics, 18(10):1638– 1649, October 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Partl:2014:CDD

[PLS^{+}14] Christian Partl, Alexander Lex, Marc Streit, Hendrik Strobelt, Anne-Mai Wassermann, Hanspeter Pfister, and Dieter Schmalstieg. ConTour: Data-driven exploration of multi-relational datasets for drug discovery. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1883–1892, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875994-
abs.html.

Pezzotti:2017:AUS

[PLvdM^{+}17] Nicola Pezzotti, Boudewijn P. F. Lelieveldt, Laurens van der Maaten, Thomas Hollt, Elmar Eisemann, and Anna Vilanova. Approximated and user steerable tSNE for progressive visual analytics. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1739–1752, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07473883-
abs.html.

Park:2011:EKD

[PLW11] Youngmin Park, Vincent Lepetit, and Woontack Woo. Extended keyframe detection with stable tracking for multiple 3D object tracking. IEEE Transactions on Visualization and Computer Graphics, 17(11):1728–1735, November 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Park:2012:HMB

[PLW12] Youngmin Park, Vincent Lepetit, and Woontack Woo. Handling motion-blur in 3D tracking and rendering for augmented reality. IEEE Transactions on Visualization and Computer Graphics, 18(9):1449–1459, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paulovich:2008:HNH

[PM08] Fernando V. Paulovich and Rosane Minghim. HiPP: a novel hierarchical point placement strategy and its application to the exploration of document collections. IEEE Transactions on Visualization and Computer Graphics, 14(6):1229–1236, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pietroni:2011:ILF

[PMCS11] Nico Pietroni, Corsini Massimiliano, Paolo Cignoni, and Roberto Scopigno. An interactive local flattening operator to support digital investigations on artwork surfaces. IEEE Transactions on Visualization and Computer Graphics, 17(12):1989– 1996, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pugmire:2007:NBI

[PMD^{+}07] David Pugmire, Laura Monroe, Carolyn Connor Davenport, Andrew DuBois, David DuBois, and Stephen Poole. NPU-based image compositing in a distributed visualization system. IEEE Transactions on Visualization and Computer Graphics, 13(4): 798–809, July/August 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Poco:2018:ERC

[PMH18] J. Poco, A. Mayhua, and J. Heer. Extracting and retargeting color mappings from bitmap images of visualizations. IEEE Transactions on Visualization and Computer Graphics, 24(1):637–646, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ponamgi:1997:IAC

[PML97] M. K. Ponamgi, D. Manocha, and M. C. Lin. Incremental algorithms for collision detection between polygonal models. IEEE Transactions on Visualization and Computer Graphics, 3(1):51–64, January/March 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0051.
pdf;   http://www.computer.
org/tvcg/tg1997/v0051abs.
htm.

Pandey:2014:PPD

[PMN^{+}14] Anshul Vikram Pandey, Anjali Manivannan, Oded Nov, Margaret Satterthwaite, and Enrico Bertini. The persuasive power of data visualization. IEEE Transactions on Visualization and Computer Graphics, 20(12):2211– 2220, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876023-
abs.html.

Papaioannou:2010:RTV

[PMP10] Georgios Papaioannou, Maria Lida Menexi, and Charilaos Papadopoulos. Real-time volumebased ambient occlusion. IEEE Transactions on Visualization and Computer Graphics, 16(5):752–762, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pujades:2019:VCR

[PMT^{+}19] Sergi Pujades, Betty Mohler, Anne Thaler, Joachim Tesch, Naureen Mahmood, Nikolas Hesse, Heinrich H. Bülthoff, and Michael J. Black. The virtual caliper: Rapid creation of metrically accurate avatars from 3D measurements. IEEE Transactions on Visualization and Computer Graphics, 25(5):1887–1897, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8648222/.

Pares:2005:ADC

[PMvWC05] Narcis Pares, Paul Masri, Gerard van Wolferen, and Chris Creed. Achieving dialogue with children with severe autism in an adaptive multisensory interaction: The “MEDIATE” Project. IEEE Transactions on Visualization and Computer Graphics, 11(6):734–743, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pfaffelmoser:2013:VVG

[PMW13] Tobias Pfaffelmoser, Mihaela Mihai, and Rudiger Westermann. Visualizing the variability of gradients in uncertain 2D scalar fields. IEEE Transactions on Visualization and Computer Graphics, 19(11):1948–1961, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paulovich:2008:LSP

[PNML08] Fernando V. Paulovich, Luis G.

Nonato, Rosane Minghim, and Haim Levkowitz. Least square projection: a fast highprecision multidimensional projection technique and its application to document mapping. IEEE Transactions on Visualization and Computer Graphics, 14(3):564– 575, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prilepov:2013:CGB

[POD^{+}13] Iuri Prilepov, Harald Obermaier, Eduard Deines, Christoph Garth, and Kenneth I. Joy. Cubic gradient-based material interfaces. IEEE Transactions on Visualization and Computer Graphics, 19(10):1687– 1699, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Preim:2009:SVE

[POM^{+}09] Bernhard Preim, Steffen Oeltze, Matej Mlejnek, Eduard Gröeller, Anja Hennemuth, and Sarah Behrens. Survey of the visual exploration and analysis of perfusion data. IEEE Transactions on Visualization and Computer Graphics, 15(2):205– 220, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Puppo:2009:RS

[PP09] Enrico Puppo and Daniele Panozzo. RGB subdivision. IEEE Transactions on Visualization and Computer Graphics, 15(2):295– 310, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Parker:1999:IRT

[PPL^{+}99] Steven Parker, Michael Parker, Yarden Livnat, Peter-Pike Sloan, Charles Hansen, and Peter Shirley. Interactive ray tracing for volume visualization. IEEE Transactions on Visualization and Computer Graphics, 5(3):238–250, July/September 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0238.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0238abs.
                                        htm.

Petronetto:2010:MHH

[PPL^{+}10] Fabiano Petronetto, Afonso Paiva, Marcos Lage, Geovan Tavares, Helio Lopes, and Thomas Lewiner. Meshless Helmholtz–Hodge decomposition. IEEE Transactions on Visualization and Computer Graphics, 16(2):338– 349, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pu:2011:SBR

[PPM^{+}11] Jiantao Pu, David S. Paik, Xin Meng, Justus E. Roos, and Geoffrey D. Rubin. Shape “break-and-repair” strategy and its application to automated medical image segmentation. IEEE Transactions on Visualization and Computer Graphics, 17(1):115–124, January 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Purchase:2012:GDA

[PPP12] Helen C. Purchase, Christopher Pilcher, and Beryl Plimmer. Graph drawing aesthetics — created by users, not algorithms. IEEE Transactions on Visualization and Computer Graphics, 18(1):81– 92, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Panozzo:2011:ACQ

[PPT^{+}11] Daniele Panozzo, Enrico Puppo, Marco Tarini, Nico Pietroni, and Paolo Cignoni. Automatic construction of quad-based subdivision surfaces using fitmaps. IEEE Transactions on Visualization and Computer Graphics, 17(10):1510–1520, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prckovska:2011:FDH

[PPvA^{+}11] Vesna Prčkovska, Tim H. J. M. Peeters, Markus van Almsick, Bart ter Haar Romeny, and Anna Vilanova. Fused DTI/HARDI visualization. IEEE Transactions on Visualization and Computer Graphics, 17(10):1407– 1419, October 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Petkov:2012:IVR

[PPZ^{+}12] Kaloian Petkov, Charilaos Papadopoulos, Min Zhang, Arie E. Kaufman, and Xianfeng David Gu. Interactive visibility retargeting in VR using conformal visualization. IEEE Transactions on Visualization and Computer Graphics, 18(7):1027–1040, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Petkov:2009:ELV

[PQF^{+}09] Kaloian Petkov, Feng Qiu, Zhe Fan, Arie E. Kaufman, and Klaus Mueller. Efficient LBM visual simulation on face-centered cubic lattices. IEEE Transactions on Visualization and Computer Graphics, 15(5):802–814, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Padilla:2017:EIB

[PQMCR17] Lace Padilla, P. Samuel Quinan, Miriah Meyer, and Sarah H. Creem-Regehr. Evaluating the impact of binning 2D scalar fields. IEEE Transactions on Visualization and Computer Graphics, 23 (1):431–440, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pajarola:2000:CPM

[PR00a] R. Pajarola and J. Rossignac.

Compressed progressive meshes. IEEE Transactions on Visualization and Computer Graphics, 6(1):79–93, January/March 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0079.
pdf;   http://www.computer.
org/tvcg/tg2000/v0079abs.

htm. See corrections [PR00b].

Pajarola:2000:CCP

[PR00b] R. Pajarola and J. Rossignac.

Corrections to “Compressed Progressive Meshes”. IEEE Transactions on Visualization and Computer Graphics, 6 (2):190–192, April/June 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/

                                        tg/books/tg2000/pdf/v0190.

pdf. See [PR00a].

Popescu:2010:GPC

[PRA^{+}10] Voicu Popescu, Paul Rosen, Laura Arns, Xavier Tricoche, Chris Wyman, and Christoph M. Hoffmann. The general pinhole camera: Effective and efficient nonuniform sampling for visualization. IEEE Transactions on Visualization and Computer Graphics, 16(5):777–790, September/October 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prassni:2010:UAG

[PRH10] Jorg-Stefan Prassni, Timo Ropinski, and Klaus Hinrichs. Uncertainty-aware guided volume segmentation. IEEE Transactions on Visualization and Computer Graphics, 16(6):1358–1365, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Perer:2006:BSF

[PS06] Adam Perer and Ben Shneiderman. Balancing systematic and flexible exploration of social networks. IEEE Transactions on Visualization and Computer Graphics, 12(5): 693–700, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patel:2012:ACL

[PS12] Ankur Patel and William A. P. Smith. Automated construction of low-resolution, texture-mapped, class-optimal meshes. IEEE Transactions on Visualization and Computer Graphics, 18(3):434– 446, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pileggi:2012:SVP

[PSBS12] Hannah Pileggi, Charles D.

Stolper, J. Michael Boyle, and John T. Stasko. Snap-Shot: Visualization to propel ice hockey analytics. IEEE Transactions on Visualization and Computer Graphics, 18(12):2819–2828, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patane:2009:MCA

[PSF09] Giuseppe Patanè, Michela Spagnuolo, and Bianca Falcidieno. A minimal contouring approach to the computation of the Reeb graph. IEEE Transactions on Visualization and Computer Graphics, 15(4):583–595, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pajarola:2004:COS

[PSG04] Renato Pajarola, Miguel Sainz, and Patrick Guidotti.

Confetti: Object-space point blending and splatting. IEEE Transactions on Visualization and Computer Graphics, 10(5):598–608, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0598.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0598.pdf.

Panse:2006:VGS

[PSKN06] Christian Panse, Mike Sips, Daniel Keim, and Stephen North. Visualization of geospatial point sets via global shape transformation and local pixel placement. IEEE Transactions on Visualization and Computer Graphics, 12(5):749–756, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Popescu:2006:SBC

[PSM06] Voicu Popescu, Elisha Sacks, and Chunhui Mei. Samplebased cameras for feed forward reflection rendering. IEEE Transactions on Visualization and Computer Graphics, 12(6):1590–1600, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pousman:2007:CIV

[PSM07] Zachary Pousman, John Stasko, and Michael Mateas. Casual information visualization: Depictions of data in everyday life. IEEE Transactions on Visualization and Computer Graphics, 13(6):1145–1152, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pohl:2012:UPE

[PSM12] Margit Pohl, Michael Smuc, and Eva Mayr. The user puzzle — explaining the interaction with visual analytics systems. IEEE Transactions on Visualization and Computer Graphics, 18(12):2908– 2916, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paulovich:2010:TPM

[PSN10] Fernando V. Paulovich, Claudio T. Silva, and Luis G. Nonato. Two-phase mapping for projecting massive data sets. IEEE Transactions on Visualization and Computer Graphics, 16(6):1281– 1290, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Paiva:2015:ASI

[PSPM15] Jose Gustavo S. Paiva, William Robson Schwartz, Helio Pedrini, and Rosane Minghim. An approach to supporting incremental visual data classification. IEEE Transactions on Visualization and Computer Graphics, 21 (1):4–17, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06840370-
abs.html.

Peng:2017:GAA

[PSR17] Chao Peng, Sandip Sahani, and John Rushing. A GPU-accelerated approach for feature tracking in timevarying imagery datasets. IEEE Transactions on Visualization and Computer Graphics, 24(10):2262–2274, October 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07779153-
abs.html.

Pahins:2017:HSL

[PSSC17] Cícero A. L. Pahins, Sean A. Stephens, Carlos Scheidegger, and João L. D. Comba. Hashedcubes: Simple, low memory, real-time visual exploration of big data. IEEE Transactions on Visualization and Computer Graphics, 23 (1):671–680, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Philip:2015:DSG

[PST^{+}15] Sujin Philip, Brian Summa, Julien Tierny, Peer-Timo Bremer, and Valerio Pascucci. Distributed seams for gigapixel panoramas. IEEE Transactions on Visualization and Computer Graphics, 21 (3):350–362, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06940309-
abs.html.

Pajer:2017:WVW

[PSTW^{+}17] Stephan Pajer, Marc Streit, Thomas Torsney-Weir, Florian Spechtenhauser, Torsten Möller, and Harald Piringer. WeightLifter: Visual weight space exploration for multicriteria decision making. IEEE Transactions on Visualization and Computer Graphics, 23(1):611–620, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peer:2017:PVG

[PT17] Andreas Peer and Matthias Teschner. Prescribed velocity gradients for highly viscous SPH fluids with vorticity diffusion. IEEE Transactions on Visualization and Computer Graphics, 23(12): 2656–2662, December 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/12/07775068-
                                        abs.html.

Pietroni:2010:AIM

[PTC10] Nico Pietroni, Marco Tarini, and Paolo Cignoni. Almost isometric mesh parameterization through abstract domains. IEEE Transactions on Visualization and Computer Graphics, 16(4):621– 635, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Polvi:2018:HGI

[PTM^{+}18] Jarkko Polvi, Takafumi Taketomi, Atsunori Moteki, Toshiyuki Yoshitake, Toshiyuki Fukuoka, Goshiro Yamamoto, Christian Sandor, and Hirokazu Kato. Handheld guides in inspection tasks: Augmented reality versus picture. IEEE Transactions on Visualization and Computer Graphics, 24(7):2118–2128, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07935524-
                                        abs.html.

Piringer:2009:MTA

[PTMB09] Harald Piringer, Christian Tominski, Philipp Muigg, and Wolfgang Berger. A multithreading architecture to support interactive visual exploration. IEEE Transactions on Visualization and Computer Graphics, 15(6):1113– 1120, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Prendinger:2011:MSA

[PUNI11] Helmut Prendinger, Sebastian Ullrich, Arturo Nakasone, and Mitsuru Ishizuka. MPML3D: Scripting agents for the 3D Internet. IEEE Transactions on Visualization and Computer Graphics, 17(5):655–668, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Purgathofer:2009:GEI

[Pur09] Werner Purgathofer. Guest Editor’s introduction: Special section on VRST. IEEE Transactions on Visualization and Computer Graphics, 15(1):4–5, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pretorius:2006:VAM

[PV06] A. Johannes Pretorius and Jarke J. Van Wijk. Visual analysis of multivariate state transition graphs. IEEE Transactions on Visualization and Computer Graphics, 12(5):685–692, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Patil:2011:DCS

[PvdBC^{+}11] Sachin Patil, Jur van den Berg, Sean Curtis, Ming C. Lin, and Dinesh Manocha. Directing crowd simulations using navigation fields. IEEE Transactions on Visualization and Computer Graphics, 17 (2):244–254, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Perin:2013:SKV

[PVF13] Charles Perin, Romain Vuillemot, and Jean-Daniel Fekete. SoccerStories: A kick-off for visual soccer analysis. IEEE Transactions on Visualization and Computer Graphics, 19(12):2506–2515, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pagendarm:1995:CCC

[PW95] Hans-Georg Pagendarm and Birgit Walter. Competent, compact, comparative visualization of a vortical flow field. IEEE Transactions on Visualization and Computer Graphics, 1(2):142–150, June 1995. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0142.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0142abs.
                                        htm.

Pineo:2012:DVO

[PW12] Daniel Pineo and Colin Ware.

Data visualization optimization via computational modeling of perception. IEEE Transactions on Visualization and Computer Graphics, 18 (2):309–320, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pilar:2013:RFP

[PW13] David H. F. Pilar and Colin Ware. Representing flow patterns by using streamlines with glyphs. IEEE Transactions on Visualization and Computer Graphics, 19(8): 1331–1341, August 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pjanic:2017:GPC

[PWG17] Petar Pjanic, Simon Willi, and Anselm Grundhofer. Geometric and photometric consistency in a mixed video and galvanoscopic scanning laser projection mapping system. IEEE Transactions on Visualization and Computer Graphics, 23(11):2430– 2439, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007213-
abs.html.

Pick:2016:DED

[PWHK16] S. Pick, B. Weyers, B. Hentschel, and T. W. Kuhlen. Design and evaluation of data annotation workflows for CAVElike virtual environments. IEEE Transactions on Visualization and Computer Graphics, 22(4):1452–1461, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pjanic:2018:SMP

[PWIG18] Petar Pjanic, Simon Willi, Daisuke Iwai, and Anselm Grundhöfer. Seamless multiprojection revisited. IEEE Transactions on Visualization and Computer Graphics, 24(11):2963–2973, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL https://ieeexplore.
                                        ieee.org/document/8497080/

.

Poullis:2009:PLS

[PY09] Charalambos Poullis and Suya You. Photorealistic large-scale urban city model reconstruction. IEEE Transactions on Visualization and Computer Graphics, 15(4): 654–669, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Polk:2014:TVT

[PYHZ14] Tom Polk, Jing Yang, Yueqi Hu, and Ye Zhao. TenniVis:

Visualization for tennis match analysis. IEEE Transactions on Visualization and Computer Graphics, 20(12):2339– 2348, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876044-
abs.html.

Palacios:2016:FSS

[PYW^{+}16] J. Palacios, H. Yeh, W. Wang, Y. Zhang, R. S. Laramee, R. Sharma, T. Schultz, and E. Zhang. Feature surfaces in symmetric tensor fields based on eigenvalue manifold. IEEE Transactions on Visualization and Computer Graphics, 22 (3):1248–1260, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pei:2007:TSM

[PZ07] Yuru Pei and Hongbin Zha.

Transferring of speech movements from video to 3D face space. IEEE Transactions on Visualization and Computer Graphics, 13(1):58–69, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Palacios:2011:IVR

[PZ11] J. Palacios and E. Zhang. Interactive visualization of rotational symmetry fields on surfaces. IEEE Transactions on Visualization and Computer Graphics, 17(7):947–955, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pajarola:2012:GEI

[PZ12] Renato Pajarola and Kun Zhou. Guest Editor’s introduction: Special section on the Eurographics Symposium on Parallel Graphics and Visualization (EGPGV). IEEE Transactions on Visualization and Computer Graphics, 18 (6):837, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pan:2017:SRD

[PZLZ17] Xiao Pan, Yuanfeng Zhou, Feng Li, and Caiming Zhang. Superpixels of RGB-D images for indoor scenes based on weighted geodesic driven metric. IEEE Transactions on Visualization and Computer Graphics, 24(10):2342– 2356, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/10/07707351-
                                        abs.html.

Qin:2014:CTF

[QCH^{+}14] Hao Qin, Menglei Chai, Qiming Hou, Zhong Ren, and Kun Zhou. Cone tracing for furry object rendering. IEEE Transactions on Visualization and Computer Graphics, 20 (8):1, August 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qi:2013:CCD

[QCT13] Meng Qi, Thanh-Tung Cao, and Tiow-Seng Tan. Computing 2D constrained Delaunay triangulation using the GPU. IEEE Transactions on Visualization and Computer Graphics, 19(5):736–748, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qu:2007:VAA

[QCX^{+}07] Huamin Qu, Wing-Yi Chan, Anbang Xu, Kai-Lun Chung, Kai-Hon Lau, and Ping Guo. Visual analysis of the air pollution problem in Hong Kong. IEEE Transactions on Visualization and Computer Graphics, 13(6):1408–1415, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1408s.
mov.

Qu:2018:KMV

[QH18] Z. Qu and J. Hullman. Keeping multiple views consistent: Constraints, validations, and exceptions in visualization authoring. IEEE Transactions on Visualization and Computer Graphics, 24(1):468– 477, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qin:2009:GEI

[Qin09] Hong Qin. Guest Editor’s introduction: Special section on shape, solid, and physical modeling. IEEE Transactions on Visualization and Computer Graphics, 15(4): 529, July/August 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qu:2004:BFS

[QK04] Huamin Qu and Arie E. Kaufman. O-buffer: a framework for sample-based graphics. IEEE Transactions on Visualization and Computer Graphics, 10(4):410– 421, July/August 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        dl/trans/tg/2004/04/v0410.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/04/
                                        v0410.pdf.

Quinn:2013:GAS

[QLLM13] Jonathan A. Quinn, Frank C. Langbein, Yu-Kun Lai, and Ralph R. Martin. Generalized anisotropic stratified surface sampling. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1143–1157, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qu:2008:PGP

[QM08] Lijun Qu and Gary W.

Meyer. Perceptually guided polygon reduction. IEEE Transactions on Visualization and Computer Graphics, 14(5):1015–1029, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Quinan:2016:VCW

[QM16] P. S. Quinan and M. Meyer.

Visually comparing weather features in forecasts. IEEE Transactions on Visualization and Computer Graphics, 22 (1):389–398, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qiao:2006:HBS

[QMK^{+}06] Wei Qiao, Michael McLennan, Rick Kennell, David Ebert, and Gerhard Klimeck. Hub-based simulation and graphics hardware accelerated visualization for nanotechnology applications. IEEE Transactions on Visualization and Computer Graphics, 12(5):1061–1068, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qin:1998:DCC

[QMV98] H. Qin, C. Mandal, and B. C. Vemuri. Dynamic Catmull–Clark subdivision surfaces. IEEE Transactions on Visualization and Computer Graphics, 4(3):215–229, July/September 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0215.
pdf;   http://www.computer.
org/tvcg/tg1998/v0215abs.
htm.

Qian:2018:RAO

[QPNK18] Long Qian, Alexander Plopski, Nassir Navab, and Peter Kazanzides. Restoring the awareness in the occluded visual field for optical seethrough head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 24(11):2936– 2946, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8456571/.

Qin:1996:DNP

[QT96] Hong Qin and Demetri Terzopoulos. D-NURBS: a physics-based framework for geometric design. IEEE Transactions on Visualization and Computer Graphics, 2 (1):85–96, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0085.
pdf;   http://www.computer.
org/tvcg/tg1996/v0085abs.
htm.

Qu:2009:FCR

[QWC^{+}09] Huamin Qu, Haomian Wang, Weiwei Cui, Yingcai Wu, and Ming-Yuen Chan. Focus + context route zooming and information overlay in 3D urban environments. IEEE Transactions on Visualization and Computer Graphics, 15(6):1547–1554, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qiu:2007:LBV

[QXF^{+}07] Feng Qiu, Fang Xu, Zhe Fan, Neophytou Neophytos, Arie Kaufman, and Klaus Mueller. Lattice-based volumetric global illumination. IEEE Transactions on Visualization and Computer Graphics, 13(6):1576–1583, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qin:2007:AT

[QY07] Xuejie Qin and Yee-Hong Yang. Aura 3D textures. IEEE Transactions on Visualization and Computer Graphics, 13(2):379– 389, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Quigley:2018:RTI

[QYH^{+}18] Ed Quigley, Yue Yu, Jingwei Huang, Winnie Lin, and Ronald Fedkiw. Realtime interactive tree animation. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1717–1727, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/05/07836345-
                                        abs.html.

Ren:2017:SSI

[RAL^{+}17] Donghao Ren, Saleema Amershi, Bongshin Lee, Jina Suh, and Jason D. Williams. Squares: Supporting interactive performance analysis for multiclass classifiers. IEEE Transactions on Visualization and Computer Graphics, 23 (1):61–70, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rocha:2017:DMR

[RASS17] Allan Rocha, Usman Alim, Julio Daniel Silva, and Mario Costa Sousa. Decalmaps: Real-time layering of decals on surfaces for multivariate visualization. IEEE Transactions on Visualization and Computer Graphics, 23 (1):821–830, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ramanarayanan:2007:CTS

[RB07] Ganesh Ramanarayanan and Kavita Bala. Constrained texture synthesis via energy minimization. IEEE Transactions on Visualization and Computer Graphics, 13(1):167– 178, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rodgers:2011:EAA

[RB11] Johnny Rodgers and Lyn Bartram. Exploring ambient and artistic visualization for residential energy use feedback. IEEE Transactions on Visualization and Computer Graphics, 17(12):2489– 2497, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rubel:2018:BSR

[RB18] O. Rübel and B. P. Bowen.

BASTet: Shareable and reproducible analysis and visualization of mass spectrometry imaging data via Open-MSI. IEEE Transactions on Visualization and Computer Graphics, 24(1):1025– 1035, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rohmer:2015:INF

[RBDG15] K. Rohmer, W. Buschel, R. Dachselt, and T. Grosch. Interactive near-field illumination for photorealistic augmented reality with varying materials on mobile devices. IEEE Transactions on Visualization and Computer Graphics, 21(12):1349– 1362, December 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rautek:2007:SLI

[RBG07] Peter Rautek, Stefan Bruckner, and Eduard Gröller. Semantic layers for illustrative volume rendering. IEEE Transactions on Visualization and Computer Graphics, 13(6):1336–1343, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1336s.
mov.

Rautek:2014:VSI

[RBGH14] Peter Rautek, Stefan Bruckner, M. Eduard Groller, and Markus Hadwiger. ViSlang: A system for interpreted domain-specific languages for scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 20(12):2388– 2396, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876040-
abs.html.

Ragan:2015:EFV

[RBK^{+}15] Eric D. Ragan, Doug A. Bowman, Regis Kopper, Cheryl Stinson, Siroberto Scerbo, and Ryan P. McMahan. Effects of field of view and visual complexity on virtual reality training effectiveness for a visual scanning task. IEEE Transactions on Visualization and Computer Graphics, 21(7):794–807, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
07/07042312-abs.html.

Randhavane:2019:FMP

[RBK^{+}19] T. Randhavane, A. Bera, K. Kapsaskis, K. Gray, and D. Manocha. FVA: Modeling perceived friendliness of virtual agents using movement characteristics. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3135–3145, November 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rusdorf:2007:RTI

[RBLW07] Stephan Rusdorf, Guido Brunnett, Mario Lorenz, and Tobias Winkler. Real-time interaction with a humanoid avatar in an immersive table tennis simulation. IEEE Transactions on Visualization and Computer Graphics, 13 (1):15–25, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Raith:2019:TFV

[RBN^{+}19] Felix Raith, Christian Blecha, Thomas Nagel, Francesco Parisio, Olaf Kolditz, Fabian Gunther, Markus Stommel, and Gerik Scheuermann. Tensor field visualization using fiber surfaces of invariant space. IEEE Transactions on Visualization and Computer Graphics, 25(1):1122– 1131, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08447439-
abs.html.

Rautenhaus:2018:VMS

[RBS^{+}18] Marc Rautenhaus, Michael Böttinger, Stephan Siemen, Robert Hoffman, Robert M. Kirby, Mahsa Mirzargar, Niklas Röber, and Rüdiger Westermann. Visualization in meteorology — a survey of techniques and tools for data analysis tasks. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3268–3296, ???? 2018.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8126857/.

Rutten:2006:AVB

[RC06] Markus Rütten and Min S.

Chong. Analyzing vortex breakdown flow structures by assignment of colors to tensor invariants. IEEE Transactions on Visualization and Computer Graphics, 12(5):1189–1196, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Robb:2015:TVH

[RCL^{+}15] A. Robb, A. Cordar, S. Lampotang, C. White, A. Wendling, and B. Lok. Teaming up with virtual humans: How other people change our perceptions of and behavior with virtual teammates. IEEE Transactions on Visualization and Computer Graphics, 21 (4):511–519, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reckziegel:2018:PTM

[RCSJ18] Martin Reckziegel, Muhammad Faisal Cheema, Gerik Scheuermann, and Stefan Janicke. Predominance tag maps. IEEE Transactions on Visualization and Computer Graphics, 24(6):1893– 1904, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/08320795-
abs.html.

Reichherzer:2018:NSM

[RCW^{+}18] Carolin Reichherzer, Andrew Cunningham, James Walsh, Mark Kohler, Mark Billinghurst, and Bruce H. Thomas. Narrative and spatial memory for jury viewings in a reconstructed virtual environment. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2917–2926, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8462799/.

Reinhard:2005:DRR

[RD05] Erik Reinhard and Kate Devlin. Dynamic range reduction inspired by photoreceptor physiology. IEEE Transactions on Visualization and Computer Graphics, 11(1):13– 24, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0013abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2005/01/
v0013.htm;   http://csdl.
computer.org/dl/trans/tg/
2005/01/v0013.pdf.

Riche:2010:UED

[RD10] Nathalie Henry Riche and Tim Dwyer. Untangling Euler diagrams. IEEE Transactions on Visualization and Computer Graphics, 16(6):1090– 1099, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ropinski:2012:UBA

[RDB^{+}12] Timo Ropinski, Stefan Diepenbrock, Stefan Bruckner, Klaus Hinrichs, and Eduard Gröller. Unified boundary-aware texturing for interactive volume rendering. IEEE Transactions on Visualization and Computer Graphics, 18(11):1942– 1955, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Raviv:2001:IDR

[RE01a] A. Raviv and G. Elber. Interactive direct rendering of trivariate B-spline scalar functions. IEEE Transactions on Visualization and Computer Graphics, 7(2):109– ??, April/June 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0109.
pdf;   http://www.computer.
org/tvcg/tg2001/v0109abs.
htm.

Rheingans:2001:VIN

[RE01b] Penny Rheingans and David Ebert. Volume illustration: Nonphotorealistic rendering of volume models. IEEE Transactions on Visualization and Computer Graphics, 7 (3):253–264, July 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2001/pdf/v0253.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0253abs.
                                        htm.

Ricks:2014:WSA

[RE14] Brian C. Ricks and Parris K. Egbert. A whole surface approach to crowd simulation on arbitrary topologies. IEEE Transactions on Visualization and Computer Graphics, 20 (2):159–171, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Raidou:2016:OEP

[REB^{+}16] R. G. Raidou, M. Eisemann, M. Breeuwer, E. Eisemann, and A. Vilanova. Orientationenhanced parallel coordinate plots. IEEE Transactions on Visualization and Computer Graphics, 22(1):589–598, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ragan:2016:CPV

[RESC16] E. D. Ragan, A. Endert, J. Sanyal, and J. Chen. Characterizing provenance in visualization and data analysis: An organizational framework of provenance types and purposes. IEEE Transactions on Visualization and Computer Graphics, 22(1):31–40, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rueda:2011:RNG

[RF11] Antonio J. Rueda and Francisco R. Feito. EL-REP: a new 2D geometric decomposition scheme and its applications. IEEE Transactions on Visualization and Computer Graphics, 17(9):1325–1336, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Robertson:2008:EAT

[RFF^{+}08] George Robertson, Roland Fernandez, Danyel Fisher, Bongshin Lee, and John Stasko. Effectiveness of animation in trend visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1325–1332, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rauber:2017:VHA

[RFFT17] Paulo E. Rauber, Samuel G.

Fadel, Alexandre X. Falcão, and Alexandru C. Telea. Visualizing the hidden activity of artificial neural networks. IEEE Transactions on Visualization and Computer Graphics, 23(1):101–110, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rieck:2018:CCP

[RFLL18] B. Rieck, U. Fugacci, J. Lukasczyk, and H. Leitte. Clique community persistence: A topological visual analysis approach for complex networks. IEEE Transactions on Visualization and Computer Graphics, 24 (1):822–831, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rolin:2019:PLM

[RFSP19] R. A. Rolin, J. Fooken, M. Spering, and D. K. Pai. Perception of looming motion in virtual reality egocentric interception tasks. IEEE Transactions on Visualization and Computer Graphics, 25(10):3042–3048, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roytman:1995:DCQ

[RG95] Evgeny Roytman and Craig Gotsman. Dynamic color quantization of video sequences. IEEE Transactions on Visualization and Computer Graphics, 1(3):274– 286, September 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0274.
pdf;   http://www.computer.
org/tvcg/tg1995/v0274abs.
htm.

Rubel:2014:FBA

[RGC^{+}14] Oliver Rubel, Cameron G. R.

Geddes, Min Chen, Estelle Cormier-Michel, and E. Wes Bethel. Feature-based analysis of plasma-based particle acceleration data. IEEE Transactions on Visualization and Computer Graphics, 20 (2):196–210, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Raidou:2019:RDS

[RGE19] Renata G. Raidou, M. Eduard Gröller, and Martin Eisemann. Relaxing dense scatter plots with pixel-based mappings. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2205–2216, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8667672/.

Raghupathi:2004:ISS

[RGF^{+}04] Laks Raghupathi, Laurent Grisoni, François Faure, Damien Marchal, Marie-Paule Cani, and Christophe Chaillou. An intestinal surgery simulator: Real-time collision processing and visualization. IEEE Transactions on Visualization and Computer Graphics, 10(6):708–718, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0708abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0708.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0708.pdf.

Rivera-Gutierrez:2014:GPA

[RGFLL14] Diego Rivera-Gutierrez, Rick Ferdig, Jian Li, and Benjamin Lok. Getting the point across: Exploring the effects of dynamic virtual humans in an interactive museum exhibit on user perceptions. IEEE Transactions on Visualization and Computer Graphics, 20 (4):636–643, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reh:2013:MNM

[RGK^{+}13] Andreas Reh, Christian Gusenbauer, Johann Kastner, M. Eduard Groller, and Christoph Heinzl. MObjects — a novel method for the visualization and interactive exploration of defects in industrial XCT data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2906– 2915, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Riehmann:2012:WKC

[RGP^{+}12] Patrick Riehmann, Henning Gruendl, Martin Potthast, Martin Trenkmann, Benno Stein, and Bernd Froehlich. WORDGRAPH: Keyword-in-context visualization for NETSPEAK’s wildcard search. IEEE Transactions on Visualization and Computer Graphics, 18(9): 1411–1423, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rajamki:2019:CCM

[RH19] J. Rajamäki and P. Hämäläinen.

Continuous control Monte Carlo tree search informed by multiple experts. IEEE Transactions on Visualization and Computer Graphics, 25 (8):2540–2553, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ritter:2006:RTI

[RHD^{+}06] Felix Ritter, Christian Hansen, Volker Dicken, Olaf Konrad, Bernhard Preim, and Heinz-Otto Peitgen. Realtime illustration of vascular structures. IEEE Transactions on Visualization and Computer Graphics, 12(5): 877–884, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rameau:2016:RTA

[RHJ^{+}16] François Rameau, Hyowon Ha, Kyungdon Joo, Jinsoo Choi, Kibaek Park, and In So Kweon. A real-time augmented reality system to seethrough cars. IEEE Transactions on Visualization and Computer Graphics, 22(11): 2395–2404, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ropinski:2009:MVV

[RHR^{+}09] Timo Ropinski, Sven Hermann, Rainer Reich, Michael Schafers, and Klaus Hinrichs. Multimodal vessel visualization of mouse aorta PET/ CT scans. IEEE Transactions on Visualization and Computer Graphics, 15(6):1515– 1522, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roberts:2016:SDU

[RHR16] J. C. Roberts, C. Headleand, and P. D. Ritsos. Sketching designs using the five designsheet methodology. IEEE Transactions on Visualization and Computer Graphics, 22 (1):419–428, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2014:IEI

[RHY14] Donghao Ren, Tobias Hollerer, and Xiaoru Yuan. iVisDesigner: Expressive interactive design of information visualizations. IEEE Transactions on Visualization and Computer Graphics, 20(12):2092– 2101, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876042-
abs.html.

Ramachandra:2011:SPM

[RHZN11] Vikas Ramachandra, Keigo Hirakawa, Matthias Zwicker, and Truong Nguyen. Spatioangular prefiltering for multiview 3D displays. IEEE Transactions on Visualization and Computer Graphics, 17 (5):642–654, May 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Raij:2007:CII

[RJD^{+}07] Andrew B. Raij, Kyle Johnsen, Robert F. Dickerson, Benjamin C. Lok, Marc S. Cohen, Margaret Duerson, Rebecca Rainer Pauly, Amy O. Stevens, Peggy Wagner, and D. Scott Lind. Comparing interpersonal interactions with a virtual human to those with a real human. IEEE Transactions on Visualization and Computer Graphics, 13 (3):443–457, May/June 2007.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rohmer:2017:NEI

[RJG17] Kai Rohmer, Johannes Jendersie, and Thorsten Grosch. Natural environment illumination: Coherent interactive augmented reality for mobile and non-mobile devices. IEEE Transactions on Visualization and Computer Graphics, 23(11):2474– 2484, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/11/08007317-
                                        abs.html.

Rahimian:2017:OCP

[RK17] Pooya Rahimian and Joseph K. Kearney. Optimal camera placement for motion capture systems. IEEE Transactions on Visualization and Computer Graphics, 23(3): 1209–1221, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/03/07778256-
                                        abs.html.

Robb:2013:LVH

[RKA^{+}13] Andrew Robb, Regis Kopper, Ravi Ambani, Farda Qayyum, David Lind, Li-Ming Su, and Benjamin Lok. Leveraging virtual humans to effectively prepare learners for stressful interpersonal experiences. IEEE Transactions on Visualization and Computer Graphics, 19(4): 662–670, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Robb:2016:DVA

[RKC^{+}16] A. Robb, A. Kleinsmith, A. Cordar, C. White, S. Lampotang, A. Wendling, and B. Lok. Do variations in agency indirectly affect behavior with others? An analysis of gaze behavior. IEEE Transactions on Visualization and Computer Graphics, 22 (4):1336–1345, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reininghaus:2011:SSB

[RKG^{+}11] Jan Reininghaus, Natallia Kotava, David Guenther, Jens Kasten, Hans Hagen, and Ingrid Hotz. A scale space based persistence measure for critical points in 2D scalar fields. IEEE Transactions on Visualization and Computer Graphics, 17(12):2045– 2052, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roy:2018:IDV

[RKG^{+}18] L. Roy, P. Kumar, S. Golbabaei, Y. Zhang, and E. Zhang. Interactive design and visualization of branched covering spaces. IEEE Transactions on Visualization and Computer Graphics, 24(1): 843–852, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Resch:2016:SPM

[RKK16] C. Resch, P. Keitler, and G. Klinker. Sticky projections — a model-based approach to interactive shader lamps tracking. IEEE Transactions on Visualization and Computer Graphics, 22(3):1291– 1301, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Riehmann:2019:VTL

[RKKF19] Patrick Riehmann, Dora Kiesel, Martin Kohlhaas, and Bernd Froehlich. Visualizing a thinker’s life. IEEE Transactions on Visualization and Computer Graphics, 25(4):1803–1816, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8334579/.

Ragan:2013:SES

[RKSB13] Eric D. Ragan, Regis Kopper, Philip Schuchardt, and Doug A. Bowman. Studying the effects of stereo, head tracking, and field of regard on a small-scale spatial judgment task. IEEE Transactions on Visualization and Computer Graphics, 19 (5):886–896, May 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rieder:2011:GBR

[RKSH11] Christian Rieder, Tim Kroeger, Christian Schumann, and Horst K. Hahn. GPU-based real-time approximation of the ablation zone for radiofrequency ablation. IEEE Transactions on Visualization and Computer Graphics, 17(12): 1812–1821, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reininghaus:2012:ECC

[RKWH12] Jan Reininghaus, Jens Kasten, Tino Weinkauf, and Ingrid Hotz. Efficient computation of combinatorial feature flow fields. IEEE Transactions on Visualization and Computer Graphics, 18(9):1563– 1573, September 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roy:2019:RFE

[RKZZ19] Lawrence Roy, Prashant Kumar, Yue Zhang, and Eugene Zhang. Robust and fast extraction of 3D symmetric tensor field topology. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1102–1111, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08453873-
                                        abs.html.

Rosenthal:2008:SSE

[RL08] Paul Rosenthal and Lars Linsen. Smooth surface extraction from unstructured point-based volume data using PDEs. IEEE Transactions on Visualization and Computer Graphics, 14(6):1531– 1546, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rind:2013:TDM

[RLA^{+}13] Alexander Rind, Tim Lammarsch, Wolfgang Aigner, Bilal Alsallakh, and Silvia Miksch. TimeBench: A data model and software library for visual analytics of timeoriented data. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2247–2256, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2019:CIC

[RLB19] Donghao Ren, Bongshin Lee, and Matthew Brehmer. Charticulator: Interactive construction of bespoke chart layouts. IEEE Transactions on Visualization and Computer Graphics, 25(1):789– 799, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440827-
abs.html.

Reininghaus:2011:FCV

[RLH11] Jan Reininghaus, Christian Lowen, and Ingrid Hotz. Fast combinatorial vector field topology. IEEE Transactions on Visualization and Computer Graphics, 17(10):1433– 1443, October 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rojas:2019:AVF

[RLK19] J. Rojas, T. Liu, and L. Kavan. Average vector field integration for St. Venant– Kirchhoff deformable models. IEEE Transactions on Visualization and Computer Graphics, 25(8):2529–2539, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2013:FFM

[RLL^{+}13] Bo Ren, Chen-Feng Li, Ming C. Lin, Theodore Kim, and Shi-Min Hu. Flow field modulation. IEEE Transactions on Visualization and Computer Graphics, 19(10): 1708–1719, October 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roman:2010:SDP

[RLM10] Pablo Roman, Maxim Lazarov, and Aditi Majumder. A scalable distributed paradigm for multi-user interaction with tiled rear projection display walls. IEEE Transactions on Visualization and Computer Graphics, 16(6):1623– 1632, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rhee:2011:SBV

[RLNN11] Taehyun Rhee, J. P. Lewis, Ulrich Neumann, and Krishna S. Nayak. Scan-based volume animation driven by locally adaptive articulated registrations. IEEE Transactions on Visualization and Computer Graphics, 17(3): 368–379, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roberts:2019:SBP

[RLS^{+}19] R. C. Roberts, R. S. Laramee, G. A. Smith, P. Brookes, and T. D’Cruze. Smart brushing for parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 25(3):1575–1590, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rong:2011:GAC

[RLW^{+}11] Guodong Rong, Yang Liu, Wenping Wang, Xiaotian Yin, Xianfeng David Gu, and Xiaohu Guo. GPU-assisted computation of centroidal Voronoi tessellation. IEEE Transactions on Visualization and Computer Graphics, 17(3): 345–356, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rufiange:2013:DVD

[RM13] Sebastien Rufiange and Michael J.

McGuffin. DiffAni: Visualizing dynamic graphs with a hybrid of difference maps and animation. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2556–2565, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ruchikachorn:2015:LVA

[RM15] Puripant Ruchikachorn and Klaus Mueller. Learning visualizations by analogy: Promoting visual literacy through visualization morphing. IEEE Transactions on Visualization and Computer Graphics, 21(9):1028–1044, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07061477-abs.html.

Regan:2017:PPR

[RM17] Matthew Regan and Gavin S. P. Miller. The problem of persistence with rotating displays. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1295–1301, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829409-
abs.html.

Ryan:2019:GPA

[RMCW19] Gabriel Ryan, Abigail Mosca, Remco Chang, and Eugene Wu. At a glance: Pixel approximate entropy as a measure of line chart complexity. IEEE Transactions on Visualization and Computer Graphics, 25(1):872– 881, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440849-
abs.html.

Rieck:2012:MDA

[RML12] Bastian Rieck, Hubert Mara, and Heike Leitte. Multivariate data analysis using persistence-based filtering and topological signatures. IEEE Transactions on Visualization and Computer Graphics, 18(12):2382– 2391, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Robertson:2009:EGC

[RMW09] Cindy M. Robertson, Blair MacIntyre, and Bruce N. Walker. An evaluation of graphical context as a means for ameliorating the effects of registration error. IEEE Transactions on Visualization and Computer Graphics, 15(2):179–192, March/April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reach:2019:SEI

[RN19] Andrew McCal Reach and Chris North. Smooth, efficient, and interruptible zooming and panning. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1421–1434, February 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08276573-
abs.html.

Roodaki:2017:SSV

[RNE^{+}17] Hessam Roodaki, Navid Navab, Abouzar Eslami, Christopher Stapleton, and Nassir Navab. SonifEye: Sonification of visual information using physical modeling sound synthesis. IEEE Transactions on Visualization and Computer Graphics, 23(11):2366– 2371, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007327-
abs.html.

Resch:2015:SSA

[RNK^{+}15] Christoph Resch, Hemal Naik, Peter Keitler, Steven Benkhardt, and Gudrun Klinker. On-site semiautomatic calibration and registration of a projectorcamera system using arbitrary objects with known geometry. IEEE Transactions on Visualization and Computer Graphics, 21(11):1211– 1220, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07164353-
abs.html.

Raghuvanshi:2009:EAS

[RNL09] Nikunj Raghuvanshi, Rahul Narain, and Ming C. Lin. Efficient and accurate sound propagation using adaptive rectangular decomposition. IEEE Transactions on Visualization and Computer Graphics, 15(5):789–801, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Robbins:1998:VSV

[Rob98]  K. A. Robbins.  Visual-

ization of scientific video data using KL decomposition. IEEE Transactions on Visualization and Computer Graphics, 4(4):330–343, October/December 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0330.
pdf;   http://www.computer.
org/tvcg/tg1998/v0330abs.
htm.

Rossignac:1999:ECC

[Ros99]  Jarek Rossignac.  Edge-

breaker: Connectivity compression for triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 5(1):47–61, January/March 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0047.
pdf;   http://www.computer.
org/tvcg/tg1999/v0047abs.
htm.

Rossignac:2011:OBL

[Ros11] Jaroslaw (Jarek) Rossignac.

Ordered Boolean List (OBL): Reducing the footprint for evaluating Boolean expressions. IEEE Transactions on Visualization and Computer Graphics, 17(9):1337– 1351, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rosenblum:2013:VRC

[Ros13] Lawrence Rosenblum. The 2012 Virtual Reality Career Award. IEEE Transactions on Visualization and Computer Graphics, 19(4): xv, April 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roth:2013:EDT

[Rot13]  Robert E. Roth.  An

empirically-derived taxonomy of interaction primitives for interactive cartography and geovisualization. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2356–2365, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rosen:2012:SNP

[RP12] Paul Rosen and Voicu Popescu. Simplification of node position data for interactive visualization of dynamic data sets. IEEE Transactions on Visualization and Computer Graphics, 18(9):1537–1548, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rhee:2017:MMR

[RPAC17] Taehyun Rhee, Lohit Petikam, Benjamin Allen, and Andrew Chalmers. MR360: Mixed reality rendering for 360^{∘}; panoramic videos. IEEE Transactions on Visualization and Computer Graphics, 23(4):1379–1388, April 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829404-
abs.html.

Rosen:2008:HQH

[RPHI08] Paul Rosen, Voicu Popescu, Christoph Hoffmann, and Ayhan Irfanoglu. A highquality high-fidelity visualization of the September 11 attack on the World Trade Center. IEEE Transactions on Visualization and Computer Graphics, 14(4):937– 947, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/04/ttg2008040937s.
zip.

Ray:1999:RCA

[RPSC99] Harvey Ray, Hanspeter Pfister, Deborah Silver, and Todd A. Cook. Ray casting architectures for volume visualization. IEEE Transactions on Visualization and Computer Graphics, 5(3):210– 223, July/September 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0210.
pdf;   http://www.computer.

                                        org/tvcg/tg1999/v0210abs.
                                        htm.

Roberts:2013:EGV

[RRD^{+}13] David J. Roberts, John Rae, Tobias W. Duckworth, Carl M. Moore, and Rob Aspin. Estimating the gaze of a virtuality human. IEEE Transactions on Visualization and Computer Graphics, 19 (4):681–690, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Roberts:2018:EVF

[RRJH18] J. C. Roberts, P. D. Ritsos, J. R. Jackson, and C. Headleand. The explanatory visualization framework: An active learning framework for teaching creative computing using explanatory visualizations. IEEE Transactions on Visualization and Computer Graphics, 24(1):791–801, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Reich:2012:ASS

[RS12] Wieland Reich and Gerik Scheuermann. Analysis of streamline separation at infinity using time-discrete Markov chains. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2140–2148, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rocha:2019:DLI

[RSA^{+}19] A. Rocha, J. D. Silva, U. R.

Alim, S. Carpendale, and M. C. Sousa. Decal-lenses: Interactive lenses on surfaces for multivariate visualization. IEEE Transactions on Visualization and Computer Graphics, 25(8):2568–2582, August 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rappoport:1996:VPF

[RSB96] Ari Rappoport, Alla Sheffer, and Michel Bercovier. Volume-preserving free-form solids. IEEE Transactions on Visualization and Computer Graphics, 2(1): 19–27, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0019.
pdf;   http://www.computer.
org/tvcg/tg1996/v0019abs.
htm.

Ragan:2017:AHR

[RSBB17] Eric D. Ragan, Siroberto Scerbo, Felipe Bacim, and Doug A. Bowman. Amplified head rotation in virtual reality and the effects on 3D search, training transfer, and spatial orientation. IEEE Transactions on Visualization and Computer Graphics, 23 (8):1880–1895, August 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07547900-
                                        abs.html.

Rungta:2013:MMA

[RSD^{+}13] Atul Rungta, Brian Summa, Dogan Demir, Peer-Timo Bremer, and Valerio Pascucci. ManyVis: Multiple applications in an integrated visualization environment. IEEE Transactions on Visualization and Computer Graphics, 19(12):2878–2885, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rodgers:2014:DAP

[RSFH14] Peter Rodgers, Gem Stapleton, Jean Flower, and John Howse. Drawing areaproportional Euler diagrams representing up to three sets. IEEE Transactions on Visualization and Computer Graphics, 20(1):1, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rungta:2016:SIS

[RSM^{+}16] A. Rungta, C. Schissler, R. Mehra, C. Malloy, M. Lin, and D. Manocha. SynCo-Pation: Interactive synthesiscoupled sound propagation. IEEE Transactions on Visualization and Computer Graphics, 22(4):1346–1355, ???? 2016. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2018:JGL

[RSOW18] Jing Ren, Jens Schneider, Maks Ovsjanikov, and Peter Wonka. Joint graph layouts for visualizing collections of segmented meshes. IEEE Transactions on Visualization and Computer Graphics, 24(9):2546–2558, September 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08031987-
abs.html.

Rungta:2018:DKI

[RSR^{+}18] Atul Rungta, Carl Schissler, Nicholas Rewkowski, Ravish Mehra, and Dinesh Manocha. Diffraction kernels for interactive sound propagation in dynamic environments. IEEE Transactions on Visualization and Computer Graphics, 24(4):1613–1622, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08307458-
abs.html.

Rubio-Sanchez:2016:CSB

[RSRDS16] M. Rubio-Sanchez, L. Raya, F. Diaz, and A. Sanchez. A comparative study between RadViz and star coordinates. IEEE Transactions on Visualization and Computer Graphics, 22(1):619–628, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rubio-Sanchez:2014:ACI

[RSS14] Manuel Rubio-Sanchez and Alberto Sanchez. Axis calibration for improving data attribute estimation in star coordinates plots. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2013–2022, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875998-
                                        abs.html.

Romero:2008:VVT

[RSSA08] Mario Romero, Jay Summet, John Stasko, and Gregory Abowd. Viz-A-Vis: Toward visualizing video through computer vision. IEEE Transactions on Visualization and Computer Graphics, 14(6):1261–1268, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rossl:2012:SEV

[RT12] Christian Rossl and Holger Theisel. Streamline embedding for 3D vector field exploration. IEEE Transactions on Visualization and Computer Graphics, 18(3): 407–420, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rushmeier:1999:GEI

[Rus99] H. Rushmeier. Guest Editor’s introduction. IEEE Transactions on Visualization and Computer Graphics, 5(2): 97, April/June 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0097.
pdf.

Rautek:2006:CV

[RVG06] Peter Rautek, Ivan Viola, and M. Eduard Gröller. Caricaturistic visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1085–1092, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rushmeier:2005:GEI

[RvWT05] Holly Rushmeier, Jarke J. van Wijk, and Greg Turk. Guest Editors’ introduction: Special section on IEEE visualization. IEEE Transactions on Visualization and Computer Graphics, 11(4):353– 354, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

                                        comp/trans/tg/2005/04/v0353.
                                        pdf.

Reniers:2008:CMC

[RvWT08] Dennie Reniers, Jarke van Wijk, and Alexandru Telea. Computing multiscale curve and surface skeletons of genus 0 shapes using a global importance measure. IEEE Transactions on Visualization and Computer Graphics, 14(2): 355–368, March/April 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rodrigues:2018:NDP

[RW18] N. Rodrigues and D. Weiskopf. Nonlinear dot plots. IEEE Transactions on Visualization and Computer Graphics, 24 (1):616–625, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rathinavel:2018:EDF

[RWBF18] Kishore Rathinavel, Hanpeng Wang, Alex Blate, and Henry Fuchs. An extended depth-at-field volumetric near-eye augmented reality display. IEEE Transactions on Visualization and Computer Graphics, 24(11): 2857–2866, November 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8456852/.

Ribicic:2013:VAS

[RWF^{+}13] Hrvoje Ribičić, Jürgen Waser, Raphael Fuchs, Gunter Bloschl, and Eduard Groller. Visual analysis and steering of flooding simulations. IEEE Transactions on Visualization and Computer Graphics, 19(6): 1062–1075, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rathinavel:2019:VOC

[RWF19] K. Rathinavel, G. Wetzstein, and H. Fuchs. Varifocal occlusion-capable optical see-through augmented reality display based on focustunable optics. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3125–3134, November 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ribicic:2012:SUS

[RWG^{+}12] Hrvoje Ribicic, Juergen Waser, Roman Gurbat, Bernhard Sadransky, and M. Eduard Gröller. Sketching uncertainty into simulations. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2255–2264, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2013:APG

[RYKL13] Zhimin Ren, Hengchin Yeh, Roberta Klatzky, and Ming C.

Lin. Auditory perception of geometry-invariant material properties. IEEE Transactions on Visualization and Computer Graphics, 19(4): 557–566, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2018:RTH

[RYL^{+}18] Bo Ren, Tailing Yuan, Chenfeng Li, Kun Xu, and Shi-Min Hu. Real-time high-fidelity surface flow simulation. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2411–2423, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07964760-
abs.html.

Reis:2008:HQR

[RZHB^{+}08] Gerd Reis, Frank Zeilfelder, Martin Hering-Bertram, Gerald E. Farin, and Hans Hagen. High-quality rendering of quartic spline surfaces on the GPU. IEEE Transactions on Visualization and Computer Graphics, 14(5):1126–1139, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/.

Rossl:2004:RVD

[RZNS04] Christian Rössl, Frank Zeilfelder, Günther Nürnberger, and Hans-Peter Seidel. Reconstruction of volume data with quadratic super splines. IEEE Transactions on Visualization and Computer Graphics, 10(4):397–409, July/ August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/04/v0397.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0397.pdf.

Rhee:2007:LCT

[RZP^{+}07] Seon-Min Rhee, Remo Ziegler, Jiyoung Park, Martin Naef, Markus Gross, and Myoung-Hee Kim. Low-cost telepresence for collaborative virtual environments. IEEE Transactions on Visualization and Computer Graphics, 13(1): 156–166, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rodgers:2012:WPE

[RZP12] Peter Rodgers, Leishi Zhang, and Helen Purchase. Wellformedness properties in Euler diagrams: Which should be used? IEEE Transactions on Visualization and Computer Graphics, 18(7): 1089–1100, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shneiderman:2006:NVS

[SA06] Ben Shneiderman and Aleks Aris. Network visualization by semantic substrates. IEEE Transactions on Visualization and Computer Graphics, 12(5):733–740, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Subramonyam:2019:SMQ

[SA19] Hariharan Subramonyam and Eytan Adar. SmartCues: A multitouch query approach for details-on-demand through dynamically computed overlays. IEEE Transactions on Visualization and Computer Graphics, 25(1): 597–607, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440833-
                                        abs.html.

Strobelt:2016:VVA

[SAB^{+}16] H. Strobelt, B. Alsallakh, J. Botros, B. Peterson, M. Borowsky, H. Pfister, and A. Lex. Vials: Visualizing alternative splicing of genes. IEEE Transactions on Visualization and Computer Graphics, 22(1):399–408, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sud:2008:RTP

[SAC^{+}08] Avneesh Sud, Erik Andersen, Sean Curtis, Ming C. Lin, and Dinesh Manocha. Realtime path planning in dynamic virtual environments using multiagent navigation graphs. IEEE Transactions on Visualization and Computer Graphics, 14(3):526– 538, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/03/ttg2008030526s.
zip.

Sherbondy:2005:ECB

[SAM^{+}05] Anthony Sherbondy, David Akers, Rachel Mackenzie, Robert Dougherty, and Brian Wandell. Exploring connectivity of the brain’s white matter with dynamic queries. IEEE Transactions on Visualization and Computer Graphics, 11(4):419– 430, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sharp:2007:PBS

[SAM^{+}07] Richard Sharp, Jacob Adams, Raghu Machiraju, Robert Lee, and Robert Crane. Physics-based subsurface visualization of human tissue. IEEE Transactions on Visualization and Computer Graphics, 13(3):620–629, May/June 2007. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sheehy:1996:SDM

[SAR96] Damian J. Sheehy, Cecil G. Armstrong, and Desmond J. Robinson. Shape description by medial surface construction. IEEE Transactions on Visualization and Computer Graphics, 2(1): 62–72, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0062.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0062abs.
                                        htm.

Stark:2005:BPI

[SAS05] Michael M. Stark, James Arvo, and Brian Smits. Barycentric parameterizations for isotropic BRDFs. IEEE Transactions on Visualization and Computer Graphics, 11(2):126–138, March/ April 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Simonetto:2016:SAB

[SAS16] P. Simonetto, D. Archambault, and C. Scheidegger. A simple approach for boundary improvement of Euler diagrams. IEEE Transactions on Visualization and Computer Graphics, 22(1):678–687, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schulz:2016:EVP

[SASS16] Hans-Jörg Schulz, Marco Angelini, Giuseppe Santucci, and Heidrun Schumann. An enhanced visualization process model for incremental visualization. IEEE Transactions on Visualization and Computer Graphics, 22(7):1830– 1842, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Satava:2013:KSV

[Sat13] Richard Satava. Keynote speaker: Virtual reality: Current uses in medical simulation and future. IEEE Transactions on Visualization and Computer Graphics, 19 (4):xii, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:2004:FVM

[SB04] Yuzhong Shen and Kenneth E. Barner. Fuzzy vector median-based surface smoothing. IEEE Transactions on Visualization and Computer Graphics, 10(3):252– 265, May/June 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0252abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0252.htm;   http://csdl.

                                        computer.org/dl/trans/tg/
                                        2004/03/v0252.pdf.

Sohn:2006:TVC

[SB06] Bong-Soo Sohn and Chandrajit Bajaj. Time-varying contour topology. IEEE Transactions on Visualization and Computer Graphics, 12(1):14– 25, January/February 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stinson:2014:FTA

[SB14] Cheryl Stinson and Doug A. Bowman. Feasibility of training athletes for highpressure situations using virtual reality. IEEE Transactions on Visualization and Computer Graphics, 20(4): 606–615, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stoppel:2017:VPI

[SB17] Sergej Stoppel and Stefan Bruckner. Vol^{2} velle: Printable interactive volume visualization. IEEE Transactions on Visualization and Computer Graphics, 23(1): 861–870, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shi:2018:MVN

[SBB^{+}18] Yang Shi, Chris Bryan, Sridatt Bhamidipati, Ying Zhao, Yaoxue Zhang, and Kwan-Liu Ma. MeetingVis: Visual narratives to assist in recalling meeting context and content. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1918–1929, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/08320320-
abs.html.

Sbert:1997:ECR

[Sbe97] M. Sbert. Error and complexity of random walk Monte Carlo radiosity. IEEE Transactions on Visualization and Computer Graphics, 3(1):23– 38, January/March 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0023.
pdf;   http://www.computer.
org/tvcg/tg1997/v0023abs.
htm.

Stengel:2015:TVF

[SBE^{+}15] M. Stengel, P. Bauszat, M. Eisemann, E. Eisemann, and M. Magnor. Temporal video filtering and exposure control for perceptual motion blur. IEEE Transactions on Visualization and Computer Graphics, 21(5):663–671, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-

Steinicke:2011:CBP

[SBHW11] Frank Steinicke, Gerd Bruder, Klaus Hinrichs, and Pete Willemsen. Change blindness phenomena for virtual reality display systems. IEEE Transactions on Visualization and Computer Graphics, 17(9): 1223–1233, September 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steinicke:2010:EDT

[SBJ^{+}10] Frank Steinicke, Gerd Bruder, Jason Jerald, Harald Frenz, and Markus Lappe. Estimation of detection thresholds for redirected walking techniques. IEEE Transactions on Visualization and Computer Graphics, 16(1):17–27, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steinicke:2011:NPP

[SBK^{+}11] F. Steinicke, G. Bruder, S. Kuhl, P. Willemsen, M. Lappe, and K. Hinrichs. Natural perspective projections for head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 17(7):888–899, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schroeder:2006:MFV

[SBM^{+}06] William J. Schroeder, François Bertel, Mathieu Malaterre, David Thompson, Philippe P. Pebay, Robert O’Bara, and Saurabh Tendulkar. Methods and framework for visualizing higher-order finite elements. IEEE Transactions on Visualization and Computer Graphics, 12(4):446– 460, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schatzschneider:2016:WTC

[SBS16] C. Schatzschneider, G. Bruder, and F. Steinicke. Who turned the clock? Effects of manipulated zeitgebers, cognitive load and immersion on time estimation. IEEE Transactions on Visualization and Computer Graphics, 22(4): 1387–1395, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sereda:2006:VBV

[SBSG06] Petr Sereda, Anna Vilanova Bartroli, Iwo W. O. Serlie, and Frans A. Gerritsen. Visualization of boundaries in volumetric data sets using LH histograms. IEEE Transactions on Visualization and Computer Graphics, 12(2): 208–218, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schafhitzel:2011:VEI

[SBV^{+}11] Tobias Schafhitzel, Kudret Baysal, Mikael Vaaraniemi, Ulrich Rist, and Daniel Weiskopf. Visualizing the evolution and interaction of vortices and shear layers in timedependent 3D flow. IEEE Transactions on Visualization and Computer Graphics, 17 (4):412–425, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stevens:2017:HSE

[SBW17] Andrew H. Stevens, Thomas Butkiewicz, and Colin Ware. Hairy slices: Evaluating the perceptual effectiveness of cutting plane glyphs for 3D vector fields. IEEE Transactions on Visualization and Computer Graphics, 23(1): 990–999, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Su:2015:DST

[SC15] Yan-Jen Su and Yung-Yu Chuang. Disambiguating stereoscopic transparency using a thaumatrope approach. IEEE Transactions on Visualization and Computer Graphics, 21(8):959–969, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://www.computer.
                                        org/csdl/trans/tg/2015/
                                        08/07055322-abs.html.

Sarikaya:2019:WDW

[SCB^{+}19] Alper Sarikaya, Michael Correll, Lyn Bartram, Melanie Tory, and Danyel Fisher. What do we talk about when we talk about dashboards? IEEE Transactions on Visualization and Computer Graphics, 25(1): 682–692, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08443395-
abs.html.

Schmalstieg:2013:VRT

[Sch13]  Dieter Schmalstieg.  The

2012 Virtual Reality Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 19(4):xvi, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steffen:2008:ISI

[SCKR08] Michael Steffen, Sean Curtis, Robert M. Kirby, and Jennifer K. Ryan. Investigation of smoothness-increasing accuracy-conserving filters for improving streamline integration through discontinuous fields. IEEE Transactions on Visualization and Computer Graphics, 14(3):680– 692, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steinhurst:2008:RPM

[SCL08] Joshua Steinhurst, Greg Coombe, and Anselmo Lastra. Reducing photon-mapping bandwidth by query reordering. IEEE Transactions on Visualization and Computer Graphics, 14(1):13–24, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shi:2012:RVR

[SCL^{+}12] Conglei Shi, Weiwei Cui, Shixia Liu, Panpan Xu, Wei Chen, and Huamin Qu. RankExplorer: Visualization of ranking changes in large time series data. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2669–2678, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Smith:2006:FSS

[SCM^{+}06] Greg Smith, Mary Czerwinski, Brian Meyers, Daniel Robbins, George Robertson, and Desney S. Tan. FacetMap: a scalable search and browse visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):797–804, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sorkine:2005:GAB

[SCOIT05] Olga Sorkine, Daniel Cohen-Or, Dror Irony, and Sivan Toledo. Geometry-aware bases for shape approximation. IEEE Transactions on Visualization and Computer Graphics, 11(2):171– 180, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sadowsky:2006:PTR

[SCT06] Ofri Sadowsky, Jonathan D.

Cohen, and Russell H. Taylor. Projected tetrahedra revisited: a barycentric formulation applied to digital radiograph reconstruction using higher-order attenuation functions. IEEE Transactions on Visualization and Computer Graphics, 12(4):461– 473, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sanderson:2010:ARP

[SCT^{+}10] Allen Sanderson, Guoning Chen, Xavier Tricoche, David Pugmire, Scott Kruger, and Joshua Breslau. Analysis of recurrent patterns in toroidal magnetic fields. IEEE Transactions on Visualization and Computer Graphics, 16(6):1431–1440, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2016:MMC

[SCYW16] F. Sun, Y. K. Choi, Y. Yu, and W. Wang. Medial meshes — a compact and accurate representation of medial axis transform. IEEE Transactions on Visualization and Computer Graphics, 22(3): 1278–1290, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sen:2011:CRR

[SD11] Pradeep Sen and Soheil Darabi. Compressive rendering: a rendering application of compressed sensing. IEEE Transactions on Visualization and Computer Graphics, 17 (4):487–499, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Spritzer:2012:DEM

[SD12] Andre Suslik Spritzer and Carla Maria Dal Sasso Freitas. Design and evaluation of MagnetViz — a graph visualization tool. IEEE Transactions on Visualization and Computer Graphics, 18(5): 822–835, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Srinivasan:2019:AVI

[SDES19] Arjun Srinivasan, Steven M. Drucker, Alex Endert, and John Stasko. Augmenting visualizations with interactive data facts to facilitate interpretation and communication. IEEE Transactions on Visualization and Computer Graphics, 25(1):672– 681, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440860-
abs.html.

Shafii:2012:TES

[SDHH12] Sohail Shafii, Scott E. Dillard, Mario Hlawitschka, and Bernd Hamann. The topological effects of smoothing. IEEE Transactions on Visualization and Computer Graphics, 18(1):160–172, January 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stahnke:2016:PPI

[SDMT16] J. Stahnke, M. Dork, B. Muller, and A. Thom. Probing projections: Interaction techniques for interpreting arrangements and errors of dimensionality reductions. IEEE Transactions on Visualization and Computer Graphics, 22(1): 629–638, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Slingsby:2009:CHL

[SDW09] Aidan Slingsby, Jason Dykes, and Jo Wood. Configuring hierarchical layouts to address research questions. IEEE Transactions on Visualization and Computer Graphics, 15(6):977–984, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Slingsby:2011:EUG

[SDW11] Aidan Slingsby, Jason Dykes, and Jo Wood. Exploring uncertainty in geodemographics with interactive graphics. IEEE Transactions on Visualization and Computer Graphics, 17(12):2545– 2554, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sakhaee:2017:SDV

[SE17] Elham Sakhaee and Alireza Entezari. A statistical direct volume rendering framework for visualization of uncertain data. IEEE Transactions on Visualization and Computer Graphics, 23(12): 2509–2520, December 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/12/07778257-
                                        abs.html.

Schmidtke:2018:CBV

[SE18] Robert Schmidtke and Kenny Erleben. Chunked bounding volume hierarchies for fast digital prototyping using volumetric meshes. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3044–3057, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8219711/.

Svakhine:2009:IID

[SEA09] Nikolai A. Svakhine, David S.

Ebert, and William M. Andrews. Illustration-inspired depth enhanced volumetric medical visualization. IEEE Transactions on Visualization and Computer Graphics, 15 (1):77–86, January/February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stoppel:2019:FVI

[SEB19] Sergej Stoppel, Magnus Paulson Erga, and Stefan Bruckner. Firefly: Virtual illumination drones for interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1204–1213, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440109-
abs.html.

Saket:2019:TBE

[SED19] B. Saket, A. Endert, and Ç. Demiralp. Task-based effectiveness of basic visualizations. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2505–2512, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stroila:2008:CAR

[SEH08] Matei Stroila, Elmar Eisemann, and John Hart. Clip art rendering of smooth isosurfaces. IEEE Transactions on Visualization and Computer Graphics, 14(1):135– 145, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Selver:2015:EBB

[Sel15] M. Alper Selver. Exploring brushlet based 3D textures in transfer function specification for direct volume rendering of abdominal organs. IEEE Transactions on Visualization and Computer Graphics, 21 (2):174–187, February 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/02/06908014-
                                        abs.html.

Sequin:2012:KST

[Seq12] Carlo H. Sequin. Keynote speaker: Taking the “virtual” out of virtual reality. IEEE Transactions on Visualization and Computer Graphics, 18(4):xi, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stylianou:2004:CLS

[SF04] Georgios Stylianou and Gerald Farin. Crest lines for surface segmentation and flattening. IEEE Transactions on Visualization and Computer Graphics, 10(5):536–544, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/05/v0536.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0536.pdf.

Schollmeyer:2014:DIR

[SF14] Andre Schollmeyer and Bernd Froehlich. Direct isosurface ray casting of nurbs-based isogeometric analysis. IEEE Transactions on Visualization and Computer Graphics, 20 (9):1, September 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schollmeyer:2019:EAA

[SF19] A. Schollmeyer and B. Froehlich.

Efficient and anti-aliased trimming for rendering large NURBS models. IEEE Transactions on Visualization and Computer Graphics, 25(3): 1489–1498, March 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sanchez:2015:STT

[SFA^{+}15] Mathieu Sanchez, Oleg Fryazinov, Valery Adzhiev, Peter Comninos, and Alexander Pasko. Space–time transfinite interpolation of volumetric material properties. IEEE Transactions on Visualization and Computer Graphics, 21 (2):278–288, February 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/02/06894217-
                                        abs.html.

Schindler:2012:LCS

[SFB^{+}12] Benjamin Schindler, Raphael Fuchs, Stefan Barp, Jürgen Waser, Armin Pobitzer, Robert Carnecky, Krešimir Matković, and Ronald Peikert. Lagrangian coherent structures for design analysis of revolving doors. IEEE Transactions on Visualization and Computer Graphics, 18(12):2159–2168, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schindler:2009:PCS

[SFBP09] Benjamin Schindler, Raphael Fuchs, John Biddiscombe, and Ronald Peikert. Predictorcorrector schemes for visualization of smoothed particle hydrodynamics data. IEEE Transactions on Visualization and Computer Graphics, 15(6):1243–1250, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Summet:2007:SEB

[SFC^{+}07] Jay Summet, Matthew Flagg, Tat-Jen Cham, James M. Rehg, and Rahul Sukthankar. Shadow elimination and blinding light suppression for interactive projected displays. IEEE Transactions on Visualization and Computer Graphics, 13(3):508–517, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saroul:2006:DPF

[SFH06] Laurent Saroul, Oscar Figueiredo, and Roger D. Hersch. Distance preserving flattening of surface sections. IEEE Transactions on Visualization and Computer Graphics, 12(1):26– 35, January/February 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steed:2016:AWE

[SFL^{+}16] Anthony Steed, Sebastian Friston, Maria Murcía Lopez, Jason Drummond, Ye Pan, and David Swapp. An ‘in the wild’ experiment on presence and embodiment using consumer virtual reality equipment. IEEE Transactions on Visualization and Computer Graphics, 22(4):1406– 1414, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sedlmair:2012:RVA

[SFMB12] Michael Sedlmair, Annika Frank, Tamara Munzner, and Andreas Butz. RelEx: Visualization for actively changing overlay network specifications. IEEE Transactions on Visualization and Computer Graphics, 18(12):2729– 2738, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schmidt:2019:PPW

[SFP^{+}19] J. Schmidt, D. Fleischmann, B. Preim, N. Brändle, and G. Mistelbauer. Popup-plots: Warping temporal data visualization. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2443–2457, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Suma:2010:ECE

[SFR^{+}10] Evan A. Suma, Samantha L. Finkelstein, Myra Reid, Sabarish V. Babu, Amy C. Ulinski, and Larry F. Hodges. Evaluation of the cognitive effects of travel technique in complex real and virtual environments. IEEE Transactions on Visualization and Computer Graphics, 16(4):690– 702, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sudarsky:1999:DSO

[SG99] Oded Sudarsky and Craig Gotsman. Dynamic scene occlusion culling. IEEE Transactions on Visualization and Computer Graphics, 5(1):13– 29, January/March 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0013.
pdf;   http://www.computer.
org/tvcg/tg1999/v0013abs.
htm.

Shaffer:2005:MRM

[SG05] Eric Shaffer and Michael Garland. A multiresolution representation for massive meshes. IEEE Transactions on Visualization and Computer Graphics, 11(2):139– 148, March/April 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Selver:2009:STF

[SG09] M. Alper Selver and Cüneyt Güzelis. Semiautomatic transfer function initialization for abdominal visualization using self-generating hierarchical radial basis function networks. IEEE Transactions on Visualization and Computer Graphics, 15(3):395– 409, May/June 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stitz:2016:TVM

[SGAS16] Holger Stitz, Samuel Gratzl, Wolfgang Aigner, and Marc Streit. ThermalPlot: Visualizing multi-attribute timeseries data using a thermal metaphor. IEEE Transactions on Visualization and Computer Graphics, 22(12):2594– 2607, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07369992-
abs.html.

Schmidt:2013:VVA

[SGB13] Johanna Schmidt, M. Eduard Groller, and Stefan Bruckner. VAICo: Visual analysis for image comparison. IEEE Transactions on Visualization and Computer Graphics, 19(12):2090–2099, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Strobelt:2019:SVV

[SGB^{+}19] Hendrik Strobelt, Sebastian Gehrmann, Michael Behrisch, Adam Perer, Hanspeter Pfister, and Alexander M. Rush. Seq2seq-Vis: A visual debugging tool for sequenceto-sequence models. IEEE Transactions on Visualization and Computer Graphics, 25 (1):353–363, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08494828-
abs.html.

Sra:2018:OPG

[SGJM18] Misha Sra, Sergio Garrido-Jurado, and Pattie Maes. Oasis: Procedurally generated social virtual spaces from 3D scanned real spaces. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3174–3187, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8067498/.

Sharko:2008:VRA

[SGM08] John Sharko, Georges Grinstein, and Kenneth A. Marx. Vectorized Radviz and its application to multiple cluster datasets. IEEE Transactions on Visualization and Computer Graphics, 14(6):1444– 1427, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Suter:2011:IMT

[SGM^{+}11] Susanne K. Suter, Jose A. Iglesias Guitian, Fabio Marton, Marco Agus, Andreas Elsener, Christoph P. E.

Zollikofer, M. Gopi, Enrico Gobbetti, and Renato Pajarola. Interactive multiscale tensor reconstruction for multiresolution volume visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):2135– 2143, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stitz:2019:KPB

[SGP^{+}19] Holger Stitz, Samuel Gratzl, Harald Piringer, Thomas Zichner, and Marc Streit. KnowledgePearls: Provenancebased visualization retrieval. IEEE Transactions on Visualization and Computer Graphics, 25(1):120–130, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL https://www.computer.
org/csdl/trans/tg/2019/
01/08440831.pdf.

Strobelt:2018:LTV

[SGPR18] H. Strobelt, S. Gehrmann, H. Pfister, and A. M. Rush. LSTMVis: A tool for visual analysis of hidden state dynamics in recurrent neural networks. IEEE Transactions on Visualization and Computer Graphics, 24(1): 667–676, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Samaraweera:2016:HTL

[SGQ16] G. Samaraweera, R. Guo, and J. Quarles. Head tracking latency in virtual environments revisited: Do users with multiple sclerosis notice latency less? IEEE Transactions on Visualization and Computer Graphics, 22(5):1630– 1636, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Silva:2006:GES

[SGR06] Claudio T. Silva, Eduard Gröller, and Holly Rushmeier. Guest editorial: Special section on Visualization 2005. IEEE Transactions on Visualization and Computer Graphics, 12(4):419– 420, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/04/v0419.
pdf.

Schloss:2019:MCM

[SGS^{+}19] Karen B. Schloss, Connor C.

Gramazio, Allison T. Silverman, Madeline L. Parker, and Audrey S. Wang. Mapping color to meaning in colormap data visualizations. IEEE Transactions on Visualization and Computer Graphics, 25 (1):810–819, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08454346-
                                        abs.html.

Seo:2000:CFA

[SH00a] Y. Seo and K. S. Hong. Calibration-free augmented reality in perspective. IEEE Transactions on Visualization and Computer Graphics, 6(4):346–??, October/ December 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0346.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0346abs.
                                        htm.

Sutton:2000:AIE

[SH00b] P. M. Sutton and C. D. Hansen. Accelerated isosurface extraction in timevarying fields. IEEE Transactions on Visualization and Computer Graphics, 6(2):98– 107, April/June 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0098.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0098abs.
                                        htm.

Segel:2010:NVT

[SH10] Edward Segel and Jeffrey Heer. Narrative visualization: Telling stories with data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1139–1148, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Spillmann:2012:RIC

[SH12] Jonas Spillmann and Matthias Harders. Robust interactive collision handling between tools and thin volumetric objects. IEEE Transactions on Visualization and Computer Graphics, 18(8): 1241–1254, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sedlmair:2014:VPS

[SHB^{+}14] Michael Sedlmair, Christoph Heinzl, Stefan Bruckner, Harald Piringer, and Torsten Moller. Visual parameter space analysis: A conceptual framework. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2161–2170, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876043-
abs.html.

Smelyanskiy:2009:MHF

[SHC^{+}09] Mikhail Smelyanskiy, David Holmes, Jatin Chhugani, Alan Larson, Douglas M. Carmean, Dennis Hanson, Pradeep Dubey, Kurt Augustine, Daehyun Kim, Alan Kyker, Victor W. Lee, Anthony D. Nguyen, Larry Seiler, and Richard Robb. Mapping highfidelity volume rendering for medical imaging to CPU, GPU and many-core architectures. IEEE Transactions on Visualization and Computer Graphics, 15(6):1563– 1570, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Selassie:2011:DEB

[SHH11] David Selassie, Brandon Heller, and Jeffrey Heer. Divided edge bundling for directional network data. IEEE Transactions on Visualization and Computer Graphics, 17(12):2354–2363, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schlemmer:2007:MIA

[SHM^{+}07] Michael Schlemmer, Manuel Heringer, Florian Morr, Ingrid Hotz, Martin Hering-Bertram, Christoph Garth, Wolfgang Kollmann, Bernd Hamann, and Hans Hagen. Moment invariants for the analysis of 2D flow fields. IEEE Transactions on Visualization and Computer Graphics, 13(6):1743–1750, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saad:2010:EVS

[SHM10] Ahmed Saad, Ghassan Hamarneh, and Torsten Moller. Exploration and visualization of segmentation uncertainty using shape and appearance prior information. IEEE Transactions on Visualization and Computer Graphics, 16(6):1366–1375, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2011:RTB

[SHR^{+}11] Xin Sun, Qiming Hou, Zhong Ren, Kun Zhou, and Baining Guo. Radiance transfer biclustering for real-time all-frequency biscale rendering. IEEE Transactions on Visualization and Computer Graphics, 17(1):64–73, January 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schulz:2011:DSI

[SHS11a] Hans-Jörg Schulz, Steffen Hadlak, and Heidrun Schumann. The design space of implicit hierarchy visualization: a survey. IEEE Transactions on Visualization and Computer Graphics, 17(4): 393–411, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schulz:2011:PBV

[SHS11b] Hans-Jörg Schulz, Steffen Hadlak, and Heidrun Schumann. Point-based visualization for large hierarchies. IEEE Transactions on Visualization and Computer Graphics, 17(5):598–611, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schmitz:2018:YSM

[SHV^{+}18] Patric Schmitz, Julian Hildebrandt, Andre Calero Valdez, Leif Kobbelt, and Martina Ziefle. You spin my head right round: Threshold of limited immersion for rotation gains in redirected walking. IEEE Transactions on Visualization and Computer Graphics, 24(4):1623–1632, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08260943-
abs.html.

Scheepens:2016:VSA

[SHVV16] R. Scheepens, C. Hurter, H. Van De Wetering, and J. J. Van Wijk. Visualization, selection, and analysis of traffic flows. IEEE Transactions on Visualization and Computer Graphics, 22(1):379–388, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sillion:1995:UHA

[Sil95] François X. Sillion. A unified hierarchical algorithm for global illumination with scattering volumes and object clusters. IEEE Transactions on Visualization and Computer Graphics, 1(3):240– 254, September 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0240.
pdf;   http://www.computer.
org/tvcg/tg1995/v0240abs.
htm.

Silva:2017:IVGa

[Sil17a] Cláudio T. Silva. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 23(1):xvi, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Silva:2017:IVGb

[Sil17b] Claudio T. Silva. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 23(4):vii, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07876894.
pdf.

Silva:2018:IVGa

[Sil18a] C. T. Silva. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 24 (1):xvi, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Silva:2018:IVGb

[Sil18b] Claudio T. Silva. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 24(4):viii, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/04/08315167.
                                        pdf.

Silva:2019:IVG

[Sil19] Claudio T. Silva. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 25(1):xviii, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08570934.
                                        pdf.

Simon:2007:UMI

[Sim07]  Andreas Simon.  Usabil-

ity of multiviewpoint images for spatial interaction in projection-based display systems. IEEE Transactions on Visualization and Computer Graphics, 13(1):26–33, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shi:2006:IES

[SJ06] Qingmin Shi and Joseph JaJa.

Isosurface extraction and spatial filtering using persistent octree (POT). IEEE Transactions on Visualization and Computer Graphics, 12(5):1283–1290, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Spencer:2009:HPM

[SJ09] Ben Spencer and Mark W.

Jones. Hierarchical photon mapping. IEEE Transactions on Visualization and Computer Graphics, 15(1):49–61, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shan:2010:GOL

[SJB10] Qi Shan, Jiaya Jia, and Michael S. Brown. Globally optimized linear windowed tone mapping. IEEE Transactions on Visualization and Computer Graphics, 16(4): 663–675, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sisneros:2007:MLC

[SJH^{+}07] Robert Sisneros, Chad Jones, Jian Huang, Jinzhu Gao, Byung-Hoon Park, and Nagiza Samatova. A multi-level cache model for run-time optimization of remote visualization. IEEE Transactions on Visualization and Computer Graphics, 13(5):991– 1003, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sagrista:2017:TAI

[SJJ^{+}17] Antoni Sagristà, Stefan Jordan, Andreas Just, Fabio Dias, Luis Gustavo Nonato, and Filip Sadlo. Topological analysis of inertial dynamics. IEEE Transactions on Visualization and Computer Graphics, 23(1):950–959, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Swan:2007:EDJ

[SJK^{+}07] J. Edward Swan II, Adam Jones, Eric Kolstad, Mark A. Livingston, and Harvey S. Smallman. Egocentric depth judgments in optical, seethrough augmented reality. IEEE Transactions on Visualization and Computer Graphics, 13(3):429–442, May/June 2007. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sunkavalli:2012:VSC

[SJK^{+}12] Kalyan Sunkavalli, Neel Joshi, Sing Bing Kang, Michael F. Cohen, and Hanspeter Pfister. Video snapshots: Creating high-quality images from video clips. IEEE Transactions on Visualization and Computer Graphics, 18(11): 1868–1879, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saha:2018:FOS

[SJL^{+}18a] Punam K. Saha, Dakai Jin, Yinxiao Liu, Gary E. Christensen, and Cheng Chen. Fuzzy object skeletonization: Theory, algorithms, and applications. IEEE Transactions on Visualization and Computer Graphics, 24(8):2298– 2314, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/08007307-
abs.html.

Stein:2018:BIP

[SJL^{+}18b] M. Stein, H. Janetzko, A. Lamprecht, T. Breitkreutz, P. Zimmermann, B. Goldlücke, T. Schreck, G. Andrienko, M. Grossniklaus, and D. A. Keim. Bring it to the pitch: Combining video and movement data to enhance team sport analysis. IEEE Transactions on Visualization and Computer Graphics, 24(1):13–22, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2014:BRM

[SJM14] Xianfang Sun, Zhongping Ji, and Weiyin Ma. Bas-relief modeling from normal images with intuitive styles. IEEE Transactions on Visualization and Computer Graphics, 20 (5):675–685, May 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sagrista:2019:GSN

[SJMS19] Antoni Sagrista, Stefan Jordan, Thomas Muller, and Filip Sadlo. Gaia Sky: Navigating the Gaia catalog. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1070–1079, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440086-
                                        abs.html.

Schaefer:2007:MDC

[SJW07] Scott Schaefer, Tao Ju, and Joe Warren. Manifold dual contouring. IEEE Transactions on Visualization and Computer Graphics, 13(3): 610–619, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:1998:NLI

[SK98] Han-Wei Shen and D. L. Kao.

A new line integral convolution algorithm for visualizing time-varying flow fields. IEEE Transactions on Visualization and Computer Graphics, 4 (2):98–108, April/June 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0098.
pdf;   http://www.computer.
org/tvcg/tg1998/v0098abs.
htm.

Sramek:1999:AFV

[SK99] Milos Sramek and Arie E.

Kaufman. Alias-free voxelization of geometric objects. IEEE Transactions on Visualization and Computer Graphics, 5(3):251–267, July/September 1999. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0251.
pdf;   http://www.computer.
org/tvcg/tg1999/v0251abs.
htm.

Sramek:2000:FRT

[SK00] M. Sramek and A. Kaufman.

Fast ray-tracing of rectilinear volume data using distance transforms. IEEE Transactions on Visualization and Computer Graphics, 6(3):236– 252, July/September 2000. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2000/pdf/v0236.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2000/v0236abs.
                                        htm.

Schultz:2010:SGS

[SK10] Thomas Schultz and Gordon L. Kindlmann. Superquadric glyphs for symmetric second-order tensors. IEEE Transactions on Visualization and Computer Graphics, 16(6):1595–1604, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schultz:2013:OBS

[SK13] Thomas Schultz and Gordon L. Kindlmann. Openbox spectral clustering: Applications to medical image analysis. IEEE Transactions on Visualization and Computer Graphics, 19(12):2100– 2108, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sifakis:2015:GEI

[SK15] Eftychios Sifakis and Vladlen Koltun. Guest Editor’s introduction: Special section on the ACM SIG-GRAPH/Eurographics Symposium on Computer Animation (SCA). IEEE Transactions on Visualization and Computer Graphics, 21(10): 1101–1102, October 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07230338.
pdf.

Sander:2016:GEI

[SK16a] Pedro V. Sander and John Keyser. Guest Editor’s introduction to the special section on the ACM Symposium on Interactive 3D Graphics and Games (I3D). IEEE Transactions on Visualization and Computer Graphics, 22 (10):2214, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schroeder:2016:VSA

[SK16b] D. Schroeder and D. F.

Keefe. Visualization-by-sketching: An Artist’s interface for creating multivariate time-varying data visualizations. IEEE Transactions on Visualization and Computer Graphics, 22(1):877–885, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sacha:2018:SGE

[SKB^{+}18] D. Sacha, M. Kraus, J. Bernard, M. Behrisch, T. Schreck, Y. Asano, and D. A. Keim.

SOMFlow: Guided exploratory cluster analysis with self-organizing maps and analytic provenance. IEEE Transactions on Visualization and Computer Graphics, 24 (1):120–130, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saket:2017:VDI

[SKBE17] Bahador Saket, Hannah Kim, Eli T. Brown, and Alex Endert. Visualization by demonstration: An interaction paradigm for visual data exploration. IEEE Transactions on Visualization and Computer Graphics, 23(1): 331–340, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Serrano:2019:MPR

[SKC^{+}19] Ana Serrano, Incheol Kim, Zhili Chen, Stephen Di-Verdi, Diego Gutierrez, Aaron Hertzmann, and Belen Masia. Motion parallax for 360 RGBD video. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1817–1827, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8661657/.

Stotko:2019:SLS

[SKH^{+}19] Patrick Stotko, Stefan Krumpen, Matthias B. Hullin, Michael Weinmann, and Reinhard Klein. SLAMCast: Largescale, real-time 3D reconstruction and streaming for immersive multi-client live telepresence. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2102–2112, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643537/.

Salama:2006:HLU

[SKK06] Christof Rezk Salama, Maik Keller, and Peter Kohlmann. High-level user interfaces for transfer function design with semantics. IEEE Transactions on Visualization and Computer Graphics, 12(5):1021–1028, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schroeder:2014:TCM

[SKK^{+}14] David Schroeder, Fedor Korsakov, Carissa Mai-Ping Knipe, Lauren Thorson, Arin M. Ellingson, David Nuckley, John Carlis, and Daniel F. Keefe. Trendcentric motion visualization: Designing and applying a new strategy for analyzing scientific motion collections. IEEE Transactions on Visualization and Computer Graphics, 20(12):2644– 2653, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875981-
                                        abs.html.

Sacha:2019:VOV

[SKKC19] Dominik Sacha, Matthias Kraus, Daniel A. Keim, and Min Chen. VIS4ML: An ontology for visual analytics assisted machine learning. IEEE Transactions on Visualization and Computer Graphics, 25 (1):385–395, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440124-
                                        abs.html.

Stolper:2014:GSG

[SKL^{+}14] Charles D. Stolper, Minsuk Kahng, Zhiyuan Lin, Florian Foerster, Aakash Goel, John Stasko, and Duen Horng Chau. GLO-STIX: Graphlevel operations for specifying techniques and interactive eXploration. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2320–2328, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875969-
                                        abs.html.

Szirmay-Kalos:2011:PIR

[SKLU^{+}11] Laszlo Szirmay-Kalos, Gabor Liktor, Tamas Umenhoffer, Balazs Toth, Shree Kumar, and Glenn Lupton. Parallel iteration to the radiative transport in inhomogeneous media with bootstrapping. IEEE Transactions on Visualization and Computer Graphics, 17 (2):146–158, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sicat:2014:SPV

[SKMH14] Ronell Sicat, Jens Kruger, Torsten Moller, and Markus Hadwiger. Sparse PDF volumes for consistent multiresolution volume rendering. IEEE Transactions on Visualization and Computer Graphics, 20(12):2417– 2426, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876002-
abs.html.

Scheuermann:1998:VNV

[SKMR98] G. Scheuermann, H. Krüger, M. Menzel, and A. P. Rockwood. Visualizing nonlinear vector field topology. IEEE Transactions on Visualization and Computer Graphics, 4 (2):109–116, April/June 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0109.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0109abs.
                                        htm.

Shah:2007:CMI

[SKP07] Musawir A. Shah, Jaakko Konttinen, and Sumanta Pattanaik. Caustics mapping: an image-space technique for real-time caustics. IEEE Transactions on Visualization and Computer Graphics, 13(2):272–280, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schlegel:2012:IDN

[SKS12] Steven Schlegel, Nico Korn, and Gerik Scheuermann. On the interpolation of data with normally distributed uncertainty for visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2305–2314, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sips:2012:VAA

[SKU^{+}12] Mike Sips, Patrick Kothur, Andrea Unger, Hans-Christian Hege, and Doris Dransch. A visual analytics approach to multiscale exploration of environmental time series. IEEE Transactions on Visualization and Computer Graphics, 18(12):2899–2907, December 2012. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steffen:2011:SMT

[SKW^{+}11] Wolfgang Steffen, Nicholas Koning, Stephan Wenger, Christophe Morisset, and Marcus Magnor. Shape: a 3D modeling tool for astrophysics. IEEE Transactions on Visualization and Computer Graphics, 17(4): 454–465, April 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shum:2012:SMC

[SKY12] Hubert P. H. Shum, Taku Komura, and Shuntaro Yamazaki. Simulating multiple character interactions with collaborative and adversarial goals. IEEE Transactions on Visualization and Computer Graphics, 18(5):741–752, May 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Seo:2014:GIG

[SKYS14] Jinwook Seo, Bohyoung Kim, Jihye Yun, and Hyunjoo Song. GazeVis: Interactive 3D gaze visualization for contiguous cross-sectional medical images. IEEE Transactions on Visualization and Computer Graphics, 20(5):726–739, May 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Servin:2008:RBC

[SL08] Martin Servin and Claude Lacoursière. Rigid body cable for virtual environments. IEEE Transactions on Visualization and Computer Graphics, 14(4):783– 796, July/August 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steed:2011:GEI

[SL11] Anthony Steed and Robert W. Lindeman. Guest Editor’s introduction: Special section on the Virtual Reality Conference (VR). IEEE Transactions on Visualization and Computer Graphics, 17(1): 1–2, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Santos:2009:VSC

[SLA^{+}09] Emanuele Santos, Lauro Lins, James Ahrens, Juliana Freire, and Claudio Silva. Vis-Mashup: Streamlining the creation of custom visualization applications. IEEE Transactions on Visualization and Computer Graphics, 15(6):1539–1546, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:2004:IVT

[SLB04] Han-Wei Shen, Guo-Shi Li, and Udeepta D. Bordoloi. Interactive visualization of three-dimensional vector fields with flexible appearance control. IEEE Transactions on Visualization and Computer Graphics, 10(4): 434–445, July/August 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/04/v0434.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0434.pdf.

Sicat:2019:DTB

[SLC^{+}19] Ronell Sicat, Jiabao Li, Junyoung Choi, Maxime Cordeil, Won-Ki Jeong, Benjamin Bach, and Hanspeter Pfister. DXR: A toolkit for building immersive data visualizations. IEEE Transactions on Visualization and Computer Graphics, 25(1): 715–725, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440858-
abs.html.

Suma:2012:ISM

[SLF^{+}12] Evan A. Suma, Zachary Lipps, Samantha Finkelstein, David M. Krum, and Mark Bolas. Impossible spaces: Maximizing natural walking in virtual environments with self-overlapping architecture. IEEE Transactions on Visualization and Computer Graphics, 18(4): 555–564, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shuai:2017:MCE

[SLG^{+}17] Liang Shuai, Chao Li, Xiaohu Guo, Balakrishnan Prabhakaran, and Jinxiang Chai. Motion capture with ellipsoidal skeleton using multiple depth cameras. IEEE Transactions on Visualization and Computer Graphics, 23(2): 1085–1098, February 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/02/07390092-
                                        abs.html.

Sajadi:2009:CSM

[SLGM09] Behzad Sajadi, Maxim Lazarov, M. Gopi, and Aditi Majumder. Color seamlessness in multi-projector displays using constrained gamut morphing. IEEE Transactions on Visualization and Computer Graphics, 15(6):1317–1326, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sheng:2019:DNR

[SLGM19] B. Sheng, P. Li, C. Gao, and K. Ma. Deep neural representation guided face sketch synthesis. IEEE Transactions on Visualization and Computer Graphics, 25(12):3216– 3230, December 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Smit:2017:PAB

[SLK^{+}17a] Noeska Smit, Kai Lawonn, Annelot Kraima, Marco DeRuiter, Hessam Sokooti, Stefan Bruckner, Elmar Eisemann, and Anna Vilanova. PelVis: Atlas-based surgical planning for oncological pelvic surgery. IEEE Transactions on Visualization and Computer Graphics, 23(1): 741–750, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Song:2017:GIV

[SLK^{+}17b] Hyunjoo Song, Jeongjin Lee, Tae Jung Kim, Kyoung Ho Lee, Bohyoung Kim, and Jinwook Seo. GazeDx: Interactive visual analytics framework for comparative gaze analysis with volumetric medical images. IEEE Transactions on Visualization and Computer Graphics, 23(1): 311–320, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schissler:2018:ACO

[SLM18] Carl Schissler, Christian Loftin, and Dinesh Manocha. Acoustic classification and optimization for multi-modal rendering of real-world scenes. IEEE Transactions on Visualization and Computer Graphics, 24(3):1246–1259, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sreng:2006:UVC

[SLMA06] Jean Sreng, Anatole Lécuyer, Christine Mégard, and Claude Andriot. Using visual cues of contact to improve interactive manipulation of virtual objects in industrial assembly/maintenance simulations. IEEE Transactions on Visualization and Computer Graphics, 12(5):1013–1020, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Servin:2011:HMW

[SLNB11] M. Servin, C. Lacoursiere, F. Nordfelth, and K. Bodin. Hybrid, multiresolution wires with massless frictional contacts. IEEE Transactions on Visualization and Computer Graphics, 17(7):970–982, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2017:EST

[SLQW17] Guodao Sun, Ronghua Liang, Huamin Qu, and Yingcai Wu. Embedding spatio-temporal information into maps by route-zooming. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1506–1519, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07420745-
abs.html.

Siegl:2017:FMN

[SLS^{+}17] Christian Siegl, Vanessa Lange, Marc Stamminger, Frank Bauer, and Justus Thies. FaceForge: Markerless non-rigid face multi-projection mapping. IEEE Transactions on Visualization and Computer Graphics, 23(11):2440–2446, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007312-
abs.html.

Sun:2010:HDR

[SLW^{+}10] Wen Sun, Yan Lu, Feng Wu, Shipeng Li, and John Tardif. High-dynamic-range texture compression for rendering systems of different capacities. IEEE Transactions on Visualization and Computer Graphics, 16(1):57–69, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Silva:1997:LSR

[SM97] C. T. Silva and J. S. B.

Mitchell. The lazy sweep ray casting algorithm for rendering irregular grids. IEEE Transactions on Visualization and Computer Graphics, 3 (2):142–157, April/June 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0142.
pdf;   http://www.computer.
org/tvcg/tg1997/v0142abs.
htm.

Schmidl:2004:FIC

[SM04] Harald Schmidl and Victor J. Milenkovic. A fast impulsive contact suite for rigid body simulation. IEEE Transactions on Visualization and Computer Graphics, 10(2): 189–197, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0189abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0189.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/02/v0189.pdf.

Slack:2006:CRD

[SM06] James Slack and Tamara Munzner. Composite rectilinear deformation for stretch and squish navigation. IEEE Transactions on Visualization and Computer Graphics, 12(5):901–908, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sajadi:2009:MVI

[SM09] Behzad Sajadi and Aditi Majumder. Markerless viewindependent registration of multiple distorted projectors on extruded surfaces using an uncalibrated camera. IEEE Transactions on Visualization and Computer Graphics, 15(6):1307–1316, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sajadi:2011:ATP

[SM11] Behzad Sajadi and Aditi Majumder. Autocalibrating tiled projectors on piecewise smooth vertically extruded surfaces. IEEE Transactions on Visualization and Computer Graphics, 17(9):1209– 1222, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sajadi:2012:AMC

[SM12] Behzad Sajadi and Aditi Majumder. Autocalibration of multiprojector CAVE-like immersive environments. IEEE Transactions on Visualization and Computer Graphics, 18 (3):381–393, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sauer:2017:STF

[SM17] Franz Sauer and Kwan-Liu Ma. Spatio-temporal feature exploration in combined particle/volume reference frames. IEEE Transactions on Visualization and Computer Graphics, 23(6): 1624–1635, June 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07864476-
abs.html.

Sadana:2014:OVT

[SMDS14] Ramik Sadana, Timothy Major, Alistair Dove, and John Stasko. OnSet: A visualization technique for large-scale binary set data. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1993–2002, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876026-
abs.html.

Shen:2006:VAL

[SMER06] Zeqian Shen, Kwan-Liu Ma, and Tina Eliassi-Rad. Visual analysis of large heterogeneous social networks by semantic and structural abstraction. IEEE Transactions on Visualization and Computer Graphics, 12(6):1427– 1439, November/December 2006. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schott:2013:AOE

[SMG^{+}13] Mathias Schott, Tobias Martin, A. V. Pascal Grosset, Sean T. Smith, and Charles D. Hansen. Ambient occlusion effects for combined volumes and tubular geometry. IEEE Transactions on Visualization and Computer Graphics, 19(6): 913–926, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sedlmair:2012:DSM

[SMM12] Michael Sedlmair, Miriah Meyer, and Tamara Munzner. Design study methodology: Reflections from the trenches and the stacks. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2431–2440, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shivashankar:2012:PCM

[SMN12] Nithin Shivashankar, Senthilnathan M., and Vijay Natarajan. Parallel computation of 2D Morse–Smale complexes. IEEE Transactions on Visualization and Computer Graphics, 18(10):1757–1770, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2016:BSE

[SMNR16] M. Sun, P. Mi, C. North, and N. Ramakrishnan. BiSet: Semantic edge bundling with biclusters for sensemaking. IEEE Transactions on Visualization and Computer Graphics, 22(1):310–319, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sajadi:2013:UPE

[SMO^{+}13] B. Sajadi, A. Majumder, M. M. Oliveira, R. G. Schneider, and R. Raskar. Using patterns to encode color information for dichromats. IEEE Transactions on Visualization and Computer Graphics, 19 (1):118–129, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schlegel:2011:EBS

[SMP11] Philipp Schlegel, Maxim Makhinya, and Renato Pajarola. Extinction-based shading and illumination in GPU volume ray-casting. IEEE Transactions on Visualization and Computer Graphics, 17(12):1795–1802, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Simeone:2017:AUM

[SMP17] Adalberto L. Simeone, Ifigeneia Mavridou, and Wendy Powell. Altering user movement behaviour in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 23(4): 1312–1321, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07835276-
abs.html.

Sedlmair:2013:EGS

[SMT13] Michael Sedlmair, Tamara Munzner, and Melanie Tory. Empirical guidance on scatterplot and dimension reduction technique choices. IEEE Transactions on Visualization and Computer Graphics, 19(12):2634–2643, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stuvel:2017:TC

[SMTT^{+}17] Sybren A. Stuvel, Nadia Magnenat-Thalmann, Daniel Thalmann, A. Frank van der Stappen, and Arjan Egges. Torso crowds. IEEE Transactions on Visualization and Computer Graphics, 23(7): 1823–1837, July 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07439844-
abs.html.

Satyanarayan:2017:VLG

[SMWH17] Arvind Satyanarayan, Dominik Moritz, Kanit Wongsuphasawat, and Jeffrey Heer. Vega-Lite: A grammar of interactive graphics. IEEE Transactions on Visualization and Computer Graphics, 23 (1):341–350, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Subramanian:1997:CDI

[SN97] K. R. Subramanian and B. F. Naylor. Converting discrete images to partitioning trees. IEEE Transactions on Visualization and Computer Graphics, 3(3):273–288, July/September 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0273.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0273abs.
                                        htm.

Singh:2010:RTR

[SN10] Jag Mohan Singh and P. J. Narayanan. Real-time ray tracing of implicit surfaces on the GPU. IEEE Transactions on Visualization and Computer Graphics, 16(2):261– 272, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Skarbez:2017:PER

[SNB^{+}17] Richard Skarbez, Solene Neyret, Frederick P. Brooks, Mel Slater, and Mary C.

Whitton. A psychophysical experiment regarding components of the plausibility illusion. IEEE Transactions on Visualization and Computer Graphics, 23(4):1369– 1378, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07833192-
abs.html.

Saraiya:2005:IBM

[SND05] Purvi Saraiya, Chris North, and Karen Duca. An insightbased methodology for evaluating bioinformatics visualizations. IEEE Transactions on Visualization and Computer Graphics, 11(4):443– 456, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schulz:2017:PGL

[SNG^{+}17] Christoph Schulz, Arlind Nocaj, Jochen Goertler, Oliver Deussen, Ulrik Brandes, and Daniel Weiskopf. Probabilistic graph layout for uncertain network visualization. IEEE Transactions on Visualization and Computer Graphics, 23 (1):531–540, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schulz:2013:DSV

[SNHS13] Hans-Jörg Schulz, Thomas Nocke, Magnus Heitzler, and Heidrun Schumann. A design space of visualization tasks. IEEE Transactions on Visualization and Computer Graphics, 19(12):2366– 2375, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saraiya:2006:IBL

[SNLD06] Purvi Saraiya, Chris North, Vy Lam, and Karen A. Duca. An insight-based longitudinal study of visual analytics. IEEE Transactions on Visualization and Computer Graphics, 12(6):1511–1522, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schissler:2016:EHB

[SNM16] C. Schissler, A. Nicholls, and R. Mehra. Efficient HRTFbased spatial audio for area and volumetric sources. IEEE Transactions on Visualization and Computer Graphics, 22 (4):1356–1366, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2014:FLD

[SNR14] Maoyuan Sun, Chris North, and Naren Ramakrishnan. A five-level design framework for bicluster visualizations. IEEE Transactions on Visualization and Computer Graphics, 20(12):1713– 1722, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875974-
abs.html.

Stanley:2017:DMB

[SO17] Andrew A. Stanley and Allison M. Okamura. Deformable model-based methods for shape control of a haptic jamming surface. IEEE Transactions on Visualization and Computer Graphics, 23(2): 1029–1041, February 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07399409-
abs.html.

Strobelt:2016:EAE

[SOK^{+}16a] H. Strobelt, D. Oelke, B. C.

Kwon, T. Schreck, and H. Pfister. Errata to “Guidelines for Effective Usage of Text Highlighting Techniques” [1]. IEEE Transactions on Visualization and Computer Graphics, 22(5): 1637, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Strobelt:2016:GEU

[SOK^{+}16b] H. Strobelt, D. Oelke, B. C.

Kwon, T. Schreck, and H. Pfister. Guidelines for effective usage of text highlighting techniques. IEEE Transactions on Visualization and Computer Graphics, 22(1): 489–498, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shafii:2013:VAV

[SOL^{+}13] Sohail Shafii, Herald Obermaier, Rodman Linn, Eunmo Koo, Mario Hlawitschka, Christoph Garth, Bernd Hamann, and Kenneth I. Joy. Visualization and analysis of vortex-turbine intersections in wind farms. IEEE Transactions on Visualization and Computer Graphics, 19(9):1579–1591, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sorger:2016:LIV

[SOL^{+}16] J. Sorger, T. Ortner, C. Luksch, M. Schwarzler, E. Groller, and H. Piringer. Lite-Vis: Integrated visualization for simulation-based decision support in lighting design. IEEE Transactions on Visualization and Computer Graphics, 22(1):290–299, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Strobelt:2009:DCT

[SOR^{+}09] Hendrik Strobelt, Daniela Oelke, Christian Rohrdantz, Andreas Stoffel, Daniel A. Keim, and Oliver Deussen. Document cards: a top trumps visualization for documents. IEEE Transactions on Visualization and Computer Graphics, 15(6):1145– 1152, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schops:2017:RTV

[SOS^{+}17] Thomas Schops, Martin R.

Oswald, Pablo Speciale, Shuoran Yang, and Marc Pollefeys. Real-time view correction for mobile devices. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2455–2462, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007219-
abs.html.

Sourin:1996:FRS

[SP96] Alexi I. Sourin and Alexander A. Pasko. Function representation for sweeping by a moving solid. IEEE Transactions on Visualization and Computer Graphics, 2 (1):11–18, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0011.
pdf;   http://www.computer.
org/tvcg/tg1996/v0011abs.
htm.

Sadlo:2007:EVL

[SP07] Filip Sadlo and Ronald Peikert. Efficient visualization of Lagrangian coherent structures by filtered AMR ridge extraction. IEEE Transactions on Visualization and Computer Graphics, 13(6):1456–1463, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1456s.
zip.

Sherbrooke:1996:AMA

[SPB96] Evan C. Sherbrooke, Nicholas M. Patrikalakis, and Erik Brisson. An algorithm for the medial axis transform of 3D polyhedral solids. IEEE Transactions on Visualization and Computer Graphics, 2 (1):44–61, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0044.
pdf;   http://www.computer.
org/tvcg/tg1996/v0044abs.
htm.

Streit:2008:SAF

[SPB08] Alexander Streit, Binh Pham, and Ross Brown. A spreadsheet approach to facilitate visualization of uncertainty in information. IEEE Transactions on Visualization and Computer Graphics, 14(1):61– 72, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Slavin:2006:TVT

[SPCJL06] Vadim Slavin, Robert Pelcovits, George Loriotl Andrew Callan-Jones, and David Laidlaw. Techniques for the visualization of topological defect behavior in nematic liquid crystals. IEEE Transactions on Visualization and Computer Graphics, 12(5):1323–1328, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Srinivasan:2018:GIS

[SPEB18] A. Srinivasan, H. Park, A. Endert, and R. C. Basole. Graphiti: Interactive specification of attribute-based edges for network modeling and visualization. IEEE Transactions on Visualization and Computer Graphics, 24 (1):226–235, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stolper:2014:PVA

[SPG14] Charles D. Stolper, Adam Perer, and David Gotz. Progressive visual analytics: User-driven visual exploration of in-progress analytics. IEEE Transactions on Visualization and Computer Graphics, 20(12):1653– 1662, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876049-
abs.html.

Smith:2007:NSS

[SPK^{+}07] Randall Smith, Richard Pawlicki, István Kókai, Jörg Finger, and Thomas Vetter. Navigating in a shape space of registered models. IEEE Transactions on Visualization and Computer Graphics, 13(6):1552–1559, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1552s.
mov.

Song:2013:AAS

[SPL^{+}13] Yi-Zhe Song, David Pickup, Chuan Li, Paul Rosin, and Peter Hall. Abstract art by shape classification. IEEE Transactions on Visualization and Computer Graphics, 19 (8):1252–1263, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schaefer:2013:MAH

[SPM^{+}13] Henry Schaefer, Magdalena Prus, Quirin Meyer, Jochen Suessmuth, and Marc Stamminger. Multiresolution attributes for hardware tessellated objects. IEEE Transactions on Visualization and Computer Graphics, 19(9): 1488–1498, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shivashankar:2016:FTB

[SPN^{+}16] Nithin Shivashankar, Pratyush Pranav, Vijay Natarajan, Rien van de Weygaert, E. G. Patrick Bos, and Steven Rieder. Felix: A topology based framework for visual exploration of cosmic filaments. IEEE Transactions on Visualization and Computer Graphics, 22(6):1745–1759, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schroder:2012:VFB

[SPO^{+}12] Simon Schroder, John A.

Peterson, Harald Obermaier, Louise H. Kellogg, Kenneth I. Joy, and Hans Hagen. Visualization of flow behavior in earth mantle convection. IEEE Transactions on Visualization and Computer Graphics, 18(12):2198– 2207, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Seo:2014:OLS

[SPP^{+}14] Byung-Kuk Seo, Hanhoon Park, Jong-Il Park, Stefan Hinterstoisser, and Slobodan Ilic. Optimal local searching for fast and robust textureless 3D object tracking in highly cluttered backgrounds. IEEE Transactions on Visualization and Computer Graphics, 20(1):99–110, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sadlo:2006:VTV

[SPS06] Filip Sadlo, Ronald Peikert, and Mirjam Sick. Visualization tools for vorticity transport analysis in incompressible flow. IEEE Transactions on Visualization and Computer Graphics, 12(5): 949–956, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sung:2002:STA

[SPW02] K. Sung, A. Pearce, and C. Wang. Spatial-temporal antialiasing. IEEE Transactions on Visualization and Computer Graphics, 8(2): 144–153, April 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0144.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0144abs.
                                        htm.

Su:2007:PEB

[SPW07] Wen-Poh Su, Binh Pham, and Aster Wardhani. Personality and emotion-based high-level control of affective story characters. IEEE Transactions on Visualization and Computer Graphics, 13(2):281– 293, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Spitz:2000:AAU

[SR00] S. N. Spitz and A. A. G. Requicha. Accessibility analysis using computer graphics hardware. IEEE Transactions on Visualization and Computer Graphics, 6(3):208–219, July/September 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0208.
pdf;   http://www.computer.
org/tvcg/tg2000/v0208abs.
htm.

Schneider:2017:VEG

[SR17] Jens Schneider and Peter Rautek. A versatile and efficient GPU data structure for spatial indexing. IEEE Transactions on Visualization and Computer Graphics, 23(1): 911–920, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Seron:2002:ASH

[SRCP02] Francisco J. Seron, Rafael Rodriguez, Eva Cerezo, and Alfredo Pina. Adding support for high-level skeletal animation. IEEE Transactions on Visualization and Computer Graphics, 8(4):360–372, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0360abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0360.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0360.pdf. See cor-

rection [SRCP03].

Seron:2003:CAS

[SRCP03] Francisco J. Seron, Rafael Rodriguez, Eva Cerezo, and Alfredo Pina. Correction to ‘Adding Support for High-Level Skeletal Animation’. IEEE Transactions on Visualization and Computer Graphics, 9(1):110, January/March 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/01/v0110.
pdf;   http://csdl.computer.
org/dl/trans/tg/2003/01/

v0110.htm. See [SRCP02].

Satyanarayan:2016:RVS

[SRHH16] A. Satyanarayan, R. Russell, J. Hoffswell, and J. Heer. Reactive Vega: A streaming dataflow architecture for declarative interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 22(1):659– 668, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stapleton:2011:IGE

[SRHZ11] Gem Stapleton, Peter Rodgers, John Howse, and Leishi Zhang. Inductively generating Euler diagrams. IEEE Transactions on Visualization and Computer Graphics, 17(1): 88–100, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schmitz:2011:ERO

[SRK^{+}11] Arne Schmitz, Tobias Rick, Thomas Karolski, Torsten Kuhlen, and Leif Kobbelt. Efficient rasterization for outdoor radio wave propagation. IEEE Transactions on Visualization and Computer Graphics, 17(2):159–170, February 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sterling:2019:AMR

[SRKL19] Auston Sterling, Nicholas Rewkowski, Roberta L. Klatzky, and Ming C. Lin. Audiomaterial reconstruction for virtualized reality using a probabilistic damping model. IEEE Transactions on Visualization and Computer Graphics, 25(5):1855–1864, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL https://ieeexplore.

ieee.org/document/8642440/

.

Shih:2019:DGF

[SRM19] Min Shih, Charles Rozhon, and Kwan-Liu Ma. A declarative grammar of flexible volume visualization pipelines. IEEE Transactions on Visualization and Computer Graphics, 25(1):1050– 1059, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440063-
abs.html.

Sun:2007:FEF

[SRML07] Xianfang Sun, Paul Rosin, Ralph Martin, and Frank Langbein. Fast and effective feature-preserving mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 13(5): 925–938, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2009:BRG

[SRML09] Xianfang Sun, Paul L. Rosin, Ralph R. Martin, and Frank C. Langbein. Bas-relief generation using adaptive histogram equalization. IEEE Transactions on Visualization and Computer Graphics, 15(4):642–653, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stott:2011:AMM

[SRMOW11] Jonathan Stott, Peter Rodgers, Juan Carlos Martinez-Ovando, and Stephen G. Walker. Automatic metro map layout using multicriteria optimization. IEEE Transactions on Visualization and Computer Graphics, 17(1):101–114, January 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Skraba:2016:CPC

[SRW^{+}16] Primoz Skraba, Paul Rosen, Bei Wang, Guoning Chen, Harsh Bhatia, and Valerio Pascucci. Critical point cancellation in 3D vector fields: Robustness and discussion. IEEE Transactions on Visualization and Computer Graphics, 22(6):1683–1693, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Slusallek:1995:VAG

[SS95] Philipp Slusallek and Hans-Peter Seidel. Vision — an architecture for global illumination calculations. IEEE Transactions on Visualization and Computer Graphics, 1 (1):77–96, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0077.
pdf;   http://www.computer.
org/tvcg/tg1995/v0077abs.
htm.

Salzbrunn:2006:SP

[SS06a] Tobias Salzbrunn and Gerik Scheuermann. Streamline predicates. IEEE Transactions on Visualization and Computer Graphics, 12(6):1601–1612, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/.

Seo:2006:KDH

[SS06b] Jinwook Seo and Ben Shneiderman. Knowledge discovery in high-dimensional data: Case studies and a user survey for the rank-by-feature framework. IEEE Transactions on Visualization and Computer Graphics, 12(3):311– 322, May/June 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schultz:2008:ECF

[SS08] Thomas Schultz and Hans-Peter Seidel. Estimating crossing fibers: a tensor decomposition approach. IEEE Transactions on Visualization and Computer Graphics, 14(6):1635–1642, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schenzel:2013:IVB

[SS13a] Peter Schenzel and Christian Stussak. Interactive visualizations of blowups of the plane. IEEE Transactions on Visualization and Computer Graphics, 19(6):978–990, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szymczak:2013:VMC

[SS13b] Andrzej Szymczak and Levente Sipeki. Visualization of Morse connection graphs for topologically rich 2D vector fields. IEEE Transactions on Visualization and Computer Graphics, 19(12):2763– 2772, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Setlur:2016:LAC

[SS16] V. Setlur and M. C. Stone. A linguistic approach to categorical color assignment for data visualization. IEEE Transactions on Visualization and Computer Graphics, 22(1): 698–707, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Srinivasan:2018:OFM

[SS18] A. Srinivasan and J. Stasko.

Orko: Facilitating multimodal interaction for visual exploration and analysis of networks. IEEE Transactions on Visualization and Computer Graphics, 24(1): 511–521, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Song:2019:WMD

[SS19] Hayeong Song and Danielle Albers Szafir. Where’s my data? Evaluating visualizations with missing data. IEEE Transactions on Visualization and Computer Graphics, 25(1): 914–924, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440857-
abs.html.

Schollmeyer:2017:EHI

[SSB^{+}17] Andre Schollmeyer, Simon Schneegans, Stephan Beck, Anthony Steed, and Bernd Froehlich. Efficient hybrid image warping for high frame-rate stereoscopic rendering. IEEE Transactions on Visualization and Computer Graphics, 23(4):1332– 1341, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07833188-
abs.html.

Sultanum:2019:DCB

[SSBC19] Nicole Sultanum, Devin Singh, Michael Brudno, and Fanny Chevalier. Doccurate: A curation-based approach for clinical text visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 142–151, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440832-
abs.html.

Sakurai:2016:IVS

[SSC^{+}16] D. Sakurai, O. Saeki, H. Carr, H. Wu, T. Yamamoto, D. Duke, and S. Takahashi. Interactive visualization for singular fibers of functions f: R^{3} → R^{2}. IEEE Transactions on Visualization and Computer Graphics, 22(1): 945–954, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Scheidegger:2008:RHI

[SSD^{+}08] Carlos E. Scheidegger, John M.

Schreiner, Brian Duffy, Hamish Carr, and Cláudio T. Silva. Revisiting histograms and isosurface statistics. IEEE Transactions on Visualization and Computer Graphics, 14(6):1659–1666, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Swan:2015:MRD

[SSE15] J. Edward Swan, Gurjot Singh, and Stephen R. Ellis. Matching and reaching depth judgments with real and augmented reality targets. IEEE Transactions on Visualization and Computer Graphics, 21(11):1289– 1298, November 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07164348-
abs.html.

Sarvghad:2019:EMS

[SSEW19] Ali Sarvghad, Bahador Saket, Alex Endert, and Nadir Weibel. Embedded merge & split: Visual adjustment of data grouping. IEEE Transactions on Visualization and Computer Graphics, 25(1): 800–809, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440855-
abs.html.

Su:2013:ECS

[SSF13] J. Su, R. Sheth, and R. Fedkiw. Energy conservation for the simulation of deformable bodies. IEEE Transactions on Visualization and Computer Graphics, 19(2):189– 200, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shiravi:2012:SVS

[SSG12] Hadi Shiravi, Ali Shiravi, and Ali A. Ghorbani. A survey of visualization systems for network security. IEEE Transactions on Visualization and Computer Graphics, 18 (8):1313–1329, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szafir:2016:LCS

[SSG16] Danielle Albers Szafir, Alper Sarikaya, and Michael Gleicher. Lightness constancy in surface visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (9):2107–2121, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stopar:2019:SEM

[SSGM19] Luka Stopar, Primoz Skraba, Marko Grobelnik, and Dunja Mladenic. StreamStory: Exploring multivariate time series on multiple scales. IEEE Transactions on Visualization and Computer Graphics, 25(4):1788–1802, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8340877/.

Seiler:2014:DDS

[SSH14] Martin Seiler, Jonas Spillmann, and Matthias Harders. Data-driven simulation of detailed surface deformations for surgery training simulators. IEEE Transactions on Visualization and Computer Graphics, 20(10):1379– 1391, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06797976-
abs.html.

Sato:1999:ARD

[SSI99] Imari Sato, Yoichi Sato, and Katsushi Ikeuchi. Acquiring a radiance distribution to superimpose virtual objects onto a real scene. IEEE Transactions on Visualization and Computer Graphics, 5(1): 1–12, January/March 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1999/pdf/v0001.
pdf;   http://www.computer.
org/tvcg/tg1999/v0001abs.
htm.

Selle:2009:RHR

[SSIF09] Andrew Selle, Jonathan Su, Geoffrey Irving, and Ronald Fedkiw. Robust highresolution cloth using parallelism, history-based collisions, and accurate friction. IEEE Transactions on Visualization and Computer Graphics, 15(2):339– 350, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sacha:2016:RUA

[SSK^{+}16] D. Sacha, H. Senaratne, B. C.

Kwon, G. Ellis, and D. A. Keim. The role of uncertainty, awareness, and trust in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 22 (1):240–249, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Saket:2014:NNL

[SSKB14] Bahador Saket, Paolo Simonetto, Stephen Kobourov, and Katy Borner. Node, node-link, and node-link-group diagrams: An evaluation. IEEE Transactions on Visualization and Computer Graphics, 20(12):2231– 2240, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876036-
abs.html.

Steed:2008:GEI

[SSL08] Anthony Steed, William Sherman, and Ming C. Lin. Guest Editor’s introduction: Special section on virtual reality. IEEE Transactions on Visualization and Computer Graphics, 14(3):485– 486, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/03/ttg2008030485.
pdf.

Streit:2012:MDD

[SSL^{+}12] Marc Streit, Hans-Jörg Schulz, Alexander Lex, Dieter Schmalstieg, and Heidrun Schumann. Model-driven design for the visual analysis of heterogeneous data. IEEE Transactions on Visualization and Computer Graphics, 18(6): 998–1010, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sambasivan:2013:VRF

[SSMG13] Raja R. Sambasivan, Ilari Shafer, Michelle L. Mazurek, and Gregory R. Ganger. Visualizing request-flow comparison to aid performance diagnosis in distributed systems. IEEE Transactions on Visualization and Computer Graphics, 19(12):2466– 2475, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2007:TVB

[SSR^{+}07] Bo Sun, Kalyan Sunkavalli, Ravi Ramamoorthi, Peter N. Belhumeur, and Shree K. Nayar. Time-varying BRDFs. IEEE Transactions on Visualization and Computer Graphics, 13(3):595– 609, May/June 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/

comp/trans/tg/2007/03/v0595s.
avi.

Saket:2018:EIG

[SSRE18] Bahador Saket, Arjun Srinivasan, Eric D. Ragan, and Alex Endert. Evaluating interactive graphical encodings for data visualization. IEEE Transactions on Visualization and Computer Graphics, 24 (3):1316–1330, March 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schreiner:2006:HQE

[SSS06] John Schreiner, Carlos Scheidegger, and Claudio Silva. High-quality extraction of isosurfaces from regular and irregular grids. IEEE Transactions on Visualization and Computer Graphics, 12(5):1205–1212, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steptoe:2013:HTO

[SSS13] William Steptoe, Anthony Steed, and Mel Slater. Human tails: Ownership and control of extended humanoid avatars. IEEE Transactions on Visualization and Computer Graphics, 19(4): 583–590, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sacha:2014:KGM

[SSS^{+}14] Dominik Sacha, Andreas Stoffel, Florian Stoffel, Bum Chul Kwon, Geoffrey Ellis, and Daniel A. Keim. Knowledge generation model for visual analytics. IEEE Transactions on Visualization and Computer Graphics, 20(12):1604– 1613, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875967-
abs.html.

Schwab:2017:BIE

[SST^{+}17] Michail Schwab, Hendrik Strobelt, James Tompkin, Colin Fredericks, Connor Huff, Dana Higgins, Anton Strezhnev, Mayya Komisarchik, Gary King, and Hanspeter Pfister. booc.io: An education system with hierarchical concept maps and dynamic non-linear learning plans. IEEE Transactions on Visualization and Computer Graphics, 23(1):571–580, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sondag:2018:STL

[SSV18] M. Sondag, B. Speckmann, and K. Verbeek. Stable treemaps via local moves. IEEE Transactions on Visualization and Computer Graphics, 24(1):729–738, January 2018. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2018:ISR

[SSW18] Yujing Sun, Scott Schaefer, and Wenping Wang. Image structure retrieval via L_{0} minimization. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2129–2139, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07938357-
abs.html.

Spillmann:2009:CN

[ST09] Jonas Spillmann and Matthias Teschner. Cosserat nets. IEEE Transactions on Visualization and Computer Graphics, 15(2):325–338, March/ April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Said:2018:IBM

[STB18] Souheil Hadj Said, Mohamed Tamaazousti, and Adrien Bartoli. Image-based models for specularity propagation in diminished reality. IEEE Transactions on Visualization and Computer Graphics, 24(7):2140–2152, July 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/07/07931582-
abs.html.

Stewart:1998:FHC

[Ste98] A. J. Stewart. Fast horizon computation at all points of a terrain with visibility and shading applications. IEEE Transactions on Visualization and Computer Graphics, 4(1): 82–93, January/March 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0082.
pdf;   http://www.computer.
org/tvcg/tg1998/v0082abs.
htm.

Stolte:2002:PSQ

[STH02] Chris Stolte, Diane Tang, and Pat Hanrahan. Polaris: a system for query, analysis, and visualization of multidimensional relational databases. IEEE Transactions on Visualization and Computer Graphics, 8(1):52–65, January 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0052.
pdf;   http://www.computer.
org/tvcg/tg2002/v0052abs.
htm.

Stolte:2003:MVU

[STH03] Chris Stolte, Diane Tang, and Pat Hanrahan. Multiscale visualization using data cubes. IEEE Transactions on Visualization and Computer Graphics, 9(2):176– 187, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0176abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0176.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0176.pdf.

Spillmann:2013:ASW

[STH13] Jonas Spillmann, Stefan Tuchschmid, and Matthias Harders. Adaptive space warping to enhance passive haptics in an arthroscopy surgical simulator. IEEE Transactions on Visualization and Computer Graphics, 19(4): 626–633, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Soni:2008:VPF

[STM08] Bela Soni, David Thompson, and Raghu Machiraju. Visualizing particle/flow structure interactions in the small bronchial tubes. IEEE Transactions on Visualization and Computer Graphics, 14(6):1412–1427, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sarvghad:2017:VDC

[STM17] Ali Sarvghad, Melanie Tory, and Narges Mahyar. Visualizing dimension coverage to support exploratory analysis. IEEE Transactions on Visualization and Computer Graphics, 23(1):21–30, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Solteszova:2012:PSS

[ŠTPV12] Veronika Šoltészová, Cagatay Turkay, Mark C. Price, and Ivan Viola. A perceptual-statistics shading model. IEEE Transactions on Visualization and Computer Graphics, 18(12):2265– 2274, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sauber:2006:MGA

[STS06] Natascha Sauber, Holger Theisel, and Hans-Peter Seidel. Multifield-graphs: an approach to visualizing correlations in multifield scalar data. IEEE Transactions on Visualization and Computer Graphics, 12(5):917–924, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schultz:2007:TVB

[STS07] Thomas Schultz, Holger Theisel, and Hans-Peter Seidel. Topological visualization of brain diffusion MRI data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1496–1503, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schultz:2010:CST

[STS10] Thomas Schultz, Holger Theisel, and Hans-Peter Seidel. Crease surfaces: From theory to extraction and application to diffusion tensor MRI. IEEE Transactions on Visualization and Computer Graphics, 16(1):109–119, January/February 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stusak:2014:ASE

[STS^{+}14] Simon Stusak, Aurelien Tabard, Franziska Sauka, Rohit Ashok Khot, and Andreas Butz. Activity sculptures: Exploring the impact of physical visualizations on running activity. IEEE Transactions on Visualization and Computer Graphics, 20(12):2201– 2210, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06888482-
abs.html.

Sturzlinger:1998:RTT

[Stü98] W. Stürzlinger. Ray-tracing triangular trimmed free-form surfaces. IEEE Transactions on Visualization and Computer Graphics, 4(3):202–214, July/September 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1998/pdf/v0202.
pdf;   http://www.computer.
org/tvcg/tg1998/v0202abs.
htm.

Shin-Ting:2012:ICR

[STYC12] Wu Shin-Ting, Clarissa Lin Yasuda, and Fernando Cendes. Interactive curvilinear reformatting in native space. IEEE Transactions on Visualization and Computer Graphics, 18(2):299–308, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sundquist:2003:DLI

[Sun03] Andreas Sundquist. Dynamic line integral convolution for visualizing streamline evolution. IEEE Transactions on Visualization and Computer Graphics, 9(3):273–282, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0273abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0273.pdf.

San-Vicente:2012:CMS

[SVAC12] Gaizka San-Vicente, Iker Aguinaga, and Juan Tomas Celigueta. Cubical massspring model design based on a tensile deformation test and nonlinear material model. IEEE Transactions on Visualization and Computer Graphics, 18(2):228–241, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sewall:2011:VTR

[SvdBLM11] Jason Sewall, Jur van den Berg, Ming C. Lin, and Dinesh Manocha. Virtualized traffic: Reconstructing traffic flows from discrete spatiotemporal data. IEEE Transactions on Visualization and Computer Graphics, 17(1):26– 37, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Skanberg:2016:RTM

[SVGR16] R. Skanberg, P. Vazquez, V. Guallar, and T. Ropinski. Real-time molecular visualization supporting diffuse interreflections and ambient occlusion. IEEE Transactions on Visualization and Computer Graphics, 22(1):718–727, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Scheidegger:2007:QCV

[SVK^{+}07] Carlos Scheidegger, Huy Vo, David Koop, Juliana Freire, and Claudio Silva. Querying and creating visualizations by analogy. IEEE Transactions on Visualization and Computer Graphics, 13(6):1560–1567, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1560s.
                                        mov.

Smit:2010:PDL

[SvLF10] Ferdi Alexander Smit, Robert van Liere, and Bernd Froehlich. A programmable display layer for virtual reality system architectures. IEEE Transactions on Visualization and Computer Graphics, 16(1):28– 42, January/February 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schneider:1998:EPC

[SvW98] B.-O. Schneider and J. van Welzen. Efficient polygon clipping for an SIMD graphics pipeline. IEEE Transactions on Visualization and Computer Graphics, 4(3):272– 285, July/September 1998. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0272.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0272abs.
                                        htm.

Silver:1997:TVT

[SW97] D. Silver and X. Wang. Tracking and visualizing turbulent 3D features. IEEE Transactions on Visualization and Computer Graphics, 3(2):129– 141, April/June 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0129.
pdf;   http://www.computer.
org/tvcg/tg1997/v0129abs.
htm.

Stasko:2006:GEI

[SW06] John Stasko and Matthew O.

Ward. Guest editorial: InfoVis 2005. IEEE Transactions on Visualization and Computer Graphics, 12(4): 535, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/04/v0535.
pdf.

Sanftmann:2012:SN

[SW12] Harald Sanftmann and Daniel Weiskopf. 3D scatterplot navigation. IEEE Transactions on Visualization and Computer Graphics, 18(11):1969– 1978, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shadoan:2013:VAH

[SW13] Rachel Shadoan and Chris Weaver. Visual analysis of higher-order conjunctive relationships in multidimensional data using a hypergraph query system. IEEE Transactions on Visualization and Computer Graphics, 19(12):2070–2079, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Subr:2017:GEI

[SW17] Kartic Subr and Li-Yi Wei. Guest Editor’s introduction to the special section on the ACM Symposium on Interactive 3D Graphics and Games (I3D). IEEE Transactions on Visualization and Computer Graphics, 23(5): 1427, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07891683.
pdf.

Sato:2000:TCB

[SWB^{+}00] Y. Sato, C.-F. Westin, A. Bhalerao, S. Nakajima, N. Shiraga, S. Tamura, and R. Kikinis. Tissue classification based on 3D local intensity structures for volume rendering. IEEE Transactions on Visualization and Computer Graphics, 6(2):160– 180, April/June 2000. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2000/pdf/v0160.
pdf;   http://www.computer.

org/tvcg/tg2000/v0160abs.
htm.

Schneider:2008:ICS

[SWC^{+}08] Dominic Schneider, Alexander Wiebel, Hamish Carr, Mario Hlawitschka, and Gerik Scheuermann. Interactive comparison of scalar fields based on largest contours with applications to flow visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1475– 1482, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Skraba:2015:RBS

[SWCR15] Primoz Skraba, Bei Wang, Guoning Chen, and Paul Rosen. Robustness-based simplification of 2D steady and unsteady vector fields. IEEE Transactions on Visualization and Computer Graphics, 21(8):930–944, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/07117431-abs.html.

Sui:2016:LWF

[SWF^{+}16] W. Sui, L. Wang, B. Fan, H. Xiao, H. Wu, and C. Pan. Layer-wise floorplan extraction for automatic urban building reconstruction. IEEE Transactions on Visualization and Computer Graphics, 22 (3):1261–1277, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shi:2014:LVL

[SWL^{+}14a] Conglei Shi, Yingcai Wu, Shixia Liu, Hong Zhou, and Huamin Qu. LoyalTracker: Visualizing loyalty dynamics in search engines. IEEE Transactions on Visualization and Computer Graphics, 20(12):1733– 1742, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876038-
abs.html.

Sun:2014:EVA

[SWL^{+}14b] Guodao Sun, Yingcai Wu, Shixia Liu, Tai-Quan Peng, Jonathan J. H. Zhu, and Ronghua Liang. EvoRiver: Visual analysis of topic coopetition on social media. IEEE Transactions on Visualization and Computer Graphics, 20(12):1753– 1762, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875992-
abs.html.

Schindler:2013:MDF

[SWR^{+}13] Benjamin Schindler, Jürgen Waser, Hrvoje Ribičić, Raphael Fuchs, and Ronald Peikert.

Multiverse data-flow control. IEEE Transactions on Visualization and Computer Graphics, 19(6):1005–1019, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Steinberger:2011:CPV

[SWS^{+}11] Markus Steinberger, Manuela Waldner, Marc Streit, Alexander Lex, and Dieter Schmalstieg. Context-preserving visual links. IEEE Transactions on Visualization and Computer Graphics, 17(12):2249– 2258, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sahner:2007:VSS

[SWTH07] Jan Sahner, Tino Weinkauf, Nathalie Teuber, and Hans-Christian Hege. Vortex and strain skeletons in Eulerian and Lagrangian frames. IEEE Transactions on Visualization and Computer Graphics, 13(5):980–990, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Scheepens:2011:CDM

[SWvdW^{+}11] Roeland Scheepens, Niels Willems, Huub van de Wetering, Gennady Andrienko, Natalia Andrienko, and Jarke J. van Wijk. Composite density maps for multivariate trajectories. IEEE Transactions on Visualization and Computer Graphics, 17(12):2518– 2527, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shi:2015:EDN

[SWW^{+}15] Lei Shi, Chen Wang, Zhen Wen, Huamin Qu, Chuang Lin, and Qi Liao. 1.5D egocentric dynamic network visualization. IEEE Transactions on Visualization and Computer Graphics, 21(5): 624–637, May 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:2017:NVA

[SWY^{+}17] Qiaomu Shen, Tongshuang Wu, Haiyan Yang, Yanhong Wu, Huamin Qu, and Weiwei Cui. NameClarifier: A visual analytics system for author name disambiguation. IEEE Transactions on Visualization and Computer Graphics, 23(1):141–150, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sauer:2017:CEL

[SXM17] Franz Sauer, Jinrong Xie, and Kwan-Liu Ma. A combined Eulerian–Lagrangian data representation for large-scale applications. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2248–2261, October 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07676418-
abs.html.

Shu:2019:DPI

[SXX^{+}19] Z. Shu, S. Xin, X. Xu, L. Liu, and L. Kavan. Detecting 3D points of interest using multiple features and stacked autoencoder. IEEE Transactions on Visualization and Computer Graphics, 25(8):2583– 2596, August 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shkurko:2018:TIR

[SYK^{+}18] Konstantin Shkurko, Cem Yuksel, Daniel Kopta, Ian Mallett, and Erik Brunvand. Time interval ray tracing for motion blur. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3225–3238, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8115176/.

Sauer:2014:TBF

[SYM14] Franz Sauer, Hongfeng Yu, and Kwan-Liu Ma. Trajectorybased flow feature tracking in joint particle/volume datasets. IEEE Transactions on Visualization and Computer Graphics, 20(12):2565– 2574, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875975-
                                        abs.html.

Sunden:2011:IPS

[SYR11] Erik Sunden, Anders Ynnerman, and Timo Ropinski. Image plane sweep volume illumination. IEEE Transactions on Visualization and Computer Graphics, 17(12):2125– 2134, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Song:2006:AVA

[SYS^{+}06] Yuyan Song, Jing Ye, Nikolai Svakhine, Sonia Lasher-Trapp, Mike Baldwin, and David Ebert. An atmospheric visual analysis and exploration system. IEEE Transactions on Visualization and Computer Graphics, 12(5):1157–1164, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sheng:2011:SAR

[SYYC11] Yu Sheng, Theodore C. Yapo, Christopher Young, and Barbara Cutler. A spatially augmented reality sketching interface for architectural daylighting design. IEEE Transactions on Visualization and Computer Graphics, 17(1):38– 50, January 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Southern:2011:MSA

[SZ11] Richard Southern and Jian J.

Zhang. Motion-sensitive anchor identification of leastsquares meshes from examples. IEEE Transactions on Visualization and Computer Graphics, 17(6): 850–856, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szymczak:2012:RMD

[SZ12] Andrzej Szymczak and Eugene Zhang. Robust Morse decompositions of piecewise constant vector fields. IEEE Transactions on Visualization and Computer Graphics, 18 (6):938–951, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szalay:2010:VCA

[Sza10] Alexander S. Szalay. VisWeek capstone address. IEEE Transactions on Visualization and Computer Graphics, 16(6):xxv–xxvi, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szafir:2018:MCD

[Sza18] D. A. Szafir. Modeling color difference for visualization design. IEEE Transactions on Visualization and Computer Graphics, 24(1):392–401, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sanyal:2009:USC

[SZB^{+}09] Jibonananda Sanyal, Song Zhang, Gargi Bhattacharya, Phil Amburn, and Robert Moorhead. A user study to compare four uncertainty visualization methods for 1D and 2D datasets. IEEE Transactions on Visualization and Computer Graphics, 15(6):1209–1218, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sanyal:2010:NTV

[SZD^{+}10] Jibonananda Sanyal, Song Zhang, Jamie Dyer, Andrew Mercer, Philip Amburn, and Robert Moorhead. Noodles: a tool for visualization of numerical weather model ensemble uncertainty. IEEE Transactions on Visualization and Computer Graphics, 16(6):1421–1430, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Stalling:1997:FDI

[SZH97] D. Stalling, M. Zöckler, and H.-C. Hege. Fast display of illuminated field lines. IEEE Transactions on Visualization and Computer Graphics, 3 (2):118–128, April/June 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0118.
pdf;   http://www.computer.
org/tvcg/tg1997/v0118abs.
htm.

Stapleton:2011:DED

[SZHR11] G. Stapleton, L. Zhang, J. Howse, and P. Rodgers. Drawing Euler diagrams with circles: The theory of piercings. IEEE Transactions on Visualization and Computer Graphics, 17(7):1020–1032, July 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schroder:2015:IBR

[SZK15] Kai Schroder, Arno Zinke, and Reinhard Klein. Imagebased reverse engineering and visual prototyping of woven cloth. IEEE Transactions on Visualization and Computer Graphics, 21(2):188– 200, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/06858070-
abs.html.

Shen:2018:RTB

[SZN^{+}18] Liang Shen, Dengming Zhu, Saad Nadeem, Zhaoqi Wang, and Arie E. Kaufman. Radiative transport based flame volume reconstruction from videos. IEEE Transactions on Visualization and Computer Graphics, 24(7):2209– 2222, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07940071-
                                        abs.html.

Sacha:2017:VID

[SZS^{+}17] Dominik Sacha, Leishi Zhang, Michael Sedlmair, John A. Lee, Jaakko Peltonen, Daniel Weiskopf, Stephen C. North, and Daniel A. Keim. Visual interaction with dimensionality reduction: A structured literature analysis. IEEE Transactions on Visualization and Computer Graphics, 23 (1):241–250, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sun:2019:EEB

[SZW^{+}19] M. Sun, J. Zhao, H. Wu, K. Luther, C. North, and N. Ramakrishnan. The effect of edge bundling and seriation on sensemaking of biclusters in bipartite graphs. IEEE Transactions on Visualization and Computer Graphics, 25(10):2983–2998, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szymczak:2013:HSM

[Szy13] Andrzej Szymczak. Hierarchy of stable Morse decompositions. IEEE Transactions on Visualization and Computer Graphics, 19(5):799–810, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:2018:SVE

[SZY^{+}18] Q. Shen, W. Zeng, Y. Ye, S. M. Arisona, S. Schubiger, R. Burkhard, and H. Qu. StreetVizor: Visual exploration of human-scale urban forms based on street views. IEEE Transactions on Visualization and Computer Graphics, 24(1):1004–1013, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tachi:2010:OPS

[Tac10] Tomohiro Tachi. Origamizing polyhedral surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(2):298–311, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tatu:2011:AAM

[TAE^{+}11] Andrada Tatu, Georgia Albuquerque, Martin Eisemann, Peter Bak, Holger Theisel, Marcus Magnor, and Daniel Keim. Automated analytical methods to support visual exploration of high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 17(5):584–597, May 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tadakuma:2005:DAM

[TAK^{+}05] Riichiro Tadakuma, Yoshiaki Asahara, Hiroyuki Kajimoto, Naoki Kawakami, and Susumu Tachi. Development of anthropomorphic multi-D.O.F. master-slave arm for mutual telexistence. IEEE Transactions on Visualization and Computer Graphics, 11(6):626–636, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Theobalt:2007:SPD

[TAL^{+}07] Christian Theobalt, Naveed Ahmed, Hendrik Lensch, Marcus Magnor, and Hans-Peter Seidel. Seeing people in different light-joint shape, motion, and reflectance capture. IEEE Transactions on Visualization and Computer Graphics, 13(4):663– 674, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Taylor:2002:VMC

[Tay02] C. J. Taylor. VideoPlus: a method for capturing the structure and appearance of immersive environments. IEEE Transactions on Visualization and Computer Graphics, 8(2):171–182, April 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0171.
pdf;   http://www.computer.
org/tvcg/tg2002/v0171abs.
htm.

Termeer:2007:CCV

[TBB^{+}07] Maurice Termeer, Javier Oliván Bescós, Marcel Breeuwer, Anna Vilanova, Frans Gerritsen, and Eduard Gröller. CoViCAD: Comprehensive visualization of coronary artery disease. IEEE Transactions on Visualization and Computer Graphics, 13(6):1632– 1639, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2007/
06/v1632s.avi.

Termeer:2008:VMP

[TBB^{+}08] Maurice Termeer, Javier Oliván Bescós, Marcel Breeuwer, Anna Vilanova, Frans Gerritsen, M. Eduard Gröller, and Eike Nagel. Visualization of myocardial perfusion derived from coronary anatomy. IEEE Transactions on Visualization and Computer Graphics, 14(6):1595–1602, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thudt:2016:VMR

[TBHC16] A. Thudt, D. Baur, S. Huron, and S. Carpendale. Visual mementos: Reflecting memories with personal data. IEEE Transactions on Visualization and Computer Graphics, 22 (1):369–378, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Treib:2012:TVT

[TBR^{+}12] Marc Treib, Kai Burger, Florian Reichl, Charles Meneveau, Alex Szalay, and Rudiger Westermann. Turbulence visualization at the terascale on desktop PCs. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2169–2177, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tang:2009:QDM

[TC09a] Kai Tang and Ming Chen. Quasi-developable mesh surface interpolation via mesh deformation. IEEE Transactions on Visualization and Computer Graphics, 15(3): 518–528, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Teyseyre:2009:OSV

[TC09b] Alfredo R. Teyseyre and Marcelo R. Campo. An overview of 3D software visualization. IEEE Transactions on Visualization and Computer Graphics, 15(1):87–105, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tak:2013:ESS

[TC13] S. Tak and A. Cockburn. Enhanced spatial stability with Hilbert and Moore treemaps. IEEE Transactions on Visualization and Computer Graphics, 19(1):141–148, January 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tierny:2017:JFS

[TC17] Julien Tierny and Hamish Carr. Jacobi fiber surfaces for bivariate Reeb space computation. IEEE Transactions on Visualization and Computer Graphics, 23(1):960–969, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tam:2013:RPC

[TCL^{+}13] Gary K. L. Tam, Zhi-Quan Cheng, Yu-Kun Lai, Frank C. Langbein, Yonghuai Liu, David Marshall, Ralph R. Martin, Xian-Fang Sun, and Paul L. Rosin. Registration of 3D point clouds and meshes: A survey from rigid to nonrigid. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1199–1217, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tarini:2006:AOE

[TCM06] Marco Tarini, Paolo Cignoni, and Claudio Montani. Ambient occlusion and edge cueing for enhancing real time molecular visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1237–1244, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tikhonova:2010:VPN

[TCM10] Anna Tikhonova, Carlos D. Correa, and Kwan-Liu Ma. Visualization by proxy: a novel framework for deferred interaction with volume data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1551–1559, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Taylor:2012:GMR

[TCM^{+}12] Micah Taylor, Anish Chandak, Qi Mo, Christian Lauterbach, Carl Schissler, and Dinesh Manocha. Guided multiview ray tracing for fast auralization. IEEE Transactions on Visualization and Computer Graphics, 18(11):1797– 1810, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

tenCaat:2007:DET

[tCMR07] Michael ten Caat, Natasha M.

Maurits, and Jos B. T. M. Roerdink. Design and evaluation of tiled parallel coordinate visualization of multichannel EEG data. IEEE Transactions on Visualization and Computer Graphics, 13 (1):70–79, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

tenCaat:2008:DDV

[tCMR08] Michael ten Caat, Natasha M.

Maurits, and Jos. B. T. M. Roerdink. Data-driven visualization and group analysis of multichannel EEG coherence with functional units. IEEE Transactions on Visualization and Computer Graphics, 14(4):756–771, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tang:2009:IIC

[TCYM09] Min Tang, Sean Curtis, Sung-Eui Yoon, and Dinesh Manocha. ICCD: Interactive continuous collision detection between deformable models using connectivitybased culling. IEEE Transactions on Visualization and Computer Graphics, 15(4): 544–557, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tal:1995:VGA

[TD95] Ayellet Tal and David Dobkin. Visualization of geometric algorithms. IEEE Transactions on Visualization and Computer Graphics, 1(2): 194–204, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0194.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0194abs.
                                        htm.

Tennekes:2014:TCC

[TdJ14] Martijn Tennekes and Edwin de Jonge. Tree Colors: Color schemes for treestructured data. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2072–2081, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875961-

abs.html. See errata [TdJ15].

Tennekes:2015:ETC

[TdJ15] Martijn Tennekes and Edwin de Jonge. Errata to “Tree Colors: Color Schemes for Tree-Structured Data”. IEEE Transactions on Visualization and Computer Graphics, 21 (1):136, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06966879-

abs.html. See [TdJ14].

Tan:2019:PPB

[TDLG19] J. Tan, S. DiVerdi, J. Lu, and Y. Gingold. Pigmento: Pigment-based image analysis and editing. IEEE Transactions on Visualization and Computer Graphics, 25(9): 2791–2803, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tang:2018:ACC

[TDM^{+}18] Fan Tang, Weiming Dong, Yiping Meng, Xing Mei, Feiyue Huang, Xiaopeng Zhang, and Oliver Deussen. Animated construction of Chinese brush paintings. IEEE Transactions on Visualization and Computer Graphics, 24(12):3019–3031, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8113507/.

Tierny:2012:IQR

[TDN^{+}12] Julien Tierny, Joel Daniels II, Luis Gustavo Nonato, Valerio Pascucci, and Claudio T. Silva. Interactive quadrangulation with Reeb atlases and connectivity textures. IEEE Transactions on Visualization and Computer Graphics, 18(10):1650– 1663, October 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tchiboukdjian:2010:BMP

[TDR10] Marc Tchiboukdjian, Vincent Danjean, and Bruno Raffin. Binary mesh partitioning for cache-efficient visualization. IEEE Transactions on Visualization and Computer Graphics, 16(5): 815–828, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tong:2016:VDS

[TEC^{+}16] Xin Tong, John Edwards, Chun-Ming Chen, Han-Wei Shen, Chris R. Johnson, and Pak Chung Wong. Viewdependent streamline deformation and exploration. IEEE Transactions on Visualization and Computer Graphics, 22 (7):1788–1801, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Teran:2006:DSA

[TF06] Rachel Weinstein Joseph Teran and Ron Fedkiw. Dynamic simulation of articulated rigid bodies with contact and collision. IEEE Transactions on Visualization and Computer Graphics, 12 (3):365–374, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Turkay:2011:BDD

[TFH11] Cagatay Turkay, Peter Filzmoser, and Helwig Hauser. Brushing dimensions — a dual visual analysis model for high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 17(12):2591–2599, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tominski:2012:ISV

[TFJ12] Christian Tominski, Camilla Forsell, and Jimmy Johansson. Interaction support for visual comparison inspired by natural behavior. IEEE Transactions on Visualization and Computer Graphics, 18(12):2719–2728, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tierny:2018:TT

[TFL^{+}18] J. Tierny, G. Favelier, J. A.

Levine, C. Gueunet, and M. Michaux. The Topology ToolKit. IEEE Transactions on Visualization and Computer Graphics, 24(1): 832–842, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Takahashi:2009:AML

[TFO09] Shigeo Takahashi, Issei Fujishiro, and Masato Okada. Applying manifold learning to plotting approximate contour trees. IEEE Transactions on Visualization and Computer Graphics, 15(6):1185– 1192, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Talbot:2012:EMS

[TGH12] Justin Talbot, John Gerth, and Pat Hanrahan. An empirical model of slope ratio comparisons. IEEE Transactions on Visualization and Computer Graphics, 18(12):2613– 2620, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tricoche:2011:VTS

[TGS11] Xavier Tricoche, Christoph Garth, and Allen Sanderson. Visualization of topological structures in area-preserving maps. IEEE Transactions on Visualization and Computer Graphics, 17(12):1765– 1774, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tierny:2009:LSV

[TGSP09] Julien Tierny, Attila Gyulassy, Eddie Simon, and Valerio Pascucci. Loop surgery for volumetric meshes: Reeb graphs reduced to contour trees. IEEE Transactions on Visualization and Computer Graphics, 15(6):1177– 1184, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tuceryan:1995:CRP

[TGW^{+}95] Mihran Tuceryan, Douglas S.

Greer, Ross T. Whitaker, David E. Breen, Chris Crampton, Eric Rose, and Klaus H. Ahlers. Calibration requirements and procedures for a monitor-based augmented reality system. IEEE Transactions on Visualization and Computer Graphics, 1(3):255– 273, September 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0255.
pdf;   http://www.computer.
org/tvcg/tg1995/v0255abs.
htm.

Taimouri:2013:VSM

[TH13] Vahid Taimouri and Jing Hua.

Visualization of shape motions in shape space. IEEE Transactions on Visualization and Computer Graphics, 19(12):2644–2652, December 2013. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thagard:2011:VKA

[Tha11]  Paul Thagard.  VisWeek

keynote address. IEEE Transactions on Visualization and Computer Graphics, 17(12): xxiii, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Trotts:1999:STM

[THJ99] Issac J. Trotts, Bernd Hamann, and Kenneth I. Joy. Simplification of tetrahedral meshes with error bounds. IEEE Transactions on Visualization and Computer Graphics, 5(3):224–237, July/ September 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0224.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0224abs.
                                        htm.

Tanahashi:2015:EFG

[THM15] Y. Tanahashi, C. Hsueh, and K. Ma. An efficient framework for generating storyline visualizations from streaming data. IEEE Transactions on Visualization and Computer Graphics, 21(6): 730–742, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Traore:2019:IOF

[THT19] Michael Traore, Christophe Hurter, and Alexandru Telea. Interactive obstruction-free lensing for volumetric data visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 1029–1039, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08467383-
abs.html.

Turk:2014:MES

[THV^{+}14] Matthew Turk, Tobias Hollerer, Jonathan Ventura, Chris Sweeney, and Steffen Gauglitz. Model estimation and selection towards unconstrained real-time tracking and mapping. IEEE Transactions on Visualization and Computer Graphics, 20(6):825– 838, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tobiasz:2009:LCC

[TIC09] Matthew Tobiasz, Petra Isenberg, and Sheelagh Carpendale. Lark: Coordinating colocated collaboration with information visualization. IEEE Transactions on Visualization and Computer Graphics, 15(6):1065–1072, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tsukamoto:2015:RCC

[TIK15] Jun Tsukamoto, Daisuke Iwai, and Kenji Kashima. Radiometric compensation for cooperative distributed multiprojection system through 2-DOF distributed control. IEEE Transactions on Visualization and Computer Graphics, 21(11):1221–1229, November 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/11/07164338-
                                        abs.html.

Takeda:2016:IRC

[TIS16] S. Takeda, D. Iwai, and K. Sato. Inter-reflection compensation of immersive projection display by spatiotemporal screen reflectance modulation. IEEE Transactions on Visualization and Computer Graphics, 22(4): 1424–1431, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2019:ETV

[TIW^{+}19] Jun Tao, Martin Imre, Chaoli Wang, Nitesh V. Chawla, Hanqi Guo, Gokhan Sever, and Seung Hyun Kim. Exploring time-varying multivariate volume data using matrix of isosurface similarity maps. IEEE Transactions on Visualization and Computer Graphics, 25(1):1236– 1245, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440120-
abs.html.

Taher:2017:IUD

[TJW^{+}17] Faisal Taher, Yvonne Jansen, Jonathan Woodruff, John Hardy, Kasper Hornbæk, and Jason Alexander. Investigating the use of a dynamic physical bar chart for data exploration and presentation. IEEE Transactions on Visualization and Computer Graphics, 23 (1):451–460, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tamm:2014:HRS

[TK14] Georg Tamm and Jens Kruger. Hybrid rendering with scheduling under uncertainty. IEEE Transactions on Visualization and Computer Graphics, 20(5):767–780, May 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tory:2006:VTP

[TKAM06] Melanie Tory, Arthur E. Kirkpatrick, M. Stella Atkins, and Torsten Möller. Visualization task performance with 2D, 3D, and combination displays. IEEE Transactions on Visualization and Computer Graphics, 12(1):2–13, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Turkay:2017:DPI

[TKBH17] Cagatay Turkay, Erdem Kaya, Selim Balcisoy, and Helwig Hauser. Designing progressive and interactive analytics processes for high-dimensional data analysis. IEEE Transactions on Visualization and Computer Graphics, 23(1): 131–140, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tam:2017:AMH

[TKC17] Gary K. L. Tam, Vivek Kothari, and Min Chen. An analysis of machine- and human-analytics in classification. IEEE Transactions on Visualization and Computer Graphics, 23(1):71–80, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thom:2016:CTS

[TKE16] Dennis Thom, Robert Krüger, and Thomas Ertl. Can Twitter save lives? A broadscale study on visual social media analytics for public safety. IEEE Transactions on Visualization and Computer Graphics, 22(7):1816– 1829, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Taguchi:2009:TLT

[TKTN09] Yuichi Taguchi, Takafumi Koike, Keita Takahashi, and Takeshi Naemura. TransCAIP: a live 3D TV system using a camera array and an integral photography display with interactive control of viewing parameters. IEEE Transactions on Visualization and Computer Graphics, 15(5):841–852, September/October 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tricoche:2008:ICL

[TKW08] Xavier Tricoche, Gordon Kindlmann, and Carl-Fredrik Westin. Invariant crease lines for topological and structural analysis of tensor fields. IEEE Transactions on Visualization and Computer Graphics, 14(6):1627–1634, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tam:2007:DMR

[TL07] Gary K. L. Tam and Rynson W. H. Lau. Deformable model retrieval based on topological and geometric signatures. IEEE Transactions on Visualization and Computer Graphics, 13(3):470–482, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thalmann:2011:GEI

[TL11] D. Thalmann and B. Lok. Guest Editors’ introduction: Special section on the ACM Symposium on Virtual Reality and Software Technology (VRST 2009). IEEE Transactions on Visualization and Computer Graphics, 17 (7):873–874, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tsai:2010:RTP

[TLC^{+}10] Yao-Yang Tsai, Wen-Chieh Lin, Kuangyou B. Cheng, Jehee Lee, and Tong-Yee Lee. Real-time physics-based 3D biped character animation using an inverted pendulum model. IEEE Transactions on Visualization and Computer Graphics, 16(2):325– 337, March/April 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2012:SAL

[TLD^{+}12] Yubo Tao, Hai Lin, Feng Dong, Chao Wang, Gordon Clapworthy, and Hujun Bao. Structure-aware lighting design for volume visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2372– 2381, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Talbot:2010:EWA

[TLH10] Justin Talbot, Sharon Lin, and Pat Hanrahan. An extension of Wilkinson’s algorithm for positioning tick labels on axes. IEEE Transactions on Visualization and Computer Graphics, 16(6):1036– 1043, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Turkay:2012:RFG

[TLLH12] Cagatay Turkay, Arvid Lundervold, Astri Johansen Lundervold, and Helwig Hauser. Representative factor generation for the interactive visual analysis of high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 18(12):2621– 2630, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tzeng:2005:ISA

[TLM05] Fan-Yin Tzeng, Eric B. Lum, and Kwan-Liu Ma. An intelligent system approach to higher-dimensional classification of volume data. IEEE Transactions on Visualization and Computer Graphics, 11 (3):273–284, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tan:2008:FRR

[TLQ^{+}08] Ping Tan, Stephen Lin, Long Quan, Baining Guo, and Harry Shum. Filtering and rendering of resolutiondependent reflectance models. IEEE Transactions on Visualization and Computer Graphics, 14(2):412– 425, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tong:2017:GVD

[TLS17] Xin Tong, Cheng Li, and Han-Wei Shen. GlyphLens: Viewdependent occlusion management in the interactive glyph visualization. IEEE Transactions on Visualization and Computer Graphics, 23(1): 891–900, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tory:2004:HFV

[TM04] Melanie Tory and Torsten Möller. Human factors in visualization research. IEEE Transactions on Visualization and Computer Graphics, 10(1):72–84, January/ February 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0072abs.

htm;   http://csdl.computer.
org/dl/trans/tg/2004/01/
v0072.pdf.

Tanahashi:2012:DCO

[TM12] Yuzuru Tanahashi and Kwan-Liu Ma. Design considerations for optimizing storyline visualizations. IEEE Transactions on Visualization and Computer Graphics, 18(12):2679– 2688, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Talvas:2015:ACV

[TMDO15] A. Talvas, M. Marchal, C. Duriez, and M. A. Otaduy. Aggregate constraints for virtual manipulation with soft fingers. IEEE Transactions on Visualization and Computer Graphics, 21(4):452– 461, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tateosian:2010:TTG

[TMH^{+}10] Laura Tateosian, Helena Mitasova, Brendan Harmon, Brent Fogleman, Katherine Weaver, and Russel Harmon. TanGeoMS: Tangible geospatial modeling system. IEEE Transactions on Visualization and Computer Graphics, 16(6):1605–1612, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Terziman:2013:PMC

[TMM^{+}13] Leo Terziman, Maud Marchal, Franck Multon, Bruno Arnaldi, and Anatole Lecuyer. Personified and multistate camera motions for firstperson navigation in desktop virtual reality. IEEE Transactions on Visualization and Computer Graphics, 19(4): 652–661, April 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2013:UAS

[TMWS13] Jun Tao, Jun Ma, Chaoli Wang, and Ching-Kuang Shene. A unified approach to streamline selection and viewpoint selection for 3D flow visualization. IEEE Transactions on Visualization and Computer Graphics, 19(3): 393–406, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thomas:2011:SSF

[TN11] Dilip Mathew Thomas and Vijay Natarajan. Symmetry in scalar field topology. IEEE Transactions on Visualization and Computer Graphics, 17(12):2035– 2044, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thomas:2013:DSS

[TN13] Dilip Mathew Thomas and Vijay Natarajan. Detecting symmetry in scalar fields using augmented extremum graphs. IEEE Transactions on Visualization and Computer Graphics, 19(12):2663– 2672, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Thomas:2014:MSD

[TN14] Dilip Mathew Thomas and Vijay Natarajan. Multiscale symmetry detection in scalar fields by clustering contours. IEEE Transactions on Visualization and Computer Graphics, 20(12):2427– 2436, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875976-
abs.html.

Thomas:2011:LCS

[TNB11] D. M. Thomas, V. Natarajan, and G. P. Bonneau. Link conditions for simplifying meshes with embedded structures. IEEE Transactions on Visualization and Computer Graphics, 17(7): 1007–1019, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tuttle:2010:PSS

[TNS10] Claurissa Tuttle, Luis Gustavo Nonato, and Claudio Silva. PedVis: a structured, space-efficient technique for pedigree visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1063–1072, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tan:2017:LBS

[TNT17] David Joseph Tan, Nassir Navab, and Federico Tombari. Looking beyond the simple scenarios: Combining learners and optimizers in 3D temporal tracking. IEEE Transactions on Visualization and Computer Graphics, 23(11): 2399–2409, November 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007238-
abs.html.

Tierny:2012:GTS

[TP12] Julien Tierny and Valerio Pascucci. Generalized topological simplification of scalar fields on surfaces. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2005–2013, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Trimm:2012:VSH

[TRd12] David Trimm, Penny Rheingans, and Marie desJardins. Visualizing student histories using clustering and composition. IEEE Transactions on Visualization and Computer Graphics, 18(12):2809– 2818, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tang:2019:IEC

[TRL^{+}19] Tan Tang, Sadia Rubab, Jiewen Lai, Weiwei Cui, Lingyun Yu, and Yingcai Wu. iStoryline: Effective convergence to hand-drawn storylines. IEEE Transactions on Visualization and Computer Graphics, 25(1):769– 778, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440828-
abs.html.

Tu:2007:VCH

[TS07] Ying Tu and Han-Wei Shen.

Visualizing changes of hierarchical data using treemaps. IEEE Transactions on Visualization and Computer Graphics, 13(6):1286–1293, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tu:2008:BFS

[TS08] Ying Tu and Han-Wei Shen. Balloon focus: a seamless multi-focus+context method for treemaps. IEEE Transactions on Visualization and Computer Graphics, 14(6): 1157–1164, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Talbot:2014:FEP

[TSA14] Justin Talbot, Vidya Setlur, and Anushka Anand. Four experiments on the perception of bar charts. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2152–2160, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876021-
                                        abs.html.

Tominski:2012:SBV

[TSAA12] Christian Tominski, Heidrun Schumann, Gennady Andrienko, and Natalia Andrienko. Stacking-based visualization of trajectory attribute data. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2565–2574, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Teran:2005:CSS

[TSB^{+}05] Joseph Teran, Eftychios Sifakis, Silvia S. Blemker, Victor Ng-Thow-Hing, Cynthia Lau, and Ronald Fedkiw. Creating and simulating skeletal muscle from the visible human data set. IEEE Transactions on Visualization and Computer Graphics, 11(3):317– 328, May/June 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tory:2009:CDL

[TSD09] Melanie Tory, Colin Swindells, and Rebecca Dreezer. Comparing dot and landscape spatializations for visual memory differences. IEEE Transactions on Visualization and Computer Graphics, 15(6):1033–1040, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Turkay:2014:ASD

[TSH^{+}14] Cagatay Turkay, Aidan Slingsby, Helwig Hauser, Jo Wood, and Jason Dykes. Attribute signatures: Dynamic visual summaries for analyzing multivariate geographical data. IEEE Transactions on Visualization and Computer Graphics, 20(12):2033– 2042, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/

                                        csdl/trans/tg/2014/12/06875987-
                                        abs.html.

Tan:2007:AAT

[TSLR07] Desney Tan, Greg Smith, Bongshin Lee, and George Robertson. AdaptiviTree: Adaptive tree visualization for tournament-style brackets. IEEE Transactions on Visualization and Computer Graphics, 13(6):1113–1120, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1113s.
                                        mov.

Tory:2007:SDC

[TSW^{+}07] Melanie Tory, David Sprague, Fuqu Wu, Wing Yan So, and Tamara Munzner. Spatialization design: Comparing points and landscapes. IEEE Transactions on Visualization and Computer Graphics, 13(6):1262–1269, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tanaka:2005:TOV

[TT05] Kenji Tanaka and Susumu Tachi. TORNADO: Omnistereo video imaging with rotating optics. IEEE Transactions on Visualization and Computer Graphics, 11(6):614–625, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/06/v0614s.
zip.

Teets:2010:UCF

[TTR10] Jay M. Teets, David P. Tegarden, and Roberta S. Russell. Using cognitive fit theory to evaluate the effectiveness of information visualizations: an example using quality assurance data. IEEE Transactions on Visualization and Computer Graphics, 16(5): 841–853, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tsang:2010:EVS

[TTS10] Hoi Ying Tsang, Melanie Tory, and Colin Swindells. eSeeTrack — visualizing sequential fixation patterns. IEEE Transactions on Visualization and Computer Graphics, 16(6):953–962, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Turban:2017:EVE

[TUG17] Laura Turban, Fabrice Urban, and Philippe Guillotel. Extrafoveal video extension for an immersive viewing experience. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1520–1533, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07404057-
abs.html.

Toet:2014:PVU

[TvET14] Alexander Toet, Jan van Erp, and Susanne Tak. The perception of visual uncertainty representation by non-experts. IEEE Transactions on Visualization and Computer Graphics, 20(6):935–943, June 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2018:SAG

[TW18] Jun Tao and Chaoli Wang. Semi-automatic generation of stream surfaces via sketching. IEEE Transactions on Visualization and Computer Graphics, 24(9):2622– 2635, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08038807-
abs.html.

Torsney-Weir:2017:PIR

[TWBBM17] Thomas Torsney-Weir, Steven Bergner, Derek Bingham, and Torsten Moller. Predicting the interactive rendering time threshold of Gaussian process models with HyperSlice. IEEE Transactions on Visualization and Computer Graphics, 23(2):1111– 1123, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07414480-
abs.html.

Tao:2018:SFG

[TWC^{+}18] Jun Tao, Chaoli Wang, Nitesh V. Chawla, Lei Shi, and Seung Hyun Kim. Semantic flow graph: A framework for discovering object relationships in flow fields. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3200–3213, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8118164/.

Theisel:2005:TMT

[TWHS05] Holger Theisel, Tino Weinkauf, Hans-Christian Hege, and Hans-Peter Seidel. Topological methods for 2D timedependent vector fields based on stream lines and path lines. IEEE Transactions on Visualization and Computer Graphics, 11(4):383– 394, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tian:2019:RHO

[TWMZ19] H. Tian, C. Wang, D. Manocha, and X. Zhang. Realtime hand-object interaction using learned grasp space for virtual environments. IEEE Transactions on Visualization and Computer Graphics, 25(8): 2623–2635, August 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2014:DFF

[TWSK14] Jun Tao, Chaoli Wang, Ching-Kuang Shene, and Seung Hyun Kim. A deformation framework for focus+context flow visualization. IEEE Transactions on Visualization and Computer Graphics, 20(1):42–55, January 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Torsney-Weir:2011:TPP

[TWSM^{+}11] Thomas Torsney-Weir, Ahmed Saad, Torsten Moller, Hans-Christian Hege, Britta Weber, and Jean-Marc Verbavatz. Tuner: Principled parameter finding for image segmentation algorithms using visual response surface exploration. IEEE Transactions on Visualization and Computer Graphics, 17(12):1892– 1901, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tao:2016:VAP

[TWSS16] J. Tao, C. Wang, C. K. Shene, and R. A. Shaw. A vocabulary approach to partial streamline matching and exploratory flow visualization. IEEE Transactions on Visualization and Computer Graphics, 22 (5):1503–1516, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tong:2018:MLH

[TYL^{+}18] Qianqian Tong, Zhiyong Yuan, Xiangyun Liao, Mianlun Zheng, Tianchen Yuan, and Jianhui Zhao. Magnetic levitation haptic augmentation for virtual tissue stiffness perception. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3123–3136, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8103791/.

Takahashi:2006:OFA

[TYSN06] Shigeo Takahashi, Kenichi Yoshida, Kenji Shimada, and Tomoyuki Nishita. Occlusionfree animation of driving routes for car navigation systems. IEEE Transactions on Visualization and Computer Graphics, 12(5):1141– 1148, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tong:2013:SIC

[TZC13] Ruo-Feng Tong, Yun Zhang, and Ke-Li Cheng. StereoPasting: Interactive composition in stereoscopic images. IEEE Transactions on Visualization and Computer Graphics, 19 (8):1375–1385, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Tong:2012:SFH

[TZL^{+}12] Jing Tong, Jin Zhou, Ligang Liu, Zhigeng Pan, and Hao Yan. Scanning 3D full human bodies using kinects. IEEE Transactions on Visualization and Computer Graphics, 18 (4):643–650, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ujitoko:2019:MFR

[UBH19] Yusuke Ujitoko, Yuki Ban, and Koichi Hirota. Modulating fine roughness perception of vibrotactile textured surface using pseudohaptic effect. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1981–1990, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642446/.

Unger:2018:USS

[UDSL18] A. Unger, N. Dräger, M. Sips, and D. J. Lehmann. Understanding a sequence of sequences: Visual exploration of categorical states in lake sediment cores. IEEE Transactions on Visualization and Computer Graphics, 24(1):66– 76, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ullrich:2012:HPM

[UK12] Sebastian Ullrich and Torsten Kuhlen. Haptic palpation for medical simulation in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 18 (4):617–625, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Usher:2018:VRV

[UKF^{+}18] W. Usher, P. Klacansky, F. Federer, P. T. Bremer, A. Knoll, J. Yarch, A. Angelucci, and V. Pascucci. A virtual reality visualization tool for neuron tracing. IEEE Transactions on Visualization and Computer Graphics, 24 (1):994–1003, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Uijong:2019:YMP

[UKW19] Ju Uijong, June Kang, and Christian Wallraven. You or me? Personality traits predict sacrificial decisions in an accident situation. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1898–1907, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8651471/.

Ushizima:2012:ATD

[UMW^{+}12] Daniela Ushizima, Dmitriy Morozov, Gunther H. Weber, Andrea G. C. Bianchi, James A. Sethian, and E. Wes Bethel. Augmented topological descriptors of pore networks for material science. IEEE Transactions on Visualization and Computer Graphics, 18(12):2041– 2050, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Uffinger:2013:TDV

[USE13] Markus Uffinger, Filip Sadlo, and Thomas Ertl. A timedependent vector field topology based on streak surfaces. IEEE Transactions on Visualization and Computer Graphics, 19(3):379– 392, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Unger:2012:VAC

[USKD12] Andrea Unger, Sven Schulte, Volker Klemann, and Doris Dransch. A visual analysis concept for the validation of geoscientific simulation models. IEEE Transactions on Visualization and Computer Graphics, 18(12):2216– 2225, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ueng:1996:ESS

[USM96] Shyh-Kuang Ueng, Christopher Sikorski, and Kwan-Liu Ma. Efficient streamline, streamribbon, and streamtube constructions on unstructured grids. IEEE Transactions on Visualization and Computer Graphics, 2(2): 100–110, June 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0100.
pdf;   http://www.computer.
org/tvcg/tg1996/v0100abs.
htm.

Ueng:1997:CSV

[USM97] Shyh-Kuang Ueng, C. Sikorski, and Kwan-Liu Ma. Outof-core streamline visualization on large unstructured meshes. IEEE Transactions on Visualization and Computer Graphics, 3(4):370–380, October/December 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0370.
pdf;   http://www.computer.
org/tvcg/tg1997/v0370abs.
htm.

Vanegas:2012:AEM

[VAB12] Carlos A. Vanegas, Daniel G. Aliaga, and Bedrich Benes. Automatic extraction of Manhattan-world building masses from 3D laser range scans. IEEE Transactions on Visualization and Computer Graphics, 18(10):1627– 1637, October 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vanegas:2009:VSU

[VABW09] Carlos A. Vanegas, Daniel G. Aliaga, Bedrich Benes, and Paul Waddell. Visualization of simulated urban spaces: Inferring parameterized generation of streets, parcels, and aerial imagery. IEEE Transactions on Visualization and Computer Graphics, 15 (3):424–435, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanAlmsick:2011:GBR

[vAPP^{+}11] Markus van Almsick, Tim H. J. M. Peeters, Vesna Prčkovska, Anna Vilanova, and Bart ter Haar Romeny. GPU-based ray-casting of spherical functions applied to high angular resolution diffusion imaging. IEEE Transactions on Visualization and Computer Graphics, 17(5): 612–625, May 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Varshney:2001:GEI

[Var01] Amitabh Varshney. Guest Editor’s introduction: Special issue on Visualization 2000. IEEE Transactions on Visualization and Computer Graphics, 7(3):193–194, July 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0193.
pdf;   http://www.computer.
org/tvcg/tg2001/v0193abs.
htm.

Ventura:2014:GLM

[VARS14] Jonathan Ventura, Clemens Arth, Gerhard Reitmayr, and Dieter Schmalstieg. Global localization from monocular SLAM on a mobile phone. IEEE Transactions on Visualization and Computer Graphics, 20(4):531–539, April 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vasa:2016:EIM

[Vas16] Libor Vasa. On the efficiency of image metrics for evaluating the visual quality of 3D models. IEEE Transactions on Visualization and Computer Graphics, 22(8):1987– 1999, August 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vantzos:2017:FTF

[VAW^{+}17] Orestis Vantzos, Omri Azencot, Max Wardeztky, Martin Rumpf, and Mirela Ben-Chen. Functional thin films on surfaces. IEEE Transactions on Visualization and Computer Graphics, 23(3): 1179–1192, March 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/03/07558232-
                                        abs.html.

Vasa:2013:ECI

[VB13] Libor Vasa and Guido Brunnett. Exploiting connectivity to improve the tangential part of geometry prediction. IEEE Transactions on Visualization and Computer Graphics, 19(9):1467–1475, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vuillemot:2018:SVM

[VB18] R. Vuillemot and J. Boy. Structuring visualization mockups at the graphical level by dividing the display space. IEEE Transactions on Visualization and Computer Graphics, 24(1):424–434, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Volante:2016:EVH

[VBC^{+}16] M. Volante, S. V. Babu, H. Chaturvedi, N. Newsome, E. Ebrahimi, T. Roy, S. B. Daily, and T. Fasolino. Effects of virtual human appearance fidelity on emotion contagion in affective inter-personal simulations. IEEE Transactions on Visualization and Computer Graphics, 22(4):1326– 1335, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vaillant:2017:MCP

[VBK17] Joris Vaillant, Karim Bouyarmane, and Abderrahmane Kheddar. Multi-character physical and behavioral interactions controller. IEEE Transactions on Visualization and Computer Graphics, 23(6):1650–1662, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/06/07433460.
html.

Volmer:2018:CPS

[VBV^{+}18] Benjamin Volmer, James Baumeister, Stewart Von Itzstein, Ina Bornkessel-Schlesewsky, Matthias Schlesewsky, Mark Billinghurst, and Bruce H. Thomas. A comparison of predictive spatial augmented reality cues for procedural tasks. IEEE Transactions on Visualization and Computer Graphics, 24(11):2846– 2856, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8493594/.

Vehlow:2016:VDH

[VBW16] Corinna Vehlow, Fabian Beck, and Daniel Weiskopf. Visualizing dynamic hierarchies in graph sequences. IEEE Transactions on Visualization and Computer Graphics, 22(10): 2343–2357, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Veras:2017:OHV

[VC17] Rafael Veras and Christopher Collins. Optimizing hierarchical visualizations with the minimum description length principle. IEEE Transactions on Visualization and Computer Graphics, 23(1):631– 640, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vo:2007:SST

[VCL^{+}07] Huy T. Vo, Steven P. Callahan, Peter Lindstrom, Valerio Pascucci, and Cláudio T. Silva. Streaming simplification of tetrahedral meshes. IEEE Transactions on Visualization and Computer Graphics, 13(1):145–155, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Valette:2008:GRT

[VCP08] Sebastien Valette, Jean Marc Chassery, and Rémy Prost. Generic remeshing of 3D triangular meshes with metricdependent discrete Voronoi diagrams. IEEE Transactions on Visualization and Computer Graphics, 14(2):369– 381, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanderCorput:2014:EPM

[vdCvW14] Paul van der Corput and Jarke J. van Wijk. Effects of presentation mode and pace control on performance in image classification. IEEE Transactions on Visualization and Computer Graphics, 20(12):2301– 2309, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875979-
abs.html.

vandenElzen:2016:RSP

[vdEHBV16] S. van den Elzen, D. Holten, J. Blaas, and J. J. Van Wijk. Reducing snapshots to points: A visual analytics approach to dynamic network exploration. IEEE Transactions on Visualization and Computer Graphics, 22(1):1–10, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vandenElzen:2014:DNV

[vdEHBvW14] Stef van den Elzen, Danny Holten, Jorik Blaas, and Jarke J. van Wijk. Dynamic network visualization with extended massive sequence views. IEEE Transactions on Visualization and Computer Graphics, 20(8): 1087–1099, August 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vandenElzen:2014:MNE

[vdEvW14] Stef van den Elzen and Jarke J. van Wijk. Multivariate network exploration and presentation: From detail to overview via selections and aggregations. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2310–2319, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06875972-
                                        abs.html.

vanderZwan:2016:CUR

[vdZCT16] Matthew van der Zwan, Valeriu Codreanu, and Alexandru Telea. CUBu: Universal real-time bundling for large graphs. IEEE Transactions on Visualization and Computer Graphics, 22(12):2550– 2563, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07374742-
abs.html.

Vasilakis:2013:DFA

[VF13]  Andreas-Alexandros  Vasi-

lakis and Ioannis Fudos. Depth-fighting aware methods for multifragment rendering. IEEE Transactions on Visualization and Computer Graphics, 19(6):967–977, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vickers:2013:UVF

[VFR13] Paul Vickers, Joe Faith, and Nick Rossiter. Understanding visualization: A formal approach using category theory and semiotics. IEEE Transactions on Visualization and Computer Graphics, 19(6): 1048–1061, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Viola:2006:IDF

[VFSG06] Ivan Viola, Miquel Feixas, Mateu Sbert, and Meister Eduard Gröller. Importancedriven focus of attention. IEEE Transactions on Visualization and Computer Graphics, 12(5):933–940, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vonFunck:2008:SSI

[vFWTS08] Wolfram von Funck, Tino Weinkauf, Holger Theisel, and Hans-Peter Seidel. Smoke surfaces: an interactive flow visualization technique inspired by real-world flow experiments. IEEE Transactions on Visualization and Computer Graphics, 14(6):1396– 1403, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Veas:2012:EOT

[VGKS12] Eduardo Veas, Raphael Grasset, Ernst Kruijff, and Dieter Schmalstieg. Extended overview techniques for outdoor augmented reality. IEEE Transactions on Visualization and Computer Graphics, 18 (4):565–572, April 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanGoethem:2015:ECS

[vGMSW15] Arthur van Goethem, Wouter Meulemans, Bettina Speckmann, and Jo Wood. Exploring curved schematization of territorial outlines. IEEE Transactions on Visualization and Computer Graphics, 21(8):889–902, August 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
08/07035078-abs.html.

vanGoethem:2014:SSS

[vGRSW14] Arthur van Goethem, Andreas Reimer, Bettina Speckmann, and Jo Wood. Stenomaps: Shorthand for shapes. IEEE Transactions on Visualization and Computer Graphics, 20(12):2053–2062, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876003-
abs.html.

VanderPlas:2016:SRD

[VH16] S. VanderPlas and H. Hofmann. Spatial reasoning and data displays. IEEE Transactions on Visualization and Computer Graphics, 22(1): 459–468, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vinkler:2016:PDH

[VHBS16] Marek Vinkler, Vlastimil Havran, Jiří Bittner, and Jiří Sochor. Parallel on-demand hierarchy construction on contemporary GPUs. IEEE Transactions on Visualization and Computer Graphics, 22 (7):1886–1898, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vines:2014:VIF

[VHLL14] Mauricio Vines, Ben Houston, Jochen Lang, and Won-Sook Lee. Vortical inviscid flows with two-way solidfluid coupling. IEEE Transactions on Visualization and Computer Graphics, 20(2): 303–315, February 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanHam:2011:MPC

[vHMM^{+}11] Frank van Ham, Raghu Machiraju, Klaus Mueller, Gerik Scheuermann, and Chris Weaver. Message from the Paper Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 17(12): x, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanHam:2009:SSC

[vHP09] Frank van Ham and Adam Perer. “Search, Show Context, Expand on Demand”: Supporting large graph exploration with degree-of-interest. IEEE Transactions on Visualization and Computer Graphics, 15(6):953–960, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanHam:2008:POU

[vHR08] Frank van Ham and Bernice Rogowitz. Perceptual organization in usergenerated graph layouts.

IEEE Transactions on Visualization and Computer Graphics, 14(6):1333–1339, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanHam:2002:IVS

[vHvdWvW02] Frank van Ham, Huub van de Wetering, and Jarke J. van Wijk. Interactive visualization of state transition systems. IEEE Transactions on Visualization and Computer Graphics, 8(4):319–329, October/December 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0319abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0319.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0319.pdf.

vanHam:2009:MTP

[vHWV09] Frank van Ham, Martin Wattenberg, and Fernanda B. Viegas. Mapping text with phrase nets. IEEE Transactions on Visualization and Computer Graphics, 15(6):1169– 1176, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Viola:2018:PCA

[VI18] Ivan Viola and Tobias Isenberg. Pondering the concept of abstraction in (illustrative) visualization. IEEE Transactions on Visualization and Computer Graphics, 24(9): 2573–2588, September 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/09/08025425.
                                        pdf.

Visell:2015:FPA

[Vis15] Y. Visell. Fast physically accurate rendering of multimodal signatures of distributed fracture in heterogeneous materials. IEEE Transactions on Visualization and Computer Graphics, 21 (4):443–451, April 2015. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vrotsou:2009:AIV

[VJC09] Katerina Vrotsou, Jimmy Johansson, and Matthew Cooper. ActiviTree: Interactive visual exploration of sequences in event-based data using graph similarity. IEEE Transactions on Visualization and Computer Graphics, 15(6):945–952, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vrotsou:2015:SMS

[VJN^{+}15] Katerina Vrotsou, Halldor Janetzko, Carlo Navarra, Georg Fuchs, David Spretke, Florian Mansmann, Natalia Andrienko, and Gennady Andrienko. SimpliFly: A methodology for simplification and thematic enhancement of trajectories. IEEE Transactions on Visualization and Computer Graphics, 21 (1):107–121, January 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06851202-
abs.html.

Viola:2005:IDF

[VKG05] Ivan Viola, Armin Kanitsar, and M. Eduard Gröller. Importance-driven feature enhancement in volume visualization. IEEE Transactions on Visualization and Computer Graphics, 11(4):408– 418, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vonLandesberger:2016:CLQ

[vLBB16] Tatiana von Landesberger, Dennis Basgier, and Meike Becker. Comparative local quality assessment of 3D medical image segmentations with focus on statistical shape model-based algorithms. IEEE Transactions on Visualization and Computer Graphics, 22(12):2537– 2549, December 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2016/12/07331662-
                                        abs.html.

vonLandesberger:2016:MVA

[vLBR^{+}16] T. von Landesberger, F. Brodkorb, P. Roskosch, N. Andrienko, G. Andrienko, and A. Kerren. MobilityGraphs: Visual analysis of mass mobility dynamics via spatiotemporal graphs and clustering. IEEE Transactions on Visualization and Computer Graphics, 22(1):11–20, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanLiere:2003:GVG

[vLdL03] Robert van Liere and Wim de Leeuw. GraphSplatting: Visualizing graphs as continuous fields. IEEE Transactions on Visualization and Computer Graphics, 9(2):206– 212, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/02/v0206abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/02/
                                        v0206.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2003/02/v0206.pdf.

vonLandesberger:2017:VSR

[vLFR17] Tatiana von Landesberger, Dieter W. Fellner, and Roy A. Ruddle. Visualization system requirements for data processing pipeline design and optimization. IEEE Transactions on Visualization and Computer Graphics, 23(8):2028– 2041, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/08/07552504-
abs.html.

Viau:2010:FPS

[VMCJ10] Christophe Viau, Michael J.

McGuffin, Yves Chiricota, and Igor Jurisica. The FlowVizMenu and parallel scatterplot matrix: Hybrid multidimensional visualizations for network exploration. IEEE Transactions on Visualization and Computer Graphics, 16(6):1100–1108, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vemuri:1997:FGS

[VML97] B. C. Vemuri, C. Mandal, and S.-H. Lai. A fast Gibbs sampler for synthesizing constrained fractals. IEEE Transactions on Visualization and Computer Graphics, 3(4):337– 351, October/December 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0337.
pdf;   http://www.computer.
org/tvcg/tg1997/v0337abs.
htm.

Valsangkar:2019:EFC

[VMN^{+}19] A. A. Valsangkar, J. M. Monteiro, V. Narayanan, I. Hotz, and V. Natarajan. An exploratory framework for cyclone identification and tracking. IEEE Transactions on Visualization and Computer Graphics, 25(3):1460– 1473, March 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Volino:2006:RTA

[VMT06] Pascal Volino and Nadia Magnenat-Thalmann. Realtime animation of complex hairstyles. IEEE Transactions on Visualization and Computer Graphics, 12(2):131– 142, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vrotsou:2019:EVS

[VN19] K. Vrotsou and A. Nordman. Exploratory visual sequence mining based on patterngrowth. IEEE Transactions on Visualization and Computer Graphics, 25(8):2597– 2610, August 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Valette:2004:WBM

[VP04a] Sébastien Valette and Rémy Prost. Wavelet-based multiresolution analysis of irregular surface meshes. IEEE Transactions on Visualization and Computer Graphics, 10(2):113–122, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2004/
02/v0113abs.htm;   http:
//csdl.computer.org/dl/
trans/tg/2004/02/v0113.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0113.pdf.

Valette:2004:WBP

[VP04b] Sébastien Valette and Rémy Prost. Wavelet-based progressive compression scheme for triangle meshes: Wavemesh. IEEE Transactions on Visualization and Computer Graphics, 10(2):123–129, March/ April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2004/
02/v0123abs.htm;   http:
//csdl.computer.org/dl/
trans/tg/2004/02/v0123.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0123.pdf.

Verma:2004:CFV

[VP04c] Vivek Verma and Alex Pang.

Comparative flow visualization. IEEE Transactions on Visualization and Computer Graphics, 10(6):609– 624, November/December 2004. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/comp/trans/tg/2004/
                                        06/v0609abs.htm;   http:
                                        //csdl.computer.org/dl/
                                        trans/tg/2004/06/v0609.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/06/
                                        v0609.pdf.

VanGelder:2009:UPA

[VP09] Allen Van Gelder and Alex Pang. Using PVsolve to analyze and locate positions of parallel vectors. IEEE Transactions on Visualization and Computer Graphics, 15(4): 682–695, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vergne:2011:ISD

[VPB^{+}11] Romain Vergne, Romain Pacanowski, Pascal Barla, Xavier Granier, and Christophe Schlick. Improving shape depiction under arbitrary rendering. IEEE Transactions on Visualization and Computer Graphics, 17(8):1071– 1081, August 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanPelt:2010:EMB

[vPBB^{+}10] Roy van Pelt, Javier Olivan Bescos, Marcel Breeuwer, Rachel E. Clough, M. Eduard Gröller, Bart ter Haar Romenij, and Anna Vilanova. Exploration of 4D MRI blood flow using stylistic visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1339–1347, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanPelt:2011:IVP

[vPBB^{+}11] Roy van Pelt, Javier Olivan Bescos, Marcel Breeuwer, Rachel E. Clough, M. Eduard Gröller, Bart ter Haar Romenij, and Anna Vilanova. Interactive virtual probing of 4D MRI blood-flow. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2153–2162, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vasilakis:2015:BEM

[VPF15] A.-A. Vasilakis, G. Papaioannou, and I. Fudos. k^{+}- buffer: An efficient, memoryfriendly and dynamic k-buffer framework. IEEE Transactions on Visualization and Computer Graphics, 21(6): 688–700, June 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanPelt:2010:IVV

[vPVvdW10] Roy van Pelt, Anna Vilanova, and Huub van de Wetering. Illustrative volume visualization using GPU-based particle systems. IEEE Transactions on Visualization and Computer Graphics, 16(4):571– 582, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vonRadziewsky:2017:ESR

[vRKEE17] Philipp von Radziewsky, Thomas Kroes, Martin Eisemann, and Elmar Eisemann. Efficient stochastic rendering of static and animated volumes using visibility sweeps. IEEE Transactions on Visualization and Computer Graphics, 23(9):2069– 2081, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/09/07562447-
                                        abs.html.

Vehlow:2013:VFO

[VRW13] Corinna Vehlow, Thomas Reinhardt, and Daniel Weiskopf. Visualizing fuzzy overlapping communities in networks. IEEE Transactions on Visualization and Computer Graphics, 19(12):2486– 2495, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vasa:2011:PCC

[VS11] Libor Vasa and Vaclav Skala. A perception correlated comparison method for dynamic meshes. IEEE Transactions on Visualization and Computer Graphics, 17(2):220– 230, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vinayagamoorthy:2008:ICP

[VSS08] Vinoba Vinayagamoorthy, Anthony Steed, and Mel Slater. The impact of a character posture model on the communication of affect in an immersive virtual environment. IEEE Transactions on Visualization and Computer Graphics, 14(5): 965–982, September/October 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

VanHorn:2008:APS

[VT08] R. Brooks Van Horn III and Greg Turk. Antialiasing procedural shaders with reduction maps. IEEE Transactions on Visualization and Computer Graphics, 14(3): 539–550, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/03/ttg2008030539s.
zip.

vanTeylingen:1997:VDV

[vTRvdM97] R. van Teylingen, W. Ribarsky, and C. van der Mast. Virtual data visualizer. IEEE Transactions on Visualization and Computer Graphics, 3(1): 65–74, January/March 1997. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1997/pdf/v0065.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1997/v0065abs.
                                        htm.

Vliegen:2006:VBD

[VvWvdL06] Roel Vliegen, Jarke J. van Wijk, and Erik-Jan van der Linden. Visualizing business data with generalized treemaps. IEEE Transactions on Visualization and Computer Graphics, 12(5): 789–796, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanWijk:2006:VV

[vW06] Jarke J. van Wijk. Views on visualization. IEEE Transactions on Visualization and Computer Graphics, 12(4): 421–433, July/August 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanWijk:2011:GEI

[vW11] Jarke J. van Wijk. Guest Editor’s introduction: Special section on the IEEE Symposium on Visual Analytics Science and Technology (VAST). IEEE Transactions on Visualization and Computer Graphics, 17(5):555–556, May 2011.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Varshney:2012:GEI

[VW12] Amitabh Varshney and Chris Wyman. Guest Editors’ introduction: Special section on the symposium on interactive 3D graphics and games (I3D). IEEE Transactions on Visualization and Computer Graphics, 18(10):1589–1590, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanWijk:2014:VRM

[vW14] Jarke J. van Wijk. Visualization of regular maps: The chase continues. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2614–2623, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06887357-
abs.html.

vanWijk:2006:VSS

[vWC06] Jarke J. van Wijk and Arjeh M. Cohen. Visualization of Seifert surfaces. IEEE Transactions on Visualization and Computer Graphics, 12(4): 485–496, July/August 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Viegas:2009:PVW

[VWF09] Fernanda B. Viegas, Martin Wattenberg, and Jonathan Feinberg. Participatory visualization with Wordle. IEEE Transactions on Visualization and Computer Graphics, 15(6):1137–1144, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

vanWijk:2004:GEI

[vWMT04] Jarke J. van Wijk, Robert J. Moorhead II, and Greg Turk. Guest Editors’ introduction: Special section on IEEE Visualization. IEEE Transactions on Visualization and Computer Graphics, 10(4):369– ??, July/August 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/04/v0369.
pdf;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0369.htm.

vanWijk:2004:MSV

[vWN04] Jarke J. van Wijk and Wim A. A. Nuij. A model for smooth viewing and navigation of large 2D information spaces. IEEE Transactions on Visualization and Computer Graphics, 10(4):447– 458, July/August 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
dl/trans/tg/2004/04/v0447.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/04/
v0447.pdf.

vanWalsum:1996:FEI

[vWPSP96] Theo van Walsum, Frits H.

Post, Deborah Silver, and Frank J. Post. Feature extraction and iconic visualization. IEEE Transactions on Visualization and Computer Graphics, 2(2):111–119, June 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1996/pdf/v0111.
pdf;   http://www.computer.
org/tvcg/tg1996/v0111abs.
htm.

Viegas:2007:MSV

[VWvH^{+}07] Fernanda B. Viegas, Martin Wattenberg, Frank van Ham, Jesse Kriss, and Matt McKeon. ManyEyes: a site for visualization at Internet scale. IEEE Transactions on Visualization and Computer Graphics, 13(6):1121–1128, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Valdez:2018:PAE

[VZS18] A. C. Valdez, M. Ziefle, and M. Sedlmair. Priming and anchoring effects in visualization. IEEE Transactions on Visualization and Computer Graphics, 24(1):584–594, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wilkinson:2006:HDV

[WAG06] Leland Wilkinson, Anushka Anand, and Robert Grossman. High-dimensional visual analytics: Interactive exploration guided by pairwise views of point distributions. IEEE Transactions on Visualization and Computer Graphics, 12(6):1363–1372, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wenger:2012:VAN

[WAG^{+}12] Stephan Wenger, Marco Ament, Stefan Guthe, Dirk Lorenz, Andreas Tillmann, Daniel Weiskopf, and Marcus Magnor. Visualization of astronomical nebulae via distributed multi-GPU compressed sensing tomography. IEEE Transactions on Visualization and Computer Graphics, 18(12):2188– 2197, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wah:2014:VCM

[Wah14] Benjamin Wah. On view consistency in multi-server distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 20(10):1428– 1440, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06636307-
abs.html.

Wald:2012:FCS

[Wal12] Ingo Wald. Fast construction of SAH BVHs on the Intel Many Integrated Core (MIC) architecture. IEEE Transactions on Visualization and Computer Graphics, 18 (1):47–57, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wall:2019:HAV

[WAM^{+}19] Emily Wall, Meeshu Agnihotri, Laura Matzen, Kristin Divis, Michael Haass, Alex Endert, and John Stasko. A heuristic approach to valuedriven evaluation of visualizations. IEEE Transactions on Visualization and Computer Graphics, 25(1):491– 500, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08454343-
abs.html.

Wang:2006:BRS

[Wan06] Charlie C. L. Wang. Bilateral recovering of sharp edges on feature-insensitive sampled meshes. IEEE Transactions on Visualization and Computer Graphics, 12(4):629– 639, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:CLP

[Wan08] Charlie Wang. Computing length-preserved free boundary for quasi-developable mesh segmentation. IEEE Transactions on Visualization and Computer Graphics, 14 (1):25–36, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2011:ABO

[Wan11] Charlie C. L. Wang. Approximate Boolean operations on large polyhedral solids with partial mesh reconstruction. IEEE Transactions on Visualization and Computer Graphics, 17(6):836–849, June 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ahn:2014:TTN

[wAPS14] Jae wook Ahn, Catherine Plaisant, and Ben Shneiderman. A task taxonomy for network evolution analysis. IEEE Transactions on Visualization and Computer Graphics, 20(3):365– 376, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ware:2009:QTS

[War09] Colin Ware. Quantitative texton sequences for legible bivariate maps. IEEE Transactions on Visualization and Computer Graphics, 15(6):1523–1530, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:VFD

[WASQ18] Yong Wang, Daniel Archambault, Carlos E. Scheidegger, and Huamin Qu. A vector field design approach to animated transitions. IEEE Transactions on Visualization and Computer Graphics, 24(9):2487– 2500, September 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08031015-
abs.html.

Wu:2018:IOA

[WAWS18] J. Wu, N. Aage, R. Westermann, and O. Sigmund. Infill optimization for additive manufacturing — approaching bone-like porous structures. IEEE Transactions on Visualization and Computer Graphics, 24(2):1127– 1140, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wong:2005:DIB

[WB05] Wingo Sai-Keung Wong and George Baciu. Dynamic interaction between deformable surfaces and nonsmooth objects. IEEE Transactions on Visualization and Computer Graphics, 11(3):329– 340, May/June 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weigle:2008:CPB

[WB08] Chris Weigle and David Banks. A comparison of the perceptual benefits of linear perspective and physicallybased illumination for display of dense 3D streamtubes. IEEE Transactions on Visualization and Computer Graphics, 14(6):1723–1730, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2014:EVH

[WBA^{+}14] Yanxiang Wu, Sabarish V. Babu, Rowan Armstrong, Jeffrey W. Bertrand, Jun Luo, Tania Roy, Shaundra B. Daily, Lauren Cairco Dukes, Larry F. Hodges, and Tracy Fasolino. Effects of virtual human animation on emotion contagion in simulated interpersonal experiences. IEEE Transactions on Visualization and Computer Graphics, 20 (4):626–635, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wood:2014:MBS

[WBD14] Jo Wood, Roger Beecham, and Jason Dykes. Moving beyond sequential design: Reflections on a rich multichannel approach to data visualization. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2171–2180, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06875966-
abs.html.

Wood:2011:BDN

[WBDS11] Jo Wood, Donia Badawood, Jason Dykes, and Aidan Slingsby. BallotMaps: Detecting name bias in alphabetically ordered ballot papers. IEEE Transactions on Visualization and Computer Graphics, 17(12):2384– 2391, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiskopf:2006:EIV

[WBE^{+}06] Daniel Weiskopf, Marc Borchers, Thomas Ertl, Martin Falk, Oliver Fechtig, Regine Frank, Frank Grave, Andreas King, Ute Kraus, Thomas Muller, Hans-Peter Nollert, Isabel Rica Mendez, Hanns Ruder, Tobias Schafhitzel, Sonja Schar, Corvin Zahn, and Michael Zatloukal. Explanatory and illustrative visualization of special and general relativity. IEEE Transactions on Visualization and Computer Graphics, 12(4):522– 534, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wiley:2004:CHB

[WBH04] David F. Wiley, Martin Bertram, and Bernd Hamann. On a construction of a hierarchy of best linear spline approximations using a finite element approach. IEEE Transactions on Visualization and Computer Graphics, 10(5):548–563, September/October 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/05/v0548s.
pdf;   http://csdl.computer.
org/dl/trans/tg/2004/05/
v0548.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/05/v0548.pdf.

Walker:2016:TSE

[WBJ16] J. Walker, R. Borgo, and M. W. Jones. TimeNotes: A study on effective chart visualization and interaction techniques for time-series data. IEEE Transactions on Visualization and Computer Graphics, 22(1):549–558, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ward:2007:SHM

[WBK^{+}07] Kelly Ward, Florence Bertails, Tae-Yong Kim, Stephen R. Marschner, Marie-Paule Cani, and Ming C. Lin. A survey on hair modeling: Styling, simulation, and rendering. IEEE Transactions on Visualization and Computer Graphics, 13(2):213–234, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:EVP

[WBK^{+}08] Yi Wang, Doug Bowman, David Krum, Enylton Coalho, Tonya Smith-Jackson, David Bailey, Sarah Peck, Swethan Anand, Trevor Kennedy, and Yernar Abdrazakov. Effects of video placement and spatial context presentation on path reconstruction tasks with contextualized videos. IEEE Transactions on Visualization and Computer Graphics, 14(6):1755–1762, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weber:2007:TLT

[WBP07] Gunther Weber, Peer-Timo Bremer, and Valerio Pascucci. Topological landscapes: a terrain metaphor for scientific data. IEEE Transactions on Visualization and Computer Graphics, 13(6):1416–1423, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1416s.
                                        zip.

Wang:2009:IEV

[WC09] Cong Wang and Yi-Jen Chiang. Isosurface extraction and view-dependent filtering from time-varying fields using persistent time-octree (PTOT). IEEE Transactions on Visualization and Computer Graphics, 15(6):1367–1374, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See errata [WC10].

Wang:2010:EIE

[WC10] Cong Wang and Yi-Jen Chiang. Errata to “Isosurface Extraction and View-Dependent Filtering from Time-Varying Fields Using Persistent Time-Octree (PTOT)”. IEEE Transactions on Visualization and Computer Graphics, 16(2):350–351, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See [WC09].

Wang:2011:FMM

[WC11] Yu-Shuen Wang and Ming-Te Chi. Focus+context metro maps. IEEE Transactions on Visualization and Computer Graphics, 17(12):2528– 2535, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2013:VAD

[WC13] Yu-Ting Wu and Yung-Yu Chuang. VisibilityCluster: Average directional visibility for many-light rendering. IEEE Transactions on Visualization and Computer Graphics, 19(9):1566–1578, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2017:EGS

[WCA^{+}17] Yanhong Wu, Nan Cao, Daniel Archambault, Qiaomu Shen, Huamin Qu, and Weiwei Cui. Evaluation of graph sampling: A visualization perspective. IEEE Transactions on Visualization and Computer Graphics, 23(1):401– 410, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Westerteiger:2012:IRD

[WCB^{+}12] Rolf Westerteiger, Tracy Compton, Tony Bernadin, Eric Cowgill, Klaus Gwinner, Bernd Hamann, Andreas Gerndt, and Hans Hagen. Interactive retro-deformation of terrain for reconstructing 3D fault displacements. IEEE Transactions on Visualization and Computer Graphics, 18(12):2208–2215, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:ECP

[WCB^{+}18] Y. Wang, X. Chu, C. Bao, L. Zhu, O. Deussen, B. Chen, and M. Sedlmair. EdWordle: Consistency-preserving word cloud editing. IEEE Transactions on Visualization and Computer Graphics, 24 (1):647–656, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:UAV

[WCC^{+}18] X. Wang, J. K. Chou, W. Chen, H. Guan, W. Chen, T. Lao, and K. L. Ma. A utility-aware visual approach for anonymizing multiattribute tabular data. IEEE Transactions on Visualization and Computer Graphics, 24 (1):351–360, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:GVI

[WCC^{+}19] Xumeng Wang, Wei Chen, Jia-Kai Chou, Chris Bryan, Huihua Guan, Wenlong Chen, Rusheng Pan, and Kwan-Liu Ma. GraphProtector: A visual interface for employing and assessing multiple privacy preserving graph algorithms. IEEE Transactions on Visualization and Computer Graphics, 25(1):193– 203, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440807-
abs.html.

Weng:2019:STB

[WCD^{+}19] Di Weng, Ran Chen, Zikun Deng, Feiran Wu, Jingmin Chen, and Yingcai Wu. SRVis: Towards better spatial integration in ranking visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 459–469, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08456575-
abs.html.

Wang:2019:OCA

[WCG^{+}19] Yunhai Wang, Xin Chen, Tong Ge, Chen Bao, Michael Sedlmair, Chi-Wing Fu, Oliver Deussen, and Baoquan Chen. Optimizing color assignment for perception of class separability in multiclass scatterplots. IEEE Transactions on Visualization and Computer Graphics, 25(1): 820–829, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2019/01/08440853-
                                        abs.html.

Wickham:2010:GII

[WCHB10] Hadley Wickham, Dianne Cook, Heike Hofmann, and Andreas Buja. Graphical inference for infovis. IEEE Transactions on Visualization and Computer Graphics, 16(6):973–979, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Waters:2006:UDI

[WCJ06] Kenneth W. Waters, Christopher S. Co, and Kenneth I. Joy. Using difference intervals for time-varying isosurface visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1275–1282, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2009:IVO

[WCQ^{+}09] Yingcai Wu, Ka-Kei Chung, Huamin Qu, Xiaoru Yuan, and S. C. Cheung. Interactive visual optimization and analysis for RFID benchmarking. IEEE Transactions on Visualization and Computer Graphics, 15(6):1335–1342, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Walny:2011:VTA

[WCR^{+}11] Jagoda Walny, Sheelagh Carpendale, Nathalie Henry Riche, Gina Venolia, and Philip Fawcett. Visual thinking in action: Visualizations as used on whiteboards. IEEE Transactions on Visualization and Computer Graphics, 17(12):2508–2517, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wenskovitch:2018:TSC

[WCR^{+}18] J. Wenskovitch, I. Crandell, N. Ramakrishnan, L. House, S. Leman, and C. North. Towards a systematic combination of dimension reduction and clustering in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 24(1):131–141, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2018:SMS

[WCS^{+}18] Yingcai Wu, Zhutian Chen, Guodao Sun, Xiao Xie, Nan Cao, Shixia Liu, and Weiwei Cui. StreamExplorer: A multi-stage system for visually exploring events in social streams. IEEE Transactions on Visualization and Computer Graphics, 24(10):2758– 2772, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/10/08074775-
                                        abs.html.

Wen:2016:CRC

[WCW^{+}16] P. C. Wen, W. C. Cheng, Y. S. Wang, H. K. Chu, N. C. Tang, and H. Y. M. Liao. Court reconstruction for camera calibration in broadcast basketball videos. IEEE Transactions on Visualization and Computer Graphics, 22 (5):1517–1526, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2011:EVE

[WCZ^{+}11] Yunhai Wang, Wei Chen, Jian Zhang, Tingxing Dong, Guihua Shan, and Xuebin Chi. Efficient volume exploration using the Gaussian mixture model. IEEE Transactions on Visualization and Computer Graphics, 17(11):1560– 1573, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wood:2008:SOT

[WD08] Jo Wood and Jason Dykes. Spatially ordered treemaps. IEEE Transactions on Visualization and Computer Graphics, 14(6):1348–1355, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiss:2009:SHL

[WD09] Kenneth Weiss and Leila De Floriani. Supercubes: a highlevel primitive for diamond hierarchies. IEEE Transactions on Visualization and Computer Graphics, 15(6):1603– 1610, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiss:2010:IHE

[WD10] Kenneth Weiss and Leila De Floriani. Isodiamond hierarchies: an efficient multiresolution representation for isosurfaces and interval volumes. IEEE Transactions on Visualization and Computer Graphics, 16(4):583– 598, July/August 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weber:2007:TCV

[WDC^{+}07] Gunther H. Weber, Scott E.

Dillard, Hamish Carr, Valerio Pascucci, and Bernd Hamann. Topology-controlled volume rendering. IEEE Transactions on Visualization and Computer Graphics, 13(2): 330–341, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/02/v0330s.
mov.

Wu:2008:DEA

[WDC08] Chunlin Wu, Jiansong Deng, and Falai Chen. Diffusion equations over arbitrary triangulated surfaces for filtering and texture applications. IEEE Transactions on Visualization and Computer Graphics, 14(3):666–679, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wall:2018:PRD

[WDC^{+}18] E. Wall, S. Das, R. Chawla, B. Kalidindi, E. T. Brown, and A. Endert. Podium: Ranking data using mixedinitiative visual analytics. IEEE Transactions on Visualization and Computer Graphics, 24(1):288–297, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wood:2007:IVE

[WDSC07] Jo Wood, Jason Dykes, Aidan Slingsby, and Keith Clarke. Interactive visual exploration of a large spatio-temporal dataset: Reflections on a geovisualization mashup. IEEE Transactions on Visualization and Computer Graphics, 13(6):1176–1183, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2016:SHR

[WDW16] J. Wu, C. Dick, and R. Westermann. A system for highresolution topology optimization. IEEE Transactions on Visualization and Computer Graphics, 22(3):1195– 1208, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weaver:2009:CVF

[Wea09]  Chris Weaver.  Conjunc-

tive visual forms. IEEE Transactions on Visualization and Computer Graphics, 15(6):929–936, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weaver:2010:CFV

[Wea10] Chris Weaver. Cross-filtered views for multidimensional visual analysis. IEEE Transactions on Visualization and Computer Graphics, 16(2): 192–204, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiskopf:2003:ICT

[WEE03] Daniel Weiskopf, Klaus Engel, and Thomas Ertl. Interactive clipping techniques for texture-based volume visualization and volume shading. IEEE Transactions on Visualization and Computer Graphics, 9(3):298–312, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0298abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0298.pdf.

Wenger:2014:EFF

[Wen14] Rephael Wenger. Exploring flow fields using spacefilling analysis of streamlines. IEEE Transactions on Visualization and Computer Graphics, 20(10):1392– 1404, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06767149-
abs.html.

Wong:2006:GSV

[WFC^{+}06] Pak Chung Wong, Harlan Foote, George Chin, Jr., Patrick Mackey, and Ken Perrine. Graph signatures for visual analytics. IEEE Transactions on Visualization and Computer Graphics, 12(6):1399–1413, November/December 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:PDA

[WFC^{+}18] Yunhai Wang, Kang Feng, Xiaowei Chu, Jian Zhang, Chi-Wing Fu, Michael Sedlmair, Xiaohui Yu, and Baoquan Chen. A perceptiondriven approach to supervised dimensionality reduction for visualization. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1828–1840, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07920403-
abs.html.

Weissenbock:2019:DVL

[WFG^{+}19] Johannes Weissenbock, Bernhard Frohler, Eduard Groller, Johann Kastner, and Christoph Heinzl. Dynamic volume lines: Visual comparison of 3D volumes through space-filling curves. IEEE Transactions on Visualization and Computer Graphics, 25(1):1040– 1049, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440038-
abs.html.

Wald:2007:IIR

[WFKH07] Ingo Wald, Heiko Friedrich, Aaron Knoll, and Charles D. Hansen. Interactive isosurface ray tracing of timevarying tetrahedral volumes. IEEE Transactions on Visualization and Computer Graphics, 13(6):1727–1734, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wald:2005:FIR

[WFM^{+}05] Ingo Wald, Heiko Friedrich, Gerd Marmitt, Philipp Slusallek, and Hans-Peter Seidel. Faster isosurface ray tracing using implicit KD-trees. IEEE Transactions on Visualization and Computer Graphics, 11(5):562–572, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wong:2006:GGV

[WFM^{+}06] Pak Chung Wong, Harlan Foote, Patrick Mackey, Ken Perrine, and George Chin, Jr. Generating graphs for visual analytics through interactive sketching. IEEE Transactions on Visualization and Computer Graphics, 12(6):1386– 1398, November/December 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wong:2012:SFV

[WFM^{+}12] Pak Chung Wong, Harlan Foote, Patrick Mackey, George Chin, Zhenyu Huang, and Jim Thomas. A spacefilling visualization technique for multivariate small-world graphs. IEEE Transactions on Visualization and Computer Graphics, 18(5): 797–809, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Waser:2010:WL

[WFR^{+}10] Jürgen Waser, Raphael Fuchs, Hrvoje Ribicic, Benjamin Schindler, Gunther Bloschl, and Eduard Gröller. World lines. IEEE Transactions on Visualization and Computer Graphics, 16(6):1458– 1467, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2016:IBB

[WFS^{+}16] Jinglu Wang, Tian Fang, Qingkun Su, Siyu Zhu, Jingbo Liu, Shengnan Cai, Chiew-Lan Tai, and Long Quan. Image-based building regularization using structural linear features. IEEE Transactions on Visualization and Computer Graphics, 22(6):1760– 1772, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Windhager:2019:VCH

[WFS^{+}19] Florian Windhager, Paolo Federico, Günther Schreder, Katrin Glinka, Marian Dörk, Silvia Miksch, and Eva Mayr. Visualization of cultural heritage collection data: State of the art and future challenges. IEEE Transactions on Visualization and Computer Graphics, 25(6):2311– 2330, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8352050/.

Wang:2017:GCR

[WFW^{+}17] Zhe Wang, Nivan Ferreira, Youhao Wei, Aarthy Sankari Bhaskar, and Carlos Scheidegger. Gaussian cubes: Realtime modeling for visual exploration of large multidimensional datasets. IEEE Transactions on Visualization and Computer Graphics, 23(1): 681–690, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wongsuphasawat:2012:EFF

[WG12] Krist Wongsuphasawat and David Gotz. Exploring flow, factors, and outcomes of temporal event sequences with the outflow visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2659–2668, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Willi:2016:STP

[WG16] Simon Willi and Anselm Grundhöfer. Spatio-temporal point path analysis and optimization of a galvanoscopic scanning laser projector. IEEE Transactions on Visualization and Computer Graphics, 22(11):2377–2384, ????

2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Williams:2008:VCP

[WGC^{+}08] David Williams, Sören Grimm, Ernesto Coto, Abdul Roudsari, and Haralambos Hatzakis. Volumetric curved planar reformation for virtual endoscopy. IEEE Transactions on Visualization and Computer Graphics, 14(1):109– 119, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weinstein:2008:IBC

[WGF08] Rachel Weinstein, Eran Guendelman, and Ronald Fedkiw. Impulse-based control of joints and muscles. IEEE Transactions on Visualization and Computer Graphics, 14(1):37–46, January/February 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/01/ttg2008010037s.
zip.

Wang:2008:CDI

[WGM^{+}08] Lujin Wang, Joachim Giesen, Kevin T. McDonnell, Peter Zolliker, and Klaus Mueller. Color design for illustrative visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1739–1754, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Waltemate:2018:IAP

[WGR^{+}18] Thomas Waltemate, Dominik Gall, Daniel Roth, Mario Botsch, and Marc Erich Latoschik. The impact of avatar personalization and immersion on virtual body ownership, presence, and emotional response. IEEE Transactions on Visualization and Computer Graphics, 24(4):1643–1652, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/04/08263407-
                                        abs.html.

Wang:2007:ILD

[WGS07a] Chaoli Wang, Antonio Garcia, and Han-Wei Shen. Interactive level-of-detail selection using image-based quality metric for large volume visualization. IEEE Transactions on Visualization and Computer Graphics, 13(1): 122–134, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wiebel:2007:CLF

[WGS07b] Alexander Wiebel, Christoph Garth, and Gerik Scheuermann. Computation of localized flow for steady and unsteady vector fields and its applications. IEEE Transactions on Visualization and Computer Graphics, 13(4):641– 651, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Willett:2013:IRE

[WGS^{+}13] Wesley Willett, Shiry Ginosar, Avital Steinitz, Bjorn Hartmann, and Maneesh Agrawala. Identifying redundancy and exposing provenance in crowdsourced data analysis. IEEE Transactions on Visualization and Computer Graphics, 19(12):2198– 2206, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:DVA

[WGSY19] Junpeng Wang, Liang Gou, Han-Wei Shen, and Hao Yang. DQNViz: A visual analytics approach to understand deep Q-networks. IEEE Transactions on Visualization and Computer Graphics, 25(1): 288–298, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08454905-
abs.html.

Wang:2018:GVA

[WGYS18] Junpeng Wang, Liang Gou, Hao Yang, and Han-Wei Shen.

GANViz: A visual analytics approach to understand the adversarial game. IEEE Transactions on Visualization and Computer Graphics, 24(6):1905–1917, June 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/06/08320546-
abs.html.

Wang:2019:DDV

[WGZ^{+}19] Junpeng Wang, Liang Gou, Wei Zhang, Hao Yang, and Han-Wei Shen. DeepVID: Deep visual interpretation and diagnosis for image classifiers via knowledge distillation. IEEE Transactions on Visualization and Computer Graphics, 25(6):2168– 2180, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8667661/.

Wang:2009:NWM

[WH09] Yu-Ping Wang and Shi-Min Hu. A new watermarking method for 3D models based on integral invariants. IEEE Transactions on Visualization and Computer Graphics, 15(2):285– 294, March/April 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wickham:2011:PP

[WH11] Hadley Wickham and Heike Hofmann. Product plots. IEEE Transactions on Visualization and Computer Graphics, 17(12):2223–2230, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:PBR

[WH18] Beibei Wang and Nicolas Holzschuch. Point-based rendering for homogeneous participating media with refractive boundaries. IEEE Transactions on Visualization and Computer Graphics, 24(10): 2743–2757, October 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08093692-
abs.html.

Willett:2007:SWI

[WHA07] Wesley Willett, Jeffrey Heer, and Maneesh Agrawala. Scented widgets: Improving navigation cues with embedded visualizations. IEEE Transactions on Visualization and Computer Graphics, 13(6):1129– 1136, November/December 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2014:SMC

[WHFL14] Jiang Wang, Zhe Huang, Hongbo Fu, and Rynson W. H. Lau. Structured mechanical collage. IEEE Transactions on Visualization and Computer Graphics, 20(7): 1076–1082, July 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2015:EDM

[WHK15] He Wang, Edmond S. L. Ho, and Taku Komura. An energy-driven motion planning method for two distant postures. IEEE Transactions on Visualization and Computer Graphics, 21(1): 18–30, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06824787-
abs.html.

Wyman:2016:FTI

[WHL16] Chris Wyman, Rama Hoetzlein, and Aaron Lefohn. Frustum-traced irregular Zbuffers: Fast, sub-pixel accurate hard shadows. IEEE Transactions on Visualization and Computer Graphics, 22 (10):2249–2261, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:VVA

[WHLS19] J. Wang, S. Hazarika, C. Li, and H. Shen. Visualization and visual analysis of ensemble data: a survey.

IEEE Transactions on Visualization and Computer Graphics, 25(9):2853–2872, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2014:GFE

[WHM14] Tim Wu, Alice Hung, and Kumar Mithraratne. Generating facial expressions using an anatomically accurate biomechanical model. IEEE Transactions on Visualization and Computer Graphics, 20(11): 1519–1529, November 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06872553-
abs.html.

Walny:2018:ARV

[WHP^{+}18] J. Walny, S. Huron, C. Perin, T. Wun, R. Pusch, and S. Carpendale. Active reading of visualizations. IEEE Transactions on Visualization and Computer Graphics, 24 (1):770–780, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wartell:2002:GCA

[WHR02] Z. Wartell, L. F. Hodges, and W. Ribarsky. A geometric comparison of algorithms for fusion control in stereoscopic HTDs. IEEE Transactions on Visualization and Computer Graphics, 8(2):

129–143, April 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg2002/pdf/v0129.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2002/v0129abs.
                                        htm.

Wei:2019:MDG

[WHX^{+}19] M. Wei, J. Huang, X. Xie, L. Liu, J. Wang, and J. Qin. Mesh denoising guided by patch normal co-filtering via kernel low-rank recovery. IEEE Transactions on Visualization and Computer Graphics, 25(10):2910–2926, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:LGS

[WHZ^{+}18] Y. Wang, F. Han, L. Zhu, O. Deussen, and B. Chen. Line graph or scatter plot? Automatic selection of methods for visualizing trends in time series. IEEE Transactions on Visualization and Computer Graphics, 24(2): 1141–1154, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wood:2012:SRI

[WII^{+}12] Jo Wood, Petra Isenberg, Tobias Isenberg, Jason Dykes, Nadia Boukhelifa, and Aidan Slingsby. Sketchy rendering for information visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2749– 2758, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wilkinson:2012:EAA

[Wil12]  Leland Wilkinson.  Ex-

act and approximate areaproportional circular Venn and Euler diagrams. IEEE Transactions on Visualization and Computer Graphics, 18 (2):321–331, February 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wilkinson:2018:VBD

[Wil18] L. Wilkinson. Visualizing big data outliers through distributed aggregation. IEEE Transactions on Visualization and Computer Graphics, 24 (1):256–266, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:IHR

[WJ08] Qin Wang and Joseph JaJa.

Interactive high-resolution isosurface ray casting on multicore processors. IEEE Transactions on Visualization and Computer Graphics, 14 (3):603–614, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wald:2017:OCR

[WJA^{+}17] I. Wald, G. P. Johnson, J. Amstutz, C. Brownlee, A. Knoll, J. Jeffers, J. Günther, and P. Navratil. OSPRay — a CPU ray tracing framework for scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 23(1): 931–940, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Willett:2017:EDR

[WJD17] Wesley Willett, Yvonne Jansen, and Pierre Dragicevic. Embedded data representations. IEEE Transactions on Visualization and Computer Graphics, 23(1): 461–470, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Westermann:2001:TPS

[WJE01] Rüdiger Westermann, Christopher Johnson, and Thomas Ertl. Topology-preserving smoothing of vector fields. IEEE Transactions on Visualization and Computer Graphics, 7(3):222–229, July 2001. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0222.
pdf;   http://www.computer.
org/tvcg/tg2001/v0222abs.
htm.

Won:2013:USI

[WJR^{+}13] Joong-Ho Won, Yongkweon Jeon, J. K. Rosenberg, Sungroh Yoon, G. D. Rubin, and S. Napel. Uncluttered singleimage visualization of vascular structures using GPU and integer programming. IEEE Transactions on Visualization and Computer Graphics, 19 (1):81–93, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wattenberg:2006:DSD

[WK06] Martin Wattenberg and Jesse Kriss. Designing for social data analysis. IEEE Transactions on Visualization and Computer Graphics, 12(4): 549–557, July/August 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2013:LSV

[WK13] Lei Wang and A. E. Kaufman. Lighting system for visual perception enhancement in volume rendering. IEEE Transactions on Visualization and Computer Graphics, 19 (1):67–80, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wiemker:2013:RST

[WKB^{+}13] Rafael Wiemker, Tobias Klinder, Martin Bergtholdt, Kirsten Meetz, Ingwer C.

Carlsen, and Thomas Bulow. A radial structure tensor and its use for shapeencoding medical visualization of tubular and nodular structures. IEEE Transactions on Visualization and Computer Graphics, 19(3): 353–366, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2007:CVC

[WKCB07] Yi Wang, David M. Krum, Enylton M. Coelho, and Doug A. Bowman. Contextualized videos: Combining videos with environment models to support situational understanding. IEEE Transactions on Visualization and Computer Graphics, 13(6):1568–1575, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1568s.
mpg.

Wood:2019:DEL

[WKD19] Jo Wood, Alexander Kachkaev, and Jason Dykes. Design exposition with literate visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 759–768, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440080-
abs.html.

Wu:2017:DMV

[WKI^{+}17] Kui Wu, Aaron Knoll, Benjamin J. Isaac, Hamish Carr, and Valerio Pascucci. Direct multifield volume ray casting of fiber surfaces. IEEE Transactions on Visualization and Computer Graphics, 23 (1):941–949, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiler:2003:HBV

[WKME03] Manfred Weiler, Martin Kraus, Markus Merz, and Thomas Ertl. Hardwarebased view-independent cell projection. IEEE Transactions on Visualization and Computer Graphics, 9(2):163– 175, April/June 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0163abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0163.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0163.pdf.

Wu:2005:PES

[WKSS05] Bing Wu, Roberta L. Klatzky, Damion Shelton, and George D. Stetten. Psychophysical evaluation of in-situ ultrasound visualization. IEEE Transactions on Visualization and Computer Graphics, 11(6):684–693, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Willemsen:2006:RNM

[WKW06] Peter Willemsen, Joseph K. Kearney, and Hongling Wang. Ribbon networks for modeling navigable paths of autonomous agents in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 12(3): 331–342, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wenger:2004:IVR

[WKZL04] Andreas Wenger, Daniel F. Keefe, Song Zhang, and David H. Laidlaw. Interactive volume rendering of thin thread structures within multivalued scientific data sets. IEEE Transactions on Visualization and Computer Graphics, 10(6):664–672, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0664abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0664.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0664.pdf.

Wang:2008:CSE

[WL08] Yu-Shuen Wang and Tong-Yee Lee. Curve-skeleton extraction using iterative least squares optimization. IEEE Transactions on Visualization and Computer Graphics, 14(4):926–936, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiss:2016:AMT

[WL16] K. Weiss and P. Lindstrom.

Adaptive multilinear tensor product wavelets. IEEE Transactions on Visualization and Computer Graphics, 22 (1):985–994, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Waldner:2014:AFG

[WLB^{+}14] Manuela Waldner, Mathieu Le Muzic, Matthias Bernhard, Werner Purgathofer, and Ivan Viola. Attractive flicker — guiding attention in dynamic narrative visualizations. IEEE Transactions on Visualization and Computer Graphics, 20(12):2456– 2465, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876019-
abs.html.

Weiss:2019:FSP

[WLD^{+}19] T. Weiss, A. Litteneker, N. Duncan, M. Nakada, C. Jiang, L. Yu, and D. Terzopoulos. Fast and scalable position-based layout synthesis. IEEE Transactions on Visualization and Computer Graphics, 25(12):3231– 3243, December 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2011:FMP

[WLDW11] Chenglei Wu, Yebin Liu, Qionghai Dai, and Bennett Wilburn. Fusing multiview and photometric stereo for 3D reconstruction under uncalibrated illumination. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1082–1095, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:NAN

[WLF^{+}19] Qianwen Wang, Zhen Li, Siwei Fu, Weiwei Cui, and Huamin Qu. Narvis: Authoring narrative slideshows for introducing data visualization designs. IEEE Transactions on Visualization and Computer Graphics, 25(1): 779–788, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

csdl/trans/tg/2019/01/08444072-
abs.html.

Wei:2012:FVC

[WLH^{+}12] Jin Wei, Chen-Feng Li, Shi-Min Hu, Ralph R. Martin, and Chiew-Lan Tai. Fisheye video correction. IEEE Transactions on Visualization and Computer Graphics, 18(10):1771–1783, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2017:LFA

[WLHD17] Yuwang Wang, Yebin Liu, Wolfgang Heidrich, and Qionghai Dai. The light field attachment: Turning a DSLR into a light field camera using a low budget camera ring. IEEE Transactions on Visualization and Computer Graphics, 24(10):2357– 2364, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07744676-
abs.html.

Wang:2013:STO

[WLHL13] Yu-Shuen Wang, Feng Liu, Pu-Sheng Hsu, and Tong-Yee Lee. Spatially and temporally optimized video stabilization. IEEE Transactions on Visualization and Computer Graphics, 19(8):1354–1361, August 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Walny:2012:UPT

[WLJ^{+}12] Jagoda Walny, Bongshin Lee, Paul Johns, Nathalie Henry Riche, and Sheelagh Carpendale. Understanding pen and touch interaction for data exploration on interactive whiteboards. IEEE Transactions on Visualization and Computer Graphics, 18(12):2779– 2788, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2005:LFM

[WLL^{+}05] Lifeng Wang, Stephen Lin, Seungyong Lee, Baining Guo, and Heung-Yeung Shum. Light field morphing using 2D features. IEEE Transactions on Visualization and Computer Graphics, 11(1):25– 34, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2005/01/v0025abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2005/01/
                                        v0025.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2005/01/v0025.pdf.

Wang:2012:RTP

[WLL^{+}12] Kexiang Wang, Xin Li, Bo Li, Huanhuan Xu, and Hong Qin. Restricted trivariate polycube splines for volumetric data modeling. IEEE Transactions on Visualization and Computer Graphics, 18 (5):703–716, May 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2016:TFP

[WLL^{+}16] Xiting Wang, Shixia Liu, Junlin Liu, Jianfei Chen, Jun Zhu, and Baining Guo. TopicPanorama: A full picture of relevant topics. IEEE Transactions on Visualization and Computer Graphics, 22(12): 2508–2521, December 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07374750-
abs.html.

Wu:2015:DMS

[WLLC15] Yu-Ting Wu, Tzu-Mao Li, Yu-Hsun Lin, and Yung-Yu Chuang. Dual-matrix sampling for scalable translucent material rendering. IEEE Transactions on Visualization and Computer Graphics, 21 (3):363–374, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/03/06994841-
abs.html.

Wu:2013:VVR

[WLLM13] Yingcai Wu, Xiaotong Liu, Shixia Liu, and Kwan-Liu Ma. ViSizer: A visualization resizing framework. IEEE Transactions on Visualization and Computer Graphics, 19 (2):278–290, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2004:LBM

[WLMK04] Xiaoming Wei, Wei Li, Klaus Mueller, and Arie E. Kaufman. The lattice-Boltzmann method for simulating gaseous phenomena. IEEE Transactions on Visualization and Computer Graphics, 10(2): 164–176, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2004/02/v0164abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/02/
                                        v0164.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2004/02/v0164.pdf.

Wang:2019:FRS

[WLMP19] Lili Wang, Xinglun Liang, Chunlei Meng, and Voicu Popescu. Fast ray-scene intersection for interactive shadow rendering with thousands of dynamic lights. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2242–2254, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8341814/.

Wu:2018:IIV

[WLS^{+}18] Y. Wu, J. Lan, X. Shu, C. Ji, K. Zhao, J. Wang, and H. Zhang. iTTVis: Interactive visualization of table tennis data. IEEE Transactions on Visualization and Computer Graphics, 24(1):709–718, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:VAF

[WLS^{+}19] Hong Wang, Yafeng Lu, Shade T. Shutters, Michael Steptoe, Feng Wang, Steven Landis, and Ross Maciejewski. A visual analytics framework for spatiotemporal trade network analysis. IEEE Transactions on Visualization and Computer Graphics, 25 (1):331–341, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440040-
abs.html.

Wang:2017:MRC

[WLSL17] Junpeng Wang, Xiaotong Liu, Han-Wei Shen, and Guang Lin. Multi-resolution climate ensemble parameter analysis with nested parallel coordinates plots. IEEE Transactions on Visualization and Computer Graphics, 23(1):81– 90, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:LBR

[WLSW08] Wencheng Wang, Jing Li, Hanqiu Sun, and Enhua Wu. Layer-based representation of polyhedrons for point containment tests. IEEE Transactions on Visualization and Computer Graphics, 14(1):73– 83, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:FVD

[WLT08] Yu-Shuen Wang, Tong-Yee Lee, and Chiew-Lan Tai. Focus+context visualization with distortion minimization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1731–1738, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:SFD

[WLT18a] Xiaojun Wang, Shiguang Liu, and Yiying Tong. Stain formation on deforming inelastic cloth. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3214–3224, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8245822/.

Wang:2018:OMM

[WLT^{+}18b] Yangang Wang, Yebin Liu, Xin Tong, Qionghai Dai, and Ping Tan. Outdoor markerless motion capture with sparse handheld video cameras. IEEE Transactions on Visualization and Computer Graphics, 24(5): 1856–1866, May 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/05/07896626-
abs.html.

Wang:2017:GEI

[WLW17] Lili Wang, Ming Lin, and Enhua Wu. Guest Editors introduction: Special section on the ACM Symposium on Virtual Reality Software and Technology 2015. IEEE Transactions on Visualization and Computer Graphics, 23 (3):1207–1208, March 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/03/07835794.
pdf.

Wang:2018:SFH

[WLW^{+}18] Weiming Wang, Yong-Jin Liu, Jun Wu, Shengjing Tian, Charlie C. L. Wang, Ligang Liu, and Xiuping Liu. Support-free hollowing. IEEE Transactions on Visualization and Computer Graphics, 24(10):2787– 2798, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2018/10/08082529-
                                        abs.html.

Wan:2010:EMI

[WLWL10] Liang Wan, Xiaopei Liu, Tien-Tsin Wong, and Chi-Sing Leung. Evolving mazes from images. IEEE Transactions on Visualization and Computer Graphics, 16(2): 287–297, March/April 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2014:OVA

[WLY^{+}14] Yingcai Wu, Shixia Liu, Kai Yan, Mengchen Liu, and Fangzhao Wu. Opinion-Flow: Visual analysis of opinion diffusion on social media. IEEE Transactions on Visualization and Computer Graphics, 20(12):1763– 1772, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876032-
                                        abs.html.

Wu:2010:TAE

[WLZM10] Keqin Wu, Zhanping Liu, Song Zhang, and Robert J. Moorhead II. Topologyaware evenly spaced streamline placement. IEEE Transactions on Visualization and Computer Graphics, 16(5): 791–801, September/October 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ward:2005:GEI

[WM05] Matthew O. Ward and Tamara Munzner. Guest Editors’ introduction: Special section on InfoVis. IEEE Transactions on Visualization and Computer Graphics, 11 (4):431, July/August 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/04/v0431.
pdf.

Wang:2008:SAV

[WM08] Chaoli Wang and Kwan-Liu Ma. A statistical approach to volume data quality assessment. IEEE Transactions on Visualization and Computer Graphics, 14(3):590– 602, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2013:SDW

[WM13a] Bing Wang and Klaus Mueller.

SketchPadN-D: WYDIWYG sculpting and editing in highdimensional space. IEEE Transactions on Visualization and Computer Graphics, 19(12):2060–2069, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2013:EBE

[WM13b] C. C. L. Wang and D. Manocha.

Efficient boundary extraction of BSP solids based on clipping operations. IEEE Transactions on Visualization and Computer Graphics, 19(1): 16–29, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2016:VCA

[WM16] J. Wang and K. Mueller. The visual causality analyst: An interactive interface for causal reasoning. IEEE Transactions on Visualization and Computer Graphics, 22(1): 230–239, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:SVE

[WM18] B. Wang and K. Mueller. The Subspace Voyager: Exploring high-dimensional data along a continuum of salient 3D subspaces. IEEE Transactions on Visualization and Computer Graphics, 24(2):1204– 1222, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wyman:2019:IAT

[WM19] Chris Wyman and Morgan McGuire. Improved alpha testing using hashed sampling. IEEE Transactions on Visualization and Computer Graphics, 25(2):1309– 1320, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08010336-
abs.html.

Wongsuphasawat:2016:VEA

[WMA^{+}16] K. Wongsuphasawat, D. Moritz, A. Anand, J. Mackinlay, B. Howe, and J. Heer. Voyager: Exploratory analysis via faceted browsing of visualization recommendations. IEEE Transactions on Visualization and Computer Graphics, 22 (1):649–658, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Woo:2012:FDD

[WMGE12] Insoo Woo, Ross Maciejewski, Kelly P. Gaither, and David S. Ebert. Featuredriven data exploration for volumetric rendering. IEEE Transactions on Visualization and Computer Graphics, 18(10):1731–1743, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Whitaker:2013:CBM

[WMK13] Ross T. Whitaker, Mahsa Mirzargar, and Robert M. Kirby. Contour boxplots: A method for characterizing uncertainty in feature sets from simulation ensembles. IEEE Transactions on Visualization and Computer Graphics, 19(12):2713– 2722, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Williams:1998:HAV

[WMS98] P. L. Williams, N. L. Max, and C. M. Stein. A high accuracy volume renderer for unstructured data. IEEE Transactions on Visualization and Computer Graphics, 4(1):37– 54, January/March 1998. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1998/pdf/v0037.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1998/v0037abs.
                                        htm.

Wan:2011:SSD

[WMWL11] Liang Wan, Shue-Kwan Mak, Tien-Tsin Wong, and Chi-Sing Leung. Spatiotemporal sampling of dynamic environment sequences. IEEE Transactions on Visualization and Computer Graphics, 17(10):1499–1509, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wan:2009:IVT

[WOCH09] Yong Wan, Hideo Otsuna, Chi-Bin Chien, and Charles Hansen. An interactive visualization tool for multichannel confocal microscopy data in neurobiology research. IEEE Transactions on Visualization and Computer Graphics, 15(6):1489–1496, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wonka:2016:ACD

[Won16] Peter Wonka. Automatic constraint detection for 2D layout regularization. IEEE Transactions on Visualization and Computer Graphics, 22 (8):1933–1944, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2017:TPN

[WOO17] He Wang, Jan Ondrej, and Carol O’Sullivan. Trending paths: A new semanticlevel metric for comparing simulated and real crowd data. IEEE Transactions on Visualization and Computer Graphics, 23(5):1454– 1464, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/05/07797248-
abs.html.

Wang:2016:IMM

[WP16a] Yu-Shuen Wang and Wan-Yu Peng. Interactive metro map editing. IEEE Transactions on Visualization and Computer Graphics, 22(2):1115– 1126, February 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2016:MFV

[WP16b] M. L. Wu and V. Popescu. Multiperspective focus+context visualization. IEEE Transactions on Visualization and Computer Graphics, 22(5): 1555–1567, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2018:EVA

[WP18] Meng-Lin Wu and Voicu Popescu. Efficient VR and AR navigation through multiperspective occlusion management. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3069–3080, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8123949/.

Williams:2019:ERG

[WP19] N. L. Williams and T. C. Peck. Estimation of rotation gain thresholds considering FOV, gender, and distractors. IEEE Transactions on Visualization and Computer Graphics, 25(11):3158– 3168, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Williams:2011:AEQ

[WPB^{+}11] Sean Williams, Mark Petersen, Peer-Timo Bremer, Matthew Hecht, Valerio Pascucci, James Ahrens, Mario Hlawitschka, and Bernd Hamann. Adaptive extraction and quantification of geophysical vortices. IEEE Transactions on Visualization and Computer Graphics, 17(12): 2088–2095, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2013:ADP

[WPC^{+}13] Rui Wang, Minghao Pan, Weifeng Chen, Zhong Ren, Kun Zhou, Wei Hua, and Hujun Bao. Analytic double product integrals for allfrequency relighting. IEEE Transactions on Visualization and Computer Graphics, 19(7):1133–1142, July 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weyrich:2005:RDS

[WPG05] Tim Weyrich, Hanspeter Pfister, and Markus Gross. Rendering deformable surface reflectance fields. IEEE Transactions on Visualization and Computer Graphics, 11(1):48– 58, January/February 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2005/01/v0048abs.

                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2005/01/
                                        v0048.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2005/01/v0048.pdf.

Wittenbrink:1996:GVU

[WPL96] Craig M. Wittenbrink, Alex T. Pang, and Suresh K. Lodha. Glyphs for visualizing uncertainty in vector fields. IEEE Transactions on Visualization and Computer Graphics, 2 (3):266–279, September 1996. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0266.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0266abs.
                                        htm.

Wang:2009:TSS

[WPS^{+}09] Taowei David Wang, Catherine Plaisant, Ben Shneiderman, Neil Spring, David Roseman, Greg Marchand, Vikramjit Mukherjee, and Mark Smith. Temporal summaries: Supporting temporal categorical searching, aggregation and comparison. IEEE Transactions on Visualization and Computer Graphics, 15(6):1049–1056, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Woodring:2016:SEA

[WPS^{+}16] J. Woodring, M. Petersen, A. Schmeiber, J. Patchett, J. Ahrens, and H. Hagen. In situ eddy analysis in a high-resolution ocean climate model. IEEE Transactions on Visualization and Computer Graphics, 22(1):857–866, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wyman:2006:IDI

[WPSH06] Chris Wyman, Steven Parker, Peter Shirley, and Charles Hansen. Interactive display of isosurfaces with global illumination. IEEE Transactions on Visualization and Computer Graphics, 12(2):186– 196, March/April 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2011:PCP

[WPZ^{+}11] Huai-Yu Wu, Chunhong Pan, Hongbin Zha, Qing Yang, and Songde Ma. Partwise crossparameterization via nonregular convex hull domains. IEEE Transactions on Visualization and Computer Graphics, 17(10):1531–1544, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2016:EVA

[WPZ^{+}16] Y. Wu, N. Pitipornvivat, J. Zhao, S. Yang, G. Huang, and H. Qu. egoSlider: Visual analysis of egocentric network evolution. IEEE Transactions on Visualization and Computer Graphics, 22(1): 260–269, January 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2007:ITF

[WQ07] Yingcai Wu and Huamin Qu. Interactive transfer function design based on editing direct volume rendered images. IEEE Transactions on Visualization and Computer Graphics, 13(5):1027–1040, September/October 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2007:SBB

[WQS07] Huawei Wang, Kaihuai Qin, and Hanqiu Sun. \sqrt{3}- subdivision-based biorthogonal wavelets. IEEE Transactions on Visualization and Computer Graphics, 13(5): 914–925, September/October 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2018:TSF

[WQZ^{+}18] Xiangzhi Wei, Siqi Qiu, Lin Zhu, Ruiliang Feng, Yaobin Tian, Juntong Xi, and Youyi Zheng. Toward supportfree 3D printing: A skeletal approach for partitioning models. IEEE Transactions on Visualization and Computer Graphics, 24(10):2799– 2812, October 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08086214-
abs.html.

Whited:2011:BMD

[WR11] Brian Whited and Jaroslaw (Jarek)

Rossignac. Ball-Morph: Definition, implementation, and comparative evaluation. IEEE Transactions on Visualization and Computer Graphics, 17 (6):757–769, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Waser:2011:NRC

[WRF^{+}11] Jürgen Waser, Hrvoje Ribičić, Raphael Fuchs, Christian Hirsch, Benjamin Schindler, Günther Blöschl, and Eduard Gröller. Nodes on ropes: a comprehensive data and control flow for steering ensemble simulations. IEEE Transactions on Visualization and Computer Graphics, 17(12): 1872–1881, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wolf:2019:FMM

[WRHR19] D. Wolf, M. Rietzler, L. Hnatek, and E. Rukzio. Face/On: Multi-modal haptic feedback for head-mounted displays in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3169–3177, November 2019.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wagner:2010:RTD

[WRM^{+}10] Daniel Wagner, Gerhard Reitmayr, Alessandro Mulloni, Tom Drummond, and Dieter Schmalstieg. Real-time detection and tracking for augmented reality on mobile phones. IEEE Transactions on Visualization and Computer Graphics, 16(3):355– 368, May/June 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wilde:2019:RSF

[WRT19] Thomas Wilde, Christian Rossi, and Holger Theisel. Recirculation surfaces for flow visualization. IEEE Transactions on Visualization and Computer Graphics, 25(1): 946–955, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440089-
                                        abs.html.

Wischgoll:2001:DVC

[WS01] T. Wischgoll and G. Scheuermann. Detection and visualization of closed streamlines in planar flows. IEEE Transactions on Visualization and Computer Graphics, 7(2):165– ??, April/June 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2001/pdf/v0165.
pdf;   http://www.computer.
org/tvcg/tg2001/v0165abs.
htm.

Wang:2006:LMV

[WS06a] Chaoli Wang and Han-Wei Shen. LOD map — a visual interface for navigating multiresolution volume visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5):1029– 1036, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Woodring:2006:MVT

[WS06b] Jonathan Woodring and Han-Wei Shen. Multi-variate, time varying, and comparative visualization with contextual cues. IEEE Transactions on Visualization and Computer Graphics, 12(5): 909–916, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Woodring:2009:MTA

[WS09] Jonathan Woodring and Han-Wei Shen. Multiscale time activity data exploration via temporal clustering visualization spreadsheet. IEEE Transactions on Visualization and Computer Graphics, 15(1):123–137, January/ February 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2016:AVA

[WSA^{+}16] Y. Wang, Q. Shen, D. Archambault, Z. Zhou, M. Zhu, S. Yang, and H. Qu. AmbiguityVis: Visualization of ambiguity in graph layouts. IEEE Transactions on Visualization and Computer Graphics, 22(1):359–368, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Whang:1995:ORA

[WSC^{+}95] Kyu-Young Whang, Ju-Won Song, Ji-Woong Chang, Ji-Yun Kim, Wan-Sup Cho, Chong-Mok Park, and Il-Yeol Song. Octree-R: an adaptive octree for efficient ray tracing. IEEE Transactions on Visualization and Computer Graphics, 1(4):343– 349, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1995/pdf/v0343.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1995/v0343abs.
                                        htm.

Walker:2013:EFP

[WSD^{+}13] Rick Walker, Aiden Slingsby, Jason Dykes, Kai Xu, Jo Wood, Phong H. Nguyen, Derek Stephens, B. L. William Wong, and Yongjun Zheng. An extensible framework for provenance in human terrain visual analytics. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2139–2148, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weiskopf:2007:TBV

[WSE07] Daniel Weiskopf, Tobias Schafhitzel, and Thomas Ertl. Texture-based visualization of unsteady 3D flow by real-time advection and volumetric illumination. IEEE Transactions on Visualization and Computer Graphics, 13(3):569– 582, May/June 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wagner:2019:KKA

[WSH^{+}19] M. Wagner, D. Slijepcevic, B. Horsak, A. Rind, M. Zeppelzauer, and W. Aigner. KAVAGait: Knowledgeassisted visual analytics for clinical gait analysis. IEEE Transactions on Visualization and Computer Graphics, 25 (3):1528–1542, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wilkie:2012:TGD

[WSLL12] David Wilkie, Jason Sewall, Ming C. Lin, and Ming C. Lin. Transforming GIS data into functional road models for large-scale traffic simulation. IEEE Transactions on Visualization and Computer Graphics, 18(6):890–901, June 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wong:2009:NVT

[WSM^{+}09] Pak Chung Wong, Kevin Schneider, Patrick Mackey, Harlan Foote, George Chin, Jr., Ross Guttromson, and Jim Thomas. A novel visualization technique for electric power grid analytics. IEEE Transactions on Visualization and Computer Graphics, 15 (3):410–423, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2011:BCF

[WSPVJ11] Bei Wang, Brian Summa, Valerio Pascucci, and Mikael Vejdemo-Johansson. Branching and circular features in high dimensional data. IEEE Transactions on Visualization and Computer Graphics, 17(12):1902–1911, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2009:BSR

[WSS09] Hao Wang, Carlos E. Scheidegger, and Cláudio T. Silva. Bandwidth selection and reconstruction quality in pointbased surfaces. IEEE Transactions on Visualization and Computer Graphics, 15(4): 572–582, July/August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weinkauf:2007:CSP

[WSTH07] Tino Weinkauf, Jan Sahner, Holger Theisel, and Hans-Christian Hege. Cores of swirling particle motion in unsteady flows. IEEE Transactions on Visualization and Computer Graphics, 13(6):1759–1766, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1759s.
zip.

Wang:2016:MFP

[WSW16] Z. Wang, H. Seidel, and T. Weinkauf. Multi-field pattern matching based on sparse feature sampling. IEEE Transactions on Visualization and Computer Graphics, 22 (1):807–816, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wongsuphasawat:2018:VDG

[WSW^{+}18] K. Wongsuphasawat, D. Smilkov, J. Wexler, J. Wilson, D. Mané, D. Fritz, D. Krishnan, F. B. Viégas, and M. Wattenberg. Visualizing dataflow graphs of deep learning models in TensorFlow. IEEE Transactions on Visualization and Computer Graphics, 24(1): 1–12, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2007:STL

[WSY07] Huamin Wang, Mingxuan Sun, and Ruigang Yang. Space–time light field rendering. IEEE Transactions on Visualization and Computer Graphics, 13(4):697– 710, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2017:STC

[WSYM17] Wenguan Wang, Jianbing Shen, Yizhou Yu, and Kwan-Liu Ma. Stereoscopic thumbnail creation via efficient stereo saliency detection. IEEE Transactions on Visualization and Computer Graphics, 23(8):2014– 2027, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07544591-
                                        abs.html.

Weinkauf:2010:SLT

[WT10a] Tino Weinkauf and Holger Theisel. Streak lines as tangent curves of a derived vector field. IEEE Transactions on Visualization and Computer Graphics, 16(6):1225– 1234, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2010:LSF

[WT10b] Chunlin Wu and Xuecheng Tai. A level set formulation of geodesic curvature flow on simplicial surfaces. IEEE Transactions on Visualization and Computer Graphics, 16(4):647–662, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ware:2019:MMF

[WTB^{+}19] C. Ware, T. L. Turton, R. Bujack, F. Samsel, P. Shrivastava, and D. H. Rogers. Measuring and modeling the feature detection threshold functions of colormaps. IEEE Transactions on Visualization and Computer Graphics, 25(9):2777–2790, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2009:GGA

[WTL^{+}09] Jyh-Da Wei, Ming-Hung Tsai, Gen-Cher Lee, Jeng-Hung Huang, and Der-Tsai Lee. GeoBuilder: a geometric algorithm visualization and debugging system for 2D and 3D geometric computing. IEEE Transactions on Visualization and Computer Graphics, 15(2):234–248, March/April 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2019:BRM

[WTP^{+}19] Mingqiang Wei, Yang Tian, Wai-Man Pang, Charlie C. L. Wang, Ming-Yong Pang, Jun Wang, Jing Qin, and Pheng-Ann Heng. Bas-relief modeling from normal layers. IEEE Transactions on Visualization and Computer Graphics, 25(4):1651–1665, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8322258/.

Wiebel:2007:GSL

[WTS^{+}07] Alexander Wiebel, Xavier Tricoche, Dominic Schneider, Heike Jaenicke, and Gerik Scheuermann. Generalized streak lines: analysis and visualization of boundary induced vortices. IEEE Transactions on Visualization and Computer Graphics, 13(6):1735–1742, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2007/06/v1735s.
                                        zip.

Weinkauf:2011:SFF

[WTVP11] Tino Weinkauf, Holger Theisel, Allen Van Gelder, and Alex Pang. Stable feature flow fields. IEEE Transactions on Visualization and Computer Graphics, 17(6):770– 780, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:SSR

[WTW^{+}08] Lu Wang, Changhe Tu, Wenping Wang, Xiangxu Meng, Bin Chan, and Dongming Yan. Silhouette smoothing for real-time rendering of mesh surfaces. IEEE Transactions on Visualization and Computer Graphics, 14(3):640– 652, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/03/ttg2008030640s.
avi.

Wu:2016:MIT

[Wu16] Enhua Wu. Multiphase interface tracking with fast semi-Lagrangian contouring. IEEE Transactions on Visualization and Computer Graphics, 22 (8):1973–1986, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wattenberg:2008:WTI

[WV08] Martin Wattenberg and Fernanda B. Viégas. The word tree, an interactive visual concordance. IEEE Transactions on Visualization and Computer Graphics, 14(6):1221– 1228, November/December 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wiebel:2012:WWY

[WVFH12] Alexander Wiebel, Frans M. Vos, David Foerster, and Hans-Christian Hege. WYSI-WYP: What you see is what you pick. IEEE Transactions on Visualization and Computer Graphics, 18(12):2236– 2244, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2007:IBC

[WW07] Chung-Ming Wang and Ren-Jie Wang. Image-based color ink diffusion rendering. IEEE Transactions on Visualization and Computer Graphics, 13(2):235–246, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wan:2013:ROS

[WWB^{+}13] Min Wan, Yu Wang, E. Bae, Xue-Cheng Tai, and Desheng Wang. Reconstructing open surfaces via graph-cuts. IEEE Transactions on Visualization and Computer Graphics, 19 (2):306–318, February 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2014:DRI

[WWC^{+}14] Lisheng Wang, Pai Wang, Liuhang Cheng, Yu Ma, Shenzhi Wu, Yu-Ping Wang, and Zongben Xu. Detection and reconstruction of an implicit boundary surface by adaptively expanding a small surface patch in a 3D image. IEEE Transactions on Visualization and Computer Graphics, 20(11):1490– 1506, November 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/11/06767138-
abs.html.

Wang:2019:IBA

[WWF^{+}19] Yunhai Wang, Zeyu Wang, Chi-Wing Fu, Hansjorq Schmauder, Oliver Deussen, and Daniel Weiskopf. Image-based aspect ratio selection. IEEE Transactions on Visualization and Computer Graphics, 25 (1):840–849, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440843-
abs.html.

Wong:2003:GVA

[WWFT03] Pak Chung Wong, Kwong Kwok Wong, Harlan Foote, and Jim Thomas. Global visualization and alignments of whole bacterial genomes. IEEE Transactions on Visualization and Computer Graphics, 9(3):361– 377, July/September 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2003/03/v0361abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2003/03/
                                        v0361.pdf.

Wan:2007:IEC

[WWL07] Liang Wan, Tien-Tsin Wong, and Chi-Sing Leung. Isocube: Exploiting the Cubemap hardware. IEEE Transactions on Visualization and Computer Graphics, 13(4):720– 731, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2010:OIV

[WWL^{+}10] Yingcai Wu, Furu Wei, Shixia Liu, Norman Au, Weiwei Cui, Hong Zhou, and Huamin Qu. OpinionSeer: Interactive visualization of hotel customer feedback. IEEE Transactions on Visualization and Computer Graphics, 16(6):1109– 1118, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2011:FPV

[WWLM11] Yu-Shuen Wang, Chaoli Wang, Tong-Yee Lee, and Kwan-Liu Ma. Featurepreserving volume data reduction and focus+context visualization. IEEE Transactions on Visualization and Computer Graphics, 17(2):171– 181, February 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2016:PTB

[WWS^{+}16] Tongshuang Wu, Yingcai Wu, Conglei Shi, Huamin Qu, and Weiwei Cui. PieceStack: Toward better understanding of stacked graphs. IEEE Transactions on Visualization and Computer Graphics, 22(6): 1640–1651, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2018:RSM

[WWS^{+}18] Y. Wang, Y. Wang, Y. Sun, L. Zhu, K. Lu, C. W. Fu, M. Sedlmair, O. Deussen, and B. Chen. Revisiting stress majorization as a unified framework for interactive constrained graph visualization. IEEE Transactions on Visualization and Computer Graphics, 24(1):489–499, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:CIR

[WWW^{+}19] Feng Wang, Ingo Wald, Qi Wu, Will Usher, and Chris R. Johnson. CPU isosurface ray tracing of adaptive mesh refinement data. IEEE Transactions on Visualization and Computer Graphics, 25 (1):1142–1151, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08493612-
                                        abs.html.

Wu:2014:OSA

[WWY14] Ruobing Wu, Wenping Wang, and Yizhou Yu. Optimized synthesis of art patterns and layered textures. IEEE Transactions on Visualization and Computer Graphics, 20(3): 436–446, March 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2019:VEA

[WWYP19] Lili Wang, Jian Wu, Xuefeng Yang, and Voicu Popescu. VR exploration assistance through automatic occlusion removal. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2083–2092, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642365/.

Wang:2004:EEB

[WWYS04] Bin Wang, Wenping Wang, Huaiping Yang, and Jiaguang Sun. Efficient example-based painting and synthesis of 2D directional texture. IEEE Transactions on Visualization and Computer Graphics, 10 (3):266–277, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/03/v0266abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/03/
v0266.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/03/v0266.pdf.

Wang:2018:TRT

[WWZ^{+}18] Yunhai Wang, Zeyu Wang, Lifeng Zhu, Jian Zhang, Chi-Wing Fu, Zhanglin Cheng, Changhe Tu, and Baoquan Chen. Is there a robust technique for selecting aspect ratios in line charts? IEEE Transactions on Visualization and Computer Graphics, 24 (12):3096–3110, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8239850/.

Wang:2019:SAF

[WWZ^{+}19] Yunhai Wang, Yanyan Wang, Haifeng Zhang, Yinqi Sun, Chi-Wing Fu, Michael Sedlmair, Baoquan Chen, and Oliver Deussen. Structureaware fisheye views for efficient large graph exploration. IEEE Transactions on Visualization and Computer Graphics, 25(1):566– 575, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2019/01/08440835-
                                        abs.html.

Wang:2013:MES

[WX13] Wencheng Wang and Guofu Xie. Memory-efficient singlepass GPU rendering of multifragment effects. IEEE Transactions on Visualization and Computer Graphics, 19(8): 1307–1316, August 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2017:SDR

[WX17] Jun Wang and Kai Xu. Shape detection from raw LiDAR data with subspace modeling. IEEE Transactions on Visualization and Computer Graphics, 23(9):2137– 2150, September 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/09/07557019-
                                        abs.html.

Wu:2008:HTA

[WXC^{+}08] Qing Wu, Tian Xia, Chun Chen, Hsueh-Yi Sean Lin, Hongcheng Wang, and Yizhou Yu. Hierarchical tensor approximation of multi-dimensional visual data. IEEE Transactions on Visualization and Computer Graphics, 14(1): 186–199, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2017:CVT

[WXJD17] Yutong Wang, Xiaowei Xue, Xiaogang Jin, and Zhigang Deng. Creative virtual tree modeling through hierarchical topology-preserving blending. IEEE Transactions on Visualization and Computer Graphics, 23(12):2521– 2534, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/12/07775115-
abs.html.

Wang:2014:SOF

[WXKP14] Lili Wang, Naiwen Xie, Wei Ke, and Voicu Popescu. Second-order feed-forward rendering for specular and glossy reflections. IEEE Transactions on Visualization and Computer Graphics, 20 (9):1, September 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2019:FVS

[WXW^{+}19] Yingcai Wu, Xiao Xie, Jiachen Wang, Dazhen Deng, Hongye Liang, Hui Zhang, Shoubin Cheng, and Wei Chen. ForVizor: Visualizing spatio-temporal team formations in soccer. IEEE Transactions on Visualization and Computer Graphics, 25(1): 65–75, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440804-
                                        abs.html.

Wen:2017:RTE

[WXY17] Quan Wen, Feng Xu, and Jun-Hai Yong. Real-time 3D eye performance reconstruction for RGBD cameras. IEEE Transactions on Visualization and Computer Graphics, 23(12):2586– 2598, December 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/12/07790904-
                                        abs.html.

Wu:2016:TVE

[WXZ^{+}16] W. Wu, J. Xu, H. Zeng, Y. Zheng, H. Qu, B. Ni, M. Yuan, and L. M. Ni. Tel-CoVis: Visual exploration of co-occurrence in urban human mobility based on Telco data. IEEE Transactions on Visualization and Computer Graphics, 22(1):935–944, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2019:RTC

[WY19a] Kui Wu and Cem Yuksel. Real-time cloth rendering with fiber-level detail. IEEE Transactions on Visualization and Computer Graphics, 25(2):1297– 1308, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/07993021-
abs.html.

Wyman:2019:GEI

[WY19b] Chris Wyman and Cem Yuksel. Guest Editor’s introduction to the special section on I3D. IEEE Transactions on Visualization and Computer Graphics, 25(2): 1269, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08594691.
pdf.

Wang:2014:VES

[WYL^{+}14] Zuchao Wang, Tangzhi Ye, Min Lu, Xiaoru Yuan, Huamin Qu, Jacky Yuan, and Qianliang Wu. Visual exploration of sparse traffic trajectory data. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1813–1822, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876014-
abs.html.

Wang:2019:ISR

[WYL^{+}19] Y. Wang, D. Yan, X. Liu, C. Tang, J. Guo, X. Zhang, and P. Wonka. Isotropic surface remeshing without large and small angles. IEEE Transactions on Visualization and Computer Graphics, 25(7):2430–2442, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2008:IDT

[WYM08] Chaoli Wang, Hongfeng Yu, and Kwan-Liu Ma. Importance-driven time-varying data visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1547–1554, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2012:VFU

[WYM12] Yingcai Wu, Guo-Xun Yuan, and Kwan-Liu Ma. Visualizing flow of uncertainty through analytical processes. IEEE Transactions on Visualization and Computer Graphics, 18(12):2526–2535, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2015:BNF

[WYP^{+}15] Mingqiang Wei, Jinze Yu, Wai-Man Pang, Jun Wang, Jing Qin, Ligang Liu, and Pheng-Ann Heng. Bi-normal filtering for mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 21(1):43– 55, January 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/01/06822598-
abs.html.

Wen:2008:EUD

[WZ08] Zhen Wen and Michelle Zhou.

Evaluating the use of data transformation for information visualization. IEEE Transactions on Visualization and Computer Graphics, 14(6):1309–1316, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2015:PSM

[WZC^{+}15] Hongjian Wang, Naiyu Zhang, Jean-Charles Creput, Julien Moreau, and Yassine Ruichek. Parallel structured mesh generation with disparity maps by GPU implementation. IEEE Transactions on Visualization and Computer Graphics, 21(9):1045–1057, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07061525-abs.html.

Wei:2004:LBF

[WZF^{+}04] Xiaoming Wei, Ye Zhao, Zhe Fan, Wei Li, Feng Qiu, Suzanne Yoakum-Stover, and Arie E. Kaufman. Latticebased flow field modeling.

IEEE Transactions on Visualization and Computer Graphics, 10(6):719–729, November/December 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/06/v0719abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/06/
v0719.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/06/v0719.pdf.

Wang:2012:MDA

[WZK12] Lei Wang, Xin Zhao, and Arie E. Kaufman. Modified dendrogram of attribute space for multidimensional transfer function design. IEEE Transactions on Visualization and Computer Graphics, 18(1): 121–131, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wan:2004:ISR

[WZQK04] Ming Wan, Nan Zhang, Huamin Qu, and Arie E. Kaufman. Interactive stereoscopic rendering of volumetric environments. IEEE Transactions on Visualization and Computer Graphics, 10(1):15– 28, January/February 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0015abs.
htm;   http://csdl.computer.

org/dl/trans/tg/2004/01/
v0015.pdf.

Wang:2013:VTJ

[WZvdW13] Zuchao Wang, Junping Zhang, and Huub van de Wetering. Visual traffic jam analysis based on trajectory data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2159– 2168, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2005:CTM

[WZW^{+}05] Daniel Wang, Yuru Zhang, Yuhui Wang, Yuan-Shin Lee, Peijun Lu, and Yong Wang. Cutting on triangle mesh: Local model-based haptic display for dental preparation surgery simulation. IEEE Transactions on Visualization and Computer Graphics, 11(6):671–683, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2009:ACA

[XA09] Yi Xu and Daniel G. Aliaga.

An adaptive correspondence algorithm for modeling scenes with strong interreflections. IEEE Transactions on Visualization and Computer Graphics, 15(3):465–480, May/June 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2010:MRM

[XA10] Yi Xu and Daniel G. Aliaga. Modeling repetitive motions using structured light. IEEE Transactions on Visualization and Computer Graphics, 16(4):676–689, July/August 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2019:HMS

[XC19] Kaoji Xu and Guoning Chen. Hexahedral mesh structure visualization and evaluation. IEEE Transactions on Visualization and Computer Graphics, 25(1):1173– 1182, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440081-
                                        abs.html.

Xie:2014:VVA

[XCH^{+}14] Cong Xie, Wei Chen, Xinxin Huang, Yueqi Hu, Scott Barlowe, and Jing Yang. VAET: A visual analytics approach for e-transactions time-series. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1743–1752, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876015-
                                        abs.html.

Xie:2019:SBM

[XCZ^{+}19a] X. Xie, X. Cai, J. Zhou, N. Cao, and Y. Wu. A semantic-based method for visualizing large image collections. IEEE Transactions on Visualization and Computer Graphics, 25(7): 2362–2377, July 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2019:MCR

[XCZ^{+}19b] Weipeng Xu, Avishek Chatterjee, Michael Zollhöfer, Helge Rhodin, Pascal Fua, Hans-Peter Seidel, and Christian Theobalt. Mo^{2}Cap^{2}: Real-time mobile 3D motion capture with a cap-mounted fisheye camera. IEEE Transactions on Visualization and Computer Graphics, 25(5): 2093–2101, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643070/.

Xu:2011:CCP

[XDN11] Di Xu, Colin Doutre, and Panos Nasiopoulos. Correction of clipped pixels in color images. IEEE Transactions on Visualization and Computer Graphics, 17(3): 333–344, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xia:1997:ART

[XESV97] J. C. Xia, J. El-Sana, and A. Varshney. Adaptive realtime level-of-detail-based rendering for polygonal models. IEEE Transactions on Visualization and Computer Graphics, 3(2):171– 183, April/June 1997. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1997/pdf/v0171.
pdf;   http://www.computer.
org/tvcg/tg1997/v0171abs.
htm.

Xie:2004:IRU

[XF04] Zhiyong Xie and Gerald E. Farin. Image registration using hierarchical B-splines. IEEE Transactions on Visualization and Computer Graphics, 10(1):85–94, January/February 2004. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/01/v0085abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/01/
v0085.pdf.

Xu:2019:MGS

[XFZ^{+}19] P. Xu, H. Fu, Y. Zheng, K. Singh, H. Huang, and C. Tai. Model-guided 3D sketching. IEEE Transactions on Visualization and Computer Graphics, 25(10):2927– 2939, October 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xia:2019:TSA

[XGS^{+}19] X. Xia, Y. Guan, A. State, P. Chakravarthula, K. Rathinavel, T. Cham, and H. Fuchs. Towards a switchable AR/ VR near-eye display with accommodation-vergence and eyeglass prescription support. IEEE Transactions on Visualization and Computer Graphics, 25(11):3114– 3124, November 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xin:2012:ECE

[XHF12] Shi-Qing Xin, Ying He, and Chi-Wing Fu. Efficiently computing exact geodesic loops within finite steps. IEEE Transactions on Visualization and Computer Graphics, 18 (6):879–889, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiao:2018:SAT

[XHL18] Yi Xiao, Tze-Yiu Ho, and Chi-Sing Leung. Summed area tables for cube maps. IEEE Transactions on Visualization and Computer Graphics, 24(10):2773–2786, October 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/10/08064721-
abs.html.

Xie:2007:EIV

[XHT^{+}07] Xuexiang Xie, Ying He, Feng Tian, Hock-Soon Seah, Xianfeng Gu, and Hong Qin. An effective illustrative visualization framework based on photic extremum lines (PELs). IEEE Transactions on Visualization and Computer Graphics, 13(6):1328–1335, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2007/06/v1328s.
mov.

Xu:2008:SPC

[XJF^{+}08] Kun Xu, Yun-Tao Jia, Hongbo Fu, Shimin Hu, and Chiew-Lan Tai. Spherical piecewise constant basis functions for all-frequency precomputed radiance transfer. IEEE Transactions on Visualization and Computer Graphics, 14(2):454– 467, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/02/ttg2008020454s.
avi.

Xu:2018:FMM

[XLC^{+}18] Lan Xu, Yebin Liu, Wei Cheng, Kaiwen Guo, Guyue Zhou, Qionghai Dai, and Lu Fang. FlyCap: Markerless motion capture using multiple autonomous flying cameras. IEEE Transactions on Visualization and Computer Graphics, 24(8):2284– 2297, August 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/07983006-
abs.html.

Xiao:2011:FEN

[XLND11] Chunxia Xiao, Meng Liu, Yongwei Nie, and Zhao Dong. Fast exact nearest patch matching for patch-based image editing and processing. IEEE Transactions on Visualization and Computer Graphics, 17(8):1122–1134, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2010:ITF

[XLS10] Lijie Xu, Teng-Yok Lee, and Han-Wei Shen. An information-theoretic framework for flow visualization. IEEE Transactions on Visualization and Computer Graphics, 16(6):1216–1224, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2019:ERD

[XLZ^{+}19a] W. Xu, H. Liang, Y. Zhao, T. Zhang, D. Yu, and D. Monteiro. Errata to RingText: Dwell-Free and Hands-Free Text Entry for Mobile Head-Mounted Displays Using Head Motion. IEEE Transactions on Visualization and Computer Graphics, 25(7):2513, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See [XLZ^{+}19b].

Xu:2019:RDF

[XLZ^{+}19b] Wenge Xu, Hai-Ning Liang, Yuxuan Zhao, Tianyu Zhang, Difeng Yu, and Diego Monteiro. RingText: Dwell-free and hands-free text entry for mobile head-mounted displays using head motions. IEEE Transactions on Visualization and Computer Graphics, 25(5):1991–2001, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642443/. See

errata [XLZ^{+}19a].

Xu:2019:TSL

[XLZ19c] X. Xu, C. Liu, and Y. Zheng. 3D tooth segmentation and labeling using deep convolutional neural networks. IEEE Transactions on Visualization and Computer Graphics, 25(7):2336–2348, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2017:VVD

[XMRC17] Panpan Xu, Honghui Mei, Liu Ren, and Wei Chen. ViDX: Visual diagnostics of assembly line performance in smart factories. IEEE Transactions on Visualization and Computer Graphics, 23(1):291–300, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiao:2011:EEP

[XNT11] Chunxia Xiao, Yongwei Nie, and Feng Tang. Efficient edit propagation using hierarchical data structure. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1135–1147, August 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2012:USC

[XRP^{+}12] Kai Xu, Chris Rooney, Peter Passmore, Dong-Han Ham, and Phong H. Nguyen. A user study on curved edges in graph visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2449–2456, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiao:2018:MAT

[XST^{+}18] Robert Xiao, Julia Schwarz, Nick Throm, Andrew D. Wilson, and Hrvoje Benko. MR-Touch: Adding touch input to head-mounted mixed reality. IEEE Transactions on Visualization and Computer Graphics, 24(4):1653– 1660, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08263123-
abs.html.

Xiong:2017:CMV

[XSZ^{+}17] Guanglei Xiong, Peng Sun, Haoyin Zhou, Seongmin Ha, Briain o Hartaigh, Quynh A. Truong, and James K. Min. Comprehensive modeling and visualization of cardiac anatomy and physiology from CT imaging and computer simulations. IEEE Transactions on Visualization and Computer Graphics, 23(2):1014–1028, February 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07399768-
abs.html.

Xiang:2011:SCE

[XTY^{+}11] Dehui Xiang, Jie Tian, Fei Yang, Qi Yang, Xing Zhang, Qingde Li, and Xin Liu. Skeleton cuts — an efficient segmentation method for volume rendering. IEEE Transactions on Visualization and Computer Graphics, 17(9):1295– 1306, September 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2015:FWP

[XWL^{+}15] Chunxu Xu, Tuanfeng Y.

Wang, Yong-Jin Liu, Ligang Liu, and Ying He. Fast Wavefront Propagation (FWP) for computing exact geodesic distances on meshes. IEEE Transactions on Visualization and Computer Graphics, 21(7):822–834, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
07/07050361-abs.html.

Xu:2013:VAT

[XWW^{+}13] Panpan Xu, Yingcai Wu, Enxun Wei, Tai-Quan Peng, Shixia Liu, Jonathan J. H. Zhu, and Huamin Qu. Visual analysis of topic competition on social media. IEEE Transactions on Visualization and Computer Graphics, 19(12): 2012–2021, December 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xie:2019:VAF

[XXM19a] Cong Xie, Wei Xu, and Klaus Mueller. A visual analytics framework for the detection of anomalous call stack trees in high performance computing applications. IEEE Transactions on Visualization and Computer Graphics, 25(1):

215–224, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440859-
abs.html.

Xu:2019:EEB

[XXM^{+}19b] Ke Xu, Meng Xia, Xing Mu, Yun Wang, and Nan Cao. EnsembleLens: Ensemble-based visual exploration of anomaly detection algorithms with multidimensional data. IEEE Transactions on Visualization and Computer Graphics, 25 (1):109–119, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440102-
abs.html.

Xia:2018:LEA

[XYC^{+}18] J. Xia, F. Ye, W. Chen, Y. Wang, W. Chen, Y. Ma, and A. K. H. Tung. LDSScanner: Exploratory analysis of low-dimensional structures in high-dimensional datasets. IEEE Transactions on Visualization and Computer Graphics, 24(1):236–245, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2013:BVM

[XYGL13] Huanhuan Xu, Wuyi Yu, Shiyuan Gu, and Xin Li. Biharmonic volumetric mapping using fundamental solutions.

IEEE Transactions on Visualization and Computer Graphics, 19(5):787–798, May 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xie:2016:TMR

[XYS^{+}16] Ke Xie, Feilong Yan, Andrei Sharf, Oliver Deussen, Hui Huang, and Baoquan Chen. Tree modeling with real treeparts examples. IEEE Transactions on Visualization and Computer Graphics, 22(12): 2608–2618, December 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2016/12/07368927-
abs.html.

Xu:2014:IMP

[XZB14] Hongyi Xu, Yili Zhao, and Jernej Barbic. Implicit multibody penalty-based distributed contact. IEEE Transactions on Visualization and Computer Graphics, 20(9):1266–1279, September 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xie:2017:VAA

[XZM17] Cong Xie, Wen Zhong, and Klaus Mueller. A visual analytics approach for categorical joint distribution reconstruction from marginal projections. IEEE Transactions on Visualization and Computer Graphics, 23(1):51–60, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2017:AAS

[XZX^{+}17] Xuemiao Xu, Linyuan Zhong, Minshan Xie, Xueting Liu, Jing Qin, and Tien-Tsin Wong. ASCII art synthesis from natural photographs. IEEE Transactions on Visualization and Computer Graphics, 23(8):1910– 1923, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07491376-
                                        abs.html.

Young:2007:DCA

[YAE07] Sean D. Young, Bernard D. Adelstein, and Stephen R. Ellis. Demand characteristics in assessing motion sickness in a virtual environment: Or does taking a motion sickness questionnaire make you sick? IEEE Transactions on Visualization and Computer Graphics, 13(3):422–428, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yi:2007:TDU

[YaKSJ07] Ji Soo Yi, Youn ah Kang, John Stasko, and Julie Jacko. Toward a deeper understanding of the role of interaction in information visualization.

IEEE Transactions on Visualization and Computer Graphics, 13(6):1224–1231, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2018:CAC

[Yan18]  W. Yang.  Context-aware

computer aided inbetweening. IEEE Transactions on Visualization and Computer Graphics, 24(2):1049–1062, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2019:PDH

[YBW^{+}19] Hyeonseung Yu, Mojtaba Bemana, Marek Wernikowski, Michał Chwesiuk, Okan Tarhan Tursun, Gurprit Singh, Karol Myszkowski, Radosław Mantiuk, Hans-Peter Seidel, and Piotr Didyk. A perceptiondriven hybrid decomposition for multi-layer accommodative displays. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1940–1950, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8643583/.

Yan:2014:LRR

[YBZW14] Dong-Ming Yan, Guanbo Bao, Xiaopeng Zhang, and Peter Wonka. Low-resolution remeshing using the localized restricted Voronoi diagram. IEEE Transactions on Visualization and Computer Graphics, 20(10):1418– 1427, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/10/06832586-
                                        abs.html.

Yang:2014:LOL

[YC14] Xin Yang and Kwang-Ting Cheng. Learning optimized local difference binaries for scalable augmented reality on mobile devices. IEEE Transactions on Visualization and Computer Graphics, 20(6): 852–865, June 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2012:IGL

[YCHZ12] Xiaoru Yuan, Limei Che, Yifan Hu, and Xin Zhang. Intelligent graph layout using many users’ input. IEEE Transactions on Visualization and Computer Graphics, 18(12):2699–2708, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yao:2012:RBL

[YCLJ12] Chih-Yuan Yao, Ming-Te Chi, Tong-Yee Lee, and Tao Ju. Region-based line field design using harmonic functions. IEEE Transactions on Visualization and Computer Graphics, 18(6):902–913, June 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yen:2008:SRU

[YCLL08] Chung-Ren Yen, Ming-Te Chi, Tong-Yee Lee, and Wen-Chieh Lin. Stylized rendering using samples of a painted image. IEEE Transactions on Visualization and Computer Graphics, 14(2):468– 480, March/April 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2008/02/ttg2008020468s.
pdf.

Yang:2016:ECI

[YCZ^{+}16] Ming Yang, Hongyang Chao, Chi Zhang, Jun Guo, Lu Yuan, and Jian Sun. Effective clipart image vectorization through direct optimization of bezigons. IEEE Transactions on Visualization and Computer Graphics, 22(2):1063– 1075, February 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2014:GIW

[YDC^{+}14] Bowen Yu, Harish Doraiswamy, Xi Chen, Emily Miraldi, Mario Luis Arrieta-Ortiz, Christoph Hafemeister, Aviv Madar, Richard Bonneau, and Claudio T.

Silva. Genotet: An interactive Web-based visual exploration framework to support validation of gene regulatory networks. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1903–1912, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/12/06876028-
                                        abs.html.

Yoghourdjian:2016:HQU

[YDG^{+}16] V. Yoghourdjian, T. Dwyer, G. Gange, S. Kieffer, K. Klein, and K. Marriott. High-quality ultra-compact grid layout of grouped networks. IEEE Transactions on Visualization and Computer Graphics, 22 (1):339–348, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2017:MMG

[YDGM17] Yalong Yang, Tim Dwyer, Sarah Goodwin, and Kim Marriott. Many-to-many geographically-embedded flow visualisation: An evaluation. IEEE Transactions on Visualization and Computer Graphics, 23(1):411–420, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2019:ODF

[YDJ^{+}19] Yalong Yang, Tim Dwyer, Bernhard Jenny, Kim Marriott, Maxime Cordeil, and Haohui Chen. Origindestination flow maps in immersive environments. IEEE Transactions on Visualization and Computer Graphics, 25 (1):693–703, January 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440844-
abs.html.

Yoghourdjian:2018:GTI

[YDK^{+}18] Vahan Yoghourdjian, Tim Dwyer, Karsten Klein, Kim Marriott, and Michael Wybrow. Graph thumbnails: Identifying and comparing multiple graphs at a glance. IEEE Transactions on Visualization and Computer Graphics, 24 (12):3081–3095, ???? 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8249874/.

Yalcin:2016:ARS

[YEB16] M. A. Yalcin, N. Elmqvist, and B. B. Bederson. AggreSet: Rich and scalable set exploration using visualizations of element aggregations. IEEE Transactions on Visualization and Computer Graphics, 22(1):688–697, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yalcin:2018:KRE

[YEB18] Mehmet Adil Yalcin, Niklas Elmqvist, and Benjamin B. Bederson. Keshif: Rapid and expressive tabular data exploration for novices. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2339–2352, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/08/07968338-
                                        abs.html.

Yu:2012:ESA

[YEII12] Lingyun Yu, Konstantinos Efstathiou, Petra Isenberg, and Tobias Isenberg. Efficient structure-aware selection techniques for 3D point cloud visualizations with 2DOF input. IEEE Transactions on Visualization and Computer Graphics, 18(12):2245–2254, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2016:CEE

[YEII16] L. Yu, K. Efstathiou, P. Isenberg, and T. Isenberg. CAST: Effective and efficient user interaction for context-aware selection in 3D particle clouds. IEEE Transactions on Visualization and Computer Graphics, 22(1):886–895, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yagel:1995:GVR

[YESK95] Roni Yagel, David S. Ebert, James N. Scott, and Yair Kurzion. Grouping volume renderers for enhanced visualization in computational fluid dynamics. IEEE Transactions on Visualization and Computer Graphics, 1(2): 117–132, June 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0117.
pdf;   http://www.computer.
org/tvcg/tg1995/v0117abs.
htm.

Yang:2019:VEC

[YFC^{+}19] Y. Yang, X. Fu, S. Chai, S. Xiao, and L. Liu. Volumeenhanced compatible remeshing of 3D models. IEEE Transactions on Visualization and Computer Graphics, 25(10):2999–3010, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2001:EOR

[YFM01] Yizhou Yu, Andras Ferencz, and Jitendra Malik. Extracting objects from range and radiance images. IEEE Transactions on Visualization and Computer Graphics, 7(4): 351–364, October 2001. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/

                                        tg/books/tg2001/pdf/v0351.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg2001/v0351abs.
                                        htm.

Yu:2018:PTE

[YFZ^{+}18] Difeng Yu, Kaixuan Fan, Heng Zhang, Diego Monteiro, Wenge Xu, and Hai-Ning Liang. PizzaText: Text entry for virtual reality systems using dual thumbsticks. IEEE Transactions on Visualization and Computer Graphics, 24(11):2927– 2935, November 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8456570/.

Yuan:2019:TUD

[YGFX19] M. Yuan, L. Gao, H. Fu, and S. Xia. Temporal upsampling of depth maps using a hybrid camera. IEEE Transactions on Visualization and Computer Graphics, 25 (3):1591–1602, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2013:PBD

[YGV^{+}13] Yin Yang, Xiaohu Guo, Jennell Vick, Luis G. Torres, and Thomas F. Campbell. Physics-based deformable tongue visualization. IEEE Transactions on Visualization and Computer Graphics, 19(5):811–823, May 2013.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2009:SPP

[YGX^{+}09] Xiaoru Yuan, Peihong Guo, He Xiao, Hong Zhou, and Huamin Qu. Scattering points in parallel coordinates. IEEE Transactions on Visualization and Computer Graphics, 15(6):1001–1008, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yin:2019:SAG

[YHH^{+}19] Kangxue Yin, Hui Huang, Edmond S. L. Ho, Hao Wang, Taku Komura, Daniel Cohen-Or, and Hao Zhang. A sampling approach to generating closely interacting 3D pose-pairs from 2D annotations. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2217–2227, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8353147/.

Yeh:2017:IHR

[YHJ^{+}17] Chih-Kuo Yeh, Shi-Yang Huang, Pradeep Kumar Jayaraman, Chi-Wing Fu, and Tong-Yee Lee. Interactive high-relief reconstruction for organic and double-sided objects from a photo. IEEE Transactions on Visualization and Computer Graphics, 23(7):1796–1808, July 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/07/07482721-
abs.html.

Yao:2017:MVM

[YHL^{+}17] Chih-Yuan Yao, Shih-Hsuan Hung, Guo-Wei Li, I-Yu Chen, Reza Adhitya, and Yu-Chi Lai. Manga vectorization and manipulation with procedural simple screentone. IEEE Transactions on Visualization and Computer Graphics, 23(2):1070– 1084, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07399427-
abs.html.

Yang:2008:TLF

[YHLJ08] Ruigang Yang, Xinyu Huang, Sifang Li, and Christopher Jaynes. Toward the light field display: Autostereoscopic rendering via a cluster of projectors. IEEE Transactions on Visualization and Computer Graphics, 14(1):84– 96, January/February 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yan:2008:SDU

[YHMY08] Han-Bing Yan, Shimin Hu, Ralph R. Martin, and Yong-Liang Yang. Shape deformation using a skeleton to drive simplex transformations. IEEE Transactions on Visualization and Computer Graphics, 14(3):693– 706, May/June 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2019:CJV

[YHR^{+}19] F. Yang, L. T. Harrison, R. A.

Rensink, S. L. Franconeri, and R. Chang. Correlation judgment and visualization features: A comparative study. IEEE Transactions on Visualization and Computer Graphics, 25(3):1474–1488, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2007:VRD

[YHW^{+}07] Jing Yang, Daniel Hubball, Matthew O. Ward, Elke A. Rundensteiner, and William Ribarsky. Value and relation display: Interactive visual exploration of large data sets with hundreds of dimensions. IEEE Transactions on Visualization and Computer Graphics, 13(3):494–507, May/June 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yeh:2015:CHM

[YJL^{+}15] Chih-Kuo Yeh, Pradeep Kumar Jayaraman, Xiaopei Liu, Chi-Wing Fu, and Tong-Yee Lee. 2.5D cartoon hair modeling and manipulation. IEEE Transactions on Visualization and Computer Graphics, 21 (3):304–314, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/03/06910280-
                                        abs.html.

Yang:2008:OSP

[YKL^{+}08] Yong-Liang Yang, Junho Kim, Busan Feng Luo, Shi-Min Hu, and Xianfeng Gu. Optimal surface parameterization using inverse curvature map. IEEE Transactions on Visualization and Computer Graphics, 14(5):1054– 1066, September/October 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kwon:2012:SSS

[yKL12] Ji yong Kwon and In-Kwon Lee. The squash-and-stretch stylization for character motions. IEEE Transactions on Visualization and Computer Graphics, 18(3):488– 500, March 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yeo:1995:VRD

[YL95] Boon-Lock Yeo and Bede Liu. Volume rendering of DCT-based compressed 3D scalar data. IEEE Transactions on Visualization and Computer Graphics, 1(1): 29–43, March 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0029.
pdf;   http://www.computer.
org/tvcg/tg1995/v0029abs.
htm.

Yoon:2006:MLB

[YL06] Sung-Eui Yoon and Peter Lindstrom. Mesh layouts for block-based caches. IEEE Transactions on Visualization and Computer Graphics, 12(5):1213–1220, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yao:2008:AGI

[YL08] Chih-Yuan Yao and Tong-Yee Lee. Adaptive geometry image. IEEE Transactions on Visualization and Computer Graphics, 14(4): 948–960, July/August 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/.

Yang:2016:MAE

[YL16] Shan Yang and Ming C.

Lin. MaterialCloning: Acquiring elasticity parameters from images for medical applications. IEEE Transactions on Visualization and Computer Graphics, 22(9):2122– 2135, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yin:2018:SSC

[YL18] Q. Yin and S. Liu. Sounding solid combustibles: Nonpremixed flame sound synthesis for different solid combustibles. IEEE Transactions on Visualization and Computer Graphics, 24(2):1179– 1189, February 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Young:2019:ITR

[YLC^{+}19] Jacob Young, Tobias Langlotz, Matthew Cook, Steven Mills, and Holger Regenbrecht. Immersive telepresence and remote collaboration using mobile and wearable devices. IEEE Transactions on Visualization and Computer Graphics, 25(5): 1908–1918, May 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8642375/.

Yu:2014:CAP

[YLG^{+}14] Zongqiao Yu, Lin Lu, Yanwen Guo, Rongfei Fan, Mingming Liu, and Wenping Wang. Content-aware photo collage using circle packing. IEEE Transactions on Visualization and Computer Graphics, 20 (2):182–195, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yoon:2012:VPB

[YLK12] Jong-Chul Yoon, In-Kwon Lee, and Henry Kang. Video painting based on a stabilized time-varying flow field. IEEE Transactions on Visualization and Computer Graphics, 18 (1):58–67, January 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. See erratum [Ano14h].

Yeh:2012:SED

[YLL^{+}12] I-Cheng Yeh, Wen-Chieh Lin, Tong-Yee Lee, Hsin-Ju Han, Jehee Lee, and Manmyung Kim. Social-event-driven camera control for multicharacter animations. IEEE Transactions on Visualization and Computer Graphics, 18(9):1496–1510, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yeh:2011:TBM

[YLSL11] I-Cheng Yeh, Chao-Hung Lin, Olga Sorkine, and Tong-Yee Lee. Template-based 3D model fitting using dualdomain relaxation. IEEE Transactions on Visualization and Computer Graphics, 17 (8):1178–1190, August 2011.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2012:VOM

[YLX^{+}12] Fei Yang, Qingde Li, Dehui Xiang, Yong Cao, and Jie Tian. A versatile optical model for hybrid rendering of volume data. IEEE Transactions on Visualization and Computer Graphics, 18 (6):925–937, June 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2012:RBR

[YLY^{+}12] Hongchuan Yu, Tong-Yee Lee, I-Cheng Yeh, Xiaosong Yang, Wenxi Li, and Jian J. Zhang. An RBF-based reparameterization method for constrained texture mapping. IEEE Transactions on Visualization and Computer Graphics, 18(7):1115–1124, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2013:PVE

[YLZ^{+}13] Jing Yang, Yujie Liu, Xin Zhang, Xiaoru Yuan, Ye Zhao, Scott Barlowe, and Shixia Liu. PIWI: Visually exploring graphs based on their community structure. IEEE Transactions on Visualization and Computer Graphics, 19(6): 1034–1047, June 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2017:PFS

[YML^{+}17] Tao Yang, Ralph R. Martin, Ming C. Lin, Jian Chang, and Shi-Min Hu. Pairwise force SPH model for realtime multi-interaction applications. IEEE Transactions on Visualization and Computer Graphics, 24(10):2235– 2247, October 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/10/07932108-
abs.html.

Yamane:2003:NMA

[YN03] Katsu Yamane and Yoshihiko Nakamura. Natural motion animation through constraining and deconstraining at will. IEEE Transactions on Visualization and Computer Graphics, 9(3):352–360, July/September 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/03/v0352abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/03/
v0352.pdf.

Yost:2006:PSV

[YN06] Beth Yost and Chris North.

The perceptual scalability of visualization. IEEE Transactions on Visualization and Computer Graphics, 12(5): 837–844, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2011:LTA

[YNBH11] Qizhi Yu, Fabrice Neyret, Eric Bruneton, and Nicolas Holzschuch. Lagrangian texture advection: Preserving both spectrum and velocity field. IEEE Transactions on Visualization and Computer Graphics, 17(11):1612– 1623, November 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2006:HVH

[YNCP06] Xiaoru Yuan, Minh X. Nguyen, Baoquan Chen, and David H. Porter. HDR VolVis: High dynamic range volume visualization. IEEE Transactions on Visualization and Computer Graphics, 12(4): 433–445, July/August 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2015:LGC

[YNM15] Liming Yang, Jean-Marie Normand, and Guillaume Moreau. Local geometric consensus: A general purpose point pattern-based tracking algorithm. IEEE Transactions on Visualization and Computer Graphics, 21(11): 1299–1308, November 2015.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/11/07165652-
abs.html.

Ynnerman:2019:VTA

[Ynn19]  Anders Ynnerman.  The

2018 Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 25 (1):xxix, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08570931.
pdf.

Yamamoto:2006:ETD

[YNYH06] Akio Yamamoto, Shuichi Nagasawa, Hiroaki Yamamoto, and Toshiro Higuchi. Electrostatic tactile display with thin film slider and its application to tactile telepresentation systems. IEEE Transactions on Visualization and Computer Graphics, 12(2): 168–177, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2005:RBP

[YON05] M. L. Yuan, S. K. Ong, and A. Y. C. Nee. Registration based on projective reconstruction technique for augmented reality systems. IEEE Transactions on Visualization and Computer Graphics, 11 (3):254–264, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2006:RUN

[YON06] M. L. Yuan, S. K. Ong, and A. Y. C. Nee. Registration using natural features for augmented reality systems. IEEE Transactions on Visualization and Computer Graphics, 12(4):569– 580, July/August 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yong:2014:TPW

[Yon14] Jun-Hai Yong. Towards photo watercolorization with artistic verisimilitude. IEEE Transactions on Visualization and Computer Graphics, 20(10): 1451–1460, October 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/10/06732968-
                                        abs.html.

Yamanaka:2013:SPR

[YOS13] Daiki Yamanaka, Yutaka Ohtake, and Hiromasa Suzuki. The sinogram polygonizer for reconstructing 3D shapes. IEEE Transactions on Visualization and Computer Graphics, 19(11):1911–1922, November 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2013:LCB

[YPI13] Ying Yang, N. Peyerimhoff, and I. Ivrissimtzis. Linear correlations between spatial and normal noise in triangle meshes. IEEE Transactions on Visualization and Computer Graphics, 19(1):45– 55, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2017:SAS

[YPRI17] Ying Yang, Ruggero Pintus, Holly Rushmeier, and Ioannis Ivrissimtzis. A 3D steganalytic algorithm and steganalysis-resistant watermarking. IEEE Transactions on Visualization and Computer Graphics, 23(2):1002– 1013, February 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/02/07399411-
abs.html.

Ye:2017:MEH

[YQK^{+}17] Tian Ye, Siyuan Qi, James Kubricht, Yixin Zhu, Hongjing Lu, and Song-Chun Zhu. The Martian: Examining human physical judgments across virtual gravity fields. IEEE Transactions on Visualization and Computer Graphics, 23(4):1399–1408, April 2017.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07829434-
abs.html.

Yao:1995:HSC

[YR95] Chengfu Yao and Jon G. Rokne. Hybrid scan-conversion of circles. IEEE Transactions on Visualization and Computer Graphics, 1(4):311– 318, December 1995. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg1995/pdf/v0311.
pdf;   http://www.computer.
org/tvcg/tg1995/v0311abs.
htm.

Yadav:2018:MDB

[YRP18] Sunil Kumar Yadav, Ulrich Reitebuch, and Konrad Polthier. Mesh denoising based on normal voting tensor and binary optimization. IEEE Transactions on Visualization and Computer Graphics, 24 (8):2366–2379, August 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/08/08012522-
abs.html.

Yadav:2019:RHF

[YRP19] Sunil Kumar Yadav, Ulrich Reitebuch, and Konrad Polthier. Robust and high fidelity mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 25(6): 2304–2310, June 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8344461/.

Yuan:2013:DPM

[YRWG13] Xiaoru Yuan, Donghao Ren, Zuchao Wang, and Cong Guo. Dimension projection matrix/tree: Interactive subspace visual exploration and analysis of high dimensional data. IEEE Transactions on Visualization and Computer Graphics, 19(12):2625– 2633, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yemez:2003:MRT

[YS03] Yücel Yemez and Francis Schmitt. Multilevel representation and transmission of real objects with progressive octree particles. IEEE Transactions on Visualization and Computer Graphics, 9(4):551– 569, October/December 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0551abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0551.pdf.

Yu:2017:VWB

[YS17] Bowen Yu and Cláudio T. Silva. VisFlow — Webbased visualization framework for tabular data with a subset flow model. IEEE Transactions on Visualization and Computer Graphics, 23(1): 251–260, January 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2017:BHB

[YSD^{+}17] Xinsong Yang, Lei Shi, Madelaine Daianu, Hanghang Tong, Qingsong Liu, and Paul Thompson. Blockwise human brain network visual comparison using NodeTrix representation. IEEE Transactions on Visualization and Computer Graphics, 23(1):181–190, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yoon:2005:QVC

[YSGM05] Sung-Eui Yoon, Brian Salomon, Russell Gayle, and Dinesh Manocha. Quick-VDR: Out-of-core view-dependent rendering of gigantic models. IEEE Transactions on Visualization and Computer Graphics, 11(4):369– 382, July/August 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2010:FDT

[YSI^{+}10] Lingyun Yu, Pjotr Svetachov, Petra Isenberg, Maarten H. Everts, and Tobias Isenberg. FI3D: Direct-touch interaction for the exploration of 3D scientific visualization spaces. IEEE Transactions on Visualization and Computer Graphics, 16(6):1613–1622, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Song:2007:DPS

[ySKK07] Oh young Song, Doyub Kim, and Hyeong-Seok Ko. Derivative particles for simulating detailed movements of fluids. IEEE Transactions on Visualization and Computer Graphics, 13(4):711– 719, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yeh:2013:DSG

[YSL^{+}13] Chih-Kuo Yeh, Peng Song, Peng-Yen Lin, Chi-Wing Fu, Chao-Hung Lin, and Tong-Yee Lee. Double-sided 2.5D graphics. IEEE Transactions on Visualization and Computer Graphics, 19(2):225– 235, February 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yoo:2012:TIM

[YSS^{+}12] Sang Wook Yoo, Joon-Kyung Seong, Min-Hyuk Sung, Sung Yong Shin, and Elaine Cohen. A triangulationinvariant method for anisotropic geodesic map computation on surface meshes. IEEE Transactions on Visualization and Computer Graphics, 18(10):1664–1677, October 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yan:2004:MSH

[YSZ04] Jingqi Yan, Pengfei Shi, and David Zhang. Mesh simplification with hierarchical shape analysis and iterative edge contraction. IEEE Transactions on Visualization and Computer Graphics, 10(2): 142–151, March/April 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2004/02/v0142abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2004/02/
v0142.htm;   http://csdl.
computer.org/dl/trans/tg/
2004/02/v0142.pdf.

Yue:2019:BIV

[YSZ^{+}19] Xuanwu Yue, Xinhuan Shu, Xinyu Zhu, Xinnan Du, Zheqing Yu, Dimitrios Papadopoulos, and Siyuan Liu. BitExTract: Interactive visualization for extracting Bitcoin exchange intelligence. IEEE Transactions on Visualization and Computer Graphics, 25(1):162– 171, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440044-
abs.html.

Yngve:2002:RCI

[YT02] Gary Yngve and Greg Turk.

Robust creation of implicit surfaces from polygonal meshes. IEEE Transactions on Visualization and Computer Graphics, 8(4):346– 359, October/December 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2002/04/v0346abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2002/04/
v0346.htm;   http://csdl.
computer.org/dl/trans/tg/
2002/04/v0346.pdf.

Yan:2016:NOR

[YW16] Dong-Ming Yan and Peter Wonka. Non-obtuse remeshing with centroidal Voronoi tessellation. IEEE Transactions on Visualization and Computer Graphics, 22(9): 2136–2144, ???? 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2012:HSB

[YWSC12] Hongfeng Yu, Chaoli Wang, Ching-Kuang Shene, and Jacqueline H. Chen. Hierarchical streamline bundles. IEEE Transactions on Visualization and Computer Graphics, 18(8):1353–1367, August 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2019:PEC

[YWV^{+}19] H. Yang, B. Wang, N. Vesdapunt, M. Guo, and S. B. Kang. Personalized exposure control using adaptive metering and reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 25(10):2953– 2968, October 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yan:2014:USM

[YWW14] Dong-Ming Yan, Johannes Wallner, and Peter Wonka. Unbiased sampling and meshing of isosurfaces. IEEE Transactions on Visualization and Computer Graphics, 20(11):1579–1589, November 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

                                        URL http://csdl.computer.
                                        org/csdl/trans/tg/2014/
                                        11/06811174-abs.html.

Yuan:2010:SMV

[YXG^{+}10] Xiaoru Yuan, He Xiao, Hanqi Guo, Peihong Guo, Wesley Kendall, Jian Huang, and Yongxian Zhang. Scalable multi-variate analytics of seismic and satellite-based observational data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1413–1420, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2013:BAM

[YXG^{+}13] Yin Yang, Weiwei Xu, Xiaohu Guo, Kun Zhou, and Baining Guo. Boundary-aware multidomain subspace deformation. IEEE Transactions on Visualization and Computer Graphics, 19(10):1633– 1645, October 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2015:SPD

[YXM^{+}15] Hongfeng Yu, Jinrong Xie, Kwan-Liu Ma, Hemanth Kolla, and Jacqueline H. Chen. Scalable parallel distance field construction for large-scale applications. IEEE Transactions on Visualization and Computer Graphics, 21(10):1187– 1200, October 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/10/07072474-
abs.html.

Ying:2013:IAP

[YXSH13] Xiang Ying, Shi-Qing Xin, Qian Sun, and Ying He. An intrinsic algorithm for parallel Poisson disk sampling on arbitrary surfaces. IEEE Transactions on Visualization and Computer Graphics, 19(9): 1425–1437, September 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2019:SES

[YYD^{+}19] M. Yang, J. Ye, F. Ding, Y. Zhang, and D. Yan. A semi-explicit surface tracking mechanism for multi-phase immiscible liquids. IEEE Transactions on Visualization and Computer Graphics, 25(10):2873–2885, October 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2018:SRF

[YYFX18] L. Yang, Q. Yan, Y. Fu, and C. Xiao. Surface reconstruction via fusing sparsesequence of depth images. IEEE Transactions on Visualization and Computer Graphics, 24(2):1190–1203, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yan:2017:DVS

[YYR17] Zhixin Yan, Mao Ye, and Liu Ren. Dense visual SLAM with probabilistic surfel map. IEEE Transactions on Visualization and Computer Graphics, 23(11):2389– 2398, November 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/11/08007295-
abs.html.

Yan:2006:MPM

[YYSZ06] Jingqi Yan, Xin Yang, Pengfei Shi, and David Zhang. Mesh parameterization by minimizing the synthesized distortion metric with the coefficientoptimizing algorithm. IEEE Transactions on Visualization and Computer Graphics, 12 (1):83–92, January/February 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yu:2016:CIT

[YYT16] Lap-Fai Yu, Sai-Kit Yeung, and D. Terzopoulos. The clutterpalette: An interactive tool for detailing indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 22(2):1138–1148, February 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2016:FIO

[YYY16] Y. Yang, X. Yang, and S. Yang. A fast iterated orthogonal projection framework for smoke simulation. IEEE Transactions on Visualization and Computer Graphics, 22(5):1492–1502, ????

2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zagajac:1996:FME

[Zag96]  Jovan Zagajac.  A fast

method for estimating discrete field values in early engineering design. IEEE Transactions on Visualization and Computer Graphics, 2 (1):35–43, March 1996. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1996/pdf/v0035.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1996/v0035abs.
                                        htm.

Zheng:2011:IFC

[ZAM11] Ziyi Zheng, Nafees Ahmed, and Klaus Mueller. iView: a feature clustering framework for suggesting informative views in volume visualization. IEEE Transactions on Visualization and Computer Graphics, 17(12):1959– 1968, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2006:LCP

[ZBB^{+}06] Lingxiao Zhao, Charl Botha, Javier Bescos, Roel Truyen, Frans Vos, and Frits Post. Lines of curvature for polyp detection in virtual colonoscopy. IEEE Transactions on Visualization and Computer Graphics, 12(5):885–892, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

zuBerge:2014:PBF

[zBBKN14] Christian Schulte zu Berge, Maximilian Baust, Ankur Kapoor, and Nassir Navab. Predicate-based focus-and-context visualization for 3D ultrasound. IEEE Transactions on Visualization and Computer Graphics, 20(12): 2379–2387, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876031-
abs.html.

Zinsmaier:2012:ILD

[ZBDS12] Michael Zinsmaier, Ulrik Brandes, Oliver Deussen, and Hendrik Strobelt. Interactive level-of-detail rendering of large graphs. IEEE Transactions on Visualization and Computer Graphics, 18(12): 2486–2495, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2017:VSE

[ZBG^{+}17] Henan Zhao, Garnett W.

Bryant, Wesley Griffin, Judith E. Terrill, and Jian Chen. Validation of SplitVectors encoding for quantitative visualization of large-magnitude-range vector fields. IEEE Transactions on Visualization and Computer Graphics, 23(6):1691–1705, June 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/06/07429760-
                                        abs.html.

Zordan:2014:CRD

[ZBMY14] Victor Zordan, David Brown, Adriano Macchietto, and KangKang Yin. Control of rotational dynamics for ground and aerial behavior. IEEE Transactions on Visualization and Computer Graphics, 20(10):1356– 1366, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2014/10/06832615-
                                        abs.html.

Zurdo:2013:AWE

[ZBO13] J. S. Zurdo, J. P. Brito, and M. A. Otaduy. Animating wrinkles by example on nonskinned cloth. IEEE Transactions on Visualization and Computer Graphics, 19(1): 149–158, January 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2013:INR

[ZBZ^{+}13] Jiahua Zhang, George Baciu, Dejun Zheng, Cheng Liang, Guiqing Li, and Jinlian Hu. IDSS: A novel representation for woven fabrics. IEEE Transactions on Visualization and Computer Graphics, 19 (3):420–432, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2003:SSS

[ZC03] Caixia Zhang and Roger Crawfis. Shadows and soft shadows with participating media using splatting. IEEE Transactions on Visualization and Computer Graphics, 9 (2):139–149, April/June 2003. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/02/v0139abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/02/
v0139.htm;   http://csdl.
computer.org/dl/trans/tg/
2003/02/v0139.pdf.

Zheng:2006:IAT

[ZC06] Jianmin Zheng and Yiyu Cai.

Interpolation over arbitrary topology meshes using a twophase subdivision scheme. IEEE Transactions on Visualization and Computer Graphics, 12(3):301–310, May/June 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2012:FDA

[ZCCB12] Jian Zhao, Fanny Chevalier, Christopher Collins, and Ravin Balakrishnan. Facilitating discourse analysis with interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 18(12):2639– 2648, December 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2013:IEI

[ZCCB13] Jian Zhao, Christopher Collins, Fanny Chevalier, and Ravin Balakrishnan. Interactive exploration of implicit and explicit relations in faceted datasets. IEEE Transactions on Visualization and Computer Graphics, 19(12):2080– 2089, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:ITE

[ZCD19] Yixuan Zhang, Kartik Chanana, and Cody Dunne. IDMVis: Temporal event sequence visualization for Type 1 diabetes treatment decision support. IEEE Transactions on Visualization and Computer Graphics, 25(1):512– 522, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440834-
abs.html.

Zheng:2019:STS

[ZCF^{+}19] Y. Zheng, Y. Chen, G. Fei, J. Dorsey, and E. Wu. Simulation of textile stains. IEEE Transactions on Visualization and Computer Graphics, 25(7):2471–2481, July 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zou:2015:PRP

[ZCFL15] Changqing Zou, Shifeng Chen, Hongbo Fu, and Jianzhuang Liu. Progressive 3D reconstruction of planarfaced manifold objects with DRF-based line drawing decomposition. IEEE Transactions on Visualization and Computer Graphics, 21(2): 252–263, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/06891368-
abs.html.

Zhang:2012:IPT

[ZCJH12] Fang-Lue Zhang, Ming-Ming Cheng, Jiaya Jia, and Shi-Min Hu. ImageAdmixture: Putting together dissimilar objects from groups. IEEE Transactions on Visualization and Computer Graphics, 18(11):1849–1857, November 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2008:IWM

[ZCL08] Song Zhang, Stephen Correia, and David H. Laidlaw. Identifying whitematter fiber bundles in DTI data using an automated proximity-based fiberclustering method. IEEE Transactions on Visualization and Computer Graphics, 14(5):1044–1053, September/October 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2009:VIS

[ZCL09] Jian Zhang, Chaomei Chen, and Jiexun Li. Visualizing the intellectual structure with paper-reference matrices. IEEE Transactions on Visualization and Computer Graphics, 15(6):1153– 1160, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:MNE

[ZCL^{+}19] Jie Zhang, Junjie Cao, Xiuping Liu, He Chen, Bo Li, and Ligang Liu. Multi-normal estimation via pair consistency voting. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1693–1706, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8340177/.

Zhao:2011:EAT

[ZCPB11] Jian Zhao, Fanny Chevalier, Emmanuel Pietriga, and Ravin Balakrishnan. Exploratory analysis of timeseries with ChronoLenses. IEEE Transactions on Visualization and Computer Graphics, 17(12):2422–2431, December 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2014:FVA

[ZCW^{+}14] Jian Zhao, Nan Cao, Zhen Wen, Yale Song, Yu-Ru Lin, and Christopher Collins. #FluxFlow: Visual analysis of anomalous information spreading on social media. IEEE Transactions on Visualization and Computer Graphics, 20(12):1773– 1782, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876013-
abs.html.

Zhou:2019:ELS

[ZCW19] Bo Zhou, Yi-Jen Chiang, and Cong Wang. Efficient local statistical analysis via pointwise histograms in tetrahedral meshes and curvilinear grids. IEEE Transactions on Visualization and Computer Graphics, 25(2):1392– 1406, February 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/02/08267140-
abs.html.

Zhang:2009:VCA

[ZCZ^{+}09] Song-Hai Zhang, Tao Chen, Yi-Fei Zhang, Shi-Min Hu, and Ralph R. Martin. Vectorizing cartoon animations. IEEE Transactions on Visualization and Computer Graphics, 15(4):618–629, July/ August 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zirr:2018:MEF

[ZD18] T. Zirr and C. Dachsbacher. Memory-efficient on-the-fly voxelization and rendering of particle data. IEEE Transactions on Visualization and Computer Graphics, 24(2): 1155–1166, February 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:SDF

[ZDH^{+}19] Juyong Zhang, Bailin Deng, Yang Hong, Yue Peng, Wenjie Qin, and Ligang Liu. Static/ dynamic filtering for mesh geometry. IEEE Transactions on Visualization and Computer Graphics, 25(4): 1774–1787, April 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8327892/.

Zhang:2009:RDI

[ZDJ^{+}09] Guofeng Zhang, Zilong Dong, Jiaya Jia, Liang Wan, Tien-Tsin Wong, and Hujun Bao. Refilming with depth-inferred videos. IEEE Transactions on Visualization and Computer Graphics, 15(5): 828–840, September/October 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2003:VDT

[ZDL03] Song Zhang, Çagatay Demiralp, and David H. Laidlaw. Visualizing diffusion tensor MR images using streamtubes and streamsurfaces. IEEE Transactions on Visualization and Computer Graphics, 9(4):454–462, October/December 2003. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2003/04/v0454abs.
htm;   http://csdl.computer.
org/dl/trans/tg/2003/04/
v0454.pdf.

Zhang:2013:RTV

[ZDM13] Yubo Zhang, Zhao Dong, and Kwan-Liu Ma. Real-time volume rendering in dynamic lighting environments using precomputed photon mapping. IEEE Transactions on Visualization and Computer Graphics, 19(8):1317– 1330, August 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2005:DSB

[ZDW^{+}05] Kun Zhou, Peng Du, Lifeng Wang, Yasuyuki Matsushita, Jiaoying Shi, Baining Guo, and Heung-Yeung Shum. Decorating surfaces with bidirectional texture functions. IEEE Transactions on Visualization and Computer Graphics, 11(5):519–528, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2018:IDC

[ZDZ18] Shuang Zhao, Fredo Durand, and Changxi Zheng. Inverse diffusion curves using shape optimization. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2153–2166, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07962259-
                                        abs.html.

Zheng:2008:IVD

[ZEC08] Zhi Zheng, Prakash Edmond, and Tony Chan. Interactive view-dependent rendering over networks. IEEE Transactions on Visualization and Computer Graphics, 14 (3):576–589, May/June 2008. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zeng:2014:VMP

[ZFA^{+}14] Wei Zeng, Chi-Wing Fu, Stefan Muller Arisona, Alexander Erath, and Huamin Qu. Visualizing mobility of public transportation system. IEEE Transactions on Visualization and Computer Graphics, 20(12):1833– 1842, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876029-
abs.html.

Zheng:2011:BNF

[ZFAT11] Youyi Zheng, Hongbo Fu, Oscar Kin-Chung Au, and Chiew-Lan Tai. Bilateral normal filtering for mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 17(10):1521– 1530, October 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2017:GMD

[ZFL17] Junsong Zhang, Jiepeng Fan, and Zhenshan Luo. Generating multi-destination maps. IEEE Transactions on Visualization and Computer Graphics, 23(8):1964– 1976, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/

                                        csdl/trans/tg/2017/08/07530932-
                                        abs.html.

Zhang:2019:VBF

[ZFS^{+}19] Hui Zhang, Steffen Frey, Holger Steeb, David Uribe, Thomas Ertl, and Wenping Wang. Visualization of bubble formation in porous media. IEEE Transactions on Visualization and Computer Graphics, 25(1):1060– 1069, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08445644-
                                        abs.html.

Zhang:2019:UMM

[ZFSL19] J. Zhang, Z. Fan, D. Sun, and H. Liao. Unified mathematical model for multilayermultiframe compressive light field displays using LCDs. IEEE Transactions on Visualization and Computer Graphics, 25(3):1603–1614, March 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2006:IPB

[ZG06] Yuan Zhou and Michael Garland. Interactive pointbased rendering of higherorder tetrahedral data. IEEE Transactions on Visualization and Computer Graphics, 12(5):1229–1236, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zagorchev:2012:CAI

[ZG12] Lyubomir G. Zagorchev and A. Ardeshir Goshtasby. A curvature-adaptive implicit surface reconstruction for irregularly spaced points. IEEE Transactions on Visualization and Computer Graphics, 18(9):1460–1473, September 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2017:AGG

[ZGB^{+}17] Jian Zhao, Michael Glueck, Simon Breslav, Fanny Chevalier, and Azam Khan. Annotation graphs: A graphbased visualization for metaanalysis of data based on userauthored annotations. IEEE Transactions on Visualization and Computer Graphics, 23 (1):261–270, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zgraggen:2017:HPV

[ZGC^{+}17] Emanuel Zgraggen, Alex Galakatos, Andrew Crotty, Jean-Daniel Fekete, and Tim Kraska. How progressive visualizations affect exploratory analysis. IEEE Transactions on Visualization and Computer Graphics, 23(8):1977– 1987, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/08/07563865-
                                        abs.html.

Zhang:2018:DLB

[ZGH^{+}18] J. Zhang, H. Guo, F. Hong, X. Yuan, and T. Peterka. Dynamic load balancing based on constrained K–D tree decomposition for parallel particle tracing. IEEE Transactions on Visualization and Computer Graphics, 24(1): 954–963, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2011:DPO

[ZGHG11] Kun Zhou, Minmin Gong, Xin Huang, and Baining Guo. Data-parallel octrees for surface reconstruction. IEEE Transactions on Visualization and Computer Graphics, 17 (5):669–681, May 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2018:SHA

[ZGI^{+}18] J. Zhao, M. Glueck, P. Isenberg, F. Chevalier, and A. Khan. Supporting handoff in asynchronous collaborative sensemaking using knowledgetransfer graphs. IEEE Transactions on Visualization and Computer Graphics, 24(1): 340–350, January 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhu:2018:DDA

[ZGM18] Jie Zhu, Yanwen Guo, and Han Ma. A data-driven approach for furniture and indoor scene colorization. IEEE Transactions on Visualization and Computer Graphics, 24(9):2473–2486, September 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/09/08039524-
abs.html.

Zhang:2014:PPE

[ZGW^{+}14] Chenxi Zhang, Jizhou Gao, Oliver Wang, Pierre Georgel, Ruigang Yang, James Davis, Jan-Michael Frahm, and Marc Pollefeys. Personal photograph enhancement using Internet photo collections. IEEE Transactions on Visualization and Computer Graphics, 20(2):262–275, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:HTM

[ZGW^{+}19] J. Zhang, M. Gui, Q. Wang, R. Liu, J. Xu, and S. Chen. Hierarchical topic model based object association for semantic SLAM. IEEE Transactions on Visualization and Computer Graphics, 25(11): 3052–3062, November 2019.

CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2007:SDH

[ZH07] Hui Zhang and Andrew Hanson. Shadow-driven 4D haptic visualization. IEEE Transactions on Visualization and Computer Graphics, 13(6):1688–1695, November/December 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2014:MHA

[Zha14] Nan Zhang. Memory-hazard-aware K-buffer algorithm for order-independent transparency rendering. IEEE Transactions on Visualization and Computer Graphics, 20 (2):238–248, February 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2018:LSM

[ZHC18] Lingyun Zhao, Miles Hansard, and Andrea Cavallaro. Layered scene models from single hazy images. IEEE Transactions on Visualization and Computer Graphics, 24(7): 2167–2179, July 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/07/07938419-
                                        abs.html.

Zhao:2007:VSH

[ZHF^{+}07] Ye Zhao, Yiping Han, Zhe Fan, Feng Qiu, Yu-Chuan Kuo, Arie E. Kaufman, and Klaus Mueller. Visual simulation of heat shimmering and mirage. IEEE Transactions on Visualization and Computer Graphics, 13(1):179– 189, January/February 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2012:EEC

[ZHF12] Lei Zhang, Hua Huang, and Hongbo Fu. EXCOL: an EXtract-and-COmplete layering approach to cartoon animation reusing. IEEE Transactions on Visualization and Computer Graphics, 18(7): 1156–1169, July 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2019:EDS

[ZHF^{+}19] Qian Zhou, Georg Hagemann, Dylan Fafard, Ian Stavness, and Sidney Fels. An evaluation of depth and size perception on a spherical fish tank virtual reality display. IEEE Transactions on Visualization and Computer Graphics, 25(5):2040–2049, May 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8642347/.

Zou:2011:APG

[ZHGH11] Guangyu Zou, Jiaxi Hu, Xianfeng Gu, and Jing Hua. Authalic parameterization of general surfaces using Lie advection. IEEE Transactions on Visualization and Computer Graphics, 17(12):2005– 2014, December 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zou:2009:IGS

[ZHL^{+}09] Guangyu Zou, Jing Hua, Zhaoqiang Lai, Xianfeng Gu, and Ming Dong. Intrinsic geometric scale space by shape diffusion. IEEE Transactions on Visualization and Computer Graphics, 15(6):1193– 1200, November/December 2009. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2013:DCA

[ZHLB13] Muyang Zhang, Jin Huang, Xinguo Liu, and Hujun Bao. A divide-and-conquer approach to quad remeshing. IEEE Transactions on Visualization and Computer Graphics, 19(6):941–952, June 2013. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zollmann:2014:FAR

[ZHLR14] Stefanie Zollmann, Christof Hoppe, Tobias Langlotz, and Gerhard Reitmayr. FlyAR: Augmented reality supported micro aerial vehicle navigation. IEEE Transactions on Visualization and Computer Graphics, 20(4): 560–568, April 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2016:SLA

[Zho16]  Kun Zhou.  Simultane-

ous localization and appearance estimation with a consumer RGB-D camera. IEEE Transactions on Visualization and Computer Graphics, 22(8):2012–2023, August 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2016/
08/07321825.pdf.

Zhang:2007:SVS

[ZHQ^{+}07] Guofeng Zhang, Wei Hua, Xueying Qin, Tien-Tsin Wong, and Hujun Bao. Stereoscopic video synthesis from a monocular video. IEEE Transactions on Visualization and Computer Graphics, 13(4):686–696, July/ August 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2007:ITF

[ZHT07] Eugene Zhang, James Hays, and Greg Turk. Interactive tensor field design and visualization on surfaces. IEEE Transactions on Visualization and Computer Graphics, 13(1):94–107, January/February 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhu:2005:URA

[Zhu05]  Yuanchen Zhu.  Uniform

remeshing with an adaptive domain: a new scheme for view-dependent level-of-detail rendering of meshes. IEEE Transactions on Visualization and Computer Graphics, 11 (3):306–316, May/June 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2011:RTS

[ZHX^{+}11] L. Zhang, Y. He, J. Xia, X. Xie, and W. Chen. Realtime shape illustration using Laplacian lines. IEEE Transactions on Visualization and Computer Graphics, 17 (7):993–1006, July 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zang:2015:GAI

[ZHZ15] Yu Zang, Hua Huang, and Lei Zhang. Guided adaptive image smoothing via directional anisotropic structure measurement. IEEE Transactions on Visualization and Computer Graphics, 21(9):1015–1027, September 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
09/07055305-abs.html.

Zhou:2007:TRS

[ZJH07] Qian-Yi Zhou, Tao Ju, and Shi-Min Hu. Topology repair of solid models using skeletons. IEEE Transactions on Visualization and Computer Graphics, 13(4):675– 685, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2011:MIH

[ZJH^{+}11] Guofeng Zhang, Hanqing Jiang, Jin Huang, Jiaya Jia, Tien-Tsin Wong, Kun Zhou, and Hujun Bao. Motion imitation with a handheld camera. IEEE Transactions on Visualization and Computer Graphics, 17(10):1475–1486, October 2011. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2015:PSA

[ZJX^{+}15] Yandan Zhao, Xiaogang Jin, Yingqing Xu, Hanli Zhao, Meng Ai, and Kun Zhou. Parallel style-aware image cloning for artworks. IEEE Transactions on Visualization and Computer Graphics, 21(2): 229–240, February 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/02/06893032-
                                        abs.html.

Zhang:2006:IPB

[ZK06] Haitao Zhang and Arie Kaufman. Interactive point-based isosurface exploration and high-quality rendering. IEEE Transactions on Visualization and Computer Graphics, 12(5):1267–1274, September/October 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2007:ICD

[ZK07] Xinyu Zhang and Young J. Kim. Interactive collision detection for deformable models using streaming AABBs. IEEE Transactions on Visualization and Computer Graphics, 13(2):318–329, March/ April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ziemkiewicz:2008:SIV

[ZK08] Caroline Ziemkiewicz and Robert Kosara. The shaping of information by visual metaphors. IEEE Transactions on Visualization and Computer Graphics, 14(6):1269–1276, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ziemkiewicz:2010:LAP

[ZK10] Caroline Ziemkiewicz and Robert Kosara. Laws of attraction: From perceptual forces to conceptual similarity. IEEE Transactions on Visualization and Computer Graphics, 16(6):1009– 1016, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2012:SCM

[ZK12] Xinyu Zhang and Young J.

Kim. Simple culling methods for continuous collision detection of deforming triangles. IEEE Transactions on Visualization and Computer Graphics, 18(7):1146–1155, July 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2014:SCD

[ZK14a] Xinyu Zhang and Young J.

Kim. Scalable collision detection using p-partition fronts on many-core processors. IEEE Transactions on Visualization and Computer Graphics, 20(3):447–456, March 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zimmer:2014:ZRF

[ZK14b] Henrik Zimmer and Leif Kobbelt. Zometool rationalization of freeform surfaces. IEEE Transactions on Visualization and Computer Graphics, 20(10):1461– 1473, October 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/10/06747390-
abs.html.

Zenner:2017:SWS

[ZK17] Andre Zenner and Antonio Kruger. Shifty: A weightshifting dynamic passive haptic proxy to enhance object perception in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 23(4):1285– 1294, April 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/04/07833030-
abs.html.

Ziegler:2007:FHS

[ZKG07] Remo Ziegler, Peter Kaufmann, and Markus Gross. A framework for holographic scene representation and image synthesis. IEEE Transactions on Visualization and Computer Graphics, 13(2): 403–415, March/April 2007. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://bell.computer.org/
dlcomments/.

Zigelman:2002:TMU

[ZKK02] G. Zigelman, R. Kimmel, and N. Kiryati. Texture mapping using surface flattening via multidimensional scaling. IEEE Transactions on Visualization and Computer Graphics, 8(2):198–207, April 2002. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://dlib.computer.org/
tg/books/tg2002/pdf/v0198.
pdf;   http://www.computer.
org/tvcg/tg2002/v0198abs.
htm.

Zibrek:2018:ERA

[ZKM18] Katja Zibrek, Elena Kokkinara, and Rachel Mcdonnell. The effect of realistic appearance of virtual characters in immersive environments — does the character’s personality play a role? IEEE Transactions on Visualization and Computer Graphics, 24(4):1681–1690, April 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08267290-
abs.html.

Zanbaka:2005:CPV

[ZLB^{+}05] Catherine A. Zanbaka, Benjamin C. Lok, Sabarish V. Babu, Amy C. Ulinski, and Larry F. Hodges. Comparison of path visualizations and cognitive measures relative to travel technique in a virtual environment. IEEE Transactions on Visualization and Computer Graphics, 11(6):694–705, November/December 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2019:EMD

[ZLC^{+}19] Ying Zhao, Feng Luo, Minghui Chen, Yingchao Wang, Jiazhi Xia, Fangfang Zhou, Yunhai Wang, Yi Chen, and Wei Chen. Evaluating multi-dimensional visualizations for understanding fuzzy clusters. IEEE Transactions on Visualization and Computer Graphics, 25(1): 12–21, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08440829-
                                        abs.html.

Zhang:2014:DDS

[ZLD^{+}14] Xiaopeng Zhang, Hongjun Li, Mingrui Dai, Wei Ma, and Long Quan. Data-driven synthetic modeling of trees. IEEE Transactions on Visualization and Computer Graphics, 20 (9):1, September 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zheng:2016:EIR

[ZLDM16] Youyi Zheng, Han Liu, Julie Dorsey, and Niloy J. Mitra. Ergonomics-inspired reshaping and exploration of collections of models. IEEE Transactions on Visualization and Computer Graphics, 22 (6):1732–1744, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2006:GDP

[ZLG^{+}06] Qingshan Zhang, Zicheng Liu, Baining Guo, Demetri Terzopoulos, and Heung-Yeung Shum. Geometry-driven photorealistic facial expression synthesis. IEEE Transactions on Visualization and Computer Graphics, 12(1):48–60, January/February 2006. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
comp/trans/tg/2006/01/v0048s.
avi.

Zhang:2018:DTR

[ZLK^{+}18] Jingxin Zhang, Eike Langbehn, Dennis Krupke, Nicholas Katzakis, and Frank Steinicke. Detection thresholds for rotation and translation gains in 360^{∘} video-based telepresence systems. IEEE Transactions on Visualization and Computer Graphics, 24(4): 1671–1680, April 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2018/04/08314105-
abs.html.

Zhang:2018:CSG

[ZLZY18] Mohan Zhang, Hongwei Lin, Kang Zhang, and Jinhui Yu. Computer simulation and generation of moving sand pictures. IEEE Transactions on Visualization and Computer Graphics, 24(12): 3058–3068, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8141977/.

Zhang:2013:LDG

[ZM13] Yubo Zhang and Kwan-Liu Ma. Lighting design for globally illuminated volume rendering. IEEE Transactions on Visualization and Computer Graphics, 19(12):2946– 2955, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2017:QVI

[ZM17] Yifan Zhang and Ross Maciejewski. Quantifying the visual impact of classification boundaries in choropleth maps. IEEE Transactions on Visualization and Computer Graphics, 23(1):371–380, January 2017. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zeng:2010:SPC

[ZMG^{+}10] Wei Zeng, Joseph Marino, Krishna Chaitanya Gurijala, Xianfeng Gu, and Arie Kaufman. Supine and prone colon registration using quasiconformal mapping. IEEE Transactions on Visualization and Computer Graphics, 16(6):1348–1357, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zachow:2009:VEN

[ZMH^{+}09] Stefan Zachow, Philipp Muigg, Thomas Hildebrandt, Helmut Doleisch, and Hans-Christian Hege. Visual exploration of nasal airflow. IEEE Transactions on Visualization and Computer Graphics, 15(6):1407–1414, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2019:VAL

[ZMT^{+}19] Zhiguang Zhou, Linhao Meng, Cheng Tang, Ying Zhao, Zhiyong Guo, Miaoxin Hu, and Wei Chen. Visual abstraction of large scale geospatial origin-destination movement data. IEEE Transactions on Visualization and Computer Graphics, 25(1): 43–53, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440039-
abs.html.

Zhang:2015:VCA

[ZMZM15] Zhiyuan Zhang, Kevin T. Mc-Donnell, Erez Zadok, and Klaus Mueller. Visual correlation analysis of numerical and categorical data on the correlation map. IEEE Transactions on Visualization and Computer Graphics, 21 (2):289–303, February 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2015/02/06881685-
abs.html.

Zhang:2016:UVE

[ZNZX16] Qing Zhang, Yongwei Nie, Ling Zhang, and Chunxia Xiao. Underexposed video enhancement via perceptiondriven progressive fusion. IEEE Transactions on Visualization and Computer Graphics, 22(6):1773–1785, ???? 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ziemkiewicz:2013:HVL

[ZOC^{+}13] Caroline Ziemkiewicz, Alvitta Ottley, R. Jordan Crouser, Ashley Rye Yauilla, Sara L. Su, William Ribarsky, and Remco Chang. How visualization layout relates to locus of control and other personality factors. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1109–1121, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zheng:2019:GDS

[ZPG19] J. X. Zheng, S. Pawar, and D. F. M. Goodman. Graph drawing by stochastic gradient descent. IEEE Transactions on Visualization and Computer Graphics, 25(9): 2738–2748, September 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zheng:2005:TLT

[ZPP05] Xiaoqiang Zheng, Beresford N. Parlett, and Alex Pang. Topological lines in 3D tensor fields and discriminant Hessian factorization. IEEE Transactions on Visualization and Computer Graphics, 11(4):395–407, July/August 2005. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2004:NPM

[ZPS04] Yu Zhang, Edmond C.

Prakash, and Eric Sung. A new physical model with multilayer architecture for facial expression animation using dynamic adaptive mesh. IEEE Transactions on Visualization and Computer Graphics, 10(3):339–352, May/June 2004. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://csdl.computer.
org/comp/trans/tg/2004/
03/V0339-supp.pdf;   http:
//csdl.computer.org/comp/
trans/tg/2004/03/v0339abs.

                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2004/03/
                                        v0339.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2004/03/v0339.pdf.

Zwicker:2002:ES

[ZPvBG02] Matthias Zwicker, Hanspeter Pfister, Jeroen van Baar, and Markus Gross. EWA splatting. IEEE Transactions on Visualization and Computer Graphics, 8(3):223–238, July/September 2002. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        comp/trans/tg/2002/03/v0223abs.
                                        htm;   http://csdl.computer.
                                        org/dl/trans/tg/2002/03/
                                        v0223.htm;   http://csdl.
                                        computer.org/dl/trans/tg/
                                        2002/03/v0223.pdf.

Zhang:2011:OPR

[ZQS11] Nan Zhang, Huamin Qu, and Robert Sweet. Orientationpreserving rod elements for real-time thin-shell simulation. IEEE Transactions on Visualization and Computer Graphics, 17(6):822– 835, June 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2018:BVE

[ZSCC18] J. Zhao, M. Sun, F. Chen, and P. Chiu. BiDots: Visual exploration of weighted biclusters. IEEE Transactions on Visualization and Computer Graphics, 24(1):195–204, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2013:APM

[ZSG^{+}13] Xin Zhao, Zhengyu Su, Xianfeng David Gu, Arie Kaufman, Jian Sun, Jie Gao, and Feng Luo. Area-preservation mapping using optimal mass transport. IEEE Transactions on Visualization and Computer Graphics, 19(12):2838– 2847, December 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2016:GBC

[ZSL^{+}16] C. Zhang, T. Schultz, K. Lawonn, E. Eisemann, and A. Vilanova. Glyph-based comparative visualization for diffusion tensor fields. IEEE Transactions on Visualization and Computer Graphics, 22 (1):797–806, January 2016. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2007:TSD

[ZSTR07] Howard Zhou, Jie Sun, Greg Turk, and James M. Rehg. Terrain synthesis from digital elevation models. IEEE Transactions on Visualization and Computer Graphics, 13(4):834– 848, July/August 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://bell.computer.org/
                                        dlcomments/;   http://csdl.
                                        computer.org/comp/trans/
                                        tg/2007/04/v0834s.avi.

Zhou:1999:ESV

[ZT99] Yong Zhou and Arthur W. Toga. Efficient skeletonization of volumetric objects. IEEE Transactions on Visualization and Computer Graphics, 5(3):196–209, July/ September 1999. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://dlib.computer.org/
                                        tg/books/tg1999/pdf/v0196.
                                        pdf;   http://www.computer.
                                        org/tvcg/tg1999/v0196abs.
                                        htm.

Zhou:2009:ATF

[ZT09] Jianlong Zhou and Masahiro Takatsuka. Automatic transfer function generation using contour tree controlled residue flow model and color harmonics. IEEE Transactions on Visualization and Computer Graphics, 15(6):1481–1488, November/December 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zheng:2012:DSS

[ZTA12] Youyi Zheng, Chiew-Lan Tai, and Oscar Kin-Chung Au. Dot scissor: a single-click interface for mesh segmentation. IEEE Transactions on Visualization and Computer Graphics, 18(8):1304– 1312, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zafeiriou:2005:BRW

[ZTP05] Stefanos Zafeiriou, Anastasios Tefas, and Ioannis Pitas. Blind robust watermarking schemes for copyright protection of 3D mesh objects. IEEE Transactions on Visualization and Computer Graphics, 11(5):596–607, September/October 2005. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2017:CPS

[ZTZ17] Yizhong Zhang, Yiying Tong, and Kun Zhou. Coloring 3D printed surfaces by thermoforming. IEEE Transactions on Visualization and Computer Graphics, 23(8):1924– 1935, August 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2017/08/07536161-
abs.html.

Zheng:2013:PHS

[ZTZX13] Youyi Zheng, Chiew-Lan Tai, Eugene Zhang, and Pengfei Xu. Pairwise harmonics for shape analysis. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1172–1184, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zinke:2007:LSF

[ZW07] Arno Zinke and Andreas Weber. Light scattering from filaments. IEEE Transactions on Visualization and Computer Graphics, 13(2):342– 356, March/April 2007. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhou:2018:IPP

[ZW18] Liang Zhou and Daniel Weiskopf. Indexed-points parallel coordinates visualization of multivariate correlations. IEEE Transactions on Visualization and Computer Graphics, 24(6): 1997–2010, June 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2018/06/07911335-
                                        abs.html.

Zhang:2013:FWI

[ZWA^{+}13] Zhiyuan Zhang, Bing Wang, Faisal Ahmed, I. V. Ramakrishnan, Rong Zhao, Asa Viccellio, and Klaus Mueller. The five Ws for information visualization with application to healthcare informatics. IEEE Transactions on Visualization and Computer Graphics, 19(11):1895– 1910, November 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zmuda:2013:OCE

[ZWBH13] Michael A. Zmuda, Joshua L.

Wonser, Eric R. Bachmann, and Eric Hodgson. Optimizing constrained-environment redirected walking instructions using search techniques. IEEE Transactions on Visualization and Computer Graphics, 19(11):1872– 1884, November 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2016:DPD

[ZWC^{+}16] Ran Zhang, Shiwei Wang, Xuejin Chen, Chao Ding, Luo Jiang, Jie Zhou, and Ligang Liu. Designing planar deployable objects via scissor structures. IEEE Transactions on Visualization and Computer Graphics, 22(2):1051– 1062, February 2016. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2018:SVA

[ZWC^{+}18] X. Zhao, Y. Wu, W. Cui, X. Du, Y. Chen, Y. Wang, D. L. Lee, and H. Qu. SkyLens: Visual analysis of skyline on multi-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 24(1):246–255, January 2018. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2012:KVE

[ZWJZ12] Hui Zhang, Jianguang Weng, Lin Jing, and Yiwen Zhong. KnotPad: Visualizing and exploring knot theory with fluid Reidemeister moves. IEEE Transactions on Visualization and Computer Graphics, 18(12):2051–2060, December 2012. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2019:IIR

[ZWLC19] Xun Zhao, Yanhong Wu, Dik Lun Lee, and Weiwei Cui. iForest: Interpreting random forests via visual analytics. IEEE Transactions on Visualization and Computer Graphics, 25(1):407– 416, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2019/01/08454906-
                                        abs.html.

Zheng:2013:PBD

[ZWM13] Lin Zheng, Yingcai Wu, and Kwan-Liu Ma. Perceptuallybased depth-ordering enhancement for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 19(3): 446–459, March 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:MMA

[ZWM^{+}19] Jiawei Zhang, Yang Wang, Piero Molino, Lezhi Li, and David S. Ebert. Manifold: A model-agnostic framework for interpretation and diagnosis of machine learning models. IEEE Transactions on Visualization and Computer Graphics, 25(1):364– 373, January 2019. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://www.computer.org/
csdl/trans/tg/2019/01/08440091-
abs.html.

Zhang:2014:VDM

[ZWR14] Hui Zhang, Jianguang Weng, and Guangchen Ruan. Visualizing 2-dimensional manifolds with curve handles in 4D. IEEE Transactions on Visualization and Computer Graphics, 20(12):2575– 2584, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876027-
abs.html.

Zhang:2017:PPB

[ZWS^{+}17] Fang-Lue Zhang, Jue Wang, Eli Shechtman, Zi-Ye Zhou, Jia-Xin Shi, and Shi-Min Hu. PlenoPatch: Patchbased plenoptic image manipulation. IEEE Transactions on Visualization and Computer Graphics, 23(5): 1561–1573, May 2017. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://www.computer.org/
                                        csdl/trans/tg/2017/05/07414488-
                                        abs.html.

Zhang:2012:DSM

[ZWW^{+}12] Yizhong Zhang, Huamin Wang, Shuai Wang, Yiying Tong, and Kun Zhou. A deformable surface model for real-time water drop animation. IEEE Transactions on Visualization and Computer Graphics, 18(8):1281– 1289, August 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2013:PSP

[ZWZ^{+}13] Wenyao Zhang, Yi Wang, Jianfeng Zhan, Beichen Liu, and Jianguo Ning. Parallel streamline placement for 2D flow fields. IEEE Transactions on Visualization and Computer Graphics, 19(7): 1185–1198, July 2013. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2015:VMD

[ZWZD15] Huayan Zhang, Chunlin Wu, Juyong Zhang, and Jiansong Deng. Variational mesh denoising using total variation and piecewise constant function space. IEEE Transactions on Visualization and Computer Graphics, 21(7):873–886, July 2015. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

URL http://www.computer.
org/csdl/trans/tg/2015/
07/07029103-abs.html.

Zhang:2018:MRT

[ZX18] Hao Zhang and Feng Xu.

MixedFusion: Real-time reconstruction of an indoor scene with dynamic objects. IEEE Transactions on Visualization and Computer Graphics, 24(12):3137– 3146, ???? 2018. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

https://ieeexplore.ieee.
org/document/8241434/.

Zheng:2010:VVV

[ZXM10] Ziyi Zheng, Wei Xu, and Klaus Mueller. VDVR: Verifiable volume visualization of projection-based data. IEEE Transactions on Visualization and Computer Graphics, 16(6):1515–1524, November/December 2010. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2009:ATA

[ZYLL09] Eugene Zhang, Harry Yeh, Zhongzang Lin, and Robert S.

Laramee. Asymmetric tensor analysis for flow visualization. IEEE Transactions on Visualization and Computer Graphics, 15(1):106–122, January/February 2009. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2014:VAP

[ZYM^{+}14] Jiawan Zhang, E. Yanli, Jing Ma, Yahui Zhao, Binghan Xu, Liting Sun, Jinyan Chen, and Xiaoru Yuan. Visual analysis of public utility service problems in a metropolis. IEEE Transactions on Visualization and Computer Graphics, 20(12): 1843–1852, December 2014. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876008-
abs.html.

Zhou:2008:VDM

[ZZBW08] Kaichi Zhou, Eugene Zhang, Jiřì Bittner, and Peter Wonka. Visibility-driven mesh analysis and visualization through graph cuts. IEEE Transactions on Visualization and Computer Graphics, 14(6):1667–1674, November/December 2008. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2011:IMC

[ZZC11] Juyong Zhang, Jianmin Zheng, and Jianfei Cai. Interactive mesh cutting using constrained random walks. IEEE Transactions on Visualization and Computer Graphics, 17 (3):357–367, March 2011. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zgraggen:2014:PDA

[ZZD14] Emanuel Zgraggen, Robert Zeleznik, and Steven M. Drucker. PanoramicData: Data analysis through pen & touch. IEEE Transactions on Visualization and Computer Graphics, 20(12):2112– 2121, December 2014. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

http://csdl.computer.org/
csdl/trans/tg/2014/12/06876039-
abs.html.

Zhao:2012:CMF

[ZZG^{+}12] Xin Zhao, Wei Zeng, Xianfeng David Gu, Arie E. Kaufman, Wei Xu, and Klaus Mueller. Conformal magnifier: a focus + context technique with local shape preservation. IEEE Transactions on Visualization and Computer Graphics, 18(11):1928– 1941, November 2012. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2019:IMS

[ZZH19] Lei Zhang, Qingzhuo Zheng, and Hua Huang. Intrinsic motion stability assessment for video stabilization. IEEE Transactions on Visualization and Computer Graphics, 25(4):1681–1692, April 2019. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        https://ieeexplore.ieee.
                                        org/document/8322185/.

Zhang:2015:BRG

[ZZL^{+}15] Yu-Wei Zhang, Yi-Qi Zhou, Xue-Lin Li, Hui Liu, and Li-Li Zhang. Bas-relief generation and shape editing through gradient-based mesh deformation. IEEE Transactions on Visualization and Computer Graphics, 21(3): 328–338, March 2015. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306. URL

                                        http://csdl.computer.org/
                                        csdl/trans/tg/2015/03/06975236-
                                        abs.html.

Zheng:2006:SST

[ZZM06] Jiang Yu Zheng, Yu Zhou, and Panayiotis Mili. Scanning scene tunnel for city traversing. IEEE Transactions on Visualization and Computer Graphics, 12(2): 155–167, March/April 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhang:2010:VML

[ZZSS10] Nan Zhang, Xiangmin Zhou, Yunhe Shen, and Robert Sweet. Volumetric modeling in laser BPH therapy simulation. IEEE Transactions on Visualization and Computer Graphics, 16(6):1405– 1412, November/December 2010. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.